A Bibliography of Publications about the Java Programming Language, 2000–2009

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

08 August 2024 Version 2.178

Abstract

This bibliography records books about the Java Programming Language and related software.

Title word cross-reference

#55 [Och09a]. #56 [Och09b]. #57 [Och09c]. #58 [Och09d].

1 [Lia03b]. $14.95 [Bal03c, Ano03b]. 2 [BDRV01, BBGP01, MD00, MCLC02, Tre03]. $29.95 [Ano00b]. 3 [Ano01o, Ano02m, Bar00c, BE02, CWWS03, CN03a, Che03a, CF02, CE01, FMA02, GV05, GP05, Hit03, HJF06, JHSL03, MD00, Nik03, PFJ05, Sei09, SQG^{+}05, WBS01, WWSL02, Yah01]. $34.95 [Ano00c]. $39.99 [Kuc06]. $52.50 [Ano01a]. $74.99 [Mil08]. $75.00 [Cha05a]. $79.95/£ [Azi06]. $83.95 [Ano04e]. $99 [Kro00a]. ^{(}R) [LS04a]. ^{T M} [Bla03, Cza00, IKY^{+}00b, IKY^{+}00a, MZB00, QGC00, Win02, vdPE02]. G [CiLH01]. ≫ [Rum01]. k [dCG^{+}02]. ≪ [Rum01]. m [BO09]. Cl(4, 1) [Hit03]. mc [BO09]. μ [vdPE02]. μoνo2πoλυ [Lik04a]. N [Rol08b]. Ω [BO09].

-D [MCLC02]. -Machine [CiLH01]. -pure [Ano03-31]. -Queens [Rol08b]. -space [dCG^{+}02]. -valued [Yah01]. -Wire [Lia03b].

.INI [Mey03]. .NET [Cha05a, SKS08, Ano02r, Ano05e, Apr05, Bar03c, BHW05, Bri05, Bro09, FLMS06, GS05a, HF06, HJR^{+}03, LN04, LAT04, Lut03b, Lyk02, Men03, PE06, SM04b, Stu07, Way03, Zhu04, Ano04o, DHR^{+}01, Kil03b]. .NET-to-Java [Apr05].

/Java [Och09c, Och09d, Och09a, Kum04, Kum05]. /MOM [DJLT01].

0 [Bal03c, Cha05a, Che05, Pet06]. 0-262-69276-7 [Bal03c]. 0-521-52583-7 [Cha05a]. 0-521-77477-2 [Pet06]. 0-521-89308-9 [Che05]. 0-7506-6496-7 [Dud06]. ’01 [Ano00a, Ano01b, Ano01g, USE01c, USE01b]. ’02 [USE02]. ’05 [ACM05, Chr05].

1 [AF03, Ano03-31, CCC^{+}04, Kuc06, She03]. 1-2-3 [Ano00h]. 1-59059-503-3 [Kuc06]. 1-85233-704-4 [Azi06]. 1.2 [CG01]. 1.4 [WMC04]. 1.5 [Ano03-36, Ano04p, S.04a, KHKH01, Lan04, S.04b]. 10 [Ano03-36]. 10-Gigabit [Ano03-36]. 10.4-4 [YMP^{+}05]. 100 [Mar01b]. 10G [Ano04-29, KM07]. 13 [Cow01]. 19005-1 [ISO05]. 1Og [Ano05i, Ano05i]. 1st [Ano01b, Mil08].

2 [Ano00e, Ano01m, Ano05i, Aus00, Ber00a, BC01, Bir01, BS00a, BH03, CL03a, CI01, DS00a, DDS02, DD02a, Gab07, Gig00, Goo03b, HS00a, Haw02, HC01a, HC02, HC03, JRN00, KT00, KCF01, Knu01b, Lad01, LG99, LG00a, Lit00, LRO02, Lut00, Pet06, RTVH01, SC01a, SO00, Sch01, Sha00b, Swa01b, WCS00, WN01, vdL02]. 2.0 [Ano00m, Ano00n, GAG06, KL07, NPRC01, Rao02, Sch03b, Tul02, Wal03c, WMM04]. ’2000 [ACM00b, ACM00a, Ano00n, GHM^{+}01, Kro00a, Kro00b]. ’2001 [ACM01d, ACM01b, Ano01e, Pap05]. ’2001/PERFORMANCE [ACM01d]. 2002 [GAR03]. 2002-21-0002 [San02b]. 2003 [ACM03b]. 2004 [ACM04]. 2004Q2 [Ano04-35]. 2005 [Car06, Gla06, ISO05, Won05]. 2007 [SM07]. 2008 [LL08a]. 21 [AJ01b]. 25th [SBH^{+}04]. 27.99/US$44.95 [Dud06]. 2D [Har00b, Gea00, Rod01]. 2k [USE00b]. 2nd [Ano02b, Feu02, GDC^{+}04, Mas01, Zen02, USE02].

3 [DC09, Ell06, KK03a, Kuc06, Lia00a, Lia00c, MMBAS04, Sch00b]. 3.0 [Ano05k, CSFS00, Hei01, WA04]. 3.1 [Ano04j, See04]. 30 [AGG02]. 310-025 [HS00a]. 32 [SOK^{+}04]. 32-Bit [Ano02p, Ano02j, VED06, Whi03a]. 32bit [XX05]. 390 [DBC^{+}00, GEAS00]. 3D [SRD00, WG02, BL04, SML06, WSVX03, XAN07]. 3D-Molecular [BL04]. 3D-Molekulvisualisierung [BL04]. 3rd [ACM06].

4 [Ano00m, Lia02, Lia03a, SC05, Wal02a]. 45-degree [TP08]. 45.00/£ [Azi06]. 4847-51 [Bus02b]. 4th [GRR05].

5 [Cur07, Hef07, HTY^{+}03, IEE02b]. 5.0 [Won04]. 5.6 [Ano00m]. 500 [Pra03]. 5029-90 [ZAVT03]. 5033-55 [MF03]. 5367-05 [HBX^{+}04]. 5434-19 [CHMB04]. 5684-20 [VVG^{+}05].

6 [Ano04-36, KWM^{+}08, Tan07]. 6.0 [Ano00m, Lia00b]. 6.1 [Nyb02]. 61499 [TSL^{+}04]. 63.50 [Ano04e]. 64 [IKN03]. 64-bit [Ano02j, BWLR06, VED06, VED07]. 6th [USE01a].

7 [Ano01n, Bal03c, Cha05a, Mil08]. 7th [USE00b].

8 [Ano03c, Ano03x]. 819.315 [Sib00]. 8i [DHMT00].

9 [Che05]. 9075-13 [ISO08]. 95

[BW01b, BW04, GD00, Wel03]. 978 [Mil08]. 978-1-4302-0973-7 [Mil08].

A-1 [ISO05]. A.NET [Men03]. A/V [ZP03]. A300 [YKS^{+}02]. Abaco [Ano01o]. Abbotsbrook [Ano00k]. Abrupt [HJ00]. Abstract [BDT04, BD02, Dro01a, GSW00, JR05, LM02, PL05, SSV05, BDL^{+}08, DC09, Dil00, KPH^{+}09, SCWL08, WB01, WBF^{+}06, Wit00, vMV05]. AbstractCollection [Hui02]. Abstracted [PDV01]. Abstraction [BS04, CP04, CP01, DGGD08, LH08b, LG00b, PB08, Soo09, ZR07]. Abstractions [CD03]. Academic [Ber05a]. academically [CR02b]. academically-diverse [CR02b]. accelerated [BHDS09]. Accelerates [Ano03-37]. Accelerating [OOOiM05]. Acceleration [DEK^{+}03, Ano03-46, JMP09]. Accelerator [Ano02c, KMOS03, DPT^{+}02]. Access [AK01, Ano02s, CCSA02, Gun01, HD02, KPK02, Kro00b, OWR04, Smi01b, SCLV04, Ano03-42, GB01, HO03, HO07, MF03, NC04a, Oi08, PH00a, RR01, Sch04a, KT01a]. Accessibility [CFGL05, CY02, CHUB08]. accessible [Rob00b]. accessors [TJ00]. According [TSL^{+}04]. Accounting [Lai08, SAWW01, BH04b, HB08]. Accrual [FBR^{+}03]. Accurate [ZSCC06, Bin06, CM02, ZR07]. achieve [Ano03-50]. Achieving [WW09, WC00a]. Achilles [XSaJ08b]. ACL2 [LM04, Moo03a]. ACLU [Bar01c]. ACM [ACM00b, ACM04, ACM05, CNB00, IEE02a, Jac04b, LL08a, Fox00a, Fox00b, Fox00c, Fox01a, Fox05, RBC^{+}05, RBC^{+}06]. ACM/IFIP/USENIX [Jac04b]. ACM/USENIX [ACM05]. acme [AGST04a, AGST04b]. Acquisition [Lin03a]. Acronyms [Bar01a]. Across [Nat00, KLS00, PWC00, SGW01, TM07]. Act [Atk01]. Actel [Ano02n]. Action [BK05a, CPJ05, FF05, Rei03, Ric06a, WRO04, HD03c, Man05, WB05, WB08]. Action-Demonstration [Rei03]. Active [SLC03b, Ham07, New01, XX04].

ActiveScaffold [STB08]. ActiveState [Ano00m, Ano00n, Ano01m]. ActiveX [Wil04a]. activities [Bow07]. Activity [AH04b, Bar09, CQX^{+}09, Ren00, TBM09]. Activity-based [Bar09, TBM09]. ActorFoundry [BNO03]. ad [SM01a]. Ada [BD01b, Bro03a, BW03a, BW03b, Bro04, Bro05, BA07b, BW01b, BW04, CVW03, Car06, GD00, KPPÉR06, Lam03, MH09, Och09c, Och09d, Och09b, Och09a, Och09e, Pot04, San02a, San03, SC01b, Swa07, Ten00, Wel03, Wil06]. Ada95 [KK03b, NMH^{+}02]. Adabas [DHMT00]. Adaptable [SMCS04, BIB05]. Adaptation [BR01d, ONRV08, RW04, WSM06]. Adaptec [Ano03-36]. Adapter [Ano02q]. adapters [Apt02]. Adapting [AG05, DH00, EKEL01, JMSG02, Kon03, LBJ05]. adaption [AK09]. Adaptive [AFG^{+}00, FOS^{+}04, KDH^{+}06, KM02, LBJ02, OL01, PSZ^{+}07, QH03, WHKS01, Wol01a, ZK04a, Gra04, NC05, SVY09, ZSCC06]. Add [Bar01b, WS01c, Ano04-27, CFL05b]. added [ZJ03]. Adding [NHY^{+}04, vRS05, Ano03x, ABL08, KdJNNV09, TE05]. Addition [Dau01]. Address [LCHY03, And01, Ano03g]. Adds [Ano00m, Ano02m, Ano03-38, Ano03-40, Ano02v, Sur04a]. Administration [Ano01o]. administrator [Pan04]. Adobe [Ano02t, CDH07]. Adopting [BN03]. adoption [Ano03w]. advance [SCH05]. Advanced [AWS^{+}09, BZ05, Ber00a, BF02, Bur02, CY04, DF03, DDS02, Dud06, FR02, Gea01, Hei03b, HC02, KC00, Lan05b, LZ04, LCHY03, NC05, Pro01, Rod01, SS00b, Top00, ADT03, Aus00, BZ07, BVD01, OHL^{+}05, Ano01m, NIS00]. Advances [LBQ00, Ano04w]. Advantages [Bro03a, Lex02]. adventures [Lab09]. Advice [Mor03b]. aerial [HHM04]. AES [Dra00, SL00, Bro02b]. Aether [Ano01m]. affect [RVZ04]. affecting [PJ05]. affects [Eng00]. again [Rol05]. against [BSPF01, BSB^{+}03, MP05, Pre03]. Age [Thi02, MFH01]. Agent [BIB05, Bru02, Det01, FVK01, LL01a, RC01, RB01, VB01a, VHL01, Vrb03, ACZ05, MJ00, SSC00]. agent-based [MJ00]. agent-oriented [ACZ05]. Agents [BIB05, CWHB03, CY03, ES06, IKKW01, Jon02, Liu03, NP01, SSM03, Sat04, SV02, AHN02, BWLP01, BB01, CFL05b, CFL05a, ESPP01]. Agere [Ano02t]. aggregate [TGO00]. aggressive [MGM^{+}06]. Agile [SH06]. Agilent [Ano04b]. agility [Way05]. Aglets [Jon02]. Agreement [Bar01b]. AGVs [YHL01]. ahead [CSFS00, HKS^{+}07, HKM^{+}09, JPB^{+}08]. ahead-of-time [HKS^{+}07, HKM^{+}09, JPB^{+}08]. AI [Lut03a, MJ00]. Aid [NLC03]. Aided [Kog04, KNG02, ZG04]. aim [WVMN05]. aimed [Way03]. Air [CDH07]. AJA [BIB05]. AJAX [DV07, CPJ05, Cur07, Fit07, GAG06, JF06, Mah06, McL06a, MGB^{+}09, Mor08a, Ols07, Per06, Ski07]. AjaxScope [KL07]. Ajents [ICB00]. AJIS [Och09b]. al. [Fox01d]. ALAT [LCHY03]. Alfonse [Har00e, Har01b]. Algebra [CCR00, GGHvdG01, BB05, Gam00, LFG00]. Algebraic [HD03a, Tra00b, Fei01, HRD08b]. Algorithm [ABG02, Bar00a, Bar01b, Bar01c, EGLZ02, LSW08, TT01, ZX05, BS07, EKEL01, GGL^{+}08, JFH00, LPH06, LH07, Nau02, RV05, VIPCUF08, SA02]. Algorithms [All00c, BH02a, BGadH06, BP05, GT97, GT04, GT06, GT10, KC01, Ler03, LPSY04, Lut01, Lut03b, Mas01, MH00a, Par04a, PGM^{+}05, RS01, Sch02, Sed03, SL00, TCM^{+}00, ZT02, BV05, CCT01, Dro01b, GT01, MCHN05, NM02, OG05, Pre00b, Sah00, WB01, WM00b, Wu05, dCG^{+}02, vdBDS00, Lut02]. Alias [WGW04, Woo05]. aliased [BA07a]. aliasing [FYD^{+}08, Gad03, MF07a, NA07]. Alice [DC09, LS08c, Pau08, Sei09]. alignment [CCSB04]. alleviate [Apr05].

Allocation [CCM05, KMEA04, SGF^{+}02, YLL^{+}07, ZSZ^{+}09, CGS^{+}03, EFJM07]. Allocator [LMK06, QH03]. Allow [KFLN04, OJ09]. Allowing [RTJ00]. almost [BR06b, BK05b, Duc08, PT09b]. almost-whole [BK05b]. alnoite [INM05]. Along [Pau03]. alpha [BD03a]. alpha-Methyl [BD03a]. Altera [Ano02s]. Altering [TSDNP02]. Alternative [CF03, LR04, MLG^{+}02b, Ano05b]. Alternatives [SLB^{+}02, Swa01a]. although [Ano05n]. Altia [Ano02q, MD00]. Alto [ACM01b]. am [Lex02]. Amazon [LAT04]. among [Ano04b, BA09, MT07, TS01]. amp [Ano03i]. AMPS [Lin03a]. Analyse [Wol03a, Wol03b, Zus03, Ano04c]. Analyser [PL05]. analyses [BS09, LPH01, MRR02]. Analysing [BD02, Sch04a, PV06]. Analysis [Ano01h, Ano02o, Ano02p, Ano03-40, ASB^{+}04, AW03, BCMT03, Bar01b, BHJR05, CHS01, CC04, Dra00, FCMR04, FMR05, GNYZ05, GS05b, Hec07, HJR^{+}03, Hol06, HWB03, JRN00, KOO08, KC01, KMS04, KK03b, KPK02, KP01, Laz07, LYC02, LH03b, Liu04, LFH03, Mac05, Mor03c, MOS07, NT01, PCC01, RWL07, RST^{+}04, RCR06, RMR03, RMR04, RKG04, SR05, SF01, SR06, SK00, She03, SPR^{+}03, SCLV04, SBA01, SM02b, TH02, Way05, Wei01, Wol03b, WGW04, Woo05, XC01, Zus03, dL05, ACM01a, ABLU00, Ano03-34, Ano03-35, Ano05k, BGH^{+}06, Bla03, BGNM04, BS00b, BPSH05, BGED04, CM05a, Cha06, CRL01, CTF03, CGS^{+}03, Cor00, DH08, DV01, EKVM07, GW08, GPW03, HEJ09, JCYC04, JPSN09, JKH^{+}04, KGH^{+}05, KH00, LH08a, LH08b, LPH02, LSW07, LFG00, MBED06, MSG01, Mas00, MP05, MRR05]. analysis [MLM^{+}08, Mur05, NK06, NC04a, Off00, PH00c, RV05, RSS^{+}04, RSD01, RMR01, RJGH06, SBAD01, SAB08, SGK09, SK08, SS08, ST00a, SGSB05, SB06b, TM07, TPF^{+}09, Uni03, Ano04c, Ano05k, DHPW01, MVM07]. analytical [TCC02]. Analyzer [Ano02m, Ano03-37, Ano03-39, Ano03-48, Ano03-35, DZHS03]. Analyzing [Li02, PV08, TCM^{+}00]. anatomic [Woo03]. anatomist [ZAVT03]. anatomy [GV05, GP05]. Anchor [MSK09]. Anders [Bar01a]. Andersen [LPH06]. Anderson [Ano04-29]. Andrew [Ano00d, Che05]. Andrews [Tra00b]. ANEJOS [SM01a]. Angle [Uni02, Ano02g]. Angles [Col02]. animated [BDFL04, HG08]. Animating [Gri02b]. Animation [DMU02, Pau03, JFH00, MMBAS04, VIPCUF08]. Animations [Soj03a, ABL07, Hu03]. animator [Gri03]. annotated [MMU04, RMR01]. Annotating [JK00]. Annotation [FL01, TT08, ANH00]. annotation-aware [ANH00]. annotations [Jac04a, Kic04, SD04]. Announcement [Ano00a]. Announces [Ano03-38, Ano03-39, Ano03-35, Ano03-36]. Annual [SBH^{+}04, USE00a]. Anomalous [HWM01]. Anomaly [SBAD01]. Anonymity [Bar01a, VV05]. ANSI [Oiw09]. ANSI-C [Oiw09]. Anspruchsvolle [Ste08b]. answer [Bur02]. Ant [Mor03b, Mor03b, HL02a, Hol05, NP03, PL03, TB02, ZK05]. Anthology [AE06, EA06, For06]. Anti [Ano00k]. Anti-Virus [Ano00k]. antipatterns [BPSH05]. Antonio [USE01a]. ANTS [Way03]. Anwendungen [Ano03s, Wol03b, Wol03a, Zus03]. Any [Pre03, CAF04]. Anything [McG03b]. Anytime [DJLT01]. Anywhere [DJLT01, Ano03-44]. AOP [TTPN08]. AP [DHRH05]. Apache [Gab07, GW00, Gou01, HJL00]. Apart [Lut00]. APDU [PvdBJ01]. API [Mil08, Zea00b, Ano03o, Ano03-34, BC00, EM04, Fit07, Gag02, Gea00, GGH^{+}03, Hap02, Har00b, HFL03, Hoh03, LS00, MP01b, MWM01, PvdBJ01, Rap03, RG00, Rou02b, SRD00, Tul08, VLMO09, WG02, Wal02a]. APIs [Ano02r, BKT03, BBGP01, Kon03, KKT04, Sun01]. APL [BL02b]. aplicaciones [Ano04-33]. App [Ano03-40, Vau03a, Way05]. Appajodu [Bar03a]. AppDev [Ano08, Pra08, BI07]. appeared [PPJ03]. AppForce [Ano03-35]. AppForge [Ano02o]. Appgen [Ano00k]. Apple [Ano01k]. Apples [Lut00, BNK^{+}07]. Applet [ACL03, Bar00a, BRL03, DMP05, Fre05, GKMZ04, GKW04, Hol04a, Iva02, MH00a, RT02, Ros00, TC03, ZFK04, Ano01d, Ano02v, CMS05, EGST08, GM02, Hu03, Rob07b, YL03]. Applet-Based [RT02]. Applets [And04, BF03, BL04, DK02, EH04, Hei03a, IKKM03, MdB01, Mos05a, RKK03, SSL02, Ano00f, Ano03e, Bis03, Fre01, Goo03b, HWM01, MR00a, Mls04, Moo03b, BL04]. Appliance [Kro00b, Ano03-34]. applicability [Man01]. Application [Ano00d, Ano01h, Ano01i, Ano01l, Ano01k, Ano01m, Ano01n, Ano01o, Ano02n, Ano02o, Ano02q, Ano04-37, Ano05i, BKT03, Ber05b, Bru05c, BG02, CF02, Cza00, DFL00, FOS^{+}04, GKM01, GW00, GM03, GMM00, HHK^{+}01, HK02a, HF00, Hon05, HCB04b, II04a, Ish01, JWC03, KSK04a, KK00, KK03a, KX04, Lia00c, MF01b, NZM03, Pip03, RCR06, Ren00, RT02, RC01, RW04, ESGS00, SM01b, Sta01, TCF^{+}03, TS02, TEM^{+}01, VWS^{+}05, Wan03a, ZS01b, ZX05, deC04, vdBJP01, Ano00c, Ano00g, Ano02e, Ano02w, Ano03-35, Che03c, CLM^{+}07, DLL03, Fei01, FL04, Gab07, GN01a, HSD04, Hef07, IK04, JDJ^{+}06, Kag09, KGH^{+}05, Kre01, KKT04, LSK^{+}02, LLS^{+}08, Mer04, PC08, Rem01, Roc01, Rol08b, SL06, SM03a, SD04, TABP07, Tre03, Tro04a, Tro04b, WAB^{+}04, XSaJ08b, ZS01a, ZR07, ZAVT03]. application [dMSAV08, Zea00b]. application-layer [Ano03-35, IK04]. Application-Monitoring [Ano02n]. Application-Specific [ZS01b, ZS01a]. Applications [AR03a, AA02b, Ano00k, Ano02q, Ano02t, Ano03s, Ano03-28, Ano03-37, Ano04d, AFT^{+}00, Bar03a, Bar05, Ben00c, Ber00a, BL02a, Bou01, BFM^{+}02a, BFM^{+}02b, BFS^{+}03, BRC03, BJK07, BSPF01, CW04a, CFLL03a, CI01, CM05b, Cer02, Cha03, CL03b, CGR00, CCB09, CGRR04, Cox01b, Des01, Dmi04, ET01, Fel03, FDTL02, Feu02, Fox00d, Fox03a, Fox03b, FGLS04, FBS04, GCB^{+}00, GAR04, GRR05, HE03, Joh03, KNY03, Kod04, Kro00a, KKK04, LLMK03, LR04, LS03, LD03, Mah04b, MSR03, MS03, MSSJ00, NMH^{+}02, PKF02, Ric06a, RS00b, RLR00, SAFG03, SK04, SGF^{+}02, SSS02, TSL03, Tor01, VKK^{+}01, WXW^{+}05, Wan05, WVE^{+}00, WHKS01, Yua03, Zea00a, dFR04, AU02, AK01, ASS^{+}05, Ano03-50, Ano03-51, Ano04f, Apr05, ABC^{+}07, Aus00, Bar02a, BDP02, BPSH05, BALP01, BALP06, Bre02, BVD01, BFW^{+}03, BSB^{+}03, Bur01b]. applications [BGED04, CV03, CB04, CHMB04, CLM^{+}09, CHL^{+}00, Cla04, CMLC06, CBGM03, DFW04, Die00, DBC^{+}00, DJLT01, DM07, ET07, Eng00, FTD03, FT06, FMRW05, FLWW04, GCRD04, Goo03b, GJ09, Gro02c, GAR03, HG08, HAL02c, HF06, Has02, Hig03, HD03c, ICB00, KK04a, KT00, KL07, Las02, LS00, LCFL04, LCZ04, LHFL07, Man01, MR09, MP05, McL02a, MGB^{+}09, MAJC03, Mor08a, NR06, NC04a, Gal02, NP03, Pet05, PNKN04, Ree02, Ric01, Rod01, Röß06, Sah00, San04a, SML06, SCBH09, SYAS05, SAB^{+}06, SW06, SKP^{+}02, ST00b, TT08, TPF^{+}09, WGSD07, Wea07, ZSZ^{+}09, vHMB08, Lut03c, Cal00a]. applicazioni [Pel03]. Applied [SAFG03, SM02a, Ano02o, Lut03b]. Applikationen [Ste08a]. Applying [AA02a, DF03, Lut03a, MS01]. Apprentice [KB04a]. Apprentice-Based [KB04a]. Approach [BO08, BB03, BRL03, CD01b, DJLT01, DFL00, FP03, JHJX04, KVK^{+}04, KM02, KS02b, PC04, QHV02, SD08, YDWL04, ABLU00, AW00, BP01c, BL02b, CFS09, CCKP06, CF04a, DMKN02, Fei01, Gra04, Gri08, HKI08, HL02b, HNZS03, LFM09, MSR09, MR09, SV05, SML06, SHM09, VN00, Vir03, BHS07, Lut02]. Approaches [AJMJS02, BLPV04, Egy01, Lam03, MMG01a, PH04, AHN02, BDT01, HB09]. Appropriate [Ron01, PHM^{+}01]. approximate [GEG07, GE08]. Apps [Ano03d, Ano03-38, Apr03, WA04, Ano03y, Ano03-30, Ano04d, Ano05i]. Apptivity [Ano00m]. Apress [Kuc06, Mil08]. April [Ano01g, NIS00, Uni01, USE01c]. Aprisa [Ano02q]. ARANEA [MCLDP01]. Arbitrary [GHM^{+}01]. Arc [Ano00n]. Architect [Mil08, Tul08, CR02a]. Architectural [ACN02, GHH01, JR02, AAAG^{+}05, Chr05, RVJ^{+}01]. Architecture [AA02b, BCH02, BALV03, BFS^{+}03, CQ05, Cha05a, DS09, EGLZ02, Gol00, Hsu01, Hua03, IKKW01, JLV02, KFLN04, KM04a, KR03, LMG00, LMG01, Lut02, MWL00, MB03, MTSM03, Rot02, SSB03, WFGK03, ZCQS04, AGST04a, AGST04b, Ano04y, AZ02, Apt02, CvE00, Che00, GCARPC^{+}01, GEAS00, Hub02, Ibb02, IKN03, Lee03, MAWW^{+}01, McL02a, PSS01, RB04, Swa07, WWJ07, Zhu04, Lut02, NT01, vdPE02]. Architectures [ABM^{+}03, Bru05c, CB04, HECR00, LR04, Par05, SAWW01, Ano02j, BWLR06, RJGH06]. Archives [RC01]. Archiving [Ano01i]. ArchJava [ACN02, AGST04a, AGST04b]. Aren’t [BHP^{+}01]. argumentation [CHMB04]. arguments [Lan04]. Arithmetic [Cow01, Dar01b, Fig00, MOS07, Win02]. ARLEQUIN [Sta01]. ARM [Ano03-38, DGMY06]. Aroma [Sur01]. ARP [Zdr09]. Array [Bur03, PH02, QHV02, Ano02j, BWLR06, CM05a, LGFM05]. ArrayLists [JT04]. Arrays [All00a, LK01, MMG01a, SF01, MMG03, JT04]. Arrival [Wat02]. arrow [GE08]. arrow-type [GE08]. arrows [KHFS09]. Art [BGP00, For04b, Mar05, Cha03]. article [Zus03]. Artikel [Wol03a, Zus03]. As-if-serial [ZK09]. Ascend [Ano01n]. Aside [SK04]. ASM [Zam03a]. ASM-based [Zam03a]. ASP [Kro00b]. ASP.NET [OBr05]. Aspect [KH01, Kic03, PSDF01, FB07, KKG09, LFM09]. Aspect-Oriented [Kic03, PSDF01, KH01, FB07, KKG09, LFM09]. AspectJ [HK02b, HZS08, Kic03, Mil05, PWBK07, RZW01, ACH^{+}05, BTV06]. Aspects [Hsu01, Ano02e, BLLB08, FB07]. assembler [MSU08]. assemblies [LCC09]. Assembly [Ano03-30, BD01a, Juo07, VS06]. Assertion [JSSM04, ÁdBdRS05]. assertion-based [ÁdBdRS05]. Assertions [BFMW04, Moo06]. assess [SCL^{+}08]. Assessing [CLP06, JFH00, Lut01, Mer04]. Assessment [Ano01k, BK01b, KWK03, SASZ03, Bro07, DMP09, Eng04, Eng06, ER09, HTSW07, SDF00]. Asset [Kro00a, GS00b, SDF00]. assignment [Djo09, GPF08, Liu08]. Assignments [LBD^{+}03, Par04b, Ros02b, Hel07b, Mor02, OJJ00]. assist [BC04, KKM^{+}06]. Assistance [FOS^{+}04, SFM^{+}07]. Assistant [FL01, Ano03-36]. Assisted [BCDdS02, Tre02c]. associated [San04a]. Associates [Ano01h, Ano02o]. Associating [VTD06]. Association [Ano00j, STB08]. Assurance [KKL^{+}04, KVK^{+}04]. assured [GHS05]. Astronomy [Bar01b, ZGB03]. Astrophysics [CO07]. Asynchronous [BBC07, BHR02, BW03a, BW03b, Hoh03, JPJ05, SM01c, Tddd03, vLSM01, Ano03k]. ATA [Ano03-36]. ATE [SFP03]. Atinav [Ano02m]. atlases [ZAVT03]. ATM [Zea00a]. Atomic [Ano03-39, HPS02, KKO02, BBA08, MBS^{+}08, RD06, San00, WMRT^{+}05]. atomicity [FFLQ08, NRS^{+}07, SMSAT08]. ATOMOS [CMC^{+}06]. Attached [Ano02m]. Attack [GM05c, Zdr09]. Attacks [LN02, Zdr09, MP05, SW06]. Attention [RCdBL02]. attract [PB06]. Attraktivität [Sel03]. attribute [CY02, NP07]. attribute-grammar [CY02]. attribute-oriented [NP07]. Attributes [Kic04, PQVR^{+}01]. audio [Lin00]. auditing [LAHC06]. Audits [Ano05k]. Aug [HRD08a]. Augmented [RPJ04, Wel03]. August [AGG02, Gho01, SBH^{+}04, Tra00b, USE00d, USE02]. Ausdrücke [SKS08]. Ausfallsicherheit [DHMT00]. Austin [IEE02b, USE00b]. Authentication [Cim02, EM03, Str01, SJ05]. Authoring [Ano01i, SL04, WDSD02]. authorship [DS04]. autoboxing [Lan04]. AutoCAD [Ano02m]. AutoCAD-to-PDF [Ano02m]. AutoGraL [BDRV01]. automata [FW02, Gri02b, LJ08, WW06]. Automate [Par00, Pau03]. Automated [Ano02n, Ano03-41, BDJ^{+}01b, BFMT00, CCR00, DH04a, DRV02, DC03b, Eng04, GN01a, HKK^{+}01, KF00, KY03a, KP01, MS03, BGNM04, BKM02, Eng06, ER09, HTSW07]. Automatic [AGMM00, Car06, CA04, CQX^{+}09, Ebe02, MdB01, MS00b, OS02, PP02b, PWN04, SMES01, SLC03a, SD01b, SD03b, TS02, UL08, WML02, ZR07, AC01, CLM^{+}07, CLM^{+}09, CS04, Fel03, Hel07b, KLS00, SB07, TABP07]. Automatically [Mor02]. Automating [Apr03, Kah06a]. Automation [AA04, PGM^{+}05, Ano05a, Cla04, HMD04]. Automatisierungssysteme [Ano05a]. automaton [Gri03]. automotive [BDRV01]. autonomous [EL04]. Auxiliary [vON02a, vON02b]. av [HJL00]. availability [KS01a]. Available [Ano03-41, DJLT01, GM02]. AVal [NP07]. Avanti [Ano03a]. Avatars [CF02]. Avinash [Ano04e]. avionics [ABC^{+}07]. await [San00]. Aware [Bar05, CHV01, RP03b, dFR04, ANH00, EQT07, HEJ09, Oga09, XSaJ08a, Zea00a]. Awareness [Bar05, ST09]. AWT [Rod01, WWJ07, WW09]. AWT/Swing [WWJ07, WW09]. AXe [Ano00j]. AXi [Ano00j]. AXIS [BI02, For04b]. Ayres [Fox01b, Fox01d].

B [BR01c, Req03, TRVH03, YWZ03]. B/S [YWZ03]. Babylon [vHMB08]. Back [GDC^{+}04, Reg06]. Backstop [MKKC08]. Backup [DHMT00]. Bad [BHP^{+}01, BNK^{+}07, MLM^{+}08, PWN04]. bad-smell [PWN04]. Balancing [Atk01, Gou01, FJ05a, FT06, GJ09, MRC03]. Baltimore [IEE02a]. ban [Gen00]. Bandera [HD01]. Bandwidth [KFN04, CM02]. bandwith [JH03]. banking [Van04]. Bantam [CL08]. BAOBAB [DG02]. BAPI [Sch00b]. barely [Mur07]. barrier [BKO09]. BASCOM [Ano00i]. base [Ano04-27]. Based [AA04, ABG02, AG03b, ABM^{+}03, AR03a, AL04b, Ano01h, Ano01k, Ano01o, Ano02p, Ano04-34, AAA^{+}04, BH02a, Bal03a, Ben00c, BNO03, BCH02, BL03, BLW00, BK01b, CLCC02, Che03a, CQX^{+}09, CiLH01, CBD01, CKKH03, CGRR04, DYH05, DK02, Ebe02, EXA^{+}05, EGLZ02, EM03, FSBP03, FVK01, FGLS04, GGG03, Gös03, GLS02, HD02, HHKS03, HK02a, Hit03, HJF06, HD03b, HL03b, Hua03, JSSM04, KM04b, Kie01, KM02, KB04a, KS04, Kum04, Kun02, KS02b, LL01a, LKL^{+}03, Li03, Lia03b, Lik04a, LHS04a, Liu03, MB03, MCLC02, MS01, MLG02a, Meh02, MSF03, NP01, NPRC01, NLFA02, N^{+}00, Omm01, PDCL02, PGM^{+}05, RM04, Ran02, Ren00, RT02, RKK03, Rum01, RP03b, SDPM04, SAWW01, SR06, SO02, SSS05, SRJS08, SL04, SSE05, TS01, TMG03, TFL^{+}04, TC04, TT01, VT01, VWS^{+}05, VB01a, Vrb03]. Based [WS01b, WXW^{+}05, WL04, WK02, YWZ03, YHL01, YHL04, ZL05, ZCQS04, ZYC03, ZK04b, ZX05, ZT02, dFR04, vLSM01, ÁdBdRS05, AK01, ACZ05, Ano00g, Ano00i, Ano01p, Ano03k, Ano03l, Ano03n, Ano03-29, Ano03-35, Ano03-36, Ano04n, Ano04-32, Ano05a, AZ02, Bak00, Bar09, BP01c, BD04, BR06a, BHM^{+}07, BDFL04, BKM02, BSBR03, BJ04, BKY^{+}03, BCR03b, CB04, CCT01, CW03b, CM02, CHB03, CCKP06, CMR05, CR02b, CL08, Cul00, DPT^{+}02, DLL03, DZHS03, EKEL01, EL04, Esp06, Est01, Fal00a, Fal00b, FMA02, FF00, FW02, Fre07, FL04, FCW01, FLWW04, GES^{+}09, GW08, GV05, GP05, GKL08, GW00, GE08, Gra04, Ham07, HLT09, HL03a, Hel07b, HKI08, HE03, Hon05, HKF00, HNZS03, HBH01, HdS^{+}05, HS02b, Ish01, IH01, JLV02, JT04, JFH00, JCP^{+}05, JH03, JKKL04, JMP09, JHSL03]. based [Kag09, KHMW05, KT01a, KLL03, Kro00a, Lab09, Lex02, LH04, LH08a, LH08b, LRW01, Li02, Li04, LCZ04, LMK06, Lik04b, LSK^{+}02, LW03, LYL^{+}04, LLS^{+}08, LAL02, LSW07, ML09, Mam01, MJ00, MAJC03, MM04, NK06, NIKN06, NHY^{+}04, NC04a, NC05, NKBM01, NMKB03, NZM03, OBr05, Oga09, Oi05, Oi06, Oi08, ONRV08, PSS01, PFS05, QH03, Rad06, RSS^{+}04, Röß06, Sam04, SM01a, SDF00, Sci07, Sha04, SGK09, SG02, SRW^{+}00, SS08, SB06b, SCFP00, SCH05, SYN03, SYN06, SD04, ST00b, TCF^{+}03, TSL03, Tre02c, TBM09, VDPC01, VDPC03, VN00, Vog03, WAF00, WAB^{+}04, Wen05, Wit00, Woo03, XP04, XAN07, YdOLS^{+}05, Zam03a, Zea00b, ZP03, ZLG08, dH05, dCG^{+}02, dGNv04, vNMW^{+}05, vNMKB05, vdSPP05, Ano02h, HKHK03, MAWW^{+}01]. basert [HJL00]. Basic [All00b, Ano01i, Ano01o, JP00, Bel02, MSK09, Ano04f, HM02]. Basics [CWH01, BMS02, LO03b, Reg06, ZCR^{+}06]. basierten [Lex02]. Basis [SSM03, CHL07, Way03, Ano01h, Ano01o]. Batting [Bar00a]. Battle [VN03, Vau03b]. Baudis [IEE03a]. BC [LL08a]. BDD [LH04, LH08a, LH08b]. BDD-based [LH04, LH08a, LH08b]. Be [Pet03, Sch03a, KS07, Rei00b, Rei00c]. BEA [Ano03-34, Ano04i]. Bean [BR01c, Ano02k, WCD^{+}01]. Beans [BR01c, Rao02, Sch03b, Ano02k, EK01, KMSL03, Pro01]. Beats [Bar01b]. because [Ano03f]. Becomes [Gee05]. becoming [Pay04]. Beefs [Ano05p]. been [Hun03a]. Before [Lut00, GKM01]. Beginner [Bro03b, Pol01]. beginners [Wis06]. Beginning [Bar03b, Hoo05, SB06a, WMC04, BMS02, Gol04a, Lar01, PRR02, Ska00, Ano01a]. Behavior [BP01c, BAJ01, DeP03a, GBED04, VKK^{+}01, YLW04, GS00c, HSD04, KL07, KH00, Oi08, SSGS01]. Behavioral [FLF01, LBR06]. Behaviors [SQG^{+}05, BCV03]. Behaviour [Hig04, BE02]. Behavioural [NT01, WS01c]. Behind [Lut03c]. Beispiel [Lex02]. Bell [Fox01b, Mer04]. BEM [Nik03]. Benchmark [Bar01c, DHPW01, GKM01, SBO01, ZS01b, BSW^{+}00, Eng00, GPW03, GPW05, Wan02a]. Benchmarking [BSPF01, BSB^{+}03, KS02b, BGH^{+}06, ZS01a]. Benchmarks [Ano03-38, Ano03g, BDF^{+}00, BGH^{+}06, KPH^{+}09, LJN^{+}00]. beneath [INM05]. Benefits [GD00, JFH00, LH08a]. Best [ACM01e, CMS03a, FCW01, Lut03b, OBr05, PSS01, SM01a, Sch03a, Way05, Eck02, FLMS06, Pan09, Ree03]. Bet [Lyk02]. Betriebsmanagementsystems [Lex02]. Betriebssystem [Ano04v]. Better [Gri06, MW05, PH02, TG04, Wel03]. Bettis [Fox01b]. Between [Pot04, Wan05, ASS03, AHKR01, BDJdS02, BF02, CF04a, CF04b, Lin01, LZZ03, NK03, QM09b, SCH05, Urb09]. Beyond [Tat05, Gag02]. biased [RD06]. Bible [WCS00, Goo01a, Goo01b]. Bibliography [Bee00]. Big [Hor02a, Hor02b, Hor05]. BigDecimal [CBD04, Sun02]. Bill [Gla06]. Binaries [JMSG02]. Binary [GEAS00, Jam01, PH00a]. Binding [Ano01o, Ano02t, CLL03, McL02b, dGNv04]. binds [Ano05i]. BioconX [Ano01n].

Bioinformatics [SHK^{+}03, CB04, KS04]. BioLayoutJava [GCEO05]. biological [HNZS03, THMT03]. Biomechanical [Eng00]. Biometric [Ano01n, EM03]. BIOMODULE [HPH03]. Biopathway [NDS^{+}02]. Birkhäuser [Pap05]. Birrell [MDJ05]. Bishop [Fox01b]. bison [Kag09]. bison/flex [Kag09]. Bit [Ano02p, Ano02j, BWLR06, VED06, VED07, Whi03a, ZFK04]. bits [Eub05]. Bitter [Tat02]. Bjarki [Fox01b]. Black [Hol00c]. BlackBerry [Ano02m]. Blaxxun [Ano00n]. bloat [XAM^{+}09]. Block [CCW02, TCM^{+}00]. blocking [HL03a]. Blocks [Pet03, TSL^{+}04, BBA08, EK03]. blowing [BVPE06]. Blue [CSFS00]. BlueJ [Hag00a, KR00, PH03, PHBM05, XSD07]. blueprint [Mur00, Pas04]. Bluetooth [Ano00m, Ano01j, Ano02m, Ano02n, Ano03o, Ano05a, BKT03, KKT04, VV05, WCCL05]. Bluetooth-Kommunikation [Ano05a]. Blunders [SLB^{+}02]. Board [Bar01b]. Bob [Bet02]. Body [RJFG03]. Bogavich [Fox01b]. Bohnenkamp [Ano08]. Bologna [FPA^{+}06]. Booch [Lam03]. Book [Ano00b, Ano00c, Ano00d, Ano01a, Ano03b, Ano04e, Ano08, Azi06, Bal03c, Bar03a, Bro02a, Cal00a, Cha03, Dud06, GS00b, Hec07, Hol00c, Laz07, Mar05, Mas01, Mil08, Mor03b, Omm01, Pap05, Pap00, Tha00, dL05, Hol06, Tha06]. Books [BALV03, Lut00, Lut01]. Bookshelf [BALV03, DFL00, LRO02, Lut02, Lut03a, Lut03c, Lut03b, Wil00b, Wil00c, Wil00d, Wil01b, Wil03a, Wil03b, Wil03d, Wil03c, FMHH^{+}00, Har02]. Borland [Ano00m, Ano00n, Ano01m, Ano03c, Ano05c]. Borneo [Dar01a]. Bose [GKMZ04]. Boston [AGG02]. Both [OBr05, Ano04g]. Bottleneck [BGED04, BWW^{+}03]. bounded [Rob00a]. Bounds [QHV02, Ano02j, BWLR06, LGFM05]. Bourne [Ano00k]. Bradenbaugh [Ano00c]. Braille [AJB^{+}04]. brain [ZAVT03].

Branch [LBJ02, LBJ05]. branch-target [LBJ05]. branches [LTOT07]. Brand [Lut02]. Brand-Name [Lut02]. Brave [Ano03d]. breadth [Ano05o]. breaks [BAL^{+}01]. Breeze [Ano02t]. brew [Ano03i, Ano03-46]. Brewing [Ols01]. Brian [Cha03]. Bridge [ASS03, Ano02p, HR00, Men03, Ano04c, Ano04r, Ano01i]. Bridges [Ano04f]. Bridging [ACM04, Tre05]. Briefs [Gar00, Lea00b, Pau01, Pau03]. Brightest [Lut03b]. bring [Ano05o]. Bringing [Moo02, UCJ^{+}04]. brings [Ano05k]. Bristol [Ano01h]. Broadcom [Ano00m, Ano03-36]. broaden [Ano04-27]. broken [Mil09, SC08]. Broker [HR00]. Brownian [GKW04]. browser [Ano03-36, Lab09, NM02, YCIS07]. browser-based [Ano03-36, Lab09]. browsers [Ano03e]. BrowserShield [RDW^{+}07]. Browsersoft [Way03, Wil04b]. Brucke [Ano04c]. BSP [GLC01]. BT [VV05]. BT-Crowds [VV05]. BTB [LBJ02]. Bucks [Ano00k]. budding [ML07]. Buege [Cha03]. Buffer [LBJ02, SK04, GSHO06, LBJ05, Rob00a]. Buffering [BCS07]. buffers [Ano03k]. Bug [Ano02o]. Bugs [Lut03c]. Bugzilla [PL03, ZK05]. Build [Kro00a, LRO02, PH00b, VHL01, Ano03-30, Atk00, Cla04, SML06, Way03]. Building [Ano04f, Bar02a, Cal00a, CI01, CKC^{+}02, CLM^{+}09, CK05, DBC^{+}00, GW00, Lut03a, Mar02, McL02a, Met01, Pet03, Rem01, Rod01, RS00b, SSM03, San02b, She01b, TOG^{+}05, Ano03l, Ano03w, Apt02, BDFL04, BVD01, DAK00, Fre07, Gro02c, HF06, HPB^{+}00, Hig03, Hub02, JF06, LS00, MBED06, Mor08a, Mur00, NP03, Pas04, PNKN04, SFMH01, ZABL09, HD03c]. built [Ano04f]. bulk [BDT01, RD06]. Bumgardner [Che05]. Bundles [Jac01a]. Burke [Fox01c]. burned [LAHC06]. Business [Ano00k, Ano01h, Ano01l, Ano01o, Bar01b, CI01, Lyk02, NSI03, Wan03a, Ano05i, Joh00b, KNN^{+}01, Lex02, AK01]. buys [Ano05c]. Byte [Cas02, HS02a, LTOT07, WS01c, WHW01, BCR03b]. Byte-code [LTOT07, BCR03b]. Bytecode [ADDZ05, ABH^{+}01, BBDT02, BDT04, BFG03, BD02, CN03b, Coo02, FM03, GH01, GH03, GPF05, Gam03, GS05b, GK08, KC00, KW03, Kle05b, KK05, KK04b, LN04, Ler01f, Ler01e, Ler02, Ler03, MH02, Nip01, Nip03, OKN02a, OKN02b, OKN02c, Qui03, Ros03, RW03b, SMBZ07, SD01b, SW01, SS00a, SS03, SSE05, TSDNP02, TSCI01, TCC01, ZXNH02, Ano03-31, A^{+}01, ABF03, BDLM04, BDL^{+}08, Ber00b, CFL05b, CFL05a, CY04, CSCM00, Cog03, Cog04, CMS07, EKEL01, GPF08, JCOP07, JPB^{+}08, KBV08, KR01a, Qia00, SV05, SS02, SD03b, VDMW06, WR08, Wil02]. Bytecode-to-.NET [LN04]. bytecode-to-C [JPB^{+}08]. bytecodes [TCC02].

C

[Ano00j, Ano04e, Che05, GF01, Gla06, Pap05, Pla00, AC01, Ano01h, Ano01k, Ano01m, Ano01o, Ano03-44, Ano04-30, Ano05k, Bat04, BA08, Bru05b, Bru04c, BSPF01, BSB^{+}03, FCHE02, G^{+}01, GK03, Gho04, HS01, Hin02, JPB^{+}08, Kic04, KW01b, Kum04, Kum05, LS04a, Lin01, Men03, MAJC03, Mul00, NNS03, Nil05, Oiw09, PZ00, PWH00, PM01b, Pon03, Pre03, Rei00b, Rei00c, SH03, SML06, SCBH09, Sib00, SHHS04, Ste00, SM04b, Stu07, TM07, Ten00, TP02, Tre05, Urb09, VKB01, VP05, WSP02, Wil06, Wit05]. C# [SKS08, Ano03w, Ano04f, Ano04g, Ano05b, Ano05k, Bar01a, BHW05, BHP^{+}01, BS04, BFGS05, Bro09, Bru05b, Cro01, DLE06, Ead01, G^{+}01, GS05a, GK03, Hun03a, KPPÉR06, Kic04, Lip01, Lut03a, Reg02a, Win04]. C/C [Pla00, Ano01m, Lin01, Sib00, Tre05]. CA [ACM00b, Ano00b, Ano00c, USE00a]. Cable [Ano00m]. Cache [CS06, Jol01, RHR02, Sch04c, Oi05]. Cache-conscious [CS06]. Caching [BR01c, ET01, WPN08, ET07, LR05]. Cactus [HL02a, PL03]. CAD [Ano00n, MD00]. Caja [Pot08]. Calculation [RGN07]. Calculi [BGZ00]. Calculus [Kle05a, RWH01, Ste04, ALZ01, BP03a, GK07, IPW01]. Caldera [Ano00i]. Calif [ACM01b]. California [Ano01g, USE00c, USE01c, USE02]. Call [DEK^{+}03, Dmi04, RKG04, Ano04i, Ano05n, Har01b, LYK^{+}00, MCD09, SHR^{+}00, ZR07]. Calling [Pon03, BM07, ZSCC06]. calls [BBG04, FF08, Och09b, ZFA00]. Cambridge [Ano03b, Cha05a, Che05, Gla06, Pet06]. CAMERA [NR05]. Cameras [VUPB02]. Can [Ano04r, Ben00c, BD01c, Cal00b, Gso00, Jen00a, Jol01, KKO02, Kie01, Kie02, KS07, Lai08, Mos00, Pet03, Reg02a, Sea02, Smi01b, Wra01, Ano04q, Hoh03, IN09, SC08, Ano02p]. Canada [Jac04b, LL08a]. Canceled [Coc02]. Candidate [NIS00, SL00]. Candidates [Dra00]. Cannes [AJ01a, AJ01b]. Canoo [Way05]. Capabilities [Cal00b, KAN^{+}03, Ano04-27, TS09]. Capability [HD02]. Capability-Based [HD02]. Capacity [Ano01o, CSFS00]. Capture [SCFP00, Sur01]. Capture/Replay [SCFP00]. capturing [LL01d]. Car [Fri02]. CARA [Sta04b]. Carbopolis [EXA^{+}05]. Card [ACL03, Ano03-28, Bec01c, BCE^{+}01, BML01, CMG^{+}01, CHS01, Cas02, DJ00, DMP05, ÉJD01, Fre05, HdJ01, HP04, KJ02, KM01, Ler01f, LS03, MdB01, MK01, Siv04, Ste04, TRVH03, Ano01p, Ano02v, AJ01b, DJ02, HM01a, Has02, LZ04, BM03, Ano00o, ACC^{+}01, BKH02, BL03, Che00, Eng00, HOP04, HP04, Mos05a, Mos05b, Req03]. Cardiff [Ano01i]. CardKt [GN01a]. Cards [AJ01b, BJvdB02, DJLT01, GN01a, WVE^{+}00, Ano04h, Ano04-28, AJ01a, Ler02, Ano02v, Ano03j, Che00]. CardS4 [GN01b]. care [Ano03j, LSK^{+}02]. careers [PB06]. Carl [Fox01b]. Carlo [GKMZ04, PFJ05, War02]. CartaBlanca [VDPC01, VDPC03]. Case [BCMT03, BS04, BL03, CQX^{+}09, CK05, DFL00, GGG03, HWB03, Hui02, KMSL03, MORW04, NW03, RZW01, Wan03a, BS00b, BS01, CCK^{+}08, CHL^{+}00, DAK00, ER09, GEVZ09a, HJvdB01, KPPÉR06, KBV08, Man01, Roc01, Utt06, VZGE07, VP05]. Case-Based [GGG03]. Cases [SGV04, BG05]. CAT [LS03]. Catalyst [Ano03-37]. Catch [MRB06, AH03]. Catches [Bar01b]. caught [HBM^{+}02]. Causes [RCR06]. cavity [PC03]. CBL [Gel00]. CC4J [KA02]. CCJ [NMKB03]. CD [Ano00h, FMHH^{+}00, Hal01a, Har02]. CD-ROM [Hal01a]. CDK [SHK^{+}03]. CE [Ano01j, TCM^{+}00]. cell [AZ02, MLVB05]. cellular [FW02]. Center [ACM00c, Ano02i, BL04, Lan04, Yua04]. Center-of-Gravity [BL04]. Centered [AF03]. Central [Ano00i, Ano02n, GKW04]. centralized [AHN02]. Centre [IEE03a]. centric [DV07, SHM09]. Century [Ano00j]. CEO [Ano04i]. Certificates [CMG^{+}01]. Certification [GH00, HS00a, BS00a, MMU04, MR00b]. Certified [Ano00d, CR02a, DDF^{+}03]. Certifying [SS03, CLN^{+}00, MSLL07]. Cg [Ano03-39]. CGI [Han01, HL02b]. Ch [Wan02b]. Chain [War02, Mau02, WSP02]. Chains [RKG04]. Challenge [CM04, KPH^{+}09, Lut01]. challenged [Kro00a]. Challenges [Bar01c, JKW03, KNN^{+}01]. Challenging [DFL00]. Chameleon [SVY09]. Change [RST^{+}04, RCR06, BDN05, GJ09]. Changed [McG03b]. Changes [DHRH05]. Channel [SRJS08]. Chaos [DFL00]. characteristics [PJ05]. Characterization [DS09, IEE02b, RVJ^{+}01]. characterizations [GS00c]. characterize [LJN^{+}00]. Characterizing [SSGS01]. charts [PPJ03]. Chat [BLW00]. cheat [HBM^{+}02]. Check [HD01, KKN00, QHV02, Cha06]. Checked [Gol01, KN06, PWH00]. Checker [Lut03c, SSE05]. Checking [BFG03, BD02, BDLM04, CH02, Dar07, DMP05, FF08, GV02a, KM04a, Nel04, PDV01, SL01, Ano02j, BK08, BS07, BWLR06, BA07a, DNS05, Dil00, FLL^{+}02, FFLQ08, GV02b, GV04, HP00, Hor00c, RHDB08, SV05, Sto02b, WGSD07, XJC09]. Checkmate [PWH00]. checkpoint [Eng06]. Checks [CC03, LGFM05, SB07]. Chemical [Guh07]. Chemistry [SHK^{+}03]. Chemo [SHK^{+}03]. Chemo- [SHK^{+}03]. Chianti [RST^{+}04]. Chicago [ACM05, Ano02i]. Chip [Ano00m, Won03a, Ano03-36, Ano04h]. Chipkarten [Ano04h]. Chirp [XM06]. Chockful [Coh04]. choice [Pay04]. choose [Ano04g]. CHR [Sch04d, Wol01a]. Chris [Azi06]. churn [SAB08]. CICS [Ano02a, BCCN01]. CIM [AZ02]. ciphers [MWM01]. Circuit [MLG02a]. circuits [JMS02]. Cisco [Lut02]. citizens [Ano03j]. Civil [SG03]. Cjj [TP02]. clamping [Ano03j]. CLANS [FL04]. Clara [ACM00b]. Clashes [HT03]. Class [Aki02, BC01, Bet04, BHP^{+}01, Gro02a, HR00, HT03, Hui02, KJ02, KS02a, KS01b, Men00, NLC03, PKF03, PP02c, RE01, Roe00, RMR03, RMR04, SLPO02, TH02, vdBJP01, AK09, Bee04a, Dur02, ET05, Fek02, Gad03, Hig03, HJvdB01, JK00, PZ00, PvdBJ01, PT09b, QGC00, ST00a, WBF^{+}06, Wor02]. Classbox [BDN05]. Classbox/J [BDN05]. Classes [All00e, AČMN05, Ano02n, Bac01, DeP03a, DTD04, Gut00, HD03a, HRD07, HRD08a, MPG^{+}00, vD04, Bac03, CLCM00, DHS02, Fau02, Fek08, HRD08b, LY03, MT07, Mey03, NW02b, QM09b, Ton04, Top02a]. classfile [Ano02u]. Classfiles [FC01, FS03b]. Classic [Bud01, CLZ06]. Classical [HS01, Pap05]. Classics [Wil00c]. Classloaders [FC01]. ClassLoading [PC04]. Classroom [HSSC05, Bow07, CL08, JMS02, KM04c, RC04, UCJ^{+}04]. CLDC [RTVH01]. ClearSight [Ano03-35]. CLI [Vog03]. CLI-based [Vog03]. click [Swa01b]. Client [Ano00k, HKM^{+}09, ML09, Ano04u, BHJR05, HKS^{+}07, JS01, KJBH^{+}00, KL07, KWM^{+}08, LHFL07, New01, Sha02]. Client-based [ML09]. client-server [LHFL07]. client-side [Ano04u, JS01, KL07, Wea07]. client/server [KJBH^{+}00, Sha02]. clients [HG08]. Clinical [TA04, VWS^{+}05, MF03]. Clock [BCHP08]. Clock-directed [BCHP08]. Clojure [Hal09]. clones [HKI08]. Closed [Ano04i, Les03]. Cluster [Ano00i, AFT^{+}00, BP01b, Gou01, HS00b, HRAB05, JM00, KMSB08, TTD03, WC00a, ZYZ06]. clustered [LR05]. clustering [GGL^{+}08]. Clusters [AFT01b, BF02, Dek00, FDTL02, ZYC03, FWL03, LP01a, ZLG08]. CML [WMRT^{+}05]. Co [WP04, Ano01f, KTV^{+}04, YLW08, ACM01c]. co-location [KTV^{+}04, YLW08]. co-operate [Ano01f]. Co-Routines [WP04]. Coal [RYD^{+}03]. Coalgebras [JP03]. coallocation [CS06]. Coarse [DFA03]. Coarse-Grained [DFA03]. COBOL [Ano04-37, Ano01l, Ano04o, Hor00a, Hor00b, Gla06]. Cocoon [For04b]. Codagen [Ano03-39]. Code [Ano00n, Ano01l, Ano02o, Ano02q, Ano05k, Bar03b, Bet05, BR06a, BHP^{+}01, BKLS00, BKLS01, Cas02, CDFR04, DDF^{+}03, Dmi04, FMR05, HS02a, KSK04a, KNY03, KA02, KK04b, Lai08, LBJ02, Lin03b, Mos00, SLPO02, Sea02, TYS04, TRVH03, VMMF00, WS01c, WA04, Wol03b, AY05, AY07, Ano04i, Bad00, BK08, BP01c, BDLM04, BCHP08, BCR03b, Dep03b, DC03a, DNR06, EvG04, Eub05, Gib09, GM05a, HTSW07, HKI08, ACM03a, LTOT07, LHGM09, LBJ05, MLVB05, New01, NAR08, PFJ05, PV08, RM07b, SML06, ZK04a]. code-copying [PV08]. CodeGuide [Ano02p]. Codemesh [Ano01i, Ano01k]. Coders [SAFG03]. Codes [LRSW00, RCB01, WHW01, LRW01, RCB03]. CodeWarrior [Ano00m, Ano02p, Kro00b]. CodeWeavers [Ano03-41]. CodeWizard [Ano00j]. Coding [AA02b, Hec07, Hol06, Hsu01, Laz07, Lou05, dL05, Ano05o, Ano05q, Lan04, Mur05]. coffee [BAL^{+}01]. CoG [vLH05]. cognitive [BS01]. cohesion [ML09]. ColdFire [Ano04b]. ColdFusion [Ano02t]. Collaboration [Ano01l, BC07, BF02, SEGS03, OOOiM05]. Collaborative [Che03a, CKKH03, Fox00d, SL04, JHSL03, OOOiM05]. collecting [CO04]. Collection [Ano03-41, Ano04l, PUF^{+}04, PP02c, SGF^{+}02, SHB^{+}03, ZT02, Bac07, BCM04, BALP01, BALP06, CSK^{+}02, CLN07, Fek02, HBM^{+}02, JMP09, LH07, PHV07, WK09, XSaJ08b]. Collections [All00c, NW06, NW07, PKF03, Wic03, Ano03h, Col01, FTD03, SVY09, WB01, Zuk01]. Collective [LCFkL05, NKBM01, NMKB03]. Collector [BCR03a, DKL^{+}01, MJ06, SLC03b, ZS01b, BAL^{+}01, BBYG^{+}05, DKP00, GSaC05, LP01b, LP06, WK08a, WK08c, WK08b]. collectors [MSLL07, SMTZ09]. College [Bar00a, CKMP09, Bar01b]. collision [XAN07]. Colorado [USE00d]. colour [MM04]. colour-map [MM04]. column [Hun03a]. COM [EK01, Gso00]. Combination [JKJ05]. Combinatorial [RM08]. Combine [NLFA02]. Combined [KW02]. Combining [BD02, NM02, Tho03]. Comes [LD03]. command [SW06]. commands [San00]. Commarea [Ano02a]. Commentary [Zus03]. Comments [Bee04a, NLC03]. Commerce [Che02b, IK04, Kro00b, LLMK03, Wea04, Che02b]. Commercial [HKHK03, Oes01]. Commit [BR01c]. Commodity [vLGL^{+}02, GGL^{+}08, vLFGL01]. Common [Bec00a, Bec00b, Cro01, Hun03a, Rob04c, Way03]. commons [O’B05, For04b]. Communicate [JPJ05]. Communication [Ano00k, Ano05a, CHK00, NKBM01, RWL07, SCLV04, SCH05, YK03, HPB^{+}00, LC05, LCFkL05, NMKB03, Oes01, WK08d, WC00b]. communication-oriented [HPB^{+}00]. Communications [Ano00j, Ano00n, Ano01i, GP01, Lut03b, Ano03k, GvLPF01]. CommuniGate [Ano00i]. communities [ACM04]. Community [Dob01a, Aar06, Ano03o, Gar09, PPJ03]. Compact [Ano03a, Gro02a]. compaction [KP06, WK08a, WK08b, WK08c]. Companies [Gar00, Ano03f, Ano04f, Ano04g]. companion [Fla00, Fla04b, Goo01b]. Company [Ano04-37, Ano05c]. Compaq [Ano00h]. Comparative [KX04, LAT04, SKP^{+}02, Ano04e, Ano04-30, Gho04, Mau02, SH03, SCBH09]. compare [Ano02j, KW01b]. Comparing [Dor02, Hir00, KPPÉR06, PE06]. Comparison [BW03a, BW03b, Bro05, CE01, DBH04, HJR^{+}03, MMG01a, NNS03, Pot04, Pre00a, Fre01, GPW05, JKH^{+}04, Nam08, RJGH06, STB08, SH04b, SC01b, TAW03]. Compatibility [Egy01, RFZ08]. compatible [VVG^{+}05]. competing [LOW09]. competition [BVPE06]. Competitor [Win04]. competitors [Ano05m]. Compilation [ALZ02, ADDZ05, Ano03-38, BJK07, CKK^{+}04, CCF^{+}02, DJP02, Lag03, SSM04, TP01, BGH^{+}07, CO06, CHP^{+}08, GEB08, KBV08, LST02, LYM04, MSR09, NW02b, OOK^{+}06, SYN03, SYN06]. compiled [NM00]. Compiler [ATBC^{+}03, Ano01i, Ano01l, BA01, BK01a, BRBY00, DFA03, GM00, GMM00, Hol00b, KMEA04, KNG02, LST03, Mid01, MF01a, ME00b, MMG01a, NP01, NCM03, OSM^{+}00, PVC01, Rob01c, SS03, Str02, SYN02, TOG^{+}05, YLL^{+}07, vdBJ01, AP02, BC04, CMLC06, CLN^{+}00, CL08, DGMY06, EH07, FKR^{+}00, HKS^{+}07, HKM^{+}09, IKN03, IKY^{+}00b, IKY^{+}00a, ITK^{+}03, Jia04, JPB^{+}08, KN06, KWM^{+}08, LOW09, LYK^{+}00, MGM^{+}06, OOK^{+}06, Oiw09, SL07, SBMG00, Siv02, SYK^{+}01, SYN03, SOK^{+}04, SYK^{+}05, SOT^{+}00, THL03]. Compiler-Cooperative [MF01a]. Compilers [NIEH04, Sch03a, SSM04, dSC06, CHP^{+}08, LMK08, SYN06, WB00, XM06]. Compiling [ABH^{+}01, Bot03, BK05b, CiLH01, PH02, SBCK03, SS02, A^{+}01]. Complement [RW03a]. Complete [DD02a, Edw00, Pew00, PL05, II04b, LO00b, LJN^{+}00, PS01, Sch01, She01a, Tay02, WMM04]. completed [VLMO09]. Completeness [SS03]. completion [KR01a]. Complex [McG04, PG00, Cog04, Ear03, EKVM07, Jam01]. Complexity [Ano04j, CRL01, DFL00, GPS03, Ano04r, Chr05, Sub08]. Compliant [Ano01l, Ano03-38, BFS^{+}04, CF00, Goo03b, TP02]. Component [AR03a, AA02b, Ano03-41, EK01, Hal02b, Hei01, HT03, Joh00a, KMSL03, KM02, KS02b, MS01, NT01, ONRV08, Ren00, RAC^{+}02, SC07, TEM^{+}01, TFL^{+}04, VDPC01, Ano04a, BCL^{+}06, GW01, JS01, LS06, PSS01, Rou02a, Sha00b, SGK09, TM08, VDPC03, WML02, Wit00]. Component-Based [AR03a, KM02, KS02b, MS01, Ren00, TFL^{+}04, SGK09, VDPC03, Wit00]. Components [Ano01n, BH03, CV01, Gso00, HRE^{+}05, Hyu05, LRSW00, NK03, SSS02, Tul02, WCD^{+}01, ZX05, Ano02w, Ano03-30, Ano03-35, Git00, JF06, Joh00b, KS09, LRW01, LHS03, LSW07, MFH01, PHM^{+}01, TJ00, Tre03, VMWD05, WF04, YKB02]. Composing [BLW09]. Composite [YE04].

Composition [PKF02, WCD^{+}01, KS09, NQM06, SRW^{+}00, TM08, dM04]. Compositional [ADDZ05, BR06b]. comprehensibility [HCMM00, SH04b]. Comprehensive [ASCE03, Goo02a, QHV02, Gos00b, LO03a, MR00b, NM02]. Compression [Bar00a, CKV^{+}02, Pau03, SMBZ07, CKV^{+}03, CSCM00, Coo05]. Compressor [KP06]. Compromise [Lai08, RFZ08]. Computation [Ano01n, CKK^{+}04, CBD04, NZ01, SvR01, TC03, FLWW04, Nor00, PT09a, vRKS01, vRKS03, SM07, Tra00b]. Computation/Compilation [CKK^{+}04]. Computational [DFT03, Lut01, RCB01, SM07, Thi02, RCB03]. Computations [KT01b, GS04, NNS03]. Computer [ACM00b, ACM01d, Ano00h, Ano00i, Ano00j, Ano00k, Bar01a, Bar01b, CCR00, Coc02, GKM03, Ges07, GS08, HMRM03, Hsu01, Kog04, LH02, Lut02, MDS04, Rob04b, Sav01, SG00, SdSK05, XX05, ZG04, AWS^{+}09, BC07, BR02, BS01, CFGL05, CKMP09, CF04b, DW07, FFB^{+}00, FCHE02, Fro07, Gol04b, Hel07a, Ibb02, Juo07, KMR02, ML07, MJ00, Rad06, Ras00, Rio02, Rob04c, RVZ04, Sco02, SSC00, TCF^{+}03, Tre02c, VVV04, Ano01h, Ano01k, Ano02o, Lut02]. Computer-Aided [ZG04]. computer-assisted [Tre02c]. Computers [BB03, Roj00, SPS^{+}02]. Computing [ACM00c, ACM01c, ACM04, ACM06, ANN01, Art00, Azi06, BC00, Bar01b, BP01b, BBHL01, BGadH06, CM01, CCFG00, Cha00a, CLL03, CT00, CSK00, Fox03a, GK03, GP01, GSC^{+}00, GMM00, HS00b, HRAB05, Hor03, HBD04, Kro00a, LBQ00, Lut01, MWL00, Mak03, NPRC01, NC04b, Pap05, PBG^{+}01, SMBZ07, Ste01, Vog03, WFGK03, Wil03b, WGW04, Woo05, Yan05, AG05, AGG02, Bar09, Cha00b, ESPP01, FJ05a, FWL03, FPA^{+}06, GvLPF01, HS01, HLT09, KHBB01, KMSB08, LP05, Lau01, LAL02, MI01, MMG00b, MMG^{+}00a, MMG^{+}02, Nau02, NC05, PSZ^{+}07, PB06, RR02, SMS00, SHHS04, TDB00, VP05, dGNv04, GS00b, Pap00]. Compuware [Ano03-40, Ano03-39, Ano02n, Ano03-36, Ano04j, Ano05c, See04]. Concept [ÁMdBdRS02, CY01b, MSK09, ST00a]. conception [FTD03]. conceptions [ET05]. Concepts [Bar03b, Bur03, JBMP03, PSS01, vLH05, Gag02, Gol04b, Hor03, NR05, Sch04a, Ses08, She01a, SCS01, SK08, SM03b, TB00b, VZGE07, ZJ03]. concepts-first [Gol04b]. Concerns [MVM07, SPS^{+}02, RM07b, WBGM05]. Concierge [RA07]. Conclusive [SGV04]. concrete [DC09]. Concurrency [DSBH03, GPB^{+}06, GSW00, IJ03, KFLN04, MSV05, RS00a, RSH01, Wel02, Zha05, BA04, BA08, Bog01, FR02, HL06, LSW07, Rob03, WJH06, Yan02, YKB02]. Concurrent [CX01a, CWY01, HD01, Lea00a, Lut03c, Meh02, MMK04, OK04, Par04a, RH04, SJG03, WHBS01, Wel04, BBYG^{+}05, Bar01d, BP01c, BFN^{+}09, Cor00, GHS05, JPS^{+}08, KP06, LHS03, LSW07, RZW01, RH07, SBAD01, San04a, San00, Sen08, WK08a, WK08b, WK08c, WCC04, Yah01, Ano01k]. Condensation [GKMZ04]. condition [Jac04a, Yan02]. Conditional [NA07]. Conference [ACM00a, ACM00b, ACM01b, ACM01d, ACM04, ACM05, Ano01b, Ano02b, Ano02i, AJ01b, Cha00a, CNB00, IEE02a, Jac04b, NIS00, SM07, SY^{+}05, SBH^{+}04, Uni01, USE00b, USE00a, USE01a, ACM06, Ano04-31, ACM00a, Fox00a, Fox00b, Fox00c, Fox01a, Fox05]. Confessions [Mil08, Tul08]. Confidence [BF03, JS01]. Configurable [RP03b, Sat04, TP01, BDRV01]. Configuration [CSK00, Han05a, RTVH01, Sin00, Ano05a, PC03]. Confined [II04a, VB01b]. confinement [ZPV03]. Conformal [Hit03]. Conformance [LBR00]. Congrès [IEE03a]. connect [Sha00a]. Connected [RTVH01, SMES01, MS00b]. Connection [Jen00b, MD00, Tre02b, Uni01, Li04]. connections [Ano02f]. connectivity [Urb09]. Connector [Han05a, Apt02]. connectors [Apt02]. Conquer [vNKB01]. Conquering [Gol00]. cons [Ano04-38]. conscientious [FB07]. conscious [CS06]. conservative [Nau02]. Conservatively [Reg00]. consideration [Emu04]. Considered [Ams02, SD08, ACFG01, Our02]. considering [Ano02k]. Consistency [AL04a, ABH^{+}00, GS00c]. consistent [WW09]. console [Rem01]. Consortium [Bar01b, DV01]. constituent [RHR02]. Constrained [RWH01, BNV08, CKV^{+}03, RA07, ZK04a]. ConstrainedJava [GNB04]. Constraint [RM04, SJG03, WS01b, Wol01a, TP08]. Constraint-Based [RM04, WS01b]. Constraints [DTD04, Sun01, Ano02u, RMR01, VTD06]. construct [SAB^{+}06]. constructed [Fle00]. Constructing [BB01, JC04, RLR00, GHBG^{+}03a]. Construction [Gar00, Hon05, Kaf00, LN04, CMS03b, Mor08a, ZR07]. Constructive [Stu01, Boe05]. constructors [SI09]. Constructs [Won04, LS08c]. Consumer [Ano00i]. Consumption [BCR03a, SKS03, BNV08, FFB^{+}00, VED07]. Contained [Ano03a]. Container [HRD07, HRD08a]. Containers [Hin02, WP00b]. Contemporary [Lut03b]. Content [Ano01m, Men00, Rap03, SLB^{+}02, Fer07, Lot02, Tho03, ZJ03]. Contention [XSaJ08a]. Contention-aware [XSaJ08a]. Contest [Bar00a]. Context [ABM^{+}03, Bar05, BML01, CHS01, DJLT01, vLSM01, BM07, LH08a, LPH01, LPH06, SM01c, SB06b, Tro04a, Tro04b, WM00a, ZSCC06]. Context-Aware [Bar05]. context-insensitive [LPH01]. context-sensitive [LH08a, SB06b]. context-sensitivity [LPH06]. Contexts [JMSG02]. contextual [TM08]. Continuing [Coc02]. continuous [TCC02]. contours [Nik03]. contract [XJC09]. Contraction [PH02]. contracts [FLF01, GHBG^{+}03a]. contribute [Ano04i]. Control [Ano00j, Ano01h, BH04b, BALV03, BP05, BW03a, BW03b, CHHC04, DS00c, HD02, Hol04a, HBD04, JC04, KK03a, Kog04, LH03a, MD00, NMH^{+}02, OWR04, PDCL02, SDPM04, Sur01, Tim03, ZD02, BWLP01, BHV01, BHR02, CVW03, DPT^{+}02, FJ05a, FR02, GB01, HCMM00, HO03, HO07, HB08, LZ04, NC04a, PSZ^{+}07, PH00a, RPB^{+}09, WSVX03, YL03, YKB02, ZP03, dM04]. control-flows [dM04]. Controlled [NAR08]. controller [AZ02, XM06]. Controllers [New04]. Controlling [Ano03e, BCR03a, BALP01, BALP06, Kro00a, Pot08, BDN05]. controls [Hu03]. Controversy [Bru04b, Bru05a]. Convenient [BKŁ01]. Convention [ACM00c]. conventions [DC03a]. convergence [BD01b, GEAS00]. Convergent [Hub02]. Conversion [Lik04a, AC01, Ano03-36, Lik04b, YTY00]. Convert2Java [AC01]. converter [Kil03a]. Converting [DKTE04, vD04]. Cookbook [Ano00d, Dar01c, Dar03, Hol04c, BC03, Dar04, EL09, Goo03a, Goo07, Mil05, O’B05, Per04, Sig05, Ano00c]. cool [Ano04-29, Eub05]. Cooling [GKM03]. cooperated [TCSC04]. cooperation [BVPE06]. Cooperative [BCM05, MF01a]. Coordination [ABM^{+}03, BGZ00, CGR00, DGGD08, WK08d]. copies [XAM^{+}09]. Coping [ABV00, San04a]. Copolymerization [BD03a]. Copying [HM01b, Oga09, PV08]. Coq [ACL03]. CORBA [ASS03, BVD01, DLL03, Des01, Die01, DHR^{+}01, EF02, EK01, GCARPC^{+}01, Hou00, JHSL03, KSK04b, LRSW00, LRW01, MSR03, NMH^{+}02, P^{+}98, Rao01a, Rao01b, RJFG03, TEM^{+}01, Won05, ZYC03, Zhu03, CSFS00, SAWW01]. CORBA-based [SRW^{+}00]. CORBA/Java [DLL03]. CORBA/Java-based [DLL03]. Core [ACM01e, Atk00, Bag02, Edw00, Edw01, GH07, Gle02, Hal00, HB01, Hal01a, HC00, HC01a, HC02, HC03, JR05, Lut03c, MP01a, Muc02, Top00, Top02a, TVMB03, WBS01, ALZ01, BP03a, CMP^{+}07, HN00, IPW01, SCB09, SSP07, WBF^{+}06, ZSZ^{+}09, GH04]. Corel [Ano03-41]. Cores [AAA^{+}04]. Cores-Based [AAA^{+}04]. Corfu [SM07]. Corner [Bro03b, Cha00a, BG05]. cornering [PWH00]. Corpora [CHHC04]. Corporate [Bro00, HAL02c, Bar03a]. Corporation [Ano00h, Ano00i, Ano00j, Ano00k, Ano00j, Ano01h, Ano04-29]. Corpus [Wei01, Mas00]. correct [AAD^{+}07, BBA08, CY01b]. Correcting [HMRM03]. Correction [BHP^{+}01, TEM^{+}01]. Correctly [Coh02]. Correctness [BRL03, DJ00, DJ02, Fre05, KC01, GHBG^{+}03a, GHBG^{+}03b]. Correspondence [BDJdS02, Mur05, Rei00c, dL05, Hec07, Hol06, Laz07]. Cosimulation [Ano03-38]. Cost [SSM04, NSI03]. Cost-Effective [SSM04]. Costs [RWC^{+}03]. could [Ano02l, Ano04u]. Counter [PDV01]. Counter-examples [PDV01]. counterevasion [MV09]. Counterpoint [Hor00a, Hor00b]. Counters [Ano03-40]. counting [JMP09, LP01b, LP06]. Coupled [VDPC01, PK00, VDPC03]. coupling [CD08, KKG09]. Course [BLPV04, CWH01, DD02a, DK02, Edw00, Hal01a, Hei03a, HTY^{+}03, LS04b, Pew00, And02, Bar01d, BZ07, BVPE06, CKMP09, CR02b, GEVZ09b, Gou06, LO00b, LO03a, LP05, LHS04b, Mau02, Moo02, MB05, PHBM05, RVZ04, SC01a, SL07, TBM09, Wan02a, ZJ03, ZCR^{+}06]. Courses [ES05a, JT04, SS07, DV07, ES05b, ET02, GEVZ09a, Hel07a, HKF00, MS05, VIPCUF08, vTNC08]. Courseware [JWC03, DUK02, Hel07a, JFH00]. court [Ano03z]. Coverage [KA02, VMWD05, Gat03, SM01d]. Covert [Kal04]. COW [BMR02]. CPU [Ano02c, BH04a, BH04b, HB08]. CPU-Management [BH04a]. CPU/DSP [Ano02c]. CR-2000-210329 [Nat00]. craft [Way05]. Cram [Ano00d]. crash [SC01a]. Crawford [Ano00b]. Create [LAB^{+}00, Esq04]. created [Ano00g]. Creating [Bro02a, BKLS00, BKLS01, Fer07, Lew00, Mey03, SGF^{+}02, Wal03a, HP02, Och09b]. Creation [Ano01m, Ano03p, ABL07, Bos04, FTD03, Tre02c]. Creator [Ano04-35, Sur04b]. Cresce [Pel03]. CRF [MS00a]. crickets [XM06]. criteria [VDMW06]. Critical [Gar00, Bro07, San04a]. Criticality [CW04a]. critics [Ano05h]. CRL [vdPE02]. Cross [Ano01h, Ano02o, Ano02q, BSMV09, JR02, Gri02b, ITK^{+}03, II04b, Och09c, OOOiM05, WK08d]. Cross-Architectural [JR02]. Cross-Platform [Ano01h, Ano02o, Ano02q, Gri02b, ITK^{+}03]. Cross-profiling [BSMV09]. cross-project [OOOiM05]. cross-reference [II04b]. cross-runtime [WK08d]. Crosscut [Kic04]. Crosscutting [MVM07]. CrossOver [Ano03-41]. Crowds [VV05, VV05]. Crowds-Style [VV05]. Crowned [Bar00a]. CRUD [STB08]. Cruncher [Mak03]. crunching [Wil05]. Cryptographic [WBL01]. Cryptography [LDM04, Gal02, SJ05, Wei04, Bis03, Hoo05, Nis03]. Crystal [Ano00j]. CS [DHRH05, AF03, Bru04b, Bru05a, HKF00, HM02, SdSK05, BR01c]. CS-1 [AF03]. CS0 [EBG^{+}05, Ree01]. CS1 [BCM05, Bec01a, CC02, CR02b, CLP06, CH06, Djo09, Fit09, GEVZ09a, GEVZ09b, Gao00, GL08, Gri00, Hun03b, LBD^{+}03, LH02, LS08c, LRD09, MRB06, MB05, Mur07, NSS^{+}05, Reg00, Reg02a, Reg06, Rou02a, Sch00a, VZGE07, WVMN05, WN05]. CS2 [CTLW03, CH06, Hun03b, KB04b, LM06, LH02, NM02, Reg02a, Reg06, WKB02]. CSFS [HYX05]. CSO [OJJ00]. CSP [MORW04, WAF02]. CSP-OZ [MORW04]. CSS [Goo02a, II04b]. Cup [Nis02a]. Curiosity [Way03]. Curl [Ano01i]. Current [SS00a]. curricula [Cha00b, Cha00a]. Curriculum [CBD01, BS01, CKMP09, GCF^{+}01, HM02, MB05]. curve [Mer04]. Custom [Han01, Lut03b, Roe00, Ano02e, Apt02, Wei02b]. Customizable [PKF02, CL08]. Customization [DTD04]. customized [MBED06]. Cut [LN02]. Cut-&-Paste [LN02]. Cutting [Ano04j]. CVS [PL03, ZK05]. Cyber [WWSL02]. Cybercourt [Pau01]. Cybernet [Ano00h]. Cyberspace [CF02]. cyberTech [PB06]. cyberTech-ITEST [PB06]. Cycle [AH04b, Gat03, KS09, LH07]. cycles [MT07]. cyclone [Mor03c].

D [MD00, Ano01o, Ano02m, Bar00c, BDRV01, BBGP01, BE02, CWWS03, CN03a, Che03a, CF02, CE01, FMA02, GV05, GP05, Hit03, HJF06, JLV02, JHSL03, MD00, MCLC02, Nik03, PFJ05, Sei09, SQG^{+}05, Tre03, WBS01, WWSL02]. D-Enabled [WWSL02]. D-SOL [JLV02]. D/ [MD00]. DaCapo [BGH^{+}06]. Daikon [NE04]. Dallas [ACM00c, CNB00]. Dan [Cal00a, Bar03a]. Danny [Fox01b, Fox01d]. d’applications [FTD03]. Darkstar [Bur07]. dash [Ano04z]. dashboards [BDRV01]. Data [AR03b, And02, Ano00k, Ano01o, Ano02r, Ano02t, Arm04, Bar01c, BH03, BW01a, CF03, CP01, CP04, CNB00, CD01c, CE01, Col01, Dro01b, EVS07, Fel04, Fox00d, Fox01b, Fox01d, GT97, GT01, GT04, GT06, GT10, GS04, Hec07, Hir07, HJF06, Hol06, HS02b, JR03, KC01, Laz07, Lin01, LZZ03, Liu04, Lut00, Lut03a, MD00, Mai03, Pre00b, Sah00, SK00, Smi01b, SCLV04, TGV^{+}01, TVMB03, Uni02, Vil08, W^{+}04, Wan04, Wan05, Wei02a, WL04, WP00a, Wil05, WF00, WF02, dL05, Ano02g, Ano03-29, Ano03-42, Ano04c, Aye01, BST00, Bai03, BCHP08, BDE^{+}03, Bud01, Bus02b, CFKL00, CHMB04, CZ02, CS06, CLN07, CHJB07, DJ01, EKVM07, Fal00a, Fal00b, Fek02, Fry08, GEVZ09a, HCB04a, Hub01, KMSB08, KF00, LO00a, Mad01, MR06, McL02b, MSK09, Mur05, NM02]. data [PHBM05, PRB07, Sal04, SBAD01, San04b, SML06, SFMH01, SB07, Tre03, VTD06, WSVX03, WB01, ZKR08, dCG^{+}02, vRS05, Mas01]. Data-Access [SCLV04]. Data-Binding [Ano01o, Ano02t]. data-flow [BCHP08]. Data-gathering [Fel04]. data-intensive [SFMH01]. data-member [KF00]. Database [Ano00n, Ano01i, Ano02q, Ano03-40, Bir01, ISO08, KW02, LL08a, PH03, Ree00, Rog03, Sea02, SO02, YWZ03, Yua02, AR08, AYWM08, DLL03, DFW04, FMA02, Li04, LC04, Mer00, Moo02, Gal02, Pan04, Ree03, Ric01, Sci07, WGSD07, WAB^{+}04]. databases [CZ01, Cha02, DSCU01]. dataflow [SFMH01]. datalog [dMSAV08]. DataScan [RSD01]. date [Bee00]. Datenbanken [DHMT00]. David [Ano00b]. DAVIS [HS02b, NHY^{+}04]. days [CL03a]. DB [Ano03-42]. DB2 [DHMT00, Ano03-42]. DBA [Lut03a]. DCT [Whi03a]. Deadlines [BD01c]. deadlocks [JPSN09, PRB07]. Deal [Ano04k]. Death [Nil05]. Debues [Ano03-41]. Debug [LHGM09, OS02]. debuggability [OOK^{+}06]. Debugger [Ano00i, Ano01j, Ano02n, IKKW01, RB01, ZYC03, RM07a]. Debugging [Hor00c, KY03a, KY03b, KKJY04, Meh02, MLM^{+}08, RCdBL02, SFM^{+}07, BRBY00, HRD08b, LHGM09, MKKC08, PTP07, Ste05, THL03]. Debuts [Ano02t, Ano04b]. Decaf [Bar01c]. decentralized [ML00, RPB^{+}09]. Decimal [BJvdB02, Cow01, SKC09]. Decision [Ano03-40, GKM01, PWC00]. Decision-Support [Ano03-40]. Declarative [BTV06, Cal04, DSBH03, Fab02, RS00a, RSH01, BS09, HL06, RPP07]. Declaratively [RP03b]. Decompiling [Kal04, MH02, Nol04]. Decomposing [BDL^{+}08]. decomposition [Soo09]. deconstruct [Way05]. decoupled [Uni03]. Decoupling [JC04]. Deduction [CCR00, GN01a]. Deductive [ÁdBdRS08]. Deep [LM04, TTS^{+}08, Ano05k, Lut03b]. DeepJava [KS07]. Default [Dau01, SJG03, CR06]. defects [AVY08]. defends [Ano03-34]. defense [CHMB04, Ano03-40]. Defensive [BDJdS02]. definition [BFGS05, BTV06, SSB01, SSP07]. Definitive [BGG^{+}03, Goo02a, MC04, TB02, BD03c, BD07, Fla02c, Fla06, Gar09, Hol05]. degree [TP08]. Deisgn [Ano02s]. delayed [FX07]. Delegate [Lip01]. delineation [Woo03]. Deliver [WA04, Tre03]. Delivering [JRH05]. Delivers [Ano02s]. Delivery [Ano01n, Ano08, Pra08, BI07]. Delphi [TEM^{+}01, Hei01]. delve [Way03]. Demand [Ano03f, SGSB05, Ano03e]. Demand-driven [SGSB05]. demanding [Man01]. Demise [Got06]. Demo [GM03]. demographics [Die00]. Demonstration [Kun02, Rei03, BLN06, DUK02, RRP02]. demonstrations [Ell00]. Denver [ACM01c, Gho01, USE00d]. Department [BHP^{+}01]. dependability [AAAG^{+}05]. Dependence [RH04, SF01, XC01, Zha05]. Dependencies [RAC^{+}04]. Dependency [SGK09]. Dependent [Bil03, ADR09, PG03b]. deploy [Cla04]. deployed [AVY08]. deploying [NP03]. Deployment [Ano01m, PKF02, PKF03, RAC^{+}04, TP01, AAB^{+}05, LS06, OBr05, RK02]. depth [Ano05o]. Derived [BCS07]. Deriving [HWB03]. Desarrollo [Ano04-33]. Descrambling [Lut00]. described [Hun03a]. describing [Woo04]. Description [Rei03]. Descriptors [RGN07]. Design [AF03, ASS03, ABG02, ACM01e, AR03a, Ano01h, Ano01l, Ano01m, Ano01n, Ano02o, Ano02p, Ano02q, Ano03-37, Ano03-38, Ano03-40, Ano03-41, BTS^{+}00, Bar00a, Bec00a, Bec00b, BKY^{+}03, Cha05a, CKKH03, Cim02, Coo00, CS02, CS03, DYH05, DHRH05, Dud06, DLS^{+}01, GS08, GLS02, HK02b, Hol00b, IKY^{+}00b, JJ02b, Kaf00, KT04, KSC^{+}00, KPKL03, KC01, Kog04, KWM^{+}08, KX04, Lam03, LL01b, Li04, LC04, Lut03a, LAB^{+}00, Mah06, Met02, Mil08, NW03, NK03, NSS^{+}05, Omo03, PGM^{+}05, RWH01, Rou02a, SG02, Sma07, SCLV04, SP03, SYK^{+}05, Sun01, SM02b, Sur01, TCSC02, USE00c, WS01a, WLW^{+}03, WHBS01, Wel02, WK02, ZG04, ZYC03, Ano02k, Ano03-35, AT01, BCM05, BD04, Bil03, BV05, BC04, CMS06, CK03b, CLZ06, DWH01, DC03a, DCA04, DNR06, FWL03, FFSB04, Gab07, Gao00, Ges07, HTSW07, Hun00]. design [Ing09, JMS02, JHSL03, KHMW05, Kno02, LO00a, Lan05a, Lan05b, Lea00a, LBR06, LL00, LL03, LL01c, LG00b, LFG00, MWM01, MB05, NH02, Oi05, Pan09, Pre00b, RV05, RRP01, SL07, SJ01, SSP07, Tul08, WC08, Wol01b, ZP03, Zhu04, Ano01m, Ano02q, CMLC06, CMP^{+}07, Lut03b, GS00b]. design-code [HTSW07]. design-first [MB05]. Design-Time [SCLV04]. Designed [BR01d, Ano04j, San04a]. Designing [AA02b, GHM^{+}01, Gro02c, HP02, KT00, Lut00, RM00, TGCF08, ALZ03, PC03, Sha01, Bro02a]. designs [HBR00]. Desk [Kro00b, II04b]. Desktop [Ano03-41, WGC09, AH04a, Ano00b, FFC02, Fla02a, Fla05b, HG08, OW00, Top02b, LTOT07]. desukutoppu [SM04b]. Desupport [DHR^{+}01]. detect [MP05]. detected [NE04]. Detecting [BCE^{+}01, Bog00, FJ01, AVY08, HT06, JPSN09]. Detection [Ano02o, CD01c, CD01b, AFF06, FF00, FF09, HWM01, LMK08, NAW06, NA07, PWN04, Rei05, SBAD01, XAN07]. determine [GMM09]. Deterministic [LSW08, SW01, BAD^{+}09]. Deugo [Pet06]. Dev [Ano00m]. Develop [Cha03, KSK04a, Les03, SL06, SL07, SSS02, Ano03f, Fek08, PCC00, San00]. Developed [VWS^{+}05, Ano03n, Ano03o, RM08]. Developer [Ano03-38, AM02, Bar01b, BRL03, NRV00, SH06, Ada05, Ano04-27, Bro01, GT05, Gig00, MOL05, MCG03a, MF04, RG05, Swe06, TGL05, PKC01, Cal00a]. Developer-Oriented [BRL03]. Developers [CDH07, Col02, Dar01c, Dar03, MKF06, Ano03-30, BS00a, Coh04, HG07a, HG07b, KM07, Nis03, Ses08, Wil04b]. Developing [AU02, BH04c, BBV03, Cha03, CCB09, GW01, HRD08b, LC05, Lut03c, Lut03b, Man01, Pet05, Ree02, Ric06a, RYD^{+}03, SV02, SG03, Tor01, Tul02, Wei02b, WR00, YAA07, Yua03, HG08, HL02b, Knu01b, Gal02, Pay04, Roc01]. Development [Ano00k, Ano00n, Ano01h, Ano01i, Ano01j, Ano01l, Ano01k, Ano01m, Ano01n, Ano01o, Ano02h, Ano02m, Ano02n, Ano02q, Ano02r, Ano02s, Ano03p, Ano03-38, Ano03-39, Ano05c, AGS01, Ber00a, Ber05b, Bir01, BDJ^{+}01b, Bro00, Cas02, CN03a, DF03, DeP03a, DYH05, Fab02, FK00, Gat03, GS08, Gun01, HHK^{+}01, HK02a, HF00, HTY^{+}03, HD03b, Kim02, Kog04, KW02, Kro00a, Kro00b, LL01a, Lia00c, Lin03a, MD00, Mah04b, MS01, Mor03b, Mos05a, NSI03, Pip03, SLB^{+}02, SAWW01, SSS05, SHK^{+}03, TCF^{+}03, Wan03a, Zen02, Ano03-30, Ano03-36, Ano04j, Ano04q, Ano04r, Ano04u, Ano04x, Ano04-29, ACC^{+}01, BGH^{+}06, BFMT00, BS01, BCR03b, CSFS00, DS00a, For04b, Gar09, Hal02b, Hef07, Jia00, JHA^{+}05, KS09, Lak02, LT02, LM06, LG00b, Mau02, Mer04, MF03, NSS^{+}05, OBr05, Rob00b, Tay02]. development [WWJ07, Wil06, Wis06, You02, vTNC08, HL04, Mar05]. Developments [Ano04-27, JP04]. Développement [BCR03b]. Develops [Ano01j]. Device [Ano02p, Ano03-37, MD00, RTVH01, SQG^{+}05]. Devices [Ano01j, AAA^{+}04, Bar03a, Bat03, BL02a, CKK^{+}04, Gib01, Hac01, KK05, Kro00a, SSB03, SLC03b, TP01, Tui04, dFR04, CC01, CT03, GSaC05, HAL02c, Kon03, Lea02, Pay04, RA07, RTVH01, Sha00a, Tre02b, TBM09, Whi03a, YMP^{+}05, Yua04]. devirtualization [IKY^{+}00a]. DHTML [BHP^{+}01, Fre01, Gil00b, Goo03a, Goo07, Lan05a, NLFA02]. Diagnosing [Eth01, MS03]. Diagram [CQX^{+}09, MLG02a]. Diagram-Based [CQX^{+}09]. Diagrams [AH04b, BLL06, DH04b, IKKM03, OS02, HCMM00]. Dialect [Bac01, BST00, Bac03]. dialogue [OHL^{+}05]. DICOM [PFS05, Kon04]. DicoSE [PFS05]. Didactic [FSBP03]. Diego [USE00c, USE00a]. dielectric [KM08]. Dienste [Sig04]. differences [Ano05e]. Different [BLPV04, LZZ03, Ano02k, CC02, DM07, KS09]. differential [LS04a]. Difficulties [WVMN05]. difficulty [BBS04]. Diffraction [Uni02, Ano02g]. Digital [AAA^{+}04, Bar00a, Eff00, EGST08, GMW^{+}02, Kro00a, Lin00, Lut01, Lut03c, MD00, Pau03, SBH^{+}04, VUPB02, WVE^{+}00, Ano03g, Hal01a, LYL^{+}04, Mls04, Per01, Rad06, CM02, Lut03c, SA02]. Digitizer [MD00]. Dimensional [Bur03, BW01a, WBGM05]. Dimensionality [Vil08]. dinosaur [Lab09]. diode [PC03, EBG^{+}05]. Direct [LSW08]. Directed [AHR02, BCHP08, BKO09, ACM03a, Sen08, OKN06]. Directing [KHFS09]. Directives [BKO00]. DirectJ [BBGP01]. directly [Ano03n]. directories [HW00]. directory [LS00]. directory-enabled [LS00]. disassembler [MSU08]. DisASTer [OG05]. Disasters [Lut03a]. discardable [Sto01a]. discontinuous [TCC02]. Discovering [HD03a, HRD07, HRD08a]. Discovery [DC03b, EH04, Eng00]. Discrete [Ano01n, CWZ04, JLV02, KW02, MCLC02, Gar01, PCC00]. Discrete-Event [Ano01n, Gar01]. Discussion [G^{+}01, Bru04b, Bru05a]. disequilibrium [DZHS03]. disk [Rob04a]. DisMedJava [BG02]. Dispatch [ACGL01, DLS^{+}01, ZD02, BH02b, CLCM00, MFRW09, MPTN08]. Dispatching [Fei04, Och09c]. Display [Ano02n, SQG^{+}05, AWE04, Ano03-50, CWS04]. display-independent [Ano03-50]. Displaying [ZAVT03]. Dissection [PM01b, PM00]. Distance [HL03b, SS07, SV02, ET02, LW03, MAWW^{+}01, PC08]. distance-learning [ET02]. Distinctness [PCC01]. Distinguished [ABH^{+}01]. distribuées [FTD03]. Distributed [AJMJS02, ABH^{+}01, BMR02, BBM04, BCS02, BD03b, Bet04, BCH02, Bir01, BF02, Dd01b, BM04, BLL06, BFM^{+}02a, BFM^{+}02b, BFS^{+}03, BG02, CCFG00, Cer02, CLL03, CKKH03, CGR00, Des01, DS00c, Die01, ET01, ESS02, FSS06, FJ01, FDTL02, FC01, FGLS04, FP03, FBS04, FMMd03, GS00b, GAR04, GRR05, Gun01, HR00, HRE^{+}02, HRE^{+}05, HE03, HWB04, Hyu05, IEE03b, Ish01, JLV02, JSSM04, Jia04, JPJ05, JRN00, KAN^{+}03, KGMO04, KMSL03, MB03, MSF03, MSS00, MKM^{+}06, PKF02, Par04a, PP02b, PP02a, PC08, RWL07, RM04, Sch02, SV02, SSS02, SL01, SBA01, SM02b, TSCI01, TMG03, TS04, Tor01, WFGK03, WTV03, WTV05, WK02, YE04, Zhu03, ZWL03, And01, A^{+}01, AFT01a, BDP02, Bog01, BVD01, BFW^{+}03, ET07, ESS04, FJ05a, FT06, Gro02c, GAR03, GW01, HW00, IH01]. distributed [ICB00, Jen01, Lau01, LLdA08, Mer04, MDJ05, NB00, NB01, OG05, Pap00, PV03b, RZW01, RR02, RJGH06, Sto02b, dGNv04, vHMB08, FTD03, Gil00c]. Distributing [Bar01b, McG04, PWC00, SSL02]. Distribution [Ano00k, Ano00n, Ano02o, KM01, Bog01, TS09]. Disturbances [Wat02]. DITTO [SB07]. diverse [CR02b]. Divide [vNKB01]. Divide-and-Conquer [vNKB01]. dividing [Ano05f]. DJ [OL01]. DMC [Mar01b]. DNA [Ano03-37]. Do [BH03, Coh02, Cox01a, HCMM00, HL00, Jac01a, Jen00b, Jen02b, KKO02, NLC03, PH00b, Rao02, Rei00a, Wei01, Win01, Yua02, Ano04g, Mas00, OPS^{+}02]. Document [Ano00n, Ano01i, Gal01, ISO05, Sha04, Sto01b, TMF05, YLM^{+}05]. document-level [Sto01b]. Documentation [HRD07, HRD08a, Luk04, GMM09, Hoh03]. Documents [BK01b, Tre02c]. Does [Hag02, RVZ04, Hug02, San04a, San04b]. Doesn’t [MKS^{+}03]. Doke [Gla06]. DOLFIN.COM [Ano00k]. DOM [GSWZ08, Goo02a, Har03, Lan05a]. Domain [BBDT02, HZS08, Sto02a]. Domain-specific [HZS08]. Domains [HZC^{+}04, PCC01]. Dominant [Gee05, Oga09]. dominant-thread-based [Oga09]. Domino [LZZ03, Tam00]. dotplots [BRU04a]. dotter [BRU04a]. down [Ano03j]. downtime [Ano04d]. Draft [Cow01]. drag [Ber06]. Drawing [BH02a]. dream [Rob04c]. Drive [Lin03b, BGH^{+}07]. Driven [DK03, DFL00, Pip03, CC02, DHS02, Hub02, RDW^{+}07, SPG07, SGSB05]. Driver [Ano00k, Ano02n, Rao02]. drives [Ano04-39]. drizzle [EBG^{+}05]. DrJava [ACS02]. drop [Ber06]. Droplet [Ano01h]. DSA [SA02]. DSM [ABH^{+}00, KBVP07, SNOM01, VHBB01, VHBB03]. DSP [SASZ03, Ano02c, Ano03-38, Ano03-40, GSV02, SASZ03]. Dual [EGLZ02, Ano03k, OBr05]. dual-platform [OBr05]. Duane [Zen02]. Duke [Ano05d]. Dumb [BHP^{+}01]. d’un [BCR03b]. During [DeP03a, RCdBL02, BAJ01, Gad03, JJ02a, LYC02, Uni03]. dwarf [Ano00i]. Dwight [Pet06]. dying [Pau08]. Dylan [GI00]. DynaMetrics [SS08]. Dynamic [ATBC^{+}03, Ano00i, ASB^{+}04, Bar03c, Bec01c, Ber00b, BCH02, BPSH05, CHJB07, DHPW01, Dmi04, Dro01a, DDHV03, EGLZ02, FT06, GSHO06, Goo02a, GJ09, Har00d, IKKM03, Joh00a, JCKS04, KNG02, LK01, LMK06, MPG^{+}00, MMK04, Mos05b, OL01, OWR04, Rei05, RJFG03, RKG04, SMSAT08, She01b, SK08, SSS05, SHM09, TYS04, TT01, WR08, WK09, ZD02, ZX05, ZHC04, Atk00, BCV03, BCV09, BWW^{+}03, Bro02a, BGH^{+}07, CO06, CO04, CD08, CLS00, CH06, DGMY06, DLE06, FF09, FC00, GES^{+}09, GV05, GP05, GPW03, HP02, HCB04a, JMK^{+}08a, JMK^{+}08b, JMK^{+}08c, JPSN09, LC05, MP05, MKM^{+}06, Mur00, OKN01, Pas04, PWH00, RDW^{+}07, SBAD01, SAB08, SYK^{+}01, SYK^{+}05, SYN06, Tho03, TAW03, Tre03, Wea07]. dynamic-reconfigurable [LC05]. Dynamically [BL02a, CO03b, CO03a, NM00, NW02b, NE04, WGSD07]. Dynamicity [GDC^{+}04]. Dynamics [KW02, RCB01, Vor01, RCB03]. dynamische [Ste08a].

e-AMPS [Lin03a]. e-business

[KNN^{+}01, Ano01h, Ano01l, Wan03a].

E-Commerce

[Che02b, Che02b, Kro00b, LLMK03]. e-Government [LS03]. E-Grind [Lut00]. E-Mail [Pau01]. e-payment [Has02]. e-services [SGW01]. E-smart [AJ01b]. E-Speak [AM02]. E2 [Ano03-48]. E410 [Ano00h]. Eager [KS02a, NC05]. eaLib [RS01]. Early [EM04, NW03, BWC^{+}05, CVW03, CMS06, MS05, PFJ05]. Earth [IEE03a, Wat02]. earthquakes [JJ02a, Uni03]. easier [Ano05q, Lan04]. Easing [LP01a, WM00a]. Easy [Apr05, CN03b, Esq04, GF01, Sun01, Vor01, Ano05b, Tre03]. Easy-to-Use [CN03b, Ano05b]. EBay [Ano04-27]. Echtzeit [Ano03s, Ano04l]. Echtzeit-Anwendungen [Ano03s]. Echtzeittaugliches [Ano05l]. eclipse [CT05, Fre07, Ano05o, AL04c, Bur05, Gee05, Hol04d, Hol04c, JRH05, MKF06, Pil04, WA04, ZK05]. eclipse-based [Fre07]. eclipses [Ano03-44]. Eclpss [Wen05]. economic [CC01]. Economics [Rob01c]. Economy [Lut01]. Ecosystem [San02b, Wen05]. Ecrix [Ano00h]. ed [Feu02, Mas01, Nis03]. Edge [LR04, Mar01a]. Edge-Server [LR04]. edit [Way05]. Editing [Ano00n, PH00a, SCWL08]. Edition [Ano00d, Ano00h, CI01, KC01, Yan03, For06, Gig00, KCF01, Knu01b, Lad01, Mar01a, Mil08, RTVH01, Sha00b, Wut00, Zen02, Ano02l, Ano04-33, Mer04]. Editor [Kro00b, TCM^{+}00, Ano04q, Ber06, CCSB04, DG02, KK00, THMT03, Pil04]. Editorial [Fox00a, Fox00b, Fox00c]. EDK [Ano02s]. EDO [OKN06]. Education [CQ05, EH04, EXA^{+}05, SD08, SV02, Chr00, DW07, KPN02, LYL^{+}04, Mah04a, MAWW^{+}01, PHM^{+}01, PC08, Rob04c, SSC00, SdSK05, VS06, YL03, DC09]. education-oriented [VS06]. Educational [BD04, MJ00, CHB03, NB00, NB01, Rob00b]. EE [Hef07, FLMS06]. EEMBC [Ano03g]. eEMU [Ano00j]. Effect [SR05, SSV05, BP03a, BAD^{+}09, GEVZ09a, MRR02]. Effective [AAD^{+}01, Blo01, Blo08, CSK00, FYD^{+}08, GH03, Goo02b, KKN00, KKN06, KPN02, Lew00, MFSL02, NAW06, New05, Ruf00, Sat02, SSM04, SM01d, CM05a, Cal00a, SNO^{+}07, TPF^{+}09]. effectively [Coh04]. Effectiveness [ITK^{+}03, SKS01b, Gri03, LLdA08, TE04]. Effects [BP03a, MD00, vON02a, vON02b, HG08]. Effexis [Way05]. efficacy [Emu04]. Efficiency [Ten00]. Efficient [ACGL01, ACFG01, ASB^{+}04, BFG02, BAdMS08, BHDS09, CCC^{+}04, CN03b, CC03, ET01, GH01, GEK01, HIBP04, JPB^{+}08, KY03b, KC03, LYM04, MVV^{+}01, MMK04, NK03, RHDB08, SF01, SKS01a, TP01, TS04, WP04, YLL^{+}07, vNKB01, vNMKB05, AVY08, BHK^{+}04, BDE^{+}03, CR07, DAK00, EKVM07, EGKP02, FWL03, FF09, Gam00, GSaC05, KTV^{+}04, LOW09, LH07, NAR08, OGA^{+}01, PT09a, PHN00, SMSAT08, WC00b, ZYZ06, ZSCC06, vNMW^{+}05, vMV05]. Efficiently [JMSG02]. Effort [BAJ01, KK04a]. EIC [Sak01]. Eighteenth [Uni01]. Eignen [Wol03b]. Eikonal [SGV04]. Einführung [Lex02]. Einsatz [HMD04]. Einstein [GKMZ04]. Einstieg [Ste08b]. EJB [EF02, EK01, GKM01, GM05b, LL01d, Mar01a, NP03, Rao02, SB03a, TEM^{+}01, Tul02]. EJVM [CC01]. Ektron [Ano03-36]. elaboration [KR01a]. Electromagnetic [HKHK03]. electromagnetics [CHB03]. Electronic [Bar01c, CH02, HL03b, ISO05, Lin03a, Wea04, Sha04]. Electronics [DK02]. Elegance [Ten00]. Element [KW02, MCLC02, MAJC03, NNS03]. Elements [Che05, GS00b, VAB^{+}00, Bai00]. Elevated [BD03a]. Eliminate [Bar01b]. Eliminating [RD06, Ano02j]. Elimination [KKN00, LGFM05, QHV02, ASCE03, KKN06, VED07]. Elsevier [Dud06]. elusive [Coh04]. Embarcadero [Ano02q]. embarqué [BCR03b]. Embedded [Ano00l, Ano00i, Ano01h, Ano01k, Ano01m, Ano01n, Ano01o, Ano02o, Ano02q, Ano02s, Ano03-33, Ano03-38, AAA^{+}04, BL02a, Cas02, CKV^{+}02, CSFS00, CCF^{+}02, DEK^{+}03, DJP02, DYH05, DS09, DS00c, DFT03, Fri02, JKJ05, KPKL03, KFLN04, KFN04, KMOS03, KC03, Leh01, Leh02, Lut02, New04, Nis02a, Nis02b, Pot04, SMK02, Sal06, SMBZ07, SBCK03, SK04, SLC03b, SSA03, TGB^{+}04, TFL^{+}04, Uma02, Wri03, XX05, Ano03-35, Ano03-44, BNV08, BLN06, Caa00, CC01, CG02, CSK^{+}02, CT03, CSCM00, DGMY06, GSaC05, HKS^{+}07, HKM^{+}09, Ive03a, Jia04, JPB^{+}08, LMK08, Nis03, Pel03, RTJ00, RK02, SKP^{+}02, WLW^{+}03, XM06, Yua04, Zar02, ZABL09, Ano01j, Ano02p, Ano03-33, Lut02]. embedded-C [Ano03-44]. Embedded-Systemen [Ano03-33]. Embedding [Bur01b, Cal04, CW04b, LM04]. Embedix [Ano00h, Ano00i]. Embryonic [Ras03]. emerging [LSK^{+}02, ZSZ^{+}09]. eMiner [LL01a]. EMJ [Ano00i]. emotion [Bea05]. Emphasize [JT04]. emphasizing [Gar09, MS05]. Empirical [DMP09, Pre00a, SYN02, BBS04, CMS07, CLN07, Gri03, MT07]. Empirix [Ano03-39]. Employing [DK02]. Employment [HMD04]. Empress [DHMT00]. Emulation [Ano03-37]. emulator [VVV04]. emWare [Ano02p]. Enable [Yan05, Coh04]. Enabled [CKK^{+}04, GSV02, KPKL03, MWL00, RAC^{+}04, Tui04, WWSL02, WH01, ZCQS04, Cul00, HYX05, LS00, LCFL04, RB04, Sak01, SGW01, YHL04]. Enables [MD00]. Enabling [Ano02t, DH08, Hei03a, KHBB01, PR03, Thi02, WC00b]. Encapsulation [Fle01, Rot05, TSL^{+}04, KT01a, MF07a]. Encoding [Wic03, BDE^{+}03]. Encrypting [RC01]. Encryption [NIS00, ZFK04]. End [Ano00i, Ano00k, HECR00, SBCK03, Ano03f, Ano04x, CSCM00, IK04]. End-to-End [Ano00i, IK04]. Ended [OSM^{+}00]. Energy [CKV^{+}02, CKK^{+}04, KTV^{+}04, VKK^{+}01, BNV08, CSK^{+}02, FFB^{+}00, GSaC05]. Energy-efficient [KTV^{+}04]. enforcement [GB01]. Enforcing [RW03b, SMAT^{+}07, AAAG^{+}05]. engagement [SMS^{+}04]. Engine [AGH05a, Ano00n, Ano03-40, Hab04, NM02]. Engineer [Ano00d]. Engineering [BLL06, CQ05, Cha05a, DDDM04, Fox03a, GDC^{+}04, GAR04, GRR05, Kal04, Lut03c, RKK03, SD08, SPS^{+}02, Sib00, SM07, ACM01a, BCS09, DBH04, FLWW04, GAR03, Kes04, MORW08, Nam08, Ril02, Ril03, SML06, SKM01, TMF05, Zhu04]. Engineers [Cha00c, SC02a, BB00a, Lau04, Bur02]. Engines [Ebe02, Pau03, ZT02]. English [Coo05]. Enhance [CQ05, EH04, Rob00b, SPBE09]. Enhanced [Ano02n, KPKL03, LMK08, TCC01, CMS05, CY01a, CY01b, Lan04, LJ08]. Enhancement [Ano02q, BAJ01, MFSL02]. Enhances [Ano03-39, Ano03-34, Ano03-35, Ano03-36]. Enhancing [HBD04, KFN04, KS01a, KB04a, KSK04b, Nat00, RPJ04, SE04, ST09, TS09]. Enhydra [You02]. enjoyable [Lan04]. ensuring [Req03]. Enterasys [Kro00b]. entering [SCWL08]. Enterprise [AA02b, Ano01m, Ano02l, Ano04-36, Ano04-37, Ano05f, Ano05o, Arr01, Azi06, Bar03a, Ber00a, BH03, BMH06, CR02a, CI01, Cha03, Eck02, Fab02, FCF02, FFC02, HM00, Hig03, JFt00, KMSL03, LLMK03, Mer04, MF01b, Par05, PNKN04, Ric06a, RAC^{+}02, SPBE09, Yua03, Yus04, AU02, Ano00b, FMHH^{+}00, HAL02c, LYC02, McL02a, Moo02, Sha00b, Tro04b, XLG03, XOWM06, AA02b, Ano02k, Ano02q, Ano03-37, BCCN01, BR01c, Bro02b, CMS03a, FC06, HL03c, Jor02, KNN^{+}01, LR04, LR05, Ler01a, MS01, MH00b, MH01, MH04, MHB06, NT01, New05, Nyb02, Pro01, Ric06b, RAJ02, Sch03b, TJ00, Tre01, Tro04a, YAA07]. Enterprise-Secure [Cha03]. Entertainment [Ano00h, Lea02]. Entities [JPJ05]. entitled [CY01b]. Entity [BR01c]. entornos [Ano04-33]. Entropy [GKM03]. enum [Lan04]. Enums [TCM^{+}00]. Environment [Ais03, Ano01h, Ano01i, Ano01l, Ano01k, Ano01m, Ano01n, Ano01o, Ano02m, Ano02p, Ano02q, Ano03-39, Art00, AAA^{+}04, AGS01, BC00, Bal03a, BCH02, BGadH06, BH03, BK01a, CW04a, Che03a, CR05, CSK00, CEG^{+}03, DT02, FMMd03, GHH01, GGG03, HD02, HK02a, HWB04, HL03b, LLMK03, LL01a, LZZ03, MD00, Meh02, PP02b, PP02a, RWL07, SDPM04, SAWW01, SV02, SFP03, SSS05, WK02, YE04, dBdd04, ADT03, ABLU00, ACS02, AAB^{+}05, Ano00g, Ano03q, Ano03-30, Ano03-36, ACC^{+}01, BBBD01, BHJR05, BGNM04, CC01, CSK^{+}02, CR02b, ET02, ESS04, Fei07, GCRD04, GJ04, Gol04a, HLT09, HT06, HKF00, IH01, ICB00, JCP^{+}05, KK00, KNN^{+}01, LHGM09, Man01, OBr05, Rio02, SRW^{+}00, SKM01, WCCL05, WSP02, ZYZ06, vNMW^{+}05, vTNC08, Dau01, GGHvdG01]. Environmental [EXA^{+}05, RT02]. Environments [ACM05, ATBC^{+}03, GP03, HHK^{+}01, KM02, SMBZ07, SM01b, SBA01, BE02, CKV^{+}03, KdJNNV09, KM04c, LR05, PSZ^{+}07, SM03a, ESGS00]. ENVY [PKC01]. ENVY/Developer [PKC01]. EPerl [Wit05]. Epi [FB07]. Epi-aspects [FB07]. eQ [Way03]. equals [Coh02]. equation [LS04a]. Equator [Ano01n]. equipment [Ano04-32]. Equivalence [SP03]. Era [DDDM04, GDC^{+}04]. Eric [Fox01c, Mor03b]. Errata [HRD08a]. Error [HBM^{+}02, Hol04a, KdJNNV09, RSS^{+}04, Sma07, vdSPP05]. Error-free [HBM^{+}02]. Errors [CMB^{+}01, HMRM03, KY03b, BNK^{+}07, MKKC08, PWH00]. ESC [CH02, CK05, FL01, NE04, Won05]. ESC/Java [CH02, CK05, FL01, NE04]. ESC/Java2 [CK05]. Escape [Bla03, CGS^{+}03]. eServer [Ano00i]. eServer.group [Ano00j]. Esmertec [Ano04z]. essay [Bea05]. essence [SW06, Wam02]. Essential [AE06, Ano00k, Lan00, Lut03c, ZK05, Dur02, EA06, Goo01b]. Essentials [Ana01, Cer02, PR02, WMC04, Hor03, PM00]. Establish [Jen00b]. Establishing [BWLP01, FX07, VDMW06]. Estimating [SKS03, SC02b]. Estimation [BAJ01, Kro00a, BG03, KK04a, SYAS05]. etc [CM05c]. Ethernet [Ano03-36]. EtherShare [Ano00h]. Etnus [Ano00i]. Euclidean [Hit03]. EuroClimHist [Fel04]. Evaluate [VHL01]. Evaluating [ER09, FVK01, LH08a, LPH02, LPH06, SAFG03, WP03, ZS01b, GM02, LPH01, TE04]. Evaluation [BBG04, BLW00, GSC^{+}00, HdJ01, HS02a, LHS04a, PL01b, SHB^{+}03, TTD03, Vrb03, dSC05, All03, AHN02, BBBD01, BCM05, Bel02, GBE07, GEB08, Gri03, IKY^{+}00b, LH05, MI01, MCHN05, Nor00, SH03, SZ00, SYK^{+}05, SKP^{+}02, TGO00, Zea00b]. Evaluator [Kun02]. Evasion [MV09]. even [DA04]. Evenet [GHM^{+}01]. Evening [DHWH03]. Event [Ano01n, Bru02, Che02a, Che03b, CWZ04, JLV02, KF05, dH05, CC02, Gar01, KBP^{+}03, KLS00, Pal02, PCC00, Soo01]. Event-based [dH05]. event-driven [CC02]. event-handling [KBP^{+}03]. Eventrons [SAB^{+}06]. Events [Hou00]. Everybody [Dar01b]. everyday [Wil05]. Everything [Ron01]. Everywhere [Ano00h]. Evidence [INM05]. Evidential [Lut01]. Evolution [AZ02, ESS02, JM00, SOK^{+}04, Aki02, GHS05, GBCW00, Sak01, WM00a]. Evolutionary [Lut03b, RS01, Ton04, FLWW04]. evolvable [Gra04]. evolve [OJ09]. Evolving [Lut03b, Vau03a]. Exact [CBD04]. Exam [Ano00d, GM02, HS00a, BS00a, DHRH05]. examines [Ano04-29, Nis03]. Example [BLPV04, ER01, Hal01b, JFt00, KHKH01, Lea02, Lex02]. Examples [Ano08, Bur03, Dar01c, Dar03, Pra08, Ros02b, BI07, BLN06, Fla00, Fla04a, Fla04b, Goo01b, PDV01]. Excel [Ano01n]. Excellent [Cha05b, GT00]. Excelsior [MLG^{+}02b]. Exception [Jac01b, JC04, SM04a, BS00b, JCYC04, JPB^{+}08, LYM04, Och09d, OKN01, Ste05, SC01b, ZK09, OKN06]. Exception-Directed [OKN06]. Exceptional [WN08]. Exceptions [ÁdBdRS08, AHKR01, Gol01, GCH00, SK00, AH03, ALZ01, CRL01, RM00]. Exchange [LZZ03]. Exchanging [Lin01]. excitable [FCHE02]. Exclusion [Bro05]. execJS [Sto01a]. Executable [BDJ^{+}01a, BL03, MP01c]. Executables [BHP^{+}01]. executes [Ano03-31].

Executing [CCC^{+}06, FGLS04]. Execution [ACM05, ABH^{+}01, BL02a, Dd01b, Coo02, GH01, Gam03, GR07, GPS03, HWB03, KFN04, PV04, DJM^{+}02, SW01, TSCI01, WTV03, vLSM01, AYWM08, AAB^{+}05, A^{+}01, BBBD01, BALP01, BALP06, ESS04, GCARPC^{+}01, GK05, KTV^{+}04, MR00a, PG03a, Rob07a, SM01c, XSaJ08a]. Execution-State [WTV03]. executions [NM00]. exercise [BVPE06]. Exile [Ano00j]. Existing [BDT01]. ExoLab [Ano01o]. exotasks [ABI^{+}07, ABI^{+}09]. exotic [GS05a]. ExoVM [TABP07]. expanders [WSM06]. Expansion [KK04b]. Experience [Ano01c, BHW05, CKC^{+}02, Fre07, LS04b, Oes01, Ren02, CVW03, CLP06, GCF^{+}01, LHS04b, Mah04a, SMS^{+}04, TGCF08, XSD07]. Experienced [BBL03]. Experiences [BN03, BHK^{+}04, HPB^{+}00, MKS^{+}03, TE04, dSC06, CMP^{+}07, OJJ00, SFMH01]. Experiment [CW04b, GKM03, Man01, WAB^{+}04]. Experimental [CCW02, KK03b, SH04b, dSC05, BCM05, BGNM04, OMK04]. Experimentation [Hun05, Rob00a, Rob01a]. Experiments [BR01d, GKW04, HCMM00]. Expert [Dep03b, Dob01a, VWS^{+}05]. explicit [AY05, AY07]. Exploding [YWZ03]. Exploitation [GGL^{+}08, OGA^{+}01]. Exploiting [BS04, CFL05b, DFA03, Pan09, TCC01, YLW04, ZJ03, KKM^{+}06, Lot02]. Exploration [Rob02]. Explorer [Nas04, HSD04, Way03]. Exploring [AH04a, AHKR01, BW01a, Cav02a, CF04a, CHUB08, KHMW05, CKMP09, DJ01]. Exposed [Cha03]. Express [DJ01]. Expressing [FDTL02]. Expression [Sun01, Vel01, DJ01, GV05, GP05, Stu07]. Expressions [Hab04, Hei03b, Zam03b, AOMC07, Kah06a, Mor02, SM04b, Stu07]. Expressive [CWY01, HS08, MFRW09, WP03, BLW09, SC07]. Extend [Ano03x, Cal00b, Wra01]. Extended [FLL^{+}02, KGMO04, Nel04, OK04, PC03, Ano01j]. Extender [BP01a]. Extending [BCV03, BH05b, CT03, CMS03b, HSB09, JCKS04, LPH01, LS08a, YTY00, New01]. Extends [Ano03-39, Ano03-40, Kro00b, Ano03-36]. extensibility [Gri06, IV07, MRC03]. Extensible [DA02, EH07, HWB04, NCM03, dBdd04, BFN^{+}09, BTV06, DCA04, GSHO06, GB01, HCB04a, NP07, RSD01, Sal04, SEdM08]. Extension [ALZ00, Ano00m, AGS01, BDJ^{+}01b, CKC^{+}02, OWR04, Par00, TBSN01, XX05, ALZ03, BH02b, KKN06, LH04, LS08b, vRKS01]. Extensions [Ano02o, BG04a, Gle02, Per02, Rot02, Tre04, Wei04, Ano02j, Ano04b, BDT01, New01, vRKS03, Ang01, JM00, Kre01]. extra [Ano03x]. extracted [WF04]. Extracting [RK02, ST00b, TSL03, Dep03b]. Extraction [BO05, DS04, TSL^{+}02, WL04, WML02, WIC08]. Extreme [NP03, BC03, HL02a]. Eye [Ano05c].

F [Laz07]. Fab [McG04]. Fabric [MD00]. face [Apr05]. Faces [W^{+}04, Ano03-43, Ber04a, GH04, GH07, Cha05b, D^{+}04, Kur04, Man05]. faceted [SPBE09]. FaceTime [Ano02r]. facilitating [Ren02]. Facilities [AGS01]. facility [Rob00b, CVW03]. facto [Egy01]. factor [ZSZ^{+}09, Ano02t]. Factors [BBS04]. factory [Ano05g, Ano01i]. Facts [BALV03, Wil03b]. Fail [She01b]. Fail-Over [She01b]. Failure [RCR06]. Failures [Bar01b, LSW07]. Faithful [Kle05a]. Fall [Lut00]. Fallacies [Wil03b]. families [FL04, QM09b]. family [Ano03-36, DMKN02, Kic04]. Fan [MVM07]. Fan-In [MVM07]. Fantasies [BALV03]. FAQs [AL04c]. Farlye [Ano00b]. fashioned [MFH01]. Fast [Dic01, KMEA04, MZB00, Red01, SGV04, ABL07, CWWS03, Sib00].

Faster [Kie02, TG04, WA04, Rei00b, Rei00c]. FastTrack [FF09]. fatally [Pug00]. Fault [Ano01n, FK03, TMG03, GK08]. Fault-Tolerant [FK03, TMG03]. Favorite [LAB^{+}00]. Fe [ACM00a]. Feasible [KSK04a, PDV01]. FeatherTrait [LS08a, LS08b]. Featherweight [BKMS04, BCV09, IPW01, Stu01, ZPV03, LST02, LS08b]. Feature [MD00, AWE04, CWS04]. Features [BW03a, BW03b, Bro05, Cav02a, HC02, KSK04b, vLGL^{+}02, Lan04, VN00, WC08]. features-including [Lan04]. featuring [And01, Las02]. February [USE00b, USE01a]. Feedback [AHR02, BKO09, ACM03a, KdJNNV09]. Feedback-Directed [AHR02, BKO09, ACM03a]. Feel [Kro00a]. Feeling [Bea05]. Feinberg [Ano00d]. FEM [HKHK03, Nik03]. FEM-Based [HKHK03]. FEM/BEM [Nik03]. Ferris [Fox01b]. Fetch [OKN02b, OKN02c, OKN02a]. Few [Lea00b]. FGPA [Ano02n]. Fibonacci [Bee04b]. Fickle [AAD^{+}01, AAD^{+}07]. FIDJI [GAR04, GRR05, GAR03]. Field [SG03]. fields [UL08, Zen02]. Fighting [HT03, Pau01]. File [Ano02m, KJ02, BDT01, HYX05, ISO05, Sto01b, Sto01a]. files [JK00, Way03]. Filesystems [WBL01]. Fill [Ano04m]. Filter [Ano03h, JMM03]. Filtering [MSF03, OOOiM05, RDW^{+}07]. filters [KM08]. Filthy [HG08]. Final [Dra00, Nat00, RBC^{+}06, UL08]. finalizes [Ano03-36]. Financial [MD00]. Find [PH00b, XAM^{+}09]. Finding [HZC^{+}04, PDV01, TT01, VMMF00]. fine [PH00a, RPB^{+}09]. fine-grained [PH00a, RPB^{+}09]. Fingerprinting [FS03b]. fingerprints [DS04]. Finite [KW02, Cor00, DH00, Gri02b, Gri03, MAJC03, NNS03, WW06]. finite-state [Cor00, DH00]. Finread [Ano03-51]. Fionn [Hec07, Hol06]. fires [Ano05h]. Firewall [ÉJD01]. FireWire [Ano01j]. Firm [BG04a]. First [ACM05, Ano03-38, JT04, Ano03-35, AWS^{+}09, AJ01a, BSB04, BSB08, Bel02, Edm09, FFSB04, Gol04b, Gri08, KR00, LP05, LS08c, MS05, MB05, Mor08b, Rad06, Ras00, Rio02, Rou02a, Sei09, SB03a, SB03b, SB05, SHB^{+}03, Ano01j, Ano02p, HR04b]. first-year [Edm09, Rio02]. Fit [CCM05]. Fits [Uni02, Ano02g, Gro02a]. Fitting [Bus02a, Bus02b]. Five [Lut03c, Lut03c]. Fix [TEM^{+}01, SC08]. Fixed [CBD04]. Fixing [BBDT02, Lut00]. fixpoint [Qia00]. FLAME [GGHvdG01]. Flanagan [Ano00b]. Flapjax [MGB^{+}09]. Flash [Ano02p, ST06, Ano03x, Won03a]. Flash-Based [Ano02p]. flavor [Ano03i]. flawed [Pug00]. flawless [GS00b, Pap00]. Flaws [LAB^{+}00]. fledged [Ano04-32]. flex [Kag09]. flexibility [Gar09, GJ09]. Flexible [ABG^{+}08, BK01b, CMG^{+}01, CEG^{+}03, JMP09, JCKS04, KGMO04, KS01b, MK01, PSDF01, SPB01, SSV05, TTPN08, TOG^{+}05, DLE06, HvE02, HLM06, IV06, LM06, PT09a, TGCF08, ZABL09, vNMW^{+}05]. Flight [BN03, ABI^{+}07]. Flight-Like [BN03]. Flipper [Ano00j]. Floating [CBD04, Dar01b, Fig00, SKC09]. Floating-Point [Dar01b, Fig00, SKC09]. flop [MMG00b]. Florence [IEE03b]. Flow [BCE^{+}01, GS05b, JC04, Liu04, SK00, ABF03, BDLM04, BCHP08, CCKP06, CMJL09, Li02, LZ04, LPH01, MP05, Nau02, RPB^{+}09, SBAD01, WMRT^{+}05, XAM^{+}09, DSBH03]. flow-based [CCKP06]. flow-insensitive [LPH01]. flowcharts [CM05c]. flows [dM04]. fluff [For06]. Fluid [RCB01, RCB03]. Fly [CD01b, DKL^{+}01, Gar00, DKP00, LP01b, LP06]. Flyby [KSC^{+}00]. Flyer [Wil00b]. Focus [Leh01, Leh02, RCdBL02]. focuses [Ano03q]. Folding [EGLZ02, KC00, TCC01, EKEL01, Oi06, TCC02, TCSC02, TCSC04, YCFX09]. fonts [Ano03x]. foolish [Rol08a]. Force [Ano03-39, RBC^{+}05, RBC^{+}06]. Ford [Mar05]. Forecast [Wat02]. foreign [FF08]. Forge [Ler01a, Ler01b, Ler01c, Ler01d]. fork [Rob02]. form [Ano02p, GPF08]. Formal [ALZ02, AOMC07, AW03, BDJ^{+}01a, BDJdS02, Bec01c, BML01, BL03, Cas02, CH02, Che02a, Che03b, CHK^{+}04, DEJ^{+}01, DEL^{+}02, ELM^{+}04, FCMR04, FMR05, LDE^{+}02, MP01b, MP01c, Mos05a, vdPE02, PvdBJ01, Str02, Zam03a, Zam03b, vdBJP01, BTV06, EL01, LYC02, LS06, MORW08, QGC00, BCR03b, GGHvdG01]. Formalisation [Jac01b, Mos05b]. Formalising [AY05, AY07]. Formalism [JV04]. Formalization [TH02]. Formalizations [Ler03]. Formalizing [Ber01c, HM01a, RW03a, SSD^{+}03, ZHC04]. Formally [Sta04b, Ste04, HOP04]. format [ISO05]. Formation [CF02]. Formats [LUH^{+}05]. Formatted [All00d]. formel [BCR03b]. FORMI [KDH^{+}06]. forms [AOMC07, KM07]. formulas [SCWL08]. Forte [Ano01n, Ano02m]. Fortify [Ano05k]. Fortran [BSPF01, BSB^{+}03, FCHE02, LP05, LS04a, SD01b, SD03b]. Fortune [Pra03, Wan03a]. Forum [Ano03-43, Reg02b, DHPW01, GPW03]. Forward [Way05]. Forwarders [AHN02]. found [MMN09]. Foundation [Gut00, Top02a, Ano01i, Way03]. Foundations [BA08, LL01b, Stu01, Die01, LL00, LL03, LL01c]. Four [Ano03k, Ano05d]. Four-way [Ano03k]. Fourth [Ano03-41, Fro07, USE00c]. Fourth-Generation [Ano03-41]. FPGA [Ano02s, Sch04b]. FPGAs [Ano02p]. FPV [CWWS03]. FRACTAL [BCL^{+}06]. Fragment [RMR03, RMR04]. Fragmentation [BCR03a, SC02b]. Fragmented [KDH^{+}06]. Frame [GKMZ04, Gui08]. Framelets [PK00]. FrameMaker [Ano02t]. Framework [ACD^{+}04, AA02a, ALZ02, Ano01o, Bar05, BP01b, BH04a, CM05b, Che03a, DHR^{+}01, EFG^{+}03, Fig00, FP03, GH01, GR07, GHH01, Hun05, Ish01, Kro00a, KS01b, LMV02, LCS04, Mil08, MK01, MF03, NSI03, NCM03, OSM^{+}00, ONRV08, PL05, PQVR^{+}01, RAC^{+}04, RS01, RP03b, SLPO02, SAFG03, SV02, SG03, TMG03, VHL01, WS01a, WH01, Wic03, ABL07, ACZ05, ANMM06, Ano03h, Ano04-29, BDE^{+}03, CV03, CY02, CO04, CR07, Col01, CTLW03, CLZ06, DHS02, DW07, FT00, Gar09, Gri00, HCB04a, HLM06, Hu03, HD03c, Kag09, KKM^{+}06, LO00a, Lau01, Lea05, LJ07, LS06, LRD09, MSU08, MSLL07, NZM03, PV06, PSS01, RB04, SC07, SJ01, SYK^{+}01, SD04, TDB00, Tre02c, Tro04a, Tro04b, Wen05, Yua04, ZS01a, AK01, Bar05, HF00, JHA^{+}05, Spi03b, TA04]. Framework [Tre02b, Tul08]. framework-based [ACZ05]. Frameworks [Ber05b, CC02, DFL00, HHK^{+}01, HHKS03, Ric06a, Jia00, KK00, NP02, PK00, TM08, dM04]. France [AJ01a, AJ01b, IEE03a]. Francisco [USE02, CHL^{+}00, Joh00b]. Frappé [Cou01]. fraud [Ano03j]. Free [AS03, Ano00n, Ano02s, Ano03-37, EXA^{+}05, Sta04a, Ano04q, BR01b, HBM^{+}02, Ano01i]. Freedom [Bar01c]. Freely [GM02]. frees [Ano05i]. French [BCR03b, FTD03]. frequency [SAB^{+}06]. Frequent [Wil00b]. Fresnel [SGV04]. Friedman [Ano00d]. front [Ano03f, Ano03q, Ano04x, Kon03]. front-end [Ano03f, Ano04x]. FrontEnd [Jor02]. Frontiers [ACM06]. Froschzucht [YAW02]. FT [TMG03]. FT-Java [TMG03]. FTfJP [CHK^{+}04]. Full [MP01b, Mor03b, Ste04, ZKR08, Ano04-32, Oiw09]. full-fledged [Ano04-32]. Fully [Fig00, JR05]. Fun [Bee04b, MRB06]. Function [TSL^{+}04, FF08]. Functional [Dd01b, CiLH01, Cou01, GCEO05, Set03, BR01d, Dek06, HD02, VP05, ZKR08]. Functionality [Guh07, Ano03x, Coh04, GB01]. functions [Ano05f, BR06b, NHY^{+}04, SY04]. Fundamental [VZGE07]. Fundamentals [Ano00h, Gil01, HC00, HC03, LO03a, Mad01, WP00a, Dei08]. funkbasierter [Ano05a]. Funny [LAB^{+}00]. Further [Nor00, Gat03]. Fury [McG03b]. fusion [CHMB04, Man01]. Future [CM04, Fri02, Leh02, Pau01, AWS^{+}09]. Futures [PSH04, WJH05, ZK09]. fuzzing [GKL08]. Fuzzy [Dor02, SPBE09].

G [Ano00d]. G&D [Ano01p]. G.lite [Ano00i]. gadgets [Ano03i]. Gains [Ano02c]. Game [Bur07, DHR^{+}01, GS08, RM08, Ros02b, Dav05, DW07, Gui08, LM06, Sei09, Swe06, WWJ07, BGNM04, Sco03]. game-frame [Gui08]. Games [BBV03, LH02, RM08, Bre02, Fro08, Ges07, LRD09, SdSK05, Sel03]. gap [Ano04r]. Garage [Pra03]. Garbage [Ano04l, Ano04s, BCR03a, DKL^{+}01, MJ06, PUF^{+}04, SGF^{+}02, SLC03b, SHB^{+}03, XSaJ08b, ZS01b, ZT02, BAL^{+}01, Bac07, BBYG^{+}05, BCM04, BALP01, BALP06, CSK^{+}02, DKP00, GSaC05, HBM^{+}02, JMP09, LP01b, LP06, MSLL07, PHV07, SMTZ09]. Garden [MSK09]. Gas [PDCL02]. Gate [Way03]. Gateway [Ano02r, Yua04]. Gateways [RAC^{+}04, CG02]. gathering [Fel04, HNZS03]. Gaussian [Ano00h]. GC [HM01b, Oga09, SKS01b]. GCC [BHP^{+}01]. GCJ [Bot03, Sal06]. Gear [Ano00h]. Geeks [Ive03b]. Gem [Och09c, Och09d, Och09b, Och09a]. GemIdent [HKL09]. Gemplus [Ano02d, CH02]. Gems [Deu00, Pet06]. Gene [Wil00d, DJ01, GV05, GP05, SD04, CSFS00]. General [WP00b, BDE^{+}03, MSLL07]. General-Purpose [WP00b]. Generalization [SLPO02, UL08]. Generalized [KKG09, HNZS03, KdJNNV09]. generalized-LR [KdJNNV09]. Generate [Sea02, Ano03h]. generated [BRU04a, CMS06, KdJNNV09, Ren02, WGSD07].

Generating [HHK^{+}01, HHKS03, HBM^{+}06, Jen02a, KNY03, Nik03, MCLDP01]. Generation [Ano01l, Ano03-41, BM04, BL03, CF00, CQX^{+}09, Ebe02, EFN^{+}01, GM05c, HKS02, KK04b, MdB01, PV04, SMCS04, SSS05, TRVH03, VPK04, Ano02a, Ano04-28, BI02, BCHP08, Car06, EFN^{+}02, HZS08, ACM03a, JA01, Pay04, Yam04]. Generational [MJ06, DKP00, WK08a, WK08b, WK08c]. Generative [CM05b, Sch04d, GST05]. Generator [Ano02q, Bri02, LRSW00, PSW07, vMV05, EGKP02, For04a, vdSPP05]. generators [Cle01a, Cle01b]. Generic [ABH^{+}00, DKTE04, GK03, PNCB06, SM04a, Wad00, BGNM04, CO04, CR07, SH03, Tor01, AC06, Tre02b]. Genericity [AR08]. Generics [Bat04, Gho04, MPO08, NW06, NW07, vD04, IV06, RFZ08]. Genomic [NDS^{+}02]. gentle [TV08]. gentler [Fry03]. gently [BB00a]. geographic [HL02b]. geography [LYL^{+}04]. geolocation [MV09]. Geometry [Bar00a, KM04c]. Geoscience [IEE03a]. Geospatial [HJF06]. German [Ano03s, Ano03-33, Ano04c, Ano04h, Ano04l, Ano04v, Ano05a, BL04, HMD04, Lex02, Sig04, Wol03b, Zus03]. get [Ano03-32, HBM^{+}02, Hoh03, IN09]. Gets [Ano03r]. getter [Hug02]. Getting [Ell06, LAHC06]. Gigabit [Ano03-36]. gInstall [Ano03-38]. GIS [XP04]. give [Har01b]. gives [Ano04-29]. GJ [IPW01, Wad00]. Glassfish [Hef07]. Glenn [Fox01b]. Global [Ano00i, Uni01, EL04, FWL03, MBS^{+}08, NIKN06]. Globus [SC05]. Gluecode [Ano04m]. GmbH [Ano00h]. GNAT [Och09b, Shi03a]. GNAT-AJIS [Och09b]. GNOME [Pet05]. Go [Bar03a, XAM^{+}09, HAL02c]. Goes [Bar03a, Kic04, Pau01, Ano04g]. Going [SCL^{+}08]. GoJava [Wis06]. Goldilocks [EQT07]. Good [Pre03, Zen02, Cro08, MLM^{+}08]. Goodrich [Mas01]. Google [Fit07]. Gopher [Mam01]. Gosling [Hol04b]. Government [LS03, LAB^{+}00]. GPIB [Tim03]. GPS [Hon05]. grade [Fro07]. grading [Hel07b, Mor02]. Grained [DFA03, PH00a, RPB^{+}09]. Grammar [GKL08, CY02]. Grammar-based [GKL08]. Grammars [SB00]. Grande [ACM01b, DHPW01, Fox00a, Fox00b, Fox00c, Fox01a, Fox05, GPW03, Ano01c, Fox03a, Fox03b, GPW05, SBO01, WG01]. Grande-ISCOPE [Fox05]. Grande/ISCOPE [ACM01b]. grandmother [Hol04b]. Grant [TCM^{+}00]. Granting [TCM^{+}00, HG07b]. Graph [Ano00j, BH02a, CCW02, CDFR04, Dmi04, JC04, CMS05, CCT01, Wu05, ZR07, ZABL09]. Graphic [Gea00]. Graphical [Ano03l, ACR01, LM06, MCLC02, Sco03, AWE04, BE02, CWS04, DSCU01, HG08, LP05, Las02]. Graphically [Uni02, Ano02g]. Graphics [Ano02q, Ano03-41, Ano08, BI07, CN03a, MCLDP01, Par04c, Par04b, Pra08, Sch00a, BDRV01, BBGP01, Gou06, Har00b, MRB06, MJ00, PC08, SML06, Ano02m]. Graphing [Ano01m]. Graphs [BH02a, Wal02b, ABG^{+}08]. Gravity [BL04]. Gray [Che05]. grayscale [Woo03]. Greasemonkey [Pil05]. Great [BR02, SLB^{+}02, Ano01i]. Greece [SM07, SBH^{+}04]. Greek [Lik04a, Lik04b]. Green [Ano01j, Ano01k, SKP^{+}02]. Gregory [Che05]. Grehan [Fox01b]. Grid [vLSM01, vLGL^{+}02, AG05, HdS^{+}05, YdOLS^{+}05, vLFGL01, ABG02, AG03a, AG03b, BBC07, Bal03a, CLL03, GvLPF01, Hua03, HBD04, JF05, LTOT07, LCFL04, Tui04, Wal03a, WXW^{+}05, YAA07, ZCQS04, vNMW^{+}05, vNMKB05]. Grid-Based [vLSM01]. Grid-enabled [LCFL04]. Grids [VDPC01, VDPC03, GR07]. Grind [Lut00]. Gripper [ZG04]. gritty [Way03]. Groovy [AK09]. Grossenmasse [Wol03b]. Group [Ano00h, Ano00j, BCMT03, BW03c, DL02, SBH^{+}04, KK00, Oes01, Ano01o, Dob01a]. Groups [BBC07, CF02]. groupware [KK00, Ano04n]. Groupwork [Bow07]. grow [Eng00]. Growing [BK03]. Grows [Ano05f]. growth [BALP01, BALP06]. Gsm [Cim02]. Guarantee [Hag02]. Guaranteeing [BD03b, Fre05]. Guarantees [PSM01a, MSG01, PSM03]. Guava [BST00]. GUI [Kon03, Ano04a, BH04c, BK03, Bri02, Che02a, Che03b, Eng04, Hei03a, KW01a, TETPQ08]. GUI-like [KW01a]. guidance [HSB09]. Guide [AM02, Azi06, Blo01, BGG^{+}03, Bru03, CR02a, Cal03, CDH07, HS00a, HL03c, LG99, LG00a, Lut03a, Mak03, ME00a, MC04, Nas04, NRV00, Pau03, Red01, Spi03a, Spi03b, TB02, Wei04, Ana01, Bec04, BS00a, BD03c, BD07, Bro01, Bur05, Cal00a, CD01a, Che00, EFO08, Est02, Fla02c, Fla06, Gar09, Gig00, Hag00b, Har03, Hol05, Jor02, LL08b, MD06, MCG03a, Mer04, MR00b, New00, PM01a, Pol01, Sik03, Spe02, Tay02, Tha00, Tha06]. Guidelines [KR01b, Lut00, Rou02a]. Guiding [Ros02b]. guild [Gui08]. GUIs [Les03, MA05, PRR02, Röß06]. Gumbie [Bri02]. gut [SKS08]. Guys [Pra03]. GVis [ZCQS04].

h [MAWW^{+}01]. Hacking [Cha03]. Hacks [AE06, MA05, EA06, Per06, Pil05]. Half [Lut02]. Hall [Hal01a]. Halstead [Wol03b, Wol03b]. Halstead-Lange [Wol03b]. Halstead-Metrik [Wol03b]. Hand [WBL01]. Handbook [LRO02, JPC00]. Handheld [CD03, Pau01]. Handheld-to-Handheld [Pau01]. Handhelds [Ano02o]. Handle [Cox01a]. Handling [BM03, Che02a, Che03b, SM04a, Wol01a, BHJR05, BS00b, JPB^{+}08, KBP^{+}03, LYM04, Och09d, OKN01, Pal02, SMTZ09, Ste05, SC01b, ZK09]. Hands [BBHL01, Ana01]. Hands-On [BBHL01, Ana01]. handset [Ano03n]. handy [Mer04, Suo04]. HANDY-STANDARD [Suo04]. Hans [Pap05]. happen [Gen00]. Harassment [TCM^{+}00]. Hard [Eng00, Fre08, NK03, TGB^{+}04, SAB^{+}06]. Hardcore [Gol00, Sim04a, Sim04b]. Hardgrave [Gla06]. Hardware [Ano01m, Ano03-38, HT06, HIBP04, Hsu01, KKN00, LMK06, MD00, NRS^{+}07, SLC03b, WHW01, BHDS09, BGED04, GGL^{+}08, IN09, JMS02, JMP09, KKM^{+}06, Oi05, Oi06, Oi08, SPG07, TCSC04]. hardware-assist [KKM^{+}06]. Hardware-in-the-Loop [Ano03-38]. hardware-translation [Oi06, Oi08]. Hardy [Pap05]. Harkey [Bar03a]. Harman [Mar01b]. Harmful [Ams02, SD08, GEVZ09a, Our02]. harmless [ACFG01]. Harness [KS01b, MSS00]. Harnessing [EFO08, SQG^{+}05]. Hartstone [Wan02a]. Harvey [Ano00d]. Hashing [SSS05, CHL07, Duc08]. Haskell [Fre07, PT09b, XJC09]. hasn’t [Moo03b]. Hatcher [Mor03b]. HAVi [Lea02]. HBE [Ano00k]. HBench [ZS01b, ZS01a]. HDM [KY03a]. HDT [KKJY04]. Head [BSB04, BSB08, FFSB04, MD00, McL06a, Mor08b, SB03a, SB03b, SB05, Ano03w, Ano04g, Rob04a]. headaches [Ano03o, Apr05, Wan02b]. header [VED07]. Headless [Yua04]. healing [GK05]. Health [HE03, Ano03j, LSK^{+}02]. health-care [Ano03j]. Heap [CKV^{+}03, SKS01a, SKS03, BALP01, BALP06, CH08, KF00, LLS^{+}08, ST06]. Heaps [DGK^{+}03]. heart [Mer04]. Heat [GKM03, ZK04b]. Heavy [Ano00h]. heel [XSaJ08b]. Held [HR04b, MFRW07, SBH^{+}04]. HELIOS [Ano00h]. Helix [Ano03-37]. Help [Kro00b, Ano04q, HPH03, Men03]. helpful [VVV04]. helps [Ano03-30, Way03]. HERCULE [Ren00]. Here [Mer04]. Heterogeneity [Zhu03]. Heterogeneous [AJMJS02, BCS02, CCC^{+}04, KM02, RLR00, SMS00, SRJS08, CCK^{+}08, GCARPC^{+}01, SGW01, ZYZ06, ZLG08]. Heuristic [Coo05, GV02a]. Heuristics [GV04, Sch03a, GV02b, LMK08]. Hibernate [BK05a, Ell06, EFO08, WACBL03]. Hickory [Ano02i]. HIDOORS [MLJH04]. Hierarchical [PHV07, WDSD02]. Hierarchically [LFP04]. hierarchies [AK09, PZ00, ST00a]. hierarchy [Ano02k, KF00]. High [ACM00c, ACM01c, ACM04, BC00, BBHL01, BDT01, BW01a, BA01, CW03a, CT00, CEG^{+}03, Fig00, GP03, GGH^{+}03, GMM00, HWB04, HCB04b, IJ03, KMOS03, KWK03, Lau03, LMG01, LRSW00, Lut03a, MLG^{+}02b, PBG^{+}01, PS03, RCB01, RCB03, RB01, SD01a, Vil08, Vog03, WGW04, Woo05, Ano03f, Ano04b, AGG02, Bar02a, BFGS05, BSW^{+}00, CMS03b, Chr05, Dob01b, Gam00, G^{+}01, GBCW00, HF06, KCSL00, KHBB01, KWK05, Lau01, LCFL04, LMG00, LAL02, MI01, MMG^{+}00a, MMG^{+}02, PC08, SAB^{+}06, SPGV07, WW09, PL01a]. High-dimensional [BW01a]. High-Dimensionality [Vil08]. high-frequency [SAB^{+}06]. High-Integrity [HWB04, Dob01b]. High-Level [Fig00, RB01, BFGS05, CMS03b]. High-Performance [BBHL01, BA01, CEG^{+}03, GP03, GGH^{+}03, KMOS03, Lau03, LMG01, PS03, RCB01, SD01a, WGW04, Woo05, BDT01, RCB03, AGG02, Bar02a, HF06, KHBB01, LCFL04, LMG00, LAL02, MI01, MMG^{+}00a, PL01a]. high-performing [GBCW00]. High-Tech [Lut03a]. high-throughput [SPGV07]. Higher [BO05, BO08, MPO08, Nik03]. higher-order [Nik03]. highlighting [SPBE09]. highly [TGCF08]. Hills [Ano01j, Ano01k]. hindered [Ano03w]. HIPPI [Ano00k]. Historians [Fel04]. historical [MWM01]. history [Nis03]. hjelp [HJL00]. HLA [McG04]. Hoare [GSWZ08, HJ00, vON02a, RWH01, vO01, vON02b]. Hobby [LAB^{+}00]. Hoboken [Ano04e]. hoc [SM01a]. Hogging [Bar01a]. HOL [RW03a, Sch04a, ZHC04, vO01]. Hold [GM05c]. Holm [Fox01b]. Home [AA04, Ano00m, Ano05j, Lea02, LSK^{+}02]. Homepage [Dar01a]. Homework [GM02]. Homework/ [GM02]. Hong [Uni01]. hook [Kic04]. hope [CAF04]. Hopes [Bar01b]. hospitals [Bar09]. hostile [HWM01]. Hosting [PKF02]. HostML [Ano00j]. Hot [Ano04o, Ano04p, S.04a, S.04b, CS06, LAHC06, LMK08]. HotSpot [GM00]. Hotspots [WG01]. HotSpot^{T M} [KWM^{+}08, PVC01, RB01]. Hotswapping [Dmi04]. Houdini [FL01]. hours [AK00, WMM04]. HP [CFLL03a, CFLL03b, LCFL04]. HPC [Ano03-38, BCS07, SCB09]. HPC.NET [Vog03]. HPJava [CF03, LCFkL05]. HPM [BGH^{+}07]. HPM-sampling [BGH^{+}07]. HTML [AL04b, AF02, Goo02a, GT00, II04b, Knu01a, MDS04, RDW^{+}07, TB00b, ZJ03]. HTTP [Ano03k, SRJS08]. Huffman [Wic03]. Huge [BHP^{+}01]. Human [LH03a]. Human-in-the-Loop [LH03a]. Humidity [Lia03b]. Humming [Pau03]. Hunt [Azi06]. Hunting [Lut03c]. Hybrid [LMK06, XAN07, RB04]. HYDRA [War02]. hyogen [SM04b]. Hyperformix [Ano01n]. Hyperion [A^{+}01].

I/O [All00b, Ano03k, BDT01, Gri00, Har06, VT01, WC00a, WC00b]. IA [Ano00h, IKN03, SOK^{+}04]. IA-32 [SOK^{+}04]. IA-64 [IKN03]. IAPPGA [Wu05]. Iava [Ric00]. Ibis [Bal03a, vNMW^{+}05]. IBM [Ano00h, Ano04i, GEAS00, SKC09, SOT^{+}00, Yus04]. ICANN [Bar01c]. ICCMSE [SM07]. ICE [BC04]. ICE/TTM [BC04]. ICETM [BC04]. Iconic [CM05c]. ICT [Ano03m]. ID [Ano03-28, Ano04t, GM05c]. IDE [Ano02p, Ano01i, Ano01l, Ano01n, Ano02n, Ano02q, Ano03-37, Ano04-29, Bur05, CH06, Fre07, Gee05, HCB04a, MKF06, PH03, PHBM05, RC04, Sur04a, VN03, Vau03b, WKB02]. idea [Ano04i, ABL07]. ideas [BR02, Eub05, WKB02, BHP^{+}01]. Identification [SPR^{+}03, WG01, DS04]. Identifier [vdBJP01, CDF05]. Identifying [HMRM03, LSW08, MVM07, PHM^{+}01, RCR06, HKI08]. identity [Ano05f]. IDEs [Ano05d, Gat03, MKS^{+}03, OPS^{+}02]. Idiom [LG99, LG00a, KKM^{+}06]. idioms [PZ00]. IEC [ISO08, TSL^{+}04]. IEEE [ACM04, IEE02b, Fig00]. IEEE/ACM [ACM04]. If [Mer04, ZK09]. IFIP [Jac04b]. IGARSS [IEE03a]. Igniting [ACM03b]. Ignition [CVW03]. ihre [Ano04l]. II [Ano00h, Fox01b, Ang00b, Dei08, HC02, PDCL02]. III [Ano00j, Ano00m]. iJADE [LL01a, LL01a]. ILE [HKF00]. Ilea [TM07]. Illegal [BCE^{+}01, HT06]. Illinois [ACM05]. Illuminating [BLPV04]. illustrate [AYWM08]. Illustrated [SDPM04]. Illustrating [Hol04a]. Illustration [GKW04]. ILP [RTJ00]. ILS [Ano03a]. im [BL04, Ano02r]. Image [Bur03, BG02, CE01, HKL09, Lau03, MWL00, RLR00, SU03, SAFG03, YWZ03, Ano03-36, Bos04, Eff00, Hun03b, KGH^{+}05, MM04, MF03, RSD01, Sam04, WN05, XAN07, dCG^{+}02]. image-based [Sam04, XAN07]. Image-Processing [SU03]. ImageJ [MM04]. images [Woo03]. imaging [HBX^{+}04, Rod01, dGNv04, Bur02]. Immersive [Lut03a]. immutability [TE05]. Impact [BNV08, RST^{+}04, RCR06, Rob01c, SKS03, BCM04, CD08, LPH06]. imperative [Ras00, ZKR09]. Implement [CZ02, Coh02, Gso00, Zhu03]. Implementation [ASS03, AAA^{+}04, BFG02, BKH02, BR01a, BO09, BNO03, BKY^{+}03, CWHB03, CS02, CHK00, DHRH05, DLS^{+}01, Gle02, GLS02, HK02b, JR02, JJ02b, KT04, KPKL03, KM04a, KMOS03, LPSY04, Mam01, MLVB05, MSS00, NK03, Oiw09, Omo03, PL05, RS01, SG02, SNOM01, Sur01, TGB^{+}04, USE00c, VHBB01, WXW^{+}05, Zea00a, ZYC03, ACFG01, Ano04l, AP02, AFT01a, ANH00, Bes01, BV05, BC04, CHMB04, CMLC06, Die01, DCA04, FDR04, FLWW04, Gab07, HdS^{+}05, IKY^{+}00b, JH03, KBVP07, Kon04, Lan00, LH08a, Li04, LY03, LC04, OG05, Oes01, Sig04, SH04b, VVG^{+}05, VHBB03, Vir03, WLW^{+}03, WM00b, YdOLS^{+}05, ZP03, ZFK04]. Implementations [HdJ01, Hir00, SS00a, CZ01, DMP09, JS01, LLdA08, SZ00, WCC04, WF00, WF02]. Implemented [Sch04d, YKS^{+}02, PSW07, Tor01]. Implementierung [Ano04l]. Implementing [ABH^{+}00, AFT01b, BP05, CLCC02, Dic01, DKL^{+}01, GGH^{+}03, GEK01, Hin02, HOP04, IJ03, LDM04, MBMZ01, NS01b, NIEH04, OHL^{+}05, Pot04, RSH01, Rou02b, SP03, WP04, WKB02, AGST04a, AGST04b, ANMM06, BHK^{+}04, HW00, HLM06, Lut03b]. implications [AR08, RVJ^{+}01]. Implicit [BWLR06, BH05c, WM00a]. Implicit-signal [BH05c]. Implicitly [AHKR01]. import [All00a, All00b, All00c, All00d, All00e, All00f, Lan04]. importance [BC07]. Imported [Mac05]. Improve [LBJ02, Pau03, RT02, Ano02l, Bar01d, D^{+}00, HCMM00, KF00, LBJ05]. improved [Wel06]. Improvements [GCB^{+}00, Vau03a]. Improving [AAAG^{+}05, BJK07, Cog03, CCB^{+}01, JMK^{+}08a, JMK^{+}08b, JMK^{+}08c, MS00a, Pau01, OOK^{+}06]. IMS [Ano03-42]. In-lining [SYN02]. inalambricos [Ano04-33]. inAspect [ASS^{+}05]. Inc. [Ano00i, Wan03a]. InCert [Ano01n]. incinerator [Lex02]. include [Ano03z]. includes [Gar09, SML06, SM01d]. Including [CK05, Des01, HL02a, Lan04]. Inclusive [DW07]. Incorporating [Kod04, LJ08, Tre03]. Increase [GKM03]. increases [Ano04-31]. Increasing [JS01, WCK^{+}07]. incremental [BBYG^{+}05, KP06]. incrementalisation [WPN08]. incrementalization [SB07]. independence [ADR09]. Independent [DHPW01, DS09, FSS06, LN04, SBB05, TS01, Ano03l, Ano03-50, GPW03, PG03b, PG03a]. InDesign [Kah06a, Kah06b]. indirect [JMK^{+}08a, JMK^{+}08b, JMK^{+}08c]. indirection [LGFM05]. individual [LW03]. Indoor [dFR04]. Inductive [AddS03a, Moo06]. Indus [JRH05, RH07]. Industrial [AA02a, HMD04]. Industrieautomation [HMD04]. Industry [Ano03n, Bar01a, DFL00, Ano02w, Reg02b, UCJ^{+}04]. inefficiencies [KOO08]. Inference [AS03, CHS01, Ebe02, WS01b, BAdMS08, BP03a, FFLQ08, GF07, SC08, UL08, dMSAV08]. Inferred [MCD09]. Inferring [MF07a, TT08]. informaticas [Ano04-33]. Informatics [Guh07]. Information [Ano02r, DTD04, Gal01, GS05b, Hac01, ISO08, Kro00a, LN04, RTVH01, SPS^{+}02, SKS03, TA04, Ano03-29, AT01, ABF03, BDLM04, CO04, CMJL09, Dep03b, Ham07, HNZS03, Li02, MP05, RPB^{+}09, WMRT^{+}05]. information-flow [Li02]. Informix [DHMT00, Ano00n, Har00d]. Infotainment [Bat03]. Infragistics [Ano03-41]. Infrastructure [Bar05, BA01, DA02, Tui04, VHL01, BG03, Bro09, Joh00b, LM06]. inheritance [Ano02k, BLV03, DMP09, Lyo02, Mor02, PB08, TB00a, WSP02]. INIDP04 [LDM04]. initial [Jen01, Utt06]. Initialization [Ber01c, BS00c, KS02a, QM09a]. initiative [PB06]. Injecting [CFL05a]. injection [GK08, SW06]. Inlet [PDCL02]. Inline [GH03]. Inline-Threaded [GH03]. inlining [LH05]. Inner [All00e]. Innovation [ACM03b, Lut03b, McG03b]. Inprise [Ano00m]. Inprise/Borland [Ano00m].

Input [MD00, SRJS08, VPK04, PT01]. inputs [SMTZ09]. ins [Ano05o, DHMT00, FS03a]. Insecurity [Lai08]. insensitive [LPH01]. Insertion [Zdr09]. Insight [IEE02a]. Insightful [SPS^{+}02]. Inspection [SG03, Cha06]. inspired [TDB00]. Installation [Ano03-40, DHMT00]. Installations [Kro00a]. Installer [Ano01h]. Installing [EXA^{+}05]. InstallShield [Ano00h, Ano01h, Ano02p, Ano03-40]. Instant [Tre00, Tre01]. instantiation [AC06, Ano01l]. Instantiations [Ano02o]. Instruction [AHKR01, KC00, LFH03, Oi06, Sch04c, XX05, Ano02j, AWS^{+}09, Emu04, Sco02, YCFX09]. Instructional [NLFA02]. Instructions [HPS02, Ano03-31, KKM^{+}06]. instrument [Bus02b]. Instrumentation [GNYZ05, BP01c, BWW^{+}03, CO04, YCIS07]. Instruments [HL03b]. insurance [Ano01p]. Integer [BK08, Win02, YTY00]. integer-reference [YTY00]. Integral [Jac03, Kun02, RW03a]. Integrate [Zhu03]. Integrated [Ano00h, Ano01k, Ano02p, CDH07, GPF05, Hel07a, IKN03, LKL^{+}03, Sta01, ACC^{+}01, JCP^{+}05, NM02, Rio02, ZKR09, Ano01j, Ano02t]. Integrates [Ano04-37, Ano04o]. Integrating [AL04b, HL04, KDH^{+}06, MORW08, NE04, PT09a, SJG03, TA04, WSVX03, YE04, BHW05, LHFL07]. Integration [AGH05a, Ano01k, Ano02r, Cha05a, DF03, GF01, Kun02, LFM09, MF01b, SM01b, SM03a, Zhu04, ACZ05, Ano02l, Ano04-27, DOR05, FLMS06, HNZS03, RB04, dCG^{+}02]. Integration-Ready [Cha05a, Zhu04]. Integrity [Ano02s, CW03a, HWB04, KWK03, Dob01b, KWK05]. Intel [BHP^{+}01, CMP^{+}07]. Intelligence [Lut01, Lut03c, WL04, Lut03a]. Intelligent [Ano02n, Ano02p, LL01a, Lut03b, MLG02a, SV02, Ano05k, BB01, Kim02]. IntelliJ [Ano03-37]. intensive [SFMH01]. intent [AAAG^{+}05]. inter [TM07]. inter-language [TM07]. interact [EGD03]. Interaction [AHKR01, Hei03b, JV04, WP04, Ano01d, LYC02, Rob02]. INteractive [ESGS00, BW01a, BLN06, DK02, GLS02, Hit03, HKL09, Kro00b, LS04b, NLFA02, Soj03b, Tra00a, Uni02, Vor01, ZGB03, ZCQS04, ABL07, Ano02g, BD04, BG04b, CHB03, Est01, GJ04, Gol04a, JFH00, Knu01a, LW03, LHS04b, LRD09, MAJC03, MSK09, Rob06, Sei09, SM03b, Tha00, Tha06, Ano00n, Ano02m]. interactivity [KW01a]. interactomes [CMS05]. interaktive [Ste08a]. Interception [CW04b]. Interceptors [NMMS01]. Interdisciplinary [Fel04]. Interdomain [Lut02]. interests [Djo08]. Interface [ACGL01, AČMN05, Ano02o, BFM^{+}02b, CGRR04, Hel07b, KSC^{+}00, KM01, MCLC02, OS02, Ros00, SH04a, Sco03, TDB00, VUPB02, Wil00a, YHGL01, Zea00b, AJMJS05, Ano02a, Ano02k, Ano03l, Bak00, BRU04a, BK00, CFKL00, CvE00, CMS05, CHS^{+}05, DSCU01, Gam00, HTSW07, KOB01, Kon04, LBR06, PFJ05, PT01, PFS05, AMJS05, HG07b, MCLDP01, PZ00, VL00]. Interface-based [Hel07b, Bak00]. Interfaces [Alb03, All00e, Bar00c, BKLS00, Gut00, NK03, Sch03b, TT01, ACFG01, Kon03, WML02, BKLS01, LS08a]. Interfacing [LAT04, ASS^{+}05, Och09a]. Interference [RH04, KM08, Kle05a]. intermediate [Ano03k, vTNC08]. intermediate/proxy [Ano03k]. Internal [Ano00i, SC02b]. internals [Sci07]. International [ACM00a, ACM00b, ACM01d, ACM05, Ano00i, Ano00k, Ano02i, AJ01b, CNB00, GAR04, GRR05, HR04b, IEE02b, IEE03a, Jac04b, SM07, SY^{+}05, SBH^{+}04, Tra00b, Uni01, AJ01a, GAR03, ACM03a, YLM^{+}05, Ano01o]. Internationalization [Ish01, Jac01a, DC01, Röß06]. Internet [Ano00i, BL04, LS03, Ano03-37, Bar01a, Bar01c, BL04, BKY^{+}03, Chr00, CSK00, CCB09, CE01, CK05, EM03, Hol04a, HL02b, JF06, Knu01a, Kro00a, KPN02, LL01a, MV09, NPRC01, Gal02, Ric01, RJFG03, Sat04, SEGS03, TS01, Wea07, Wil00a]. Internet-challenged [Kro00a]. Internet/client [Wea07]. Internet/client-side [Wea07]. InternetBeans [For04b]. InterNetwork [Ano01o]. interop [Ano03o]. Interoperability [DHR^{+}01, FJ05b, TEM^{+}01, Ano03o, Ano04w, FLMS06, Men03]. Interplanetary [Wat02]. Interposition [XLG03]. interpret [HPH03]. Interpretation [BDT04, BD02, GH03, MD00, PL05, SSV05, BDL^{+}08]. Interpreter [GEK01, OKN02b, OKN02c, SMK02, OKN02a, PT09a, Ric00]. Interpreters [CGEN03, EGKP02, WB00]. Interpreting [Han05b]. Interprocedural [NR06, WIC08]. InterProlog [Cal04]. Interruptible [LKM06]. Interruptlets [CCB^{+}01]. Interscience [Ano04e]. intersection [NQM06]. Interval [LL01d]. Intervals [BF03]. Intervoice [Ano03-35]. IntraLinux [Ano00i]. Intranet [Ano03-37]. Intrinsic [KFLN04]. Introduce [RP03a, LS08c]. Introduces [Ano01k, Ano01m, Ano01o, Ano02m, Ano02q, Ano03-39, Gil01]. Introducing [Ano02e, Hac01, Soo09, CC02, DMKN02, GM08, Gri00, NR05, SD03a, Sto01b, Sto01a, ZJ03]. Introduction [ANN01, AW00, Bar00b, Bis03, BA07b, CO07, DWH01, Goo03b, Knu01a, Lia00a, Lia00b, Lia01, Lia02, Lia03a, Sav01, Zen02, Bes01, Bro09, Coo01, Eff00, Gar01, Gol04b, GT00, Hun02, KMR02, MR06, NH02, Och09a, Rad06, Ril02, Ril03, RVZ04, TV08, WB01, Wu01, Lex02]. Introductory [DK02, ES05a, HMRM03, MDS04, Rob04b, Bar02b, BVPE06, CFGL05, ES05b, ET02, Gel00, LDB^{+}03, SCS01]. Introspection [BO05, WWMG06]. intrusion [HWM01]. Intuitive [Ano01h]. iNUX [Ano00i].

Invariant [PV04, SB07]. invariants

[FX07, NE04]. invasively [Ren00]. inventor [CY01b, Hol04b]. inverse [GEG07]. inverses [GE08]. Inverted [KK03a, SDPM04]. Invest [Wan03a]. Investigating [GSW00, JKKL04, Lut01, MFRW07]. investigation [BP01c, CLN07, HTSW07, PJ05]. investment [Ano02w]. Invitation [SG00]. Invited [LD03]. Invocation [JO03, MK01, Tddd03, PM01a, AV05, NMMS01]. invocations [IH01]. Invokeinterface [ACFG01]. Involving [CK05]. IO [PR04]. Iomegas [Ano02m]. IONA [Ano01m]. Iopsis [Ano01n]. IP [CD01a, Cal03, CF00, KSC^{+}00, Lut03b]. iPES [DK02]. IPP [Est01]. iPro [Ano02f]. IPv6 [Ano01j]. IQ2 [Ano00i]. IRI [MAWW^{+}01]. IRI-h [MAWW^{+}01]. Iris [KK00]. IronGrid [Ano03-36, Ano03-41]. irreconcilable [Tan07]. Irrelevant [Spi05]. Isabelle [RW03a, Sch04a, Str02, vO01]. Isabelle/HOL [RW03a, Sch04a, vO01]. ISAPI [YWZ03]. ISBN [Azi06, Bal03c, Cha05a, Dud06, Kuc06, Mil08, Pet06]. Ischia [ACM06]. ISCOPE [ACM01b, Fox05]. Islands [INM05]. Isn’t [Ron01, Ano05n, Yua04]. ISO/IEC [ISO08]. isolated [BKO09]. Isolation [ACL03, BHL00, DMP05, Cza00, SMAT^{+}07]. ISSAC [Tra00b]. Issue [Bak00, Dek00, Fox00a, Fox00b, Fox00c, Fox01a, Fox05, HR04b, Ano01p, EL01]. Issues [AJMJS02, CK05, Liu03, McG04, MSSJ00, NK03, Bro07, GEAS00, Mor03c]. ISVs [Apr05]. Italy [IEE03b, ACM06]. Iterable [LM02]. iteration [Qia00]. iterators [LKM06]. ITEST [PB06]. iTunes [Rog03]. IUC18 [Uni01]. Iverson [Ano08]. ivory [Reg02b]. IVR [Ano00k]. iXj [BG04b].

J [Gil00a, Goo03b, Lia00b, SASZ03, APA04, BDN05, DV01, DJ01, LS03, SMCS04, TS02, TS09]. J# [GS05a]. J& [NQM06]. J-CAT [LS03]. J-DSP [SASZ03]. J-Express [DJ01]. J-Orchestra [TS09, TS02]. J.A.D.E. [Dau01]. j.MD [VWS^{+}05]. J2EE [Azi06, Cha03, AU02, ACM01e, Ano03-36, Ano03-40, Bar02a, BG03, CR02a, CI01, CK03b, DF03, Fry03, HK02a, Hap02, Hub02, HL03c, Jol01, JCKS04, JDJ^{+}06, Jor02, Lai03, MS01, Mer04, NC04a, OBr05, PPJ03, PNKN04, WMC04, Wal03b]. J2ME [Vir05, Yan03, Ano02m, Ano03m, IK04, KM04c, Muc02, Pir02, RTVH01, Top02b, UCJ^{+}04, Utt06, Yua03, Wri03]. J2SE [Utt06]. J3DV [FMA02]. Jabiru [SQG^{+}05]. JAC [HL06, KT01a, PSDF01]. Jackie [Ano08]. JADE [SV02, DK03]. JAFARDD [EGLZ02]. Jaguar [WC00b]. JAI [Rod01, Bur02]. Jakarta [BDHdS01, Cav02b, CK03a, Cav04, Ler01d, O’B05, Sig05]. Jakarta-Tomcat [Ler01d]. Jalapeño [AAB^{+}00, AFG^{+}00, NS01b]. Jalview [CCSB04]. Jam [ALZ00, ALZ03]. JamaicaVM [Ano04l]. JaMake [BK01a]. James [Hol04b]. JaMP [KBVP07]. Janet [BKLS00, BKLS01, BKŁ01]. JANIS [Ano03-29]. January [USE01a]. Janus [Ada06]. Japanese [Ano00i]. Japlo [Esp06]. JaRec [Chr01, GCRD04]. Jaroslav [Mil08]. Jarrix [Ano00j]. JaRTS [Gle02]. JAS [KS01a]. JASMINE [ESGS00, SEGS03]. Jasp [NHY^{+}04]. Jass [BFMW04]. JastAdd [EH07]. JATOON [dS02]. JaTS [SGV04]. JAVA [Lex02, ACM01b, Ahm01, Ano00a, Ano00h, Ano00k, Ano01b, Ano01g, Ano01n, Ano02b, Ano02h, Ano02k, Ano03c, Ano03s, Ano03-27, Ano03-37, Ano03-33, Ano04c, Ano04h, Ano04l, Ano04-36, Ano04-35, Ano05a, Ano08, Azi06, BIB05, Bal03c, Bar03a, Bee00, Cal00a, Cha00a, Cha05a, Cha03, Che02b, CY01b, DHMT00, Dob01a, DFL00, Dud06, Fox00a, Fox00b, Fox00c, Fox01a, Fox05, Fox01b, Fox01d, GP01, GS00b, GDB02, GAR04, GRR05, Hec07, HRD08a, Hep04, Hol06, ISO08, INM05, JRH05, KT01b, Kuc06, Laz07, Ler01e, Lut03c, Mar05, MLJH04, Mil08, Mor03b, NK02, NP03, Omm01, Pap05, Pap00, Pet06, Pro01, RBC^{+}05, RBC^{+}06, Rum01, Sch03b, SML06, Sig04, Sim04b, SvR01, Ste08a, SKS08, SOT^{+}00, Sun02, Sur04a, Sur04b, USE01b, USE02, VLMO09, Wal02a, Wol03a, Wol03b, Zus03, dL05, AA02a]. JAVA [AL04b, Ano04-34, BMR02, BM03, BB01, CCR00, Fre01, Gal01, Gos00a, HP00, Hon05, HZC^{+}04, KKK04, LN02, LFP04, MZ04, MMU04, MLG02a, MSS00, NH02, OPS^{+}02, PFS05, PC03, Rog03, RWC^{+}03, Suo04, WAB^{+}04, WBL01, ZK04b, Zhu03, dSC05, AFF06, ÁMdB00, ÁMdBdRS02, AddS03a, AddS03b, ÁdBdRS05, ÁdBdRS08, ANN01, AF03, Ada05, AS03, AY05, AY07, AU02, dS02, Aki02, AJMJS02, AJMJS05, AA04, AMJS05, AL04a, AR08, Alb03, ADT03, ASCE03, AK01, ASS03, ABV00, ABLU00, ASS^{+}05, ACD^{+}04, AWE04, AC01, ACS02, AH03, AC06, AGH05a, APA04, ACGL01, ACFG01, ABG02, AG03a, AG03b, AG05, AČMN05, ABM^{+}03, ACZ05, Ams00, Ams02, AR03a, AR03b, Ana01, ALZ00, ALZ01, AAD^{+}01, AZ01, ALZ02, ALZ03, AZ04, ADDZ05, AAD^{+}07, And02, AF02, And04, ACL03, Ang01, Ano00e]. Java [Ano00f, Ano00g, Ano00l, Ano00n, Ano00o, Ano01d, Ano01f, Ano01h, Ano01i, Ano01j, Ano01l, Ano01k, Ano01m, Ano01o, Ano01p, Ano02a, Ano02c, Ano02d, Ano02e, Ano02f, Ano02g, Ano02h, Ano02j, Ano02m, Ano02n, Ano02o, Ano02p, Ano02q, Ano02l, Ano02r, Ano02s, Ano02t, Ano02u, Ano02v, Ano02w, Ano03a, Ano03e, Ano03f, Ano03g, Ano03h, Ano03k, Ano03l, Ano03m, Ano03n, Ano03o, Ano03p, Ano03q, Ano03s, Ano03r, Ano03w, Ano03i, Ano03t, Ano03u, Ano03v, Ano03x, Ano03y, Ano03z, Ano03-30, Ano03-28, Ano03-29, Ano03-31, Ano03-34, Ano03-35, Ano03-36, Ano03-33, Ano03-38, Ano03-32, Ano03-39, Ano03-40, Ano03-41, Ano03-42, Ano03-44, Ano03-43, Ano03-45, Ano03-47, Ano03-46, Ano03-48, Ano03j, Ano03-49, Ano03-50, Ano03-51, Ano03-52, Ano04d, Ano04b, Ano04c, Ano04f, Ano04j, Ano04g]. Java [Ano04h, Ano04i, Ano04k, Ano04l, Ano04m, Ano04n, Ano04q, Ano04o, Ano04p, Ano04r, Ano04s, Ano04t, Ano04u, Ano04w, Ano04v, Ano04x, Ano04y, Ano04z, Ano04-27, Ano04-28, Ano04-29, Ano04-30, Ano04-31, Ano04-33, Ano04-32, Ano04a, Ano04-37, Ano04-38, Ano04-39, Ano05a, Ano05c, Ano05b, Ano05d, Ano05g, Ano05f, Ano05e, Ano05h, Ano05i, Ano05j, Ano05k, Ano05l, Ano05m, Ano05o, Ano05n, Ano05p, Ano05q, ABH^{+}00, ABH^{+}01, A^{+}01, AP02, ABL08, Apr03, Apr05, AZ02, Apt02, AM02, AJB^{+}04, AH04b, AFT^{+}00, AFT01a, AFT01b, ABC^{+}07, Arm04, AGH00, AHKR01, AGG02, AHR02, AW00, Arr01, ASB^{+}04, Art00, AGMM00, AAA^{+}04, Atk01, ACR01, ACC^{+}01, AJ01a, ABI^{+}07, ABG^{+}08, Aus00, AGS01, ABF03, AV05, AW03, Aye01, ANH00, S.04a, BP01a, BHL00, BTS^{+}00, BH05a, BST00, BAL^{+}01, Bac01]. Java [BFG02, BCR03a, Bac03, BKMS04, BD03a, Bad00, BKH02, BH02a, BC07, Bag02, Bai00, Bai03, BC00, Bak00, BH02b, BCS07, Bal03a, BKT03, BCMT03, Bal02, BK08, Bar00a, Bar01a, Bar01b, BBDT02, BDT04, Bar05, Bar02a, BBBD01, Bar03b, Bar00b, Bar02b, Bar03c, Bar00c, BBM04, BFMW04, BI02, BS07, Bat03, Bat04, BAF03, BFN^{+}06, BDF^{+}00, Bea05, BP01c, Bec01a, Bec01c, Bee04a, Bee04b, BR01a, BP02, BCS02, BO05, BO08, BO09, BDRV01, BBGP01, BBG04, BHJR05, BBL03, BBS04, BZ05, BZ07, BN03, Ben00a, Ben00b, Ben00c, BNO03, Ber00a, BDN05, Ber02b, BB05, BD02, BDLM04, BHDS09, Ber00b, BF03, BM01, Ber05b, Bes01, BC01, BDP02, BCV03, BD03b, BLV03, Bet04, Bet05, BCV09, BCE^{+}01, BD04, BCH02, BP03a, BR02, BVPE06, BHV01, BL02a, BH04a, BH04b, BH05b]. Java [Bin06, BR06a, BSMV09, Bir01, BBHL01, BB00a, BB00b, Bis03, BHW05, BSH^{+}01, BGH^{+}06, Bla03, Blo01, BG05, Blo08, BAD^{+}09, Bod04, Boe05, Bog00, Bog01, BG04a, BL04, BI07, BF02, BV05, BML01, Bol00, BALV03, BDT01, BDFL04, BGadH06, BHP^{+}01, BS00a, Boo00, BS00c, BS00b, BS04, BPSH05, BG04b, Bos04, Bot03, BH03, Bou01, BHK^{+}04, BOT02, BM04, BL03, BDJ^{+}01b, BS00d, BR01b, BKM02, BSBR03, BBV03, BA09, BW01a, BAJ01, BWW^{+}03, BR01c, BALP01, BALP06, BD01a, BLW00, BP01d, BP03b, Bre02, BJvdB02, BA01, BLL06, Bri05, BP05, BRU04a, Bro01, Bro00, BVD01, BH02c, Bro03a, Bro03b, BW03a, BW03b, Bro04, Bro05, BF04, Bru05b, BB03, BCL^{+}06, Bru03, Bru02, Bru04c, Bru05c, Bru06, BFMT00, BKY^{+}03, BKLS00, BKLS01, BKŁ01, BFM^{+}02a, BFM^{+}02b]. Java [BFS^{+}03, BFW^{+}03, BFS^{+}04, BLPV04, Bud00, Bud01, BRC03, BK01a, BK05b, BJK07, BK01b, Bul00, BSW^{+}00, BK00, BKO00, BSPF01, BSB^{+}03, BL02b, BCR03b, BRL03, Bur03, Bur01a, Bur01b, BC03, Bur02, BW01b, BW03c, BW04, Bur07, BE02, Bus02a, Bus02b, BGED04, CAF04, CFL05b, CFL05a, CL03a, CM05a, CW03a, CW04a, Cal04, Cal01, Cal02, Cal00b, CD01a, Cal03, CWH01, CMG^{+}01, CWWS03, CCC^{+}06, CCFG00, CHS01, CV01, CHV01, CV03, CGJ^{+}00, CFKL00, CFLL03a, CFLL03b, CP01, CP04, CGEN03, Cas02, CH02, CI01, Cav02a, CM05b, CLCC02, CWHB03, CB04, CR06, Cha00b, CWS04, CY02, CY04, CHMB04, CA04, CYH04, CQ05, CvE00, CC01, CC04, CMS05, Cha06, Cha00c, CJ02, CRL01, CZ01, Cha02, Cha03, Che00, CCT01, CX01a, CX01b, Che02a, CZ02, Che02b, CCW02, CG02, CSK^{+}02, CKV^{+}02]. Java [CN03a, CT03, Che03b, CLL03, CKV^{+}03, CY03, CO03a, CO03b, Che03c, Che03a, CW03b, CW04b, CM04, CHHC04, CCC^{+}04, CKK^{+}04, CWZ04, CM05c, CR05, CHL07, CCK^{+}08, CQX^{+}09, CM02, CHB03, CTF03, CY01a, CWY01, CKC^{+}02, Chi00, CN03b, CiLH01, CGS^{+}03, CCM05, CH08, CMS03a, CHL^{+}00, CMS03b, CKM04, Chr05, Chr01, CD01c, CD01b, Chr00, CBD01, CT00, CSK00, CKKH03, CL03b, CGR00, CLS00, CV08, CDF05, CMR05, CCSB04, CSFS00, Cla04, CSCM00, CF02, Cle01a, Cle01b, CLCM00, Coc02, CE01, CG01, Cog03, CHK^{+}04, Cog04, Coh02, Coh04, CGM06, CK05, CLN^{+}00, Col02, CCF^{+}02, CMS07, Col01, CGRR04, CR02b, CF04a, Coo02, Coo00, Cor00, CL08, CDFR04, CS02, CS03, CC03, CBGM03, CLN07, Cou01, CBD04, Cox01a, Cox01b, CCB^{+}01, CLP06, CHUB08, CCSA02, CS04, CHK00, Cul00]. Java [CLZ06, Cza00, D^{+}00, DS00a, DH08, DWH01, DHS02, DHPW01, DH04a, DGGD08, DT02, Dar01c, Dar03, Dar04, Dar07, Dau01, Dav05, DDDM04, DeP03a, DS00b, DK03, DTD04, DEK^{+}03, DDF^{+}03, DGMY06, DDS02, DD02a, DD02b, DD03, DD07, Dei08, DC01, Dek00, Dek06, DPT^{+}02, DJP02, DRV02, DL02, DYH05, DJ00, DJ02, DOR05, Dep03b, DC03a, DMU02, DS09, Des01, DC03b, Deu00, DiM04, DS00c, DFT03, Dib02, Die00, Die01, DMP05, DSCU01, DUK02, Dil00, DBC^{+}00, DAK00, DZHS03, DS04, DP08, Djo08, Dmi02, Dmi04, Dob01a, Dob01b, DV01, DKP00, DKL^{+}01, DGK^{+}03, DKTE04, DJLT01, DCA04, DA04, Dra00, DM07, DSBH03, DK02, Dro01a, DEJ^{+}01, DEL^{+}02, DLE06, Dro01b, DHWH03, DHRH05, DDHV03, DH04b, DHR^{+}01, Dun02, DMKN02, Dur02, DLS^{+}01, DG02]. Java [Dwe00a, Dwe00b, DJ01, Ead01, Ear03, EH04, ET01, ET07, Ebe02, EF02, Eck00, ET05, Eck02, EL02, EFN^{+}01, EFN^{+}02, EFG^{+}03, Edm09, EGD03, Eff00, Egy01, EvG02, EvG04, EXA^{+}05, EL01, ESS02, ELM^{+}04, EM04, EH07, EKEL01, EGLZ02, EFO08, Ell00, EQT07, EL04, ES05a, ÉJD01, EK01, ET02, Emu04, EK03, Eng02, Eng00, EKM00, ESS04, EGST08, Esp06, Esq04, Eub05, Eug06, EM03, ESPP01, FSS06, Fal00a, Fal00b, FMA02, FWL03, FFB^{+}00, FCF02, FC06, FCMR04, Fau02, Fei04, Fei01, FBR^{+}03, Fek08, FR02, Fel03, Fel04, FDTL02, FTD03, FT06, FCHE02, Fer07, FL02, FSBP03, Feu02, FVK01, FLMS06, FKR^{+}00, FMHH^{+}00, Fla00, FFCM00, FF00, FL01, FLL^{+}02, FFC02, Fla02a, Fla04a, Fla04b, Fla05a, Fla05b, FFLQ08, Fle03]. Java [Fle00, Fle01, FC01, FR00, FDR04, For04b, FF05, FS03a, Fox00d, Fox00e, Fox03a, Fox03b, Fox01c, Fox02, Fre05, FW02, Fre08, Fre04, FM03, FGLS04, Fri02, FL04, FK03, Fro08, Fry03, FRMW04, FMRW05, FP03, FOS^{+}04, FS03b, FLWW04, FBS04, FJ05b, FMMd03, GK07, Gad03, Gag02, GH01, GH03, GPF05, GPF08, GKM03, GKMZ04, GKW04, Gam00, Gam03, G^{+}01, Gar00, GNYZ05, GS01, Gar01, GCB^{+}00, Gat03, Gea00, GW08, Gee05, GS05b, GI00, GCRD04, GBED04, GBE07, GEB08, GK03, GV05, GP05, GJ04, GvLPF01, GP03, GGH^{+}03, Gho01, Gho04, GK08, Gib01, Gig00, GM05a, GM08, Gil00a, Gil00c, Gil01, Git00, Gle02, GHH01, GSV02, GPB^{+}06, Gol01, Gol04a, GGG03, GMW^{+}02, GS00c, GPS03, GCARPC^{+}01, GHM^{+}01, GDC^{+}04, GT97, GT01, GT04, GT06, GT10, Goo02b]. Java [Goo00, Goo03b, GM02, GN01a, GN01b, GJSB00, GJSB05, Got06, GW00, GEG07, GE08, Gra04, GH00, GF07, GHS05, GJ09, GEK01, GPW03, GPW05, GM00, GSaC05, Gri02a, Gri00, GV02a, GV02b, GV04, Gro02a, Gro02b, Gro02c, GM03, Gso00, GBCW00, GLC01, GAR03, GLS02, GS04, GW01, GCH00, GMM00, GSW00, GMT02, GM05c, Gut00, HG08, Hab04, Hac01, Hag00a, Hag00b, Hag02, HD02, HHK^{+}01, HHKS03, Hal02b, HG07a, HM00, Ham02, Han05a, HS00a, HKS02, HK02b, HJL00, Han05b, Hap02, HR00, HHM04, Har00a, Har00b, HS01, HKK^{+}01, HAL02c, HLT09, Har00c, Har03, Har04, Har06, HS00b, Har00d, HBR00, HL03a, HF06, HJL^{+}01, HM01a, HdJ01, Has02, HRAB05, HD01, HFL03, HL06, HSD04, HR04a, HR04b, HvE02, Haw02, HL04, Hef07, HMD04, Hei03a, Hei03b]. Java [HWM01, Hel07b, HCMM00, HD03a, HRD07, HRD08b, HL00, Hep04, HJR^{+}03, HW00, HPH03, HS05, HN00, HRE^{+}02, HRE^{+}05, HL02a, Hig03, HKI08, HT06, HIBP04, Hig04, HKHK03, Hir00, HG07b, Hit02, Hit03, HT03, HE03, Hoh03, HTY^{+}03, Hol04a, Hol04b, HJ01, HKL09, Hol00b, Hol00a, Hol00c, HD03b, HKS^{+}07, HKM^{+}09, Hoo05, Hor00a, HC00, Hor00b, HC01a, Hor02a, Hor02b, HC02, Hor03, HC03, Hor05, HKF00, HS02a, HPS02, HMRM03, HSSC05, HSB09, HWB03, HWB04, HYX05, HL02b, HL03b, HNZS03, HBX^{+}04, HBH01, Hub01, HOP04, HP04, HdS^{+}05, HCB04b, Hug02, HS02b, HJ00, HJvdB01, Hui02, HBD04, HB08, Hun00, Hun02, HL03c, Hun03a, HT04, Hun05, HC01b, HD03c, Hyd00, Hyu05, IKKM03, IPW01, IKKW01, IKN03, ISF06, IN09, IS03, II04a, Ish01, IKY^{+}00b, IKY^{+}00a, ITK^{+}03]. Java [IJ03, Iva02, Ive03a, Ive03b, IH01, ICB00, Jac01a, JR02, JP00, Jac01b, JP01, JLV02, JP03, Jac03, JKW03, JP04, JV04, Jac04a, JT04, JM00, JO03, JPC00, JR05, Jen00a, Jen00b, Jen02b, Jen01, JCP^{+}05, JSSM04, JA01, JH03, Jia00, JHJX04, Jia04, JWC03, JJ02a, JMS02, JBMP03, JKKL04, JCOP07, JC04, JCYC04, Joh03, JHA^{+}05, Joh06, JMSG02, Jol01, JK00, Jon02, JR03, JMM03, JPJ05, JHSL03, JJ02b, JKJ05, JPB^{+}08, Juo07, JRN00, JKH^{+}04, KK04a, Kaf00, KPPÉR06, KSK04a, Kal01, Kal04, KGH^{+}05, KOB01, KMR02, KT04, Kan02, KDH^{+}06, KF05, KHMW05, KT00, KPKL03, KKO02, KOO08, KKN06, KJBH^{+}00, KCSL00, KAN^{+}03, KGMO04, KCF01, Kes04, KFLN04, KFN04, KM04a, KM04b, Kic03, Kic04, KHBB01, Kie01, Kie02, Kil03a, Kil02, Kil03b, KC00, KH00, Kim02, KJ02]. Java [KTV^{+}04, KKL^{+}04, KVK^{+}04, KMEA04, KMOS03, Kin00, KC01, KM08, KMS04, KMSL03, Kle05a, Kle05b, KN06, KS01a, KBVP07, KK05, KNY03, KT01a, KA02, KR01a, Kno02, Knu01b, KM02, KK04b, Kod04, KW01a, KK03a, Kog04, KR00, KR01b, KB04a, KW02, Kon04, Kon03, KK03b, KM04c, KWM^{+}08, KLL03, KY03a, KY03b, KKJY04, KNN^{+}01, KPK02, KS02a, KS04, KC03, Kre01, KBP^{+}03, KW01b, KM01, KSK04b, Kro00a, KLS00, KNG02, KKT04, Kum04, Kum05, Kun02, KP01, KX04, KS01b, KS02b, KWK03, KWK05, LMV02, Lad01, Lag03, Lai08, Lai01, Lak02, LO00b, LO00a, LO03a, LO03b, Lam03, LP05, LSW08, Lan00, Lan04, Lan05b, LG99, LG00a, Lar01, LTOT07, Las02, LLMK03, Lau03, Lau04, LBR00, LP01a, Lau01, LBD^{+}03, Law02, Lea00a, Lea02, LST02, LST03, Lea00b, LDE^{+}02, LBR06, LS00]. Java [LYK^{+}00, LL01a, LT02, LH03a, LKL^{+}03, LYM04, LCFL04, LN04, LS04a, LC05, LJ07, LMK08, Leh02, LFM09, Ler01d, Ler01f, Ler02, Ler03, Les03, LP01b, LP06, LMG00, LL00, LB00, LL01b, LL03, LL01c, LH03b, LH04, LH05, LRSW00, LRW01, Li02, LBJ02, Li03, LZ04, Li04, LCS04, LCZ04, LBJ05, Lia00a, Lia00b, Lia00c, LPH01, Lia01, Lia02, Lia03a, LPH06, Lia03b, LL08b, Lik04a, Lik04b, LS03, LAT04, LLCF08, Lin03a, LHS04a, LHFL07, LSK^{+}02, Lin00, LDM04, Lin01, Lin03b, LS08a, LS08b, LG00b, Lit00, LM02, LY03, LZZ03, LW03, Liu03, LPSY04, Liu04, LYL^{+}04, LM04, Liu08, LAL02, LLdA08, LD03, LRO02, LHS03, LSW07, LHS04b, LS04b, LH02, Lot02, LEW^{+}02, LEW^{+}03, LLK03, LC04, LGFM05, LUH^{+}05, Luk04, LFH03, Lut00, Lut01, Lut02, Lut03a, Lut03c, Lut03b, Lyk02, LAB^{+}00, Lyo02, MWL00, MF07a, MVV^{+}01, MD00, Mac05]. Java [Mad01, MBED06, MS00a, MSG01, Mah02, Mah04a, MDS04, Mah04b, MB03, Mai03, Mak03, ML09, MPG^{+}00, MR00a, MAWW^{+}01, Mam01, Man01, MP01a, MPA05, MCLDP01, MR09, Mar01a, Mar00, MLVB05, Mar02, MZB00, MBMZ01, MdB01, MCLC02, Mas00, MI01, MCG03a, McC00a, McC00b, McC00c, McC00d, McC00e, McC00f, McC01a, McC01b, MFH01, McG04, MTSM03, McG03b, McK01, McL00, McL01a, McL01b, McL02a, McL02b, MF04, McL06b, McL07, Meh02, ME00a, MT07, Men00, Men03, Mer04, Mer00, Met01, Met02, MSF03, Mey03, Mid01, MH02, MF01a, MFSL02, MLG^{+}02b, MRR02, MRR05, MJ00, MAJC03, MSR03, MFRW09, Mil09, MS03, MH00a, Mls04, MMK04, MKM^{+}06, MSV05, MORW04, MORW08, MHC01, MK01, MM04, MC06, MP01c, Moo03a, Moo03b, MR02, MMG00b, MMG^{+}00a]. Java [MMG01a, MMG^{+}02, MMG03, Mor00, MWM01, Mor03a, MF03, MF01b, MB05, Mor02, Mos00, Mos05a, Mos05b, MR00b, Mul00, MKF06, MSSJ00, MKS^{+}03, Mur05, MJ06, NW06, NW07, NDS^{+}02, NK06, NAW06, NSI03, NHY^{+}04, NR06, NP01, NMMS01, Nar05, NW02a, Nas04, NRV00, Nau02, NPRC01, NC05, NLFA02, NKBM01, NMKB03, Nel04, NC04b, NW03, N^{+}00, New05, NM00, New01, New04, NW02b, NS01b, NB00, NB01, Gal02, NS03, NAR08, NK00, NK05, NZM03, NNS03, Nik03, NK03, Nil05, NIEH04, NE04, Nip03, NMH^{+}02, Nis02a, Nis02b, Nis03, NP07, Nol04, Nor00, NLC03, NCM03, OBr05, OHL^{+}05, Oak01, OW04, Och09c, Och09d, Och09b, Och09a, OJJ00, OS02, Oes01, OMK04, OKN01, OKN06, OKN02a, OKN02b, OKN02c, OSM^{+}00, Oi05, Oi06, Oi08, ONRV08]. Java [dOHS^{+}03b, OGA^{+}01, Ols07, Ols01, OK04, Omo03, OKK04, OL01, Our02, OWR04, OOM^{+}07, PKF02, PKF03, PDCL02, PV03a, PVC01, Pal02, PL01a, Pan04, PH00a, PSM01a, PSM01b, PSM03, PT09a, PTML09, Par04a, PPJ03, PL01b, PP02b, PP02a, PC04, Par04c, Par04b, PZ00, Par00, Par05, PDV01, PV04, PH03, PH04, PE06, Pau01, Pau03, DJM^{+}02, PSDF01, PL03, Pay04, PV03b, PR03, Pel03, PH00b, PSW07, PGM^{+}05, PRB07, Per02, Per04, Pet03, Pet05, Pew00, PUF^{+}04, PG00, PHN00, PBG^{+}01, PV06, PCC00, PWN04, Pil04, PG03a, Pip03, PNKN04, PFJ05, Pla00, PM00, PM01b, PCC01, PL05, PQVR^{+}01, Pon03, PWC00, PNCB06, Pot04, Pra03, PSH04, Pra08, Pre03, Pre00b, Pri01, PP02c, PR04, PJ09, Pug00, PS03, QGC00, Qia00, QHV02, QH03, Qui03, RRP00, RFZ08]. Java [RTJ00, RVJ^{+}01, RM07a, RWL07, RHR02, RP03a, RV05, RS00a, RSH01, RM04, Ran03, Ran02, RH04, RH07, Rao00a, Rao00b, Rao00c, Rao00d, Rao00e, Rao00f, Rao01a, Rao01b, Rao02, Rap03, RRP01, RWZ09, RW03a, RK02, Red01, Ree02, Ree00, Ree03, Reg00, Reg02a, Rei00a, RR02, Rei00b, Rei00c, Rei03, Rem01, RST^{+}04, RCR06, Ren00, RE01, Ren02, Req03, RWH01, RT02, RM08, Ric01, RMHC09, Ric06b, Ric00, RTVH01, RCB01, Ril02, RCB03, Ril03, RSD01, Rob00b, Rob07a, Rob04b, Rob06, RM00, Roc01, Rod01, RJFG03, RPJ04, RB04, Roe00, RKK03, RCdBL02, Ron01, RR01, Ros02a, Ros00, RVZ04, Ros02b, RS00b, RPP07, Röß06, RC01, Rot02, Rot05, RMR01, RMR03, RMR04, RKG04, RJGH06, RW03b, Ruf00, RYD^{+}03, RAC^{+}04, RGN07, RLR00, RS01, RP03b, RW04, SMK02, S.04b, ESGS00]. Java [SMCS04, Saf02, SU03, SGV04, Sah02a, Sah02b, Sah00, Sah01, SH03, SA02, SSM03, Sak01, SR05, Sal04, SBAD01, Sal06, SSD^{+}03, SM01a, SC01a, SLPO02, SC02a, SDPM04, San02a, San03, San04a, San00, SV05, San02b, SMBZ07, SJG03, SF01, SD01a, SC07, Sat02, SL07, Sav01, SEdM08, Sch00a, SO00, Sch01, Sch03a, Sch04a, SH04a, SLB^{+}02, SG00, Sch03c, Sch04b, Sch04c, SD08, ST04, Sch02, Sch04d, SM04a, SLC03a, SBCK03, SBB05, Sch00b, SPS^{+}02, Sci07, Sco03, Sea02, Sed03, See04, SAWW01, SE04, Sel03, SAFG03, SBMG00, Ses00, Ses02, Ses05, SS07, Set03, SCBH09, SCB09, SFMH01, SYAS05, SKS01b, SKS01a, SKS03, SB07, Sha00a, Sha00b, SY04, SJ01, Sha01, Sha04, SPB01, SR06, SSB03, SK00, SCS01, SG02, SM01b, SM03a, She01b, SRW^{+}00]. Java [SK04, Shi03a, Shi00, Shi03b, SEGS03, SM01c, SSM04, SSGS01, SGF^{+}02, Sib00, SW01, SB03b, SB05, Sig04, Sik03, SMS00, SV02, Sim04a, Sim04b, SK08, SFP03, Siv02, Siv04, SSV05, Ska00, SC02b, Sla00, Sma08, Smi01a, Smi01b, SBO01, SC08, SO02, SH04b, SNOM01, SSS02, SSS05, Soo01, SMS^{+}04, SC05, SRD00, SASZ03, Spe02, Spi03b, Spi05, SPGV07, SGSB05, SB06b, SLC03b, SPR^{+}03, SCLV04, Sta04a, SM01d, SZ00, Sta00, Sta01, SSB01, SS03, Sta04b, SHHS04, Ste01, SHB^{+}03, SS00b, SHK^{+}03, SM02a, Ste05, Ste04, SL00, SP03, SL01, Sto02b, Str02, SSP07, SC01b, SSA03, SQG^{+}05, Str01, SM04b, Stu07, Stu01, SBA01, SCH05, SJ05, SYK^{+}01, SYN02, SYN03, SOK^{+}04, SYK^{+}05, SD04, SRJS08, SHR^{+}00, Sun01, SKP^{+}02, SL04, SG03, SSL02]. Java [SM02b, Sur01, Sur04a, Sur04b, SSE05, Swa01a, Swa01b, SKM01, TTD03, TGB^{+}04, TGV^{+}01, Tam00, TC03, TM07, TYS04, TSL^{+}04, TBSN01, TSDNP02, TTPN08, Tat02, TG04, Tat05, TRVH03, TSCI01, Tddd03, Tay02, TA04, TB00a, TS01, Ten00, TP01, TDB00, Thi02, TMG03, Tho03, TOG^{+}05, TCF^{+}03, TS02, TS04, TS09, Tim03, TSL^{+}02, TSL03, TCC01, TCC02, TCSC02, TCSC04, TP02, Top02a, Top03, Tor01, TH02, TFL^{+}04, Tra00a, Tre05, Tre02a, Tre02b, Tre03, Tre04, Tre02c, THMT03, TC04, TE05, TCM^{+}00, Tui04, Tul08, TZ01, TT01, TVMB03, USE01c, Uni02, Uni03, Uma02, UL08, Urb09, Utt06, VV05, VT01, Van04, VVG^{+}05, VWS^{+}05, VDPC01, VDPC03, VUPB02, VN03, Vau03a, Vau03b, VKB01, VHBB01, VHBB03, Vel01, VED06, VED07, VAB^{+}00, VMMF00, Vie03]. Java [VKK^{+}01, Vil00, Vil08, VB01a, VHL01, VMWD05, VDMW06, Vir05, VN00, Vir03, VPK04, VL00, VB01b, VP05, Vrb03, Wad00, WG01, WACBL03, WCS00, WG02, Wal03a, Wam02, WS01a, WS01b, WWSL02, Wan02a, Wan03a, WLW^{+}03, WSVX03, Wan03b, Wan03c, Wan04, WXW^{+}05, Wan05, WWJ07, WR08, WW09, War02, WF04, WB00, WB01, WFGK03, Way03, Way05, Wea00, WP04, Wea07, WGC09, WCCL05, WVMN05, WVE^{+}00, Wei02a, Wei04, Wei01, WJH05, WJH06, WS01c, WHBS01, WAF02, Wel02, WP03, Wel03, Wel04, WCC04, Wel06, WC00a, WC00b, WD00, WL04, Wen05, WTV03, WTV05, WM00b, Whi03a, Whi03b, WW06, WH01, Wic03, WP00a, Wil02, Wil01a, Wil04a, WA04, Wil06, WPN08, WDSD02, Wil04b, Wil05, Win01, WR00, WK02, Win02, Win04, WN01, WHW01, Wis06, WF00, WF02, Wit05, Wol01a, Wol04]. Java [Wol03b, Won03a, Won03b, Won04, Won05, WGW04, Woo05, Woo02, Woo03, Woo04, Wra01, WWMG06, WP00b, Wu01, Wu05, Wut00, XSaJ08a, XSaJ08b, XP04, XAN07, XSD07, XC01, XZ03, XX04, XX05, XYC05, Yah01, Yam04, Yan02, Yan05, YKS^{+}02, YL03, Yan03, YDWL04, YME05, YLL^{+}07, YWZ03, YHL01, YHL04, YHGL01, YdOLS^{+}05, YK03, YE04, YMP^{+}05, YCFX09, You02, YLW04, YLW08, Yua02, Yua03, Yua04, YAW02, YTY00, ZCR^{+}06, ZFA00, Zam03a, Zam03b, Zar02, ZW08, Zea00a, Zea00b, ZD02, ZS01a, ZGB03, ZG04, ZL05, ZYZ06, ZR07, ZLG08, ZK09, ZXNH02, ZPV03, ZCQS04, Zha05, ZSZ^{+}09, ZFK04, ZYC03, ZX05, ZT02, ZWL03, ZAVT03, Zhu04, Zuk01, ZHC04, dH05, dSC06, dCG^{+}02, dGNv04, deC04, dD01a, dM04, dOHS^{+}03a, dBdd04, dFR04, vHMB08, vNKB01, vNMW^{+}05, vNMKB05, vRKS01, vRKS03, vRS05, vdBJ01, vMV05]. Java [vdL02, vdSPP05, vD04, vLSM01, vLFGL01, vLGL^{+}02, vLH05, vO01, Ano04e, Gla06, Mas01, Ano00b, Ano03b, Ano01a]. Java-Anwendungen [Wol03a, Zus03]. Java-Applets [BL04, DK02]. Java-Applikationen [Ste08a]. Java-based [Lex02, ZK04b, PFS05, WAB^{+}04, MAWW^{+}01, ABG02, AG03b, Ano01o, Bal03a, CKKH03, CGRR04, EM03, FSBP03, FVK01, FGLS04, GLS02, HL03b, JSSM04, Li03, Lik04a, MB03, MCLC02, NPRC01, PDCL02, PGM^{+}05, SRJS08, SL04, TS01, TMG03, VT01, VB01a, Vrb03, WXW^{+}05, WK02, YHL04, ZCQS04, ZT02, dFR04, AK01, Ano00g, Ano01p, Ano03k, Ano03-29, Ano04n, Ano04-32, AZ02, BR06a, BDFL04, BKY^{+}03, BCR03b, CB04, CCT01, CM02, CHB03, CR02b, CL08, DPT^{+}02, DLL03, DZHS03, EL04, Fal00a, Fal00b, FMA02, FLWW04, GW08, Gra04, HL03a, HE03, HKF00, HdS^{+}05, HS02b, JT04, JCP^{+}05, JKKL04, KHMW05, Lik04b, LYL^{+}04, NHY^{+}04, NC05, NZM03, ONRV08, Röß06, Sci07, Sha04, SG02, SD04, Tre02c, Wen05, Woo03, YdOLS^{+}05, Zea00b, ZP03, dCG^{+}02, dGNv04, vNMW^{+}05, vNMKB05, vdSPP05]. JAVA-basierten [Lex02]. Java-Card [MdB01]. Java-Compliant [Ano01l]. Java-Component-based [VDPC01]. Java-DSP [SASZ03]. Java-Embedded [KFN04]. Java-Enabled [CKK^{+}04, GSV02, KPKL03, MWL00, RAC^{+}04, Tui04, Sak01]. Java-Games [Sel03]. Java-implemented [PSW07]. Java-Interface [VUPB02]. Java-like [KN06, CHK^{+}04, ELM^{+}04, AZ01, AZ04, ADDZ05, DGGD08, DEL^{+}02]. Java-Lösung [Ano04h]. Java-MaC [KKL^{+}04, KVK^{+}04, SSD^{+}03]. Java-MOP [CR05]. Java-Native [JKJ05]. Java-Oriented [BFS^{+}04, FJ05b, TFL^{+}04]. Java-Powered [AJB^{+}04]. Java-Programs [AGS01]. JAVA-Ring [WBL01]. Java-Scripting [KS04]. Java-Software [Ano04v]. Java-Specific [VKB01]. Java-Systeme [Wol03b]. Java-Technologie [Ano03-27]. Java-Technologien [Ano03s]. Java-tekhnologiiu [Saf02]. Java-to-JVM [SS03]. JAVA-Triggers [AA02a]. Java-XML [Lin03a]. java.* [All00a, All00b, All00c, All00d, All00e, All00f]. java.math [Cow01]. java.net [Gag02]. Java.nio [PS03]. Java.RMI [PM01a]. java.util.concurrent [Lea05]. java.util.regex [Hab04]. Java/ [SDPM04]. Java/C [Ano01k]. Java/C# [BS04]. Java/CGI [HL02b]. Java/CORBA [GCARPC^{+}01, LRSW00, LRW01, SRW^{+}00]. Java/CORBA-based [SRW^{+}00]. JAVA/JAVACARD [MMU04]. Java/Jini [AGG02, Gho01]. Java/JVM [BS00b]. Java/R [HLT09]. Java/R-based [HLT09]. Java/SQL [Ebe02]. Java2 [CK05]. Java3D [HJF06, Vor01]. JavaBean [FCW01, RAC^{+}02]. JavaBean-based [FCW01]. JavaBeans [BMH06, AA02b, BCCN01, Bro02b, DL00, Fab02, Jor02, JFt00, LYC02, LR04, LR05, Ler01a, Ler01b, MS01, MH00b, MH01, MH04, MHB06, Nyb02, PSS01, RAJ02, TJ00, Tre01, Tro04a, Tro04b, WF04, WCD^{+}01, XLG03, XOWM06, YAA07]. JavaBeans^{T M} [NT01]. JavaCard [AJ01a, MMU04, BDJ^{+}01a, BDHdS01, BDJdS02, BCDdS02, Jac01c, MP01b, PvdBJ01, vdBJP01]. Javacards [Cim02]. JavaCC [Kod04]. JavaCloak [RE01]. JavaFAN [FCMR04, FMR05]. JavaFX [CCB09, Ste08a, Ste08b, Wea07, WGC09]. JavaGrande [PBG^{+}01]. JavaHelp [Lew00]. JavaLog [ACZ05]. Javalon [Ano03-31]. Javalon-1 [Ano03-31]. JavaML [Bad00]. Javana [MBED06]. JavaNOW [TDB00]. JavaNws [KW01b]. JavaOne [Ano01e, Leh01]. JavaOS [HPB^{+}00]. JavaParty [PH00c]. JavaPod [BR01d]. JavaPSL [FJ01]. Javari [TE05]. JavaScript [Ano00d, Sto01b, Sto01a, Bro02a, AE06, AF02, Ang06, BMS02, CMJL09, Coo01, Cro08, DD02c, Doe06, EA06, Eic05, Est02, Fla02c, Fla02b, Fla06, Gab07, Gar09, Gen00, GW02, Gil00b, Goo01a, Goo01b, Goo02a, Goo03a, Goo07, Gos00b, GT00, Har00d, HP02, HRM00, II04b, Jen02a, Joh00a, Kah06b, KHFS09, KHKH01, Knu01a, Lab09, Lan05a, MJ01, MDS04, McF08, McK01, Mor08b, Mur00, NS01a, Pas04, Pol01, Pot08, PS01, Pow07, Ree01, She01a, Soj03b, SM03b, Tam00, Tha00, Tha06, TEM^{+}01, TB00b, Wat02, Woo01, YCIS07, ZJ03, Zdr09, CDH07, Ano00c]. JavaServer [W^{+}04, Zen02, AK00, Ber01a, Ber01b, Ber02a, Ber04a, Ber04b, Cha05b, D^{+}04, DBH04, FK00, Gea01, GH04, GH07, Hal00, Hal01a, Hal02a, Jor02, Kur04, Ler01c, Man05, Pek00, Tre00, Wal03c, Zen02, WMM04]. JavaSpaces [BP01b, BGZ00, Hal01b, NZ01, vdPE02]. JavaStat [HLT09]. JavaStatSoft [WC08]. JavaSymphony [FJ05a, JF05]. Java^{T M} [LMG01, SMES01, Caa00, MSU08, BD01b, CF00, CHS^{+}05, Dar01b, AGH05b, BD01c, Dic01, RB01, vD00, BHR02]. JAVAVIS [OS02]. javax.crypto [Win01]. javax.XXL [vdBDS00]. Javelin [NPRC01]. Javia [CvE00]. JaVis [Meh02]. Javiva [TZ01]. JaViz [KJBH^{+}00]. Javy [GGG03]. Jawa [BRC03]. JAWIRO [SE04]. JAWS [Ano00i]. JAXP [Gri02a, Har03]. Jazzing [San04b]. JBits [AAA^{+}04]. JBoss [MD06, RG05]. JBSP [GLC01]. JBuilder [Ano00m, Ano03c, Lia00a, Lia00c, Lia02]. JCAF [Bar05]. JCanvas [Ano01l]. JCASim [FW02]. jcc [SJG03]. JCCM [CMG^{+}01]. Jcluster [ZYZ06]. JCOD [DJP02]. JComboBox [Wra01]. JCrasher [CS04]. JCS [Ano04r]. JCSP.net [WAF02]. JDBC [Ano03-36, Bal02, Bal03b, DUK02, Kie01, ME00a, P^{+}98, Ree00, Spe02]. JDBC-Based [Kie01]. JDeveloper [KM07, Ano04-29]. JDI [OS02]. JDO [Ano02q]. JDOM [Har03]. JDotter [BRU04a]. JDS [AH04a]. JDSL [TGV^{+}01]. Jeannie [HG07b]. Jedd [LH04]. Jeff [Cha05a, Coc02]. Jeli [Rob00b]. Jeliot [MMBAS04]. Jelly [Gös03]. Jenuity [vTNC08]. jeopardy [Ber05a]. Jeroo [SD03a]. JERPA [ET02]. Jerry [Ano00c]. JESSICA [MWL00]. JET [MLG^{+}02b]. JetBrains [Ano03-37]. JetForm [Ano00i]. JEWL [Eng04]. jFAST [WW06]. JFC [Gol00, Top02a]. JFLAP [LJ08]. JGAP [CCT01]. JGC [ZS01b]. JGraph [BH02a]. jGRASP [CH06]. jHISC [HFL03]. Jiazzi [MFH01]. JICC4 [Cha00a, Cha00b]. Jim [Ano00b]. JINEXT [FJ05b]. JINI [Edw00, YHL01, AGG02, Edw01, ER01, Gho01, Hua03, JJ02b, Kum01, Kum02, Nat00, New00, OW00, Sha00a, WA01, ZP03]. Jini-Based [Hua03]. Jini/Java [ZP03]. Jini/Java-based [ZP03]. JISGA [Hua03]. JIT [OSM^{+}00, Sch03a, TP01, THL03, dSC05]. JIT-compiler [THL03]. JIVE [GJ04, Rei03]. JJ [EKM00]. JKarelRobot [BS01]. jLab [PT09a, PTML09]. JMatch [LM02]. Jmeter [PL03]. JMFA [Ano02g, Uni02]. JML [CK05, JP01, Jac04a, LBR06, MMU04, PvdBJ01, TE04, vdBJ01]. JML-JUnit [TE04]. JMM [Kle05a]. JMM-Faithful [Kle05a]. JmmSolve [Sch04d]. jMonitor [KF05]. jMoped [SSE05]. JMS [HMD04, Ano02f, MHZG06, RG00, Rou02b, Yus04]. JMT [BCS09]. JMX [JM00]. JNDI [LS00]. JNI [GF01, NS01b, SCH05]. JnJVM [TGCF08]. JNuke [ASB^{+}04]. Job [Ano02q]. JOGR [DL02]. John [Fox01b, Azi06]. Johnson [Gla06]. JoiN [HdS^{+}05, YdOLS^{+}05]. Joint [ACM01d, CF04b, YHGL01]. jointly [SBH^{+}04]. Jolt [Ano03r, SAB08]. JOMP [BK00]. JOP [Sch03c]. JOPI [AJMJS05, AMJS05]. Journeyman [Bec00a, Bec00b]. Joy [Ano05i]. jPHYDIT [JCP^{+}05]. JPolicy [OWR04]. JR [KGMO04, OK04]. Jr. [JR05]. jRapture [SCFP00]. jRate [CS02]. JRE [Ano03e]. Jrpm [CO03a, CO03b]. JRT [ISO08]. JRuby [EL09]. JSBricks [BBBD01]. JSE [BP01a]. JServ [GW00]. JSetL [RPP07]. JSF [JF06]. JSP [Ano05k, BSB04, BSB08, Bro01, Bru03, Goo00, Har01a, M^{+}00, Mar01a, NP03, Per04, Roc01, Spi03a, Tay02, Wei02b]. JSR [Cow01]. jStar [DP08]. JSTL [Spi03a]. JTL [CGM06]. JTRON [Hac01]. JUDO [CLS00]. Juggernaut [Lut01]. July [AGG02, HR04b, IEE03a, Sib00]. jump [WG02]. jump-start [WG02]. Jumpin [Wol04]. jumps [JMK^{+}08a, JMK^{+}08b, JMK^{+}08c]. June [ACM00b, ACM01a, ACM01b, ACM05, Ano01e, Ano02i, LL08a, SY^{+}05, USE00a]. Juniper [Lut02]. JUnit [Bec04, For04b, Goe01, HL02a, HT04, Lou05, NP03, PL03, RS05, TE04, WACBL03, ZK05, Alb03]. Jurassic [INM05]. Just [Bar01a, Jia04, KMEA04, KNG02, ME00b, SSM04, SOT^{+}00, SYN02, Vel01, YLL^{+}07, dSC06, vdL02, For06, GES^{+}09, ITK^{+}03, LYK^{+}00, LYM04, LMK08, OOK^{+}06, SYK^{+}01, SYN03, SOK^{+}04, SYK^{+}05, Swa01b, Yua04, IKN03, IKY^{+}00b, IKY^{+}00a]. Just-In-Time [KNG02, dSC06, Jia04, KMEA04, ME00b, SSM04, SOT^{+}00, SYN02, YLL^{+}07, GES^{+}09, ITK^{+}03, LYK^{+}00, LYM04, LMK08, OOK^{+}06, SYK^{+}01, SYN03, SOK^{+}04, SYK^{+}05, IKN03, IKY^{+}00b, IKY^{+}00a]. JVM [Ano00a, Ano01b, Ano01g, USE01c, USE01b, USE02, And01, Ano02e, Ano03-38, AFG^{+}00, BNV08, BFN^{+}09, Dd01b, BS00b, CMB^{+}01, CG01, DBC^{+}00, DA02, FMR05, GD00, HO03, HO07, Lan02, LM04, Moo03a, PG03b, SBB05, SS02, SD01b, SD03b, SS00a, SS03, Sub08, Won03a, ZS01b, ZWL03]. JVM98 [GPW05]. JVML [Ber01c]. JVMPI [DeP03a]. JVMs [San04b, ZK04a, DAK00]. JWAVE [Ano00i]. JWS [BJ04, SO02]. JX [WFGK03]. JXP4BIGI [HNZS03]. JXTA [CY03, OGT02]. Jython [PR02, Bri02, Hig03].

Kafer [ZXNH02]. Kaffemik [And01].

KaffeOS [BHL00, BH05a]. Kak [Ano04e]. Kamiwaai [Hit03]. Kardon [Mar01b]. Karel [Bec01a, Ber06]. Kava [Bac01, Bac03, WS01c]. Kaveri [JRH05, RH07]. KDE [Ano00n]. keen [Ano03f]. Keep [Pau03, RFZ08]. Kelly [Fox01b]. Kemma [Kro00b]. KenyaEclipse [CT05]. Kernel [DS00c, BL02a]. Kevin [Dud06]. KeY [BHS07, SSS05, NM02, Gal02]. Killed [Way03]. Killer [Bar01a, Dav05, MA05, Hun03a]. kind [MPO08]. kinds [San04a]. Kinetic [SO02, BJ04]. King [Ano01a, Bar00a]. Kirchberg [GAR03, GAR04, GRR05]. Kit [Ano00k, Ano00m, Ano01j, Ano01m, Ano01o, Ano02p, Ano02r, Ano02s, BRC03, SHK^{+}03, Ano04-27, Kil03a, Mor08a, WMM04, vLFGL01, vLGL^{+}02, vLH05]. KLAVA [BDP02]. Klient [HJL00]. Knell [Nil05]. Know [Dar01b, Fit09, Pan04]. Knowledge [Cha05a, Han05a, OOOiM05, RVZ04, Zhu04]. KnowledgeKinetics [HL04]. knows [Ano05n]. Kodok [YAW02]. Kolb [Zen02]. Komfort [Ano03-27]. Kommentar [Wol03a, Zus03]. Kommunikation [Ano05a]. Konfiguration [Ano05a, DHMT00]. Kong [Uni01]. Konrad [Roj00]. Korat [BKM02]. KRAKATOA [MMU04]. Krause [Ano00d]. Kris [Ano00b]. kurz [SKS08]. KYZO [Ano00k].

lab [Rad06, Rou02a]. lab-based [Rad06]. label [ML00]. Labor [TCM^{+}00]. Laboratories [SDPM04, VWS^{+}05]. Laboratory [Dor07, FSBP03, SASZ03, And02, BMS02, Rio02, Wea04]. Labs [Les03]. Laminar [RPB^{+}09]. LAN [Ano02t]. Lange [Wol03b]. Language [Ano01n, Ano01o, AGH00, AGH05b, Bil03, Blo01, CFLL03b, Dar01a, Dar01b, DDDM04, Dmi02, FM03, FMMd03, GDC^{+}04, Gös03, Gos00a, GJSB00, GMM00, HKK^{+}01, ISO08, JP01, JR05, JSSM04, KSC^{+}00, Kod04, KWK03, McK01, MMG01a, OK04, Par00, Sat02, Set03, Ste01, Ste00, Sun01, Vel01, VVV04, Wan04, WCD^{+}01, Won04, Ana01, Ano03h, Ano03w, Bad00, Bel02, BD01a, Bro09, BFMT00, CMC^{+}06, CR06, CMS06, CGM06, DM07, FCHE02, GJSB05, Hag00b, Ham02, HRM00, Juo07, KdJNNV09, KN06, LBR06, LCFkL05, LLK03, MF07b, MF09, MGB^{+}09, MSSJ00, Och09e, OJ09, PRB07, Rob04c, Ses08, SCH05, Swe06, TM07, VTD06, VS06, WAF00, WB00, ZKR09, Bee00, Way05, WCD^{+}01, WPN08]. language-based [WAF00]. Language-Dependent [Bil03]. Language-Specific [Dmi02]. Languages [AZ01, AZ04, ADDZ05, Fig00, Kil02, Pre00a, Pre03, Spi05, Wil06, Ano04g, AOMC07, BCHP08, Bro07, BW01b, BW04, Cro01, DGGD08, DH00, GES^{+}09, GS05a, HZS08, Hun03a, ISO08, JMK^{+}08a, JMK^{+}08b, JMK^{+}08c, Mau02, MSK09, Nam08, OJ09]. Lano [Dud06]. Lantronix [Ano00i]. Large [GP01, KT01b, McG04, MS03, CVW03, CHP^{+}08, CHL^{+}00, Die00, DG02, NZM03, OSH04, Req03, SCBH09, Wol03b, ZYZ06]. Large-Scale [GP01, KT01b, McG04, CHP^{+}08, CHL^{+}00, NZM03, SCBH09, ZYZ06]. Larkin [Bar03a]. Larne [Cal00a]. laser [PC03]. latching [MRB06]. latency [ABI^{+}09]. latent [BLLB08]. Latest [Ano02q, Whi03a]. LaTTe [YLL^{+}07, Ano01c]. Launches [Ano01k, Ano02q, Ano03-38, Ano02d, Ano03g]. launching [PC08]. Lava [Ano00i]. Law [GKM03, Wil03c, SPS^{+}02]. Layer [BCS07, JO03, Ano03-35, IK04]. layered [XOWM06]. Layman [Cha03]. layout [Ano03-50, KF00]. layouts [Hir07]. Layton [Ano02m]. Lazy [CiLH01, CCM05, Dek06, FC00]. LCH [Ano04y]. LDA [DZHS03]. LDAP [WD00]. Leaders [Ano01f]. leading [HD03c]. Leads [Ano03-38]. Leak [BM09]. LeakBot [MS03]. Leaks [HL00, MS03, BM08, DS00b, Wan03c]. leap [Mer04]. Learn [Ano02h, Smi01a, Ano05n]. Learned [DHRH05, Fit09, PE06]. Learning [CQ05, Cha03, Cha05b, DH04a, FOS^{+}04, HL03b, IEE03a, KB04a, Kum04, Les03, Mah02, NK00, NK02, NK05, PGM^{+}05, Pow07, SS07, SV02, TC04, WF00, BC07, BCM05, BBS04, CT05, ET02, Emu04, For04a, Ham07, MSK09, NSS^{+}05, Pan09, Rio02, VVV04, WF02]. Lecturelets [Cul00]. lectures [Cul00]. led [CF04a]. Legacy [BHP^{+}01, LRSW00, TSCI01, BKŁ01, LRW01, TT08]. LegacyJ [Ano01l]. legislation [Per01]. LEGO [Bag02, Bar02b, FL02, JCOP07, Wol01b]. Legos [LBD^{+}03]. LEGO^{T M} [LDB^{+}03]. Lehr [Ste08b]. Lehr-Programm [Ste08b]. Lemmatizer [Gal01]. lengths [Wol03b]. Lenguaje [Ano04-33]. Less [WA04]. Lessons [DHRH05, McG04, PE06, Kic04]. lets [Ano04f, Wil04b]. Letters [BHP^{+}01, DHR^{+}01, KSC^{+}00, LAB^{+}00, SLB^{+}02, SPS^{+}02, TEM^{+}01, TCM^{+}00]. Level [Ano01m, Fig00, GBED04, IJ03, RB01, SPR^{+}03, BFGS05, CMS03b, EGD03, GPW05, KS07, OGA^{+}01, ST09, Sto01b, vTNC08]. levels [BS01]. leverage [Urb09]. Leveraging [San02b]. liberated [KS07]. Libra [Ano00k]. Libranet [Ano00k]. Libraries [BHP^{+}01, CN03a, DKTE04, PP02c, CTLW03, Eub05, Fek02, HN00, Hig03, Wei02b]. Library [Ano01h, Ano01o, CKC^{+}02, DTD04, FFCM00, GMW^{+}02, Gro02a, GLC01, JSSM04, KF05, MMG01a, Pon03, RGN07, SHK^{+}03, TGV^{+}01, TSL03, WHKS01, Ano03l, BDRV01, Boe05, Fro08, HJvdB01, Lau04, LYL^{+}04, Mur07, RK02, RPP07, ST00b, War02, ZR07, vdBDS00, Aki02, CGG02, WACBL03]. Library-based [TSL03, ST00b]. life [Gat03, KS09]. lifecycle [LYC02]. lifetime [HBM^{+}06]. lifetimes [ISF06]. Ligands [HZC^{+}04]. light [HB08]. light-weight [HB08]. Lighter [TG04]. Lightweight [Bac01, BA05, BG04a, DJP02, HS00b, MS03, Ran02, Ric06a, Ros03, YME05, ZPV03, ZWL03, ACS02, Bac03, Bod04, BV05, CH06, Gar09, HCB04a, SAB08, vRS05, vTNC08]. Like [BN03, CHK^{+}04, ELM^{+}04, AZ01, AZ04, ADDZ05, BK00, BKO00, CGJ^{+}00, DGGD08, DEL^{+}02, Fei04, KOB01, KW01a, KN06]. LIMaS [WAB^{+}04]. Limit [GKW04, Ano04g]. limitations [BHJR05, HN00]. Limited [JMSG02, KK05, RTVH01, CH08]. limiting [ZSZ^{+}09]. LIMS [RB04]. Lin [Fox01b]. Linda [BGZ00, TDB00, WCC04, Wel06]. Line [MD00, SASZ03, BCS02, GM02, San04b, CM02]. Linear [Bar01b, GGHvdG01, Gam00, LFG00, OOM^{+}07, VDPC01]. Lineo [Ano00h, Ano00i]. Lines [Wol03b, Chr05]. lines-of-code [Wol03b]. Lines-of-Code-Metrik [Wol03b]. Linguistics [Wei01, Mas00]. linguists [Ham02]. lining [SYN02]. Link [AA02a, Ano03-30]. linkage [DZHS03]. linked [CZ02, DMU02, ZKR08]. Linking [Dro01a, FC01, MORW04, DLE06, FC00]. Linux [Ano00h, Ano00i, Ano00j, Ano00k, DHMT00, AH04a, Ano00d, Ano00j, Ano00n, Ano01k, Ano01m, Ano01n, Ano01o, Ano02o, Ano02p, Ano03x, Ano03-35, Ano03-39, Ano04-32, Gab07, HKS02, Hir00, Kro00a, Leh01, Leh02, MD00, She03, SKP^{+}02, Tim03, YKS^{+}02]. Linux-based [Ano00i]. Linux/Java [HKS02, YKS^{+}02]. Linux/RT [Ano00h]. Linux/Unix [Gab07, Ano03x]. Liskov [Lam03]. Lisp [Kic04, Nar05]. List [Rol05, Bru04b, Bru05a, Coo05]. listing [MDJ05]. lists [DMU02]. Literate [Dwe00a, Sah02a, Sah02b]. Lithium [DT02]. lithosphere [INM05]. Litigation [McG03b]. Little [Ano00k, Kic04, Vel01, Men03, Wil04b]. Littrow [PC03]. Live [Ben00c, NIKN06]. live-range [NIKN06]. LiveLessons [Dei08]. Liveness [SKS03]. LKH [PR03]. LLC [Ano00j, Ano00k]. Load [Ano01o, Ano02m, Chi00, Gou01, LCHY03, FJ05a, FT06]. load-balancing [FT06]. Load-Testing [Ano02m]. Load-Time [Chi00]. loaded [NW02b]. Loader [BC01, BHP^{+}01, KS01b, WBF^{+}06]. Loaders [Roe00]. Loading [Dro01a, TH02, ZHC04, LY03, QGC00]. Loads [BOT02]. LOC [Wol03b, Wol03b]. LOC-Metrik [Wol03b]. Local [DGK^{+}03, GSWZ08, HR00, Oi08, Sch03b, Whi03b, BAdMS08, KTV^{+}04, Oi05, SV05]. Locales [All00d]. Locality [PH00c, SGF^{+}02, FJ05a]. Localized [MAJC03]. Locating [KY03b, AHN02]. Location [ABM^{+}03, Hon05, Pau01, dFR04, BWW^{+}03, KTV^{+}04, YLW08]. Location-Aware [dFR04]. Location-Based [ABM^{+}03, Hon05]. Lock [EFJM07, KKO02, OKK04, MBS^{+}08]. locking [AFF06, RD06]. Locks [ACR01, BKMS04, Dic01, KKO02]. Loftus [Azi06]. log [SS06]. log-synchronization [SS06]. logging [Rob00b, Rob03]. Logic [Bec01c, BM03, Cal04, HJ00, JP01, Lut03c, Mos05b, vON02a, ONRV08, Qui03, vON02b, IS03, Mls04, PB08, Yah01, vO01]. Logical [DJ00, KY03b, DJ02]. logistic [CO06]. Loki [Ano00h]. Long [Kic04, ISO05, LM06, LW03]. long-distance [LW03]. long-term [ISO05]. longer [Coh04]. LOOJ [BF04]. Look [EM04, Hun03a, Kro00a, SK04, CZ01]. Looks [Ano04m, Nis03]. Lookup [DJ00, DJ02]. LOOM [BF04]. Loop [Ano03-38, AGMM00, LH03a, MFSL02, XZ03, OGA^{+}01, vdBJ01]. loop-level [OGA^{+}01]. loops [Lan04]. loosely [PK00]. løsning [HJL00]. lost [MMN09]. Lösung [Ano03-33, Ano04h]. lot [Cro01, Hun03a]. Loton [Fox01b]. Lotus [Ano01i, Ano04n, Gar00, LZZ03]. Loughran [Mor03b]. Lovers [Ano03i]. Low [ABI^{+}09, BG04a, NSI03, SBCK03, CSCM00]. Low-cost [NSI03]. Low-End [SBCK03, CSCM00]. Low-latency [ABI^{+}09]. LR [KdJNNV09]. Ltd [Ano00i, Ano00j, Ano00k]. Ltd. [Ano00k, Ano01h]. LTL [Bod04]. luck [Hol04b]. Luna [HvE02]. Luxembourg [GAR03, GAR04, GRR05]. Luxembourg-Kirchberg [GAR03, GAR04, GRR05]. LVDS [Ano02p]. LynuxWorks [Ano02o].

M [Fox01c, IK04, USE01c]. m-commerce [IK04]. M20 [Ano00h]. M7 [Ano05o]. MA [Ano03b]. Mac [SML06, KKL^{+}04, KVK^{+}04, SSD^{+}03, Ano00m, Ive03b]. Machine [Ano00a, Ano01b, Ano01g, Ano02b, BOT02, CW03a, CF00, CiLH01, DHPW01, GM00, SSB03, SHB^{+}03, USE01c, USE01b, USE02, VL00, WM00b, WF00, AAB^{+}00, AFT01a, ABC^{+}07, ANH00, DBC^{+}00, EGKP02, Fal00a, Fal00b, GCARPC^{+}01, GPW03, GBCW00, Kim02, KN06, MSG01, MS00b, Oi08, Req03, SCEG08, WF02, YME05, YTY00, BD01a, BP01d, BP03b, Caa00, Cza00, DCA04, DLS^{+}01, FFB^{+}00, FK03, GGG03, HM01a, HWB03, HB08, Ive03a, JR02, JDJ^{+}06, JJ02b, Juo07, LMG00, LMG01, MSR09, Men03, MP01c, Oi05, Oi06, PRB07, Ran02, RB01, SMK02, SH04a, SMES01, Shi03a, Siv04, SSB01, SM02b, Sur01, WWMG06, vD00]. machine-checked [KN06]. Machines [BDJdS02, DEK^{+}03, G^{+}01, GSW00, SD01a, Vog03, vLSM01, ABL08, CH08, Cra06, DGMY06, EGD03, PV08, RHR02, TGCF08, VED07, BHDS09, CT03, MLG^{+}02b, SM01c, VED06, ZS01a]. Macmillan [Ano00k]. Macromedia [Ano02r, Ano02t]. macros [Kic04]. Made [Apr05, GF01, PR04, DW07]. MaDViWorld [FP03]. Magnetic [Gar00, VP05, dGNv04]. Magnusson [Ano00b]. MAI [KK03a]. MAI-17-3 [KK03a]. Mail [Bar01c, Pau01]. Mail4Me [Ple02]. mailing [Bru04b, Bru05a]. Mainsoft [Ano04f, Apr05]. mainstream [Swe06]. maintenance [Wol03b]. MainWin [OBr05]. majors [Gou06]. Make [Dmi02, Kie02, WVE^{+}00, Ano05q, Lan04]. Makes [Spi05]. Making [Bou01, YLM^{+}05, GKM01, Mer04, PWC00]. Malaita [INM05]. Malicious [Zdr09]. man [Pau08]. Manage [Ano03y, Jol01, Men00]. manageability [MW05]. manageable [Lee03]. Managed [ATBC^{+}03, CEG^{+}03, GK05, WK09]. Management [AA02a, Ano00h, Ano00j, Ano00n, Ano01n, Ano02m, Ano02p, Ano02s, Ano02t, BHL00, BKH02, BH04a, BH05b, CLCC02, CNB00, CKKH03, HIBP04, HTY^{+}03, JM00, JHJX04, JCKS04, KLL03, Kre01, Lut03b, MF01a, Per02, Rei00a, SMES01, SAWW01, Tre04, WS01a, YDWL04, YLW04, Ano05f, BHDS09, BSBR03, CH08, CHS^{+}05, Fer07, GSHO06, ISO05, JH03, KS09, Lex02, LLS^{+}08, MS00b, Mer00, OHL^{+}05, SJ01, Sha01, SGW01, Tro04a, Tro04b, Wol01b, ZP03, Lut03c]. Manager [Kro00a, Lag03, LRO02, HS05, Oga09]. Managers [Ros02a, Ano03-50, Coh04]. Managing [Lut00, Mer04]. MandrakeSoft [Ano00j]. maniacs [FL02]. Manipulating [GK05, DSCU01]. Manipulation [TSDNP02, CFL05b, CFL05a]. manual [CLN07, McF08]. Manufacturing [CKKH03, LRO02, AZ02]. Many [Lea00b, Mid01, Ano03-43, Cro01, Hug02, Kic04, San04a]. Map [Yua02, LDB^{+}03, MM04]. Maple [And04, Ano01n, Kun02, LP05, LS04a]. Mapping [FMMd03, HBR00, YLL^{+}07, WK08a, WK08c, WK08b]. MapXtreme [HD03b]. MapXtreme/Java [HD03b]. Marching [SGV04]. MARIAN [GMW^{+}02]. Mark [Fox01b, Vau03a, Zen02]. Market [San02b, Ear03]. Marketing [Lut03a]. marking [BGNM04]. Markov [War02]. Markup [JSSM04, WCD^{+}01, Bad00, YLM^{+}05]. Marmot [FKR^{+}00]. MARS [VS06, Ano04-39]. marshaling [CFKL00]. mart [SL06]. Marty [Hal01a]. mash [GMM09]. mash-ups [GMM09]. Masked [QM09a]. mass [Wol03b]. Massachusetts [AGG02]. Massively [FP03, HdS^{+}05, YdOLS^{+}05]. Mastering [D^{+}04, GDB02, PKC01, RAJ02, HL02a]. Masters [Lut00, Sim04b]. Mastery [Mls04]. Matching [Dwe00b, FR00, LM02]. Materials [NLFA02, Soj03b]. math [Fos03]. Mathematica [LP05]. Mathematical [Ano01n, SCWL08]. Mathematics [EH04, CF04a, CF04b]. mathematics/computer [CF04b]. MathML [Ano02i]. MathType [Ano02q]. MathWorks [Ano01h]. Matlab [SDPM04, LS04a]. Matlab-Based [SDPM04]. Matrices [LUH^{+}05]. matrix [GS04]. Matthew [Fox01b]. mature [Ras03]. Maven [MOL05, PL03]. Max [Ano00k]. May [ACM00a, ACM06, CNB00, Sch03a, Gen00]. Maya [BH02b]. Maze [RRP02]. McJava [KT04]. McMaster [Bar00a]. MD [IEE02a]. MDA [Dud06, Lan05b, MLJH04]. MDD [Ano01o]. me [Har01b]. means [Ano02u, Nis03, PH00c]. Measure [Mos00, KKG09, Van04]. Measurement [ACM00b, ACM01d, Ano02s, Ano02t, BOT02, FSBP03, Ano04c, CM02, FWR^{+}05, NM00]. Measurements [ACM00b]. Measuring [WK02]. Mechanic [Ano00m]. Mechanics [RKK03]. Mechanism [BM03, BL03, Jac01b, KC00, KM01, XZ03, CY01a, CY01b, FT06, New01, TCSC02, WAF00]. Mechanisms [BAF03, ET07, Fei01, RWL07]. media [Ano03g, FCHE02]. Medical [BG02, CE01, Mam01, VWS^{+}05, Bar09, HBX^{+}04, Pay04, SML06]. Meet [BD01c].

Meeting [BKY^{+}03, Lut01, SBH^{+}04]. Meets [Bet02, PPJ03]. megaflops [MMG00b]. mehr [Ano03-27]. melody [PT01]. member [KF00]. members [Bru04b, Bru05a]. Membrane [NC04b]. Memory [AW03, BMR02, BR01a, BG04a, CMB^{+}01, CKV^{+}02, CCM05, CC03, DC03b, GNYZ05, GPS03, HL00, HIBP04, JMSG02, Jol01, KH00, KK05, LMK06, MPA05, Mid01, MF01a, MS03, Pau01, SMES01, Sch04d, SLC03b, SCLV04, VKK^{+}01, YLW04, BHDS09, BA08, BM08, BSBR03, CCC^{+}06, CSK^{+}02, CKV^{+}03, Che03c, CH08, DS00b, GS00c, HLM06, KOO08, KTV^{+}04, KF00, LLS^{+}08, LLdA08, MS00a, MS00b, NR05, Oga09, Oiw09, PV03b, PWH00, Pug00, SSGS01, SC02b, ST06, VED07, Wan03c, WK08a, WK08b, WK08c, WK08d, YLW08]. memory-constrained [CKV^{+}03]. memory-hierarchy [KF00]. memory-limited [CH08]. Memory-Reference [CC03]. memory-safe [Oiw09]. MEMS [Ano02r]. mental [MFRW07]. Mercury [Ano02m]. merging [HKI08]. Merlin [Ano00k, HBM^{+}06]. Mersenne [Luk04]. Mesh [MH00a, WHKS01]. meshes [MCLDP01]. Message [ASS03, Ano02f, BC00, CGG02, DK03, GR07, JO03, JPJ05, KP01, PS03, Rao02, RMHC09, Sak01, SBA01, TTD03, TA04, YHGL01, CGJ^{+}00, Hap02, Har00e, MHC01, NMKB03, SZ00, Bak00, TDB00]. Message-Driven [DK03]. Message-Driver [Rao02]. Message-Passing [TTD03, SZ00]. Messaging [AGH05a, HMD04, Hoh03, YHL04, Yus04, Ano02f, Bru06, Hap02]. Messdaten [Ano04c]. Meta [Fab02, HZS08]. meta-AspectJ [HZS08]. Metacomputer [ESPP01]. Metacomputing [ES06, Gam03]. metadata [Ano02k, Lan04]. metadata-make [Lan04]. MetaJ [dBdd04]. metalocking [BS07]. metaphor [Mil09]. Metaprogramming [dBdd04, Kic04, TTPN08]. MetaWare [Ano01m]. Methacrylate [BD03a]. Methacrylate/ [BD03a]. Method [AV05, CO06, CSK00, Coh02, DEK^{+}03, DJ00, Fei04, GBED04, KSK04a, NMMS01, SGV04, SSS05, SP03, SYN02, Tddd03, TT01, Wan05, ZL05, Ano02j, BBG04, BS00b, DJ02, GPW05, IH01, JJ02a, LSW07, MORW08, OOM^{+}07, PM01a, Sha04, SHR^{+}00, Uni03, Wor02]. Method-Level [GBED04, GPW05]. Method-specific [CO06]. Methodology [KNY03, BZ05, KH00]. Methods [ACGL01, BO08, Bog00, BML01, Cas02, GGHvdG01, vON02a, RS05, SM07, vON02b, Bes01, FDR04, Hug02, Vir03]. Methyl [BD03a]. Metric [Wol03b, HKI08, SS08]. metric-based [HKI08, SS08]. Metrics [Lut03c, SDF00, DDHV03, ML09, Wol03b]. Metrik [Wol03b]. Metronome [BCR03a]. Metrowerks [Ano02p, Ano03-35, Kro00b]. Mexico [ACM00a]. Michael [Mas01]. Michigan [Pau01]. Micro [Ano04-33, BL02a, Eng00, GM05a, Yan03, Gig00, Knu01b, RTVH01, Gar00]. Micro-kernel [BL02a]. microarchitectural [EGD03]. microarchitectures [NW02a]. microarray [Sal04, WAB^{+}04]. MICROBE [KS02b]. Microbenchmarking [Bru05b]. microbenchmarks [BBBD01]. Microcontroller [BP05, PUF^{+}04, RWC^{+}03, KBP^{+}03]. Microfibril [Uni02, Ano02g]. Microprocessor [Ran02]. Microscope [Ano03-39]. Microsoft [Ano02t, Ano03w, Ano03z, Ano03-36, Ano04f, Ano04g, Bar01c, DA04, Hun03a, Kil03a, Lia00b]. Microsystems [Ano02o, Ano05m, Van04]. Middle [Thi02, Mer04]. Middleware [ACD^{+}04, Ano00l, BD03b, CM05b, CLL03, CS03, HCB04b, Jac04b, JKJ05, JRN00, Kro00a, Zhu03, Ano05m, KHMW05, ZLG08, vHMB08, Jac04b]. MIDIet [Ano03p]. MIDP [RTVH01, Muc02, Tui04]. might [OBr05]. Mightier [Fos03]. mighty [Ano04-32]. MigraTEC [Ano01o]. Migration [Ano01o, CLL03, IKKW01, LLMK03, Sat02, XLG03, ZWL03, vLSM01, KLS00, MR09, SM01c, ZLG08]. Mike [Fox01b, Bar03a]. Mileage [BKH02]. Miles [Wil00b]. milling [Kim02]. million [Ano03j]. MIMD [KAN^{+}03]. Mind [Bar01c]. MINDSTORMS [Bar02b, EBG^{+}05, Bag02, FL02, JCOP07, LDB^{+}03]. Mine [RYD^{+}03]. MiniJava [Rob01b]. minimal [IPW01, Sco02]. minimise [Ano04d]. Mining [CHHC04, LL01a, WF00, Lot02, MR06, WF02]. MiniSQL [DHMT00]. Minolta [Ano00n]. MIPS [Ano04z, VS06]. Mirrors [CP04, CP01]. MISC [Sco02]. mise [Ano03m]. Misfeldt [Che05]. missed [PE06]. missile [CHMB04]. missing [McF08]. mission [Ano04-39]. Mistakes [Bec00a, Bec00b]. Mitchell [Fox01b]. Mix [Nis02b]. Mixed [CW04a, LHGM09]. mixed-environment [LHGM09]. Mixin [Bet04, KT04]. Mixin-Types [KT04]. Mixing [KBV08, NHY^{+}04, Wil04a]. Mixins [ALZ00, ALZ03]. MJ [CBGM03]. MKS [Ano03-40]. MM04 [CCC^{+}04]. MM04-1 [CCC^{+}04]. MobCon [CM05b]. Mobil [RTVH01]. Mobile [Ano00m, Ano01i, Ano01j, Ano01o, Ano02m, Ano02o, Bar03a, BCH02, BR06a, Bou01, BRC03, CM05b, CY03, CKK^{+}04, CCK^{+}08, ES06, FVK01, FGLS04, Hac01, IKKW01, Jon02, KSK04a, Law02, MD00, MR02, NP01, RC01, SSM03, SMBZ07, Sat04, Sig04, VB01a, WGC09, XX04, Yam04, YKS^{+}02, Yua03, dFR04, AHN02, Ano03-35, Ano04-32, BDP02, CW03b, EL04, Eng00, ESPP01, FC00, HAL02c, ICB00, LC04, New01, Tre02b, YMP^{+}05, vHMB08, Pau03, Sel03, Sig04]. mobile-code [New01]. mobile-platform [Ano03-35]. MobileRMI [AV05]. Mobilised [Par05]. Mobility [Bet04, Bet05, CWHB03, CGR00, GCB^{+}00, RP03b, RW04, AY05, AY07, AV05, BHK^{+}04].

MobJeX [RP03b]. Modal [GN01b, GN01a]. Model [Ano01o, Bac01, BFG02, BFG03, BS07, BD02, BM04, Bus02a, DL02, Dil00, Dro01a, GV02a, GV02b, Han05a, HD01, HP00, Hit03, JKJ05, LFP04, Lin03a, Lut03c, MPA05, MP01c, PDV01, RAC^{+}02, SA02, Sch04d, SCLV04, SL01, Sto02b, TS01, TCC01, TC04, VT01, Zam03a, Zha05, ZX05, ABG^{+}08, Bac03, BA08, BCL^{+}06, Bus02b, DLL03, DLE06, Gho04, GV04, GMM09, GM05b, HPH03, Hub02, JPS^{+}08, JJ02a, JF05, KN06, LL01d, MS00a, ML00, PG03a, PSS01, Pug00, RRP01, Req03, RHDB08, SV05, Soo01, TCSC04, Tor01, Uni03, WSVX03, WSP02, EK01, Lut03c]. Model-Check [HD01]. Model-checking [Sto02b]. model-driven [Hub02]. Modeler [Ano01n, Ano02m, Ing09]. Modeling [ACM00b, ACM01d, AGST04a, AGST04b, Ano01l, Ano01m, Ano01n, BD03a, CL03b, DFL00, FJ01, HECR00, JP01, JPJ05, MD00, NDS^{+}02, PP02c, TTD03, Aki02, Ano03q, BCS09, CR06, Fau02, Wen05, XOWM06]. Modelling [Che02a, Che03b, HdJ01, BJ04]. Models [Ais03, AW03, BBM04, HWB03, KX04, Mid01, RWH01, SPB01, SO02, Ste01, Bar02b, Cor00, KLS00, MFRW07]. Modem [Ano00i, Ano00m, Ano03-37]. Modern [AP02, CO07, GMW^{+}02, SM07, Lan05a]. modest [LS08b]. modification [Ano02e, Ano02u, Siv02]. Modular [BA07a, DJP02, DA02, BAF03, BCHP08, BFGS05, CLCM00, DCA04, FC00, Gri06, KdJNNV09, MRC03, MFRW09, MOS07]. modularity [DNR06]. module [CHB03, CBGM03, SSP07]. Modules [AZ01, YL03]. MoJo [NW02b]. Moka [dD01a]. Molecular [BL04, RGN07, Vor01, JCP^{+}05]. Molecule [Ber02b]. Molecule-oriented [Ber02b]. Molekulvisualisierung [BL04]. MOM [DJLT01]. Monad [JP00, SM04a]. monads [JP03]. Monetary [Arm04]. Money [LAB^{+}00]. Monitor [Bar00a, CWY01, Lia03b, Ano04d, CY01b, Cla04, IN09, Rob01a, VVG^{+}05]. Monitoring [Ano02n, Ano03-40, BCS02, BFM^{+}02a, BFM^{+}02b, BFS^{+}03, BFW^{+}03, BFS^{+}04, CR05, CCSA02, FBS04, FJ05b, HR04a, KF05, RT02, KL07, MC06, SPG07, WSVX03]. Monitors [AddS03a, Bec01b, Dic01, BH05c, BGED04, KPPÉR06, YME05]. monopoly [Lik04b]. Monotonic [Lik04a, Lik04b]. Monte [GKMZ04, PFJ05, War02]. Monte-Carlo [PFJ05]. Monterey [Ano01g, USE01c]. Mood [Lut01]. MOP [CHV01, CR05, CR07]. Moped [SSE05]. MOPs [CV01]. Morgen [Ano04c]. Morning [DHWH03]. Moronic [Lut03a]. Morphing [OBr05]. MorphJ [HS08]. mosaics [Bos04]. Most [TT01, Ano03-31]. Mostly [KKO02, BBYG^{+}05]. Motif [Ano00h]. Motion [Ano04-34]. motivated [Djo08]. Motivating [BVPE06]. motivation [Ges07]. Motocoder [Ano03-38]. Motorola [Ano02p, Ano03m, Ano03-37, Ano03-38]. move [Ano04f]. moves [CSFS00]. Moving [Law02, Lut03b]. MP [PS03]. MP3 [Li03]. MPEG [Wal02a]. MPEG-4 [Wal02a]. MPEGlets [Wal02a]. MPI [TDB00, CGJ^{+}00, CFKL00, CLL03, GR07, GGL^{+}08, LRW01, Rol08b]. MPI-based [LRW01]. MPI-like [CGJ^{+}00]. MPJ [BC00, CGJ^{+}00]. MPLS [XZ03]. MPU [Uma02]. MR [dCG^{+}02]. MS [LHFL07]. MS-Windows [LHFL07]. MSIL [LN04]. MSXML [TEM^{+}01, Hei01]. much [Way03]. much-needed [Way03]. Müllverbrennungsanlage [Lex02]. Multi [BIB05, CWHB03, Chr01, DL02, DOR05, Det01, DJLT01, DLS^{+}01, GN01a, LLMK03, MSSJ00, Och09e, RJFG03, VHL01, Bus02b, EFG^{+}03, FWL03, FDR04, GCRD04, GM05b, KS07, LJ07, MF07b, MF09, SCB09, SSC00, Sto02b, ZSZ^{+}09, JDJ^{+}06]. Multi-Agent [BIB05, Det01, VHL01, SSC00]. Multi-application [GN01a]. Multi-applications [DJLT01]. Multi-Body [RJFG03]. multi-core [SCB09, ZSZ^{+}09]. Multi-Dispatch [DLS^{+}01]. multi-instrument [Bus02b]. Multi-language [MSSJ00, Och09e, MF07b, MF09]. multi-level [KS07]. multi-methods [FDR04]. Multi-modal [GN01a]. Multi-Model [DL02]. Multi-paradigm [DOR05]. multi-server [GM05b]. Multi-tasking [JDJ^{+}06]. Multi-threaded [CWHB03, Chr01, EFG^{+}03, GCRD04, Sto02b]. multi-threading [FWL03]. Multi-tier [LLMK03]. multi-tiers [LJ07]. Multiagent [MSF03]. Multiagent-Based [MSF03]. multiapplication [HT06]. Multibody [KW02]. Multicast [Lut02, PR03, SBA01, Oes01]. multicastable [Nat00]. Multicasting [Lut02]. multicore [Sub08]. Multidimensional [MMG01a, MMG03]. MultiGen [Ano02m]. MultiGen-Paradigm [Ano02m]. MultiJava [CLCM00, CMLC06, MRC03]. Multilanguage [GD00, Sha02]. Multiline [Cox01a]. Multimedia [JWC03, dOHS^{+}03b, SEGS03, SL04, WVE^{+}00, WDSD02, dOHS^{+}03a, Ell00, FT00]. Multiparadigm [GvLPF01]. multiplatform [Sha02]. multiplatform/multilanguage [Sha02]. Multiple [CDNS07, FC01, MPTN08, TA04, BH02b, BHJR05, BLV03, BRU04a, CLCM00, DMP09, Fek02, KM08, Lyo02, MI01, Siv02, TB00a, WW09]. multiple-dispatch [BH02b]. Multiprocessor [MJ06, BAL^{+}01]. Multiprotocol [CGG02]. Multithread [LCS04]. Multithreaded [AddS03b, ÁdBdRS08, ABH^{+}00, ABH^{+}01, BP05, CC04, CT00, DRV02, EFN^{+}01, EFN^{+}02, FSS06, LB00, MP01a, PUF^{+}04, ÁdBdRS05, A^{+}01, BPSH05, KBP^{+}03, MC06, NR06, XSaJ08a, Yan02].

Multithreading [ÁMdBdRS02, BLPV04, GEG07, GE08, PV06, San04a]. multithreading-based [GE08]. Multitracer [Woo03]. multiuser [Sci07, ESGS00]. Murphy [SPS^{+}02]. Murtagh [Hec07, Hol06, Laz07]. Music [Li03, Per01]. Musicomputation [CKMP09]. Musings [SLB^{+}02]. must [Ano03z, NA07]. Mutable [BV05]. mutation [CTF03, OMK04]. mutators [MSLL07]. Mutual [Bro05]. MX [Ano02r, Ano02t]. My [Kie01, Kie02, Sea02]. MyEclipse [Ano05o]. MyFaces [STB08]. MySQL [DHMT00, Gab07, HJL00, Har01a, HF06, MCG03a]. Myths [Ano04s, BCM04].

N [Ano01a, Mar05]. Name [HT03, Lut02, Way05]. Naming [Ano02k, KM04a, Fei01]. Nanda [Fox01b]. NanoJava [vON02a, vON02b]. Nanotechnogy [Ano03-39]. NASA [Nat00]. NASA/CR [Nat00]. NASA/CR-2000-210329 [Nat00]. NASO [LPSY04]. National [Ano03-28, Ano02p, CVW03]. Native [BKLS00, BKLS01, HG07b, JKJ05, KNY03, PZ00, FS03a]. natively [Ano03-31]. naturally [Rol05]. Nautilus [FMMd03]. navigate [Eng00]. navigation [SPBE09]. Need [BH03, Fit09]. needed [Way03]. needs [OBr05, Pan04]. nelle [Pel03]. Nested [SCB09, NQM06, TGO00]. Net [Bar00a, Bel02, Jen00b, Lea00b, NDS^{+}02]. NetAdvantage [Ano03-41]. NetBeans [BGG^{+}03, Sur04a]. NetCONNECT [Ano00i]. Netfinity [Ano00h]. NetMAX [Ano00h]. Nets [LH03a, WDSD02, Bar01d]. NetSys [Ano00j]. Netware [JWC03]. Netweaver [Ano04-31]. Network [Ano00n, Ano01o, Ano02m, BB05, BC01, CM01, CLCC02, Coc02, ES05a, GS00b, Gil01, GCEO05, JHJX04, JBMP03, KLL03, Kro00a, MSF03, RLR00, Sat04, YDWL04, Ano03k, Ano03-34, ES05b, Har00c, Har04, HYX05, JMS02, LAL02, RR02, Sha00a, XOWM06]. Network-based [Kro00a, LAL02]. Networked [CT00, CT03]. Networking [ACM00c, ACM01c, ACM04, Ano00m, Gar00, JBMP03, SS00b, WAF02, Yan03, Ano03-32, Gag02, Tre02b, Zea00b]. Networks [BCS07, CCC^{+}04, GHM^{+}01, JKKL04, Lut00, Lut02, Nat00, SRJS08, Zea00a, dS02, CCK^{+}08, CM02, GCARPC^{+}01, JA01, OOOiM05, SM01a, TDB00, TBM09, Ano03-35, Kro00b]. NetworX [Ano00h]. Neural [Bar00a, GHM^{+}01, dS02]. neuroimages [VP05]. neutral [Per01]. NeuVis [Ano01l]. Never [Way03]. new-age [MFH01]. Newmark [JJ02a, Uni03]. News [Ano00l, Bar00a, Bar01a, Bar01b, Bar01c, CSFS00, Coc02, Eng00, Gar00, Got06, Lea00b, Pau01, Pau03, VN03]. Newton [GKM03]. NEXIQ [Ano02n]. Next [CF00, Fre04, HKS02, Yam04, BI02, JA01, Swe06]. Next-Generation [HKS02, Yam04]. NEXTGEN [SC07]. nically [Van04]. Niftiness [Par04d]. Nifty [Par04b]. Nijmegen [JP04]. Niklaus [BGP00]. NINJA [MMG^{+}01b, MMG^{+}02]. Ninth [USE00d]. NIO [Hit02, Rog03]. NIST [Dra00, Fal00a, Fal00b]. Nitin [Fox01b]. NitroX [Ano05o]. nitty [Way03]. nitty-gritty [Way03]. nixes [Ano04i]. NJ [Ano04e]. No [Ano03-30, For06, Ano02j, Ano03-44, Coh04, PT09b]. nodes [Ano03k]. Nolan [Ano00k]. Non [BR01d, CR06, HD02, Kle05a, Nat00, Ren00, VDPC01, WBL01, BBS04, Gou06, Sha00a]. Non-Cryptographic [WBL01]. Non-functional [BR01d, HD02]. Non-interference [Kle05a]. Non-invasively [Ren00]. non-Java [Sha00a]. Non-linear [VDPC01]. non-majors [Gou06]. Non-multicastable [Nat00]. non-novice [BBS04]. Non-null [CR06]. nonintrusive [BAL^{+}01]. nonlinear [VDPC03]. nonoperational [GS00c]. nonprocedural [Fau02]. NoodleGlue [Tre05]. Normal [JC04]. normalization [KBV08]. Norton [Ano01a]. Norway [SY^{+}05]. Notation [AR03a]. Note [Mam01, SSL02, TCC01, CY01b]. notebook [Ada05, GT05, MOL05, MF04, RG05, TGL05]. Nothing [DA04]. Notification [ASS03]. Novel [XX05]. Novell [Ano02m]. November [ACM00c, ACM01c, ACM03b, ACM04, GAR03, GAR04, GRR05, IEE02a, IEE02b]. Novice [ET05, WMC04, BBS04, CMS06, HB09, MFRW07, MLM^{+}08, PJ05, SB06a, SCL^{+}08, Soo09]. novices [BC07, SFM^{+}07]. NQL [Ano01n]. NT [Jen00a, Str01]. Nu [DNR06]. nuclear [Ano03-29, Man01]. Null [KKN00, BNK^{+}07, CR06]. NUMA [Ano00h, Oga09]. NUMA-aware [Oga09]. NUMA-Q [Ano00h]. Number [Mak03, Ano04g, Jam01]. Numbers [Dor02, Lut02, PG00]. Numeric [Wil03b, LP05]. Numerical [Ano01o, GKW04, GMM00, HRE^{+}02, HRE^{+}05, Mak03, Ste01, Bes01, Lau04, LFG00, MMG^{+}00a, MMG^{+}02]. Numerics [Ano00i, Ano01m, Ano01o, Ano02r]. Nuon [Bet02]. NuSphere [Ano01m]. Nutshell [Che02b, FCF02, OGT02, Ano00b, FC06, Fla00, FFC02, Fla02a, Fla04a, Fla04b, Fla05a, Fla05b, Har02, Top02b, Top03]. Nützen [Lex02]. Nvidia [Ano03-39]. NY [NIS00].

O [All00b, Ano03k, BDT01, Gri00, Har06, VT01, WC00a, WC00b]. Obfuscation [FS03b, SSM03, CY04, CDF05]. Object [AF03, AMJS05, Bac01, BFG02, BBC07, Bar00b, BHS07, Bes01, BB00b, BP01d, CHS01, CFKL00, CX01b, DDDM04, DL02, DFL00, ET01, EvG04, Gar01, GCB^{+}00, GDC^{+}04, Gun01, HS00b, HJR^{+}03, HJ01, Ing09, Ish01, JO03, Jia00, JRN00, Kaf00, Kal01, Kil02, Kil03b, Las02, LK01, LFH03, McK01, NDS^{+}02, NKBM01, OS02, PH03, PH04, RV05, RP03b, RW04, Sam04, SR06, SK04, SP03, USE01a, Vil00, WH01, Wic03, YHGL01, YLW08, ZL05, AJMJS05, Ano04e, Ano04-30, AW00, Bac03, BCV03, BA05, BP03b, Bud00, BRBY00, CZ01, CHP^{+}08, CF04a, CF04b, CH06, CHJB07, Die00, DSCU01, DMP09, DNR06, ET07, ET05, FX07, FWL03, Fei07, For04a, Gel00, GL08, HBM^{+}06, Hir07, Hun00, Hun02, ISF06, JPS^{+}08, JMK^{+}08a, JMK^{+}08b, JMK^{+}08c, KTV^{+}04, KR01b, LYC02, LT02, LH05]. object [LG00b, LS08c, LCC09, LFG00, MRR02, MRR05, MSK09, Mor00, MWM01, Mor03a, MH09, Nam08, NMKB03, NH02, NSS^{+}05, Off00, Pre00b, QM09a, RRP01, Ras03, Ril02, Ril03, SD03a, SML06, SAB08, SS08, ST06, ST00b, VTD06, VED07, VZGE07, Wam02, Wan03b, WSM06, WML02, Wor02, Wu01, Yan02, HRM00, LFM09]. Object-based [Ish01, NKBM01, Sam04, NMKB03]. Object-JavaScript [HRM00]. Object-orientation [BB00b]. Object-Oriented [Bar00b, BHS07, CX01b, DDDM04, GDC^{+}04, HS00b, JO03, Kaf00, Kal01, Kil02, Kil03b, LFH03, McK01, PH03, USE01a, Wic03, Bes01, EvG04, Gar01, HJ01, Ing09, Jia00, Las02, RV05, Ano04e, Ano04-30, AW00, Bud00, CHP^{+}08, CF04b, DSCU01, DMP09, Fei07, Gel00, GL08, Hir07, Hun00, JPS^{+}08, JMK^{+}08a, JMK^{+}08b, JMK^{+}08c, LT02, LG00b, Mor00, MWM01, Mor03a, Nam08, NH02, Off00, Pre00b, RRP01, Ras03, SD03a, SML06, SS08, ST00b, VTD06, Wam02, Wan03b, WML02, Wor02, Wu01, Yan02, LFM09]. Object-Passing [AMJS05, AJMJS05]. ObjectFX [Ano01h]. Objective [Urb09]. Objects [ACD^{+}04, ACR01, Bar03b, BBM04, BCH02, BF02, BRC03, CCM05, Git00, HRE^{+}02, JR03, KDH^{+}06, KR00, LS08c, NW03, PRR02, RP03a, Smi01b, TVMB03, YE04, YLW04, Yua02, Ano03-42, Ano04e, Ano04-30, BA07a, ESS04, GK07, HW00, IS03, IH01, JMM03, KF00, Kno02, Mai03, MR09, MR02, Rou02a, Woo04, XX04, W^{+}04, XLG03]. objects-first [Rou02a]. oblivious [CHL07]. Observation [Wil03d, SCFP00]. observation-based [SCFP00]. Observations [GHS05, SPS^{+}02]. Observed [Wan04]. Obtaining [AFT^{+}00, KCSL00, OOM^{+}07]. OC [Ano03-40]. oceanic [INM05]. OCL [RWH01, Rum01]. OCL-Constrained [RWH01]. OCL-Syntax [Rum01]. Octera [Ano03-31]. October [IEE03b, Jac04b, USE00c]. off [San04b]. off-line [San04b]. Offensive [BDJdS02]. offering [Kic04]. Offers [Ano01h, Ano01o, Ano03-37, Gar00, Ano02f, Ano03-36, Ano04f, Ano05b, Apr05, Way03]. Office [Ano00h, Ano00j, MD00, Ano03-35, Ano03-41]. Official [AL04c, Cog03]. Offloading [CKK^{+}04]. Offs [CKK^{+}04]. oft [Rol08a]. often [Hun03a]. Ogg [Li03]. ohne [Ano04v]. Old [Wil00c, MFH01]. old-fashioned [MFH01]. Older [SHB^{+}03]. Older-first [SHB^{+}03]. OMIS [BFS^{+}04]. Omnicore [Ano02p, Ano01o, Ano03-38]. OmniLinux [Ano00h]. omniscient [PTP07]. On-Card [Ler01f, Ano02v]. On-Line [SASZ03, BCS02, GM02]. On-the-Fly [CD01b, DKL^{+}01, Gar00, DKP00, LP01b, LP06]. One [Lia03a, LDM04]. One-Time [LDM04]. Online [Ano02q, AHR02, CQ05, Hoh03, Kum05, LAHC06, Pau03, SPG07, SPB01, TC04, Bow07, Hel07a, SCWL08, Wu05, ZJ03, BJ04, LS03]. Only [Ano03i, Bog00, Dil00, KPH^{+}09, SCWL08, Wit00]. onto [MRB06]. Ontong [INM05]. OO [Car06, Gri08]. OOD [AF03]. OoLALA [LFG00]. OOP [Ada06, BVPE06, Mad01, WP00a]. OOPtutor [Gel00]. OPAC [GMW^{+}02]. Open [AJMJS02, Ano00h, Ano00k, Ano01i, Ano01o, Ano02t, Ano03a, Bar01b, Egy01, GGH^{+}03, HE03, KR03, Kuc06, Mam01, Nas04, OSM^{+}00, SHK^{+}03, TBSN01, WACBL03, YLL^{+}07, Ano04i, Ano04-38, CG02, CLCM00, Eub05, FT00, HL02a, Liu08, MM04, Sta00, Sto02a, Vir05, Yua04, ZK05, CEG^{+}03, Pra03, SFP03]. Open-Ended [OSM^{+}00]. Open-Source [Ano01o, SHK^{+}03, YLL^{+}07, Mam01, Ano04i, Eub05, Liu08]. OpenCable [deC04]. OpenCard [HF00]. OpenDesk.com [Ano00k]. OpenGL [Ano03-36, XYC05]. OpenJIT [OSM^{+}00]. OpenLinux [Ano00i]. OpenML [Bar01a]. OpenMP [BK00, BKO00, KOB01, KBVP07]. OpenMP-like [BK00, BKO00, KOB01]. OpenOffice [CGRR04]. OpenOffice.org [Ano02t, Ano03-35]. OpenPath [Ano01i]. opens [Ano03-51]. OpenSML1.Net [Kil02]. opensource [Sur04a]. operate [Ano01f]. Operating [Ano01k, Ano04v, BTS^{+}00, LRO02, Per01, TFL^{+}04, USE00c, WFGK03, Ano03-44, Ano04-32, Lab09, NB00, NB01, Rob02]. Operational [ÉJD01, MF07b, MF09, Siv04, CVW03, FCW01, Moo06]. Operations [KKO02, SPB01, SW01, RD06, TCC02, TCSC04]. Operations-Research [SPB01]. operators [Ano03n]. opinion [Our02]. Opportunistic [BP01b]. opportunities [HKI08, LH05, SSGS01]. Opportunity [CM04]. OPT [FCW01]. optimal [TCSC02, See04]. optimalen [DHMT00]. OptimalJ [See04, Ano04j]. optimisation [dMSAV08]. Optimising [ACH^{+}05, YK03]. Optimization [AHR02, JRN00, KC00, KJ02, OKN02b, OKN02c, Rob01c, WH01, Zar02, AFG^{+}00, BBG04, BKO09, GCARPC^{+}01, ACM03a, MGM^{+}06, OKN01, OKN02a, PH00c, SMSAT08, SYK^{+}01, WCCL05, OKN06]. Optimizations [AR03b, VHBB01, YLW04, dSC06, CGS^{+}03, CLS00, IKY^{+}00b, ITK^{+}03, LAHC06, LOW09, SPG07, SSGS01, SYK^{+}05, VHBB03]. Optimized [Sch03c, BBGP01]. Optimizing [GCH00, LHS04a, OKN04, PQVR^{+}01, SMK02, VKB01, CHP^{+}08, FKR^{+}00]. Options [BR01c, KHMW05]. Opts [Bar01c]. OPUS [MSR03, Ros02a]. OpusJava [Lau01]. Oracle [DHMT00, Ano00n, Ano02s, Ano04-29, Ano05i, Bal02, Col02, KM07, Lak02, Lut03a, Pri01, Tho03, Wan03a]. Oranges [Lut00]. ORB [Won05]. Orcale [Ano05i]. Orchestra [TS02, TS09]. Order [BO08, Mam01, BO05, Nik03]. ordering [SMAT^{+}07]. Ordinary [LS04a]. O’Reilly [Ano00b, Ano00c]. organization [Juo07]. organizer [MS00b, SMES01]. ORGS [LS03]. orientation [BB00b, Hun02, KR01b, MH09]. Oriented [Ano02t, Bar00b, BHS07, BFS^{+}04, BRL03, CX01b, CR05, DDDM04, FJ05b, GDC^{+}04, HS00b, Hua03, JO03, JHJX04, Kaf00, Kal01, Kic03, Kil02, Kil03b, LFH03, McK01, PH03, PSDF01, SBA01, TFL^{+}04, USE01a, Wel02, Wic03, YDWL04, YHGL01, ACZ05, Ano04e, Ano04-30, AW00, Ber02b, Bes01, Bud00, CHP^{+}08, CF04a, CF04b, DSCU01, DMP09, EvG04, Fei07, FB07, Gar01, Gel00, GL08, HPB^{+}00, Hir07, HJ01, Hun00, Ing09, JPS^{+}08, Jia00, JMK^{+}08a, JMK^{+}08b, JMK^{+}08c, KH01, KKG09, Las02, LT02, LG00b, LFG00, MSK09, Mor00, MWM01, Mor03a, Nam08, NH02, NP07, Off00, Pre00b, RV05, RRP01, Ras03, SD03a, SML06, SS08, Swa07, ST00b, VTD06, VZGE07, VS06, Wam02, Wan03b, WML02, Wor02, Wu01, Yan02, LFM09]. origin [BNK^{+}07]. OriginLab [Ano01m]. Orsay [DPT^{+}02]. orthogonality [RFZ08]. Orthogonally [LMG01, MBMZ01, LMG00, MZB00]. OS/390 [DBC^{+}00]. OSDI [USE00c]. OSGi [Fri02, TV08, VVG^{+}05, Yua04]. OSGi-compatible [VVG^{+}05]. Oslo [SY^{+}05]. Other [Ano04s, Wil03c, Ano03h, Ano04b, BA07b, Mai03, STB08, SCH05]. Ott [SNO^{+}07]. Our [LAB^{+}00, dSC06]. Out-of-Process [RB01]. outil [FTD03]. outline [HBH01, Hub01]. Outlines [ÁMdB00, AddS03a]. Output [Ano08, BI07, Pra08]. Overcoming [CDF05]. Overflows [BK08]. overhead [OKN04]. Overheads [VKB01, LYK^{+}00, LLdA08]. overlapping [GV05, GP05]. overloading [BCV09]. Overview [AJMJS02, Dob01a, HR04b, Kum02, Ler01e, MLG^{+}02b, NB00, PB06, RB04, SOT^{+}00, Kum01, Rob01b]. own [SML06]. Ownership [BSBR03, CDNS07, PNCB06]. Oy [Ano00h]. OZ [MORW04].

P [APA04]. P2P [Coc02, Fle03, GR07, GGL^{+}08, PC04]. P2P-MPI [GGL^{+}08]. P3 [DC03a]. PA [ACM04]. PACAP [BCE^{+}01]. Pacific [Ano03-39]. Package [Bet04, Bet05, Men00, Win01, ZGB03, AK09, BDP02, BKŁ01, KW01a, MM04, Röß06, Sch04a, Wu05]. package/access [Sch04a]. Packages [And04, ZFA00]. Packeteer [Ano02n, Ano03-37]. PaCMAn [ESPP01]. pact [DA04]. Pad [LDM04]. Page [LMK06]. Page-based [LMK06]. PageRank [TMF05]. Pages [Ang00a, Ang00b, Ben00b, Ber02a, FK00, Hal00, Hal02a, Kan02, Ler01c, Pek00, Tre00, Wal03c, WMM04, Zen02, Ano00b, Ano00c, Ano01a, Ano03b, Ano04e, Ber01a, Ber01b, Ber04b, Gea01, Goo00, HP02, Jor02, Mur00, Pas04, Tha00, Tha06, AK00, DUK02, DBH04, Hal01a, Liu04, Sah01, Wut00, Zen02, Bro02a]. pagination [STB08]. pain [Ang06]. Paintbrush [EH04]. paired [Ano03k]. pairwise [FL04, LFM09]. Palm [Ano00n, Ano00n, MS00b, SMES01]. Palo [ACM01b]. Pan [Ano05n]. Panda [Ano03-34]. Panel [G^{+}01, MD00, Kon03]. Pantziarka [Ano05n]. Paper [ABH^{+}01, LD03, CY01b, Dmi04]. Papers [HR04b, GAR03, GAR04, AJ01a, GRR05]. paradigm [CF04a, CF04b, DOR05, FJ05a, GEVZ09a, Rob07b, VZGE07, Ano02m]. Paradigms [Swa01a]. paralel [FTD03]. Parallel [Aar06, AJMJS02, Ano00i, BGadH06, BKO00, CM01, CCFG00, CF03, CFLL03b, DT02, DK03, DL02, FJ01, Gam03, GCB^{+}00, GR07, GP01, Hyu05, KK03b, LK01, LCC09, MSM05, NPRC01, SM01b, SY^{+}05, SBO01, SCLV04, WFGK03, WHKS01, YHL01, YHGL01, vNKB01, ADT03, Bak00, BBYG^{+}05, BAD^{+}09, ESPP01, FJ05a, FLWW04, Gam00, GGL^{+}08, GEG07, GE08, HdS^{+}05, ICB00, KOB01, KP06, LP01a, MVV^{+}01, NC05, NZM03, Rol08b, SCBH09, SM03a, SMS00, TDB00, WK08a, WK08b, WK08c, Wen05, YdOLS^{+}05, ZYZ06, vHMB08]. parallèles [FTD03]. Parallelism [DFA03, FDTL02, SPR^{+}03, TCC01, BA09, FJ05a, OGA^{+}01, SCB09, XSaJ08a]. Parallelization [AGMM00, CA04, Fel03, WP00b]. Parallelizing [CO03b, CO03a]. Parameterized [AS03, BBM04, MRR02, MRR05, BR01b, HSB09, TP08]. Parameters [BO08, BW03c, BO09, LL01d]. Parametric [CAF04, VN00, CCKP06, IV06, Vir03]. Parasite [SSL02]. ParaSoft [Ano00j, Kro00b, Ano02n, Ano03-34]. Parent [Hig04]. Paring [BALV03]. Paris [HR04b]. Parkinson [Wil03c]. Parser [SG02, Car06, LLK03, vdSPP05, Way05]. Parsers [Met01]. Parsing [Par00, KdJNNV09]. Part [Ang00a, Bec00a, Bec00b, ISO05, ISO08, Ang00b, Lan04, She03]. Partial [HS02a, LHS04a, PL01b, DH08, LS04a]. particle [MLVB05]. particle-in-cell [MLVB05]. Partition [LLS^{+}08]. Partition-based [LLS^{+}08]. Partitioning [TS02, TP08, CLM^{+}07, CLM^{+}09, Sto02a]. parts [Cro08]. Passing [AMJS05, BC00, GR07, JPJ05, PS03, TTD03, TDB00, YHGL01, AJMJS05, Bak00, CGJ^{+}00, NMKB03, SZ00, Vir03]. passion [Pau08]. Password [Ano01o]. Paste [LN02]. PASTE’01 [ACM01a]. PastSet [PV03b]. Patching [Kal04]. Path [KNG02, CHL07, EL04, IV07, MCD09]. PathExplorer [HR04a, HR04b]. PathFinder [HP00, VPK04]. pathways [THMT03]. Pattern [Dwe00b, FR00, HHKS03, HK02a, HK02b, LM02, SP03, WBGM05, BR06b]. Pattern-Based [HHKS03, HK02a]. Pattern-Matching [FR00]. Patterns [ACM01e, BALV03, CHHC04, Coo00, DF03, GS08, Lut03a, Mah06, MSM05, NW03, NS03, SM02a, Bil03, CK03b, DS00b, FLMS06, FFSB04, GV05, GP05, Ges07, GM05a, Jia00, Lan00, Lea00a, Met02, Pre00b, WC08, Lut03a]. Paul [Ano00k]. pay [San04b]. payment [Has02]. PC [Ano00n, GEVZ09b, MD00]. PCs [Ano04t]. PDA [GW08]. PDAs [Ano02q]. PDF [ISO05, Ano02m, ISO05, Soj03a, Soj03b, Sto01b, Sto01a]. PDF/A [ISO05]. PDF/A-1 [ISO05]. PDS [AAB^{+}05]. PDZ [HZC^{+}04]. PE [Way03]. Peace [DA04]. Pearls [Ano00d]. Peck [Vie03]. pedagogic [ACS02]. Pedagogical [RRP00, Gri00, Ras00, Ras03]. Peer [CY03, GR07, MSF03]. Peer-to-Peer [CY03, GR07, MSF03]. Peers [Tui04]. Pekowsky [Cal00a]. pen [ABL07]. Pencel [Ano02o]. Pendulum [KK03a, SDPM04]. Pentium [Ano00m]. Perceptions [BBL03]. Perfect [Duc08]. PerfectBACKUP [Ano00k]. Perforce [Ano03-39]. PERFORMANCE [ACM01d, ACM00c, ACM01c, ACM04, ABG02, Ano01j, Ano02o, Ano02l, Ano03-41, BC00, BCMT03, BBHL01, BLW00, BA01, Bul00, CMS03a, CT00, CEG^{+}03, CS02, CS03, CCB^{+}01, Dra00, FJ01, GCB^{+}00, GP03, GGH^{+}03, GMM00, HECR00, HM00, HSD04, HS05, HN00, HCB04b, JR02, JRN00, KMOS03, KK03b, LG99, LG00a, Lau03, LMG01, LRSW00, McC00a, McC00b, McC00c, McC00d, McC00e, McC00f, McC01a, McC01b, MLG^{+}02b, Mos00, MSSJ00, NM00, PBG^{+}01, PS03, RWL07, Red01, RCB01, SD01a, SM01b, SPR^{+}03, SL00, SBA01, SM02b, TTD03, Vog03, WGW04, Woo05, XOWM06, Zea00a, Zea00b, ZS01b, ABLU00, Ano00l, Ano03t, Ano03y, Ano03-36, AGG02, Bar02a, BCS09, Bil03, BCM04, BDT01, BSW^{+}00, BGED04, CHL^{+}00, Coh04, CMP^{+}07, DAK00, Emu04, FWR^{+}05, Gam00, G^{+}01, GBE07, GEB08, GM02, GEG07]. performance [HF06, IN09, JJ02a, JMK^{+}08a, JMK^{+}08b, JMK^{+}08c, JK00, JKH^{+}04, KCSL00, KHBB01, KF00, KW01b, LAHC06, Lau01, LCFL04, LMG00, LAL02, LL01d, MAWW^{+}01, MLVB05, MI01, MHZG06, MMG^{+}00a, MMG^{+}02, MW05, NNS03, PJ05, PG03b, PV08, RHR02, RCB03, SPG07, SS02, SCBH09, Shi00, Shi03b, SKP^{+}02, TAW03, Uni03, WW09, Ano01j, Ano02q, PL01a]. Performing [Ano03-39, GBCW00]. perICS [ZW08]. perimeters [Ano03-34]. peripheral [Kon03]. Peripherals [Ano03-32]. Periscope [Pay04]. perk [Won05]. Perks [Won04]. Perl [Ano00m, SKS08, AF02, Ano00m, Ano01m, Cro01, Han01, HF06, Jen02a, MSR03, Pre03, SM04b, Stu07, Tan07, Wit05]. permissions [Nau02]. Persistence [ACD^{+}04, Ano02q, Atk01, PH04, WH01, ZL05, Bog01, BHK^{+}04, EFO08, WIC08, Woo04, Ano01l]. Persistence-Enabled [WH01]. Persistent [BH03, Bou01, MBMZ01, SMES01, AR08, LMG00, MZB00, MS00b, ST06, LMG01]. Personal [Ano00i, YKS^{+}02]. personalized [HSB09]. PersonalJava [Kro00b]. Perspective [BBL03, GP03, HdJ01, JP04, VKK^{+}01, DBH04, FPA^{+}06, Swe06, WBF^{+}06]. Pervasive [Yan05, AGG02, Ano03-40]. Perverse [Rol08a]. petaflops [CSFS00]. Peter [Ano03b, Bal03c]. Petri [Bar01d, LH03a, WDSD02]. PEVM [LMG00, LMG01]. Phase [GBED04, NK06]. Phase-based [NK06]. phases [KS09, RHR02, Rei05]. philosophers [Rob01a]. Phoenix [ACM03b]. Phone [Yam04]. Phones [Law02, Bre02, LC04]. Photogenics [Ano00k]. PHP [DHMT00, SKS08, Atk00, Cur07, HF06, SM04b, Stu07]. PHP5 [Gab07]. Phrasebooks [CCR00]. phylogenetic [DG02]. phylogeny [JCP^{+}05]. Physical [PGM^{+}05]. Physics [CBD01, VDPC01, VDPC03]. Physlets [CBD01]. picture [Ear03]. piece [Ano03h]. Pierre [IEE03a]. pilot [CKMP09]. pipe [Rob02]. pipe-fork [Rob02]. Pipeline [MSR03]. Pipelined [DFA03]. Pitfalls [MH02, BG05, D^{+}00, San04a]. Pittsburgh [ACM04]. PizzaBox [Ano00k]. PKI [Hoo05]. PL [KM07]. PL/SQL [KM07]. placement [AWS^{+}09]. plagiarism [Gib09]. Planar [ZG04]. Planet [Ano01k]. Planning [BALV03, EL04]. plapackJava [Gam00]. Plateau [INM05]. Platform [Ano00n, Ano00o, Ano01h, Ano01j, Ano01k, Ano01m, Ano02o, Ano02q, Ano03-38, Bag02, BDJ^{+}01a, BCDdS02, Bir01, BR01d, CI01, CN03a, CY03, CT00, DF03, DHPW01, DYH05, Dib02, FSS06, Gar00, GPW03, HKS02, HE03, IKKW01, JJ02b, KT00, KAN^{+}03, KJ02, Lai03, LN04, LRO02, MS01, NDS^{+}02, PSM01b, PTML09, Sun02, Vrb03, WMC04, WGC09, Ano03-35, Ano05q, Aus00, Cal01, CCT01, CHS^{+}05, DDS02, Eng00, FLWW04, Git00, Gri02b, Hal02b, Hap02, ITK^{+}03, KL07, LCZ04, LY03, OBr05, OG05, Pay04, PG03b, PG03a, Pir02, RA07, Ric00, RTVH01, Sha00b, Van04, CEG^{+}03, deC04]. Platform-Independent [FSS06]. Platforms [HKHK03, Kro00b, LZZ03, Ano04f, HKM^{+}09, MI01, SGW01, SOK^{+}04, WW09, ZSZ^{+}09]. Platinum [Lad01]. play [Bre02, Mor08a]. Player [Li03]. playground [MR00a]. Please [Ano03-52]. Plotting [ZGB03].

Plug [Ano05o, DHR^{+}01, HL00, Jen02b, FS03a, Kag09, Mor08a]. plug-and-play [Mor08a]. Plug-In [Jen02b, DHR^{+}01, Kag09]. Plug-ins [Ano05o, FS03a]. pluggable [ANMM06]. plugin [MM04]. PlugSys [Ano00k]. plus [Ano04-38]. Pnuts [KSC^{+}00, McC00g]. POC [TCC01, TCC02]. Pocket [CDH07, Fla02b, Bal03b, Bec04, Ber01b, Bur05, CK03a, FFB^{+}00, LL08b, Stu07]. PODS’08 [LL08a]. Point [Dar01b, Fig00, Ols01, SKC09]. Pointer [KSC^{+}00, KKN00, TCM^{+}00]. pointers [PWH00]. Points [CC04, LH03b, RMR01, BS09, CRL01, LH08a, LPH01, MRR02, MRR05, SGSB05, SB06b]. Points-to [CC04, LH03b, RMR01, BS09, CRL01, LH08a, LPH01, MRR02, MRR05, SGSB05, SB06b]. Poisoning [Zdr09]. POJOs [Ric06a, SB06a]. PolarLake [Ano02q]. policies [BLW09, GSHO06, KPPÉR06]. Policy [RWC^{+}03, GB01, JH03]. policy-based [JH03]. Polish [Vir05]. Polyglot [NCM03]. polygons [TP08]. Polymorphic [ADDZ05]. Polymorphism [RMR03, RMR04, BWC^{+}05, CAF04, VN00]. Polytonic [Lik04a, Lik04b]. Pool [Jol01, Wil00d, Li04]. Pooling [Vil00]. Poon [Fox01b]. Popkin [Ano01n]. popular [MHZG06]. Port [Han05a]. Port-and-Connector [Han05a]. Portability [JR02, SQG^{+}05, Wan02b]. Portable [BHV01, BH04a, BH04b, Bin06, CGRR04, Gle02, HWB03, MD00, RS00b, RW04, SMK02, SNOM01, TS04, VB01a, ABI^{+}07, ABI^{+}09, GCRD04, LHGM09, MZB00, WWJ07, ZAVT03, Ano03-33]. Portal [Kro00a, Ano04-39, LYL^{+}04]. portals [YAA07]. portals/portlets [YAA07]. Portfolio [Ano02s, Est01]. Porting [Apr05, Caa00, Shi03a, TCM^{+}00]. Portions [CK05]. Portlet [Hep04]. Portlets [Vie03, YAA07]. position [Dmi04]. Positioning [dFR04]. posium [USE01c].

POSIX [BW01b, BW04]. Post [DDDM04, GDC^{+}04]. Post-Java [DDDM04, GDC^{+}04]. poster [Bar01d, Hag00a, Soo01]. PostgreSQL [DHMT00, HTY^{+}03]. Potential [HZC^{+}04, Lea00b, BA09]. pour [FTD03]. Power [Ano00h, Bag02, DK02, Gar00, WP03, CMP^{+}07, RRP00, RRP01, Sma08, Way05]. Powered [AJB^{+}04]. powerful [CFS09]. PowerPC [Ano01o]. PowerWindows [Ano00k]. pp [Dud06, Azi06]. Practical [Bru03, Cal03, DFL00, Hag00b, LT02, Lut02, Mor03b, Pot04, RS05, Spi03a, Spi03b, SHR^{+}00, TSL^{+}02, Tul08, Wei04, WF00, BS00b, CD01a, CZ01, DP08, Eff00, Gar01, MD06, RPB^{+}09, Sik03, Spe02, Tha00, Tha06, WF02, Mil08]. Practice [CI01, GPB^{+}06, LST03, Mah04a, Rap03, SHB^{+}03, Bla03, Gib09, Hor02b, Mls04, MPTN08, UCJ^{+}04, ZABL09]. Practices [ACM01e, CMS03a, RT02, SH06, Eck02, FLMS06, Ree03]. Practicing [CLS00]. practitioners [Hun00]. Pragmatic [Cla04, GAG06, HT04]. pre [CKMP09, Jac04a]. pre-college [CKMP09]. pre-condition [Jac04a]. preassembled [Ano03-30]. Precise [WS01b, FF09]. Precisely [Ses02, Ano03u, Ano03v, Ses05, Bal03c, Ano03b]. Precision [LST03, LPH02, OKN04]. preconditioning [GEG07]. preconditions [CFS09]. predicate [MFRW09]. predicates [BKM02]. predication [JMK^{+}08a, JMK^{+}08b, JMK^{+}08c]. Predictability [LBJ02, LBJ05]. Predictable [Sch04c]. Predicting [Wat02]. Prediction [ABG02, CCF^{+}02, ISF06, JFH00, WK09, XOWM06]. Predictive [SS06]. Preference [Ish01]. Preferences [TCM^{+}00]. prefetching [CM05a]. Prefuse [EVS07]. Preliminary [LBR06, Gri03]. Prelude [Soo01]. Premature [Got06]. premium [Ano03y]. Preparation [GH03]. prepare [PB06]. prepass [IKN03]. Preprocessing [BO08]. Preprocessor [BO09, DC03a]. Presence [FC01, GCH00, SK00, CRL01, FYD^{+}08, FC00, LGFM05]. Presentation [Rum01, SL04, Ano04e, Ano04-30, You02]. presentations [BDFL04, Ano05j]. presenza [Pel03]. preservation [ISO05]. Preserving [LST03, SGF^{+}02, CHP^{+}08, DNR06, LST02]. Press [Ano03b, Bal03c, Cha05a, Che05, Gla06, Pet06]. Pretenuring [BSH^{+}01, BHM^{+}07]. prevalence [Ano03w]. preventing [PRB07]. Prevention [XZ03]. preview [Ano03-34]. priced [Ano04-29]. Prices [Pra03]. Primed [Ano05i]. Primer [Lut03c, PM01b, GAG06, MR00b]. Primitive [Our02, SW01]. Primitives [TTD03, Ano03l]. Princeton [Ano01i]. Principal [AZ04]. Principle [BH04b, LLK03, Ada06]. Principled [SD08, Bai03, Gri08, Kic04]. Principles [Juo07, LL08a, Ric01, Bai00, BH04c, Gra04, Jia00, Lea00a, Ril02, Ril03]. Printers [Ano03-32]. PrismTech [Ano02q]. Privacy [BD03b, ML00]. Prize [Bar01b]. Pro [Ano00i, JF06, Vir05, WGC09]. ProActive [XLG03]. Probabilistic [BM07, SGV04, CHMB04]. Probe [Ano01j]. Prober [Ano02r]. Problem [CP04, MLG02a, SS00a, TC04, CP01, HB09, HL03a, HSB09, LO00b, LP05, Mor00, Mor03a, Sla00, Wei02a]. Problem-Based [TC04]. problem-tracing [HSB09]. Problems [Eth01, FJ01, Lea00b, McL01b, MH02, SvR01, SHHS04, Utt06, BS00c, CG01, CLZ06, Hub01, Wil05]. procedural [VZGE07]. procedure [FCW01, HF06]. procedures [Ano03-42]. Proceedings [ACM00b, ACM01b, ACM04, IEE02a, ACM03a, IEE03b, SM07, USE00c, USE00d, USE00b, USE01c, USE01a, USE02, ACM00a, AJ01b, IEE03a, Tra00b, ACM00b, ACM05, ACM06, Ano01g, CNB00, LL08a, SY^{+}05, SBH^{+}04, ACM01d, Jac04b]. Process [BALV03, BGZ00, CLL03, CKKH03, DeP03a, DS00c, JV04, Lea00b, Pau03, RB01, WP04, Wel02, GMM09, Hun00, Joh00b, Kno02, MORW08, Rob02, VVV04, YL03, Dob01a, FPA^{+}06]. Process-Interaction [JV04]. Processes [BHL00, Aki02]. Processing [Boo00, Bru04c, BFS^{+}04, Bur03, BW03c, BG02, EGLZ02, Har03, Kod04, KC03, RLR00, SU03, Sat04, SY^{+}05, SSL02, Bur01b, Eff00, EvG04, Hun03b, KMSB08, MM04, Rol05, Sar03, WN05, dGNv04, vdBDS00]. Processor [Ano02s, EGLZ02, KFN04, LFH03, Sch03c, Sch04c, SLC03b, Won03a, Aar06, Ano03-31, KHMW05, RTJ00, SKC09, Whi03a, YMP^{+}05, YCFX09]. Processors [KFLN04, Omo03, BSMV09, DGMY06, EKEL01, OKN04, TCSC02, TCSC04, WB00]. Product [Kro00b, Mac05, See04, Vie03, Ano03-36, Ano04f]. Production [FOS^{+}04, RT02, SB00]. Productivity [Ano01l, Ano02t, Ano02d, LJ07, OBr05]. Products [Ano00h, Ano00i, Ano00j, Ano00k, Ano00m, Ano00n, Ano01h, Ano01i, Ano01j, Ano01l, Ano01k, Ano01m, Ano01n, Ano01o, Ano02m, Ano02n, Ano02o, Ano02p, Ano02q, Ano02r, Ano02s, Ano02t, Ano03-34, Ano03-35, Ano03-36, Ano03-37, Ano03-38, Ano03-39, Ano03-40, Ano03-41, Kro00a, Kro00b, MD00, Ano01i]. Professional [Aye01, Azi06, FFCM00, GS01, JHA^{+}05, M^{+}00, PL03, WMC04, Gig00, RC04, SB06a, Ahm01, Ano02p, Che02b, Fox01b, Fox01d]. professor [GEVZ09b]. Profile [BHM^{+}07, BG04a, DTD04, KNG02, NIKN06, RTVH01, Dob01b, KWK05, San04b]. Profile-based [BHM^{+}07, NIKN06]. Profiler [SH04a, VL00, Way03]. profiles [LOW09]. Profiling [Ano01h, Ano03-40, Dmi04, Kro00b, PWBK07, SKS01a, Bin06, BSMV09, KJBH^{+}00, LPH02, MCD09, SK08, XAM^{+}09, ZSCC06]. Proglets [Edm09]. Program [ACM01a, BM03, BAJ01, CCW02, CHHC04, Cle01a, Cle01b, EFN^{+}01, GNYZ05, Han05b, HKK^{+}01, HS02a, HZC^{+}04, HJ00, HB08, Jac01c, JKW03, JP04, JRH05, KK03b, KKJY04, Kro00b, LL01b, LG00b, LM04, MD00, MSG01, MCLC02, MMBAS04, NLC03, OS02, Rob01c, RCdBL02, Uni02, Zam03a, Ano02g, Ano03-45, Ano05k, BBS04, Cal02, CT05, DDS02, DD02b, DD03, DD07, DNS05, DS04, EFN^{+}02, GHBG^{+}03a, GHBG^{+}03b, Gri02b, HCMM00, HPH03, HZS08, JPSN09, LO00a, LL00, LL03, LL01c, LH08b, Li02, MBED06, MCLDP01, MGM^{+}06, NE04, PC03, RRP02, RSD01, SLC03a, SMTZ09, SRW^{+}00, SK08, Smi01a, ST09, WN08]. Programm [Ste08b]. Programmable [JBMP03, JKKL04, KAN^{+}03, MD00]. programmed [Emu04]. Programmer [BBL03, HS00a, Mak03, RS05, SO00, Tre02b, Way03, Wil00b, Wil00c, Wil00d, Wil01b, Wil03a, Wil03b, Wil03d, Wil03c, Bai03, Che00, ET05, II04b, Jor02, MJ01, MR00b, New00, San04a, Woo01]. programmering [HJL00]. Programmers [Bro04, Bru03, Cal03, Gla06, Spi03a, Spi03b, Wei04, BBS04, BB00b, BS00a, BMS02, CD01a, Dur02, Gol04a, HB09, MFRW07, Mul00, SCL^{+}08, Sik03, Soo09, Spe02, MSU08]. Programming [ABV00, Ano00d, Ano00k, Ano01m, Ano02h, Ano03-39, Ano04-30, AT01, AGH00, AGH05b, Atk00, BIB05, BBC07, Bag02, Bal03a, BKT03, Bal02, Bar03a, Bar05, Bar00b, Bee00, BO05, BM01, Blo01, Bul00, BKO00, Cal04, CF03, CFLL03b, Cav02b, Cav04, CG02, CR05, CWY01, CT00, CMR05, Cou01, DH04a, DT02, Dar01b, DL02, Dib02, Dmi02, Dwe00a, Esp06, Fab02, FL02, Fig00, Fle00, FMMd03, GD00, GK03, Gil00c, GLC01, Hal09, Ham02, HR00, HKK^{+}01, HdJ01, Hei03a, HMRM03, HBH01, ISO08, JT04, Kal01, KGMO04, Kic03, Kin00, Kum04, KWK03, LBD^{+}03, LB00, Lia00a, Lia00b, Lia01, LAB^{+}00, MZ04, MDS04, Mas00, MSM05, NRV00, N^{+}00, OK04, OL01, Par04a, PSDF01, P^{+}98, Pre00a, Qui03, RWL07, RTVH01, RVZ04, Ros02b, SU03, SC02a, San02a, SJG03, Sav01]. Programming [Sch00b, Sco03, Ses00, Ses08, SS07, Set03, SFP03, Sla00, SSS05, SC05, Ste01, Ste00, Sub08, Swa01a, Tam00, Top00, WB00, Wei01, XYC05, YHGL01, Zea00b, vNMKB05, ADT03, ACZ05, Ana01, AF02, Ano01a, Ano03h, Ano03-50, Ano04e, Ano04g, Ano04-38, Ano05j, Ano05q, AW00, AJ01a, AJ01b, ABI^{+}07, ABG^{+}08, ABI^{+}09, BC07, Bai00, Bak00, Bar01d, BAF03, Bee04b, BZ05, Ber02b, BD04, BVPE06, BH04c, BMS02, BVD01, Bud00, BC03, BW01b, BW04, Cal01, CMC^{+}06, CM05c, CMS06, CC02, Chr00, Dav05, Dek06, DMKN02, DH00, Edm09, Ell00, ET02, Est01, FJ05a, Fei07, For04a, Gel00, Gou06, GJ09, GST05, GDB02, Hag00b, HB01, HAL02c, Har00c, Har04, Har00d, HF06, Hel07b, HL02a, Hig03, Hol04b, HJ01, Hor02b, HC01b, Hyd00, JPS^{+}08, JF05]. programming [Kag09, KOB01, KH01, Knu01a, KS07, KKT04, Kum05, Kur04, LO00b, Lar01, Las02, LP01a, LDB^{+}03, Lea00a, Lea02, LCFL04, LZ04, Lia02, Lia03a, LCFkL05, LLCF08, Liu08, LCC09, MVV^{+}01, MS05, Mau02, MGB^{+}09, MSK09, MMG^{+}00a, Mor02, NP03, NH02, Nis03, NP07, Och09e, OJ09, PJ05, Pir02, PM00, Pri01, Ran03, Ree00, RR02, Ril02, RPP07, Sah02a, Sah02b, SH03, San03, SD03a, Sei09, SY04, SCS01, ST09, SM03b, SAB^{+}06, SPGV07, Sta00, Swe06, TP08, TB00b, Utt06, WACBL03, Wam02, Wan03b, Wel04, WD00, Woo02, Wu01, Yan02, ZJ03, ZK05, vNMW^{+}05, vTNC08, Ano01h, Ano02h, Gil01, Omm01, Ano04e]. Programs [AR03b, AH04b, AGS01, Bec01c, Dd01b, BM04, BAJ01, CA04, CC04, CX01a, CX01b, CO03b, CQX^{+}09, CiLH01, Chr01, CD01b, CHK^{+}04, CCF^{+}02, DRV02, DKTE04, DEJ^{+}01, DEL^{+}02, EvG02, ESS02, ELM^{+}04, FJ01, FCMR04, GR07, GV02a, GCH00, GMT02, HR04a, KM04b, Kie01, KKL^{+}04, KVK^{+}04, KY03a, KY03b, KKJY04, LDE^{+}02, LCS04, LFP04, Lin01, LFH03, Lut03a, Meh02, MMK04, PL01b, PP02b, PP02a, PDV01, PV04, DJM^{+}02, PH02, PCC01, Qui03, RM04, RH04, RWZ09, RST^{+}04, RCR06, Rot05, SMCS04, SR05, SK00, SCLV04, SL01, TP01, WG01, WHBS01, WP00b, XC01, YK03, ZW08, ZXNH02, Zha05, AH03, Ano02e, Ano03h, Ano03-44, BP01c, BR01b, BA09, BK05b, CCC^{+}06, CY02, CO03a, CTF03, CDF05, Coh04, CMS07, CF04b, Cor00, D^{+}00, DH08, Dar07, Dil00, Dob01b, EFG^{+}03]. programs [EGD03, EL01, Eng04, ER09, FCHE02, FC00, GHS05, GV02b, GV04, HP00, Hel07b, Hir07, Jac04a, JPS^{+}08, JJ02a, KPH^{+}09, KCSL00, Kes04, KH00, KLS00, LTOT07, LFM09, LPH06, ML09, MMU04, MF07b, MF09, MKM^{+}06, MSV05, MC06, NK06, NR06, Nau02, NAR08, PH00a, PWN04, RH07, RM00, SBAD01, San00, Sen08, SC02b, Sto02b, TETPQ08, TS09, TZ01, Uni03, VMWD05, Wan03c, WF04, Wor02, XSaJ08a, Yah01, YLW08, Zar02, ZKR09, dH05]. Progress [CK05, Wit00, Yan03, KPN02, Mls04, RVZ04, Ano00m]. Progressive [Djo09, TGO00]. Project [Ano05p, Bar01b, BALV03, CY03, Kro00a, Lin03a, MLJH04, Ano05h, Cla04, Eub05, Joh00b, Kim02, Lab09, LM06, MMG^{+}01b, MWM01, NM02, OOOiM05, PB06, Sha02, Wol01b, Ple02]. Projectors [MD00]. Projects [PH04, Ses00, Ano03h, Ano05c, Djo08, WN05]. Prolog [ACZ05, DOR05, Sch04d, TT01, ZT02]. Prolog-to-Java [TT01]. promotion [LCHY03]. Proof [ÁMdB00, AddS03a, AddS03b, ÁdBdRS08, FC00, FC01, GKW04, ÁdBdRS05, Coh04, ZKR09]. Proof-Outlines [ÁMdB00]. proofing [CHL07]. Propagate [LPSY04]. Properties [ACL03, BD02, BR01d, Fre05, HD01, Mos05b, RW03b, TC03, IS03, MF07a, Yah01]. proposal [DV01, Jen01]. Proposed [BC00, Bar01b, CG01]. Proprietary [BCS07, Egy01]. pros [Ano04-38]. Prospects [SvR01]. protect [San04a]. protected [Ano00f]. Protecting [ML00]. Protection [SLB^{+}02, HvE02, RR01]. protein [Ano01d, CWWS03, FL04, GV05, GP05]. protein-protein [Ano01d]. Proteus [CGG02]. Protocol [Cim02, CMR05, CHK00, GS00b, LC05, Gun01, HOP04]. Protocols [GSC^{+}00, BRBY00]. Prototype [AG03a, Ang06, BCE^{+}01, RP06, vdBDS00]. prototyping [LSK^{+}02]. PROVA [KS04]. provenance [GMM09]. provenly [AAD^{+}07]. Prover [Ber01c, DNS05]. provide [Kic04, GHBG^{+}03b]. Provider [LDM04]. Providers [KP01]. provides [Way03]. Providing [FJ05b, KdJNNV09, PH00a, PSM01a, PSM03, HCB04a]. Proving [GN01b, Moo03a]. ProWorks [Ano00j]. Proxies [Bar03c, PSH04, RE01, Eug06, Ren02]. Proxy [BCH02, Eth01, NW02b, Ano03k, Ros00]. ProxySource [Ano01l]. Pruning [RH04, BM09]. PSEs [SRW^{+}00]. PTIDES [ZABL09]. Pty [Ano00i, Ano00j]. Public [Cow01, Gal02]. Publications [Bee00]. Publish [Hou00, LPSY04, RG00, Rou02b, Tho03]. Publish-Propagate [LPSY04]. Publish/Subscribe [Rou02b]. Publishing [Ano00k, Pew00, Sha04]. Pure [GW02, Goo00, Lit00, Ano03n, Ano03-31, CW03b, VDPC03]. pure-Java [VDPC03]. Purity [SR05]. Purpose [WP00b]. Purse [CH02]. Push [Ano02l, Coc02]. Put [Way05]. puts [Ano03-44]. Putting [CSFS00, Gun01]. puzzlers [BG05]. Puzzles [Ros02b]. PVS [Jac03]. Pylons [Gar09]. Python [SML06, SKS08, Ang00a, Ang00b, Ano00n, Ano01l, Gar09, GL08, HF06, Hig03, MSR03, Pre03, Rad06, Rem01, SM04b, Stu07, Wil05].

Q [Ano00h, Ano03-30]. Q&A

[Bru02, Cal00b, Coh02, Cox01a, EKM00, Fox00e, Gol01, Gso00, Hag02, HL00, Jac01a, Jen00a, Jen00b, Jen02b, Jol01, Kie01, Kie02, Lai01, McK01, Mos00, PH00b, Rao02, Rei00a, Sea02, Smi01b, Str01, Tra00a, Vil00, Win01, Wra01, Yua02, dD01a]. Q-Link [Ano03-30]. QA [Coh04]. QL [ISO08]. QoS [PSM01a, PSM01b, Zea00a]. QoS-aware [Zea00a]. qualifier [GF07]. Qualitative [RJGH06, MLM^{+}08]. Quality [Ano01k, CLN07, Pau03, BWLP01, PSM03, PC08]. Quantification [WG01]. Quantifying [FFB^{+}00]. Quantitative [Lut02, RJGH06]. Quantum [Pap05, SPS^{+}02, HS01]. quasi [SBMG00]. quasi-static [SBMG00]. Queens [Rol08b]. queries [SPBE09, TGO00, WGSD07]. Query [WPN08, AYWM08, PFS05, WIC08, dMSAV08, vdBDS00]. Querying [ACD^{+}04, Ano02k]. Quest [Ano03-35]. Questioning [MLG02a]. Questions [Lea00b, SLB^{+}02, SPS^{+}02, Bur02, HSB09]. queues [SLS09]. queuing [KPPÉR06, XOWM06]. Quick [Vor01, Ano00b, FFC02, Fla02a, Fla05b, OW00, RP06, Top02b]. quickly [PPJ03]. Quicksilver [SBMG00]. QuickTime [Ada05]. quietly [Ano03o]. quirky [MLM^{+}08]. Quiz [GM02]. Quiz/Exam [GM02]. QVM [AVY08].

r [KM01, Guh07, Mur05, Nar05, Sch00b, Urb09, Hec07, Laz07, dL05, Hol06]. R-based [HLT09]. R/3 [Sch00b]. R134a [TC03]. R3 [APA04]. Race [AS03, CD01c, CD01b, Sen08, Yan02, AFF06, BR01b, CSFS00, EQT07, FF00, FF09, NAW06, NA07]. Race-Free [AS03, BR01b]. Raced [LOW09]. races [BST00, PRB07]. RAD [Ano02o]. radical [Reg00]. radio [Ano05a]. radio-based [Ano05a]. radiolysis [PFJ05]. RAGE [PSW07]. RAID [Ano03-36]. Rails [HG07a]. RakPak [Ano00h]. Ralph [Ano00d]. RAM [Gar00]. Rambutan [Sah02a, Sah02b]. Random [PSW07, Sen08, Bee04a]. randomized [JPSN09]. Randy [Cha03]. range [NIKN06]. ranked [SPBE09]. Rapid [Ano01l, Ano01m, Lia00c, NSI03, TCF^{+}03, Gar09, KdJNNV09]. RapidStream [Kro00b]. rational [CBGM03, Ano00n, Ano02q, Ano02r]. rationale [CMLC06]. Rave [Ano00j]. Ravenscar [CW04a, Dob01b, KWK05]. Ray [Uni02, Ano02g]. Raytheon [Ano01o]. RCX [Wol01b]. RDF [Ebe02]. Reachability [LCS04]. Reaching [Gar00]. reacted [PPJ03]. Reactive [Cou01, Sto02a]. Read [Bog00, Ano00f]. Read-Only [Bog00]. Ready [Ano04b, Cha05a, JM00, RH04, DW07, Zhu04]. ready-made [DW07]. Real [APA04, Ano01i, Ano02m, Ano03s, Ano03-52, BCR03a, BR01a, BN03, BG04a, BD01c, BD01b, Bro03a, Bro03b, BW03a, BW03b, Bro04, Bro05, BW01b, BW03c, BW04, CW03a, Cav02a, CKC^{+}02, CS02, CS03, CC03, DC03b, Dib02, FBR^{+}03, FCHE02, GKM03, GKMZ04, GKW04, Gle02, Gos00a, Har00a, HIBP04, Hig04, HWB04, HCB04b, JKJ05, KM08, KNY03, KM02, KK03a, KBP^{+}03, Kro00b, LD03, MB03, McL01b, MLJH04, NK03, PV03a, PSM01a, PSM01b, PUF^{+}04, Pot04, San02a, San03, She03, SLC03b, SH06, Sun01, TGB^{+}04, TSL^{+}04, Uma02, Wan04, WP03, Wel03, Won05, ABC^{+}07, ABI^{+}07, ABI^{+}09, Bol00, BSBR03, BHR02, BH02c, CY01b, DV01, HT06, Ive03a, Jen01, JPSN09, KPH^{+}09, KWK05, PSM03, PHV07, San04a, SAB^{+}06, Wan02a, WLW^{+}03, Wel04, ZABL09, Ano03s, Dob01a, KSK04b, PL03, She03]. Real-Time [APA04, Ano01i, Ano02m, Ano03s, Ano03-52, BR01a, BN03, BG04a, BD01c, Bro03a, Bro03b, BW03a, BW03b, Bro04, Bro05, BW03c, CW03a, Cav02a, CKC^{+}02, CS02, CS03, DC03b, Dib02, FBR^{+}03, GKM03, GKMZ04, Gle02, Har00a, HIBP04, Hig04, HWB04, KNY03, KM02, KK03a, Kro00b, LD03, MB03, MLJH04, NK03, PV03a, PSM01b, PUF^{+}04, Pot04, SLC03b, Sun01, TGB^{+}04, TSL^{+}04, Uma02, Wan04, WP03, Wel03, Won05, BCR03a, BD01b, BW01b, BW04, CC03, FCHE02, JKJ05, KM08, KBP^{+}03, PSM01a, San02a, San03, She03, ABC^{+}07, ABI^{+}07, ABI^{+}09, Bol00, BSBR03, BHR02, BH02c, DV01, HT06, Ive03a, Jen01, KPH^{+}09, KWK05, PSM03, PHV07, San04a, SAB^{+}06, Wan02a, WLW^{+}03, Wel04, ZABL09, Ano03s, Dob01a, KSK04b, She03]. Real-World [McL01b]. realisieren [Sig04]. realities [BCM04]. Reality [RPJ04, HL02b, Ano04l]. Realization [Che03c, DYH05, LZZ03, LW03, SY04, XZ03, CW03b]. Realizations [RWH01]. really [Fit09]. RealNetworks [Ano03-37]. reals [Boe05]. Realtime [Ano04l, Bac07, Ano02f]. Reasoning [ACN02, BDHdS01, HP04, GSWZ08, Jac04a]. rebiasing [RD06]. Recipes [RS05, FG05]. recoded [Ano03-45]. Recognition [MD00, KKM^{+}06]. Recompilation [KNG02, THL03]. reconciling [Tan07]. Reconfigurable [MH00a, LC05]. Reconfiguration [RAC^{+}02]. Reconsidered [OKK04]. Reconstruction [SGV04, dCG^{+}02]. Record [Ano03-39, BHP^{+}01, Chr01, GCRD04, HPH03]. Record-Performing [Ano03-39]. Record/Replay [Chr01, GCRD04]. recording [BW04]. Records [HTY^{+}03]. Recovery [DHMT00, KdJNNV09]. Recurrence [CM05a]. recursion [VIPCUF08]. Recursive [FR00, XC01]. Red [Ano00d, Bar00a, Ano03x, Way03]. Redesigning [MDS04]. reduce [BALP01, BALP06, Cor00, LLdA08]. Reduced [XX05, VED07]. Reduced-Instruction-Set-Computer [XX05]. Reducing [LYK^{+}00, CSK^{+}02]. Reduction [CKV^{+}02, Vil08, KOO08, RSS^{+}04, TABP07]. redundant [Tro04a, Tro04b]. redux [Dor07]. Reed [Gla06]. Reentrant [ÁMdBdRS02]. Refactoring [Wic03, HKI08, OJ09, TT08, TTS^{+}08]. Reference [Ano01j, Ano02p, Ano03-37, CC03, Fla02b, Goo02a, Lut03c, SO00, WGW04, Woo05, Bal03b, Ber01b, CK03a, DS00b, Dur02, FFC02, Fla02a, Fla05b, GK07, Hap02, II04b, JMP09, LS00, LP01b, LP06, LPH02, MJ01, MDJ05, OW00, PS01, RP06, Sch01, Stu07, Top02b, TE05, Woo01, YTY00, Ano00b]. reference-counting [LP06]. reference-counting-based [JMP09]. Reference-Set [WGW04, Woo05]. References [Ams00, SR06, CR06, HT06]. Refinement [SB06b, WHKS01, KPPÉR06]. Refinement-based [SB06b]. Reflecting [RE01]. Reflection [BK01b, Chi00, DFT03, Fei04, FF05, PL01b, Par00, TT01, WS01c, HS08, Mor02]. Reflections [Ben00b, Ben00c, CV01, Ben00a]. Reflective [Dwe00b, OSM^{+}00, TBSN01, CV03, FDR04, VN00]. Reflex [TBSN01]. refreshing [Ano04a]. Refrigerant [TC03]. Region [QH03, BSBR03, SYN03, SYN06, SD04]. Region-based [QH03, BSBR03, SYN03, SYN06]. Regions [DC03b]. Register [KMEA04, YLL^{+}07, LCHY03]. registers [JK00, SCEG08]. Registries [Tre02a]. Regression [HJL^{+}01, CO06, OSH04]. Regrowing [OJ09]. Regular [Hab04, Stu07, AOMC07, Kah06a, Mor02, SM04b]. Reguläre [SKS08]. regulatory [SD04]. Rehashable [LBJ02]. Reification [BL03, VB01a, CV08]. Rekeying [PR03]. relance [Ano03-47]. Related [CL03b, ME00a, BBS04, RD06]. relational [LH04]. Relations [DJ00, LH08b, DJ02]. Relationship [CMS06, DL02]. Relationships [GCEO05, CHUB08]. Relaxed [Dic01, MRC03]. Relaxed-Locks [Dic01]. Release [Ano05i, Bar01b, Ano03-29, Ano05n]. Released [Ano00n, Bar01a, Bar01c]. Releases [Ano00n, Ano01i, Ano01k, Ano01n, Ano01o, Ano02n, Ano02o, Ano03-37, Ano03-39, Ano03-40, Ano03-41, Kro00b, Ano03-34, Ano03-35, Ano03-36, Ano04n, Ano04u]. relevance [Gao00]. reliability [WN08]. Reliable [BL02a, IEE03b, SBA01, Ano02f, NRS^{+}07, Oes01]. Relief [Bar01a]. Relocation [ZX05]. remain [Ano05e]. remains [Ano03f]. ReMLab [FSBP03]. remodularization [CD08]. Remote [Ano01o, Ano03-42, AV05, CE01, CCSA02, FSBP03, IEE03a, KK03a, LH03a, NMMS01, Rob00b, SDPM04, SAFG03, Tddd03, WXW^{+}05, ZYC03, Ano02k, GCARPC^{+}01, IH01, JS01, LY03, MR00a, PM01a, Rob03, WSVX03]. remotely [KL07]. removal [Ruf00, SAB08]. Removing [PL01b, Tro04a, Tro04b]. renaming [CDF05, SEdM08]. rendering [WW09]. Renesas [Whi03a]. reorganizing [Ano05m]. repair [EKVM07, vdSPP05]. Replace [Reg02a]. replacement [GSHO06, NAR08]. replacing [Utt06]. Replay [Chr01, OOK^{+}06, SBB05, SCFP00, GCRD04, GEB08]. replicated [IH01]. Replication [KMSL03, LPSY04]. Report [Ano01b, Ano02b, Cha00a, DV01, LS04b, Nat00, RBC^{+}05, Fre07, KPN02, LHS04b, RBC^{+}06, SMS^{+}04]. Reporting [Ano02n, BNK^{+}07]. reports [GCF^{+}01]. Repositioning [TYS04]. repository [Fal00a, Fal00b, SFM^{+}07]. Representation [BJvdB02, RCdBL02, SPB01, WGW04, Woo05, ADR09, MGM^{+}06]. representations [Sam04]. represented [PB06]. Representing [Han05a, RM07b]. Request [BFS^{+}04]. Requirements [GSC^{+}00, KSK04a, KK05, LSK^{+}02, LFH03]. requiring [Ano02f]. ReRAGs [NIEH04]. Research [Ano00a, Ano01b, Ano01h, Ano01g, Ano02b, Ano02q, AJ01b, Che03a, CW03b, DLL03, Fel04, GH01, Gar00, HL04, HD03b, KLL03, SPB01, SSL02, TCC01, USE01c, USE01b, USE02, ZL05, Kim02, XP04]. Researchers [Coc02, Pau01, Pau03, Ham02]. Reservation [EGLZ02, KKO02, LS03, OKK04]. Resolution [RAC^{+}04, SHR^{+}00]. resonance [VP05, dGNv04]. Resource [Ano02r, Ano02u, BHL00, BH05b, Goo02a, HBD04, Jac01a, JCKS04, RP03b, Sur01, TS01, VB01a, BNV08, BHV01, CHS^{+}05, RA07, VVG^{+}05, ZK04a]. resource-constrained [BNV08, RA07, ZK04a]. Resources [KS01b, Rob04b, Ano00f, Ano04g, New01, PSZ^{+}07, Pan09]. respectability [Van04]. restore [Van04]. Restricted [RCdBL02, ABG^{+}08]. Restructuring [YK03]. result [SPBE09]. Results [HL04]. ResultSet [Ano03-42]. Resurrecting [Rob07b]. Rethinking [Ree01]. Retrieval [Gal01]. return [Ano04u, Siv02]. reusability [Sma07]. reusable [DSCU01]. Reuse [BS04, RE01, AK09, Fle01, Gib09, WM00a, YLW08]. Rev [Ano05o]. Revelation [Dmi04]. Reverse [BLL06, Coo02, Kal04, Kes04, SKM01]. Review [Ano00b, Ano00c, Ano01a, Ano03b, Ano04e, Ano08, Azi06, Bal03c, Bar03a, BALV03, Bro02a, Cal00a, Cha05a, Cha03, Che05, Cow01, DHRH05, Dud06, Fox01d, Gil00c, Gla06, Hec07, Hol06, Kuc06, Laz07, Mar05, Mas01, Mil08, Mor03b, Omm01, Pap05, Pap00, Pet06, See04, dL05, Ano02h, Che02b, Feu02, Sur04a, Zen02]. Reviewer [Ano03-41]. Reviews [Ano00d, Ano03-41, GS00b]. Revised [GAR04, GRR05, Lut03c, AJ01a, GAR03]. Revises [Ano01o]. Revisited [vON02a, vON02b, MDJ05]. Revisiting [SMBZ07]. Revocation [WJH06]. Rewriting [RW03b, WS01c]. Rexx [Pre03].

Rhody [Fox01b]. RIA [Ano00j, WGC09]. ribosomal [JCP^{+}05]. Rich [CCB09, Yua04, HG08, JF06, Wea07]. Richard [Gla06]. Rick [Fox01b]. Ridge [Ano02i]. RidgeRun [Ano01m]. rifarensu [SM04b]. right [KT01a]. Rights [KPK02]. Rigorous [Fig00, LAB^{+}00, GBE07, GEB08]. RIM [Ano02m]. Ring [WBL01]. RISC [Whi03a]. Risks [BR06a, Cha03, Mer04]. RM1U [Ano00j]. RM1U-AXe [Ano00j]. RM2U [Ano00j]. RM2U-AXi-C [Ano00j]. RMI [AY05, AY07, AG03a, AG05, CW04b, CCC^{+}04, CCK^{+}08, ET01, ET07, EK01, GSC^{+}00, Gro02b, Gro02c, JKH^{+}04, KDH^{+}06, MVV^{+}01, Mar02, PHN00, SJ01, Sha01, SR06, WS01a, WCCL05, YK03]. RMI-Based [SR06]. RNA [JCP^{+}05]. road [LDB^{+}03]. Robert [Kuc06]. Roberto [Mas01]. robocode [Liu08]. Robot [Ano04-34, CCSA02, Bec01a, CW03b, XM06]. robots [EL04, Eng00, GCF^{+}01, JCOP07, LDB^{+}03, Wol01b]. Robust [CM01, GR07, Ste05, WC00a, BFN^{+}09, Gou06, RM00]. Robustness [FRMW04, FMRW05, CS04]. Role [LAB^{+}00, CTLW03, NC04a, Sha01]. role-based [NC04a]. Roles [SE04, CFL05b, CFL05a, ST04]. Rollover [Lea00b]. ROM [Hal01a]. Rose [Ano03-41]. roster [Sur04a]. Round [Dra00]. Roundup [Vie03]. Router [Ano01j, HHM04]. Routines [ISO08, Pon03, WP04, LS04a]. Routing [Lut02, HHM04]. RPC [All03, Cer02]. RPM [Men00]. RSA [Ano02s]. RT [Ano00h, Ano03-43, Dob01a]. RT-Java [Dob01a]. RTAI [Ano00j]. RTEL [Ano00i]. RTL [WHW01]. RTS [Wil06]. RTSJ [Ano03-38, TSL^{+}04, Wel03]. RTSJ-Compliant [Ano03-38]. Ruby [SKS08, Stu07]. Ruined [Ano00j]. Rule [CMR05, Esp06, Hig04, KS04]. Rule-Based [KS04, CMR05, Esp06]. RuleML [Ebe02]. rules [Ano03z, Dun02, Fle00]. Run [Ano03-44, CA04, GNYZ05, KKL^{+}04, KVK^{+}04, LH05, RW03b, VHBB03, Bre02, CC01, Gad03, Hor00c]. Run-Time [CA04, GNYZ05, KVK^{+}04, RW03b, KKL^{+}04, LH05, VHBB03, CC01, Hor00c]. Running [BH02a, HKHK03, Cal02, NAR08]. runs [Ano04-32]. Runtime [ATBC^{+}03, Ais03, ABH^{+}00, BH05b, CKM04, CEG^{+}03, CD03, FSS06, HR04b, KF05, LLCF08, MPG^{+}00, Shi03a, TP01, TOG^{+}05, VHBB01, AVY08, AK09, BH05a, BLW09, Bod04, CFL05b, CFL05a, CR07, EQT07, ACM03a, LLdA08, MKKC08, RVJ^{+}01, Ren02, SS08, WK08d, XAM^{+}09, dH05, CDH07]. Runtimes [Han05b, GK05, WK09]. rush [McL06a]. RV01 [HR04b].

s [Ano02o, KSC^{+}00, Ste00, YWZ03]. S4 [GMM00]. SA2 [Bro07]. SABER [RSS^{+}04]. SableSpMT [PV06]. SableVM [GH01]. Safe [AC06, LBR00, MPG^{+}00, Mos05a, Vel01, WJH05, WHBS01, AFF06, BSBR03, DGGD08, Fek08, HS08, Oiw09, SAB^{+}06, WK08d, Win02]. Safety [Hag02, San02a, Bro07, CG01, FF08, HM01a, MSG01, San03, San04a, Yah01, Yan02]. safety-critical [Bro07, San04a]. SAFKASI [WAF00]. Sale [Ols01]. Salesman [Bar01c, TCM^{+}00]. SALT [Ano03-35]. SALT-based [Ano03-35]. SAML [JSSM04]. sampling [Bin06, BGH^{+}07]. SAMRAI [WHKS01]. Sams [AK00, CL03a, WMM04]. San [USE00c, USE00a, USE01a, USE02, CHL^{+}00, Joh00b]. Sandia [Bar00a]. Santa [ACM00a, ACM00b]. SAP [AK01, Ano04-31, Sch00b]. Sapphire [HM01b]. SAS [Ano00i, Ano08, BI07, Pra08, Ano08]. SAT [KM04b]. Satin [vNKB01, vNMKB05]. Satisfaction [SS07]. SavaJe [Ano03n]. saving [D^{+}00]. SAX [Har03]. SAX2 [TEM^{+}01, Hei01]. Says [Bar01a, Ano03o, Ano04-27]. SC2000 [ACM00c]. SC2001 [ACM01c]. SC2002 [IEE02a]. SC2003 [ACM03b]. Scala [Sub08]. Scalability [AFT^{+}00, Bul00, BG03, Coh04]. Scalable [CM01, Det01, KLL03, MJ06, PTP07, SD01a, SLS09, Tor01, WC00a, Bar02a, Cal00a, DAK00, GW01, IV07, LLCF08, NQM06]. Scale [GP01, KT01b, McG04, CHP^{+}08, CHL^{+}00, KMSB08, NZM03, SCBH09, WMRT^{+}05, ZYZ06]. Scaling [Joh03, JDJ^{+}06, LH03b, OSH04]. scannerless [KdJNNV09]. Scanning [VMMF00]. Scans [Ano03-40]. Scene [MD00, Wal02b, PPJ03]. Schaum [HBH01, Hub01]. Scheduled [KNY03]. Scheduler [Ano02q, RB04, XSaJ08a]. schedulers [HL03a]. Scheduling [AHKR01, FBR^{+}03, KMEA04, Lin03a, NP01, RWC^{+}03, VT01, IKN03, KBP^{+}03, LTOT07, NC05, Rob04a]. Schema [Ebe02, Lut03a]. Schemas [Lut03a]. Scheme [FS03b, LPSY04, Ano03-44, IV06, SS02]. Schemes [CFLL03b]. SchlumbergerSema [Ano02v]. School [Bar03a, BGP00]. Schwerpunkt [BL04]. Science [Bar01a, Bar01b, Coc02, DFL00, Fox03a, HMRM03, Lut03c, Rob04b, Sav01, SG00, SM07, Thi02, AWS^{+}09, BR02, BS01, CFGL05, CKMP09, CF04b, DW07, Fro07, Gol04b, Hel07a, KMR02, Rad06, Ras00, Rio02, Rob04c, RVZ04, SSC00, Ano02q]. sciences [PB06, Ran03, Woo02]. Scientific [Art00, BJK07, BSPF01, GK03, GSC^{+}00, GAR03, KT01b, LBQ00, Lut03c, NZ01, PTML09, PH02, SvR01, VP05, BBBD01, BB00b, BSB^{+}03, Esq04, FCHE02, LP05, PT09a, SML06, SHHS04, vRKS01, vRKS03, GAR04, GRR05]. Scientists [Cha00c, BB00a, Lau04, ML07]. SCM [Ano03-39]. scope [BDN05]. Scoped [BR01a, DC03b, GNYZ05, WSM06]. scoring [SPBE09]. Scotland [Tra00b]. Scratch [ML07, Sah01]. Script [Got06, Lai01, WGC09, Wea07]. scriptaculous [Ang06]. Scripting [Ano01n, Gös03, Kah06b, KS04, McC00g, PTML09, Pre03, Rem01, Spi05, Tra00a, BFN^{+}09, DM07, Han01, PT09a, Ric00, Wea07]. Scripts [BL03]. Scrutinized [GM03]. SDE [Ano02p, Way05]. SDK [Ano00h, CG01, Ano01h, Jon02]. SDL [KPKL03]. SE [Sun02]. Sealed [ZFA00]. Seamless [HR00]. Sean [Fox01b]. Search [AGH05a, BWW^{+}03, Cal00b, Lut03a, Pau03, STB08, SPBE09, BV05, Fit07, Fry03, NM02, Rob04c, WF04]. Searches [Pau01]. searching [Lee03]. Sebastopol [Ano00b, Ano00c]. sEc [SMK02]. Second [Ano00d, Ano00n]. secret [Gal02]. Secrets [Sim04b, TEM^{+}01]. section [KGH^{+}05]. Secure [Ang01, BL02a, Cha03, CLM^{+}07, DDF^{+}03, Feu02, LS03, MR00a, Mar02, Mos05a, PR03, SSM03, WVE^{+}00, WBL01, vD00, Ano00g, ABF03, BAF03, BDLM04, CLM^{+}09, II04a, PNKN04]. securities [Ano02w]. Security [Ais03, Ano00i, Ano01n, Ano01o, Ano02r, Ano05k, BD02, BR06a, BML01, CV01, CHV01, FVK01, GN01a, HOP04, HBD04, JSSM04, KSC^{+}00, KNN^{+}01, Kro00b, LKL^{+}03, Liu03, LRO02, Mos05b, PNKN04, RC01, Rot02, SPS^{+}02, USE00d, VMMF00, WFGK03, Wea00, WBL01, Yan03, AJ01a, AJ01b, BLW09, CV03, GS01, HS05, IK04, JPC00, Oak01, PE06, WAF00, YCIS07, Ano02s, Feu02]. Security-Aware [CHV01]. seeks [Ano05m]. seems [DA04]. Seetoft [Bal03c]. Segmentation [HKL09]. Seiki [SM04b]. Seismic [SGV04]. Select [Joh00a]. Selected [HR04b, GRR05]. Selecting [GKM01]. selection [HJL^{+}01, LOW09, SVY09, SMTZ09]. Selective [CCF^{+}02, DGMY06]. Self [Ano03a, BH04b, DDF^{+}03, FOS^{+}04, SI09, Ano04a, Emu04, GK05, Woo04]. Self-accounting [BH04b]. Self-Adaptive [FOS^{+}04]. Self-certified [DDF^{+}03]. Self-Contained [Ano03a]. self-describing [Woo04]. self-efficacy [Emu04]. self-healing [GK05]. sell [Ano03n].

Semantic [KS04, TMF05, SSP07]. semanticist [SNO^{+}07]. Semantics [BDJ^{+}01a, ÉJD01, HEJ09, JP00, JR05, MP01a, TSDNP02, Zam03b, Ber00b, BFGS05, JP03, MF07b, MF09, MBS^{+}08, Moo06, Siv04, ZK09]. Semantics-aware [HEJ09]. semester [LM06]. semesters [OJJ00]. Semi [Fel03, AC01]. Semi-automatic [Fel03, AC01]. Semiconductor [Ano02p]. Seminar [DK02, Hal01a, KR00]. sense [Way03]. Sensing [IEE03a, SAFG03, WXW^{+}05]. Sensitive [CC04, LH08a, SB06b]. sensitivity [LPH06, MRR02, MRR05]. sensor [TBM09, WSVX03]. Separate [ALZ02]. Separating [GB01]. Separation [PB08, WBGM05]. September [AJ01a, SM07, SBH^{+}04]. September19 [AJ01b]. September19-21 [AJ01b]. Sequence [Bar01b, BLL06, NMH^{+}02, OS02, AWE04, CWS04]. Sequences [GH03, JCP^{+}05]. Sequential [CO03b, Gam03]. serial [ZK09, Ano03-36]. Serialization [BP01d, HJR^{+}03, WTV03, WTV05, BHK^{+}04, BP03b, CFKL00, PHN00]. serialized [Woo04]. Series [Azi06, BMS02]. serve [OBr05]. Server [Ang00a, Ang00b, Ano00j, Ano00k, Ano00n, Ano01i, Ano01l, Ano02h, Ano03-37, Ano03-38, Ano05i, Bar01c, Ben00b, Bul00, CCB^{+}01, DUK02, Eth01, Goo00, GW00, HECR00, JCKS04, Kan02, LR04, Ler01d, Liu04, N^{+}00, Nyb02, Omm01, PVC01, RS00b, Sah01, Wut00, AHN02, Ano02a, BDF^{+}00, BHJR05, Cal00a, Cal01, CG02, DBC^{+}00, DAK00, FMRW05, GM05b, GW01, HJL00, Hef07, IH01, KJBH^{+}00, KS01a, LHFL07, LLS^{+}08, Sha02, Tre03, XSaJ08b, Ano02h, Ano03-37, Bur07, SPBE09]. Server-Based [N^{+}00, Ano02h]. Server-Side [Ano02h, Bul00, Ler01d, Cal00a, Cal01, Tre03]. Servers [Ano02m, Ano03-39, GKM01, Joh03, Mar02, She01b, TEM^{+}01, Ano05j, BBYG^{+}05, JDJ^{+}06, MHZG06, Tro04a, Tro04b, Vau03a]. Service [AGH05a, ABM^{+}03, Bar05, CW04b, HMD04, Hoh03, Hua03, KP01, LKL^{+}03, LDM04, RAC^{+}04, SAWW01, TA04, W^{+}04, WXW^{+}05, Aar06, Ano04-27, CG02, CMS03b, FT00, Hap02, LCZ04, MHC01, MF03, PSM03, RA07, Swa07, ASS03, Ano02f, JO03, LS03, RMHC09]. Service-Oriented [Hua03, Swa07]. Serviceability [RB01]. Services [Ano00i, Ano01m, AM02, BCS02, Bru05c, Cer02, DJLT01, FRMW04, Hon05, Jen00a, JSSM04, Kan02, KR03, Lai03, LAT04, LHS04a, MTSM03, SSS02, SC05, Wal03a, Wal03b, Ano03w, Ano03-29, Ano04n, Ano04-39, CJ02, JKH^{+}04, MR09, PPJ03, SGW01, Sig04, Top03, Tro04a, Tro04b, Lut03b]. Servlet [Hin02, HC01b, Per04]. Servlets [Ben00b, Ben00c, Bro01, Cox01b, DiM04, EF02, GHH01, Hal00, Hal01a, Hal02a, Kie02, Rei00a, RS00b, BSB04, BSB08, Cal01, Har01a, Jor02, Wut00, DUK02]. SeSF [ES05a]. SeSFJava [ES05b]. Session [BH02c, GM05c, Rei00a, Bar01d, DV01, Hag00a, KR00, PT09b, Soo01, Dob01a]. Session-ID [GM05c]. Sessions [GM05c]. Sestoft [Ano03b]. Set [Ano00o, HD01, WGW04, Woo05, XX05, Ano04z, Eng00, Moo03b, Sco02, Yua04, vRKS03]. set-tops [Ano04z]. SETI [Bar01b]. Setting [Bet04, BHP^{+}01]. Setup [Ano03-38]. Seven [Pre00a, SLB^{+}02]. Seventh [LL08a]. Sfixem [AWE04, CWS04]. Sfixem-graphical [AWE04, CWS04]. SGDL [Ano01o]. SGI [Ano02r, Ano03-36, Ano03-38, Ano03-39]. Shackled [Sta04a]. Shan [Bar03a]. Shape [LAB^{+}00, BFN^{+}06, Cor00]. shapes [IEE03a]. Shared [BMR02, BHP^{+}01, CH08, Fox00d, GPS03, HS00b, SCLV04, TEM^{+}01, Che03c, ESS04, HW00, PV03b, WK08d]. Shared-Memory [SCLV04]. Shares [Ano05i]. Sharing [BHL00, CHS01, KS01b, PCC01, QM09b, TS01, LLdA08, ESGS00]. sharp [Hun03a]. Shell [VWS^{+}05]. shift [GEVZ09a]. Shimba [SKM01]. Ships [Ano01i, Ano01j, Ano01k, Ano01m, Ano01n, Ano02s, Ano03-40]. Shirts [Bar00a]. Shop [Ano00h, Bec00a, Bec00b]. Shopping [LL01a, SL06]. Short [CWH01, LS04b, CY01b, LHS04b, ZCR^{+}06]. Shortage [KSC^{+}00]. Should [Dar01b, Lai01, Lyk02]. showdown [SCEG08]. sich [Wol03b]. Sicherheitskritische [Ano05l]. Side [Ano02h, Bul00, vON02a, SR05, vON02b, Ano04u, Cal00a, Cal01, JS01, KL07, Ler01d, MRR02, SC01b, Tre03, Wea07]. side-by-side [SC01b]. side-effect [MRR02]. SIGACT [LL08a]. SIGART [LL08a]. SIGCSE [Bru04b, Bru05a, RRP02, Reg02b]. SIGCSE-members [Bru04b, Bru05a]. sight [CAF04]. SIGMETRICS [ACM00b, ACM01d]. SIGMOD [CNB00, LL08a]. SIGMOD-SIGACT-SIGART [LL08a]. Sign [JSSM04, Ano02j, KKN06]. Sign-On [JSSM04]. Signal [Ano02s, KC03, She03, BH05c, Sar03]. Signalling [BK08, KPKL03]. Signature [SA02]. Signs [Bar00a]. SIGPLAN [ACM01a]. SIGSOFT [ACM01a]. Silas [Ano02n]. Silent [Won03b]. Silicon [Ano02p, Ano03-46, Ano03-40]. Silk [Kil02, Kil03b]. SIMA [RLR00]. Similarity [BK01b, FL04]. Simple [CHV01, Cog04, KM01, Lan04, PR04, vNMKB05, KW01a, LH07, LRD09, Sci07, WKB02, Gun01]. SimpleDB [Sci07]. simpler [Ano05q]. Simplest [Sch03a]. Simplicity [BGP00, Lee03, Rob04c]. simplified [Uni03]. simplifies [Ano04x]. Simplify [Smi01b, Ano04j, DNS05]. Simplifying [Gun01]. Simulated [GKM03]. Simulating [FGLS04, Lyo02, Roj00, TB00a]. Simulation [Ano01n, Ano03-45, Ano04-34, AH04b, AAA^{+}04, CCW02, CWZ04, CCSA02, GKMZ04, JLV02, Kil02, Kil03b, LMV02, Lut02, McG04, NDS^{+}02, PP02c, RJFG03, VDPC01, WP04, WWMG06, YHL01, AYWM08, FW02, FCW01, Gar01, GM05b, LJN^{+}00, NZM03, OG05, PFJ05, PWC00, PSS01, VDPC03, Wen05, Lut03c, SO02]. Simulations [Esq04, FCHE02, HS01, Ibb02, KM08, PCC00, SHHS04, WMRT^{+}05, Pap05]. Simulator [HKHK03, KW02, NC04b, VHL01, CMP^{+}07, Rob02, Rob04a, Rob07a, SM01a, VS06, WW06]. SimulRad [PFJ05]. Sindhi [SSS05]. Single [CWZ04, Hig04, JV04, JSSM04, Lau03, MWL00, MBS^{+}08, WP04, And01, Ano03-36, GPF08]. single-chip [Ano03-36]. Single-System-Image [MWL00]. Single-Threaded [JV04]. SIP [GHH01]. Sites [Lut03b, Ano03f, Atk00, MMN09, SM03b]. situations [WN08]. Size [AR03b, KK04a]. Sized [JJ02b]. sizes [IEE03a]. Skeletons [ABG02, AG03b]. Sketching [Hit03, ABL07]. skills [Ano04o, CLP06, Ear03, Mls04]. Skin [Ano01o]. SL-A300 [YKS^{+}02]. Slate [AJB^{+}04]. Slaves [Lut00]. slaying [Lab09]. Slicer [JRH05]. Slicing [AH03, CX01a, CX01b, KKJY04, LFP04, MMK04, RH04, RH07, Li02, MKM^{+}06, NR06, SFB07, WR08]. Slim [MD00]. Slim-Line [MD00]. slope [JJ02a, Uni03]. smack [Mer04]. Small [Ano04-32, BAJ01, CCM05, JJ02b, Kro00a, SSB03, PK00]. Small-Sized [JJ02b]. Smalltalk [Bes01, EK03, Fei04, Lut01]. Smalltalk-like [Fei04]. Smart [Ano03-41, Ano03j, AJ01b, Bar00a, BJvdB02, DJLT01, GM03, Lag03, MD00, TCM^{+}00, Ano04-28, AJ01a, Ler02, RSS^{+}04, Che00]. Smartcards [CMG^{+}01, GN01b, Ano04h]. smell [PWN04]. SML [GS05a, Kil03b]. sMobile [Yam04]. Smooth [ALZ00]. SMP [KK03b, ZLG08]. Snee [Cal00a]. Sniff [Ano02s]. Sniffer [JBMP03, JKKL04]. Snowbird [ACM01a]. Snugglebug [CFS09]. SO KEEPALIVE [Fox00e]. SOAP [BI02, Cer02, DJLT01, EF02, Eng02, Gun01, Ano04-27]. sobriquets [Way05]. SoC [Ano01j]. social [OOOiM05]. Society [SPS^{+}02, Bea05]. Socket [Ang01, KW01b]. Sockets [Cal03, CD01a]. Soft [Ano03-37, KM02, NK03, PSM01a, PSM01b, Sun01, PSM03]. Softbound [Dud06]. Softech [Ano01i]. SoftQuad [Ano01m]. Software [Ano00h, Ano00i, Ano00j, Ano00k, Ano00m, Ano01h, Ano01i, Ano01j, Ano01l, Ano01k, Ano01m, Ano01n, Ano01o, Ano02m, Ano02n, Ano02p, Ano02q, Ano02r, Ano02s, Ano03-37, Ano03-40, Ano03-41, Ano03-46, Ano04v, Ano04-33, Ano05l, BHS07, BN03, BALV03, BLL06, Cha05a, DFL00, EXA^{+}05, FP03, FS03b, Gib09, HD01, Hsu01, Kaf00, KLL03, Kro00b, Lam03, LBQ00, LL01b, LMK06, LRO02, Lut03c, MD00, MKF06, Off00, RMR03, RMR04, SGV04, SLB^{+}02, SD08, SPS^{+}02, SR06, Sin00, SB00, SNOM01, SASZ03, TGB^{+}04, TSCI01, TMG03, WR00, WK02, Wol03b, ACM01a, AGST04a, AGST04b, AAB^{+}05, Ano02l, Ano03h, Ano03l, Ano03-29, Ano03-35, Ano04-32, BFN^{+}06, BWLP01, Bos04, Bro07, BFMT00, BKŁ01, Coh04, CLN07, DWH01, DS04, DBH04, Emu04, Esq04, FB07, GK08, GM02, Gra04, HJL^{+}01]. software [HLM06, HKI08, Jia00, KS09, Kon04, Lee03, LL00, LL03, LL01c, LHFL07, MORW08, MCHN05, Nam08, NRS^{+}07, NQM06, OSH04, Pan09, PHM^{+}01, PV06, RRP01, Rei05, Ril02, Ril03, Rob00b, RHDB08, San04a, Ses08, SGK09, SS08, SHM09, SKM01, TCSC04, WM00a, Wea04, Wit00, Zhu04, Ano00n, Ano01i, Ano01l, Ano01m, Ano01n, Ano01o, Ano02q, Ano02r, Ano03-35, Ano03-39, Ano03-40, Ano04v, Kro00b]. software/hardware [TCSC04]. Softwarewartung [Wol03b]. SOI [Ano02s]. SOISIC [Ano02s]. SOL [JLV02]. Solaris [Ano01k, Ano01o]. Solaris-to-Linux [Ano01o]. solid [GS00b, Pap00]. SOLO [SCL^{+}08]. Solomon [INM05]. Solr [SPBE09]. Solution [Ano00i, Ano00k, HIBP04, LKL^{+}03, PSDF01, Ano03o, Ano03-33, OBr05, SCWL08, Whi03a, YCFX09]. Solutions [Ano00h, Ano00j, Ano04h, Dar01c, Dar03, GMM00, LL01b, McL01b, CG01, D^{+}00, JA01, LL00, LL03, LL01c, OOM^{+}07, SHHS04, Swa01b, Ano02p, Lut02]. solve [WVMN05, Wil05]. Solver [SGV04]. solvers [GCARPC^{+}01, MAJC03]. solves [Wan02b]. Solving [CP04, MLG02a, CP01, DS00b, HB09, LO00b, LP05, Mor00, Mor03a, Sla00, Wei02a]. Some [Ano05q, HKHK03, CG01, Way03]. sometimes [MMN09]. Sophisticated [Kro00a, BS09]. sort [Rol05, STB08]. Sound [McG03b, SEdM08, BW04, QM09a, SC07]. soundness [Req03, RHDB08]. Sounds [Nil05]. Source [Ano00k, Ano01i, Ano01o, Ano02t, Ano03a, Ano03-37, Ano05k, Bar01b, BHP^{+}01, Egy01, Kuc06, Nas04, Pra03, SHK^{+}03, TEM^{+}01, YLL^{+}07, Ano02e, Ano04i, Ano04-38, Bad00, BP01c, BG04b, EvG04, Eub05, HL02a, KBV08, Liu08, Mam01, MM04, RM07b, SML06, ST09, Vir05, WACBL03, ZK05, Sto01b, Sto01a]. Source-Code [BHP^{+}01, BP01c]. source-level [ST09]. source-to-source [BG04b]. southern [INM05]. SP&E [CY01b]. Space [BFG02, BCR03a, Bar00a, BKY^{+}03, CD03, Hit03, Nis02a, Nis02b, SKS01a, SKS03, And01, FWL03, FWR^{+}05, dCG^{+}02, MSS00]. Space- [BFG02]. Space-Efficient [SKS01a]. Spaces [BD03b, Bow07]. Spam [MSF03]. Spar [vRKS01, vRKS03]. SPARK [LH03b]. Sparse [LUH^{+}05, dCG^{+}02]. spatial [Ran03, Woo02]. Speak [AM02]. Speaking [Van04]. Spec [Ano02q, Bar01a, GPW05]. Special [Bak00, Dek00, EL01, Fox00a, Fox00b, Fox00c, Fox01a, Fox05, HR04b, KCF01, Wut00]. specialisation [Ren02]. Specialization [PP02b, GES^{+}09, SLC03a]. Specializing [PP02a]. Specific [Dmi02, TT01, VKB01, ZS01b, Ano05f, CO06, HZS08, ZS01a]. Specification [Ano03s, Ano04l, AW03, Bar01b, BCDdS02, BS04, BL03, BDJ^{+}01b, BW03a, BW03b, Bro05, BFM^{+}02b, BW03c, CH02, FMMd03, GJSB00, Har00a, Hep04, JV04, KF05, KM04b, MP01b, vdPE02, Rot05, Sun01, WP03, YKB02, vdBJP01, Ano03-36, BA05, Bol00, BS00b, BS09, BHR02, BH02c, Cog03, Dob01a, GJSB05, Jen01, LBR06, LYC02, LG00b, PvdBJ01, QGC00, SH04b, SRD00]. Specification-Based [BL03, KM04b]. Specifications [AČMN05, HD03a, TRVH03, HRD08b, Kes04, Sha00b, WA01, Yua04]. Specifying [BJvdB02, CY02, Sta04b]. specimen [Rol08b]. SPECjvm98 [LJN^{+}00]. Spectral [Bus02a, Bus02b, Sar03, SYAS05]. speculation [NRS^{+}07]. Speculative [LCHY03, PV06]. Specview [Bus02a, Bus02b]. Speech [Ano02t, Bar01c, Cha05a, Zhu04]. Speech-Enabling [Ano02t]. SpeechStudio [Ano02s]. Speed [Ano03p, Gut00, Kie01, VKB01, Ano04b]. speeding [MRB06]. SpeedStep [Ano00m]. Speedup [CCF^{+}02]. Spezifikation [Hep04]. Spiderweb [Ano00j]. spike [Ano04u]. spikes [Ano04z]. SPIN [Lut03c]. Spineless [CiLH01]. splitting [NIKN06]. SPMD [AGS01, Sta00]. spoken [OHL^{+}05]. spot [LMK08, TBM09]. Spotless [MS00b, SMES01]. Spread [WXW^{+}05]. Spring [GT05, JHA^{+}05, TGL05, WB05, WB08]. Springer [Azi06]. Spyglass [Kro00b]. SQL [ISO08, Ano05k, Ebe02, KM07, ME00a, Tho03, Yua02]. SQL/JRT [ISO08]. SQLAlchemy [Gar09]. SQLite [Ano04-38]. SQLJ [ME00a, Pri01]. Squint [Mur07]. SRAM [Won03a]. SRec [VIPCUF08]. SSA [MGM^{+}06]. SSJ [LMV02]. SSL [ZFK04].

SSP [WBF^{+}06]. St [Tra00b]. Stability [SBA01, Rob04c]. Stack [Ano04m, CGS^{+}03, Ran02, Ano05m, Cha06, TCC02, TCSC04, SCEG08]. Stack-Based [Ran02]. Stacks [Won03a, LC05]. Stage [Gar00]. Staged [CMJL09]. stages [PFJ05]. Stalker [Ano00i]. Stand [Ano03-52]. Standard [BH05b, FSS06, Pla00, Qia00, BDLM04, Gar09, Kon03, Suo04, Fig00, NIS00, Pla00]. Standardization [Egy01]. Standards [Ano04c, Bro00, Lea00b, BA07b]. Star [Lut03a, Ano04b, Lut03a]. Starbase [Ano00n, Ano03-40]. STARC [EKVM07]. StarCore [Ano01j]. Stardock [Ano01o]. StarJIT [ATBC^{+}03]. StarNet [Ano00j]. start [Ano03w, WG02]. started [Ell06]. starter [WMM04]. Starving [Rob01a]. Stat [Nar05]. State [ADR09, GSW00, Rei00a, Sur01, WTV03, ABL08, Cor00, DGGD08, DH00, Gri03]. State-dependent [ADR09]. Statements [Zam03b]. Static [Ano01h, CHS01, CH02, Cha06, KMS04, NC04a, Nel04, NE04, PCC01, PL05, RKG04, SR06, TM08, WGSD07, Woo05, XJC09, BCV09, CD08, DH08, DMP09, EKVM07, FLL^{+}02, GPF08, HO03, HO07, HS08, Lan04, LPH02, NAW06, NA07, PH00c, SBMG00, AFF06, FFLQ08, Wol03b]. static-dynamic [CD08]. Statically [VMMF00, WSM06, Ren02]. statically-generated [Ren02]. Station [Bar00a]. stationary [UL08]. Stations [EGLZ02]. Statische [Wol03a, Zus03, Wol03b]. Statistical [HKL09, Zus03, Aki02, HLT09, NHY^{+}04]. Statistically [GBE07]. StatSoft [Ano01o]. Status [RBC^{+}05]. STDOC02 [ASS03]. STDOC09 [CL03b]. Stealth [Ano03-40]. Steam [TC03]. Steeb [Pap05]. Steering [Lut01]. Steganography [Hun05]. Stellarator [PDCL02]. step [EFO08, BDE^{+}03]. step-by-step [EFO08]. stepwise [MR09]. Steve [Mor03b]. Still [SAFG03]. Stirring [Nis02a, Wil00d]. STM [BKO09, MBS^{+}08, SMAT^{+}07]. Stochastic [LMV02, PP02c]. Stopping [HM01b]. Storage [ACM04, Ano02m, BH03, Hei03a, LUH^{+}05, VT01, HYX05]. Store [Bar01c]. stored [Ano03-42, HF06]. Stores [WH01]. Storing [ST06]. STPTP01 [CY03]. Straight [BHP^{+}01]. strangers [Urb09]. strategic [WCK^{+}07]. Strategies [ACM01e, Egy01, Goo02b, OGA^{+}01, BWW^{+}03, FLMS06, MLM^{+}08]. stratigraphic [HPH03]. strayed [Rol08a]. Stream [All00b, WDSD02, SPGV07, ZP03]. StreamFlex [SPGV07]. Streaming [KKK04]. Streamlines [Ano03-40]. Streams [Ano00k, CS06]. strengths [Ano04g]. Stress [ABV00, LAB^{+}00, ZD02]. Stress-testing [ZD02]. Strictly [BS09]. Strings [All00f, Cox01a, BV05, KOO08]. Strong [CWHB03, SMSAT08, ZFK04]. stronger [Ano03-46]. strongly [BKO09, vMV05]. Structural [Chi00, GCEO05, LBR00, GM08, GV02b, LFM09, VDMW06]. structure [CZ02, EVS07, HCMM00, HCB04a, SB07]. Structured [DT02, WHKS01, ADT03, PV03b, SSGS01, Tre02c]. Structures [Ano02s, BO09, GT97, GT04, GT06, GT10, KC01, Mas01, TGV^{+}01, WP00a, ZD02, And02, Bai03, Bud01, Col01, CHJB07, Dro01b, Fek02, GEVZ09a, GT01, GS04, Hub01, LO00a, Mad01, Mai03, NM02, PHBM05, Pre00b, Sah00, WB01, Wei02a, ZKR08, vRS05]. Struts [FG05, Cav02b, CK03a, Cav04, For04b, HD03c, Sig05, Spi03b]. STS [Ano00i]. STSimJ [CWZ04]. Student [HTY^{+}03, SS07, Djo08, ER09, Fle00, PJ05, TETPQ08, TZ01, WKB02]. student-constructed [Fle00]. student-written [TETPQ08, TZ01]. Students [HMRM03, LAB^{+}00, Ros02b, AT01, BP02, Fek08, Fle01, JCOP07, PB06, Rio02]. Studied [GKMZ04]. Studies [NW03]. Studio [Ano04-36, Ano04-35, Ano08, Lia03a, Sur04b, W^{+}04, BI07, Ano03-41, Pra08]. Study [Ano04-34, BCMT03, BS04, BL03, CR02a, CK05, HS00a, Hui02, KJ02, KMSL03, KX04, LAT04, MORW04, NMH^{+}02, RCdBL02, Sat02, SYN02, BBS04, BS00b, BA09, BS01, CCK^{+}08, CHL^{+}00, CMS07, Die00, DAK00, ER09, GEVZ09a, HJvdB01, IKY^{+}00a, KPPÉR06, KLS00, MT07, OKN01, RHR02, RZW01, Roc01, SS02, SCBH09, SMTZ09, VZGE07, VP05, vRS05]. Studying [CKK^{+}04, GHBG^{+}03a, GHBG^{+}03b, Hig04]. stuff [For06]. Stunden [Ste08b]. Stupidity [Lut03a]. Style [VV05, VAB^{+}00, KS07, Lan00, LHFL07, Ras03, Che05]. Styrene [BD03a]. Sub [SPR^{+}03]. Sub- [SPR^{+}03]. Subject [Ano04i]. Subroutines [KW03, Wil02, Cog04]. Subscribe [Hou00, RG00, Rou02b]. Subscriber [CM02]. Subscription [Ano05m]. Subset [KPKL03, Req03, TP02]. subsets [Ano03h, RK02]. Substance [Lea00b]. Subsumption [BO05]. Subsystems [VT01]. Subtleties [Lai08]. Subtype [PV03a, Duc08, KR01a]. subtyping [FLF01, IV06]. succeed [Mer04]. Succeeding [CZ01]. success [RVZ04]. Successful [HB09, Kun02, Lut03c]. such [Ano05f]. SugarCubes [BS00d]. Suitable [BBDT02, Vog03, Wol03b]. Suite [Ano01h, Ano01n, Ano02m, Ano02n, Ano02t, Ano05k, DHPW01, Kuc06, SBO01, ZS01b, Ano03-35, BBBD01, BA04, BSW^{+}00, GPW03, Sar03, Vir05, Ano01i]. suited [OOM^{+}07]. Suites [Ano05f, Ano05m, GPW05]. summary [BH02c, Dob01a]. Sun [Moo03b, TBM09, Ano03-47, Ano04g, Ano04i, Ano04r, Ano04w, Ano04x, Ano04-36, Ano04-35, Ano05f, Ano05m, CR02a, Dob01a, DA04, HS00a, Lea00b, Lia03a, Pau03, Sur04a, Sur04b, Van04, dSC06]. Super [Ano00i]. Super-Symmetric [Ano00i]. Superclasses [LSW08]. Supercomputing [ACM00a, ACM04, Ano00l]. Superinstructions [CGEN03]. superoperators [BNV08]. Supervisory [LH03a]. Support [Ano01j, Ano03-40, BMR02, BCS07, BCH02, BP01d, CA04, CCC^{+}04, CF02, DL02, DFA03, HJL00, HFL03, HIBP04, KNY03, Kro00b, MD00, MPG^{+}00, MMG01a, Rob04b, SG03, WCCL05, Ano04g, Ano04k, Ano04-31, BP03b, BCL^{+}06, BRBY00, CCK^{+}08, GK05, HT06, LCFL04, LLCF08, LHS03, Mur07, SKC09, SNO^{+}07, SFMH01, THL03, Tre02c, WK08a, WK08b, WK08c, ZLG08]. Supported [AddS03b]. supporters [Ano05h]. Supporting [Ano03-28, AGS01, CW04a, Fab02, Fig00, JSSM04, LK01, MMG03, PSM01b, TETPQ08, ADT03, Ano03e, AK09, BS01, RPP07]. Supports [Ano03-37, CLL03, Ano02l, SML06]. sure [Ano05n]. Surface [MD00]. surfaces [Nik03]. surreal [DA04]. survey [LAL02]. Surveying [Lut03b]. Susceptibility [CMB^{+}01]. SuSE [Ano01o]. SUSSMicroTec [Ano02r]. Sweet [Lan04]. Swing [Gla06, Gut00, KK03a, LEW^{+}02, LEW^{+}03, ABL08, EL02, Gol00, MA05, Top00, WWJ07, WW09, Wra01]. SwingStates [ABL08]. switch [Ano03-36]. Switching [RCdBL02]. Sy [USE01c]. Sybase [DHMT00]. Syclo [Ano01j]. Symbolic [PV04, Tra00b, LP05, Nor00]. Symmetric [Ano00i, CLCM00]. Symposium [Ano00a, Ano01b, Ano01g, IEE03a, IEE03b, LL08a, Tra00b, USE00c, USE00d, USE01b, USE02, ACM03a, Ano02b]. Synchronization [BKMS04, Bec01b, Hei03b, RM04, ASCE03, CY01a, CY01b, CGS^{+}03, MSV05, Rob00a, Rob01a, Ruf00, RD06, SS06, VTD06]. synchronization-related [RD06]. synchronize [FJ05a]. synchronizer [Lea05]. synchronous [BCHP08, Bow07, PC08, SLS09]. synchronously [PC03]. Synergetic [Ano00k]. synergies [CF04a, CF04b]. Synergistically [NLFA02]. Syntactic [BP01a, Dep03b]. Syntax [Rum01, vdSPP05, BH02b, BTV06, Gri06, vMV05]. Synthesis [AČMN05, HKK^{+}01, YKB02]. Synthesizing [WHW01]. Synthetic [SGV04]. syst [Sci07]. System [AddS03b, ÁdBdRS08, AA04, ABG02, AG03a, AG03b, Ano00n, Ano01k, Ano01n, Ano02m, Ano02r, Ano02s, Ano03-38, Ano03-39, Ano03-40, Ano04v, Ano04-37, Ano05a, ABH^{+}00, BKH02, BH02a, BLW00, BFM^{+}02a, BFS^{+}03, BFS^{+}04, CLCC02, CKV^{+}02, CO03b, CKM04, CKKH03, CK05, DH04a, DYH05, Det01, DMP05, EM03, FM03, FOS^{+}04, FBS04, Gam03, GMW^{+}02, HFL03, HTY^{+}03, HKL09, Hon05, HS02b, II04a, JPJ05, JKJ05, KK03a, Kog04, KY03b, KS01b, Lau03, LH03a, Lia03b, LZZ03, LRO02, Lut00, MWL00, MD00, MLG02a, PDCL02, Pot04, SGV04, SDPM04, SKC09, SPS^{+}02, SM01b, Shi03a, SSV05, SL04, TFL^{+}04, VWS^{+}05, VHL01, WS01a, WFGK03, YHL04, AAAG^{+}05, ÁdBdRS05, AYWM08, Ano02l, Ano03-44, Ano04-32, A^{+}01, BH05a, BCS09, BAD^{+}09, BI07, BDFL04, BR01b, Caa00, CVW03, CHMB04, CSK^{+}02, CO03a, CW03b, CBGM03, DPT^{+}02, Dep03b]. system [EL04, Emu04, Eng06, FW02, Gel00, GM05b, HJL00, HvE02, HWM01, HKI08, HO03, HO07, HYX05, Jam01, Jia04, KH00, Lan02, Lex02, LJN^{+}00, LW03, MBED06, MAWW^{+}01, MR06, MC06, NB00, NB01, OMK04, PV03b, PRB07, RZW01, Rob06, SFMH01, SJ01, Sha01, Sha04, SSC00, Sta00, SSP07, TABP07, VIPCUF08, WF04, ZABL09, dGNv04, Ano00m, Ano01o, Ano04b, Ano05f, GEAS00, Pra08, WCK^{+}07, Ano08]. System/390 [GEAS00]. systematic [NAR08]. Systeme [Wol03b]. Systemen [Ano03-33]. SystemJ [MSR09]. Systems [ACM00b, ACM01d, AJMJS02, Ano00h, Ano00i, Ano00j, Ano00k, Ano02o, Ano02s, Ano03-33, BTS^{+}00, BIB05, BCS02, BH05b, BR06a, BG04a, CDFR04, DS00c, DFT03, Dud06, FVK01, FMMd03, Gal01, GP03, HT03, IEE03b, KPKL03, KFLN04, KMOS03, KMSL03, KK03b, KC03, KWK03, LN04, Leh01, Leh02, LL08a, Lut02, Lut03c, Lut03b, MJ06, NSI03, ONRV08, Par05, Pra03, RJFG03, SBCK03, SSA03, SG03, TA04, TP01, USE00c, USE01a, VWS^{+}05, VDPC01, VB01a, VHL01, WK02, Wri03, Zhu03, AR08, ANMM06, Ano04y, Ano05a, AVY08, BNV08, Bog01, BW01b, BW04, CSCM00, Fer07, GK05, GB01, HKS^{+}07, Hub02, JPB^{+}08, KKG09, Lab09, Lan05b, LHFL07, Mer00, Moo02, NHY^{+}04, NZM03, Nis03, OSH04, OOM^{+}07, RVJ^{+}01, RK02, Ric01, Rob02, RHDB08, SCB09, SFMH01, SGK09]. systems [SS08, Sto02a, SKM01, VDPC03, WAF00, Wan02a, WCC04, Wol03b, Zar02, ACM00b, Ano01h, Ano01j, Ano01m, Ano02t, Ano03-34, Ano03-40, Ano04i, Way05]. Syware [Ano02q].

T [Mas01]. Table

[LCHY03, DHS02, FCW01]. Tables [Sea02, Yua02]. Tackle [Coc02, Sub08]. tackles [Ano03o]. TADDs [RWZ09]. tag [Wei02b]. Tagless [CiLH01]. TAI [HTY^{+}03]. TAI-18-5 [HTY^{+}03]. Tailfit [HZC^{+}04]. tailored [Ano05f]. taint [TPF^{+}09]. Taiwan [Ano01p, Ano03j]. TAJ [TPF^{+}09]. take [Mer04]. takes [ABI^{+}07, Mer04]. taking [Ang06]. tale [HW00]. Talent [Bar01a]. talk [Urb09]. Talker [AJB^{+}04]. Tally [CK05]. Tamassia [Mas01]. Taming [Fre04, Hab04, Hol00a, HSSC05, RC04]. Tamper [CHL07]. Tamper-proofing [CHL07]. Tandem [Lou05, DPT^{+}02, MSR09]. Tape [Gib01].

Tapestry [For04b]. Target [KK04b, LBJ02, LBJ05]. targeting [DGMY06]. Tascom [Kro00b]. Task [RBC^{+}05, RBC^{+}06, SPR^{+}03, ABG^{+}08, ZABL09]. Task-Level [SPR^{+}03]. Tasking [Shi03a, Ano01o, JDJ^{+}06]. Tasks [PSM01b]. TAU [SM01b, SM03a]. taxonomy [Wor02]. Taylor [Cha03]. Tcl [SML06, USE00b, Lai01, Pre03, Ros00, ZK05]. Tcl/2k [USE00b]. Tcl/Tk [USE00b, ZK05]. TCP [CD01a, Cal03, KW01b]. TCP-Socket [KW01b]. TCP/IP [CD01a, Cal03]. Teach [JBMP03, AK00, Bru04b, Bru05a, CL03a, CLZ06, Hag00a, Hun03b, WN05, WSP02, WMM04]. teacher [SMS^{+}04]. Teaches [LAB^{+}00]. Teaching [AF03, APA04, Bar02b, Bec01a, BWC^{+}05, BF03, BB03, Bur03, CR02b, DV07, ES05a, Fek02, Fek08, Fre04, GS08, GL08, GGG03, JCOP07, Lam03, Mer00, MKS^{+}03, NW03, PH03, RP03a, RKK03, SU03, Sch00a, Sch02, Sco03, Wol01b, Wu05, XSD07, Yan03, BA04, BZ05, ES05b, Gag02, Gra04, Gri08, Gri02b, KR01b, KM04c, LDB^{+}03, LW03, MB05, Pan09, RRP00, RRP01, RM08, Rob03, Sci07, Soj03b, Utt06, WVMN05, XM06]. teaching/learning [Pan09]. teacup [Joh06]. Team [Bar00a, Mer04, Bar00a]. TeamStudio [Ano03-48]. Teamware [Ano00h]. tearing [PPJ03]. Tears [HP04]. Tech [Lan04, Lut03a, Van04]. Tech-nically [Van04]. Technauts [Ano00j]. Technical [Our02, Rei00c, USE00a, BD04, MMG00b, Lut03c]. technicians [Coh04]. Technique [KK04b, MMK04, SMK02, Cog04, JPSN09, LYC02, Li02, Sto01a, SYN03, SYN06]. Techniques [BTS^{+}00, BF02, Bul00, CHK^{+}04, DEJ^{+}01, DEL^{+}02, ELM^{+}04, Kal04, KCSL00, LDE^{+}02, SSM04, TSL^{+}02, WF00, BCM05, BVD01, CY04, Coh04, Die01, EL01, GEG07, IKY^{+}00a, LLdA08, Lot02, Gal02, She01a, SCS01, SM03b, WJH06, WM00b, WF02, Sto01b]. Technological [SLB^{+}02]. Technologie [Ano03-27]. Technologien [Ano03s]. Technologies [Ano00i, Ano00k, CL03b, Fri02, Gat03, HL04, KLL03, KX04, Lia03b, ME00a, USE01a, ZL05, Cha05a, Ano04-27, AGG02, Chr00, DH00, EK01, Gho01, Jor02, TAW03, Zhu04, Ano01k, Ano01n, Ano02n, Ano02q, Ano03-30, Ano03-35, Ano03-39]. Technology [Ano00a, Ano00j, Ano01b, Ano01j, Ano01g, Ano02b, CR02a, DJP02, DYH05, Dmi02, EXA^{+}05, GS00a, KW02, Kum02, LB00, LD03, LS04b, Lut00, Muc02, Pau03, San02b, Sch04b, SSA03, USE01c, USE01b, USE02, VN03, Wan03a, WGC09, Wel03, dSC05, Ano01f, Bar02a, Bri05, Che00, CG02, Ham02, ISO08, Kic04, Kum01, LHFL07, LSK^{+}02, LW03, LHS04b, New00, PT09a, Rod01, Cha03, Ano01h]. Technology-Based [EXA^{+}05]. Ted [SPS^{+}02]. tekhnologiiu [Saf02]. Tektronix [Ano02s, Ano02n]. Telecollaboration [dOHS^{+}03b, dOHS^{+}03a]. Telecom [Ano00k, Ano02q]. telecommunications [JA01]. telegraph [SFMH01]. Telelogic [Ano01k, Ano02s, Kro00b]. Telematics [HE03, San02b]. Telephony [Ano02s, Mar00]. Telerobotics [RPJ04]. Temperature [Lia03b]. Temperatures [BD03a]. Template [SP03]. Templates [Bat04, Vel01, AK09, XOWM06]. Temporal [BNO03, IS03, SV05]. ten [Eic05]. tensor [MAJC03]. tensor-based [MAJC03]. Terabytes [IEE02a]. Teraflop [Ano00l]. teraflops [CSFS00]. term [ISO05]. terminals [Ano03-51]. Termination [HJ00]. Ternary [DH04b]. Terrain [Ano02m, OG05]. Tertiary [VT01]. Test [Ano02n, Bar01b, BL03, BDJ^{+}01b, CQX^{+}09, EFN^{+}01, MdB01, Pip03, SGV04, VPK04, Ano03-34, CSFS00, Duc08, EFN^{+}02, GKM01, HJL^{+}01, JMS02, Man01, Ano04b]. Test-Driven [Pip03]. Tester [Ano02o, Ano02t, CS04]. TestEra [KM04b, KM04a]. Testing [Alb03, Ano01o, Ano02m, Ano02n, Ano02r, Coh04, DFW04, DiM04, FRMW04, Goe01, Goo02b, KM04b, LCS04, Liu04, Lou05, Lut03c, MS05, NS03, PR04, RS05, RMR03, RMR04, SB00, BKM02, DHS02, EFG^{+}03, FMRW05, HT04, LFM09, Lin03b, LHS03, NP02, Off00, OSH04, PJ09, Sen08, Ste05, SCFP00, TE04, Ton04, VMWD05, VDMW06, Wit00, ZD02]. Tests [Coc02, Lin03b, PV03a, TETPQ08]. Texas [USE00b, USE01a, CNB00, IEE02b]. Text [All00d, AGH05a, Kro00b, Lut03a, NLFA02, Wei01, BV05, Mas00, Tho03]. Text-Based [NLFA02]. text-search [BV05]. textbook [GS00b]. textures [Nik03]. their [HG07b, IH01, MSLL07]. theKompany.com [Ano01l]. them [WVMN05]. theme [Ras03]. Theorem [Ber01c, GKW04, GN01b, DNS05]. Theorems [Moo03a]. Theoretical [SSM03]. Theory [Rap03, RM08, BLLB08, ET05, Ham07, Hub01, VVV04, ZABL09, Bla03]. There [Ano05n, Bri05, CAF04]. Thermodynamic [TC03]. these [Coh04]. they’re [MMN09]. Thin [BKMS04, SFB07]. ThinAirApp [Ano01i]. Things [Lut00, BVPE06]. Think [LAB^{+}00]. Thinking [Eck00]. Third [GAR04, NIS00]. Thomas [Fox01b]. Thorn [BFN^{+}09]. Thought [Vel01]. Thread [CC04, CWZ04, DGK^{+}03, Hag02, Hei03b, MP01c, Sat02, WP04, Whi03b, ZWL03, ABG^{+}08, BHK^{+}04, CY01a, CY01b, Fek08, Hyd00, MC06, Oga09, ZLG08, SKP^{+}02]. thread-based [ZLG08]. Thread-Local [DGK^{+}03, Whi03b]. thread-safe [Fek08]. Thread-Sensitive [CC04]. Threaded [GH03, JV04, CWHB03, Chr01, EFG^{+}03, GCRD04, Sto02b]. Threading [DHR^{+}01, FWL03]. Threads [ÁMdB00, ACR01, BLPV04, Hol00a, MZ04, PSM01a, Pet03, San04a, TS04, WTV05, BZ07, BS00d, Cal02, Lan02, OW04, PSM03, PG03a, SKP^{+}02]. Three [FVK01, MMG01a, NS03, OJJ00, CLP06]. three-year [CLP06]. Thresholds [JHJX04, YDWL04]. Throughput [MHZG06, BG03, SPGV07]. throw [AH03]. Thrown [AHKR01]. Throws [Ano03-31]. Ticket [GM03]. Tide [Wan04]. Tier [DF03, LLMK03]. tiers [LJ07]. Tiger [Fre04, Ano05n, Ano04w, MF04]. tight [Ano04g]. Tiling [PH02]. Tim [Ano04-29]. Time [APA04, Ano01i, Ano02m, Ano03s, Ano03-52, BFG02, BR01a, BN03, BNO03, BG04a, BD01c, Bro03a, Bro03b, BW03a, BW03b, Bro04, Bro05, BW03c, CW03a, Cav02a, CA04, CKC^{+}02, Chi00, CS02, CS03, DC03b, Dib02, FBR^{+}03, GKM03, GKMZ04, GKW04, GNYZ05, Gle02, Har00a, HIBP04, Hig04, HWB03, HWB04, JT04, Jia04, KVK^{+}04, KMEA04, KNY03, KM02, KK03a, Kro00b, KNG02, LDM04, LD03, MB03, MLJH04, ME00b, NK03, PV03a, PSM01b, PUF^{+}04, Pla00, Pot04, RW03b, Sch04c, SSM04, SLC03b, SCLV04, SOT^{+}00, SYN02, Sun01, TGB^{+}04, TSL^{+}04, Uma02, Wan04, Wat02, WP03, Wel03, Wil01b, Won05, YLL^{+}07, dSC06, ABC^{+}07, ABI^{+}07, ABI^{+}09, BCR03a, Bol00, BSBR03, BALP01, BALP06, BD01b, BHR02, BH02c, BW01b, BW04, CC01, CC03, D^{+}00, DV01, FCHE02, Gad03, GES^{+}09, HT06]. time [HKS^{+}07, HKM^{+}09, Hor00c, ITK^{+}03, Ive03a, Jen01, JKJ05, JPB^{+}08, KPH^{+}09, KKL^{+}04, KM08, KBP^{+}03, KWK05, LYK^{+}00, LYM04, LMK08, LH05, OOK^{+}06, PSM01a, PSM03, PHV07, San02a, San03, San04a, She03, SAB^{+}06, SYK^{+}01, SYN03, SOK^{+}04, SYK^{+}05, VHBB03, Wan02a, WLW^{+}03, Wel04, ZABL09, Ano03s, Dob01a, IKN03, IKY^{+}00b, IKY^{+}00a, KSK04b, She03]. Time-Efficient [BFG02]. time-portable [ABI^{+}07, ABI^{+}09]. time-saving [D^{+}00]. Timed [SJG03, WDSD02]. Times [SGF^{+}02]. TimeSys [Ano00h, Ano03-38]. Timing [HWB03]. Tina [SAWW01]. TINI [Wil00a]. Tipps [DHMT00]. Tips [AE06, BM01, MA05, Ano05q, EA06, Pan09]. tissue [KGH^{+}05]. TJ [PDCL02]. TJ-II [PDCL02]. tjener [HJL00]. Tk [USE00b, Ros00, ZK05]. TM [ISO08, Kic03, Ren00]. today [CZ01, Nis03]. Together [ME00a]. Tolerant [FK03, TMG03]. Tolerating [BM08]. Tom [Cal00a]. tomahawk [STB08]. Tomasulo [EKEL01]. Tomcat [BD03c, BD07, Ler01d]. Tome [Lut03c]. Tomography [SGV04]. tomorrow [Ano04c, PB06]. Tone [Lut02]. Tony [Fox01b]. Too [Wil00b, Ano04-29]. Tool [AddS03b, ABM^{+}03, AL04b, Ano00o, Ano01h, Ano01i, Ano01m, Ano01n, Ano01o, Ano02n, Ano02o, Ano02p, Ano02r, Ano02s, Ano02t, Ano03-38, Ano03-39, Ano03-40, Ano03-41, Ano04b, BIB05, BCDdS02, BCE^{+}01, BRC03, Bus02a, Cha05b, CE01, CK05, Eng00, Fel04, Goe01, HD01, HR04b, HKHK03, Jen02b, KKL^{+}04, KNY03, LHS03, MD00, Mam01, MLG02a, MS03, PR03, RST^{+}04, RPJ04, RLR00, SEGS03, VDPC01, Wat02, Yam04, YKS^{+}02, ZG04, Ano03-34, Ano03-35, Ano03-36, Ano04q, Apr05, BK08, Bod04, Bus02b, BRBY00, CTF03, Esq04, Fal00a, Fal00b, FMA02, FTD03, FL02, GV05, GP05, GST05, JHSL03, KJBH^{+}00, Kim02, MMU04, MKKC08, SD03a, SNO^{+}07, SS08, SCFP00, TZ01, VDPC03, Wis06, Woo03]. Tool-Assisted [BCDdS02]. Tool-Kit [BRC03]. Tool-Supported [AddS03b]. Toolbook [Ell00]. Toolbox [Coh04]. Toolchest [Tre02b]. Toolkit [Ano01h, Ano01n, CWZ04, CN03b, KS02b, Ros00, Sch02, SC05, TCF^{+}03, Wil01a, Wol04, ABL08, HL02b, HBX^{+}04, SML06, SYAS05, VVV04, Ano00m, Fox00d, LS03]. Toolkits [BCMT03, Ras00]. Tools [Ano00n, Ano01i, Ano01l, Ano01m, Ano01o, Ano02o, Ano02s, Ano02t, Ano03p, Ano03-38, BM01, Ber05b, BOT02, BW01a, CBD01, FJ05b, Gat03, Kuc06, LBQ00, Lut03b, LAB^{+}00, MA05, Nas04, WF00, ZK04b, ACM01a, dS02, Ano02d, Ano03-35, Ano04b, BA04, BCS09, BC04, CM02, Coh04, CGM06, EF02, Gar09, Ham07, HL02a, MBED06, OJ09, PL03, RRP00, RRP01, Sma08, ST09, Vir05, WMRT^{+}05, WF02]. Toolset [Ano01i, BDHdS01, ZK05]. Top [Bur02]. topic [Ano04p, S.04a, S.04b]. topics [BLLB08, WN05]. Topological [CD01b]. topology [EGST08]. tops [Ano04z]. Toronto [Jac04b]. TOS [NB00, NB01]. Total [Kog04]. Totally [DHR^{+}01]. TotalView [Ano00i]. Toulouse [IEE03a]. Tower [Ano00j, Reg02b]. TowerJ [Ano00j]. Trace [GES^{+}09, JR05, BDE^{+}03, HEJ09, Ing09]. Trace-based [GES^{+}09]. Trace4J [Ing09]. traces [BA09, HBM^{+}02, HBM^{+}06, WR08]. tracing [HSB09]. Tracker [MD00]. Tracking [Ano05p, BNK^{+}07, Pau01, Ren00, AWS^{+}09, WAB^{+}04]. Tracks [Bar00a]. Trade [CKK^{+}04, CD01c, CD01b]. Traditional [GS05a, Ano05i]. Training [BBHL01, DD02a, GHM^{+}01, Hal01a, LAB^{+}00, Ste08b, SMS^{+}04]. Transaction [BM03, BL03, EQT07]. transaction-aware [EQT07]. Transactional [Ano01l, CMC^{+}06, CCC^{+}06, HLM06, ST06]. Transactions [AL04a, HP04, Pro01]. Transfer [BW03a, BW03b, GKM03, ZK04b, BHR02]. Transformation [CDFR04, Wan05, BDLM04, WBGM05]. transformational [WBF^{+}06]. Transformations [AGMM00, CKM04, KMS04, SL01, BG04b, HB08, LJ08, ST09, TT08]. transition [Sib00]. Translate [SLPO02]. Translating [AH04b, CDFR04, EK03]. Translation [AAD^{+}01, CFLL03b, EGLZ02, Gar00, SD01b, AAD^{+}07, GEAS00, Oi05, Oi06, Oi08, SD03b, VN00]. translation-based [Oi05]. Translator [Ano02m, LN04, RWZ09, TSCI01, Röß06]. Translators [CN03b]. transparency [GJ09]. Transparent [Ano02q, Bet05, FK03, IKKW01, PSH04, RW04, SMCS04, ZWL03, AZ02, ST09, WK08d, WIC08].

Transparently [AFT^{+}00]. Trap [KKN00, Sta04a, SMCS04]. TRAP/J [SMCS04]. Traps [CYH04, MH02, BG05]. Trash [Bar01c]. Traveling [Bar01c, TCM^{+}00]. TrAX [Har03]. Treaty [DA04]. tree [BK03]. Treemap [KB04b]. trees [DG02, vMV05]. Treeview [Sal04]. Treewidth [GMT02]. Trends [Zdr09]. Trevor [Che05]. triangular [MCLDP01]. Tricks [AE06, EA06]. Tries [Pau03]. Trifles [Wil03d]. Triggers [AA02a]. trivial [Hug02]. True [AZ01]. trust [Ano02w]. try [Ano04g]. TS [Chr05]. TS-05 [Chr05]. TTM [BC04]. tu [DOR05]. TUG [SBH^{+}04]. Tulach [Mil08]. tuned [PC03]. Tuning [CSK^{+}02, Red01, Shi00, Shi03b]. tunneling [JKH^{+}04]. Tuple [BD03b, FWR^{+}05]. tuples [vRS05]. TurboPower [Ano02o]. Turing [CM05c]. Turning [DJLT01]. turtle [MRB06]. Tutor [GLS02]. Tutorial [CWH01, Coo00, GMM00, Kod04, BD04, Fla00, Fla04b, Hap02, Hig03, LS00, Rob06, ZCR^{+}06]. Tutorials [HHKS03]. tutoring [Emu04]. Tutors [Kum04, Kum05]. TV [Kro00b]. Twenty [LL08a]. Twenty-Seventh [LL08a]. Twister [Luk04]. Two [Ano05o, BALV03, Bur03, Lam03, Pra03, AHN02, HW00, KS07, MCHN05, NHY^{+}04, SCBH09, WBGM05, XSD07]. Two-Dimensional [Bur03, WBGM05]. Two-Guys-in-a-Garage [Pra03]. two-level [KS07]. two-year [XSD07]. Two’s [RW03a]. Two’s-Complement [RW03a]. TX [ACM00c]. TY*SecureWS [LKL^{+}03]. Type [AS03, BBDT02, CHP^{+}08, CG01, DTD04, DMP05, FF00, FM03, GF07, KR01a, LST02, LST03, MPG^{+}00, RW03a, SSV05, WS01b, dMSAV08, ANMM06, BAdMS08, BAD^{+}09, BR01b, DGGD08, FF08, GES^{+}09, GE08, HO03, HO07, Hor00c, Lan02, PRB07, PH00c, RHDB08, SI09, SC08, Vir03, WK08d]. Type-based [FF00]. type-passing [Vir03].

Type-Preserving

[LST03, CHP^{+}08, LST02]. Type-Safe [MPG^{+}00, WK08d]. typechecking [MRC03, TTS^{+}08]. Typed [BBC07, vMV05]. Types [AFF06, BCS07, FFLQ08, FR00, ISO08, II04a, Jac03, KT04, BSBR03, CCKP06, FX07, IV06, IV07, Our02, PT09b, QM09a, Siv02, VB01b, WB01]. typesafe [Lan04]. typestate [BBA08, BA07a, FYD^{+}08]. typestates [BA05]. Typing [RE01, DMP09, GM08, RR01]. Typings [AZ04]. Typography [SBH^{+}04].

Ubiquitous [TP01]. Ucigame [Fro08].

UDDI [Cer02, Tre02a]. UI

[Ano02w, Yua04]. ULT [PG03a]. ultimate [FL02]. UltraLightClient [Way05]. UML [Dud06, AU02, Ano01m, Ano01n, Ano03-39, Arr01, BLL06, CQX^{+}09, DFL00, GDB02, HBR00, Hub02, Hun00, Kes04, Kno02, Kro00b, Lan05b, LT02, Meh02, MORW04, MORW08, Ree02, SLPO02, Wam02]. UML-Based [Meh02]. Unauthorized [Ano02s]. uncaught [JCYC04]. uncertainties [LL01d]. Uncertainty [BNO03, SPB01]. undefined [BNK^{+}07]. under-represented [PB06]. undercut [Ano05m]. Undergraduate [BLPV04, YL03, Chr00, GCF^{+}01, PHM^{+}01]. Undergraduates [BBHL01, TBM09]. Understand [DeP03a]. Understanding [BFN^{+}06, BZ07, BALV03, BAJ01, Bud00, Mar00, ME00a, NLC03, ST00a, Wal02b, ZXNH02, HSD04, LJ08]. UnForm [Ano00k]. Unicode [Uni01]. Unified [AW03, BALV03, HKS02, YHL04, ABG^{+}08, Hun00]. Uniform [Bac01, Eug06, FGLS04, Bac03]. unifying [ABLU00]. Unigraphics [Eng00]. Union [TCM^{+}00]. Unique [Ano01h]. Unit [Ano02n, Lin03b, Lou05, NS03, NP02, PJ09, HT04]. Uniting [CK05]. Universal [CLCC02, VN03, Vau03b, HHM04]. universally [Yua04]. universe [Ber06].

University [Cha05a, Che05, Gla06, Pet06, Tra00b]. UNIX [Ano01k, SML06, Ano03x, Gab07]. UNIX-Based [Ano01k]. Unleash [Bag02]. Unleashed [DL00, Fle03]. unlimited [Mar01a]. unloading [ZK04a]. unlocking [XSaJ08a]. unmanned [HHM04]. Unobtrusive [Ski07]. unresolved [Ano05e]. unsafe [Win02]. Unstructured [VDPC01, MCLDP01, VDPC03]. unsuccessful [HB09]. Untangling [Ric06b]. Unveils [Ano01h, Ano02m, Ano02t, Kil03a]. up-front [Ano03q]. Update [Ano00n, PM01b, TEM^{+}01, TCM^{+}00, Ano04y, BH02c, GJ09, VDPC03]. updated [Ano02l]. Updates [Ano00n, Ano01h, Ano01i, Ano01j, Ano01l, Ano01m, Ano01n, Ano01o, Ano02m, Ano02o, Ano03-35, SHM09]. Upgrade [MD00, TT08]. upgraded [Ano03-30]. Upgrades [Ano01m, Ano02m, Ano02n, Ano02o, Ano02p, Ano02q, Ano02s, Ano03-37, Ano03-38, Ano03-39, Ano03-40, Ano03-35, Ano03-36, Ano05c]. upgrading [AV05]. Uploaded [BL02a]. Upon [TOG^{+}05]. ups [GMM09]. Upstarts [Ano03n, Coc02]. US-based [Ano03n]. USA [ACM00b, ACM00c, ACM01a, ACM05, Ano01g, Ano02i, AGG02, Gho01, IEE02a, NIS00, USE00c, USE00b, USE00a, USE01c, USE01a, USE02]. usage [BBA08]. USB [Ano03-37]. Use [Bar01d, CN03b, CK05, DKTE04, DFL00, Hac01, HKHK03, ISO05, Jen02b, KWK03, Nat00, Rob04b, Sch03b, Wan04, Way05, Win01, vD04, Ano05b, BKŁ01, GCF^{+}01, Lex02, MJ00, OPS^{+}02, Zus03]. Used [CCW02]. Useful [Pet03, Ano03h, Yua04]. USENIX [ACM05, Jac04b]. User [Ano00j, Bar00c, Gut00, MCLDP01, MCLC02, Rei00a, Ros00, Ano03l, DSCU01, Kon03]. Users [SBH^{+}04, TS01, Ano04w, YAA07]. Using [AG03a, AG03b, ACL03, Ano03-49, Ano03-50, Ano08, ABH^{+}00, AM02, BD03a, BP01b, BL02a, BBHL01, Dd01b, Boo00, BB03, BL02b, BGH^{+}07, Cas02, CH02, CQ05, CKV^{+}02, CN03a, CL03b, CK05, CGRR04, CF04b, Cor00, CLZ06, Dar01b, DeP03a, DTD04, Dmi04, DH04b, EH04, ES05a, ES05b, Fel04, FS03a, FS03b, GH03, GHH01, Gso00, GSW00, Hag00a, HD01, Hei03b, HJF06, HTY^{+}03, HM02, Hun03b, ISO08, IKKW01, JMS02, JBMP03, JKKL04, KM04a, KM04b, KMSL03, KK04b, KY03a, KKJY04, KW01b, KX04, LH03a, Les03, LH03b, LJN^{+}00, Lia00c, LS03, LAT04, Lin03a, LZZ03, Liu08, LHS04b, LS04b, Lut03a, MVM07, MP05, McG04, MKF06, NLFA02, NW03, NIEH04, OS02, PKF03, PL01b, Par00, PV04, PH03, PHBM05, PR03, PCC00, vdPE02, PQVR^{+}01, Pra08, PS03, Rao00a, Rao00b, Rao00c]. Using [Rao00d, Rao00e, Rao00f, Rao01a, Rao01b, RE01, RT02, Rob03, RJFG03, RCdBL02, RW03b, SGV04, San00, ST04, SB00, SSS02, SP03, SSL02, Swa07, TSL^{+}04, TP01, TJ00, Vor01, Wan02a, WVE^{+}00, WS01c, Whi03b, WN05, WSP02, WHKS01, YWZ03, YHL01, Yus04, Ano03k, Ano03-30, Ano03-42, Ano05q, AW00, Atk00, BKH02, Bar02a, BB01, BH04c, BI07, BJ04, BGED04, CWWS03, Car06, CO06, CHL07, CGS^{+}03, Die01, DSCU01, DUK02, DW07, DJ01, ET07, EF02, Eff00, Eng04, ER09, Gag02, Gar09, GEG07, GV02b, Har00d, HP00, Hef07, HIBP04, JFH00, Jia00, JJ02a, JCOP07, JKJ05, Juo07, KMR02, KCF01, Kim02, KTV^{+}04, Knu01a, Kon04, KM04c, Lad01, LP05, Lan05a, LAHC06, LDB^{+}03, LYC02, LC05, LH08a, LPH02, LCHY03, LHFL07, LS08c, MS00a, Mai03, MSR09, MR00a, MAJC03, Mls04]. using [MF03, ML00, Nik03, NH02, Och09b, OJJ00, Oes01, OOOiM05, PWC00, RH07, Ril02, Ril03, Rob00b, Rod01, RVZ04, RMR01, SBAD01, SCB09, SY04, SMS00, ST00a, Soj03b, TA04, Uni03, Utt06, VP05, WF04, Wat02, Wei02a, Wic03, Wil05, Wu05, Wut00, XM06, Yah01, YL03, YAA07, ZXNH02, ZFK04, ZAVT03]. Utah [ACM01a]. Utility [Ano04-37, FBR^{+}03, Fal00a, Fal00b, PSZ^{+}07]. Utilization [KW02, SSA03]. Utilizing [DL02, KKN00]. utopia [Lan05a, Ano02p]. Utopia-LVDS [Ano02p].

v [Saf02, ZP03]. v.5.7 [Ano00i]. v1.3 [Ano00j]. v1.4.0 [Sun02]. V15 [Eng00]. v4.0 [Ano00k]. v5.0 [Ano00j]. V8 [Ano03-40]. Vacuum [Ano02r]. validating [TZ01]. Validation [Ano02t, Pre03, NSS^{+}05, SSB01]. validator [NP07]. Value [Ros02b, BNK^{+}07, WCK^{+}07, ZJ03]. value-added [ZJ03]. valued [Yah01]. Vancouver [LL08a]. Vanward [Ano05p]. variable [Lan04, Oi05, Oi08]. Variables [HS00b, vON02a, Whi03b, vON02b]. Variant [IV06, IV07, CCKP06, Win02]. variation [ET05]. variety [GKM01]. variogram [Fau02]. VB [GS05a, Sur04b]. VCluster [ZLG08]. VCOM [Ano00j]. vector [HJvdB01]. ved [HJL00]. VEE [ACM05]. vehicle [HHM04]. vehicles [HHM04]. Velocity [For04b]. Vendor [Ano03-43]. Verifiable [HOP04, WHBS01, MGM^{+}06]. Verification [ÁMdBdRS02, Ano01i, BDT04, BCDdS02, BFG03, Bec01c, CMR05, DRV02, FC01, GPF05, HR04b, HJ00, Hui02, Jac01c, JKW03, JP04, Kle05b, KK05, Ler01f, Ler01e, Ler03, LM04, Mos05b, Nip03, PV04, RM04, Ros03, Rot05, SS00a, Str02, ZW08, vdBJP01, Aki02, Ano02v, ABF03, BDLM04, BDL^{+}08, Bod04, CR07, Cog03, Cog04, DP08, DH00, FYD^{+}08, FC00, GPF08, HJvdB01, KPH^{+}09, Ler02, NE04, Qia00, SSB01, TM08, Wil02, YKB02, ZKR08, dH05, BHS07]. Verified [KW03, Kle05b, Nip01, Ste04, OOM^{+}07]. Verifier [BBDT02, Ber01c, Cas02, FM03, SS03, BCR03b]. Verifiers [Nip01]. verifieur [BCR03b]. Verify [ACL03, CK05]. Verifying [BBA08, BJvdB02, GPS03, RWH01, Yah01, LSW07]. Verlag [Pap05].

Versatile [GCEO05, Yua04]. Version [Ano00i, Ano00m, Ano02p, Fre04, Goo03b, HL04, KS09, SG00, Ano00k, Ano02l, SM01d]. Versioning [MFSL02]. versions [SM01d]. Versus [Ead01, Ano04l, Hor00a, Hor00b, Ras03, SCEG08, VED06]. Very [Pet03, SSB03]. Via [JPJ05, CLM^{+}07, DJ00, DJ02, GPF08, Hor00c, HJ00, KSK04b, LM04, Mor02, NR05, PH00a, TSDNP02, ZJ03]. viability [MFRW07]. Video [Dei08, Edw00, Pau03, Pew00, Ste08b, SFM^{+}07]. Video-Training [Ste08b]. view [PHM^{+}01, SSGS01]. viewed [Fle01]. Viewer [Ano00n, CE01, RCdBL02]. viewers [CH06, CHJB07]. ViewML [Ano00j]. Viewpoints [SLB^{+}02]. Views [Bar00a, Bar01a, Bar01b, Bar01c, Coc02, BH04c]. Viosoft [Ano01n]. Virkus [Kuc06]. ’Virtual [DMKN02, ACM05, Ano00a, Ano01b, Ano01g, Ano02b, BDJdS02, BHDS09, BD01a, BP01d, BP03b, Caa00, CW03a, CF00, CT03, Che03a, CiLH01, CF02, Cra06, DHPW01, DEK^{+}03, DCA04, DLS^{+}01, FFB^{+}00, FK03, FP03, G^{+}01, GGG03, GM00, HM01a, HWB03, HB08, Ive03a, JR02, JDJ^{+}06, JJ02b, Juo07, LMG00, LMG01, MSR09, Men03, MLG^{+}02b, MP01c, vON02a, Oi05, Oi06, PRB07, Ran02, RB01, SMK02, SD01a, SH04a, SMES01, SSB03, SCEG08, Shi03a, SM01c, Siv04, SSB01, SHB^{+}03, SBA01, SM02b, Sur01, USE01c, USE01b, USE02, VL00, Vog03, WWMG06, ZS01a, vD00, vLSM01, vON02b, AAB^{+}00, AAB^{+}05, AFT01a, ABC^{+}07, ANH00, CvE00, CH08, DGMY06, Die01, DBC^{+}00, EGD03, EGKP02, GEVZ09b, GCARPC^{+}01, GPW03, GBCW00, HL02b, JK00, KN06, LYK^{+}00, MSG01, MS00b, Oi08, PV08, RHR02, Req03, SHR^{+}00]. virtual [TGCF08, VED07, WK08a, WK08b, WK08c, YME05, YTY00, Cza00, VED06]. Virtualization [Ano03-41]. virtualized [PSZ^{+}07]. Virus [Ano00k]. VisAD [HRE^{+}02, HRE^{+}05]. visibility [CHUB08]. visible [Mur07]. VisiBroker [NRV00, P^{+}98]. VisiComp [Ano02n]. vision [WM00b]. visitors [Car06]. VistaSource [Ano00j]. Visual [Ano00i, Ano01l, Ano03-50, Ano04-38, Ano05q, Bel02, GST05, Lia00b, MD00, PSW07, Pil04, RCdBL02, Ano04q, Fei07, Mur00, Pas04, RM07a, SRW^{+}00, Ano01i, Ano01m, Ano01o, Ano02r, Ano04f, Gil00a, Goo03b, HM02, OBr05]. VisualAge [Ano02a, Ano02w, SM01d]. Visualisation [GCEO05, Ibb02]. Visualisierung [Ano04c]. Visualization [Ano01h, Ano01o, Ano02r, ACR01, BL04, Bus02a, Cal02, CE01, DH04b, EvG02, HRE^{+}02, HRE^{+}05, HJF06, HS02b, IKKM03, MB03, Meh02, OS02, ZCQS04, ZK04b, Ano04c, Bus02b, CWWS03, EVS07, FMA02, GV05, GP05, GJ04, HBX^{+}04, NK06, NHY^{+}04, NR05, Rei05, Sal04, SML06, SK08, SD04]. visualizations [HCMM00, HCB04a, KB04b]. Visualize [MH00a, PFJ05, SML06]. Visualizing [DS00b, Fry08, DJM^{+}02, Rei03, Ano01d, CMS05, FL04, TZ01]. Vital [Bar00a, Kro00b]. VLaTTe [KMEA04]. VLIW [KMEA04]. VLSI [PGM^{+}05]. VM [Ano01b, Ano03-37, Cav02a, IN09, LYK^{+}00, Lia03b, SHM09, TABP07]. VM-centric [SHM09]. Vmgen [EGKP02]. VMware [Ano03-37, Ano03-41]. Voice [Lut03b]. VoiceGenie [Ano02r, Ano03-35]. VoiceXML [Ano02r, Ano03-35]. VoIP [Ano00m, Ano03-39]. vol [McL02a]. Volume [Bul00, Gea00, HC00, HC02, HC03]. Volumes [SGV04]. volumetric [Woo03]. Voronoi [IKKM03]. Vorteil [Lex02]. VOTable [KKK04]. Voting [CK05]. Voyage [Coc02]. VR [MD00]. VRML [AL04b, Ano04-34, CN03a, Die01, LLK03, MJ00, SY04]. VRML-JAVA [Ano04-34]. vs [AHN02, Bri05, GJ09, Lam03, PG03b, SKP^{+}02, VZGE07]. VSIPL [ASS^{+}05]. VTK [SML06]. Vulnerabilities [VMMF00]. Vulnerability [RDW^{+}07].

Vulnerability-driven [RDW^{+}07]. Vvedenie [Saf02]. VXA [Ano00h].

W [Ano01a]. Waba [Wil01a]. wall [ZSZ^{+}09]. Walls [CP04, CP01]. Want [LRO02, Ano04w, Hoh03]. wants [Ano03n, Ano04-27]. WAP [YHL04]. WAP-Enabled [YHL04]. WAPPEN [Kag09]. Warehousing [Lut03a]. Wari [Sco03]. Warp [BNO03]. Warps [Wil01b]. Was [Vel01, PPJ03, San04a]. waste [Lex02]. water [PFJ05]. Waterloo [Ano01n]. watermarking [MCHN05]. WAV [Li03]. Wave [HKHK03, Leh02, Ano03-51]. Way [Kic04, Ano03k, Bea05, CC02, CSFS00, DM07, Tre03]. ways [Urb09]. Wcomp [TCF^{+}03]. Weakest [Jac04a, CFS09]. weakly [MBS^{+}08]. Wearable [TCF^{+}03]. Weathering [EBG^{+}05]. Weaving [AF02, BF04]. Web [Bro02a, Cal00a, DHMT00, HJF06, Lut00, Lut03b, Mar05, SO02, Uni01, DFW04, Gar09, GP05, HJL00, HF06, Pan09, TPF^{+}09, XP04, ABM^{+}03, AL04b, Ano00n, Ano01h, Ano01i, Ano01m, Ano01o, Ano02q, Ano02s, Ano02t, Ano03f, Ano03w, Ano03-49, Ano04n, Ano04-27, Ano04-39, Ano05o, AM02, AOMC07, Atk00, Bar02a, Ben00c, Ber05b, BD04, BDFL04, BGadH06, BJ04, Bru05c, Cer02, CJ02, CCW02, CW03b, CLM^{+}07, CLM^{+}09, CMS03b, CBD01, CL03b, Cox01b, DLL03, DV07, DK02, Eng00, Est01, Est02, FK00, For04b, Fox03a, FRMW04, Gab07, GAG06, GV05, GW00, Gou06, HECR00, HHKS03, HB01, Ham07, Har00d, HL04, HP02, Hig03, Hou00, HD03c, II04b, JFH00, JSSM04, JHJX04, JKH^{+}04, Kag09, Kan02, KL07, KMSB08, KR03, KS04, Kro00a, Kum04, Kun02, KX04, Lai03, Lan05a, LL01a]. Web [Lee03, LKL^{+}03, LJ07, LAT04, LHS04a, Lot02, Lut03a, Lut03b, MMN09, MTSM03, Mur00, NS01a, NM02, PPJ03, Pas04, Pew00, Pip03, PWC00, Roc01, RB04, RKK03, RS00b, SL06, SO02, SSS02, SM03b, SW06, Tam00, Tha00, Tha06, Tho03, TAW03, Top03, Tre03, WBS01, Wal03b, Wan04, Way05, Wea00, WL04, YDWL04, YHL01, Zen02, Cul00]. Web-Based [HJF06, GP05, AL04b, Ano01h, Ano01o, Ben00c, CBD01, DK02, Kum04, Kun02, LL01a, RKK03, YHL01, BD04, BJ04, CW03b, Est01, GV05, GW00, Ham07, JFH00, Kag09]. Web-centric [DV07]. Web-enabled [RB04]. Web-scale [KMSB08]. Web-Service [ABM^{+}03, Ano04-27]. Web/Java [HL04, JHJX04, YDWL04]. Web3D [CN03a]. WebEQ [Kun02]. WebGIS [HD03b, RYD^{+}03]. WebLogic [MC04, Nyb02]. webMethods [Ano02l]. Webserver [Ano03e]. Websim99 [FCW01, PSS01, SM01a]. Website [AF02, Tay02]. WebSphere [Bro02b, W^{+}04, Yus04]. WebWork [WACBL03]. WebWorks [For04b]. weekend [SC01a]. weight [HB08]. WEKA [MR06]. well [Ano04-29]. well-priced [Ano04-29]. Wendy [Ano08]. Westbridge [Ano02s]. where [Ano05n]. whether [Mer04]. Which [JPJ05, Ano02l, Ano03n, Ano04g]. While [Ano05c]. white [Ano00i]. Whiteboard [WVE^{+}00]. whitebox [GKL08]. Whiteoak [GM08]. whole [BK05b]. Wicked [Eub05]. Wide [Lot02, NS01a, PWC00]. Wilcox [Fox01b]. wildcards [CV08]. WildPackets [Ano02m]. Wiley [Ano04e]. Will [Ano03-52, Ano04k, Ano04-27, Rei00b, Rei00c]. Willi [Pap05]. Willi-Hans [Pap05]. William [Ano00b]. Win32 [Ano00j, Bec01b]. WinDK [Ano00m]. window [Rem01]. Windows [Ano02q, Ano03z, SML06, Ano00n, Ano01h, Ano01j, Ano01o, Ano02n, Ano04-32, Joh03, Kro00a, Kro00b, LHFL07, Lin01, Tim03, Way03]. Winners [Bar01a]. Wins [Bar00a]. Wire [Lia03b]. Wired [DHR^{+}01, JKKL04]. Wireless [Ano01c, Ano01i, Ano01j, Ano01m, Ano01o, Ano02m, Ano02o, Ano02t, Bar03a, Cha05a, CCC^{+}04, CD03, Eng00, HAL02c, JKKL04, Knu01b, Kuc06, Lea00b, LCZ04, Mah02, Mah04b, Pir02, SRJS08, Tre02b, Tui04, Yan03, CCK^{+}08, GW08, KM04c, RTVH01, Vir05, Whi03a, Zhu04, Ano01j]. Wirth [BGP00]. wishes [HG07b]. Withdraws [Lea00b]. Within [BP05, WP04, GKW04, KM02, Ric00]. Without [HM01b, KKO02, Ano02e, Ano02f, Ano04v, BST00, BAL^{+}01, LAHC06]. wizard [Est02]. Wizards [Ano03-40]. WMPI [SMS00]. Wood [Ran03]. Woods [Cal00a]. word [Coo05]. WordMage [Ano00i]. WordNet [TMF05]. Work [Mls04, Pau01, Rao02, RVZ04, Yan03, Bar09, Gun01, MD06]. workarounds [D^{+}00]. Workbench [FGLS04, MSK09, Ano05o]. Workbook [Bro02b, Nyb02, Met02]. Worker [KSC^{+}00]. Workflow [JHJX04, WS01a, YDWL04, vLH05, SJ01, Sha01, SGW01]. Working [Fel04, SNO^{+}07, SH06]. Workload [IEE02b]. Workloads [DS09, DH04b, GBED04, SSGS01]. Works [MKS^{+}03, MH09, San04a]. Workshop [CCFG00, GDC^{+}04, GAR04, GRR05, HR04b, IEE02b, ACM01a, AJ01a, BZ05, GAR03]. Workshops [SY^{+}05]. Workspace [WWSL02]. workstations [TDB00]. World [Ano00j, Gos00a, Hoh03, HM01b, McL01b, PL03, SH06, SY04, Lot02, NS01a, PWC00]. Worlds [FP03, OBr05, Die01]. Worst [CCM05, HWB03]. Worst-Case [HWB03]. Would [Pau03]. Wrapper [LRSW00, FCHE02]. Wrapping [LRSW00, LRW01]. Write [Iva02, Jen00a, LH02, WA04, Ano03-44, Lan04, Wil04b]. write/run [Ano03-44]. Writer [KKK04]. Writing [Aus00, Feu02, Mam01, Men00, DM07]. written [Ano03h, KK04a, MSG01, MLVB05, TETPQ08, TZ01]. Wrong [SPS^{+}02]. WSDL [Cer02]. WSG [Gar09]. WWC [IEE02b]. WWC-5 [IEE02b]. WWW [CE01, Ibb02].

X [Ano00j, AA02a, Ano02g, Ive03b, Uni02]. X-Link [AA02a]. X-Ray [Uni02, Ano02g]. X-Win32 [Ano00j]. X.509 [SJ05]. x86 [OKN04]. Xanthi [SBH^{+}04]. XAWare [Ano02r]. XDK [Ano00n]. XDoclet [NP03, PL03, WRO04, WACBL03]. xenoliths [INM05]. XHTML [Lad01]. Xilinx [Ano02p, Ano02s, Ano03-38, Ano03-40]. XMem [WK08d]. XMI [GDB02]. XML [Cha05a, Hei01, SBH^{+}04, TEM^{+}01, Ahm01, All03, AL04b, Ano01k, Ano01m, Ano02o, Ano02q, Ano02s, Ano02t, Ano03-34, Bar01b, Boo00, BK03, Bru04c, BFMT00, BK01b, Bur01b, Cer02, CLCC02, CQ05, CZ01, CKM04, CL03b, Cle01a, Cle01b, DS00a, DSCU01, Dwe00a, Dwe00b, EF02, Fal00a, Fal00b, Fel04, Gös03, Gri02a, GDB02, Har02, Har03, Hei03a, HNZS03, KMS04, Kro00a, Lad01, LJ07, LCZ04, Lin03a, LZZ03, Mam01, McL00, McL01a, McL01b, McL02b, McL06b, McL07, MF01b, Roc01, RJFG03, SGW01, SG02, Sin00, SFP03, Tam00, Tre02c, WL04, Woo04, XP04, YLM^{+}05, Zhu04, dGNv04]. XML-Based [CLCC02, Gös03, HNZS03, Kro00a, Mam01]. XML-enabled [SGW01]. XML-Oriented [Ano02t]. XML-RPC [All03, Cer02]. XML/Java [CQ05]. XMLC [You02]. XQJ [EM04, VLMO09]. XQL [BK01b]. XQuery [EM04, VLMO09]. XRTJ [HWB04]. XScale [Ano01m, CMP^{+}07]. XSLT [Fox01c, Bur01a, Bur01b, DBH04, Fox02, NP03, Roc01, Tho03]. XSQL [Tho03]. XTREM [CMP^{+}07].

Y2K [Lea00b]. Yama [MJ06]. Year [DHRH05, AWS^{+}09, CLP06, Edm09, Ras00, Rio02, XSD07]. Years [Lut03a, Eic05, Kic04]. YesSoftware [Ano01l, Ano02q]. yield [Ano04k, WK09].

Yoix(R) [DM07]. Yorick [Pap05]. York [Ano01a, NIS00]. you’re [Mer04]. yourself [AK00, CL03a, WMM04].

Z [SH04b, WCK^{+}07]. z10 [SKC09]. zA-APs [WCK^{+}07]. ZapMedia [Mar01b]. Zap-Station [Mar01b]. ZapStation/Harman [Mar01b]. Zaurus [HKS02]. Zayante [Ano01j]. Zhuk [Cha05a]. zIIPs [WCK^{+}07]. Zondigo [Ano01o]. zum [Wol03a, Zus03]. zur [Ano05a, DHMT00]. Zuse [BHP^{+}01, Roj00].

References

Antoniu:2001:HSC

[A^{+}01] Gabriel Antoniu et al. The Hyperion system: Compiling multithreaded Java bytecode for distributed execution. Parallel Computing, 27(10):1279–1297, September 2001. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                                      tronic). URL http://www.
                                                      elsevier.com/gej-ng/10/
                                                      35/21/47/40/27/abstract.
                                                      html;   http://www.elsevier.
                                                      nl/gej-ng/10/35/21/47/40/
                                                      27/article.pdf.

Alvarez:2002:AJT

[AA02a] A. Alvarez and Y. Amghar. Applying JAVA-Triggers for X-Link management in the industrial framework. Simulation series, 34(1):131–136, 2002. CODEN SMCPAX. ISSN 0735-9276.

Anderson:2002:EJC

[AA02b] Gail Anderson and Paul Anderson. Enterprise JavaBeans Component Architecture: Designing and Coding Enterprise Applications. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2002. ISBN 0-13-035571-2. xix + 435 pp. LCCN QA76.73.J38 A47 2002. US$59.99. URL http:/

/www.phptr.com/ptrbooks/
ptr_0130355712.html;   http:
//www.sun.com/books/catalog/
anderson/.

AlAli:2004:JBH

[AA04] A. R. AlAli and M. AlRousan.

Java based home automation system. IEEE Transactions on Consumer Electronics, 50 (2):498–504, 2004. CODEN ITCEDA. ISSN 0098-3063.

Assaf:2004:IEC

[AAA^{+}04] M. H. Assaf, R. S. Abielmona, P. Abolghasem, S. R. Das, E. M. Petriu, V. Groza, and M. Sahinoglu. Implementation of embedded coresbased digital devices in JBits Java simulation environment. Lecture Notes in Computer Science, 3356:315–325, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abi-Antoun:2005:ISD

[AAAG^{+}05] Marwan Abi-Antoun, Jonathan Aldrich, David Garlan, Bradley Schmerl, Nagi Nahas, and Tony Tseng. Improving system dependability by enforcing architectural intent. ACM SIGSOFT Software Engineering Notes, 30(4):1–7, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Alpern:2000:JAV

[AAB^{+}00] B. Alpern, C. R. Attanasio, J. J. Barton, M. G. Burke, P. Cheng, J.-D. Choi, A. Cocchi, S. J. Fink, D. Grove, M. Hind, S. F. Hummel, D. Lieber, V. Litvinov, M. F. Mergen, T. Ngo, J. R. Russell, V. Sarkar, M. J. Serrano, J. C. Shepherd, S. E. Smith, V. C. Sreedhar, H. Srinivasan, and J. Whaley. The Jalapeño virtual machine. IBM Systems Journal, 39(1): 211–238, ???? 2000. CO-DEN IBMSA7. ISSN 0018-

8670. URL http://www.
almaden.ibm.com/journal/
sj/391/alpern.html.

Alpern:2005:PVE

[AAB^{+}05] Bowen Alpern, Joshua Auerbach, Vasanth Bala, Thomas Frauenhofer, Todd Mummert, and Michael Pigott. PDS: a virtual execution environment for software deployment. In ACM [ACM05], pages 175– 185. ISBN 1-59593-047-7. LCCN QA76.9.V5 I575 2005. URL http:/

/www.loc.gov/catdir/toc/

fy0611/2006530661.html.

ACM order number 548059.

Ancona:2001:ETF

[AAD^{+}01] Davide Ancona, Christopher Anderson, Ferruccio Damiani, Sophia Drossopoulou, Paola Giannini, and Elena Zucca. An effective translation of Fickle into Java. Lecture Notes in Computer Science, 2202:215–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2202/22020215.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2202/22020215.
pdf.

Ancona:2007:PCT

[AAD^{+}07] D. Ancona, C. Anderson, F. Damiani, S. Drossopoulou, P. Giannini, and E. Zucca. A provenly correct translation of Fickle into Java. ACM Transactions on Programming Languages and Systems, 29(2):13:1–13:67, April 2007. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Aaronson:2006:PPC

[Aar06] L. Aaronson. Parallel processor [community service]. IEEE Spectrum, 43(5):61–64, May 2006. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Armbruster:2007:RTJ

[ABC^{+}07] Austin Armbruster, Jason Baker, Antonio Cunei, Chapman Flack, David Holmes, Filip Pizlo, Edward Pla, Marek Prochazka, and Jan Vitek. A real-time Java virtual machine with applications in avionics. ACM Transactions on Embedded Computing Systems, 7(1):5:1–5:49, December 2007. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Avvenuti:2003:JBV

[ABF03] Marco Avvenuti, Cinzia Bernardeschi, and Nicoletta De Francesco. Java bytecode verification for secure information flow. ACM SIG-PLAN Notices, 38(12):20–27, December 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Alt:2002:ADP

[ABG02] M. Alt, H. Bischof, and S. Gorlatch. Algorithm design and performance prediction in a Java-based Grid system with skeletons. Lecture Notes in Computer Science, 2400:899–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2400/24000899.
htm;   http://link.springer-
ny.com/link/service/series/

0558/papers/2400/24000899.
pdf.

Auerbach:2008:FTG

[ABG^{+}08] Joshua Auerbach, David F.

Bacon, Rachid Guerraoui, Jesper Honig Spring, and Jan Vitek. Flexible task graphs: a unified restricted thread programming model for Java. ACM SIGPLAN Notices, 43 (7):1–11, July 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Antoniu:2000:IJC

[ABH^{+}00] Gabriel Antoniu, Luc Bougé, Philip Hatcher, Mark Mac-Beth, Keith McGuigan, and Raymond Namyst. Implementing Java consistency using a generic, multithreaded DSM runtime system. Lecture Notes in Computer Science, 1800:560–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/1800/18000560.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/1800/18000560.
pdf.

Antoniu:2001:CMJ

[ABH^{+}01] Gabriel Antoniu, Luc Bougé, Philip Hatcher, Mark Mac-Beth, Keith McGuigan, and Raymond Namyst. Compiling multithreaded Java bytecode for distributed execution (distinguished paper). Lecture Notes in Computer Science, 1900:1039–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/1900/19001039.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/1900/19001039.
pdf.

Auerbach:2007:JTF

[ABI^{+}07] Joshua Auerbach, David F. Bacon, Daniel T. Iercan, Christoph M. Kirsch, V. T. Rajan, Harald Roeck, and Rainer Trummer. Java takes flight: time-portable realtime programming with exotasks. ACM SIGPLAN Notices, 42(7):51–62, July 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Auerbach:2009:LLT

[ABI^{+}09] Joshua Auerbach, David F. Bacon, Daniel Iercan, Christoph M. Kirsch, V. T. Rajan, Harald Röck, and Rainer Trummer. Low-latency time-portable real-time programming with Exotasks. ACM Transactions on Embedded Computing Systems, 8(2):15:1–15:??, January 2009. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Adelmann:2007:IFF

[ABL07] Robert Adelmann, Tobias Bischoff, and Tobias Lauer. IDEA: a framework for the fast creation of interactive animations by pen sketching. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39 (3):291–295, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Appert:2008:SAS

[ABL08] C. Appert and M. Beaudouin-Lafon. SwingStates: adding state machines to Java and the Swing toolkit. Software —Practice and Experience, 38(11):1149–1182, September ??, 2008. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Alexander:2000:UAP

[ABLU00] W. P. Alexander, R. F. Berry, F. E. Levine, and R. J. Urquhart. A unifying approach to performance analysis in the Java environment. IBM Systems Journal, 39(1): 118–134, ???? 2000. CO-DEN IBMSA7. ISSN 0018-

8670. URL http://www.
almaden.ibm.com/journal/
sj/391/alexander.html.

Alvarez:2003:JCT

[ABM^{+}03] P. Alvarez, J. A. Banares, P. R. MuroMedrano, J. Nogueras, and F. J. Zarazaga. A Java coordination tool for Web-service architectures: The locationbased service context. Lecture Notes in Computer Science, 2604:1–14, 2003. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alexander:2000:CJP

[ABV00] Roger T. Alexander, James M. Bieman, and John Viega. Coping with Java programming stress. Computer, 33 (4):30–38, April 2000. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//dlib.computer.org/co/
books/co2000/pdf/r4030.
pdf;   http://www.computer.
org/computer/co2000/r4030abs.
htm.

Allan:2001:CSA

[AC01] V. H. Allan and X. Chen. Convert2Java: semi-automatic conversion of C to Java. Future Generation Computer Systems, 18(2):201– 211, October 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/60/31/28/abstract.
html.

Allen:2006:SIG

[AC06] Eric E. Allen and Robert Cartwright. Safe instantiation in Generic Java. Science of Computer Programming, 59(1–2):26–37, January 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Attali:2001:IDE

[ACC^{+}01] Isabelle Attali, Denis Caromel, Carine Courbis, Ludovic Henrio, and Henrik Nilsson. An integrated development environment for Java Card. Computer Networks (Amsterdam, Netherlands: 1999), 36(4):391–405, July 16, 2001. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http:/

/www.elsevier.nl/gej-ng/
10/15/22/61/28/27/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/15/22/61/28/
27/article.pdf.

Alia:2004:MFP

[ACD^{+}04] M. Alia, S. ChassandeBarrioz, P. Dechamboux, C. Hamon, and A. Lefebvre. A middleware framework for the persistence and querying of Java objects. Lecture Notes in Computer Science, 3086:291– 315, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alpern:2001:EIJ

[ACFG01] Bowen Alpern, Anthony Cocchi, Stephen Fink, and David Grove. Efficient implementation of Java interfaces: Invokeinterface considered harmless. ACM SIG-PLAN Notices, 36(11):108– 124, November 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Proceedings of the 2001 ACM SIGPLAN Conference on Object Oriented Programming, Systems, Languages and Applications (OOPSLA’01).

Alpern:2001:EDJ

[ACGL01] Bowen Alpern, Anthony Cocchi, David Grove, and Derek Lieber. Efficient dispatch of Java interface methods. Lecture Notes in Computer Science, 2110:621–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2110/21100621.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2110/21100621.
pdf.

Avgustinov:2005:OA

[ACH^{+}05] Pavel Avgustinov, Aske Simon Christensen, Laurie Hendren, Sascha Kuzins, Jennifer Lhoták, Ondřej Lhoták, Oege de Moor, Damien Sereni, Ganesh Sittampalam, and Julian Tibble. Optimising aspectJ. ACM SIGPLAN Notices, 40(6):117–128, June 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Andronick:2003:UCV

[ACL03] J. Andronick, B. Chetali, and O. Ly. Using Coq to verify Java card applet isolation properties. Lecture Notes in Computer Science, 2758:335– 351, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

ACM:2000:CPI

[ACM00a] ACM, editor. Conference proceedings of the 2000 International Conference on Supercomputing: Santa Fe, New Mexico, May 8–11, 2000. ACM Press, New York, NY 10036, USA, 2000. ISBN 1-58113-270-0. LCCN QA76.88 .I573 2000. URL

http://www.acm.org/pubs/
contents/proceedings/supercomputing/
335231.

ACM:2000:PAS

[ACM00b] ACM, editor. Proceedings ACM SIGMETRICS ’2000: International Conference on Measurement and Modeling of Computer Systems, June 17–21, 2000, Santa Clara, CA, USA: Proceedings of the International Conference on Measurements and Modeling of Computer Systems, volume 28(1) of ACM SIGMETRICS Perform. Eval. Rev. ACM Press, New York, NY 10036, USA, 2000. ISBN 1-58113-194-1. LCCN QA1 .A854724.

ACM:2000:SHP

[ACM00c]  ACM, editor.  SC2000:

High Performance Networking and Computing. Dallas Convention Center, Dallas, TX, USA, November 4–10, 2000. ACM Press and IEEE Computer Society Press, New York, NY 10036, USA and 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA,

                                        2000. URL http://www.
                                        sc2000.org/proceedings/
                                        info/fp.pdf.

ACM:2001:ASS

[ACM01a] ACM, editor. ACM SIGPLAN– SIGSOFT workshop on Program analysis for software tools and engineering: June 18–19, 2001, Snowbird, Utah, USA: PASTE’01. ACM Press, New York, NY 10036, USA, 2001. ISBN 1-58113-413-4. LCCN QA76.758 .A345 2001. Supplement to ACM SIG-PLAN Notices.

ACM:2001:PAJ

[ACM01b]  ACM, editor.  Proceed-

ings of the ACM 2001 Java Grande/ISCOPE Conference: Palo Alto, Calif., June 2– 4, 2001. ACM Press, New York, NY 10036, USA, 2001. ISBN 1-58113-359-6. LCCN QA76.9.O35 A26 2001.

ACM:2001:SHP

[ACM01c] ACM, editor. SC2001: High Performance Networking and Computing. Denver, CO, November 10–16, 2001. ACM Press and IEEE Computer Society Press, New York, NY 10036, USA and 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2001. ISBN 1-58113-293-X. LCCN ????

ACM:2001:SPJ

[ACM01d] ACM, editor. SIGMETRICS ’2001/PERFORMANCE ’2001 Joint International Conference on Measurement and Modeling of Computer Systems: proceedings, volume 29(1) of ACM SIGMET-RICS Perform. Eval. Rev. ACM Press, New York, NY 10036, USA, June 2001. ISBN 1-58113-334-0. LCCN QA76.9.E94; Internet. ACM order number 488010.

Alur:2001:CJP

[ACM01e] Deepak Alur, John Crupi, and Dan Malks. Core J2EE Patterns: Best Practices and Design Strategies. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2001. ISBN 0-13-064884-1. xxvi + 459 pp. LCCN QA76.73.J38 A454

2001. URL http://www.sun.
com/books/catalog/crupi/
index.html.

IEEE:2003:PCI

[ACM03a] ACM, editor. Proceedings of the international symposium on code generation and optimization: feedback-directed and runtime optimization. ACM Press, New York, NY 10036, USA, 2003. ISBN 0-7695-1913-X. ACM Order No.: 530033.

ACM:2003:SII

[ACM03b] ACM, editor. SC2003: Igniting Innovation. Phoenix, AZ, November 15–21, 2003. ACM Press and IEEE Computer Society Press, New York, NY 10036, USA and 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2003. ISBN 1-58113-695-1. LCCN ????

ACM:2004:SHP

[ACM04]  ACM, editor.  SC 2004:

High Performance Computing, Networking and Storage: Bridging communities: Proceedings of the IEEE/ACM Supercomputing 2004 Conference, Pittsburgh, PA, November 6–12, 2004. ACM Press and IEEE Computer Society Press, New York, NY 10036, USA and 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2004. ISBN 0-7695-2153-3. LCCN ????

ACM:2005:PFA

[ACM05]  ACM, editor.  Proceed-

ings of the First ACM/ USENIX International Conference on Virtual Execution Environments: VEE ’05: June 11–12, 2005, Chicago, Illinois, USA. ACM Press, New York, NY 10036, USA, 2005. ISBN 1-59593-047-7. LCCN QA76.9.V5 I575 2005. URL http:/

/www.loc.gov/catdir/toc/
fy0611/2006530661.html.

ACM order number 548059.

ACM:2006:PCC

[ACM06] ACM, editor. Proceedings of the 3rd conference on Computing Frontiers, May 3–5, 2006, Ischia, Italy. ACM Press, New York, NY 10036, USA, 2006. ISBN 1-59593-302-6. LCCN ???? ACM order number 104060.

Alur:2005:SIS

[AČMN05] Rajeev Alur, Pavol Černý, P. Madhusudan, and Wonhong Nam. Synthesis of interface specifications for Java classes. ACM SIGPLAN Notices, 40(1):98–109, January 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Aldrich:2002:ARA

[ACN02] Jonathan Aldrich, Craig Chambers, and David Notkin. Architectural reasoning in ArchJava. Lecture Notes in Computer Science, 2374:334– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.
                                        com/link/service/series/
                                        0558/bibs/2374/23740334.
                                        htm;   http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2374/23740334.
                                        pdf.

Attali:2001:GVJ

[ACR01] Isabelle Attali, Denis Caromel, and Marjorie Russo. Graphical visualization of Java objects, threads, and locks. IEEE Distributed Systems Online, 2(1), 2001. ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:

                                        //dsonline.computer.org/
                                        0101/features/att0101_print.
                                        htm.

Allen:2002:DLP

[ACS02] Eric Allen, Robert Cartwright, and Brian Stoler. DrJava: a lightweight pedagogic environment for Java. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1): 137–141, March 2002. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Amandi:2005:JFB

[ACZ05] Analía Amandi, Marcelo Campo, and Alejandro Zunino. JavaLog: a framework-based integration of Java and Prolog for agent-oriented programming. Computer Languages, Systems and Structures, 31 (1):17–33, April 2005. CO-DEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic).

Adamson:2005:QJD

[Ada05] Chris Adamson. QuickTime for Java: a developer’s notebook. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2005. ISBN 0-596-00822-8. xix + 233 pp. LCCN TR899 .A33 2005; TR899; TR899 .A33 2005eb. URL http://

www.oreilly.com/catalog/
9780596008222.

Adams:2006:OJP

[Ada06] Joel C. Adams. OOP and the Janus Principle. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):359– 363, March 2006. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Abraham:2005:ABP

[ÁdBdRS05] Erika Ábrahám, Frank S.

de Boer, Willem-Paul de Roever, and Martin Steffen. An assertion-based proof system for multithreaded Java. Theoretical Computer Science, 331(2–3):251–290, February 25, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Abraham:2008:DPS

[ÁdBdRS08] Erika Ábrahám, Frank S. de Boer, Willem-Paul de Roever, and Martin Steffen. A deductive proof system for multithreaded Java with exceptions. Fundamenta Informaticae, 82(4):391–463, July 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Abraham:2003:IPO

[AddS03a] E. Abraham, F. S. deBoer, W. P. deRoever, and M. Steffen. Inductive proof outlines for monitors in Java. Lecture Notes in Computer Science, 2884:155–169, 2003. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abraham:2003:TSP

[AddS03b] E. Abraham, F. S. deBoer, W. P. deRoever, and M. Steffen. A tool-supported proof system for multithreaded Java. Lecture Notes in Computer Science, 2852:1– 32, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ancona:2005:PBC

[ADDZ05] Davide Ancona, Ferruccio Damiani, Sophia Drossopoulou, and Elena Zucca. Polymorphic bytecode: Compositional compilation for Javalike languages. ACM SIG-PLAN Notices, 40(1):26–37, January 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ahmed:2009:SDR

[ADR09] Amal Ahmed, Derek Dreyer, and Andreas Rossberg. Statedependent representation independence. ACM SIG-PLAN Notices, 44(1):340– 353, January 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Aldinucci:2003:AES

[ADT03] M. Aldinucci, M. Danelutto, and P. Teti. An advanced environment supporting structured parallel programming in Java. Future Generation Computer Systems, 19 (5):611–626, July 2003. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Adams:2006:JAE

[AE06] Cameron Adams and James Edwards. The JavaScript Anthology: 101 Essential Tips, Tricks & Hacks. SitePoint, ????, 2006. ISBN 0-9752402-6-9. 592 (est) pp. US$39.95, CAN $55.95. URL http://

www.oreilly.com/catalog/
0975240269/.

Anderson-Freed:2002:WWP

[AF02] Susan Anderson-Freed. Weaving a Website: programming in HTML, JavaScript, Perl and Java. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2002. ISBN 0-13-028220-0. 712 pp. LCCN QA76.625 .A64 2002. BUY-A-BOOK gift of Michael E. Hayden thanks to Professor Frank Deremer.

Adams:2003:OCD

[AF03] J. Adams and J. Frens. Object centered design for Java: Teaching OOD in CS-1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35 (1):273–277, 2003. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Abadi:2006:TSL

[AFF06] Martin Abadi, Cormac Flanagan, and Stephen N. Freund. Types for safe locking: Static race detection for Java. ACM Transactions on Programming Languages and Systems, 28(2): 207–255, March 2006. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Arnold:2000:AOJ

[AFG^{+}00] Matthew Arnold, Stephen Fink, David Grove, Michael Hind, and Peter F. Sweeney. Adaptive optimization in the Jalapeño JVM. ACM SIG-PLAN Notices, 35(10):47– 65, October 2000. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
proceedings/oops/353171/
p47-arnold/.

Aridor:2000:TOS

[AFT^{+}00] Yariv Aridor, Michael Factor, Avi Teperman, Tamar Eilam, and Assaf Schuster. Transparently obtaining scalability for Java applications on a cluster. Journal of Parallel and Distributed Computing, 60(10):1159–1193, October 2000. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jpdc.
2000.1649;   http://www.
idealibrary.com/links/doi/
10.1006/jpdc.2000.1649/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.
2000.1649/ref.

Aridor:2001:DIV

[AFT01a] Yariv Aridor, Michael Factor, and Avi Teperman. A distributed implementation of a virtual machine for Java. Concurrency and Computation: Practice and Experience, 13(3):221– 244, March 2001. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (elec-

tronic). URL http://www3.
interscience.wiley.com/
cgi-bin/abstract/78003113/
START;   http://www3.interscience.

                                        wiley.com/cgi-bin/fulltext?
                                        ID=78003113&PLACEBO=IE.
                                        pdf.

Aridor:2001:IJC

[AFT01b] Yariv Aridor, Michael Factor, and Avi Teperman. Implementing Java on clusters. Lecture Notes in Computer Science, 2150:722–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.
                                        com/link/service/series/
                                        0558/bibs/2150/21500722.
                                        htm;   http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2150/21500722.
                                        pdf.

Alt:2003:PGS

[AG03a] M. Alt and S. Gorlatch. A prototype Grid system using Java and RMI. Lecture Notes in Computer Science, 2763:401–414, 2003. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alt:2003:USJ

[AG03b] M. Alt and S. Gorlatch. Using skeletons in a Java-based Grid system. Lecture Notes in Computer Science, 2790:742– 749, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alt:2005:AJR

[AG05] Martin Alt and Sergei Gorlatch. Adapting Java RMI for grid computing. Future Generation Computer Systems, 21 (5):699–707, May 2005. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Arnold:2002:JJT

[AGG02] Ken Arnold, Guang R. Gao, and Sudipto Ghosh, editors. Java/Jini technologies and high-performance pervasive computing: 30 July and 1 August, 2002, Boston, [Massachusetts] USA, number 4863 in SPIE proceedings series. SPIE Optical Engineering Press, Bellingham, WA, USA, 2002. ISBN 0-8194-4642-4. LCCN QA76.9.D5 J38 2002.

Arnold:2000:JPL

[AGH00] Ken Arnold, James Gosling, and David Holmes. The Java Programming Language. Addison-Wesley, Reading, MA, USA, third edition, 2000. ISBN 0-201-70433-1 (paperback). xxiv + 595 pp. LCCN QA76.73.J38 A76 2000. US$37.95.

Almquist:2005:ITS

[AGH05a] J. Almquist, I. Gorton, and J. Haack. Integration of a text search engine with a Java messaging service. Lecture Notes in Computer Science, 3437:19–30, 2005. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arnold:2005:JPL

[AGH05b] Ken Arnold, James Gosling, and David (David Colin) Holmes. The Java^{T M} Programming Language. Addison-Wesley, Reading, MA, USA, fourth edition, 2005. ISBN 0-321-34980-6. xxviii + 891 pp. LCCN QA76.73.J38 A76 2005.

Artigas:2000:ALT

[AGMM00] P. V. Artigas, M. Gupta, S. P. Mikiff, and J. E. Moreira. Automatic loop transformations and parallelization for Java. Parallel Processing Letters, 10(2/3): 153–??, September 2000. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic). URL http:

                                        //ejournals.wspc.com.sg/
                                        ppl/10/1002_03/S0129626400000160.
                                        html.

Avetisyan:2001:EJE

[AGS01] A. Avetisyan, S. Gaissaryan, and O. Samovarov. Extension of Java environment by facilities supporting development of SPMD Javaprograms. Lecture Notes in Computer Science, 2127:175– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.
                                        com/link/service/series/
                                        0558/bibs/2127/21270175.
                                        htm;   http://link.springer-
                                        ny.com/link/service/series/

0558/papers/2127/21270175.
pdf.

Aldrich:2004:MISa

[AGST04a] Jonathan Aldrich, David Garlan, Bradley Schmerl, and Tony Tseng. Modeling and implementing software architecture with acme and arch-Java. ACM SIGPLAN Notices, 39(10):6–7, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Aldrich:2004:MISb

[AGST04b] Jonathan Aldrich, David Garlan, Bradley Schmerl, and Tony Tseng. Modeling and implementing software architecture with acme and arch-Java. ACM SIGPLAN Notices, 39(10):156–157, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Allen:2003:SJP

[AH03] Matthew Allen and Susan Horwitz. Slicing Java programs that throw and catch exceptions. ACM SIG-PLAN Notices, pages 284– 294, 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Adelstein:2004:EJL

[AH04a] Tom Adelstein and Sam Hiser. Exploring the JDS Linux desktop. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2004. ISBN 0-596-00752-3 (paperback). xiv + 390 pp. LCCN QA76.76.O63 A353

2004. URL http://www.
oreilly.com/catalog/jds/
index.html.

Araujo:2004:TAC

[AH04b] W. L. F. Araujo and C. M. Hirata. Translating activity cycle diagrams to Java simulation programs. Annual Simulation Symposium, SYMPO-SIUM37:157–166, 2004. CO-DEN ???? ISSN 0272-4715.

Arnold:2001:EIB

[AHKR01] Matthew Arnold, Michael Hsiao, Ulrich Kremer, and Barbara G. Ryder. Exploring the interaction between Java’s implicitly thrown exceptions and instruction scheduling. International Journal of Parallel Programming, 29(2):111–137, April 2001. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

ipsapp009.lwwonline.com/
content/getfile/4773/20/
1/abstract.htm;   http:/
/ipsapp009.lwwonline.com/
content/getfile/4773/20/
1/fulltext.pdf;   http:
//www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=29&
issue=2&spage=111.

Ahmed:2001:PJX

[Ahm01] Kal Ahmed. Professional Java XML. Programmer to programmer. Wrox Press, Chicago, IL, USA, 2001. ISBN 1-86100-401-X. xv + 1159 pp. LCCN QA76.76.H94 P7626 2001.

Alouf:2002:FVC

[AHN02] Sara Alouf, Fabrice Huet, and Philippe Nain. Forwarders vs. centralized server: an evaluation of two approaches for locating mobile agents. ACM SIGMETRICS Perform. Eval. Rev., 30(1): 278–279, June 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Arnold:2002:OFD

[AHR02] Matthew Arnold, Michael Hind, and Barbara G. Ryder. Online feedback-directed optimization of Java. ACM SIG-PLAN Notices, 37(11):111– 129, November 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Aissi:2003:RAW

[Ais03] Selim Aissi. Runtime environment security models. Intel Technology Journal, 7 (1):60–67, February 2003. ISSN 1535-766X. URL

http://developer.intel.
com/technology/itj/2003/
volume07issue01/art05_security/
p01_abstract.htm.

Attali:2001:JSC

[AJ01a] Isabelle Attali and Thomas Jensen, editors. Java on smart cards: programming and security: first international workshop, JavaCard 2000, Cannes, France, September 14, 2000: revised papers, volume 2041 of Lecture Notes in Computer Science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 3-540-42167-X (paperback). LCCN QA76.73.J38 J3635 2000; QA267.A1 L43 no.2041. URL

http://link.springer-ny.
com/link/service/series/
0558/tocs/t2041.htm. Also

available via the World Wide Web.

Attali:2001:SCP

[AJ01b] Isabelle Attali and Thomas Jensen, editors. Smart card programming and security: International Conference on Research in Smart Cards, Esmart 2001, Cannes, France, September19-21, 2001: proceedings, volume 2140 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 3-540-42610-8 (paperback). LCCN TK7895.S62 I54 2001. URL

http://link.springer-ny.
com/link/service/series/
0558/tocs/t2140.htm.

Arato:2004:JPB

[AJB^{+}04] A. Arato, Z. Juhasz, P. Blenkhorn, G. Evans, and G. Evreinov. Java-powered Braille slate talker. Lecture Notes in Computer Science, 3118:506–513, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Al-Jaroodi:2002:OPD

[AJMJS02] Jameela Al-Jaroodi, Nader Mohamed, Hong Jiang, and David R. Swanson. An overview of parallel and distributed Java for heterogeneous systems: Approaches and open issues. Parallel and Distributed Computing Practices, 5(4):??, December 2002. CODEN ???? ISSN 1097-2803.

Al-Jaroodi:2005:JJO

[AJMJS05] Jameela Al-Jaroodi, Nader Mohamed, Hong Jiang, and David Swanson. JOPI: a Java object-passing interface. Concurrency and Computation: Practice and Experience, 17(7–8):775–795, June/July 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Annunziato:2000:STY

[AK00] Jose Annunziato and Stephanie Fesler Kaminaris. Sams teach yourself JavaServer Pages in 24 hours. Howard W. Sams, Indianapolis, IN 46268, USA, 2000. ISBN 0-672-32023-1. xiii + 433 pp. LCCN TK5105.8885.J38 A56 2001. Edited by Rosemarie Graham.

Aleksy:2001:ASB

[AK01] M. Aleksy and A. Korthaus. Access to SAP’s Business Framework from Java-based applications. Concurrency and Computation: Practice and Experience, 13(7): 621–641, June 2001. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (elec-

tronic). URL http://www3.
interscience.wiley.com/
cgi-bin/abstract/83002172/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=83002172&PLACEBO=IE.
pdf.

Axelsen:2009:GPT

[AK09] Eyvind W. Axelsen and Stein Krogdahl. Groovy package templates: supporting reuse and runtime adaption of class hierarchies. ACM SIG-PLAN Notices, 44(12):15–26, December 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Akiyama:2002:MEP

[Aki02] S. Akiyama. A modeling of evolution processes on the Java Class Library and it’s statistical verification. Report of Researches — Nippon Institute of Technology, 32(2): 145–148, 2002. CODEN ???? ISSN 0389-2514.

Alagic:2004:CJT

[AL04a] S. Alagic and J. Logan. Consistency of Java transactions. Lecture Notes in Computer Science, 2921:71–89, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ande:2004:IVJ

[AL04b] R. N. Ande and L. M.

Liebrock. Integrating VRML, JAVA, XML, and HTML in a Web-based tool. Summer Computer Simulation Conference, ??:513–519, 2004. CO-DEN ???? ISSN 0094-7474.

Arthorne:2004:OEF

[AL04c] John Arthorne and Chris Laffra. Official Eclipse 3.0 FAQs. Eclipse series. Addison-Wes-ley, Reading, MA, USA, 2004. ISBN 0-321-26838-5 (paperback). xxxiv + 386 pp. LCCN QA76.76.D47 A776 2004.

Albrecht:2003:TJI

[Alb03] Matt Albrecht. Testing Java interfaces with Junit. Dr. Dobb’s Journal of Software Tools, 28(2):24, 26, 28, February 2003. CODEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/documents/
s=7790/ddj0302c/.

Allison:2000:IJA

[All00a]  Chuck Allison.  import

java.*: Arrays. C/C++ Users Journal, 18(3):46–??, March 2000. CODEN CCU-JEX. ISSN 1075-2838.

Allison:2000:IJB

[All00b]  Chuck Allison.  import

java.*: Basic stream I/O. C/C++ Users Journal, 18 (11):58–??, November 2000. CODEN CCUJEX. ISSN 1075-2838.

Allison:2000:IJC

[All00c]  Chuck Allison.  import

java.*: Collections and algorithms. C/C++ Users Journal, 18(9):76–??, September 2000. CODEN CCUJEX. ISSN 1075-2838.

Allison:2000:IJF

[All00d]  Chuck Allison.  import

java.*: Formatted text and locales. C/C++ Users Journal, 18(7):60–??, July 2000. CODEN CCUJEX. ISSN 1075-2838.

Allison:2000:IJI

[All00e]  Chuck Allison.  import

java.*: Interfaces and inner classes. C/C++ Users Journal, 18(1):??, January 2000. CODEN CCUJEX. ISSN 1075-2838.

Allison:2000:IJS

[All00f]  Chuck Allison.  import

java.*: Strings. C/C++ Users Journal, 18(6):??, June 2000. CODEN CCUJEX. ISSN 1075-2838.

Allman:2003:EXR

[All03] Mark Allman. An evaluation of XML-RPC. ACM SIG-METRICS Perform. Eval.

Rev., 30(4):2–11, March 2003. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ancona:2000:JSE

[ALZ00] Davide Ancona, Giovanni Lagorio, and Elena Zucca. Jam — A smooth extension of Java with mixins. Lecture Notes in Computer Science, 1850:154–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/1850/18500154.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/1850/18500154.
pdf.

Ancona:2001:CCJ

[ALZ01] D. Ancona, G. Lagorio, and E. Zucca. A core calculus for Java exceptions. ACM SIG-PLAN Notices, 36(11):16–30, November 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Proceedings of the 2001 ACM SIGPLAN Conference on Object Oriented Programming, Systems, Languages and Applications (OOPSLA’01).

Ancona:2002:FFJ

[ALZ02] Davide Ancona, Giovanni Lagorio, and Elena Zucca. A formal framework for Java separate compilation. Lecture Notes in Computer Science, 2374:609–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.
                                        com/link/service/series/
                                        0558/bibs/2374/23740609.
                                        htm;   http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2374/23740609.
                                        pdf.

Ancona:2003:JDJ

[ALZ03] Davide Ancona, Giovanni Lagorio, and Elena Zucca. Jam—designing a Java extension with mixins. ACM Transactions on Programming Languages and Systems, 25(5):641–712, September 2003. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Apte:2002:WSJ

[AM02] Naresh Apte and Toral Mehta. Web Services: a Java Developer’s Guide Using E-Speak. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2002. ISBN 0-13-062338-5. 540 (est.) pp. LCCN QA76.73.J38 A68 2001. US$44.99. URL http:/

                                        /www.phptr.com/ptrbooks/
                                        ptr_0130623385.html.

Abraham-Mumm:2000:POT

[ÁMdB00] Erika Ábrahám-Mumm and Frank S. de Boer. Proofoutlines for threads in Java. Lecture Notes in Computer Science, 1877:229–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/1877/18770229.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/1877/18770229.
pdf.

Abraham-Mumm:2002:VJR

[ÁMdBdRS02] Erika Ábrahám-Mumm, Frank S.

de Boer, Willem-Paul de Roever, and Martin Steffen. Verification for Java’s reentrant multithreading concept. Lecture Notes in Computer Science, 2303:5– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2303/23030005.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2303/23030005.
pdf.

AlJaroodi:2005:JJO

[AMJS05] J. AlJaroodi, N. Mohamed, H. Jiang, and D. Swanson. JOPI: a Java Object-Passing Interface. Concurrency and Computation: Practice and Experience, 17(7/8):775–796, 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Amsterdam:2000:JR

[Ams00] Jonathan Amsterdam. Java references. Dr. Dobb’s Journal of Software Tools, 25 (2):42, 44, 46, 48, 50, 52, February 2000. CODEN DDJOEB. ISSN 1044-789X.

                                        URL http://www.ddj.com/
                                        ftp/2000/2000_02/jrefer.
                                        txt.

Amsterdam:2002:JNC

[Ams02] Jonathan Amsterdam. Java’s new considered harmful. Dr. Dobb’s Journal of Software Tools, 27(4):19–20, 22, 24, 26, April 2002. CODEN DDJOEB. ISSN 1044-789X.

                                        URL http://www.ddj.com/.

Anantharam:2001:EJP

[Ana01] Parasuram Anantharam. Essentials of the Java programming language: a hands-on guide. ACM SIGSOFT Software Engineering Notes, 26 (2):76, March 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Andersson:2001:KDJ

[And01]  Johan Andersson.  Kaf-

femik — a distributed JVM featuring a single address space. In USENIX Association [USE01c], page ?? ISBN 1-880446-11-1. LCCN QA76.73.J38 J42 2001. URL

                                        http://www.usenix.org/publications/
                                        library/proceedings/jvm01/
                                        JVM_wips/S09.pdf. Spon-

sored by the USENIX Association.

Andersen:2002:DSJ

[And02] Sandra Andersen. Data structures in Java: a laboratory course. Jones and Bartlett, Boston, MA, USA, 2002. ISBN 0-7637-1816-5. 407 pp. LCCN QA76.73.J38 A46 2002.

Anderson:2004:MPJ

[And04]  I. M. Anderson.  Maple

packages and Java applets. London Mathematical Society Lecture Note Series, 312:193– 206, 2004. CODEN ???? ISSN 0076-0552.

Angell:2000:PSPa

[Ang00a] Kirby W. Angell. Python server pages: Part 1. Dr. Dobb’s Journal of Software Tools, 25(1):44, 46– 47, 50, January 2000. CO-DEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/ftp/2000/
2000_01/psp1.txt;   http:
//www.ddj.com/ftp/2000/
2000_01/psp100.zip.

Angell:2000:PSPb

[Ang00b] Kirby W. Angell. Python server pages: Part II. Dr. Dobb’s Journal of Software Tools, 25(2):54, 57–61, February 2000. CODEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/ftp/2000/
2000_02/psp2.txt.

Angell:2001:JSS

[Ang01]  Kirby W. Angell.  The

Java Secure Socket Extensions. Dr. Dobb’s Journal of Software Tools, 26(2):21– 22, 24, 26, 28, February 2001. CODEN DDJOEB. ISSN 1044-789X. URL

http://www.ddj.com/ftp/
2001/2001_02/jsse.txt;
http://www.ddj.com/ftp/
2001/2001_02/jsse.zip.

Angus:2006:PST

[Ang06]  Chris Angus.  Prototype

and scriptaculous: taking the pain out of JavaScript. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2006. ISBN 0-596-52919-8. LCCN QA76.73.J38 A54 2006eb; QA76.73.J38. URL http://

www.oreilly.com/catalog/
9780596529192.

Azevedo:2000:AAJ

[ANH00] Ana Azevedo, Alex Nicolau, and Joe Hummel. An annotation-aware Java virtual machine implementation. Concurrency: Practice and Experience, 12(6): 423–444, May 2000. CO-DEN CPEXEI. ISSN 1040-

3108. URL http://www3.
interscience.wiley.com/
cgi-bin/abstract/72515731/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=72515731&PLACEBO=IE.
pdf.

Andreae:2006:FIP

[ANMM06] Chris Andreae, James Noble, Shane Markstrum, and Todd Millstein. A framework for implementing pluggable type systems. ACM SIGPLAN Notices, 41(10):57–74, October 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Adams:2001:JIC

[ANN01] Joel Adams, Larry R. Nyhoff, and Jeffrey Nyhoff. Java: An Introduction to Computing. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2001. ISBN 0-13-014251-4. lvi + 945 pp. LCCN QA76.73.J38 A33 2001. US$70.00. URL http:/

/www.phptr.com/ptrbooks/
esm_0130142514.html.

Anonymous:2000:AJV

[Ano00a] Anonymous. Announcement:

Java Virtual Machine Research and Technology Symposium (JVM ’01). ;login: the USENIX Association newsletter, 25(3):??, June 2000. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/events/jvm01.

Anonymous:2000:BRJa

[Ano00b] Anonymous. Book review:

Java enterprise in a nutshell: a desktop quick reference: By David Flanagan, Jim Farlye, William Crawford and Kris Magnusson. O’Reilly, Sebastopol, CA. (1999). 604 pages. $29.95. Computers and Mathematics with Applications, 39(3–4):263, February 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122100900373.

Anonymous:2000:BRJb

[Ano00c] Anonymous. Book review: JavaScript application cookbook: By Jerry Bradenbaugh. O’Reilly, Sebastopol, CA. (1999). 462 pages. $34.95. Computers and Mathematics with Applications, 39(3– 4):266, February 2000. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122100900798.

Anonymous:2000:BRL

[Ano00d] Anonymous. Book reviews: Linux Red Hat Certified Engineer Exam Cram by Andrew G. Feinberg; JavaScript Application Cookbook by Ralph Krause; Programming Pearls, Second Edition by Harvey Friedman. Linux Journal, 73:??, May 2000. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http:

//noframes.linuxjournal.
com/lj-issues/issue73/3831.
html;   http://noframes.
linuxjournal.com/lj-issues/

issue73/3833.html;   http:
//noframes.linuxjournal.
com/lj-issues/issue73/3846.
html.

Anonymous:2000:J

[Ano00e] Anonymous. Java 2. ENI Publishing, ????, 2000. ISBN 2-7460-0559-X. ???? pp. LCCN ???? US$25.

Anonymous:2000:JAR

[Ano00f]  Anonymous.  Java ap-

plets read protected resources. Network Security, 2000(9):3, September 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S135348580009005X.

Anonymous:2000:JBS

[Ano00g] Anonymous. Java-based secure application environment created. Network Security, 2000(4):6, April 1, 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1353485800886201.

Anonymous:2000:NPH

[Ano00h] Anonymous. New products:

Heavy Gear II, Loki Entertainment Software; Compaq Power Management Software, Compaq Computer Corporation; Open Motif Everywhere, Integrated Computer Solutions; VXA Rak-Pak, Ecrix Corporation; Embedix SDK, Lineo, Inc.; Gaussian on Linux NetworX, Linux NetworX, Inc.; HELIOS EtherShare 2.6, HELIOS Software GmbH; NUMA-Q E410 and Netfinity 3500 M20, Shop IBM; InstallShield Java Edition 3.5, InstallShield Software Corporation; NetMAX IA Software, Cybernet Systems Corporation; Teamware Office 5.3 for Linux, Teamware Group Oy; TimeSys Linux/RT 1.0, TimeSys Corporation; Linux 1-2-3 Fundamentals CD, OmniLinux. Linux Journal, 77: ??, September 2000. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Anonymous:2000:NPI

[Ano00i] Anonymous. New products: IQ2 NetCONNECT, BAS-COM Global Internet Services, Inc.; OpenLinux eServer 2.3, Caldera Systems, Inc.; CommuniGate Pro, Super-Symmetric Dynamic Cluster Version, Stalker Software, Inc.; Embedix Linux 1.0, Lineo, Inc.; white dwarf linux 1.0, EMJ Embedded Systems; Linux-based Internal G.lite Modem, SAS; IntraLinux, STS International, Inc.; Linux End-to-End Security Solution, JAWS Technologies; JetForm Central for Linux, JetForm Corporation; JWAVE version 3.0, Visual Numerics, Inc.; RTEL, Lantronix; Japanese Word-Mage v.5.7, Lava Software Pty. Ltd.; TotalView 4.0 Parallel Debugger, Etnus; iNUX Consumer Linux Personal Computer systems, iNUX Inc. Linux Journal, 73:??, May 2000. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Anonymous:2000:NPL

[Ano00j] Anonymous. New products:

Linux Office Solutions, Vista-Source Inc.; CodeWizard 3.1, ParaSoft; eEMU, Jarrix Systems Pty Ltd; RIA Server, Crystal Group Inc.; Exile III: Ruined World, Spiderweb Software; User Management in MandrakeSoft 7.1, MandrakeSoft, Inc.; HostML and ViewML, Century Software; Flipper Graph Control 2.0, ProWorks LLC; RTAI v1.3, RTAI; eServer.group, Technauts Inc.; VCOM on Linux, NetSys Software Group; RM1U-AXe and RM2U-AXi-C, Rave Computer Association, Inc.; TowerJ 3.5, Tower Technology Corporation; X-Win32 v5.0, StarNet Communications Corporation. Linux Journal, 76:??, August 2000. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Anonymous:2000:NPP

[Ano00k]  Anonymous.  New prod-

ucts: PerfectBACKUP+ 6.1, Merlin Software Technologies; Linux Driver for HIPPI 800, Essential Communication Corporation; Linux by Libranet, Libra Computer Systems Ltd.; Programming Development Kit, Macmillan Computer Publishing; Linux Anti-Virus Solution, DOLFIN.COM Inc.; Open-Desk.com version 1.0, HBE Software; UnForm v4.0, Synergetic Data Systems; Max for Linux, PlugSys International LLC; PizzaBox Linux Distribution, KYZO Ltd, Little Streams, The Abbotsbrook, Bourne End, Bucks; Appgen Linux Java Client and PowerWindows Applications, Appgen Business Software, Inc.; IVR Server, Open Source Telecom; Photogenics, Paul Nolan Ltd. Linux Journal, 70:??, February 2000. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Anonymous:2000:NAS

[Ano00l] Anonymous. News: Supercomputing, teraflop performance, embedded Java, middleware. IEEE Concurrency, 8(1):5–??, January/March 2000. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http:

                                        //dlib.computer.org/pd/
                                        books/pd2000/pdf/p1005.
                                        pdf.

Anonymous:2000:PBA

[Ano00m]  Anonymous.  Products:

Broadcom adds VoIP and home networking to cable modem chip; CodeWarrior 6.0 for the Mac; Inprise/Borland JBuilder 4; WinDK extension for Bluetooth; System Mechanic Mobile Toolkit; ActiveState Perl Dev Kit 2.0 for Perl 5.6; Pentium III with SpeedStep; Progress Software Apptivity version 3. Computer, 33(11): 116–117, November 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//dlib.computer.org/co/
books/co2000/pdf/ry116.
pdf.

Anonymous:2000:POR

[Ano00n] Anonymous. Products: Oracle releases XDK update; Starbase’s code editing system; Arc Second’s palm PC CAD viewer; Minolta’s network document server for Windows 2000; Borland’s Java development tools for Palm OS; Rational’s code management tools; Blaxxun Interactive’s Web communications platform tools; Informix Software’s Linux database engine; ActiveState updates free Python distribution; KDE 2.0 released. Computer, 33(12): 144–146, December 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//dlib.computer.org/co/
books/co2000/pdf/rz144.

                                        pdf.

Anonymous:2000:TSJ

[Ano00o] Anonymous. Tool set for the Java Card platform. IEEE Micro, 20(3):87, May/June 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Anonymous:2001:BRJ

[Ano01a] Anonymous. Book review: Java programming: From the beginning: By K. N. King. W. W. Norton, New York. (2000). 788 pages. $52.50. Computers and Mathematics with Applications, 41(3– 4):541, February 2001. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122101900331.

Anonymous:2001:CRJ

[Ano01b] Anonymous. Conference report: 1st Java VM Virtual Machine Research and Technology Symposium (JVM ’01). ;login: the USENIX Association newsletter, 26(5): ??, August 2001. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:2001:GLW

[Ano01c] Anonymous. The Grande Latte wireless experience. IEEE Spectrum, 38(5):9, May 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2001:JAV

[Ano01d] Anonymous. A Java applet for visualizing proteinprotein interaction. Bioinformatics, 17(7):669–671, July 2001. CODEN ???? ISSN 1367-4803 (print), 1367-4811 (electronic).

Anonymous:2001:JJ

[Ano01e] Anonymous, editor. JavaOne 2001, June 7, 2001. ????, ????, 2001. ISBN ???? LCCN ????

Anonymous:2001:LCO

[Ano01f] Anonymous. Leaders cooperate on Java technology. Card Technology Today, 13 (4):6–7, April 1, 2001. CO-DEN ???? ISSN 0965-2590.

Anonymous:2001:PJV

[Ano01g] Anonymous, editor. Proceedings of the Java Virtual Machine Research and Technology Symposium (JVM ’01) April 23–24, 2001, Monterey, California, USA. USENIX Association, Berkeley, CA, USA, 2001. ISBN 1-880446-11-1. LCCN ????

Anonymous:2001:PCP

[Ano01h]  Anonymous.  Products:

Cross-platform toolkit for Bristol Technology; Install-Shield updates Windows installer; Droplet offers unique Web application SDK; ObjectFX Corporation’s Webbased visualization software; Basis Technology updates C++ library; MathWorks unveils embedded control design suite; Intuitive Systems offers Java profiling tool; Computer Associates’ E-business development environment; Programming Research Ltd.’s static analysis tool. Computer, 34(1):130–131, January 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//dlib.computer.org/co/
books/co2001/pdf/r1130.
pdf.

Anonymous:2001:PFS

[Ano01i] Anonymous. Products: Free Software Foundation updates compiler toolset; IT Factory’s Lotus Tools Suite; Cardiff Software’s real-time document verification application; Great Bridge updates open source database; OpenPath Products’ wireless application authoring tool; Curl’s Web application development environment; ThinAirApp ships mobile Visual Basic IDE; Princeton Softech updates server archiving software; Codemesh releases Java communications environment. Computer, 34 (9):112–114, September 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//dlib.computer.org/co/
books/co2001/pdf/r9112.
pdf;   http://www.computer.
org/computer/co2001/r9112abs.

htm.

Anonymous:2001:PGH

[Ano01j]  Anonymous.  Products:

Green Hills ships StarCore development probe; Zayante’s FireWire support for Windows CE; Embedded Performance develops SoC debugger; Extended Systems updates Bluetooth development kit; First Embedded IPv6 router reference platform; Integrated Wireless Java technology; Java performance on wireless devices; Syclo’s mobile software development platform. Computer, 34(11):124–126, November 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//dlib.computer.org/co/
books/co2001/pdf/ry124.
pdf;   http://www.computer.
org/computer/co2001/ry124abs.
htm.

Anonymous:2001:PPT

[Ano01k]  Anonymous.  Products:

Planet 7 Technologies’ new XML development software; Apple Computer launches UNIX-based operating system; Codemesh releases Java/ C++ integration software for Solaris; Telelogic AB’s C++ quality assessment environment; Green Hills ships embedded integrated development platform; Concurrent Computer introduces Linux application development environment. Computer, 34(5):85–86, May 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

                                        //dlib.computer.org/co/
                                        books/co2001/pdf/r5085.
                                        pdf.

Anonymous:2001:PPS

[Ano01l] Anonymous. Products: ProxySource’s software design and collaboration application; YesSoftware’s code generation application; Persistence Software’s transactional application server; Instantiation’s Java productivity tools; JCanvas visual rapid application IDE; theKompany.com’s Python development environment; Neu-Vis updates E-business visual modeling tools; LegacyJ’s Java-compliant Cobol compiler. Computer, 34 (3):108–109, March 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

                                        //dlib.computer.org/co/
                                        books/co2001/pdf/r3108.
                                        pdf.

Anonymous:2001:PSX

[Ano01m] Anonymous. Products: Soft-Quad’s XML content creation software; OriginLab updates graphing tool; NuSphere’s enterprise Web development platform; MetaWare’s XScale programming tools; Aether Systems’ wireless development environment; Visual Numerics upgrades Java application deployment tools; C Level Design introduces C/C++ hardware design environment; ActiveState’s Perl development and deployment software; Advanced Software ships UML design tool; Borland’s Java 2 rapid application development environment; Web services application development platform; RidgeRun’s embedded Linux development kit; IONA modeling and development environment. Computer, 34(7):90–92, July 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//dlib.computer.org/co/
books/co2001/pdf/r7090.
pdf.

Anonymous:2001:PVL

[Ano01n]  Anonymous.  Products:

Viosoft’s Linux embedded development environment; Popkin Software releases development modeling suite; Iopsis Software’s Forte for Java IDE; NQL releases scripting language components; Ascend Software updates delivery management system; Excel Software ships UML design tool; Hyperformix’s discreteevent simulation modeler; In-Cert’s application fault management software; BioconX releases biometric security software; Waterloo Maple’s mathematical computation software [Maple 7]; Equator Technologies’ application development toolkit. Computer, 34(8):84–86, August 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//dlib.computer.org/co/
books/co2001/pdf/r8084.
pdf;   http://www.computer.
org/computer/co2001/r8084abs.
htm.

Anonymous:2001:PWB

[Ano01o] Anonymous. Products: Webbased remote administration tools; SGDL System’s 3D model development language kit; MigraTEC’s Solaris-to-Linux migration software; Visual Numerics updates C numerical library; Stardock’s Windows skin development software; InterNetwork’s new load capacity testing software; SuSE Linux for PowerPC; Raytheon updates network security tools; Tasking updates embedded development tools; ExoLab Group offers open-source data-binding software; Omnicore Software’s Java development environment; Basis International releases Java-based business basic; Zondigo’s wireless software development kit; MDD introduces password administration software; StatSoft revises data visualization tool; Abaco updates mobile application development framework. Computer, 34(6):90–93, June 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//dlib.computer.org/co/
books/co2001/pdf/r6090.
pdf.

Anonymous:2001:TIJ

[Ano01p] Anonymous. Taiwan to issue Java-based insurance card from G&D. Card Technology Today, 13(9):4, October 1, 2001. CODEN ???? ISSN 0965-2590.

Anonymous:2002:CCG

[Ano02a] Anonymous. CICS Commarea generation for VisualAge for Java server interface. Research Disclosure, 462:1958–1959, 2002. CO-DEN RSDSBB. ISSN 0374-4353.

Anonymous:2002:CRJ

[Ano02b] Anonymous. Conference report: 2nd Java Virtual Machine Research and Technology Symposium. ;login: the USENIX Association newsletter, 27(5):??, October 2002. CODEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
library/proceedings/javavm02/
JavaVM2002.pdf.

Anonymous:2002:CDG

[Ano02c] Anonymous. CPU/DSP gains Java accelerator. Electronic Design, 50(23):30–31, 2002. CODEN ELODAW. ISSN 0013-4872.

Anonymous:2002:GLN

[Ano02d] Anonymous. Gemplus launches new Java productivity tools. Card Technology Today, 14 (5):5–6, May 1, 2002. CO-DEN ???? ISSN 0965-2590.

Anonymous:2002:IAJ

[Ano02e] Anonymous. Introducing aspects to Java programs without a custom JVM or application source modification. Research Disclosure, 462:1907– ??, 2002. CODEN RSDSBB. ISSN 0374-4353.

Anonymous:2002:IJM

[Ano02f] Anonymous. iPro: The Java Message Service (JMS) offers reliable messaging without requiring realtime connections. PC Magazine, 21(21):80–84, 2002. CODEN PCMGEP. ISSN 0888-8507.

Anonymous:2002:JGI

[Ano02g] Anonymous. JMFA — A graphically interactive Java program that fits microfibril angle X-ray diffraction data. Research note FPL-RN-0283., United States Forest Service, ????, 2002. Shipping List number 2002-0228-M. Shipping List Date: 04/05/2002.

Anonymous:2002:LAJ

[Ano02h] Anonymous. Learn about Java server-side development and programming: a review of Server-Based Java Programming. IEEE Distributed Systems Online, 3(2), 2002.

ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:

//dsonline.computer.org/
0202/departments/bks_b.
htm.

Anonymous:2002:MIC

[Ano02i] Anonymous, editor. MathML International Conference: Hickory Ridge Conference Center, Chicago, IL, USA, June 28–30, 2000. ????, ????, 2002. ISBN ???? LCCN ????

Anonymous:2002:MES

[Ano02j] Anonymous. A method for eliminating sign extensions for array bounds checking of Java on 64-bit architectures that have no 32-bit compare instruction. Research Disclosure, 462:1954–??, 2002. CO-DEN RSDSBB. ISSN 0374-4353.

Anonymous:2002:NMD

[Ano02k] Anonymous. Naming and metadata design for querying Enterprise Java Beans considering different inheritance hierarchy on remote interface and bean interface. Research Disclosure, 462:1919– ??, 2002. CODEN RSDSBB. ISSN 0374-4353.

Anonymous:2002:PPU

[Ano02l] Anonymous. Performance push: An updated version of webMethods’ integration software supports Java 2 Enterprise Edition, which could improve system performance.

Information Week, 914:30– 31, 2002. CODEN INFWE4. ISSN 8750-6874.

Anonymous:2002:PAU

[Ano02m] Anonymous. Products: Atinav upgrades Bluetooth software suite; new 3D terrain modeler from MultiGen-Paradigm; RIM introduces BlackBerry development environment for J2ME; Mercury Interactive adds Forte for Java to load-testing system; WildPackets’ real-time wireless analyzer; Novell updates mobile file management software; Layton Graphics unveils AutoCADto-PDF translator; Iomegas network attached storage servers. Computer, 35(5): 100–101, May 2002. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//dlib.computer.org/co/
books/co2002/pdf/r5100.
pdf;   http://www.computer.
org/computer/co2002/r5100abs.
htm.

Anonymous:2002:PEB

[Ano02n] Anonymous. Products: Enhanced Bluetooth test tool from Tektronix; NEXIQ Technologies’s intelligent display software; Actel’s FGPA development IDE; Parasoft’s automated Java classes testing unit; Packeteer upgrades central reporting application; VisiComp releases Java debugger; Compuware’s driver development suite for Windows; Silas Technologies upgrades applicationmonitoring software. Computer, 35(9):82–83, September 2002. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/co/
2002/09/r9082.htm;   http:
//csdl.computer.org/dl/
mags/co/2002/09/r9082.pdf.

Anonymous:2002:PIR

[Ano02o] Anonymous. Products: Instantiations releases RAD tools for Java; Pencel updates XML mobile application interface; LynuxWorks’s new Linux distribution; Applied Microsystems’ embedded systems code tester; AppForge upgrades crossplatform design tool for handhelds; Computer Associates’ Java, XML, and wireless extensions; TurboPower;s bug detection and performance analysis tools. Computer, 35 (2):106–108, February 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//dlib.computer.org/co/
books/co2002/pdf/r2106.
pdf;   http://www.computer.
org/computer/co2002/r2106abs.
htm.

Anonymous:2002:POU

[Ano02p] Anonymous. Products: Omnicore upgrades Java IDE CodeGuide emWare’s SDE for intelligent device management; Metrowerks’ Code-Warrior for Embedded Linux; integrated software environment form Xilinx; new version of InstallShield Professional; Motorola’s 32-bit CAN reference design; Utopia-LVDS bridge reference design kit from National Semiconductor; First Silicon Solutions’ analysis tool for flash-based FPGAs. Computer, 35(11): 78–79, November 2002. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/co/
2002/11/ry078.htm;   http:
//csdl.computer.org/dl/
mags/co/2002/11/ry078.pdf.

Anonymous:2002:PPJ

[Ano02q]  Anonymous.  Products:

PrismTech’s JDO spec for transparent persistence; Altia’s graphics code generator for embedded applications; Design Science upgrades MathType for windows; PolarLake launches Enterprise XML platform for Java; Syware’s database development software for PDAs; code generator for Web application development from YesSoftware; Embarcadero Technologies upgrades crossplatform job scheduler; Performance Technologies introduces telecom adapter; Rational Software’s latest IDE enhancement; Aprisa’s online research and design environment. Computer, 35(3):97–99, March 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//dlib.computer.org/co/
books/co2002/pdf/r3097.
pdf;   http://www.computer.
org/computer/co2002/r3097abs.
htm.

Anonymous:2002:PRS

[Ano02r] Anonymous. Products: Rational Software’s .NET development software; SGI’s development APIs for visualization; Java development tool from Visual Numerics; SUSSMicroTec’s MEMS testing vacuum prober; Face-Time’s IM information security system; MX development resource kit from Macromedia; .NET data integration tool from XAWare; VoiceGenie’s VoiceXML gateway. Computer, 35(10):70– ??, October 2002. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/co/
2002/10/rx070.htm;   http:
//csdl.computer.org/dl/
mags/co/2002/10/rx070.pdf.

Anonymous:2002:PSS

[Ano02s]  Anonymous.  Products:

SOISIC ships deisgn kit for SOI structures; systems and software development tools from Telelogic; RSA Security’s Web access management system; Altera’s free embedded processor portfolio; signal integrity measurement tools from tektronix; Oracle upgrades Java development tool; Xilinx delivers EDK for FPGA processor; Westbridge’s tool to sniff unauthorized XML; SpeechStudio’s telephony development tools. Computer, 35(12):118– 119, December 2002. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/co/
2002/12/rz118.htm;   http:
//csdl.computer.org/dl/
mags/co/2002/12/rz118.pdf.

Anonymous:2002:PXO

[Ano02t]  Anonymous.  Products:

XML-oriented FrameMaker from Adobe Systems; Breeze Factor’s XML data-binding tool for Java; Microsoft’s speech-enabling tools for Web applications; Agere Systems’s wireless LAN management tool; Macromedia debuts ColdFusion MX; OpenOffice.org unveils open source productivity suite; Integrated Measurement Systems’ validation tester. Computer, 35 (6):82–83, June 2002. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/co/
2002/06/r6082.htm;   http:
//csdl.computer.org/dl/
mags/co/2002/06/r6082.pdf.

Anonymous:2002:RCJ

[Ano02u] Anonymous. Resource constraints in Java by means of a classfile modification. Research Disclosure, 463:2105, 2002. CODEN RSDSBB. ISSN 0374-4353.

Anonymous:2002:SAC

[Ano02v]  Anonymous.  Schlumberg-

erSema adds on-card applet verification to Java Cards. Card Technology Today, 14 (5):6–7, May 1, 2002. CO-DEN ???? ISSN 0965-2590.

Anonymous:2002:VJU

[Ano02w] Anonymous. VisualAge for Java UI application components for securities and investment trust industry. Research Disclosure, 463:2236– 2237, 2002. CODEN RS-DSBB. ISSN 0374-4353.

Anonymous:2003:AOS

[Ano03a] Anonymous. Avanti: Open source compact, self-contained ILS in Java. Information Retrieval and Library Automation, 38(8):7–??, 2003. CO-DEN IRLAAQ. ISSN 0020-0220.

Anonymous:2003:BRJ

[Ano03b] Anonymous. Book review:

Java precisely: By Peter Sestoft. The MIT Press, Cambridge, MA. (2002). 118 pages. $14.95. Computers and Mathematics with Applications, 46(2–3):517–518, July/August 2003. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122103900708.

Anonymous:2003:BJJ

[Ano03c] Anonymous. Borland Jbuilder Java 8. Personal computer world, 26(3):96, 2003. CO-DEN PCWODU. ISSN 0142-0232.

Anonymous:2003:BNA

[Ano03d]  Anonymous.  Brave new

apps. PC Magazine, 22 (13):114–137, 2003. CODEN PCMGEP. ISSN 0888-8507.

Anonymous:2003:CWD

[Ano03e]  Anonymous.  Controlling

Webserver demand for Java applets for browsers supporting Java 2 JRE. Research Disclosure, 466:302, 2003. CO-DEN RSDSBB. ISSN 0374-4353.

Anonymous:2003:DJR

[Ano03f] Anonymous. Demand for Java remains high because companies are keen to develop front-end Web sites. Computing (London, England: 1980), page 43, March 13, 2003. CO-DEN CPTGB5. ISSN 0144-3097.

Anonymous:2003:ELN

[Ano03g] Anonymous. EEMBC launches new benchmarks to address digital media and Java. New Electronics, 36(16):12, 2003. CODEN ???? ISSN 0047-9624.

Anonymous:2003:FFG

[Ano03h] Anonymous. “filter” — a framework to generate subsets of collections in programs written in Java programming language. it is a piece of software useful for other software projects. Research Disclosure, 466:322, 2003. CODEN RS-DSBB. ISSN 0374-4353.

Anonymous:2003:JLO

[Ano03i] Anonymous. For Java lovers:

Only gadgets to amp the flavor of your brew. Business week, 3861:110–111, 2003. CODEN BUWEA3. ISSN 0739-8395.

Anonymous:2003:TMC

[Ano03j] Anonymous. For Taiwan’s 22 million citizens, Java Smart Cards are clamping down on health-care fraud. PC Magazine, 22(17):66–67, 2003. CO-DEN PCMGEP. ISSN 0888-8507.

Anonymous:2003:FWA

[Ano03k] Anonymous. Four-way asynchronous I/O using dual paired buffers for Java-based intermediate/proxy HTTP nodes in a communications network. Research Disclosure, 466:313, 2003. CODEN RS-DSBB. ISSN 0374-4353.

Anonymous:2003:GUI

[Ano03l] Anonymous. Graphical user interface primitives independent library for building Java based software. Research Disclosure, 466:299, 2003. CO-DEN RSDSBB. ISSN 0374-4353.

Anonymous:2003:IMM

[Ano03m] Anonymous. ICT Motorola mise sur Java et J2ME. Enjeux, 233:22, 2003. CODEN ???? ISSN 0223-4866.

Anonymous:2003:IUU

[Ano03n] Anonymous. Industry upstarts: US-based SavaJe has developed a pure Java handset OS which it wants to sell to operators directly. Mobile Communications International, 100:48, 2003. CO-DEN ???? ISSN 1352-9226.

Anonymous:2003:JAT

[Ano03o] Anonymous. Java API tackles Bluetooth interop: The Java community says it has quietly developed a solution for Bluetooth interoperability headaches. Electronic Engineering Times, 1262:16–30, 2003. CODEN ???? ISSN 0192-1541.

Anonymous:2003:JDT

[Ano03p] Anonymous. Java development tools speed MIDIet creation. Electronic Design, 51 (8):62, 2003. CODEN ELO-DAW. ISSN 0013-4872.

Anonymous:2003:JEF

[Ano03q] Anonymous. Java environment focuses on up-front modeling. Application Development Trends, 10(5):34, 2003. CODEN ADTRF4. ISSN 1073-9564.

Anonymous:2003:JGJ

[Ano03r] Anonymous. Java gets a jolt. Information Week, 944: 27, 2003. CODEN INFWE4. ISSN 8750-6874.

Anonymous:2003:JEJ

[Ano03s] Anonymous. Java in Echtzeit:

Java-Technologien für Echtzeit-Anwendungen mit der Real-Time Specification for Java. (German) [Java in real time: Java for real-time applications with the Real-Time Specification for Java]. Elektronik, 52(16):66–71, 2003. CODEN EKRKAR. ISSN 0013-5658.

Anonymous:2003:JPa

[Ano03t] Anonymous. Java performance. ACM SIGPLAN Notices, 38(11):??, November 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:2003:JPb

[Ano03u] Anonymous. Java precisely.

Computers and Mathematics with Applications, 46 (2–3):517–518, 2003. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2003:JPc

[Ano03v] Anonymous. Java precisely.

Computers and Mathematics with Applications, 46 (2–3):517–518, 2003. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2003:JHS

[Ano03w] Anonymous. Java’s head start adoption of Microsoft’s C# language for building Web services is hindered by the prevalence of Java. Information Week, 966:57, 2003. CO-DEN INFWE4. ISSN 8750-6874.

Anonymous:2003:LUE

[Ano03x] Anonymous. Linux/Unix extend Red Hat 8’s functionality by adding Flash, extra fonts and Java. Personal computer world, 26(3): 202–207, 2003. CODEN PC-WODU. ISSN 0142-0232.

Anonymous:2003:MJA

[Ano03y] Anonymous. Manage Java apps for premium performance. Application Development Trends, 10(1):50–52, 2003. CODEN ADTRF4. ISSN 1073-9564.

Anonymous:2003:MMI

[Ano03z] Anonymous. Microsoft must include Java in windows, rules a US court. Computing (London, England: 1980), pages 13–15, January 9, 2003. CO-DEN CPTGB5. ISSN 0144-3097.

Anonymous:2003:JTM

[Ano03-27]  Anonymous.  Mit Java-

Technologie mehr Komfort.

VDI Z, 145(3):47–48, 2003. CODEN ???? ISSN 0042-1766.

Anonymous:2003:NIC

[Ano03-28] Anonymous. National ID card supporting Java applications. Toshiba Review, 58(8): 56–59, 2003. CODEN ???? ISSN 0372-0462.

Anonymous:2003:NRJ

[Ano03-29] Anonymous. New release of JANIS (Java-based nuclear data information services) software. Nea News, 21 (1):25, 2003. CODEN ???? ISSN 1605-9581.

Anonymous:2003:NAQ

[Ano03-30] Anonymous. No assembly: Q-Link Technologies’ upgraded Java development environment helps developers build apps using preassembled components. Information Week, 939:54–55, 2003. CODEN IN-FWE4. ISSN 8750-6874.

Anonymous:2003:OTJ

[Ano03-31] Anonymous. Octera throws a Javalon: It’s not 100%- pure Java, but the Javalon-1 processor natively executes most Java bytecode instructions. Microprocessor report, 17(3):31–37, 2003. CODEN ???? ISSN 0899-9341.

Anonymous:2003:PPG

[Ano03-32]  Anonymous.  Peripherals

printers get Java networking. Computer Weekly, pages 26– 27, March 20, 2003. CODEN COMWAA. ISSN 0010-4787.

Anonymous:2003:PLJ

[Ano03-33] Anonymous. Portable Lösung: Java in Embedded-Systemen. (German) [Portable solution: Java in embedded systems]. Elektronik Praxis, 15:36–45, 2003. CODEN ???? ISSN 0341-5589.

Anonymous:2003:PBS

[Ano03-34] Anonymous. Products: BEA Systems releases preview of Java API for XML; Panda appliance defends network perimeters; Parasoft enhances Java test and analysis tool. Computer, 36(12): 108, December 2003. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/co/
2003/12/rz108.pdf.

Anonymous:2003:PCN

[Ano03-35]  Anonymous.  Products:

ClearSight Networks releases application-layer analyzer; Intervoice announces first SALT-based components; VoiceGenie Technologies upgrades VoiceXML platform; AppForce enhances mobileplatform design software; Metrowerks upgrades tools for embedded Linux products; OpenOffice.org updates Linux office tool suite; Quest Software releases Java analysis tools. Computer, 36(11): 86–87, November 2003. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/co/
2003/11/ry086.htm;   http:
//csdl.computer.org/dl/
mags/co/2003/11/ry086.pdf.

Anonymous:2003:PCU

[Ano03-36] Anonymous. Products: Compuware upgrades J2EE development environment; Ektron releases browser-based image tool; IronGrid offers JDBC performance tool; Microsoft enhances Java conversion assistant; Broadcom announces single-chip 10-Gigabit Ethernet switch; SGI finalizes OpenGL 1.5 specification; Adaptec extends Serial ATA RAID product family. Computer, 36(9):94– 95, September 2003. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/co/
2003/09/r9094.htm;   http:
//csdl.computer.org/dl/
mags/co/2003/09/r9094.pdf.

Anonymous:2003:PJU

[Ano03-37] Anonymous. Products: Jet-Brains upgrades IntelliJ Java IDE; Catalyst’s USB analyzer supports device emulation; VMware releases Enterprise Server VM software; Motorola offers free soft modem reference design; RealNetworks releases source for Helix DNA Server; Packeteer accelerates intranet and Internet applications. Computer, 36(3): 80–81, March 2003. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/dl/mags/co/
                                        2003/03/r3080.htm;   http:
                                        //csdl.computer.org/dl/
                                        mags/co/2003/03/r3080.pdf.

Anonymous:2003:POU

[Ano03-38] Anonymous. Products: Omnicore upgrades Java development tools; TimeSys announces first RTSJ-compliant JVM; Xilinx adds hardwarein-the-loop cosimulation to DSP design tool; gInstall’s upgrades setup system platform; Motorola launches motocoder developer platform; ARM upgrades compilation tools for embedded apps; SGI server leads HPC benchmarks. Computer, 36(5): 69–70, May 2003. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/dl/mags/co/
                                        2003/05/r5069.htm;   http:
                                        //csdl.computer.org/dl/
                                        mags/co/2003/05/r5069.pdf.

Anonymous:2003:PSA

[Ano03-39] Anonymous. Products: SGI announces record-performing Linux servers; Pacific Nanotechnogy releases atomic force microscope; Nvidia extends Cg programming environment; Empirix introduces VoIP analyzer; Codagen Technologies upgrades UML tool; Perforce Software enhances SCM system; CompuWare upgrades Java development tool. Computer, 36(2):80–81, February 2003. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/co/
2003/02/r2080.htm;   http:
//csdl.computer.org/dl/
mags/co/2003/02/r2080.pdf.

Anonymous:2003:PSR

[Ano03-40] Anonymous. Products: Starbase releases decision-support software; OC Systems extends analysis tool to J2EE; InstallShield streamlines software installation app; Silicon Defense counters stealth scans; compuware upgrades Java profiling tool;’ Pervasive Software releases V8 database engine; Xilinx ships DSP design tool; MKS adds wizards to monitoring system. Computer, 36(1):112– 113, January 2003. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/co/
2003/01/r1112.htm;   http:
//csdl.computer.org/dl/
mags/co/2003/01/r1112.pdf.

Anonymous:2003:PVF

[Ano03-41]  Anonymous.  Products:

VMware’s fourth-generation desktop virtualization software; automated design reviews with Reviewer for Rose; CodeWeavers debues CrossOver Office; Corel Smart Graphics Studio now available; IronGrid’s Java performance tool; Infragistics releases NetAdvantage component collection. Computer, 36(6):98–99, June 2003. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/dl/mags/co/
                                        2003/06/r6098.htm;   http:
                                        //csdl.computer.org/dl/
                                        mags/co/2003/06/r6098.pdf.

Anonymous:2003:RAI

[Ano03-42] Anonymous. Remote access to IMS DB data using DB2 stored procedures and Java ResultSet objects. Research Disclosure, 465:118, 2003. CODEN RSDSBB. ISSN 0374-4353.

Anonymous:2003:RVF

[Ano03-43] Anonymous. RT vendor forum: The many faces of Java. Chain Store Age, 79(5):122– 123, 2003. CODEN ????

Anonymous:2003:RAS

[Ano03-44]  Anonymous.  ‘run any-

where’ scheme eclipses OS. A “write/run anywhere” system puts Java programs into embedded-C programs with no operating system. Electronic Engineering Times, 1260:49–50, 2003. CODEN ???? ISSN 0192-1541.

Anonymous:2003:SPR

[Ano03-45] Anonymous. Simulation program recoded in Java. Advanced Composites Bulletin, APR:7, 2003. CODEN ???? ISSN 0951-953X.

Anonymous:2003:SSA

[Ano03-46] Anonymous. Software, silicon acceleration brew a stronger Java. EDN, 48(6):18–19, 2003. CODEN EDNSBH. ISSN 0012-7515, 0364-6637.

Anonymous:2003:SRJ

[Ano03-47] Anonymous. Sun relance Java. Usine Nouvelle, 2876: 31, 2003. CODEN ???? ISSN 0042-126X.

Anonymous:2003:TAJ

[Ano03-48] Anonymous. TeamStudio analyzer for Java E2. Application Development Advisor, 7 (3):18, 2003. CODEN ???? ISSN 1369-4200.

Anonymous:2003:UJW

[Ano03-49] Anonymous. Using Java on the Web. PC Plus, 198: 184–187, 2003. CODEN ???? ISSN 0952-2565. Richard Cobbett gets graphical with Swing, AWT and CodeWarrior.

Anonymous:2003:VPU

[Ano03-50]  Anonymous.  Visual pro-

gramming using layout managers to achieve displayindependent Java applications. Personal computer world, 26(7):213–214, 2003.

CODEN PCWODU. ISSN 0142-0232.

Anonymous:2003:WOF

[Ano03-51] Anonymous. Wave opens Finread terminals to Java applications. Card Technology Today, 15(1):8–9, 2003. CODEN ???? ISSN 0965-2590.

Anonymous:2003:WRT

[Ano03-52] Anonymous. Will the realtime Java please stand up? Electronic Design, 51(8):61, 2003. CODEN ELODAW. ISSN 0013-4872.

Anonymous:2004:SRJ

[Ano04a] Anonymous. A self refreshing Java GUI component. Research Disclosure, 487:1467, 2004. CODEN ???? ISSN 0374-4353.

Anonymous:2004:ANS

[Ano04b] Anonymous. Agilent’s new System Ready Test debuts. tool extensions for ColdFire and Star and a new high speed Java are among the other new tools. Embedded System Engineering, 12 (6):10–11, 2004. CODEN ???? ISSN 0969-8825.

Anonymous:2004:AVM

[Ano04c] Anonymous. Analyse und Visualisierung von Messdaten: Java — die Brucke zu den Standards von Morgen. (German) [Analysis and visualization of measurement data: Java — The bridge to tomorrow’s standards]. Elektronik, 53(8):125, 2004. CO-DEN EKRKAR. ISSN 0013-5658.

Anonymous:2004:AMJ

[Ano04d] Anonymous. Applications monitor Java apps to minimise downtime. Computer Weekly, pages 30–31, November 23, 2004. CODEN ???? ISSN 0010-4787.

Anonymous:2004:BRPc

[Ano04e] Anonymous. Book review:

Programming with objects: a comparative presentation of object-oriented programming with C++ and Java: By Avinash C. Kak. Wiley– Interscience, Hoboken, NJ. (2003). 1115 pages. $83.95, £63.50. Computers and Mathematics with Applications, 47(6–7):1150, March/April 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S089812210490106X.

Anonymous:2004:BBM

[Ano04f] Anonymous. Building bridges:

Mainsoft offers a product that lets companies move applications built in Microsoft’s Visual Basic and C# to Java platforms. Information Week, 977:30–31, 2004. CODEN IN-FWE4. ISSN 8750-6874.

Anonymous:2004:CGH

[Ano04g] Anonymous. C# goes head to head with Java: When resources are tight, companies try to limit the number of programming languages they support. Microsoft’s C# and Sun’s Java both have strengths, so which do you choose? Computer Weekly, pages 48–50, November 9, 2004. CODEN ???? ISSN 0010-4787.

Anonymous:2004:CJL

[Ano04h]  Anonymous.  Chipkarten:

Java-Lösung für SmartCards. (German) [Chip cards: Java solutions for SmartCards]. Elektronik, 53(4):6, 2004. CO-DEN EKRKAR. ISSN 0013-5658.

Anonymous:2004:CSI

[Ano04i] Anonymous. Closed subject: IBM and BEA Systems call for Sun to contribute Java as open-source code, but Sun’s CEO nixes the idea. Information Week, 996:24, 2004. CO-DEN INFWE4. ISSN 8750-6874.

Anonymous:2004:CCC

[Ano04j] Anonymous. Cutting complexity: Compuware’s OptimalJ 3.1 is designed to simplify Java development. Information Week, 985:62–63, 2004. CODEN INFWE4. ISSN 8750-6874.

Anonymous:2004:DWY

[Ano04k] Anonymous. Deal will yield support for Java. Computer Weekly, pages 12–13, April 13, 2004. CODEN ???? ISSN 0010-4787.

Anonymous:2004:GCV

[Ano04l] Anonymous. Garbage collection versus Echtzeit: Die “Realtime Specification for Java” und ihre Implementierung in JamaicaVM. (German [Reality: The “realtime specification for Java” and its implementation in JamaicaVM]. Elektronik Praxis, 3:70–73, 2004. CODEN ???? ISSN 0341-5589.

Anonymous:2004:GLF

[Ano04m] Anonymous. Gluecode looks to fill in Java stack. InfoWorld, 26(40):22, 2004. CODEN INWODU. ISSN 0199-6649.

Anonymous:2004:GLR

[Ano04n] Anonymous. Groupware Lotus releases Java-based Web services. Computer Weekly, pages 12–13, February 3, 2004. CODEN ???? ISSN 0010-4787.

Anonymous:2004:HSC

[Ano04o]  Anonymous.  Hot skills:

Cobol integrates with Java and .Net. Computer Weekly, page 49, September 7, 2004. CODEN ???? ISSN 0010-4787.

Anonymous:2004:HTJ

[Ano04p] Anonymous. Hot topic: Java 1.5. The Computer Bulletin, 46(5):32, 2004. CO-DEN COBUAH. ISSN 0010-4531 (print), 1464-357X (electronic).

Anonymous:2004:HNV

[Ano04q] Anonymous. How the new visual editor in this free Java development tool can help you. Personal computer world, 27(12):183–186, 2004. CODEN PCWODU. ISSN 0142-0232.

Anonymous:2004:JDC

[Ano04r] Anonymous. Java development: Can Sun’s JCS bridge Java complexity gap? Computer Weekly, pages 26–27, October 12, 2004. CODEN ???? ISSN 0010-4787.

Anonymous:2004:JGO

[Ano04s] Anonymous. Java garbage and other myths. Electronic Design, 52(12):55, 2004. CO-DEN ELODAW. ISSN 0013-4872.

Anonymous:2004:JIP

[Ano04t] Anonymous. Java ID for PCs? Card Technology Today, 16(3):5, 2004. CODEN ???? ISSN 0965-2590.

Anonymous:2004:JRC

[Ano04u] Anonymous. Java releases could spike the return of client-side development. Msi, 22(12):8–9, 2004. CODEN ???? ISSN 1533-7758.

Anonymous:2004:JSB

[Ano04v] Anonymous. Java-Software: Mit und ohne Betriebssystem. (German) [Java software: With and without operating system]. Elektronik, 53(4):8, 2004. CODEN EKRKAR. ISSN 0013-5658.

Anonymous:2004:JSA

[Ano04w] Anonymous. Java: Sun advances with Tiger but users want interoperability. Computer Weekly, page 18, July 6, 2004. CODEN ???? ISSN 0010-4787.

Anonymous:2004:JSS

[Ano04x] Anonymous. Java: Sun simplifies front-end Java development. Computer Weekly, page 32, March 16, 2004. CO-DEN ???? ISSN 0010-4787.

Anonymous:2004:LUI

[Ano04y] Anonymous. LCH to update IT systems to Java architecture. Computer Weekly, page 5, January 20, 2004. CO-DEN ???? ISSN 0010-4787.

Anonymous:2004:MSJ

[Ano04z] Anonymous. MIPS spikes Java set-tops with a dash of Esmertec. Electronic Engineering Times, 1332:10–17, 2004. CODEN ???? ISSN 0192-1541.

Anonymous:2004:NDE

[Ano04-27]  Anonymous.  New devel-

opments: EBay wants to broaden its developer base and says it will add integration capabilities for Soap and Java Web-service technologies to its developers’ kit. Information Week, 976:28, 2004.

CODEN INFWE4. ISSN 8750-6874.

Anonymous:2004:NGJ

[Ano04-28] Anonymous. New generation Java smart cards. Card Technology Today, 16(3):10– 11, 2004. CODEN ???? ISSN 0965-2590.

Anonymous:2004:OJT

[Ano04-29] Anonymous. Oracle Jdeveloper 10G: Tim Anderson examines a cool Java IDE from Oracle Corporation that is well-priced and gives you a development framework too. Application Development Advisor, 8(5):20–25, 2004. CO-DEN ???? ISSN 1369-4200.

Anonymous:2004:POC

[Ano04-30] Anonymous. Programming with objects: a comparative presentation of objectoriented programming with C++ and Java. Computers and Mathematics with Applications, 47(6–7):1150, 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2004:SCS

[Ano04-31] Anonymous. SAP conference: SAP increases support for Java on Netweaver. Computer Weekly, pages 30–31, October 19, 2004. CODEN ???? ISSN 0010-4787.

Anonymous:2004:SMO

[Ano04-32]  Anonymous.  Small but

mighty operating system runs full-fledged Windows, Linux, and Java-based software on mobile equipment. Information Week, 979:61, 2004. CO-DEN INFWE4. ISSN 8750-6874.

Anonymous:2004:SDA

[Ano04-33] Anonymous. Software desarrollo de aplicaciones informaticas en entornos inalambricos. lenguaje Java 2 Micro Edition. Revista española de electronica, 600:48–51, 2004. CO-DEN RVEEBT. ISSN 0482-6396.

Anonymous:2004:SVJ

[Ano04-34] Anonymous. A study of VRML-JAVA based robot motion simulation. Mechanical Science and Technology, 23(1):120–122, 2004. CODEN ???? ISSN 1003-8728.

Anonymous:2004:SJSb

[Ano04-35] Anonymous. Sun Java Studio Creator 2004Q2. PC Magazine, 23(16):54, 2004. CO-DEN PCMGEP. ISSN 0888-8507.

Anonymous:2004:SJSa

[Ano04-36] Anonymous. Sun Java Studio Enterprise 6. PC Magazine, 23(9):37, 2004. CODEN PCMGEP. ISSN 0888-8507.

Anonymous:2004:UCI

[Ano04-37] Anonymous. Utility company integrates COBOL application with enterprise Java system. Computer Economics Report, 26(1):4–6, 2004. CO-DEN ???? ISSN 1054-5026.

Anonymous:2004:VPP

[Ano04-38] Anonymous. Visual programming: The pros and cons of open source, plus Java and SQLite. Personal computer world, 27(5):212, 2004. CO-DEN PCWODU. ISSN 0142-0232.

Anonymous:2004:WSJ

[Ano04-39] Anonymous. Web services Java drives Mars mission portal. Computer Weekly, page 29, March 2, 2004. CO-DEN ???? ISSN 0010-4787.

Anonymous:2005:BKJ

[Ano05a]  Anonymous.  Bluetooth-

Kommunikation unter Java: Ein System zur Konfiguration funkbasierter Automatisierungssysteme. (German) [Bluetooth communication under Java: a system for the configuration of radiobased automation systems]. Elektronik, 54(7):56–61, 2005. CODEN EKRKAR. ISSN 0013-5658.

Anonymous:2005:COE

[Ano05b] Anonymous. C# offers an easy-to-use alternative to Java. Computer Weekly, page 30, January 18, 2005. CODEN ???? ISSN 0010-4787.

Anonymous:2005:CBE

[Ano05c]  Anonymous.  Compuware,

Borland eye development while Compuware upgrades its Java development projects, Borland buys a company. InfoWorld, 27(4):17, 2005. CO-DEN INWODU. ISSN 0199-6649.

Anonymous:2005:FJI

[Ano05d] Anonymous. Four Java IDEs duke it out. InfoWorld, 27 (13):32–38, 2005. CODEN INWODU. ISSN 0199-6649.

Anonymous:2005:JND

[Ano05e] Anonymous. Java and .NET differences remain unresolved. Computer Weekly, pages 14– 15, June 28, 2005. CODEN ???? ISSN 0010-4787.

Anonymous:2005:JGS

[Ano05f]  Anonymous.  Java grows

suites: Sun’s Java Enterprise System is dividing into suites tailored to specific functions such as identity management. InfoWorld, 27(5):16–18, 2005. CODEN INWODU. ISSN 0199-6649.

Anonymous:2005:JF

[Ano05g] Anonymous. Java in the factory. Control Engineering, 52(4):46–52, 2005. CODEN CENGAX. ISSN 0010-8049.

Anonymous:2005:JPF

[Ano05h] Anonymous. Java project fires up supporters, critics. Electronic Engineering Times, 1364:8–19, 2005. CO-DEN ???? ISSN 0192-1541.

Anonymous:2005:OSJ

[Ano05i] Anonymous. Oracle AS 1Og shares the Java joy: Primed for business, Orcale Application Server 1Og Release 2 frees Java apps from traditional server binds. InfoWorld, 27(15):23–25, 2005. CODEN INWODU. ISSN 0199-6649.

Anonymous:2005:PHS

[Ano05j] Anonymous. Presentations: Home servers: Java programming. Av Magazine, 1:33, 2005. CODEN ???? ISSN 1361-3685.

Anonymous:2005:SAS

[Ano05k] Anonymous. Security audits for source code fortify source code: Analysis Suite 3.0 brings deep, intelligent security analysis to C, C++, C#, Java, JSP, and SQL program code. InfoWorld, 27 (3):24–25, 2005. CODEN IN-WODU. ISSN 0199-6649.

Anonymous:2005:SSE

[Ano05l] Anonymous. Sicherheitskritische Software: Echtzeittaugliches Java. Elektronik, 54(HEFT1):26–27, 2005. CO-DEN EKRKAR. ISSN 0013-5658.

Anonymous:2005:SSS

[Ano05m]  Anonymous.  Subscription

suites: Sun Microsystems seeks to undercut competitors by reorganizing its Java middleware stack into suites. Information Week, 1025:32– 33, 2005. CODEN INFWE4. ISSN 8750-6874.

Anonymous:2005:TTT

[Ano05n] Anonymous. TIGER, TIGER:

There’s a new release of Java and although Pan Pantziarka isn’t sure what to call it, he knows where to learn about it. Application Development Advisor, 9(1):28–29, 2005. CO-DEN ???? ISSN 1369-4200.

Anonymous:2005:TPI

[Ano05o] Anonymous. Two plug-ins rev up: Eclipse M7 NitroX and MyEclipse Enterprise Workbench bring depth and breadth to Java Web coding. InfoWorld, 27(6):26–28, 2005. CODEN INWODU. ISSN 0199-6649.

Anonymous:2005:VBJ

[Ano05p] Anonymous. Vanward beefs up Java project tracking. InfoWorld, 27(16):32–33, 2005. CODEN INWODU. ISSN 0199-6649.

Anonymous:2005:VPS

[Ano05q] Anonymous. Visual programming: Some tips on how to make using the Java platform easier and how to make coding simpler. Personal computer world, 28(3):179–182, 2005. CODEN PCWODU. ISSN 0142-0232.

Anonymous:2008:BRBe

[Ano08] Anonymous. Book review:

SAS Graphics for Java: Examples Using SAS AppDev Studio and the Output Delivery System by Wendy Bohnenkamp; Jackie Iverson. Technometrics, 50(3): 412, August 2008. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:

//www.jstor.org/stable/
25471510.

Arbe:2007:FLT

[AOMC07] José Miguel Blanco Arbe, Ana Sánchez Ortega, and Jesús Ibáñez Mártinez-Conde. Formal languages through Web forms and regular expressions. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(4):100–104, December 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp:

//ftp.math.utah.edu/pub/
mirrors/ftp.ira.uka.de/
bibliography/Misc/DBLP/
2007.bib.

Appel:2002:MCI

[AP02] Andrew W. Appel and Jens Palsberg. Modern compiler implementation in Java. Cambridge University Press, Cambridge, UK, second edition, 2002. ISBN 0-521-82060-X. 501 pp. LCCN QA76.73.J38 A65 2002.

Alonso:2004:RTT

[APA04] D. Alonso, J. A. Pastor, and B. Alvarez. Real-time teaching with Java: J P R3. Lecture Notes in Computer Science, 3292:246–255, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

April:2003:AJA

[Apr03] C. A. April. Automating Java apps. Varbusiness, 19(11):51– 52, 2003. CODEN ???? ISSN 0894-5802.

April:2005:NJP

[Apr05] C. A. April. .NET-to-Java porting made easy: Mainsoft offers a tool to alleviate the headaches ISVs face when porting applications. Varbusiness, 21(4):46, 2005. CODEN ???? ISSN 0894-5802.

Apte:2002:JCA

[Apt02] Atul Apte. Java connector architecture: building custom connectors and adapters. Howard W. Sams, Indianapolis, IN 46268, USA, 2002. ISBN 0-672-32310-9. xi + 344 pp. LCCN QA76.73.J38 A67 2002.

Amza:2003:NCB

[AR03a] C. Amza and G. Reggio. A notation for component-based design of Java applications. Lecture Notes in Computer Science, 2604:155–164, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ananian:2003:DSO

[AR03b] C. Scott Ananian and Martin Rinard. Data size optimizations for Java programs. ACM SIGPLAN Notices, 38(7):59– 68, July 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Alagic:2008:GJP

[AR08] Suad Alagić and Mark Royer. Genericity in Java: persistent and database systems implications. VLDB Journal: Very Large Data Bases, 17 (4):847–878, July 2008. CO-DEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Armstrong:2004:JMD

[Arm04] J. N. Armstrong. Java & monetary data. Dr. Dobb’s Journal of Software Tools, 29 (5):52–56, 2004. CODEN DDJOEB. ISSN 1044-789X.

Arrington:2001:EJU

[Arr01] C. T. Arrington. Enterprise Java with UML. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 2001. ISBN 0-471-38680-4 (paperback). xxii + 451 pp. LCCN QA76.73.J38 A78 2001.

Arthur:2000:JES

[Art00] John K. Arthur. Java as an environment for scientific computing. In Langtangen et al. [LBQ00], pages 179–196. CODEN LNCSA6. ISBN 3-540-66557-9 (softcover), 3-642-57172-7 (e-book). ISSN 1439-7358. LCCN QA76.6 .A336

2000. URL http://link.

springer.com/chapter/10.
1007/978-3-642-57172-5_

5/. Papers from an International Workshop on Modern Software Tools for Scientific Computing (Sci-Tools’98), Oslo, Norway, September 14–16, 1998.

Agarwal:2003:TIP

[AS03] R. Agarwal and S. D. Stoller.

Type inference for parameterized race-free Java. Lecture Notes in Computer Science, 2937:149–160, 2003. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Artho:2004:JED

[ASB^{+}04] C. Artho, V. Schuppan, A. Biere, P. Eugster, M. Baur, and B. Zweimuller. JNuke: Efficient dynamic analysis for Java. Lecture Notes in Computer Science, 3114:462–465, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aldrich:2003:CSE

[ASCE03] Jonathan Aldrich, Emin Gün Sirer, Craig Chambers, and Susan J. Eggers. Comprehensive synchronization elimination for Java. Science of Computer Programming, 47(2–3): 91–120, May/June 2003. CO-DEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Aleksy:2003:DIB

[ASS03] M. Aleksy, M. Schader, and A. Schnell. Design and implementation of a bridge between CORBA’s Notification Service and the Java Message Service (STDOC02). Proceedings of the Annual Hawaii International Conference on System Sciences, CONF36: 318, 2003. CODEN ???? ISSN 1060-3425.

Alford:2005:IIJ

[ASS^{+}05] Torey Alford, Vijay P. Shah, Anthony Skjellum, Nicholas H. Younan, and Clayborne D. Taylor. inAspect: interfacing Java and VSIPL applications. Concurrency and Computation: Practice and Experience, 17(7–8):919– 940, June/July 2005. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ariga:2001:PSI

[AT01] Taeko Ariga and Hideki Tsuiki. Programming for students of information design. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(4):59–63, December 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp:

//ftp.math.utah.edu/pub/
mirrors/ftp.ira.uka.de/
bibliography/Misc/DBLP/
2001.bib.

Adl-Tabatabai:2003:SDC

[ATBC^{+}03] Ali-Reza Adl-Tabatabai, Jay Bharadwaj, Dong-Yuan Chen, Anwar Ghuloum, Vijay Menon, Brian Murphy, Mauricio Serrano, and Tatiana Shpeisman. The StarJIT compiler: a dynamic compiler for managed runtime environments. Intel Technology Journal, 7(1):19–31, February 2003. ISSN 1535-766X. URL

http://developer.intel.
com/technology/itj/2003/
volume07issue01/art02_starjit/
p01_abstract.htm.

Atkinson:2000:CPP

[Atk00] Leon Atkinson. Core PHP Programming: using PHP to build dynamic Web sites. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, second edition, 2000. ISBN 0-13-089398-6. xxv + 769 pp. LCCN QA76.73.P22 A85 2001.

Atkinson:2001:PJB

[Atk01] Malcolm Atkinson. Persistence and Java – A balancing act. Lecture Notes in Computer Science, 1944:1– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/1944/19440001.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/1944/19440001.
pdf.

Ahmed:2002:DEJ

[AU02] Khawar Zaman Ahmed and Cary E. Umrysh. Developing enterprise Java applications with J2EE and UML. Addison-Wesley, Reading, MA, USA, 2002. ISBN 0-201-73829-5. xxiv + 330 pp. LCCN QA76.73.J38 A35 2002.

Austin:2000:WAA

[Aus00] Calvin Austin. Writing advanced applications for the Java 2 platform. Addison-Wesley, Reading, MA, USA, 2000. ISBN 0-201-71501-5. xvi + 381 pp. LCCN A76.76.A65 A95 2000. US$44.95.

Avvenuti:2005:MUJ

[AV05] Marco Avvenuti and Alessio Vecchio. MobileRMI: upgrading Java Remote Method Invocation towards mobility. Software—Practice and Experience, 35(10):939–975, August 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Arnold:2008:QER

[AVY08] Matthew Arnold, Martin Vechev, and Eran Yahav. QVM: an efficient runtime for detecting defects in deployed systems. ACM SIGPLAN Notices, 43(10):143–162, September 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Arnow:2000:IPU

[AW00] David M. Arnow and Gerald Weiss. Introduction to programming using Java: an object-oriented approach. Addison-Wesley, Reading, MA, USA, 2000. ISBN 0-201-61272-0. xxxv + 805 pp. LCCN QA76.73.J38 A77 2000. Java 2 update.

Awhad:2003:UFS

[AW03] V. Awhad and C. Wallace.

A unified formal specification and analysis of the new Java memory models. Lecture Notes in Computer Science, 2589:166–185, 2003. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alistair:2004:SGS

[AWE04] M. Chalk Alistair, Martin Wennerberg, and L. L. Sonnhammer Erik. Sfixem-graphical sequence feature display in Java. Bioinformatics, 20(15): 2488–2490, October 12, 2004. CODEN ???? ISSN 1367-4803 (print), 1367-4811 (electronic).

Astrachan:2009:APC

[AWS^{+}09] Owen Astrachan, Henry Walker, Chris Stephenson, Lien Diaz, and Jan Cuny. Advanced placement computer science: the future of tracking the first year of instruction. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):397–398, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Ahern:2005:FJR

[AY05] Alexander Ahern and Nobuko Yoshida. Formalising Java RMI with explicit code mobility. ACM SIGPLAN Notices, 40(10):403–422, October 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ahern:2007:FJR

[AY07] Alexander Ahern and Nobuko Yoshida. Formalising Java RMI with explicit code mobility. Theoretical Computer Science, 389(3):341–410, December 15, 2007. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Ayers:2001:PJD

[Aye01] Danny Ayers. Professional Java data. Wrox Press, Chicago, IL, USA, 2001. ISBN 1-86100-410-9 (paperback). xxvii + 1320 pp. LCCN QA76.73.J38 P76 2001.

Allenstein:2008:QSS

[AYWM08] Brett Allenstein, Andrew Yost, Paul Wagner, and Joline Morrison. A query simulation system to illustrate database query execution. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):493–497, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Ancona:2001:TMJ

[AZ01] Davide Ancona and Elena Zucca. True modules for Java-like languages. Lecture Notes in Computer Science, 2072:354–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2072/20720354.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2072/20720354.
pdf.

Apte:2002:ETM

[AZ02] N. Apte and I. Zeid. Evolution of transparent manufacturing: An architecture for a Java-based controller of a CIM cell. Journal of Intelligent Manufacturing, 13(2): 89–100, April 2002. CODEN ???? ISSN 0956-5515.

Ancona:2004:PTJ

[AZ04] Davide Ancona and Elena Zucca. Principal typings for Java-like languages. ACM SIGPLAN Notices, 39(1): 306–317, January 2004. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Azizi:2006:BRJ

[Azi06] Mostafa Azizi. Book review: John Hunt and Chris Loftus, Guide to J2EE: Enterprise Java. Series Springer Professional Computing. Springer (2003). ISBN 1-85233-704-4. $79.95/£45.00/£69.95. 672 pp. The Computer Journal, 49(2):250–251, March 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/full/49/
                                        2/250;   http://comjnl.
                                        oxfordjournals.org/cgi/
                                        reprint/49/2/250.

Brewster:2001:CIH

[BA01] Neil V. Brewster and Tarek S. Abdelrahman. A compiler infrastructure for highperformance Java. Lecture Notes in Computer Science, 2110:675–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.
                                        com/link/service/series/
                                        0558/bibs/2110/21100675.
                                        htm;   http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2110/21100675.
                                        pdf.

Ben-Ari:2004:STT

[BA04] Mordechai Ben-Ari. A suite of tools for teaching concurrency. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):251, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bierhoff:2005:LOS

[BA05] Kevin Bierhoff and Jonathan Aldrich. Lightweight object specification with typestates. ACM SIGSOFT Software Engineering Notes, 30(5):217– 226, September 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bierhoff:2007:MTC

[BA07a] Kevin Bierhoff and Jonathan Aldrich. Modular typestate checking of aliased objects. ACM SIGPLAN Notices, 42(10):301–320, October 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Brosgol:2007:AOS

[BA07b] Ben Brosgol and Mario Aldea.

Ada and other standards: Introduction. ACM SIGADA Ada Letters, 27(2):88–89, August 2007. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Boehm:2008:FCC

[BA08] Hans-J. Boehm and Sarita V.

Adve. Foundations of the C++ concurrency memory model. ACM SIGPLAN Notices, 43(6):68–78, June 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bradel:2009:SPP

[BA09] Borys J. Bradel and Tarek S. Abdelrahman. A study of potential parallelism among traces in Java programs. Science of Computer Programming, 74(5–6):296–313, March 1, 2009. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Bacon:2001:KJD

[Bac01] David Bacon. Kava: a Java dialect with a uniform object model for lightweight classes. In ACM [ACM01b], pages 68–77. ISBN 1-58113-359-6. LCCN QA76.9.O35

A26 2001. URL http://www.
philippsen.com/JGI2001/
camerareadyabstracts/38.
html;   http://www.philippsen.
com/JGI2001/finalpapers/
18500068.ps.

Bacon:2003:KJD

[Bac03] David F. Bacon. Kava: a Java dialect with a uniform object model for lightweight classes. Concurrency and Computation: Practice and Experience, 15(3–5):185–206, March/April 2003. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bacon:2007:RGC

[Bac07] David F. Bacon. Realtime garbage collection. ACM Queue: Tomorrow’s Computing Today, 5(1):40–49, February 2007. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Badros:2000:JML

[Bad00] Greg J. Badros. JavaML:

a markup language for Java source code. Computer Networks (Amsterdam, Netherlands: 1999), 33(1–6):159– 177, June 2000. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/
10/15/22/48/25/37/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/15/22/48/25/
37/article.pdf;   http:
//www9.org/w9cdrom/342/
342.html.

Bocchino:2009:TES

[BAD^{+}09] Robert L. Bocchino, Jr., Vikram S. Adve, Danny Dig, Sarita V. Adve, Stephen Heumann, Rakesh Komuravelli, Jeffrey Overbey, Patrick Simmons, Hyojin Sung, and Mohsen Vakilian. A type and effect system for deterministic parallel Java. ACM SIG-PLAN Notices, 44(10):97– 116, October 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bellamy:2008:ELT

[BAdMS08] Ben Bellamy, Pavel Avgustinov, Oege de Moor, and Damien Sereni. Efficient local type inference. ACM SIGPLAN Notices, 43(10):475–492, September 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bauer:2003:MSM

[BAF03] Lujo Bauer, Andrew W. Appel, and Edward W. Felten. Mechanisms for secure modular programming in Java. Software—Practice and Experience, 33(5):461–480, April 25, 2003. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bagnall:2002:CLM

[Bag02] Brian Bagnall. Core LEGO Mindstorms Programming: Unleash the Power of The Java Platform. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2002. ISBN 0-13-009364-5. LCCN TJ211 .B32 2002. US$39.99. URL http:/

                                        /www.phptr.com/ptrbooks/
                                        ptr_0130093645.html.

Bailey:2000:JEP

[Bai00] Duane A. Bailey. Java elements: principles of programming in Java. McGraw-Hill, New York, NY, USA, 2000. ISBN 0-07-228357-2. xiv + 333 pp. LCCN QA76.73.J38 B33 2000. Includes CD-ROM.

Bailey:2003:JSD

[Bai03] Duane A. Bailey. Java structures: data structures in Java for the principled programmer. McGraw-Hill, New York, NY, USA, second edition, 2003. ISBN 0-07-239909-0, 0-07-112163-3. xiv + 514 pp. LCCN QA76.73.J38 B34 2003.

Bratthall:2001:PUB

[BAJ01] Lars Bratthall, Erik Arisholm, and Magne Jørgensen. Program understanding behavior during estimation of enhancement effort on small Java programs. Lecture Notes in Computer Science, 2188:356–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2188/21880356.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2188/21880356.
pdf.

Baker:2000:SIM

[Bak00] Mark Baker. Special issue:

Message passing interfacebased parallel programming with Java. Concurrency: Practice and Experience, 12 (11):1017–1018, September 2000. CODEN CPEXEI.

ISSN 1040-3108. URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract/76000190/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=76000190&PLACEBO=IE.
                                        pdf.

Bacon:2001:JCB

[BAL^{+}01] David F. Bacon, Clement R. Attanasio, Han B. Lee, V. T. Rajan, and Stephen Smith. Java without the coffee breaks: a nonintrusive multiprocessor garbage collector. ACM SIGPLAN Notices, 36(5):92–103, May 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bales:2002:JPO

[Bal02] Donald Bales. Java Programming with Oracle JDBC. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2002. ISBN 0-596-00088-X. xv + 477 pp. LCCN QA76.73.J38 B35 2002. US$39.95. URL http:/

                                        /safari.oreilly.com/059600088X;
                                         http://www.oreilly.com/
                                        catalog/jorajdbc.

Bal:2003:IJB

[Bal03a] H. E. Bal. Ibis: a Javabased Grid programming environment. Lecture Notes in Computer Science, 2790: 4, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bales:2003:JPR

[Bal03b] Donald Bales. JDBC pocket reference. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2003. ISBN 0-596-00457-5. vi + 145 pp. LCCN QA76.73.J38 B355 2003. URL http://

www.oreilly.com/catalog/
9780596004576.

Ballance:2003:BRJ

[Bal03c] R. Ballance. Book review:

Java Precisely, by Peter Seetoft, MIT Press, 2002, $14.95, ISBN 0-262-69276-7. ACM Queue: Tomorrow’s Computing Today, 1(6): 70, September/October 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Brecht:2001:CGC

[BALP01] Tim Brecht, Eshrat Arjomandi, Chang Li, and Hang Pham. Controlling garbage collection and heap growth to reduce the execution time of Java applications. ACM SIG-PLAN Notices, 36(11):353– 366, November 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Proceedings of the 2001 ACM SIGPLAN Conference on Object Oriented Programming, Systems, Languages and Applications (OOPSLA’01).

Brecht:2006:CGC

[BALP06] Tim Brecht, Eshrat Arjomandi, Chang Li, and Hang Pham. Controlling garbage collection and heap growth to reduce the execution time of Java applications. ACM Transactions on Programming Languages and Systems, 28(5):908–941, September 2006. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Bollinger:2003:BFF

[BALV03] Terry Bollinger, John J. Abbott, Robert C. Larrabee, and Carl Vieregger. Bookshelf: Facts and fantasies: a review of two books; understanding the unified process; software architecture: Project planning and control; paring Java patterns. IEEE Software, 20 (2):80–85, March/April 2003. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:

//dlib.computer.org/so/
books/so2003/pdf/s2080.

pdf. Review of A New Kind of Science by Stephen Wolfram, Prey: A Novel by Michael Crichton, The Unified Process Explained by Kendall Scott, Effective Work Breakdown Structures by Gregory T. Haugan, Applied Java Patterns by Stephen Stelting and Olav Maassen.

Baran:2000:NVN

[Bar00a] Nicholas Baran. News and views: Neural net crowned king; smart shirts monitor vital signs; McMaster team wins college design contest; Java applet tracks space station; Sandia Red Team batting 1000; new digital geometry compression algorithm. Dr. Dobb’s Journal of Software Tools, 25(10):18, October 2000. CODEN DDJOEB. ISSN 1044-789X.

Barnes:2000:OOP

[Bar00b]  David Barnes.  Object-

Oriented Programming with Java: an Introduction. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2000. ISBN 0-13-086900-7. xxvii + 1028 pp. LCCN QA76.64 .B38 2000. US$56.

Barrilleaux:2000:UIJ

[Bar00c] Jon Barrilleaux. 3D User Interfaces With Java 3D. Manning Publications, Greenwich, CT, USA, 2000. ISBN 1-884777-90-2. xxvi + 499 pp. LCCN QA76.9.U83 B363 2001. US$45.

Baran:2001:NVA

[Bar01a] Nicholas Baran. News and views: Anonymity and the Internet; is industry hogging computer science talent?; relief from acronyms; OpenML spec released; C# not just a Java killer, says anders; and the winners are …. Dr.

Dobb’s Journal of Software Tools, 26(7):18, July 2001. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/.

Baran:2001:NVC

[Bar01b] Nicholas Baran. News and views: College Board to add Java test; astronomy project beats out algorithm for science prize; consortium to release XML business specification; open source developer agreement proposed; consortium hopes to eliminate computer failures; new linear algorithm for sequence analysis; distributing computing à la SETI catches on. Dr. Dobb’s Journal of Software Tools, 26(3):18, March 2001. CODEN DDJOEB. ISSN 1044-789X. URL http:

//publicscience.net/;
http://www.ddj.com/.

Baran:2001:NVM

[Bar01c] Nicholas Baran. News and views: Mail server benchmark released; new traveling salesman algorithm; ACLU challenges ICANN on freedom of speech; Microsoft opts for decaf; store your data on the Internet; and mind your electronic trash. Dr. Dobb’s Journal of Software Tools, 26(4):18, April 2001. CODEN DDJOEB. ISSN 1044-789X. URL

http://www.cs.wustl.edu/
;   http://www.ddj.com/;

http://www.spec.org/osg/
mail2001/docs/FAQ.html;
http:/www.research.att.
com/~alb/ALENEX01.

Barros:2001:UPN

[Bar01d] João Paulo Barros. Use Petri nets to improve your concurrent programming course (poster session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3): 192, September 2001. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Barish:2002:BSH

[Bar02a] Greg Barish. Building scalable and high-performance Java Web applications using J2EE technology. Addison-Wesley, Reading, MA, USA, 2002. ISBN 0-201-72956-3. xviii + 392 pp. LCCN QA76.73.J38 B36 2002.

Barnes:2002:TIJ

[Bar02b] David J. Barnes. Teaching introductory Java through LEGO MINDSTORMS models. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):147–151, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Barake:2003:BRE

[Bar03a] Tony Barake. Book review: The enterprise goes to school: a review of Wireless Java Programming for Enterprise Applications: Mobile Devices Go Corporate by Dan Harkey, Shan Appajodu, and Mike Larkin. IEEE Distributed Systems Online, 4(7), 2003. ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:

//dsonline.computer.org/
0307/d/bks_a.htm.

Barker:2003:BJO

[Bar03b] Jacquie Barker. Beginning Java Objects: From Concepts to Code. Apress, Berkeley, CA, USA, 2003. ISBN 1-59059-146-1. 688 (est.) pp. LCCN ????

Barrett:2003:DPJ

[Bar03c]  Tom Barrett.  Dynamic

proxies in Java and .NET. Dr. Dobb’s Journal of Software Tools, 28(7):18, 20, 22, 24, 26, July 2003. CO-DEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/ftp/2003/
2003_07/dyproxy.zip.

Bardram:2005:JCA

[Bar05] J. E. Bardram. The Java Context Awareness Framework (JCAF) — A service infrastructure and programming framework for contextaware applications. Lecture Notes in Computer Science, 3468:98–115, 2005. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bardram:2009:ABC

[Bar09] Jakob E. Bardram. Activitybased computing for medical work in hospitals. ACM Transactions on Computer-Human Interaction, 16(2): 10:1–10:??, June 2009. CO-DEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic).

Bathelt:2003:JID

[Bat03] M. Bathelt. Java infotainment devices. VDI Berichte, 1789(??):1099–1114, 2003. CODEN ???? ISSN 0083-5560.

Batov:2004:JGC

[Bat04] V. Batov. Java generics and C++ templates. C/C++ Users Journal, 22(7):16–21, 2004. CODEN CCUJEX. ISSN 1075-2838.

Bishop:2000:JGE

[BB00a] J. M. (Judith Mary) Bishop and Nigel Bishop. Java gently for engineers and scientists. International computer science series. Addison-Wes-ley, Reading, MA, USA, 2000. ISBN 0-201-34304-5. xix + 436 pp. LCCN QA76.76.J38 B57 2000.

Bishop:2000:OOJ

[BB00b] Judith Bishop and Nigel Bishop. Object-orientation in Java for scientific programmers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):357–361, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bigus:2001:CIA

[BB01] Joseph P. Bigus and Jennifer Bigus. Constructing intelligent agents using JAVA. Professional developer’s guide series. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, second edition, 2001. ISBN 0-471-39601-X. xxii + 408 pp. LCCN QA76.76.I58 B563 2001.

Bruhn:2003:ATJ

[BB03] R. E. Bruhn and P. J. Burton. An approach to teaching Java using computers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(4):94–99, 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bergstra:2005:NAJ

[BB05] J. A. Bergstra and I. Bethke. Network algebra in Java. Journal of Logic and Algebraic Programming, 62(2): 247–271, 2005. CODEN ???? ISSN 1567-8326.

Beckman:2008:VCU

[BBA08] Nels E. Beckman, Kevin Bierhoff, and Jonathan Aldrich.

Verifying correct usage of atomic blocks and typestate. ACM SIGPLAN Notices, 43(10):227–244, September 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Barisone:2001:JSM

[BBBD01] A. Barisone, F. Bellotti, R. Berta, and A. De Gloria. JSBricks: a suite of microbenchmarks for the evaluation of Java as a scientific execution environment. Future Generation Computer Systems, 18(2): 293–306, October 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/60/31/35/abstract.
html.

Baduel:2007:ATO

[BBC07] Laurent Baduel, Françoise Baude, and Denis Caromel. Asynchronous typed object groups for Grid programming. International Journal of Parallel Programming, 35 (6):573–614, December 2007. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=35&
issue=6&spage=573.

Barbuti:2002:FJB

[BBDT02] R. Barbuti, C. Bernardeschi, N. De Francesco, and L. Tesei. Fixing the Java bytecode verifier by a suitable type domain. In Genoveffa Tortora and S. K. (Shi Kuo) Chang, editors, Proceedings of the 14th international conference on Software engineering and knowledge engineering: 2002, Ischia, Italy, July 15–19, 2002, pages 377–382. ACM Press, New York, NY 10036, USA, 2002. ISBN 1-58113-556-4. LCCN QA76.758 .I573 2002.

URL http://portal.acm.
org/toc.cfm?id=568760.

Bellotti:2004:EOM

[BBG04] Francesco Bellotti, Riccardo Berta, and Alessandro De Gloria. Evaluation and optimization of method calls in Java. Software—Practice and Experience, 34(4): 395–431, April 10, 2004. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bellotti:2001:DJA

[BBGP01] Francesco Bellotti, Riccardo Berta, Alessandro De Gloria, and Andrea Poggi. DirectJ: Java APIs for optimized 2D graphics. Software—Practice and Experience, 31(3): 259–275, March 2001. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (elec-

tronic). URL http://www3.
interscience.wiley.com/
cgi-bin/abstract/76507637/

START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=76507637&PLACEBO=IE.
pdf.

Bischof:2001:HTU

[BBHL01] Christian H. Bischof, H. Martin Bücker, Jörg Henrichs, and Bruno Lang. Hands-on training for undergraduates in high-performance computing using Java. Lecture Notes in Computer Science, 1947:306– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/1947/19470306.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/1947/19470306.
pdf.

Benander:2003:PJE

[BBL03] A. C. Benander, B. A. Benander, and M. Lin. Perceptions of Java — experienced programmers’ perspective. The Journal of Computer Information Systems, 43(4):1–7, 2003. CODEN JCISE9. ISSN 0887-4417.

Barros:2004:PMD

[BBM04] T. Barros, R. Boulifa, and E. Madelaine. Parameterized models for distributed Java objects. Lecture Notes in Computer Science, 3235:43– 60, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benander:2004:FRD

[BBS04] A. Benander, B. Benander, and J. Sang. Factors related to the difficulty of learning to program in Java — an empirical study of nonnovice programmers. Information and Software Technology, 46(2):99–107, 2004. CO-DEN ISOTE7. ISSN 0950-5849 (print), 1873-6025 (electronic).

Brackeen:2003:DGJ

[BBV03] David Brackeen, Bret Barker, and Laurence Vanhelsuwe. Developing Games in Java. New Riders Publishing, Carmel, IN, USA, 2003. ISBN 1-59273-005-1. xxxii + 972 pp. LCCN QA76.76.C672 B74 2004. US$49.99.

Barabash:2005:PIM

[BBYG^{+}05] Katherine Barabash, Ori Ben-Yitzhak, Irit Goft, Elliot K. Kolodner, Victor Leikehman, Yoav Ossia, Avi Owshanko, and Erez Petrank. A parallel, incremental, mostly concurrent garbage collector for servers. ACM Transactions on Programming Languages and Systems, 27(6):1097– 1146, November 2005. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Baker:2000:MPJ

[BC00] Mark Baker and Bryan Carpenter. MPJ: a proposed Java message passing API and environment for high performance computing. Lecture Notes in Computer Science, 1800:552–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/1800/18000552.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/1800/18000552.
pdf.

Bettini:2001:JNC

[BC01] Lorenzo Bettini and Donato Cappetta. A Java 2 network class loader. Dr. Dobb’s Journal of Software Tools, 26 (2):58, 60, 62, 64, February 2001. CODEN DDJOEB. ISSN 1044-789X. URL

http://www.ddj.com/ftp/
2001/2001_02/jload.txt;
http://www.ddj.com/ftp/
2001/2001_02/jload.zip.

Burke:2003:JEP

[BC03] Eric M. Burke and Brian M.

Coyner. Java extreme programming cookbook. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2003. ISBN 0-596-00387-0. xii + 275 pp. LCCN QA76.73.J38 B873 2003.

Boyer:2004:IIT

[BC04] Truman Parks Boyer and Mohsen Chitsaz. ICETM and ICE/TTM: tools to assist in compiler design and implementation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(4):55–57, December 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp:

//ftp.math.utah.edu/pub/
mirrors/ftp.ira.uka.de/
bibliography/Misc/DBLP/
2004.bib.

Bagley:2007:CIN

[BC07] Carole A. Bagley and C. Candace Chou. Collaboration and the importance for novices in learning Java computer programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):211– 215, September 2007. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Bainbridge:2001:CEJ

[BCCN01] A. Bainbridge, J. Colgrave, A. Colyer, and G. Normington. CICS and Enterprise JavaBeans. IBM Systems Journal, 40(1):46– 67, ???? 2001. CODEN IBMSA7. ISSN 0018-8670.

URL http://www.research.
ibm.com/journal/sj/401/
bainbridge.html.

Barthe:2002:TAS

[BCDdS02] Gilles Barthe, Pierre Courtieu, Guillaume Dufay, and Simão Melo de Sousa. Tool-assisted specification and verification of the JavaCard platform. Lecture Notes in Computer Science, 2422:41–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2422/24220041.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2422/24220041.
pdf.

Bieber:2001:PPT

[BCE^{+}01] P. Bieber, J. Cazin, A. El Marouani, P. Girard, J.- L. Lanet, V. Wiels, and G. Zanon. The PACAP prototype: a tool for detecting Java card illegal flow. Lecture Notes in Computer Science, 2041:25–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2041/20410025.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2041/20410025.
pdf.

Biegel:2002:DPB

[BCH02] Gregory Biegel, Vinny Cahill, and Mads Haahr. A dynamic proxy based architecture to support distributed Java objects in a mobile environment. Lecture Notes in Computer Science, 2519:809– 826, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/
                                        service/series/0558/bibs/
                                        2519/25190809.htm;   http:
                                        //link.springer.de/link/
                                        service/series/0558/papers/
                                        2519/25190809.pdf.

Biernacki:2008:CDM

[BCHP08] Dariusz Biernacki, Jean-Louis Colaço, Gregoire Hamon, and Marc Pouzet. Clock-directed modular code generation for synchronous data-flow languages. ACM SIGPLAN Notices, 43(7):121–130, July 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bruneton:2006:FCM

[BCL^{+}06] Eric Bruneton, Thierry Coupaye, Matthieu Leclercq, Vivien Quéma, and Jean-Bernard Stefani. The FRAC-TAL component model and its support in Java. Software —Practice and Experience, 36 (11–12):1257–1284, September/October 2006. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Blackburn:2004:MRP

[BCM04] Stephen M. Blackburn, Perry Cheng, and Kathryn S. McKinley. Myths and realities: the performance impact of garbage collection. ACM SIGMETRICS Perform. Eval. Rev., 32(1):25– 36, June 2004. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Beck:2005:CLT

[BCM05] Leland L. Beck, Alexander W. Chizhik, and Amy C. McElroy. Cooperative learning techniques in CS1: design and experimental evaluation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):470–474, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Baldoni:2003:PAJ

[BCMT03] R. Baldoni, S. Cimmino, C. Marchetti, and A. Termini. Performance analysis of Java group toolkits: a case study. Lecture Notes in Computer Science, 2604:49– 60, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bacon:2003:CFS

[BCR03a] David F. Bacon, Perry Cheng, and V. T. Rajan. Controlling fragmentation and space consumption in the Metronome, a real-time garbage collector for Java. ACM SIGPLAN Notices, 38(7):81–92, July 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Burdy:2003:DFV

[BCR03b] L. Burdy, L. Casset, and A. Requet. Développement formel d’un verifieur embarqué de byte-code Java. (French) [Formal development of a Java-based bytecode verifier]. Technique et science informatiques : TSI, 22(1):33–60, 2003. CODEN TTSIDJ. ISSN 0752-4072, 0264-7419.

Bellavista:2002:JLD

[BCS02] Paolo Bellavista, Antonio Corradi, and Cesare Stefanelli. Java for on-line distributed monitoring of heterogeneous systems and services. The Computer Journal, 45(6):595–607, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        www3.oup.co.uk/computer_
                                        journal/hdb/Volume_45/Issue_
                                        06/450595.sgm.abs.html;
                                         http://www3.oup.co.uk/
                                        computer_journal/hdb/Volume_
                                        45/Issue_06/pdf/450595.
                                        pdf.

Baker:2007:BLS

[BCS07] Mark Baker, Bryan Carpenter, and Aamir Shafi. A buffering layer to support derived types and proprietary networks for Java HPC.

Scalable Computing: Practice and Experience, 8(4):343– 358, December 2007. CO-DEN ???? ISSN 1895-

1767. URL http://www.
scpe.org/vols/vol08/no4/
SCPE_8_4_02.pdf;   http:
//www.scpe.org/vols/vol08/
no4/SCPE_8_4_02.zip.

Bertoli:2009:JPE

[BCS09] Marco Bertoli, Giuliano Casale, and Giuseppe Serazzi. JMT: performance engineering tools for system modeling. ACM SIGMETRICS Perform. Eval. Rev., 36(4): 10–15, March 2009. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bettini:2003:EJD

[BCV03] L. Bettini, S. Capecchi, and B. Venneri. Extending Java to dynamic object behaviors. Electronic Notes in Theoretical Computer Science, 82(8): 1–20, 2003. CODEN ???? ISSN 1571-0661.

Bettini:2009:FJD

[BCV09] Lorenzo Bettini, Sara Capecchi, and Betti Venneri. Featherweight Java with dynamic and static overloading. Science of Computer Programming, 74(5–6):261– 278, March 1, 2009. CO-DEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Bredlau:2001:ALT

[BD01a] Carl Bredlau and Dorothy Deremer. Assembly language through the Java Virtual Machine. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):194–198, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Brosgol:2001:RTC

[BD01b] Ben Brosgol and Brian Dobbing. Real-time convergence of Ada and Java^{T M}. ACM SIGADA Ada Letters, 21(4): 11–26, December 2001. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Brosgol:2001:CJR

[BD01c] Benjamin Brosgol and Brian Dobbing. Can Java^{T M} meet its real-time deadlines? Lecture Notes in Computer Science, 2043:68– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.
                                        com/link/service/series/
                                        0558/bibs/2043/20430068.
                                        htm;   http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2043/20430068.
                                        pdf.

Bernardeschi:2002:CAI

[BD02] Cinzia Bernardeschi and Nicoletta De Francesco. Combining abstract interpretation and model checking for analysing security properties of Java bytecode. Lecture Notes in Computer Science, 2294:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2294/22940001.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2294/22940001.
pdf.

Badeen:2003:MCM

[BD03a] C. Badeen and M. A. Dube.

Modeling the copolymerization of methyl Methacrylate/ alpha-methyl styrene at elevated temperatures using Java. Polymer Reaction Engineering, 11(1):53–78, 2003. CODEN ???? ISSN 1054-3414.

Bettini:2003:JMG

[BD03b] L. Bettini and R. DeNicola. A Java middleware for guaranteeing privacy of distributed tuple spaces. Lecture Notes in Computer Science, 2604:175– 184, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brittain:2003:TDG

[BD03c] Jason Brittain and Ian F.

Darwin. Tomcat: the definitive guide. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2003. ISBN 0-596-00318-8. xiii + 303 pp. LCCN TK5105.8885.A63 B75 2003. URL http://

                                        www.oreilly.com/catalog/
                                        9780596003180.

Bieg:2004:ETD

[BD04] C. Bieg and S. Diehl. Educational and technical design of a Web-based interactive tutorial on programming in Java. Science of Computer Programming, 53(1):25–36, 2004. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Brittain:2007:TDG

[BD07] Jason Brittain and Ian F. Darwin. Tomcat: the definitive guide. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2007. ISBN 0-596-10106-6 (paperback). xvi + 476 pp. LCCN TK5105.8885.A63 B75 2008. URL http://

                                        www.oreilly.com/catalog/
                                        9780596101060.

Brown:2003:SFE

[BDE^{+}03] Rhodes Brown, Karel Driesen, David Eng, Laurie Hendren, John Jorgensen, Clark Verbrugge, and Qin Wang. STEP: a framework for the efficient encoding of general trace data. ACM SIGSOFT Software Engineering Notes, 28(1):27–34, January 2003. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Baylor:2000:JSB

[BDF^{+}00] S. J. Baylor, M. Devarakonda, S. Fink, E. Gluzberg, M. Kalantar, P. Muttineni, E. Barsness, R. Arora, R. Dimpsey, and S. J. Munroe. Java server benchmarks. IBM Systems Journal, 39(1):57–81, ???? 2000. CODEN IBMSA7. ISSN 0018-8670.

URL http://www.almaden.
ibm.com/journal/sj/391/
baylor.html.

Bonifaci:2004:JBS

[BDFL04] V. Bonifaci, C. Demetrescu, I. Finocchi, and L. Laura. A Java-based system for building animated presentations over the Web. Science of Computer Programming, 53(1):37–49, 2004. CO-DEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Barthe:2001:JTR

[BDHdS01] G. Barthe, G. Dufay, M. Huisman, and S. Melo de Sousa. Jakarta: a toolset for reasoning about JavaCard. Lecture Notes in Computer Science, 2140:2–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2140/21400002.
htm;   http://link.springer-

ny.com/link/service/series/
0558/papers/2140/21400002.
pdf.

Barthe:2001:FES

[BDJ^{+}01a] Gilles Barthe, Guillaume Dufay, Line Jakubiec, Bernard Serpette, and Simão Melo de Sousa. A formal executable semantics of the JavaCard platform. Lecture Notes in Computer Science, 2028:302– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2028/20280302.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2028/20280302.
pdf.

Bourdonov:2001:JSE

[BDJ^{+}01b] Igor B. Bourdonov, Alexey V. Demakov, Andrew A. Jarov, Alexander S. Kossatchev, Victor V. Kuliamin, Alexander K. Petrenko, and Sergey V. Zelenov. Java specification extension for automated test development. Lecture Notes in Computer Science, 2244:301–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2244/22440301.
htm;   http://link.springer-
ny.com/link/service/series/

0558/papers/2244/22440301.
pdf.

Barthe:2002:FCB

[BDJdS02] Gilles Barthe, Guillaume Dufay, Line Jakubiec, and Simão Melo de Sousa. A formal correspondence between offensive and defensive JavaCard virtual machines. Lecture Notes in Computer Science, 2294:32– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2294/22940032.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2294/22940032.
pdf.

Bernardeschi:2008:DBV

[BDL^{+}08] C. Bernardeschi, N. De Francesco, G. Lettieri, L. Martini, and P. Masci. Decomposing bytecode verification by abstract interpretation. ACM Transactions on Programming Languages and Systems, 31(1):3:1–3:63, December 2008. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Bernardeschi:2004:CSI

[BDLM04] C. Bernardeschi, N. De Francesco, G. Lettieri, and L. Martini. Checking secure information flow in Java bytecode by code transformation and standard bytecode verification. Software —Practice and Experience, 34(13):1225–1255, November 10, 2004. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bergel:2005:CJC

[BDN05] Alexandre Bergel, Stéphane Ducasse, and Oscar Nierstrasz. Classbox/J: controlling the scope of change in Java. ACM SIGPLAN Notices, 40(10):177–189, October 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bettini:2002:KJP

[BDP02] L. Bettini, R. DeNicola, and R. Pugliese. KLAVA: a Java package for distributed and mobile applications. Software —Practice and Experience, 32 (14):1365–1394, 2002. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bellotti:2001:AJG

[BDRV01] F. Bellotti, A. De Gloria, M. Risso, and A. Villamaina. AutoGraL: a Java 2D graphics library for configurable automotive dashboards. Computers and Graphics, 25(2): 259–268, April 2001. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic). URL http:/

                                        /www.elsevier.nl/gej-ng/
                                        10/13/20/57/29/33/abstract.

html;   http://www.elsevier.
nl/gej-ng/10/13/20/57/29/
33/article.pdf.

Bonachea:2001:HPF

[BDT01] Dan Bonachea, Phillip Dickens, and Rajeev Thakur. High-performance file I/O in Java: Existing approaches and bulk I/O extensions. Concurrency and Computation: Practice and Experience, 13(8–9):713–736, July/August 2001. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (elec-

tronic). URL http://www3.
interscience.wiley.com/
cgi-bin/abstract/84503223/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=84503223&PLACEBO=IE.
pdf.

Barbuti:2004:AIJ

[BDT04] R. Barbuti, N. De Francesco, and L. Tesei. Abstract interpretation for Java bytecode verification. Technical report, Dipartimento di Informatica, University of Pisa, Pisa, Italy, 2004.

Burrows:2002:JGE

[BE02] Anthony L. Burrows and David England. Java 3D, 3D graphical environments and behaviour. Software—Practice and Experience, 32(4): 359–376, April 10, 2002. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (elec-

tronic). URL http://www3.

                                        interscience.wiley.com/
                                        cgi-bin/abstract/91016434/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=91016434&PLACEBO=IE.
                                        pdf.

Beatty:2005:FYW

[Bea05] Andrew Beatty. Feeling your way in Java: An essay on society and emotion. Ethnos, 70 (1):53–78, March 2005. CO-DEN ???? ISSN 0014-1844.

Becker:2000:JSCa

[Bec00a] Pete Becker. The journeyman’s shop: Common design mistakes, part 1. C/C++ Users Journal, 18(1):??, January 2000. CODEN CCU-JEX. ISSN 1075-2838.

Becker:2000:JSCb

[Bec00b] Pete Becker. The journeyman’s shop: Common design mistakes, part 2. C/C++ Users Journal, 18(2):77–??, February 2000. CODEN CCUJEX. ISSN 1075-2838.

Becker:2001:TCK

[Bec01a] Byron Weber Becker. Teaching CS1 with Karel the robot in Java. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):50–54, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Becker:2001:SMW

[Bec01b] Thomas Becker. Synchronization monitors for Win32.

Dr. Dobb’s Journal of Software Tools, 26(12):46, 48, 50–52, 54, December 2001. CODEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/ftp/2001/
2001_12/monitor.txt;   http:
//www.ddj.com/ftp/2001/
2001_12/monitor.zip.

Beckert:2001:DLF

[Bec01c] Bernhard Beckert. A dynamic logic for the formal verification of Java card programs. Lecture Notes in Computer Science, 2041:6– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2041/20410006.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2041/20410006.
pdf.

Beck:2004:JPG

[Bec04] Kent Beck. JUnit pocket guide. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2004. ISBN 0-596-00743-4 (paperback). vi + 84 pp. LCCN QA76.76.T48 B43 2004. URL http://

www.oreilly.com/catalog/
9780596007430.

Beebe:2000:BPAa

[Bee00] Nelson H. F. Beebe. A bibliography of publications about the Java Programming Language: 2000–date. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, March 11, 2000. 13 pp. URL https:

//www.math.utah.edu/pub/
tex/bib/index-table-j.html#

java2000. This report is updated frequently.

Beebe:2004:CJR

[Bee04a] Nelson H. F. Beebe. Comments on the Java Random class. Web document, March 24, 2004.

URL https://www.math.
utah.edu/~beebe/java/random/

. This document examines Java support for random numbers, comments on its deficiencies and inefficiencies, and reports the results of two test suites.

Beebe:2004:JPF

[Bee04b] Nelson H. F. Beebe. Java programming: Fun with Fibonacci. World-Wide Web document, March 2004.

URL https://www.math.
utah.edu/~beebe/software/
java/fibonacci/. This re-

port summarizes the origin of the Fibonacci sequence, giving the full Latin text from the original book written in 1202 (not previously available on the Web). Computation of the Fibonacci sequence, and its term ratios, is implemented in about 50 different programming languages. The report comments on the relative difficulty of the task in some of those languages, and on their suitability for numerical computation. It also provides a complete floatingpoint formatted output package for Java.

Bell:2002:VBN

[Bel02] Douglas Bell. Visual basic.

net as a first language: an evaluation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(4):107–108, December 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp:

//ftp.math.utah.edu/pub/
mirrors/ftp.ira.uka.de/
bibliography/Misc/DBLP/
2002.bib.

Benson:2000:JR

[Ben00a] Brent W. Benson, Jr. Java reflections. ACM SIGPLAN Notices, 35(12):14–18, December 2000. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Benson:2000:JRJ

[Ben00b] Brent W. Benson, Jr. Java reflections: Java server pages: Servlets, inside out. ACM SIGPLAN Notices, 35(12): 14–18, December 2000. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Benson:2000:JRS

[Ben00c] Brent W. Benson, Jr. Java reflections: Servlets: Webbased applications you can live with. ACM SIGPLAN Notices, 35(3):21–24, March 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Berg:2000:AJD

[Ber00a]  Clifford J. Berg.  Ad-

vanced Java 2 Development for Enterprise Applications. Sun Microsystems Press Java series. Sun Microsystems Press, Palo Alto, CA, USA, 2000. ISBN 0-13-084875-1. xxiv + 785 pp. LCCN QA76.73.J38B478

                                        2000. URL http://www.sun.
                                        com/books/catalog/berg/.

Bertelsen:2000:DSJ

[Ber00b] Peter Bertelsen. Dynamic semantics of Java bytecode. Future Generation Computer Systems, 16(7): 841–850, May 2000. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/41/30/32/abstract.
                                        html.

Bergsten:2001:JP

[Ber01a] Hans Bergsten. JavaServer pages. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2001. ISBN 1-56592-746-X. xviii + 552 pp. LCCN TK5105.8885.J38 B37 2001. Editors: Robert Eckstein and Paula Ferguson.

Bergsten:2001:JPP

[Ber01b] Hans Bergsten. JavaServer pages pocket reference. Java series. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2001. ISBN 0-596-00231-9. v + 81 pp. LCCN QA76.73.J38.b48 2001. Server-side Java development.

Bertot:2001:FJV

[Ber01c]  Yves Bertot.  Formaliz-

ing a JVML verifier for initialization in a theorem prover. Lecture Notes in Computer Science, 2102:14– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2102/21020014.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2102/21020014.
pdf.

Bergsten:2002:JP

[Ber02a] Hans Bergsten. JavaServer Pages. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2002. ISBN 0-596-00317-X. xxii + 660 pp. LCCN TK5105.8885.J38 B47 2002. US$44.95. URL http://

                                        safari.oreilly.com/059600317X;
                                         http://www.oreilly.com/
                                        catalog/jserverpages2.

Bergstra:2002:MOP

[Ber02b]  J. Bergstra.  Molecule-

oriented programming in Java. Information and Software Technology, 44(11):617– 638, August 15, 2002. CO-DEN ISOTE7. ISSN 0950-5849 (print), 1873-6025 (electronic).

Bergsten:2004:JF

[Ber04a] Hans Bergsten. JavaServer faces. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2004. ISBN 0-596-00539-

3. xiv + 589 pp. LCCN TK5105.8885.J38 B46 2004.

Bergsten:2004:JP

[Ber04b] Hans Bergsten. JavaServer pages. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, third edition, 2004. ISBN 0-596-00563-6. xxii + 740 pp. LCCN TK5105.8885.J38 B47 2004.

Bergin:2005:AJ

[Ber05a] Joseph Bergin. Academic jeopardy. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):389, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Berzal:2005:JTF

[Ber05b] F. Berzal. Java tools and frameworks for Web application development. IEEE Distributed Systems Online, 6(5):5, May 2005. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://ieeexplore.ieee.
org/iel5/8968/31212/01453474.
pdf?isnumber=31212&prod=
JNL&arnumber=1453474&arSt=
+5&ared=+5&arAuthor=+Berzal%
2C+F.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=31212&arnumber=
1453474&count=4&index=2.

Bergin:2006:KUD

[Ber06]  Joe Bergin.  Karel uni-

verse drag & drop editor. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38 (3):307, September 2006. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Besset:2001:OOI

[Bes01] Didier H. Besset. Objectoriented implementation of numerical methods: an introduction with Java and Smalltalk. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2001. ISBN 1-55860-679-3. xxv + 766 pp. LCCN QA76.76.D47 B475 2001.

Betz:2002:BMN

[Bet02]  David Betz.  Bob meets

Nuon. Dr. Dobb’s Journal of Software Tools, 27(2): 60, 62–64, February 2002. CODEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/ftp/2002/
2002_02/bob.zip;   http:
//www.ddj.com/ftp/2002/
2002_02/bobnuon.txt.

Bettini:2004:JPC

[Bet04] L. Bettini. A Java package for class and mixin mobility in a distributed setting. Lecture Notes in Computer Science, 2952:12–22, 2004. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bettini:2005:JPT

[Bet05] L. Bettini. A Java package for transparent code mobility. Lecture Notes in Computer Science, 3409:112–122, 2005. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boian:2002:ACT

[BF02] Florian Mircea Boian and Corina Ferdean. Advanced collaboration techniques between Java objects distributed on clusters. Lecture Notes in Computer Science, 2326:259–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2326/23260259.

htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2326/23260259.
pdf.

Bertie:2003:TCI

[BF03] A. Bertie and P. Farrington.

Teaching confidence intervals with Java applets. Teaching Statistics, 25(3):70–75, 2003. CODEN ???? ISSN 0141-982X (print), 1467-9639 (electronic).

Bruce:2004:LWL

[BF04] K. B. Bruce and J. N. Foster. LOOJ: Weaving LOOM into Java. Lecture Notes in Computer Science, 3086:389– 413, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bacon:2002:STE

[BFG02] David F. Bacon, Stephen J.

Fink, and David Grove. Space- and time-efficient implementation of the Java object model. Lecture Notes in Computer Science, 2374:111– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2374/23740111.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2374/23740111.
pdf.

Basin:2003:BVM

[BFG03] David A. Basin, Stefan Friedrich, and Marek Gawkowski. Bytecode verification by model checking. Journal of Automated Reasoning, 30(3– 4):399–444, May 2003. CO-DEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A%3A1025059508087.

Borger:2005:HLM

[BFGS05] Egon Börger, Nicu G. Fruja, Vincenzo Gervasi, and Robert F. Stärk. A high-level modular definition of the semantics of C#. Theoretical Computer Science, 336(2–3):235– 284, May 26, 2005. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.
only4gurus.net/miscellaneous/
cs03.pdf.

Bubak:2002:MSD

[BFM^{+}02a] M. Bubak, W. Funika, P. Metel, R. Orłowski, and R. Wismüller. Monitoring system for distributed Java applications. Lecture Notes in Computer Science, 2330:874– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2330/23300874.
htm;   http://link.springer-
ny.com/link/service/series/

0558/papers/2330/23300874.
pdf.

Bubak:2002:TMI

[BFM^{+}02b] Marian Bubak, Włodzimierz Funika, Piotr Mȩtel, Rafał Orłowski, and Roland Wismüller. Towards a monitoring interface specification for distributed Java applications. Lecture Notes in Computer Science, 2328:315–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2328/23280315.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2328/23280315.
pdf.

Bruns:2000:ASD

[BFMT00] Glenn R. Bruns, Alan E. Frey, Peter A. Mataga, and Susan J. Tripp. Automated software development with XML and the Java language. Bell Labs Technical Journal, 5(2): 32–43, Summer 2000. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Bartetzko:2004:JJA

[BFMW04] D. Bartetzko, C. Fischer, M. Moller, and H. Wehrheim. Jass — Java with assertions. Electronic Notes in Theoretical Computer Science, 55(2): 1–15, January 2004. CODEN ???? ISSN 1571-0661.

Baxter:2006:USJ

[BFN^{+}06] Gareth Baxter, Marcus Frean, James Noble, Mark Rickerby, Hayden Smith, Matt Visser, Hayden Melton, and Ewan Tempero. Understanding the shape of Java software. ACM SIGPLAN Notices, 41(12):397–412, December 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bloom:2009:TRC

[BFN^{+}09] Bard Bloom, John Field, Nathaniel Nystrom, Johan Östlund, Gregor Richards, Rok Strniša, Jan Vitek, and Tobias Wrigstad. Thorn: robust, concurrent, extensible scripting on the JVM. ACM SIGPLAN Notices, 44 (10):117–136, October 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bubak:2003:AMS

[BFS^{+}03] M. Bubak, W. Funika, M. Smetek, Z. Kilianski, and R. Wismuller. Architecture of monitoring system for distributed Java applications. Lecture Notes in Computer Science, 2840:447–454, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bubak:2004:RPJ

[BFS^{+}04] M. Bubak, W. Funika, M. Smetek, Z. Kilianski, and R. Wismuller. Request processing in the Java-oriented OMIS compliant monitoring system. Lecture Notes in Computer Science, 3019:352– 359, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bubak:2003:MDJ

[BFW^{+}03] Marian Bubak, Włodzimierz Funika, Roland Wismüller, Piotr Metel, and Rafał Orłowski. Monitoring of distributed Java applications. Future Generation Computer Systems, 19(5): 651–663, July 2003. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Butincu:2002:DDA

[BG02] Cristian Buţincu and Dan Grigoras. DisMedJava — A distributed application for medical image processing. Lecture Notes in Computer Science, 2326:308–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2326/23260308.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2326/23260308.
pdf.

Brebner:2003:JIS

[BG03] Paul Brebner and Jeffrey Gosper. J2EE infrastructure scalability and throughput estimation. ACM SIGMET-RICS Perform. Eval. Rev., 31(3):30–36, December 2003. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Bohme:2004:LFR

[BG04a] H. Bohme and U. Golze. Lightweight firm real-time extensions for low memory profile Java systems. Lecture Notes in Computer Science, 3292:303–314, 2004. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boshernitsan:2004:IIS

[BG04b] Marat Boshernitsan and Susan L. Graham. iXj: interactive source-to-source transformations for Java. ACM SIG-PLAN Notices, 39(10):212– 213, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bloch:2005:JPT

[BG05] Joshua Bloch and Neal Gafter. Java puzzlers: traps, pitfalls, and corner cases. Addison-Wesley, Reading, MA, USA, 2005. ISBN 0-321-33678-X (paperback). xix + 282 pp. LCCN QA76.73.J38 B58 2005. URL http:/

/www.loc.gov/catdir/toc/
ecip0513/2005015278.html.

Bonorden:2006:WCE

[BGadH06] Olaf Bonorden, Joachim Gehweiler, and Friedhelm Meyer auf der Heide. A Web computing environment for parallel algorithms in Java. Scalable Computing: Practice and Experience, 7(2): 1–14, June 2006. CO-DEN ???? ISSN 1895-

1767. URL http://www.
scpe.org/vols/vol07/no2/
SCPE_7_2_01.pdf;   http:
//www.scpe.org/vols/vol07/
no2/SCPE_7_2_01.zip.

Buytaert:2004:BAJ

[BGED04] Dries Buytaert, Andy Georges, Lieven Eeckhout, and Koen De Bosschere. Bottleneck analysis in Java applications using hardware performance monitors. ACM SIG-PLAN Notices, 39(10):172– 173, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Boudreau:2003:NDG

[BGG^{+}03] Tim Boudreau, Jesse Glick, Simeon Greene, Vaughn Spurlin, and Jack J. Woehr. NetBeans: The Definitive Guide. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2003. ISBN 0-596-00280-7. xxiv + 646 pp. LCCN QA76.73.J38 N47 2003. US$44.95. URL

http://www.oreilly.com/
catalog/9780596002800;
http://www.oreilly.com/
catalog/netbeans.

Blackburn:2006:DBJ

[BGH^{+}06] Stephen M. Blackburn, Robin Garner, Chris Hoffmann, Asjad M. Khang, Kathryn S. McKinley, Rotem Bentzur, Amer Diwan, Daniel Feinberg, Daniel Frampton, Samuel Z. Guyer, Martin Hirzel, Antony Hosking, Maria Jump, Han Lee, J. Eliot B. Moss, B. Moss, Aashish Phansalkar, Darko Stefanović, Thomas Van-Drunen, Daniel von Dincklage, and Ben Wiedermann. The DaCapo benchmarks: Java benchmarking development and analysis. ACM SIG-PLAN Notices, 41(10):169– 190, October 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Buytaert:2007:UHS

[BGH^{+}07] Dries Buytaert, Andy Georges, Michael Hind, Matthew Arnold, Lieven Eeckhout, and Koen De Bosschere. Using HPM-sampling to drive dynamic compilation. ACM SIGPLAN Notices, 42(10): 553–568, October 2007. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Blumenstein:2004:EAG

[BGNM04] Michael Blumenstein, Steven Green, Ann Nguyen, and Vallipuram Muthukkumarasamy. An experimental analysis of GAME: a generic automated marking environment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36 (3):67–71, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Boszormenyi:2000:SNW

[BGP00] László Böszörményi, Jurg Gutknecht, and Gustav Pomberger, editors. The School of Niklaus Wirth: The Art of Simplicity. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2000. ISBN 1-55860-723-4. viii + 260 pp. LCCN QA76.756 .S36 2000. US$39.95.

URL http://www.mkp.com/
books_catalog/catalog.asp?
ISBN=1-55860-723-4.

Busi:2000:PCC

[BGZ00] Nadia Busi, Roberto Gorrieri, and Gianluigi Zavattaro. Process calculi for coordination: From Linda to JavaSpaces. Lecture Notes in Computer Science, 1816:198– ??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/1816/18160198.

                                        htm;   http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/1816/18160198.
                                        pdf.

Bagga:2002:JJB

[BH02a] Jay Bagga and Adrian Heinz. JGraph — A Java based system for drawing graphs and running graph algorithms. Lecture Notes in Computer Science, 2265:459– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.
                                        com/link/service/series/
                                        0558/bibs/2265/22650459.
                                        htm;   http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2265/22650459.
                                        pdf.

Baker:2002:MMD

[BH02b] Jason Baker and Wilson C. Hsieh. Maya: multipledispatch syntax extension in Java. ACM SIGPLAN Notices, 37(5):270–281, May 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Brosgol:2002:SSU

[BH02c] Ben Brosgol and Michael González Harbour. Session summary: update on the real-time specification for Java. ACM SIGADA Ada Letters, 22 (4):128–130, December 2002. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Bottcher:2003:DWN

[BH03] S. Bottcher and R. Hoeppe.

Do we need components for persistent data storage in the Java 2 enterprise environment? Lecture Notes in Computer Science, 2591:152– 165, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Binder:2004:PCM

[BH04a] W. Binder and J. Hulaas.

A portable CPU-management framework for Java. IEEE Internet Computing, 8(5):74– 83, 2004. CODEN IICOFX. ISSN 1089-7801.

Binder:2004:SAP

[BH04b] W. Binder and J. Hulaas.

Self-accounting as principle for portable CPU control in Java. Lecture Notes in Computer Science, 3263:24– 38, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bishop:2004:DPG

[BH04c] Judith Bishop and Nigel Horspool. Developing principles of GUI programming using views. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):373–377, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Back:2005:KJR

[BH05a] Godmar Back and Wilson C. Hsieh. The KaffeOS Java runtime system. ACM Transactions on Programming Languages and Systems, 27(4): 583–630, July 2005. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Binder:2005:ESJ

[BH05b] W. Binder and J. Hulaas. Extending standard Java runtime systems for resource management. Lecture Notes in Computer Science, 3437:154–169, 2005. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buhr:2005:ISM

[BH05c] Peter A. Buhr and Ashif S. Harji. Implicit-signal monitors. ACM Transactions on Programming Languages and Systems, 27(6):1270– 1343, November 2005. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Bertels:2009:EMM

[BHDS09] Peter Bertels, Wim Heirman, Erik D’Hollander, and Dirk Stroobandt. Efficient memory management for hardware accelerated Java Virtual Machines. ACM Transactions on Design Automation of Electronic Systems, 14(4):

48:1–48:??, August 2009. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Beloglavec:2005:ALM

[BHJR05] Simon Beloglavec, Marjan Heričko, Matjaž B. Jurič, and Ivan Rozman. Analysis of the limitations of multiple client handling in a Java server environment. ACM SIGPLAN Notices, 40(4):20–28, April 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bouchenak:2004:EIE

[BHK^{+}04] S. Bouchenak, D. Hagimont, S. Krakowiak, N. De Palma, and F. Boyer. Experiences implementing efficient Java thread serialization, mobility and persistence. Software— Practice and Experience, 34 (4):355–393, April 10, 2004. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Back:2000:PKI

[BHL00] Godmar Back, Wilson H.

Hsieh, and Jay Lepreau. Processes in KaffeOS: Isolation, resource management, and sharing in Java. In USENIX [USE00c], page ?? ISBN 1-880446-16-2. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/osdi2000/
back.html.

Blackburn:2007:PBP

[BHM^{+}07] Stephen M. Blackburn, Matthew Hertz, Kathryn S. Mckinley, J. Eliot B. Moss, and Ting Yang. Profile-based pretenuring. ACM Transactions on Programming Languages and Systems, 29(1): 2:1–2:57, January 2007. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Bonzini:2001:LHG

[BHP^{+}01] Paolo Bonzini, Stuart Halloway, John Penry, Oluseyi Sonaiya, Bruce E. Hogman, Greg Bissell, Michael Hobbs, and Ben Laurie. Letters: Huge GCC executables; Java class loader; Department of Dumb Ideas; setting the record straight; the legacy of C#; DHTML source-code correction; shared libraries aren’t all bad; Zuse and Intel. Dr. Dobb’s Journal of Software Tools, 26(8):10, 12, August 2001. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/.

Brosgol:2002:ATC

[BHR02] Benjamin M. Brosgol, Ricardo J. Hassan, II, and Scott Robbins. Asynchronous transfer of control in the realtime specification for Java^{T M}. ACM SIGADA Ada Letters, 22(4):95–112, December 2002. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Beckert:2007:VOO

[BHS07] Bernhard Beckert, Reiner Hähnle, and Peter H. Schmitt, editors. Verification of Object-Oriented Software. The KeY Approach, volume 4334 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. CO-DEN LNCSD9. ISBN 3-540-68977-X (print), 3-540-69061-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL

http://www.springerlink.
com/content/978-3-540-69061-

0. Foreword by K. Rustan M. Leino.

Binder:2001:PRC

[BHV01] Walter Binder, Jane G. Hulaas, and Alex Villazón. Portable resource control in Java. ACM SIGPLAN Notices, 36(11):139–155, November 2001. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Proceedings of the 2001 ACM SIGPLAN Conference on Object Oriented Programming, Systems, Languages and Applications (OOPSLA’01).

Bishop:2005:EIJ

[BHW05] Judith Bishop, R. Nigel Horspool, and Basil Worrall. Experience in integrating Java with C# and .NET. Concurrency and Computation:

Practice and Experience, 17 (5–6):663–680, April/May 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Basha:2002:ANG

[BI02] S. Jeelani Basha and Romin Irani. AXIS: the next generation of Java SOAP. Wrox Press, Chicago, IL, USA, 2002. ISBN 1-86100-715-9. v + 275 pp. LCCN QA76.76.H94 B37 2002 Bar.

Bohnenkamp:2007:SGJ

[BI07] Wendy Bohnenkamp and Jackie Iverson. SAS Graphics for Java: examples using SAS AppDev studio and the Output delivery system. SAS Press series. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 2007. ISBN 1-59047-693-X. xii + 342 pp. LCCN HA31 .B645 2007.

Badjonski:2005:AJA

[BIB05] Mihal Badjonski, Mirjana Ivanovic, and Zoran Budimac. Adaptable Java Agents (AJA) — a tool for programming of multi-agent systems. ACM SIGPLAN Notices, 40(2):17–26, February 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Billard:2003:LDP

[Bil03] Edward A. Billard. Languagedependent performance of design patterns. ACM SIG-SOFT Software Engineering Notes, 28(3):3, May 2003. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Binder:2006:PAS

[Bin06] Walter Binder. Portable and accurate sampling profiling for Java. Software—Practice and Experience, 36(6): 615–650, May 2006. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Birnam:2001:DJP

[Bir01] Stewart Birnam. Distributed Java 2 Platform Database Development. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2001. ISBN 0-13-026861-5. xxi + 276 pp. LCCN QA76.73.J38 B57 2001. US$39.99. URL http:/

/www.phptr.com/ptrbooks/
ptr_0130268615.html;   http:
//www.sun.com/books/catalog/
birnam/.

Bishop:2003:ICJ

[Bis03] David Bishop. Introduction to cryptography with Java applets. Jones and Bartlett, Boston, MA, USA, 2003. ISBN 0-7637-2207-3. xvi + 370 pp. LCCN QA76.9.A25 B565 2003.

Brett:2004:WBK

[BJ04] G. Olivier Brett and L. Snoep Jacky. Web-based kinetic modelling using JWS Online. Bioinformatics, 20(13):2143– 2144, September 1, 2004. CO-DEN ???? ISSN 1367-4803 (print), 1367-4811 (electronic).

Budimlic:2007:ICJ

[BJK07] Zoran Budimlić, Mackale Joyner, and Ken Kennedy. Improving compilation of Java scientific applications. The International Journal of High Performance Computing Applications, 21(3):251– 265, August 2007. CO-DEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (elec-

tronic). URL http://hpc.
sagepub.com/content/21/
3/251.full.pdf+html.

Breunesse:2002:SVD

[BJvdB02] Cees-Bart Breunesse, Bart Jacobs, and Joachim van den Berg. Specifying and verifying a decimal representation in Java for smart cards. Lecture Notes in Computer Science, 2422:304–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2422/24220304.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2422/24220304.
pdf.

Bull:2000:JOL

[BK00] J. M. Bull and M. E. Kambites. JOMP: an OpenMPlike interface for Java. In ????, editor, Proceedings of the ACM 2000 conference on Java Grande, pages 44–53. ACM Press, New York, NY 10036, USA, 2000.

Budimlic:2001:JJC

[BK01a] Zoran Budimlić and Ken Kennedy. JaMake: a Java compiler environment. Lecture Notes in Computer Science, 2179:201–209, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2179/21790201.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2179/21790201.
pdf.

Buhler:2001:FSA

[BK01b] Dieter Bühler and Wolfgang Küchlin. Flexible similarity assessment for XML documents based on XQL and Java reflection. Lecture Notes in Computer Science, 2070:175–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2070/20700175.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2070/20700175.
pdf.

Boshart:2003:GGX

[BK03] Mark A. Boshart and Martha J. Kosa. Growing a GUI from an XML tree. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):223, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bauer:2005:HA

[BK05a] Christian Bauer and Gavin King. Hibernate in Action. Manning Publications, Greenwich, CT, USA, 2005. ISBN 1-932394-15-X. xxiii + 408 pp. LCCN QA76.64 .B393 2005. US$44.95. URL

                                        http://www.manning.com/
                                        bauer.

Budimlic:2005:CAW

[BK05b] Zoran Budimlić and Ken Kennedy. Compiling almostwhole Java programs. Concurrency and Computation: Practice and Experience, 17 (5–6):573–588, April/May 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bapst:2008:SIO

[BK08] Frederic Bapst and François Kilchoer. Signalling integer overflows in Java: a tool for checking overflows in Java code. Dr. Dobb’s Journal of Software Tools, 33(9):54– 58, September 2008. CODEN DDJOEB. ISSN 1044-789X.

URL http://home.hefr.ch/
bapst/cojac.

Baek:2002:IMM

[BKH02] J. M. Baek, B. M. Kang, and I. S. Hong. Implementation of mileage management system using Java Card. Simulation series, 34(2):3–8, 2002. CO-DEN SMCPAX. ISSN 0735-9276.

Bubak:2001:CUL

[BKŁ01] Marian Bubak, Dawid Kurzyniec, and Piotr Łuszczek. Convenient use of legacy software in Java with Janet package. Future Generation Computer Systems, 17 (8):987–997, June 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/35/35/abstract.
html.

Bubak:2000:CJN

[BKLS00] M. Bubak, D. Kurzyniec, P. Luszczek, and V. Sunderam. Creating Java to native code interfaces with Janet. Scientific Programming, 9(1):39–50, 2000. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Bubak:2001:CJN

[BKLS01] Marian Bubak, Dawid Kurzyniec, Piotr Luszczek, and V. Sunderam. Creating Java to Native Code Interfaces with Janet. Scientific Programming, 9(1):39–50, ???? 2001. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:/

/iospress.metapress.com/
app/home/contribution.asp%
3Fwasp=f2779jvvqg63jq64qwtm%
26referrer=parent%26backto=
issue%2C4%2C6%3Bjournal%
2C2%2C9%3Blinkingpublicationresults%
2C1%2C1.

Boyapati:2002:KAT

[BKM02]  Chandrasekhar  Boyapati,

Sarfraz Khurshid, and Darko Marinov. Korat: automated testing based on Java predicates. ACM SIGSOFT Software Engineering Notes, 27 (4):123–133, July 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bacon:2004:TLF

[BKMS04] David F. Bacon, Ravi Konuru, Chet Murthy, and Mauricio J. Serrano. Thin locks: Featherweight synchronization for Java. ACM SIGPLAN Notices, 39(4):583–595, April 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bull:2000:PPJ

[BKO00] J. Mark Bull, Mark E. Kambites, and Jan Obdrzalek. Parallel programming in Java with OpenMP-like directives. In ACM [ACM00c], page 150.

URL http://www.sc2000.
org/proceedings/info/fp.
pdf.

Bronson:2009:FDB

[BKO09] Nathan G. Bronson, Christos Kozyrakis, and Kunle Olukotun. Feedback-directed barrier optimization in a strongly isolated STM. ACM SIG-PLAN Notices, 44(1):213– 225, January 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

BalaKumar:2003:BAP

[BKT03] C. Bala Kumar, Paul Kline, and Tim Thompson. Bluetooth Application Programming with the Java APIs. The Morgan Kaufmann series in networking. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2003. ISBN 1-55860-934-2. xxii + 498 pp. LCCN TK5103.3 .K86 2004. US$44.95. URL ftp:/

/uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
description/els051/2003107476.
html;   http://www.loc.
gov/catdir/toc/els051/2003107476.
html.

BuSung:2003:DIJ

[BKY^{+}03] L. BuSung, Y. C. Kiat, S. I.

Yann, L. K. Kee, and S. Wei. Design and implementation of a Java-based meeting space over Internet. Multimedia Tools and Applications, 20 (2):179–195, 2003. CODEN MTAPFB. ISSN 1380-7501.

Binder:2002:USJ

[BL02a] Walter Binder and Balázs Lichtl. Using a secure Java micro-kernel on embedded devices for the reliable execution of dynamically uploaded applications. Lecture Notes in Computer Science, 2361:125– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.
                                        com/link/service/series/
                                        0558/bibs/2361/23610125.
                                        htm;   http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2361/23610125.
                                        pdf.

Burchfield:2002:UAA

[BL02b] J. Burchfield and S. Lipovaca. Using an APL approach with Java. APL Quote Quad, 32(4):61–68, 2002. CODEN APLQD9. ISSN 0163-6006.

Bouquet:2003:RET

[BL03] F. Bouquet and B. Legeard. Reification of executable test scripts in formal specificationbased test generation: The Java Card transaction mechanism case study. Lecture Notes in Computer Science, 2805:778–795, 2003. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

BohneLang:2004:MII

[BL04] A. BohneLang and E. Lang.

3D-Molekulvisualisierung im Internet Schwerpunkt Java-Applets. (German) [3Dmolecular visualization in Internet center-of-gravity applets]. Biospektrum, 10(2): 167–169, 2004. CODEN ???? ISSN 0947-0867.

Blanchet:2003:EAJ

[Bla03] Bruno Blanchet. Escape analysis for Java^{T M}: Theory and practice. ACM Transactions on Programming Languages and Systems, 25(6): 713–775, November 2003. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Briand:2006:TRE

[BLL06] L. C. Briand, Y. Labiche, and J. Leduc. Toward the reverse engineering of UML sequence diagrams for distributed Java software. IEEE Transactions on Software Engineering, 32(9):642–663, September 2006. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
1707665.

Baldi:2008:TAL

[BLLB08] Pierre F. Baldi, Cristina V.

Lopes, Erik J. Linstead, and Sushil K. Bajracharya. A theory of aspects as latent topics. ACM SIGPLAN Notices, 43(10):543–562, September 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bruce-Lockhart:2006:IEE

[BLN06] Michael Bruce-Lockhart and Theodore S. Norvell. Interactive embedded examples: a demonstration. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3): 357, September 2006. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bloch:2001:EJP

[Blo01]  Joshua Bloch.  Effective

Java: Programming Language Guide. Addison-Wes-ley, Reading, MA, USA, 2001. ISBN 0-201-31005-8, 3-642-56735-5 (e-book). xvi + 252 pp. LCCN QA76.73.J38 B57 2001.

Bloch:2008:EJ

[Blo08]  Joshua Bloch.  Effective

Java. The Java series. Addison-Wesley, Reading, MA, USA, second edition, 2008. ISBN 0-321-35668-3 (paperback). xxi + 346 pp. LCCN QA76.73.J38 B57 2008. URL http:/

                                        /www.loc.gov/catdir/toc/
                                        fy0805/2008926278.html.

Bucker:2004:TUC

[BLPV04] H. M. Bucker, B. Lang, H. J. Pflug, and A. Vehreschild.

Threads in an undergraduate course: a Java example illuminating different multithreading approaches. Lecture Notes in Computer Science, 3044:882–891, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bettini:2003:MIJ

[BLV03] L. Bettini, M. Loreti, and B. Venneri. On multiple inheritance in Java. Kluwer International Series in Engineering and Computer Science, 732:1–15, 2003. CO-DEN ???? ISSN 0893-3405.

Breg:2000:PEJ

[BLW00] Fabian Breg, Mike Lew, and Harry A. G. Wijshoff. Performance evaluation of a Java based chat system. Lecture Notes in Computer Science, 1940:276–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/1940/19400276.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/1940/19400276.
pdf.

Bauer:2009:CER

[BLW09] Lujo Bauer, Jay Ligatti, and David Walker. Composing expressive runtime security policies. ACM Transactions on Software Engineering and Methodology, 18(3):

9:1–9:??, May 2009. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Berzal:2001:TTJ

[BM01] Fernando Berzal and Nicolás Marín. Tools and tips for Java programming. IEEE Distributed Systems Online, 2(5), 2001. ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:

                                        //dsonline.computer.org/
                                        0105/departments/bks0105a.
                                        htm.

Beckert:2003:PLH

[BM03] B. Beckert and W. Mostowski. A program logic for handling JAVA CARD’s transaction mechanism. Lecture Notes in Computer Science, 2621:246– 260, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boulifa:2004:MGD

[BM04] R. Boulifa and E. Madelaine. Model generation for distributed Java programs. Lecture Notes in Computer Science, 2952:139–152, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bond:2007:PCC

[BM07] Michael D. Bond and Kathryn S. McKinley. Probabilistic calling context. ACM SIGPLAN Notices, 42(10):97–112, October 2007. CODEN SINODQ.

ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bond:2008:TML

[BM08] Michael D. Bond and Kathryn S.

McKinley. Tolerating memory leaks. ACM SIGPLAN Notices, 43(10):109–126, September 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bond:2009:LP

[BM09] Michael D. Bond and Kathryn S.

McKinley. Leak pruning. ACM SIGPLAN Notices, 44(3):277–288, March 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Burke:2006:EJ

[BMH06] Bill Burke and Richard Monson-Haefel. Enterprise JavaBeans 3.0. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, fifth edition, 2006. ISBN 0-596-00978-X. xxv + 732 pp. LCCN QA76.73.J38 M65

2006. URL http://www.loc.
gov/catdir/enhancements/
fy0715/2006285201-d.html.

Bolignano:2001:FMC

[BML01] D. Bolignano, D. Le Métayer, and C. Loiseaux. Formal methods in context: Security and Java card. Lecture Notes in Computer Science, 2041:1– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2041/20410001.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2041/20410001.
pdf.

Baiardi:2002:JSD

[BMR02] F. Baiardi, P. Mori, and L. Ricci. A JAVA support for distributed shared memory on COW. Applied Informatics, 3: 191–195, 2002. CODEN ???? ISSN 1027-2666.

Brady:2002:JPB

[BMS02] Alyce Brady, R. C. McDowell, and Kelly Schultz. JavaScript programming basics: a laboratory series for beginning programmers. ACM Journal on Educational Resources in Computing (JERIC), 2(2): 1, June 2002. CODEN ???? ISSN 1531-4278.

Benowitz:2003:EAR

[BN03] E. G. Benowitz and A. F. Niessner. Experiences in adopting real-time Java for flight-like software. Lecture Notes in Computer Science, 2889:490–496, 2003. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bond:2007:TBA

[BNK^{+}07] Michael D. Bond, Nicholas Nethercote, Stephen W. Kent, Samuel Z. Guyer, and Kathryn S. McKinley. Tracking bad apples: reporting the origin of null and undefined value errors. ACM SIG-PLAN Notices, 42(10):405– 422, October 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Beraldi:2003:TUT

[BNO03] R. Beraldi, L. Nigro, and A. Orlando. Temporal uncertainty time warp: An implementation based on Java and ActorFoundry. Simulation, 79(10):581–597, 2003. CO-DEN SIMUA2. ISSN 0037-5497 (print), 1741-3133 (electronic).

Badea:2008:IJS

[BNV08] Carmen Badea, Alexandru Nicolau, and Alexander V. Veidenbaum. Impact of JVM superoperators on energy consumption in resourceconstrained embedded systems. ACM SIGPLAN Notices, 43(7):23–30, July 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bellia:2005:HOP

[BO05] Marco Bellia and M. Eugenia Occhiuto. Higher order programming in Java:

Introspection, subsumption and extraction. Fundamenta Informaticae, 67(1–3): 29–44, January 2005. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bellia:2008:MPP

[BO08] Marco Bellia and M. Eugenia Occhiuto. Methods as parameters: A preprocessing approach to higher order in Java. Fundamenta Informaticae, 85(1–4):35–50, September 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bellia:2009:JSI

[BO09] Marco Bellia and M. Eugenia Occhiuto. JavaΩ: The structures and the implementation of a preprocessor for Java with m and mc parameters. Fundamenta Informaticae, 93(1–3):45–64, January 2009. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bodden:2004:LLR

[Bod04] Eric Bodden. A lightweight LTL runtime verification tool for Java. ACM SIGPLAN Notices, 39(10):306–307, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Boehm:2005:CRJ

[Boe05] H. J. Boehm. The constructive reals as a Java library. Journal of Logic and Algebraic Programming, 64(1):3– 11, 2005. CODEN ???? ISSN 1567-8326.

Bogda:2000:DRO

[Bog00] Jeff Bogda. Detecting readonly methods in Java. Lecture Notes in Computer Science, 1915:143–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/1915/19150143.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/1915/19150143.
pdf.

Boger:2001:JDS

[Bog01] Marko Boger. Java in distributed systems: concurrency, distribution, and persistence. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 2001. ISBN 0-471-49838-6. xix + 393 pp. LCCN QA76.73.J38 B64 2001.

Bollella:2000:RTS

[Bol00] Gregory Bollella. The realtime specification for Java. Java series. Addison-Wesley, Reading, MA, USA, 2000. ISBN 0-201-70323-8 (paperback). xxiii + 195 pp. LCCN QA76.73.J38 R42 2000.

Boone:2000:UJX

[Boo00] Keith W. Boone. Using Java for XML processing. Markup languages: theory & practice, 2(3):321–330, Summer 2000. CODEN MLTPFG. ISSN 1099-6621. URL

                                        http://mitpress.mit.edu/
                                        catalog/item/default.asp?
                                        sid=3108FA5E-91BB-480F-
                                        9930-C1C27725EAB8&ttype=
                                        6&tid=7574.

Bossert:2004:JSC

[Bos04] O. Bossert. A Java software for creation of image mosaics. Computer Methods and Programs in Biomedicine, 75 (2):159–161, 2004. CO-DEN CMPBEK. ISSN 0169-2607 (print), 1872-7565 (electronic).

Bouchi:2002:JTM

[BOT02] A. Bouchi, R. Olejnik, and Bernard Toursel. Java tools for measurement of the machine loads. Lecture Notes in Computer Science, 2326:271– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.
                                        com/link/service/series/
                                        0558/bibs/2326/23260271.
                                        htm;   http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2326/23260271.
                                        pdf.

Bothner:2003:CJG

[Bot03] Per Bothner. Compiling Java with GCJ. Linux Journal, 105:??, January 2003. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http:

//www.linuxjournal.com/
article/4860.

Bouchenak:2001:MJA

[Bou01]  Sara Bouchenak.  Mak-

ing Java applications mobile or persistent. In USENIX [USE01a], page ?? ISBN 1-880446-12-X. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/coots01/
bouchenak.html.

Bower:2007:GAS

[Bow07] Matt Bower. Groupwork activities in synchronous online classroom spaces. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1): 91–95, March 2007. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bachrach:2001:JSE

[BP01a] Jonthan Bachrach and Keith Playford. The Java Syntactic Extender (JSE). ACM SIG-PLAN Notices, 36(11):31–42, November 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Proceedings of the 2001 ACM SIGPLAN Conference on Object Oriented Programming, Systems, Languages and Applications (OOPSLA’01).

Batheja:2001:FOC

[BP01b] Jyoti Batheja and Manish Parashar. A framework for opportunistic cluster computing using JavaSpaces. Lecture Notes in Computer Science, 2110:647–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.
                                        com/link/service/series/
                                        0558/bibs/2110/21100647.
                                        htm;   http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2110/21100647.
                                        pdf.

Bechini:2001:BIC

[BP01c] Alessio Bechini and Cosimo Antonio Prete. Behavior investigation of concurrent Java programs: an approach based on source-code instrumentation. Future Generation Computer Systems, 18(2):307– 316, October 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/60/31/36/abstract.
                                        html.

Breg:2001:JVM

[BP01d] Fabian Breg and Constantine Polychronopoulos. Java Virtual Machine support for object serialization. In ACM [ACM01b], pages 173– 180. ISBN 1-58113-359-6. LCCN QA76.9.O35 A26

                                        2001. URL http://www.

philippsen.com/JGI2001/
camerareadyabstracts/12.
html;   http://www.philippsen.
com/JGI2001/finalpapers/
18500173.ps.

Bell:2002:JS

[BP02] Doug Bell and Mike Parr.

Java for students. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, third edition, 2002. ISBN 0-13-032377-2 (paperback). xxvi + 634 pp. LCCN QA76.73.J38 B45 2002.

Bierman:2003:EEI

[BP03a] G. M. Bierman and M. J.

Parkinson. Effects and effect inference for a core Java calculus. Electronic Notes in Theoretical Computer Science, 82(8):1–26, 2003. CO-DEN ???? ISSN 1571-0661.

Breg:2003:JVM

[BP03b] Fabian Breg and Constantine D. Polychronopoulos. Java Virtual Machine support for object serialization. Concurrency and Computation: Practice and Experience, 15 (3–5):263–275, March/April 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Brinkschulte:2005:ICA

[BP05] U. Brinkschulte and M. Pacher.

Implementing control algorithms within a multithreaded Java microcontroller. Lecture Notes in Computer Science, 3432:33– 49, 2005. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boroday:2005:DAJ

[BPSH05] S. Boroday, A. Petrenko, J. Singh, and H. Hallal. Dynamic analysis of Java applications for multithreaded antipatterns. ACM SIGSOFT Software Engineering Notes, 30(4):1–7, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Beebee:2001:ISM

[BR01a] William S. Beebee, Jr. and Martin Rinard. An implementation of scoped memory for real-time Java. Lecture Notes in Computer Science, 2211:289–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2211/22110289.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2211/22110289.
pdf.

Boyapati:2001:PTS

[BR01b] Chandrasekhar Boyapati and Martin Rinard. A parameterized type system for race-free Java programs. ACM SIG-PLAN Notices, 36(11):56–69, November 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Proceedings of the 2001 ACM SIGPLAN Conference on Object Oriented Programming, Systems, Languages and Applications (OOPSLA’01).

Brebner:2001:EBB

[BR01c] Paul Brebner and Shuping Ran. Entity bean A, B, C’s: Enterprise Java beans commit options and caching. Lecture Notes in Computer Science, 2218:36– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2218/22180036.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2218/22180036.
pdf.

Bruneton:2001:EJP

[BR01d] Eric Bruneton and Michel Riveill. Experiments with JavaPod, a platform designed for the adaptation of non-functional properties. Lecture Notes in Computer Science, 2192:52–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2192/21920052.
htm;   http://link.springer-
ny.com/link/service/series/

                                        0558/papers/2192/21920052.
                                        pdf.

Biermann:2002:GIC

[BR02] Alan W. Biermann and Dietolf Ramm. Great ideas in computer science with Java. MIT Press, Cambridge, MA, USA, 2002. ISBN 0-262-02497-7 (paperback). xx + 528 pp. LCCN QA76.73.J38 B52 2002.

Binder:2006:SRJ

[BR06a] Walter Binder and Volker Roth. Security risks in Javabased mobile code systems. Scalable Computing: Practice and Experience, 7(4):1– 11, December 2006. CO-DEN ???? ISSN 1895-

                                        1767. URL http://www.
                                        scpe.org/vols/vol07/no4/
                                        SCPE_7_4_01.pdf;   http:
                                        //www.scpe.org/vols/vol07/
                                        no4/SCPE_7_4_01.zip.

Bringert:2006:PAC

[BR06b] Björn Bringert and Aarne Ranta. A pattern for almost compositional functions. ACM SIGPLAN Notices, 41(9):216–226, September 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Butkevich:2000:CTS

[BRBY00] Sergey Butkevich, Marco Renedo, Gerald Baumgartner, and Michal Young. Compiler and tool support for debugging object protocols. ACM SIGSOFT Software Engineering Notes, 25(6):50– 59, November 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Budi:2003:JJT

[BRC03] E. M. Budi, G. Roy, and G. Cole. Jawa: a Java toolkit for mobile objects applications. Lecture Notes in Computer Science, 2604:39– 48, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bretz:2002:NPP

[Bre02] E. A. Bretz. New phones play games and run Java applications. IEEE Spectrum, 39(2): 62–63, February 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Brinkmann:2002:GGG

[Bri02] Peter Brinkmann. Gumbie:

a GUI generator for Jython. Dr. Dobb’s Journal of Software Tools, 27(4):46–50, April 2002. CODEN DDJOEB. ISSN 1044-789X. URL

http://www.ddj.com/ftp/
2002/2002_04/gumbie.txt;
http://www.ddj.com/ftp/
2002/2002_04/gumbie.zip.

Briggs:2005:TMJ

[Bri05] L. L. Briggs. There’s more to Java vs. .NET than technology. Application Development Trends, 12(5):43–47, 2005. CODEN ???? ISSN 1073-9564.

Burdy:2003:JAC

[BRL03] L. Burdy, A. Requet, and J. L. Lanet. Java applet correctness: a developer-oriented approach. Lecture Notes in Computer Science, 2805:422– 439, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brookshier:2000:JSC

[Bro00]  Daniel Brookshier.  Java

Standards for Corporate Development. R&D Books, Lawrence, KS, USA; Berkeley, CA, USA, 2000. ISBN 1-929629-00-1. ???? pp. LCCN ???? US$40.

Brogden:2001:JDG

[Bro01] William B. Brogden. Java developer’s guide to Servlets and JSP. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 2001. ISBN 0-7821-2809-2. xix + 411 pp. LCCN QA76.73.J38 B7474 2001.

Brooks:2002:BRB

[Bro02a] Terrence A. Brooks. Book review: Designing with Javascript: Creating dynamic Web pages. Journal of the American Society for Information Science and Technology: JASIST, 53(11):967– 968, ???? 2002. CO-DEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic).

Brown:2002:WAW

[Bro02b] Kyle Brown. WebSphere 4.0 AEs Workbook for Enterprise JavaBeans. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, third edition, 2002. ISBN 0-596-00418-4. 186 (est.) pp. LCCN ???? US$24.95. URL http://

www.oreilly.com/catalog/
websphereworkbook.

Brosgol:2003:AJR

[Bro03a] B. M. Brosgol. Ada and Java: Real-time advantages. Embedded Systems Programming, 16(12):30, 2003. CO-DEN EYPRE4. ISSN 1040-3272.

Brosgol:2003:BCR

[Bro03b] B. M. Brosgol. Beginner’s corner: Real-time Java. Embedded Systems Programming, 16(5):39–40, 2003. CODEN EYPRE4. ISSN 1040-3272.

Brosgol:2004:RTJ

[Bro04] Benjamin M. Brosgol. Realtime Java for Ada programmers. Ada User Journal, 25 (4):220–??, December 2004. CODEN AUJOET. ISSN 1381-6551.

Brosgol:2005:CME

[Bro05] B. M. Brosgol. A comparison of the mutual exclusion features in Ada and the real-time specification for Java. Lecture Notes in Computer Science, 3555:129–143, 2005. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brosgol:2007:SLS

[Bro07] Ben Brosgol. SA2: languages for safety-critical software: issues and assessment. ACM SIGADA Ada Letters, 27(3):2, December 2007. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Brosgol:2009:ICL

[Bro09] Ben Brosgol. An introduction to the C# language and .NET infrastructure. ACM SIGADA Ada Letters, 29(3): 3–4, December 2009. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Bruno:2002:JQ

[Bru02] Eric J. Bruno. Java Q&A: So what is a Java event agent? Dr. Dobb’s Journal of Software Tools, 27(7):86–91, July 2002. CODEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/ftp/2002/
2002_07/jqa0702.txt;   http:
//www.ddj.com/ftp/2002/
2002_07/jqa0702.zip.

Brunner:2003:JPG

[Bru03] Robert Brunner. JSP: Practical Guide for Java Programmers. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2003. ISBN 1-55860-836-2. xiii + 179 pp. LCCN QA76.73.J38 B79 2003. US$19.95. URL ftp:/

/uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
description/els041/2003066501.
html;   http://www.loc.
gov/catdir/toc/els041/2003066501.
html.

Brodie:2004:JJI

[BRU04a] R. Brodie, R. L. Roper, and C. Upton. JDotter: a Java interface to multiple dotplots generated by dotter. Bioinformatics, 20(2): 279–281, 2004. CODEN ???? ISSN 1367-4803 (print), 1367-4811 (electronic).

Bruce:2004:CHT

[Bru04b]  Kim B. Bruce.  Contro-

versy on how to teach CS 1: a discussion on the SIGCSE-members mailing list. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(4):29–34, December 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp:

//ftp.math.utah.edu/pub/
mirrors/ftp.ira.uka.de/
bibliography/Misc/DBLP/
2004.bib.

Bruno:2004:CJX

[Bru04c] E. J. Bruno. C++, Java, & XML processing. C/C++ Users Journal, 22(7):6–15, 2004. CODEN CCUJEX. ISSN 1075-2838.

Bruce:2005:CHT

[Bru05a]  Kim B. Bruce.  Contro-

versy on how to teach CS 1: a discussion on the SIGCSE-members mailing list. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(2):111–117, June 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp:

                                        //ftp.math.utah.edu/pub/
                                        mirrors/ftp.ira.uka.de/
                                        bibliography/Misc/DBLP/
                                        2005.bib.

Bruckschlegel:2005:MCC

[Bru05b]  T. Bruckschlegel.  Mi-

crobenchmarking C++, C#, and Java. C/C++ Users Journal, 23(7):14–21, 2005. CODEN CCUJEX. ISSN 1075-2838.

Bruno:2005:JWS

[Bru05c] E. J. Bruno. Java Web services & application architectures. Dr. Dobb’s Journal of Software Tools, 30(2):16– 23, 2005. CODEN DDJOEB. ISSN 1044-789X.

Bruno:2006:JM

[Bru06] Eric J. Bruno. Java messaging. Charles River Media programming series. Charles River Media, Inc., 403 VFW Drive, PO Box 417, Rockland, MA 02370, USA, 2006. ISBN 1-58450-418-8. xiii + 466 pp. LCCN TK5105.74.J32 B78

2006. URL http://www.loc.
gov/catdir/toc/ecip0517/
2005022084.html.

Boone:2000:JCE

[BS00a] Barry Boone and Willam R.

Stanek. Java 2 certification exam guide for programmers and developers. McGraw-Hill, New York, NY, USA, 2000. ISBN 0-07-913740-7. xii + 687 pp. LCCN QA76.3 .B658 2000. Includes CD-ROM.

Borger:2000:PMS

[BS00b] E. Borger and W. Schulte. A practical method for specification and analysis of exception handling — a Java/JVM case study. IEEE Transactions on Software Engineering, 26(9):872–887, September 2000. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
877847.

Borger:2000:IPJ

[BS00c] Egon Börger and Wolfram Schulte. Initialization problems for Java. Software — Concepts and Tools, 19(4): 175–178, October 2000. CO-DEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s003789900003.

Boussinot:2000:JTS

[BS00d] Frédéric Boussinot and Jean-Ferdy Susini. Java threads and SugarCubes. Software— Practice and Experience, 30 (5):545–566, April 25, 2000. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract/71004433/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=71004433&PLACEBO=IE.
                                        pdf.

Buck:2001:JCS

[BS01] Duane Buck and David J. Stucki. JKarelRobot: a case study in supporting levels of cognitive development in the computer science curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33 (1):16–20, March 2001. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Borger:2004:EAS

[BS04] E. Borger and R. F. Stark. Exploiting abstraction for specification reuse. the Java/ C# case study. Lecture Notes in Computer Science, 3188:42–76, 2004. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Basu:2007:MCJ

[BS07] Samik Basu and Scott A.

Smolka. Model checking the Java metalocking algorithm. ACM Transactions on Software Engineering and Methodology, 16(3): 12:1–12:??, July 2007. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Bravenboer:2009:SDS

[BS09] Martin Bravenboer and Yannis Smaragdakis. Strictly declarative specification of sophisticated points-to analyses. ACM SIGPLAN Notices, 44(10):243–262, October 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bull:2003:BJA

[BSB^{+}03] J. M. Bull, L. A. Smith, C. Ball, L. Pottage, and R. Freeman. Benchmarking Java against C and Fortran for scientific applications. Concurrency and Computation: Practice and Experience, 15(3–5):417–430, March/April 2003. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Basham:2004:HFS

[BSB04] Bryan Basham, Kathy Sierra, and Bert Bates. Head first servlets and JSP. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2004. ISBN 0-596-00540-7, 0-596-00540-7. xxxii + 854 pp. LCCN QA76.3 .B384 2004.

Basham:2008:HFS

[BSB08] Bryan Basham, Kathy Sierra, and Bert Bates. Head first servlets and JSP. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, second edition, 2008. ISBN 0-596-51668-1. xxxii + 879 pp. LCCN QA76.3 .B377 2008eb; QA76.3; QA76.3 .B377 2008; Internet.

Boyapati:2003:OTS

[BSBR03]  Chandrasekhar  Boyapati,

Alexandru Salcianu, William Beebee, Jr., and Martin Rinard. Ownership types for safe region-based memory management in real-time Java. ACM SIGPLAN Notices, 38(5):324–337, May 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Blackburn:2001:PJ

[BSH^{+}01] Stephen M. Blackburn, Sharad Singhai, Matthew Hertz, Kathryn S. McKinely, and J. Eliot B. Moss. Pretenuring for Java. ACM SIG-PLAN Notices, 36(11):342– 352, November 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Proceedings of the 2001 ACM SIGPLAN Conference on Object Oriented Programming, Systems, Languages and Applications (OOPSLA’01).

Binder:2009:CPJ

[BSMV09] Walter Binder, Martin Schoeberl, Philippe Moret, and Alex Villazón. Cross-profiling for Java processors. Software —Practice and Experience, 39(18):1439–1465, December 25, 2009. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bull:2001:BJA

[BSPF01] Mark Bull, Lorna Smith, Lindsay Pottage, and Robin Freeman. Benchmarking Java against C and Fortran for scientific applications. In ACM [ACM01b], pages 97–

105. ISBN 1-58113-359-

6. LCCN QA76.9.O35 A26

2001. URL http://www.
philippsen.com/JGI2001/
camerareadyabstracts/30.
html;   http://www.philippsen.
com/JGI2001/finalpapers/
18500097.pdf.

Bacon:2000:GDJ

[BST00] David F. Bacon, Robert E.

Strom, and Ashis Tarafdar. Guava: a dialect of Java without data races. ACM SIGPLAN Notices, 35 (10):382–400, October 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        proceedings/oops/353171/
                                        p382-bacon/.

Bull:2000:BSH

[BSW^{+}00] J. M. Bull, L. A. Smith, M. D. Westhead, D. S. Henty, and R. A. Davey. A benchmark suite for high performance Java. Concurrency: Practice and Experience, 12(6): 375–388, May 2000. CO-DEN CPEXEI. ISSN 1040-

                                        3108. URL http://www3.
                                        interscience.wiley.com/
                                        cgi-bin/abstract/72515724/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=72515724&PLACEBO=IE.
                                        pdf.

Back:2000:TDJ

[BTS^{+}00] Godmar Back, Patrick Tullmann, Leigh Stoller, Wilson C. Hsieh, and Jay Lepreau. Techniques for the design of Java operating systems. In USENIX [USE00a], pages 197–210. ISBN 1-880446-22-7. LCCN ???? URL http:/

                                        /www.usenix.org/events/
                                        usenix2000/general/back.
                                        html.

Bravenboer:2006:DFEa

[BTV06] Martin Bravenboer, Éric Tanter, and Eelco Visser. Declarative, formal, and extensible syntax definition for aspectJ. ACM SIGPLAN Notices, 41(10):209–228, October 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Budd:2000:UOO

[Bud00] Timothy Budd. Understanding object-oriented programming with Java. Addison-Wesley, Reading, MA, USA, 2000. ISBN 0-201-61273-9. xxi + 420 pp. LCCN QA76.64 .B835 2000.

Budd:2001:CDS

[Bud01] Timothy Budd. Classic data structures in Java. Addison-Wesley, Reading, MA, USA, 2001. ISBN 0-201-70002-6. xxiv + 593 pp. LCCN QA76.73.J38 B83 2001.

Bulka:2000:JPS

[Bul00] Dov Bulka. Java Performance and Scalability, Volume 1: Server-Side Programming Techniques. Addison-Wesley, Reading, MA, USA, 2000. ISBN 0-201-70429-3. 272 (est.) pp. LCCN QA76.73.J38 B84 2000. US$34.95.

Burke:2001:JX

[Bur01a] Eric M. Burke. Java and XSLT. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2001. ISBN 0-596-00143-6. xvi + 510 pp. LCCN QA76.73.J38 B87 2001.

Burke:2001:JXE

[Bur01b] Eric M. Burke. Java and XSLT: Embedding XML processing into Java applications. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, September 2001. ISBN 0-596-00143-6. xvi + 510 pp. LCCN QA76.73.J38 B88 2001. US$39.95. URL

http://www.oreilly.com/
catalog/9780596001438;
http://www.oreilly.com/
catalog/javaxslt.

Burkhalter:2002:JTE

[Bur02] B. Burkhalter. The JAI Top 10 Engineers answer questions about Java Advanced Imaging. Advanced Imaging, 17(10):32–39, 2002. CODEN ADIMEZ. ISSN 1042-0711.

Burger:2003:TTD

[Bur03] K. R. Burger. Teaching twodimensional array concepts in Java with image processing examples. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):205–209, 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Burnette:2005:EIP

[Bur05] Ed Burnette. Eclipse IDE: pocket guide. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2005. ISBN 0-596-10065-5. ix + 117 pp. LCCN QA76.73.J38 B875 2005. URL http://

www.oreilly.com/catalog/
9780596100650.

Burns:2007:DJG

[Bur07]  Brendan Burns.  Dark-

star: The Java Game Server. O’Reilly shortcuts. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2007. ISBN 0-596-51484-0. LCCN QA76.73.J38 B87 2007eb; QA76.73.J38. URL http://

www.oreilly.com/catalog/
9780596514846.

Busko:2002:SJTa

[Bus02a] I. Busko. Specview: a Java tool for spectral visualization and model fitting. Astronomical Society of the Pacific Conference Series, 281 (??):120–124, 2002. CODEN ???? ISSN 1050-3390.

Busko:2002:SJTb

[Bus02b]  I. Busko.  Specview: a

Java tool for spectral visualization and model fitting of multi-instrument data [4847-51]. Proceedings of the SPIE — The International Society for Optical Engineering, 4847:410–418, 2002. CODEN PSISDG. ISSN 0277-786X (print), 1996-756X (electronic).

Boldi:2005:MSJ

[BV05] Paolo Boldi and Sebastiano Vigna. Mutable strings in Java: design, implementation and lightweight textsearch algorithms. Science of Computer Programming, 54 (1):3–23, January 2005. CO-DEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Brose:2001:JPC

[BVD01] Gerald Brose, Andreas Vogel, and Keith Duddy. Java programming with CORBA: advanced techniques for building distributed applications. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, third edition, 2001. ISBN 0-471-37681-7. xlii + 710 pp. LCCN QA76.73.J38 B77 2001.

Bierre:2006:MOB

[BVPE06] Kevin Bierre, Phil Ventura, Andrew Phelps, and Christopher Egert. Motivating OOP by blowing things up: an exercise in cooperation and competition in an introductory Java programming course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):354–358, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bradley:2001:IJT

[BW01a] James W. Bradley and R. Webster West. Interactive Java tools for exploring high-dimensional data.

Journal of Statistical Software, 6(1):??, 2001. CO-DEN JSSOBK. ISSN

????  URL http://www.
jstatsoft.org/v06/i01;
http://www.jstatsoft.org/
v06/i01/bradley.tar;   http:
//www.jstatsoft.org/v06/
i01/bradley/index.html;
 http://www.jstatsoft.org/
v06/i01/updates.

Burns:2001:RTS

[BW01b] Alan Burns and Andrew J.

Wellings. Real-time systems and programming languages: Ada 95, real-time Java, and real-time POSIX. International computer science series. Addison-Wesley, Reading, MA, USA, third edition, 2001. ISBN 0-201-72988-1. xvi + 738 pp. LCCN QA76.54 .B87 2001. Revised edition of Real-time systems and their programming languages, 1990.

Brosgol:2003:CATa

[BW03a] B. M. Brosgol and A. Wellings.

A comparison of the asynchronous transfer of control features in Ada and the realtime specification for Java. YCS Report 350, University of York Department of Computer Science, York, UK, 2003.

Brosgol:2003:CATb

[BW03b] B. M. Brosgol and A. Wellings.

A comparison of the asynchronous transfer of control features in Ada and the realtime specification for Java.

Lecture Notes in Computer Science, 2655:113–128, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Burns:2003:PGP

[BW03c] A. Burns and A. J. Wellings. Processing group parameters in the real-time specification for Java. Lecture Notes in Computer Science, 2889:360– 370, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Burns:2004:RTS

[BW04] Alan Burns and Andy Wellings. Real-time systems and programming languages [sound recording]: Ada 95, real-time Java and real-time Posix. TPB, Enskede, The Netherlands, 2004. 2 CD–R (61h 54m) pp. LCCN ????

                                        URL http://katalog.tpb.
                                        se/wsHitList.Asp?SCode1=
                                        TN&SearchStr1=C30785.

Bergin:2005:TPE

[BWC^{+}05] Joseph Bergin, Eugene Wallingford, Michael Caspersen, Michael Goldweber, and Michael Kolling. Teaching polymorphism early. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):342– 343, September 2005. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Benedicenti:2001:EQC

[BWLP01] Luigi Benedicenti, Victor Wei Wang, Peter Lee, and Raman Paranjape. Establishing quality control in software agents. ACM SIGAPP Applied Computing Review, 9(3):31–33, September 2001. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
570132.570137.

Bentley:2006:IAB

[BWLR06] Chris Bentley, Scott A. Watterson, David K. Lowenthal, and Barry Rountree. Implicit array bounds checking on 64-bit architectures. ACM Transactions on Architecture and Code Optimization, 3(4):502–527, December 2006. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Brear:2003:SSJ

[BWW^{+}03] D. J. Brear, T. Weise, T. Wiffen, K. C. Yeung, S. A. M. Bennett, and P. H. J. Kelly. Search strategies for Java bottleneck location by dynamic instrumentation. IEE Proceedings Software, 150(4): 235–242, 2003. CODEN ???? ISSN 1462-5970.

Benaya:2005:APJ

[BZ05] Tamar Benaya and Ela Zur. Advanced programming in Java workshop: teaching methodology. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3): 348, September 2005. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Benaya:2007:UTA

[BZ07] Tamar Benaya and Ela Zur. Understanding threads in an advanced Java course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39 (3):323, September 2007. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Chan:2004:RTS

[CA04] Bryan Chan and Tarek S. Abdelrahman. Run-time support for the automatic parallelization of Java programs. The Journal of Supercomputing, 28(1):91–117, April 2004. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:/

/ipsapp008.kluweronline.
com/IPS/content/ext/x/J/
5189/I/52/A/6/abstract.
htm.

Caamano:2000:PJS

[Caa00] Paul Caamano. Porting a JAVA^{T M} Virtual Machine to an embedded system. Thesis (M.S.), University of California, Santa Cruz, Santa Cruz, CA, USA, 2000.

Cabana:2004:PPJ

[CAF04] Brian Cabana, Suad Alagić, and Jeff Faulkner. Parametric polymorphism for Java: is there any hope in sight? ACM SIGPLAN Notices, 39(12):22–31, December 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Calarco:2000:BRB

[Cal00a] Pascal V. Calarco. Book review: Developer’s guide to the Java Web server: Building effective and scalable serverside applications, by Dan Woods, Larne Pekowsky, and Tom Snee. Journal of the American Society for Information Science, 51(10):965– 966, ???? 2000. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Calsavara:2000:JQH

[Cal00b] Alexandre Pereira Calsavara.

Java Q&A: How can I extend Java’s search capabilities? Dr. Dobb’s Journal of Software Tools, 25(12): 141–142, 144, 146, December 2000. CODEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/ftp/2000/
2000_12/jqa0012.zip.

Callaway:2001:ISS

[Cal01] Dustin R. Callaway. Inside servlets: server-side programming for the Java platform. Addison-Wesley, Reading, MA, USA, second edition, 2001. ISBN 0-201-70906-6. xxv + 881 pp. LCCN QA76.73.J38 C35 2001.

Callaway:2002:VTR

[Cal02] John Callaway. Visualization of threads in a running Java program. Thesis (M.S.), University of California, Santa Cruz, Santa Cruz, CA, USA, 2002.

Calvert:2003:TIS

[Cal03] Kenneth Calvert. TCP/IP Sockets in Java: Practical Guide for Programmers. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2003. ISBN 1-55860-685-8. xii + 116 pp. LCCN QA76.625 .C35 2002. US$15.95.

Calejo:2004:ITD

[Cal04] M. Calejo. InterProlog: Towards a declarative embedding of logic programming in Java. Lecture Notes in Computer Science, 3229:714– 717, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carlisle:2006:AOP

[Car06] Martin C. Carlisle. Automatic OO parser generation using visitors for Ada 2005. ACM SIGADA Ada Letters, 26(3):3–8, December 2006. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Casset:2002:DEV

[Cas02] Ludovic Casset. Development of an embedded verifier for Java card byte code using formal methods. Lecture Notes in Computer Science, 2391:290–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2391/23910290.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2391/23910290.
pdf.

Cavalieri:2002:ERT

[Cav02a] S. Cavalieri. Exploring realtime features of Java VM. IECON Proceedings, 3(??): 2538–2543, 2002. CODEN ???? ISSN ????

Cavaness:2002:PJS

[Cav02b] Chuck Cavaness. Programming Jakarta Struts. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2002. ISBN 0-596-00328-5. xvii + 441 pp. LCCN QA76.76.A65 C39 2003. US$39.95. URL

http://www.oreilly.com/
catalog/jakarta.

Cavaness:2004:PJS

[Cav04] Chuck Cavaness. Programming Jakarta Struts. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2004. ISBN 0-596-00651-9. xvii + 450 pp. LCCN QA76.76.A65 C39 2004. US$44.95, CAN$65.95, UK£31.95. URL http://

www.oreilly.com/catalog/
0596006519/.

Chalasani:2004:AJB

[CB04] Suresh Chalasani and Robert Barber. Architectures for Java-based bioinformatics applications. Industrial Management & Data Systems, 104 (7):578–588, July 1, 2004. CODEN ???? ISSN 0263-5577 (print), 1758-5783 (electronic).

Christian:2001:PJT

[CBD01] Wolfgang Christian, Mario Belloni, and Melissa Dancy. Physlets: Java tools for a Web-based physics curriculum. Lecture Notes in Computer Science, 2073:1061–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2073/20731061.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2073/20731061.
pdf.

Cowlishaw:2004:FFE

[CBD04] M. Cowlishaw, J. Bloch, and J. D. Darcy. Fixed, floating, and exact computation with Java’s BigDecimal. Dr. Dobb’s Journal of Software Tools, 29(7):22–27, 2004. CO-DEN DDJOEB. ISSN 1044-789X.

Corwin:2003:MRM

[CBGM03] John Corwin, David F. Bacon, David Grove, and Chet Murthy. MJ: a rational module system for Java and its applications. ACM SIG-PLAN Notices, 38(11):241– 254, November 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chang:2001:EEJ

[CC01] Da-Wei Chang and Ruei-Chuan Chang. EJVM: an economic Java run-time environment for embedded devices. Software—Practice and Experience, 31(2):129– 146, February 2001. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (elec-

tronic). URL http://www3.
interscience.wiley.com/
cgi-bin/abstract/76504935/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=76504935&PLACEBO=IE.
pdf.

Christensen:2002:FCD

[CC02] Henrik Bærbak Christensen and Michael E. Caspersen.

Frameworks in CS1: a different way of introducing event-driven programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34 (3):75–79, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Corsaro:2003:EMR

[CC03] Angelo Corsaro and Ron K. Cytron. Efficient memoryreference checks for realtime Java. ACM SIGPLAN Notices, 38(7):51–58, July 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chang:2004:TSP

[CC04] B. M. Chang and J. D. Choi. Thread-sensitive points-to analysis for multithreaded Java programs. Lecture Notes in Computer Science, 3280:945–954, 2004. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Craig:2001:IJS

[CCB^{+}01] David Craig, Steven Carroll, Fabian Breg, Dimitrios S. Nikolopoulos, and Constantine Polychronopoulos. Improving Java server performance with interruptlets. Lecture Notes in Computer Science, 2073:223–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2073/20730223.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2073/20730223.
pdf.

Clarke:2009:JDR

[CCB09] Jim Clarke, Jim Connors, and Eric Bruno. JavaFX: Developing Rich Internet Applications. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2009. ISBN 0-13-701287-X. xxiii + 359 pp. LCCN QA76.73.J38.

Chen:2004:MES

[CCC^{+}04] C. W. Chen, C. K. Chen, J. C. Chen, C. T. Ko, J. K. Lee, H. W. Lin, and W. J. Wu. MM04-1: Efficient support of Java RMI over heterogeneous wireless networks. IEEE International Conference on Communications, 3: 1391–1395, 2004. CODEN ???? ISSN 0536-1486.

Carlstrom:2006:EJP

[CCC^{+}06] Brian D. Carlstrom, Jae-Woong Chung, Hassan Chafi, Austen McDonald, Chi Cao Minh, Lance Hammond, Christos Kozyrakis, and Kunle Olukotun. Executing Java programs with transactional memory. Science of Computer Programming, 63(2):111–129, December 1, 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

ColindeVerdiere:2002:SPS

[CCF^{+}02] Vincent Colin de Verdière, Sébastien Cros, Christian Fabre, Romain Guider, and Sergio Yovine. Speedup prediction for selective compilation of embedded Java programs. Lecture Notes in Computer Science, 2491:227–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/
                                        service/series/0558/bibs/
                                        2491/24910227.htm;   http:
                                        //link.springer.de/link/
                                        service/series/0558/papers/
                                        2491/24910227.pdf.

Caromel:2000:WJP

[CCFG00] Denis Caromel, Serge Chaumette, Geoffrey Fox, and Peter Graham. Workshop on Java for parallel and distributed computing. Lecture Notes in Computer Science, 1800:526– ??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.
                                        com/link/service/series/
                                        0558/bibs/1800/18000526.
                                        htm;   http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/1800/18000526.
                                        pdf.

Chen:2008:MJR

[CCK^{+}08] Chung-Kai Chen, Cheng-Wei Chen, Chien-Tan Ko, Jenq-Kuen Lee, and Jyh-Cheng Chen. Mobile Java RMI support over heterogeneous wireless networks: a case study. Journal of Parallel and Distributed Computing, 68(11): 1425–1436, November 2008. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Chin:2006:FBAa

[CCKP06] Wei-Ngan Chin, Florin Craciun, Siau-Cheng Khoo, and Corneliu Popeea. A flow-based approach for variant parametric types. ACM SIG-PLAN Notices, 41(10):273– 290, October 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Choi:2005:JMA

[CCM05] Hyung-Kyu Choi, Yoo C.

Chung, and Soo-Mook Moon. Java memory allocation with lazy worst fit for small objects. The Computer Journal, 48(4):437–442, July 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
48/4/437;   http://comjnl.
oxfordjournals.org/cgi/
reprint/48/4/437.

Caprotti:2000:JPC

[CCR00] Olga Caprotti, Arjeh M.

Cohen, and Manfred Riem.

JAVA phrasebooks for computer algebra and automated deduction. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 34(2):33–37, June 2000. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Cruz:2002:SRA

[CCSA02] J. L. Cruz, B. Cuesta, M. Sigut, and L. Acosta. Simulation, remote access, and monitoring of a robot in Java. International Journal of Robotics and Automation, 17(3):112–119, 2002. CODEN IJAUED. ISSN 0826-8185.

Clamp:2004:JJA

[CCSB04] M. Clamp, J. Cuff, S. M. Searle, and G. J. Barton. The Jalview Java alignment editor. Bioinformatics, 20(3): 426–427, 2004. CODEN ???? ISSN 1367-4803 (print), 1367-4811 (electronic).

Chen:2001:JJB

[CCT01] Ding-Yi Chen, Tyng-Ruey Chuang, and Shi-Chun Tsai. JGAP: a Java-based graph algorithms platform. Software —Practice and Experience, 31 (7):615–635, June 2001. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (elec-

                                        tronic). URL http://www3.
                                        interscience.wiley.com/
                                        cgi-bin/abstract/78003148/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?

ID=78003148&PLACEBO=IE.
pdf.

Chen:2002:JPU

[CCW02] F. Chen, S. Chen, and X. Wu.

Java program used in block graph experimental simulation on Web. Journal- Wuhan University Natural Sciences Edition, 48(5):621–625, 2002. CODEN ???? ISSN 0253-9888.

Calvert:2001:TIS

[CD01a] Kenneth L. Calvert and Michael J. Donahoo. TCP/ IP sockets in Java: practical guide for programmers. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2001. ISBN 1-55860-685-8. 80 pp. LCCN TK5105.585 .C34 2001.

Christiaens:2001:TTA

[CD01b] Mark Christiaens and Koen De Bosschere. TRaDe: a topological approach to onthe-fly race detection in Java programs. In USENIX Association [USE01c], page ?? ISBN 1-880446-11-1. LCCN QA76.73.J38 J42 2001. URL

http://www.usenix.org/publications/
library/proceedings/jvm01/
christiaens.html. Spon-

sored by the USENIX Association.

Christiaens:2001:TDR

[CD01c] Mark Christiaens and Koen De Bosschere. TRaDe: Data race detection for Java. Lecture Notes in Computer Science, 2074:761–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2074/20740761.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2074/20740761.
pdf.

Comp:2003:RAW

[CD03] Lynn Comp and Tim Dobbing. Runtime abstractions in the wireless and handheld space. Intel Technology Journal, 7(1):68–76, February 2003. ISSN 1535-766X. URL

http://developer.intel.
com/technology/itj/2003/
volume07issue01/art07_mobile/
p01_abstract.htm.

Chern:2008:ISD

[CD08] Rick Chern and Kris De Volder. The impact of staticdynamic coupling on remodularization. ACM SIGPLAN Notices, 43(10):261–276, September 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cimato:2005:OOJ

[CDF05] S. Cimato, A. De Santis, and U. Ferraro Petrillo. Overcoming the obfuscation of Java programs by identifier renaming. The Journal of systems and software, 78(1):

60–72, October 2005. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Corradini:2004:TJC

[CDFR04] A. Corradini, F. L. Dotti, L. Foss, and L. Ribeiro. Translating Java code to graph transformation systems. Lecture Notes in Computer Science, 3256:383–398, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chambers:2007:AIR

[CDH07] Mike Chambers, Daniel Dura, and Kevin Hoyt. Adobe Integrated Runtime (Air): for Javascript Developers Pocket Guide. Adobe developer library. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2007. ISBN 0-596-51519-7. xix + 154 pp. LCCN QA76.625; QA76.625 .C42 2007eb. URL http://

www.oreilly.com/catalog/
9780596515195.

Cameron:2007:MO

[CDNS07] Nicholas R. Cameron, Sophia Drossopoulou, James Noble, and Matthew J. Smith. Multiple ownership. ACM SIG-PLAN Notices, 42(10):441– 460, October 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cocosco:2001:JIV

[CE01] Chris A. Cocosco and Alan C. Evans. Java Internet viewer: a WWW tool for remote 3D medical image data visualization and comparison. Lecture Notes in Computer Science, 2208:1415–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2208/22081415.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2208/22081415.
pdf.

Cierniak:2003:ORP

[CEG^{+}03] Michal Cierniak, Marsha Eng, Neal Glew, Brian Lewis, and James Stichnoth. The Open Runtime Platform: a flexible high-performance managed runtime environment. Intel Technology Journal, 7(1):5–18, February 2003. ISSN 1535-766X. URL

http://developer.intel.
com/technology/itj/2003/
volume07issue01/art01_orp/
p01_abstract.htm.

Cerami:2002:WSE

[Cer02] Ethan Cerami. Web Services Essentials: Distributed Applications with XML-RPC, SOAP, UDDI & WSDL. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2002. ISBN 0-596-00224-6. xiii + 288 pp. LCCN TK5105.888 .C47 2002. US$29.95. URL http:/

/safari.oreilly.com/0596002246;
 http://www.oreilly.com/
catalog/webservess.

Chelius:2000:ING

[CF00] Guillaume Chelius and Éric Fleury. An IP next generation compliant Java^{T M} virtual machine. Lecture Notes in Computer Science, 1800:528– ??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/1800/18000528.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/1800/18000528.
pdf.

Clear:2002:ACJ

[CF02] T. Clear and G. Foot.

Avatars in cyberspace — A Java 3D application to support formation of virtual groups. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):222, 2002. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Carpenter:2003:HDP

[CF03] Bryan Carpenter and Geoffrey Fox. HPJava: a data parallel programming alternative. Computing in Science and Engineering, 5(3): 60–64, May/June 2003. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        cs/2003/03/c3060abs.htm;
                                         http://csdl.computer.org/
                                        dl/mags/cs/2003/03/c3060.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cs/2003/03/
                                        c3060.pdf.

Conrad:2004:ESB

[CF04a] Marc Conrad and Tim French. Exploring the synergies between the object oriented paradigm and mathematics: a Java led approach. International Journal of Mathematical Education in Science and Technology, 35(5):733–742, September 2004. CODEN ???? ISSN 0020-739X (print), 1464-5211 (electronic).

Conrad:2004:USB

[CF04b] Marc Conrad and Tim French. Using the synergies between the object-oriented paradigm and mathematics in joint mathematics/computer science programs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3): 254, September 2004. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cohen:2005:AIC

[CFGL05] Robert F. Cohen, Alexander V. Fairley, David Gerry, and Gustavo R. Lima. Accessibility in introductory computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1): 17–21, March 2005. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Carpenter:2000:OSM

[CFKL00] Bryan Carpenter, Geoffrey Fox, Sung Hoon Ko, and Sang Lim. Object serialization for marshaling data in a Java interface to MPI. Concurrency: Practice and Experience, 12 (7):539–553, May 2000. CO-DEN CPEXEI. ISSN 1040-

3108. URL http://www3.
interscience.wiley.com/
cgi-bin/abstract/72516217/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=72516217&PLACEBO=IE.
pdf.

Cabri:2005:IRJ

[CFL05a] G. Cabri, L. Ferrari, and L. Leonardi. Injecting roles in Java agents through runtime bytecode manipulation. IBM Systems Journal, 44(1): 185–208, ???? 2005. CODEN IBMSA7. ISSN 0018-8670.

URL http://www.research.
ibm.com/journal/sj/441/
cabri.html;   http://www.
research.ibm.com/journal/
sj/441/cabri.pdf;   http:
//www.research.ibm.com/
journal/sj/441/cabri.txt.

Cabri:2005:ERB

[CFL05b] Giacomo Cabri, Luca Ferrari, and Letizia Leonardi. Exploiting runtime bytecode manipulation to add roles to Java agents. Science of Computer Programming, 54 (1):73–98, January 2005. CO-DEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Carpenter:2003:AHJ

[CFLL03a] B. Carpenter, G. Fox, H. K. Lee, and S. B. Lim. Applications of HP Java. Lecture Notes in Computer Science, 2958:147–161, 2003. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carpenter:2003:TSH

[CFLL03b] B. Carpenter, G. Fox, H. K. Lee, and S. B. Lim. Translation schemes for the HP Java parallel programming language. Lecture Notes in Computer Science, 2624:18– 32, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chandra:2009:SPA

[CFS09] Satish Chandra, Stephen J. Fink, and Manu Sridharan. Snugglebug: a powerful approach to weakest preconditions. ACM SIGPLAN Notices, 44(6):363–374, June 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Coglio:2001:TSJ

[CG01] Alessandro Coglio and Allen Goldberg. Type safety in the JVM: some problems in Java 2 SDK 1.2 and proposed solutions. Concurrency and Computation: Practice and Experience, 13(13):1153– 1171, November 2001. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (elec-

tronic). URL http://www3.
interscience.wiley.com/
cgi-bin/abstract/88011336/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=88011336&PLACEBO=IE.
pdf.

Chen:2002:POS

[CG02] Kirk Chen and Li Gong. Programming open service gateways with Java embedded server technology. Java series. Addison-Wesley, Reading, MA, USA, 2002. ISBN 0-201-71102-8. xix + 456 pp. LCCN QA76.73.J38 C4775 2002.

Casey:2003:TSJ

[CGEN03] K. Casey, D. Gregg, M. A.

Ertl, and A. Nisbet. Towards superinstructions for Java interpreters. Lecture Notes in Computer Science, 2826:329– 343, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chiu:2002:PMM

[CGG02] Kenneth Chiu, Madhusudhan Govindaraju, and Dennis Gannon. The Proteus Multiprotocol Message Library. In IEEE [IEE02a], page ?? ISBN 0-7695-1524-X. LCCN

                                        ???? URL http://www.sc-
                                        2002.org/paperpdfs/pap.
                                        pap315.pdf.

Carpenter:2000:MML

[CGJ^{+}00] Bryan Carpenter, Vladimir Getov, Glenn Judd, Anthony Skjellum, and Geoffrey Fox. MPJ: MPIlike message passing for Java. Concurrency: Practice and Experience, 12(11):1019– 1038, September 2000. CO-DEN CPEXEI. ISSN 1040-

                                        3108. URL http://www3.
                                        interscience.wiley.com/
                                        cgi-bin/abstract/76000188/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=76000188&PLACEBO=IE.
                                        pdf.

Cohen:2006:JJTa

[CGM06] Tal Cohen, Joseph (Yossi) Gil, and Itay Maman. JTL: the Java tools language. ACM SIGPLAN Notices, 41 (10):89–108, October 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ciancarini:2000:MCD

[CGR00] Paolo Ciancarini, Andrea Giovannini, and Davide Rossi. Mobility and coordination for distributed Java applications. Lecture Notes in Computer Science, 1752:402– ??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/1752/17520402.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/1752/17520402.
pdf.

Comeau:2004:UOP

[CGRR04] T. Comeau, B. Garrett, J. Richon, and F. Romelfanger. Using OpenOffice as a portable interface to Java-based applications. Astronomical Society of the Pacific Conference Series, 314:701–703, 2004. CO-DEN ???? ISSN 1050-3390.

Choi:2003:SAS

[CGS^{+}03] Jong-Deok Choi, Manish Gupta, Mauricio J. Serrano, Vugranam C. Sreedhar, and Samuel P. Midkiff. Stack allocation and synchronization optimizations for Java using escape analysis. ACM Transactions on Programming Languages and Systems, 25(6): 876–910, November 2003. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Catano:2002:FSS

[CH02] Néstor Cataño and Marieke Huisman. Formal specification and static checking of Gemplus’ electronic purse using ESC/Java. Lecture Notes in Computer Science, 2391:272–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.
                                        com/link/service/series/
                                        0558/bibs/2391/23910272.
                                        htm;   http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2391/23910272.
                                        pdf.

Cross:2006:JLI

[CH06] James H. Cross II and T. Dean Hendrix. jGRASP: a lightweight IDE with dynamic object viewers for CS1 and CS2. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):356, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Choi:2008:SHM

[CH08] Yoonseo Choi and Hwansoo Han. Shared heap management for memory-limited Java virtual machines. ACM Transactions on Embedded Computing Systems, 7(2): 13:1–13:??, February 2008. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Chalk:2000:CCC

[Cha00a]  Peter Chalk.  Conference

corner: Conference report: JICC4: Java in the Computing Curricula. ACM SIGPLAN Notices, 35(3):18– 19, March 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chalk:2000:JJC

[Cha00b] Peter Chalk. JICC4: Java in the computing curricula. ACM SIGPLAN Notices, 35(3):18–19, March 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chapman:2000:JES

[Cha00c] Stephen J. Chapman. Java for Engineers and Scientists. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2000. ISBN 0-13-919523-8. xix + 572 pp. LCCN QA76.73.J38C477 1999. US$45.00. URL http:/

/www.phptr.com/ptrbooks/
esm_0139195238.html.

Chaudhri:2002:JD

[Cha02] Akmal B. Chaudhri, editor.

Java and databases, Innovative technology series: information systems and networks. Hermes Penton Science, London, UK, 2002. ISBN 1-903996-15-5. LCCN QA76.73.J38 J38 2002. Based on revised papers submitted to the OOPSLA ’99 workshop on “Java and Databases:

Persistence Options”, held in Denver, Colorado, in November 1999.

Chavez:2003:BRH

[Cha03] Hector Zenil Chavez. Book review: How to develop enterprise-secure Java applications learning risks A review of Hacking Exposed J2EE and Java, Developing Secure Applications with Java Technology by Art Taylor, Brian Buege, and Randy Layman. IEEE Distributed Systems Online, 4(5), 2003. ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:

                                        //dsonline.computer.org/
                                        0305/d/bks_c.htm.

Chang:2005:RIR

[Cha05a] R. S. Chang. Review of Integration-Ready Architecture and Design: Software Engineering with XML, Java, .NET, Wireless, Speech, and Knowledge Technologies by Jeff Zhuk, Cambridge University Press, 2004, $75.00, ISBN 0-521-52583-7. ACM Queue: Tomorrow’s Computing Today, 3(2):58, March 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Chavez:2005:JFE

[Cha05b] H. Z. Chavez. JavaServer Faces: An excellent learning tool. IEEE Distributed Systems Online, 6(2):6, February 2005. CODEN ????

ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://ieeexplore.ieee.
org/iel5/8968/30522/01407763.
pdf?isnumber=30522&prod=
JNL&arnumber=1407763&arSt=
+6&ared=+6&arAuthor=Chavez%
2C+H.Z.;   http://ieeexplore.
ieee.org/xpls/abs_all.jsp?
isnumber=30522&arnumber=
1407763&count=5&index=4.

Chang:2006:SCA

[Cha06] Byeong-Mo Chang. Static check analysis for Java stack inspection. ACM SIGPLAN Notices, 41(3):40–48, March 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chetty:2003:IJB

[CHB03] M. Chetty, S. Hu, and J. Bennett. An interactive Javabased educational module in electromagnetics. International Journal of Electrical Engineering Education, 40(1): 79–90, 2003. CODEN ???? ISSN 0020-7209.

Chen:2000:JCT

[Che00] Zhiqun Chen. Java Card technology for Smart Cards: architecture and programmer’s guide. Java series. Addison-Wesley, Reading, MA, USA, 2000. ISBN 0-201-70329-

7. xxii + 368 pp. LCCN QA76.73.J38 C478 2000.

Chen:2002:FMJ

[Che02a] J. Chen. Formal modelling of Java GUI event handling. Lecture Notes in Computer Science, 2495:359–370, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:2002:JCN

[Che02b]  Jiadong Chen.  Java E-

commerce in a nutshell: a review of Professional Java E-Commerce. IEEE Distributed Systems Online, 3(4), 2002. ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:

                                        //dsonline.computer.org/
                                        0204/departments/bks_a.
                                        htm.

Chen:2003:RFJ

[Che03a] G. Chen. Research on framework for Java based 3D collaborative virtual environment. Computer Integrated Manufacturing Systems, 9 (SUPP):6–9, 2003. CODEN ???? ISSN 1006-5911.

Chen:2003:FMJ

[Che03b] Jessica Chen. Formal modelling of Java GUI event handling. Lecture Notes in Computer Science, 2495:359–??, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/
                                        service/series/0558/bibs/
                                        2495/24950359.htm;   http:
                                        //link.springer.de/link/

service/series/0558/papers/
2495/24950359.pdf.

Chen:2003:RAS

[Che03c] X. Chen. Realization and application of shared memory in Java. Journal — Anhui University of Technology, 20(3): 224–226, 2003. CODEN ????

Che:2005:REC

[Che05] Haoyang Che. Review of The Elements of C++ Style by Trevor Misfeldt, Gregory Bumgardner, and Andrew Gray; Cambridge University Press: Cambridge, 2004, 0-521-89308-9. ACM SIGSOFT Software Engineering Notes, 30(2):29, March 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Chen:2004:MCP

[CHHC04] D. J. Chen, C. C. Hwang, S. K. Huang, and D. T. K. Chen. Mining control patterns from Java program corpora. Journal of Information Science and Engineering, 20 (1):57–84, 2004. CODEN JI-NEEY. ISSN 1016-2364.

Chiba:2000:LTS

[Chi00] Shigeru Chiba. Load-time structural reflection in Java. Lecture Notes in Computer Science, 1850:313–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.

com/link/service/series/
0558/bibs/1850/18500313.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/1850/18500313.
pdf.

Cross:2007:DOV

[CHJB07] James H. Cross II, T. Dean Hendrix, Jhilmil Jain, and Larry A. Barowski. Dynamic object viewers for data structures. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):4–8, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Csopaki:2000:CPI

[CHK00] Gyula Csopaki, Gábor András Horváth, and Gábor Kovács. Communication protocol implementation in Java. Lecture Notes in Computer Science, 1905:254–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/1905/19050254.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/1905/19050254.
pdf.

Coglio:2004:FTJ

[CHK^{+}04] A. Coglio, M. Huisman, J. R. Kiniry, P. Muller, and E. Poll. Formal techniques for Javalike programs (FTfJP). Lecture Notes in Computer Science, 3344:76–83, 2004. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Christ:2000:SFP

[CHL^{+}00] R. Christ, S. L. Halter, K. Lynne, S. Meizer, S. J. Munroe, and M. Pasch. San Francisco performance: a case study in performance for large-scale Java applications. IBM Systems Journal, 39(1):4–20, ???? 2000. CO-DEN IBMSA7. ISSN 0018-

8670. URL http://www.
almaden.ibm.com/journal/
sj/391/christ.html.

Chen:2007:TPB

[CHL07] Hsiang-Yang Chen, Ting-Wei Hou, and Chun-Liang Lin. Tamper-proofing basis path by using oblivious hashing on Java. ACM SIGPLAN Notices, 42(2):9–16, February 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chan:2004:JIP

[CHMB04] M. Chan, T. Hansen, P. A.

Monney, and T. L. Baker. A Java implementation of the probabilistic argumentation system for data fusion in missile defense applications [5434-19]. Proceedings of the SPIE — The International Society for Optical Engineering, 5434:176–186, 2004.

CODEN PSISDG. ISSN 0277-786X (print), 1996-756X (electronic).

Chen:2008:TPC

[CHP^{+}08] Juan Chen, Chris Hawblitzel, Frances Perry, Mike Emmi, Jeremy Condit, Derrick Coetzee, and Polyvios Pratikaki. Type-preserving compilation for large-scale optimizing object-oriented compilers. ACM SIGPLAN Notices, 43(6):183–192, June 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Christian:2000:JPI

[Chr00] Wolfgang Christian. Java programming and Internet technologies for undergraduate education. Computer Physics Communications, 127 (1):16–22, May 1, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0010465500000187.

Christiaens:2001:JRR

[Chr01] Mark Christiaens. JaRec: Record/replay for multithreaded Java programs. In USENIX Association [USE01c], page ?? ISBN 1-880446-11-1. LCCN QA76.73.J38

                                        J42 2001. URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/jvm01/
                                        JVM_wips/S07.pdf. Spon-

sored by the USENIX Association.

Christensen:2005:TLJ

[Chr05] Henrik Bærbak Christensen.

TS-05: 150 lines of Java with high architectural complexity. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37 (3):396, September 2005. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Caromel:2001:CIS

[CHS01] Denis Caromel, Ludovic Henrio, and Bernard Serpette. Context inference for static analysis of Java card object sharing. Lecture Notes in Computer Science, 2140:43– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2140/21400043.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2140/21400043.
pdf.

Czajkowski:2005:RMI

[CHS^{+}05] Grzegorz Czajkowski, Stephen Hahn, Glenn Skinner, Pete Soper, and Ciarán Bryce. A resource management interface for the Java^{T M} platform. Software—Practice and Experience, 35(2):123–157, February 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cross:2008:EAV

[CHUB08] James H. Cross II, T. Dean Hendrix, David A. Umphress, and Larry A. Barowski. Exploring accessibility and visibility relationships in Java. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40 (3):103–108, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Caromel:2001:SSA

[CHV01] Denis Caromel, Fabrice Huet, and Julien Vayssière. A simple security-aware MOP for Java. Lecture Notes in Computer Science, 2192:118–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.
                                        com/link/service/series/
                                        0558/bibs/2192/21920118.
                                        htm;   http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2192/21920118.
                                        pdf.

Cattell:2001:JPB

[CI01] R. G. G. (Roderic Geoffrey Galton) Cattell and Jim Inscore. J2EE in Practice: Building Business Applications with the Java 2 Platform, Enterprise Edition. The Java series. Addison-Wesley, Reading, MA, USA, 2001. ISBN 0-201-74622-0. xvii + 309 pp. LCCN QA76.73.J38 C38 2001.

Choi:2001:CLF

[CiLH01] Kwanghoon Choi, Hyun il Lim, and Taisook Han. Compiling lazy functional programs based on the spineless tagless G-machine for the Java virtual machine. Lecture Notes in Computer Science, 2024:92–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2024/20240092.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2024/20240092.
pdf.

Cimato:2002:DAP

[Cim02]  Stelvio Cimato.  Design

of an authentication protocol for Gsm Javacards. Lecture Notes in Computer Science, 2288:355–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2288/22880355.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2288/22880355.
pdf.

Chappell:2002:JWS

[CJ02] David A. Chappell and Tyler Jewell. Java Web services.

O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2002. ISBN 0-596-00269-6. xii + 262 pp. LCCN QA76.73.J38 C4772 2002. URL http://

www.oreilly.com/catalog/
9780596002695. Using Java

in service-oriented architectures.

Cavaness:2003:JSP

[CK03a] Chuck Cavaness and Brian Keeton. Jakarta Struts pocket reference. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2003. ISBN 0-596-00519-9. viii + 134 pp. LCCN QA76.76.A65 C385 2003. URL http://

www.oreilly.com/catalog/
9780596005191.

Crawford:2003:JDP

[CK03b] William Crawford and Jonathan Kaplan. J2EE design patterns. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2003. ISBN 0-596-00427-3. xvi + 350 pp. LCCN QA76.73.J38 C74 2003. URL

http://www.oreilly.com/
catalog/9780596004279.

Cok:2005:EJU

[CK05] D. R. Cok and J. R. Kiniry. ESC/Java2: Uniting ESC/ Java and JML — progress and issues in building and using ESC/Java2, including a case study involving the use of the tool to verify portions of an Internet voting tally system. Lecture Notes in Computer Science, 3362:108– 128, 2005. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chiao:2002:EBR

[CKC^{+}02] H. T. Chiao, S. H. J. Kao, Y. S. Chang, S. T. Huang, and S. M. Yuan. Experience in building a real-time extension library for Java. Journal of Information Science and Engineering, 18(6):905– 928, 2002. CODEN JINEEY. ISSN 1016-2364.

Chen:2004:SET

[CKK^{+}04] Guangyu Chen, Byung-Tae Kang, Mahmut Kandemir, Narayanan Vijaykrishnan, Mary Jane Irwin, and Rajarathnam Chandramouli. Studying energy trade offs in offloading computation/ compilation in Java-enabled mobile devices. IEEE Transactions on Parallel and Distributed Systems, 15(9): 795–809, September 2004. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//csdl.computer.org/dl/
trans/td/2004/09/l0795.
htm;   http://csdl.computer.
org/dl/trans/td/2004/09/
l0795.pdf.

Chung:2003:JBD

[CKKH03] M. J. Chung, S. Kim, H. Kim, and H. S. Ham. A Java-based, distributed process management system for collaborative design and manufacturing. Lecture Notes in Computer Science, 2604:61–72, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Christensen:2004:RSX

[CKM04] A. S. Christensen, C. Kirkegaard, and A. Moller. A runtime system for XML transformations in Java. Lecture Notes in Computer Science, 3186:143– 157, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cole:2009:MPC

[CKMP09] Marilyn C. Cole, Evan Korth, Adam Meyers, and Sam Pluta. Musicomputation: a pilot course exploring a precollege computer science curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):381, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Chen:2002:UMC

[CKV^{+}02] G. Chen, M. Kandemir, N. Vijaykrishnan, M. J. Irwin, and W. Wolf. Using memory compression for energy reduction in an embedded Java system. Journal of Circuits Systems and Computers, 11(5): 537–556, 2002. CODEN ???? ISSN 0218-1266.

Chen:2003:HCM

[CKV^{+}03] G. Chen, M. Kandemir, N. Vijaykrishnan, M. J. Irwin, B. Mathiske, and M. Wolczko. Heap compression for memory-constrained Java environments. ACM SIG-PLAN Notices, 38(11):282– 301, November 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cadenhead:2003:STY

[CL03a] Rogers Cadenhead and Laura Lemay. Sams teach yourself Java 2 in 21 days. Sams teach yourself in 21 days. Howard W. Sams, Indianapolis, IN 46268, USA, third edition, 2003. ISBN 0-672-32370-2. xviii + 712 pp. LCCN QA76.73.J3 L44

2003. URL http://www.
java21days.com.

Chung:2003:MWA

[CL03b] S. Chung and Y. S. Lee. Modeling Web applications using Java and XML related technologies (STDOC09). Proceedings of the Annual Hawaii International Conference on System Sciences, CONF36: 322, 2003. CODEN ???? ISSN 1060-3425.

Corliss:2008:BCJ

[CL08] Marc L. Corliss and E. Christopher Lewis. Bantam: a customizable, Java-based, classroom compiler. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1): 38–42, March 2008. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Clark:2004:PPA

[Cla04]  Mike Clark.  Pragmatic

project automation: how to build, deploy, and monitor Java applications, volume 3 of Pragmatic starter kit series. Pragmatic Bookshelf, Raleigh, NC, USA, 2004. ISBN 0-9745140-3-9. xiv + 161 pp. LCCN QA76.76.D47 C53 2004. URL http://

www.oreilly.com/catalog/
9780974514031.

Cha:2002:IXB

[CLCC02] Si-Ho Cha, Jae-Oh Lee, Young-Keun Choi, and Kook-Hyun Cho. Implementing an XML-based universal network management system in Java. Lecture Notes in Computer Science, 2343:652–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/
service/series/0558/bibs/
2343/23430652.htm;   http:
//link.springer.de/link/
service/series/0558/papers/
2343/23430652.pdf.

Clifton:2000:MMO

[CLCM00] Curtis Clifton, Gary T. Leavens, Craig Chambers, and Todd Millstein. MultiJava: modular open classes and symmetric multiple dispatch for Java. ACM SIGPLAN Notices, 35(10):130–145, October 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
proceedings/oops/353171/
p130-clifton/.

Cleaveland:2001:PGJ

[Cle01a] J. Craig Cleaveland. Program generators with Java and XML. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2001. ISBN 0-13-025878-4. xxviii + 415 pp. LCCN QA76.73.J38 C54 2001. Includes CD-ROM.

Cleaveland:2001:PGX

[Cle01b] J. Craig Cleaveland. Program generators with XML and Java. The Charles F. Goldfarb series on open information management. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2001. ISBN 0-13-025878-4. xxviii + 415 pp. LCCN QA76.73.J38 C54 2001. Includes CD-ROM.

Chen:2003:GMD

[CLL03] L. Chen, C. LiWang, and F. C. M. Lau. A Grid middleware for distributed Java computing with MPI binding and process migration supports. Journal of Computer Science and Technology, 18 (4):505–514, 2003. CODEN JCTEEM. ISSN 1000-9000.

Chong:2007:SWA

[CLM^{+}07] Stephen Chong, Jed Liu, Andrew C. Myers, Xin Qi, K. Vikram, Lantian Zheng, and Xin Zheng. Secure Web application via automatic partitioning. Operating Systems Review, 41(6): 31–44, December 2007. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chong:2009:BSW

[CLM^{+}09] Stephen Chong, Jed Liu, Andrew C. Myers, Xin Qi, K. Vikram, Lantian Zheng, and Xin Zheng. Building secure Web applications with automatic partitioning. Communications of the ACM, 52(2):79–87, February 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Colby:2000:CCJ

[CLN^{+}00] Christopher Colby, Peter Lee, George C. Necula, Fred Blau, Mark Plesko, and Kenneth Cline. A certifying compiler for Java. ACM SIG-PLAN Notices, 35(5):95– 107, May 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL

                                        http://www.acm.org/pubs/
                                        articles/proceedings/pldi/
                                        349299/p95-colby/p95-colby.

pdf;   http://www.acm.org/
pubs/citations/proceedings/
pldi/349299/p95-colby/.

Counsell:2007:QMD

[CLN07] Steve Counsell, George Loizou, and Rajaa Najjar. Quality of manual data collection in Java software: an empirical investigation. Empirical Software Engineering, 12 (3):275–293, June 2007. CO-DEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (elec-

tronic). URL http://link.
springer.com/accesspage/
article/10.1007/s10664-
006-9028-y.

Crescenzi:2006:ACJ

[CLP06] Pilu Crescenzi, Michele Loreti, and Rosario Pugliese. Assessing CS1 Java skills: a threeyear experience. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3): 348, September 2006. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cierniak:2000:PJJ

[CLS00] Michał Cierniak, Guei-Yuan Lueh, and James M. Stichnoth. Practicing JUDO: Java under dynamic optimizations. ACM SIG-PLAN Notices, 35(5):13– 26, May 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL

                                        http://www.acm.org/pubs/
                                        articles/proceedings/pldi/
                                        349299/p13-cierniak/p13-
                                        cierniak.pdf;   http://
                                        www.acm.org/pubs/citations/
                                        proceedings/pldi/349299/
                                        p13-cierniak/.

Cunningham:2006:UCP

[CLZ06] H. Conrad Cunningham, Yi Liu, and Cuihua Zhang. Using classic problems to teach Java framework design. Science of Computer Programming, 59(1–2):147– 169, January 2006. CO-DEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Cappello:2001:SRN

[CM01] Peter Cappello and Dimitrios Mourloukos. A scalable, robust network for parallel computing. In ACM [ACM01b], pages 78–86. ISBN 1-58113-359-6. LCCN QA76.9.O35

                                        A26 2001. URL http://www.
                                        philippsen.com/JGI2001/
                                        camerareadyabstracts/53.
                                        html;   http://www.philippsen.
                                        com/JGI2001/finalpapers/
                                        18500078.pdf.

Cheng:2002:JBT

[CM02] L. Cheng and I. Marsic. Javabased tools for accurate bandwidth measurement of Digital Subscriber Line networks. Integrated Computer Aided Engineering, 9(4):333–344, 2002. CODEN ???? ISSN 1069-2509.

Chen:2004:JFC

[CM04] N. Chen and K. K. Ma. Java’s future: Challenge and opportunity. IT Professional, 6(4): 19–27, 2004. CODEN ???? ISSN 1520-9202.

Cahoon:2005:RAE

[CM05a] Brendon Cahoon and Kathryn S.

McKinley. Recurrence analysis for effective array prefetching in Java. Concurrency and Computation: Practice and Experience, 17(5–6):589– 616, April/May 2005. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cepa:2005:MGM

[CM05b] V. Cepa and M. Mezini. Mob-Con: a generative middleware framework for Java mobile applications. Proceedings of the Annual Hawaii International Conference on System Sciences, CONF38:283, 2005. CODEN ???? ISSN 1060-3425.

Chen:2005:IPF

[CM05c] Stephen Chen and Stephen Morris. Iconic programming for flowcharts, Java, Turing, etc. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):104–107, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Chen:2001:JSM

[CMB^{+}01] Deqing Chen, Alan Messer, Philippe Bernadat, Guangrui Fu, Zoran Dimitrijevic, David Jeun Fung Lie, Durga Mannaru, Alma Riska, and Dejan Milojicic. JVM susceptibility to memory errors. In USENIX Association [USE01c], page ?? ISBN 1-880446-11-1. LCCN QA76.73.J38 J42 2001. URL

http://www.usenix.org/publications/
library/proceedings/jvm01/

chen.html. Sponsored by the USENIX Association.

Carlstrom:2006:ATP

[CMC^{+}06] Brian D. Carlstrom, Austen McDonald, Hassan Chafi, JaeWoong Chung, Chi Cao Minh, Christos Kozyrakis, and Kunle Olukotun. The ATOMOS transactional programming language. ACM SIGPLAN Notices, 41(6):1– 13, June 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Campo:2001:JFC

[CMG^{+}01] Celeste Campo, Andrés Marm, Arturo García, Ignacio Díaz, Peter T. Breuer, Carlos Delgado, and Carlos García. JCCM: Flexible certificates for smartcards with Java card. Lecture Notes in Computer Science, 2140:34– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2140/21400034.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2140/21400034.
pdf.

Chugh:2009:SIF

[CMJL09] Ravi Chugh, Jeffrey A. Meister, Ranjit Jhala, and Sorin Lerner. Staged information flow for JavaScript. ACM SIGPLAN Notices, 44(6):50– 62, June 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Clifton:2006:MDR

[CMLC06] Curtis Clifton, Todd Millstein, Gary T. Leavens, and Craig Chambers. Multi-Java: Design rationale, compiler implementation, and applications. ACM Transactions on Programming Languages and Systems, 28(3): 517–575, May 2006. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Contreras:2007:XPP

[CMP^{+}07] Gilberto Contreras, Margaret Martonosi, Jinzhang Peng, Guei-Yuan Lueh, and Roy Ju. The XTREM power and performance simulator for the Intel XScale core: Design and experiences. ACM Transactions on Embedded Computing Systems, 6(1):4:1–4:??, February 2007. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Cirstea:2005:RBP

[CMR05] H. Cirstea, P. E. Moreau, and A. Reilles. Rule-based programming in Java for protocol verification. Electronic Notes in Theoretical Computer Science, 117(??):209–227, 2005. CODEN ???? ISSN 1571-0661.

Chow:2003:EJP

[CMS03a] Kingsum Chow, Ricardo Morin, and Kumar Shiv. Enterprise Java performance: Best practices. Intel Technology Journal, 7(1):32– 46, February 2003. ISSN 1535-766X. URL http://

                                        developer.intel.com/technology/
                                        itj/2003/volume07issue01/
                                        art03_java/p01_abstract.
                                        htm.

Christensen:2003:EJH

[CMS03b] Aske Simon Christensen, Anders Møller, and Michael I. Schwartzbach. Extending Java for high-level Web service construction. ACM Transactions on Programming Languages and Systems, 25(6):814–875, November 2003. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Chang:2005:EJG

[CMS05] Aaron N. Chang, Jason Mc-Dermott, and Ram Samudrala. An enhanced Java graph applet interface for visualizing interactomes. Bioinformatics, 21(8):1741–1742, April 15, 2005. CODEN ???? ISSN 1367-4803 (print), 1367-4811 (electronic).

Chen:2006:REP

[CMS06] Tzu-Yi Chen, Alvaro Monge, and Beth Simon. Relationship of early programming language to novice generated design. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):495–499, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Collberg:2007:ESJ

[CMS07] Christian Collberg, Ginger Myles, and Michael Stepp. An empirical study of Java bytecode programs. Software— Practice and Experience, 37 (6):581–641, May 2007. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chen:2003:DGV

[CN03a] B. Y. Chen and T. Nishita.

Development of 3D graphics and VRML libraries for Web3D platform by using Java. Systems and computers in Japan, 34(10):47– 55, 2003. CODEN SCJAEP. ISSN 0882-1666 (print), 1520-684X (electronic).

Chiba:2003:EUT

[CN03b] S. Chiba and M. Nishizawa. An easy-to-use toolkit for efficient Java bytecode translators. Lecture Notes in Computer Science, 2830:364–376, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:2000:PAS

[CNB00] Weidong Chen, Jeffery Naughton, and Philip A. Bernstein, editors. Proceedings of the 2000 ACM SIGMOD International Conference on Management of Data: May 16– 18, 2000, Dallas, Texas, volume 29(2) of SIGMOD Record (ACM Special Interest Group on Management of Data). ACM Press, New York, NY 10036, USA, 2000. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1 .A87.

Chen:2003:JSDa

[CO03a] Michael K. Chen and Kunle Olukotun. The Jrpm system for dynamically parallelizing Java programs. ACM SIGARCH Computer Architecture News, 31(2):434– 446, May 2003. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Chen:2003:JSDb

[CO03b] Michael K. Chen and Kunle Olukotun. The Jrpm system for dynamically parallelizing sequential Java programs. IEEE Micro, 23(6):26–35, November/December 2003. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:/

/csdl.computer.org/comp/
mags/mi/2003/06/m6026abs.
htm;   http://csdl.computer.
org/dl/mags/mi/2003/06/
m6026.htm;   http://csdl.
computer.org/dl/mags/mi/
2003/06/m6026.pdf.

Chawla:2004:GIF

[CO04] Anil Chawla and Alessandro Orso. A generic instrumentation framework for collecting dynamic information. ACM SIGSOFT Software Engineering Notes, 29(5):1–4, September 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Cavazos:2006:MSDa

[CO06] John Cavazos and Michael F. P. O’Boyle. Methodspecific dynamic compilation using logistic regression. ACM SIGPLAN Notices, 41(10):229–240, October 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Carroll:2007:IMA

[CO07] Bradley W. Carroll and Dale A. Ostlie. An Introduction to Modern Astrophysics. Pearson Addison-Wesley, San Francisco, CA, USA, second edition, 2007. ISBN 0-8053-0402-9, 0-321-44284-9 (paperback). ???? pp. LCCN QB461 .C35 2007.

                                        URL http://catdir.loc.
                                        gov/catdir/toc/ecip0613/
                                        2006015391.html;   http:
                                        //www.gbv.de/dms/ilmenau/
                                        toc/512485305.PDF.

Cochran:2002:NVR

[Coc02] Shannon Cochran. News and views: Researchers tackle P2P network; Java upstarts push JEFF; computer science tests canceled; the continuing voyage. Dr. Dobb’s Journal of Software Tools, 27(12):14, December 2002. CODEN DDJOEB. ISSN 1044-789X. URL http://

                                        www.ddj.com/documents/s=
                                        7718/ddj0212n/.

Coglio:2003:IOS

[Cog03] Alessandro Coglio. Improving the official specification of Java bytecode verification. Concurrency and Computation: Practice and Experience, 15(2):155–179, February 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Coglio:2004:SVT

[Cog04]  Alessandro Coglio.  Sim-

ple verification technique for complex Java bytecode subroutines. Concurrency and Computation: Practice and Experience, 16(7):

647–670, June 2004. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cohen:2002:JQH

[Coh02] Tal Cohen. Java Q&A: How do I correctly implement the equals() method? Dr. Dobb’s Journal of Software Tools, 27(5):83–84, 86, May 2002. CODEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/ftp/2002/
2002_05/jqa0502.txt.

Cohen:2004:TTT

[Coh04] F. Cohen. The testing toolbox: With these 10 tools, Java scalability, performance and functionality are no longer elusive. chockful of techniques, they enable software developers, QA technicians and IT managers to effectively proof programs. Software Development, 12(7): 36–43, 2004. CODEN ???? ISSN 1070-8588.

Collins:2001:DSJ

[Col01] William J. (William Joseph)

Collins. Data structures and the Java collections framework. McGraw-Hill, New York, NY, USA, 2001. ISBN 0-07-236964-7. xx + 716 pp. LCCN QA76.73.J38 C657 2001.

Coleman:2002:OAJ

[Col02] C. L. Coleman. Oracle angles for Java developers. E Business Advisor, 20(1):12– 13, 2002. CODEN ???? ISSN 1098-8912.

Cooper:2000:JDP

[Coo00] James W. Cooper. Java Design Patterns: a Tutorial. Addison-Wesley, Reading, MA, USA, 2000. ISBN 0-201-48539-7. xvii + 329 pp. LCCN QA76.73.J38 C658 2000.

Cooper:2001:JI

[Coo01] Brian Cooper. JavaScript: an introduction. Essential computers. Dorling Kindersley Pub., New York, NY, USA, 2001. ISBN 0-7894-8005-0 (paperback). 72 pp. LCCN QA76.73.J39 C66 2001. At head of title: Internet.

Cook:2002:REJ

[Coo02]  Jonathan J. Cook.  Re-

verse execution of Java bytecode. The Computer Journal, 45(6):608–619, ???? 2002. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

                                        www3.oup.co.uk/computer_
                                        journal/hdb/Volume_45/Issue_
                                        06/450608.sgm.abs.html;
                                         http://www3.oup.co.uk/
                                        computer_journal/hdb/Volume_
                                        45/Issue_06/pdf/450608.
                                        pdf.

Cook:2005:HCE

[Coo05] Robert P. Cook. Heuristic compression of an English word list. Software— Practice and Experience, 35 (6):577–581, May 2005. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Corbett:2000:USA

[Cor00]  James C. Corbett.  Us-

ing shape analysis to reduce finite-state models of concurrent Java programs. ACM Transactions on Software Engineering and Methodology, 9(1):51–93, January 2000. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL

http://www.acm.org/pubs/
articles/journals/tosem/
2000-9-1/p51-corbett/p51-
corbett.pdf;   http://www.
acm.org/pubs/citations/
journals/tosem/2000-9-1/
p51-corbett/.

Courtney:2001:FFR

[Cou01] Antony Courtney. Frappé:

Functional reactive programming in Java. Lecture Notes in Computer Science, 1990:29–??, 2001. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/1990/19900029.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/1990/19900029.
pdf.

Cowlishaw:2001:DAJ

[Cow01] Mike Cowlishaw. Decimal arithmetic for java.math: JSR 13 public review draft. Technical Report Version 1.33, IBM UK Laboratories, Hursley, UK, July 16, 2001. iii + 60 pp. URL

                                        http://www2.hursley.ibm.
                                        com/jsr13/jsr13spec.pdf.

Cox:2001:JQH

[Cox01a]  Brad Cox.  Java Q&A:

How do I handle multiline strings? Dr. Dobb’s Journal of Software Tools, 26(6):147–149, June 2001. CODEN DDJOEB. ISSN 1044-789X. URL http:

                                        //www.ddj.com/ftp/2001/
                                        2001_06/jqa0601.txt.

Cox:2001:WAJ

[Cox01b] Brad Cox. Web applications as Java servlets. Dr. Dobb’s Journal of Software Tools, 26(5):97–100, 102–104, May 2001. CODEN DDJOEB. ISSN 1044-789X. URL

                                        http://www.ddj.com/ftp/
                                        2001/2001_05/webapp.txt;
                                         http://www.ddj.com/ftp/
                                        2001/2001_05/webapp.zip.

Carrano:2001:DAP

[CP01] Frank M. Carrano and Janet J. Prichard. Data abstraction and problem solving with Java: walls and mirrors. Addison-Wesley, Reading, MA, USA, 2001. ISBN 0-201-70220-7. xxii + 806 pp. LCCN QA76.73.J38 C37 2001.

Carrano:2004:DAP

[CP04] Frank M. Carrano and Janet Prichard. Data Abstraction and Problem Solving with Java, Walls and Mirrors. Addison-Wesley, Reading, MA, USA, updated edition, 2004. ISBN 0-321-19717-8. xxii + 808 pp. LCCN QA76.73.J3 C37 2003. US$93.00. URL http:/

/www.aw-bc.com/catalog/
academic/product/0,4096,
0-321-19717-8,00.html.

Crane:2005:AA

[CPJ05] Dave Crane, Eric Pascarello, and Darren James. Ajax in Action. Manning Publications, Greenwich, CT, USA, 2005. ISBN 1-932394-61-3. 680 (est.) pp. LCCN ???? US$22.50 (electronic), US$44.95 (print + electronic). URL http:

//www.manning.com/books/
crane.

Chan:2005:UXJ

[CQ05] W. L. Chan and Z. Qu. Using XML/Java to enhance an online learning architecture for engineering education. International Journal of Engineering Education, 21(2):288– 296, 2005. CODEN IEEDEF. ISSN 0949-149X.

Chen:2009:UAD

[CQX^{+}09] Mingsong Chen, Xiaokang Qiu, Wei Xu, Linzhang Wang, Jianhua Zhao, and Xuandong Li. UML activity diagrambased automatic test case generation for Java programs. The Computer Journal, 52 (5):545–556, August 2009. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
52/5/545;   http://comjnl.
oxfordjournals.org/cgi/
reprint/52/5/545.

Cade:2002:SCE

[CR02a] Mark Cade and Simon Roberts. Sun Certified Enterprise Architect for J2EE Technology Study Guide. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2002. ISBN 0-13-044916-4. 220 (est.) pp. LCCN ???? US$39.99. URL http:/

/www.phptr.com/ptrbooks/
ptr_0130449164.html.

Comer:2002:TJB

[CR02b] James Comer and Robert Roggio. Teaching a Javabased CS1 course in an academically-diverse environment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):142–146, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Chen:2005:JMM

[CR05] F. Chen and G. Rosu. Java-MOP: a monitoring oriented programming environment for Java. Lecture Notes in Computer Science, 3440:546– 550, 2005. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chalin:2006:NNR

[CR06] Patrice Chalin and Frédéric Rioux. Non-null references by default in the Java modeling language. ACM SIG-SOFT Software Engineering Notes, 31(2):9:1–9:??, March 2006. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Chen:2007:MEG

[CR07] Feng Chen and Grigore Roşu.

MOP: an efficient and generic runtime verification framework. ACM SIGPLAN Notices, 42(10):569–588, October 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Craig:2006:VM

[Cra06]  Iain D. Craig.  Virtual

machines. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. ISBN 1-85233-969-1. xv + 269 pp. LCCN QA76.9.V5 C73 2006.

Chatterjee:2001:CPA

[CRL01] R. Chatterjee, B. G. Ryder, and W. A. Landi. Complexity of points-to analysis of Java in the presence of exceptions. IEEE Transactions on Software Engineering, 27(6):481–512, June 2001. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
926173.

Crowell:2001:CP

[Cro01] Talbott Crowell. C# and Perl: C# has a lot in common with many languages. Dr. Dobb’s Journal of Software Tools, 26(10):50, 52, 54, October 2001. CODEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/ftp/2001/
2001_10/csharpperl.txt.

Crockford:2008:JGP

[Cro08] Douglas Crockford. JavaScript: the good parts. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2008. ISBN 0-596-51774-2. xiii + 153 pp. LCCN QA76.73.J39 C763 2008.

Corsaro:2002:DPJ

[CS02] Angelo Corsaro and Douglas C. Schmidt. The design and performance of the jRate real-time Java implementation. Lecture Notes in Computer Science, 2519:900– 921, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/
service/series/0558/bibs/
2519/25190900.htm;   http:
//link.springer.de/link/
service/series/0558/papers/
2519/25190900.pdf.

Corsaro:2003:DPR

[CS03] Angelo Corsaro and Douglas C. Schmidt. The design and performance of real-time Java middleware. IEEE Transactions on Parallel and Distributed Systems, 14(11):1155–1167, November 2003. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL http://csdl.
computer.org/comp/trans/
td/2003/11/l1155abs.htm;
http://csdl.computer.org/
dl/trans/td/2003/11/l1155.
pdf.

Csallner:2004:JAR

[CS04] Christoph Csallner and Yannis Smaragdakis. JCrasher: an automatic robustness tester for Java. Software —Practice and Experience, 34(11):1025–1050, September 2004. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Chilimbi:2006:CCC

[CS06] Trishul M. Chilimbi and Ran Shaham. Cache-conscious coallocation of hot data streams. ACM SIGPLAN Notices, 41(6):252–262, June 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Clausen:2000:JBC

[CSCM00] Lars Ræder Clausen, Ulrik Pagh Schultz, Charles Consel, and Gilles Muller. Java bytecode compression for low-end embedded systems. ACM Transactions on Programming Languages and Systems, 22(3):471–489, 2000. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL

http://www.acm.org/pubs/
articles/journals/toplas/
2000-22-3/p471-clausen/
p471-clausen.pdf;   http:
//www.acm.org/pubs/citations/
journals/toplas/2000-22-
3/p471-clausen/.

Clark:2000:NBG

[CSFS00] David Clark, Keri Schreiner, Jennifer Ferrero, and Dale Strok. News: Blue Gene and the race toward petaflops capacity; embedded Java development moves ahead; putting teraflops to the test; Corba 3.0 on the way. IEEE Concurrency, 8(1):5– 10, January/March 2000. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http:

//dlib.computer.org/pd/

books/pd2000/pdf/p1005.
pdf.

Chung:2000:ECM

[CSK00] C.-M. Chung, P.-S. Shin, and S.-D. Kim. An effective configuration method for Java–Internet computing environment. Parallel Processing Letters, 10(1): 73–??, March 2000. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic). URL http:

//ejournals.wspc.com.sg/
ppl/10/1001/S0129626400000093.
html.

Chen:2002:TGC

[CSK^{+}02] G. Chen, R. Shetty, M. Kandemir, N. Vijaykrishnan, M. J. Irwin, and M. Wolczko. Tuning garbage collection for reducing memory system energy in an embedded Java environment. ACM Transactions on Embedded Computing Systems, 1(1):27–55, November 2002. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Christopher:2000:HPJ

[CT00] Thomas Christopher and George Thiruvathukal. High Performance Java Platform Computing: Multithreaded and Networked Programming. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2000. ISBN 0-13-016164-0. xxii + 409 pp. LCCN QA76.73.J38 C488 2000. US$49.99.

                                        URL http://www.sun.com/
                                        books/catalog/christopher/

.

Chen:2003:EJV

[CT03] D. Chen and V. Talwar. Extending Java Virtual Machines for networked embedded devices. Research Disclosure, 475:475/076, 2003. CO-DEN ???? ISSN 0374-4353.

Chatley:2005:KLP

[CT05] Robert Chatley and Thomas Timbul. KenyaEclipse: learning to program in eclipse. ACM SIGSOFT Software Engineering Notes, 30(5):245– 248, September 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Chevalley:2003:MAT

[CTF03] P. Chevalley and P. Thévenod-Fosse. A mutation analysis tool for Java programs. International Journal on Software Tools for Technology Transfer (STTT), 5(1):90– 103, November 2003. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Collins:2003:RFL

[CTLW03] William Collins, Josh Tenenberg, Raymond Lister, and Suzanne Westbrook. The role for framework libraries in CS2. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):403–404, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Culwin:2000:LWB

[Cul00] Fintan Culwin. Lecturelets:

web based Java enabled lectures. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):5–8, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Curioso:2007:AP

[Cur07] Andrew G. Curioso. Ajax with PHP 5. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2007. ISBN 0-596-51403-4. LCCN TK5105.8885.A52 C87 2007eb; TK5105.8885.A52. URL

http://www.oreilly.com/
catalog/9780596514037.

Caromel:2001:RMC

[CV01] Denis Caromel and Julien Vayssière. Reflections on MOPs, components, and Java security. Lecture Notes in Computer Science, 2072:256– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2072/20720256.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2072/20720256.
pdf.

Caromel:2003:SFR

[CV03] Denis Caromel and Julien Vayssière. A security framework for reflective Java applications. Software—Practice and Experience, 33(9): 821–846, July 25, 2003. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cimadamore:2008:RJW

[CV08] Maurizio Cimadamore and Mirko Viroli. On the reification of Java wildcards. Science of Computer Programming, 73(2–3):59–75, October 1, 2008. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Chang:2000:JJI

[CvE00] Chi-Chao Chang and Thorsten von Eicken. Javia: A Java interface to the virtual interface architecture. Concurrency: Practice and Experience, 12 (7):573–593, May 2000. CO-DEN CPEXEI. ISSN 1040-

                                        3108. URL http://www3.
                                        interscience.wiley.com/
                                        cgi-bin/abstract/72516214/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=72516214&PLACEBO=IE.
                                        pdf.

Carey:2003:NIF

[CVW03] Robert W. Carey, Paul J. Van Arsdall, and John P. Woodruff. The National Ignition Facility: early operational experience with a large Ada control system. ACM SIGADA Ada Letters, 23(1):11, March 2003. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Cai:2003:THI

[CW03a] H. Cai and A. J. Wellings. Towards a high integrity realtime Java virtual machine. Lecture Notes in Computer Science, 2889:319–334, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:2003:RPJ

[CW03b] Q. Chen and P. Y. Woo. Research on and pure Java realization of a Web-based mobile robot system. Proceedings of the American Control Conference, 1(??):615–620, 2003. CODEN PRACEO. ISSN 0743-1619.

Cai:2004:SMC

[CW04a] H. Cai and A. Wellings. Supporting mixed criticality applications in a Ravenscar– Java environment. Lecture Notes in Computer Science, 3292:278–291, 2004. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:2004:EEI

[CW04b] J. Chen and K. Wang. Experiment on embedding interception service into Java RMI. Lecture Notes in Computer Science, 2952:48–61, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Campione:2001:JTS

[CWH01] Mary Campione, Kathy Walrath, and Alison Huml. The Java Tutorial: a Short Course on the Basics. Addison-Wesley, Reading, MA, USA, third edition, 2001. ISBN 0-201-70393-9. xi + 580 pp. LCCN QA76.73.J38 C365 2001. US$44.99.

URL http://java.sun.com/
docs/books/tutorial/books/
3e/index.html.  Includes

CD-ROM.

Chakravarti:2003:ISM

[CWHB03] A. Chakravarti, X. Wang, J. Hallstrom, and G. Baumgartner. Implementation of strong mobility for multithreaded agents in Java. In Proceedings of the International Conference on Parallel Processing, pages 321–332. ????, ????, 2003. CODEN ???? ISSN 0190-3918.

Chalk:2004:SGS

[CWS04] A. M. Chalk, M. Wennerberg, and E. L. Sonnhammer. Sfixem-graphical sequence feature display in Java. Bioinformatics, 20(15): 2488–2490, 2004. CODEN ???? ISSN 1367-4803 (print), 1367-4811 (electronic).

Can:2003:FFP

[CWWS03] T. Can, Y. Wang, Y. F.

Wang, and J. Su. FPV: fast protein visualization using Java 3D. Bioinformatics, 19(8):913–922, 2003. CODEN ???? ISSN 1367-4803 (print), 1367-4811 (electronic).

Chiao:2001:MEM

[CWY01] Hsin-Ta Chiao, Chi-Houng Wu, and Shyan-Ming Yuan. A more expressive monitor for concurrent Java programming. Lecture Notes in Computer Science, 1900:1053–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/1900/19001053.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/1900/19001053.
pdf.

Chen:2004:STD

[CWZ04] W. Chen, D. Wang, and W. Zheng. A single thread discrete event simulation toolkit for Java: STSimJ. Lecture Notes in Computer Science, 3038: 1131–1137, 2004. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:2001:SCJ

[CX01a] Zhenqiang Chen and Baowen Xu. Slicing concurrent Java programs. ACM SIGPLAN Notices, 36(4):41–47, April 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chen:2001:SOO

[CX01b] Zhenqiang Chen and Baowen Xu. Slicing object-oriented Java programs. ACM SIG-PLAN Notices, 36(4):33– 40, April 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chiao:2001:ETS

[CY01a] Hsin-Ta Chiao and Shyan-Ming Yuan. An enhanced thread synchronization mechanism for Java. Software— Practice and Experience, 31 (7):667–695, June 2001. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (elec-

tronic). URL http://www3.
interscience.wiley.com/
cgi-bin/abstract/78003134/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=78003134&PLACEBO=IE.

pdf. See correction [CY01b].

Chiao:2001:RIM

[CY01b] Hsin-Ta Chiao and Shyan-Ming Yuan. The real inventor of the monitor concept: a short note to correct the SP&E paper entitled “An enhanced thread synchronization mechanism for Java”. Software—Practice and Experience, 31(14):1393, November 25, 2001. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (elec-

tronic). URL http://www3.
interscience.wiley.com/
cgi-bin/abstract/85515675/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=85515675&PLACEBO=IE.

pdf. See [CY01a].

Chan:2002:AGF

[CY02] J. T. Chan and W. Yang.

An attribute-grammar framework for specifying the accessibility in Java programs. Computer Languages, 28(2): 203–235, 2002. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Chen:2003:JMA

[CY03] R. Y. Chen and B. Yeager.

Java mobile agents on project JXTA peer-to-peer platform (STPTP01). Proceedings of the Annual Hawaii International Conference on System Sciences, CONF36:282, 2003. CODEN ???? ISSN 1060-3425.

Chan:2004:AOT

[CY04] Jien-Tsai T. Chan and Wuu Yang. Advanced obfuscation techniques for Java bytecode. The Journal of systems and software, 71(1–2):1–10, April 2004. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chan:2004:TJ

[CYH04] Jien-Tsai T. Chan, Wuu Yang, and Jing-Wei W. Huang. Traps in Java. The Journal of systems and software, 72(1):33–47, June 2004. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Chaudhri:2001:SOD

[CZ01] Akmal B. Chaudhri and Roberto Zicari. Succeeding with object databases: a practical look at today’s implementations with Java and XML. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 2001. ISBN 0-471-38384-8 (cloth). xxi + 442 pp. LCCN QA76.9.D3 C3848 2001.

Chen:2002:ILD

[CZ02] J. Chen and H. Zhao. Implement of linked data structure in Java. Journal: Beijing University of Chemical Technology, 29(5):72–74, 2002. CODEN ???? ISSN 1007-2640.

Czajkowski:2000:AIJ

[Cza00] Grzegorz Czajkowski. Application isolation in the Java^{T M} Virtual Machine. ACM SIGPLAN Notices, 35 (10):354–366, October 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/

proceedings/oops/353171/
p354-czajkowski/.

Daconta:2000:JPT

[D^{+}00] Michael C. Daconta et al.

Java pitfalls: time-saving solutions and workarounds to improve programs. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 2000. ISBN 0-471-36174-7 (paperback). xiv + 322 pp. LCCN QA76.73.J38 P3785 2000.

Dudney:2004:MJF

[D^{+}04] Bill Dudney et al., editors. Mastering JavaServer Faces. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 2004. ISBN 0-471-46207-1 (paperback). xxiii + 456 pp. LCCN TK5105.8885.J38 M37 2004. URL http://

www.loc.gov/catdir/bios/
wiley047/2004274458.html;
 http://www.loc.gov/catdir/
description/wiley042/2004274458.
html;   http://www.loc.
gov/catdir/toc/wiley041/
2004274458.html.

Doyle:2002:MEJ

[DA02] Patrick Doyle and Tarek S.

Abdelrahman. A modular and extensible JVM infrastructure. In USENIX Association [USE02], page ?? ISBN 1-931971-01-3. LCCN QA76.73.J38 J42 2002. URL

http://www.usenix.org/publications/
library/proceedings/javavm02/
doyle.html.

Doyle:2004:JPT

[DA04] T. C. Doyle and C. A. April. Java peace treaty: Nothing seems more surreal than the Sun–Microsoft pact, even now. Varbusiness, 20(10):22, 2004. CODEN ???? ISSN 0894-5802.

Dimpsey:2000:JSP

[DAK00] R. Dimpsey, R. Arora, and K. Kuiper. Java server performance: a case study of building efficient, scalable Jvms. IBM Systems Journal, 39(1): 151–174, ???? 2000. CO-DEN IBMSA7. ISSN 0018-

                                        8670. URL http://www.
                                        almaden.ibm.com/journal/
                                        sj/391/dimpsey.html.

Darcy:2001:BLH

[Dar01a] Joseph D. Darcy. Borneo language homepage. World Wide Web site., 2001. URL http:/

                                        /www.jddarcy.org/Borneo.

Darcy:2001:WEU

[Dar01b] Joseph D. Darcy. What everybody using the Java^{T M} programming language should know about floating-point arithmetic. In JavaOne 2001, June 7, 2001, page ?? ????, ????, 2001. ISBN ???? LCCN

                                        ???? URL http://java.
                                        sun.com/people/darcy/JavaOne/
                                        2001/1789darcy.pdf.  51

slides.

Darwin:2001:JCS

[Dar01c] Ian Darwin. Java Cookbook: Solutions and Examples for Java Developers. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, June 2001. ISBN 0-596-00170-3 (paperback). xxix + 850 pp. LCCN QA76.73.J38 D348 2001. US$44.95. URL

http://www.oreilly.com/
catalog/9780596001704;
http://www.oreilly.com/
catalog/javacook.

Darwin:2003:JCS

[Dar03] Ian F. Darwin. Java Cookbook: Solutions and Examples for Java Developers. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2003. ISBN 0-596-00701-9. xxiv + 829 pp. LCCN QA76.73.J38 D348 2004. US$49.95, CAN$72.95, UK£35.50. URL

http://examples.oreilly.
com/javacook2/;   http:/
/www.oreilly.com/catalog/
0596007019/.

Darwin:2004:JC

[Dar04] Ian F. Darwin. Java cookbook. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, second edition, 2004. ISBN 0-596-00701-9. xxiv + 829 pp. LCCN QA76.73.J38 D348 2004. URL http://

www.oreilly.com/catalog/
9780596007010.

Darwin:2007:CJP

[Dar07] Ian F. Darwin. Checking Java programs. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2007. ISBN 0-596-51023-3. 54 pp. LCCN QA76.73.J38eb; QA76.73.J38. URL http://

                                        www.oreilly.com/catalog/
                                        9780596510237.

Dautelle:2001:JDJ

[Dau01] Jean-Marie Dautelle. J.A.D.E.: The Java Addition to the Default Environment. Dr. Dobb’s Journal of Software Tools, 26(2):52, 54, 56, February 2001. CODEN DDJOEB. ISSN 1044-789X. URL http:

                                        //www.ddj.com/ftp/2001/
                                        2001_02/jade.txt.

Davison:2005:KGP

[Dav05]  Andrew Davison.  Killer

game programming in Java. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00730-2. xxiv + 969 pp. LCCN QA76.76.C672 D38 22005.

Dillenberger:2000:BJV

[DBC^{+}00] D. Dillenberger, R. Bordawekar, C. W. Clark, D. Durand, D. Emmes, O. Gohda, S. Howard, M. F. Oliver, F. Samuel, and R. W. St. John. Building a Java virtual machine for server applications: The JVM on OS/390. IBM Systems Journal, 39(1): 194–210, ???? 2000. CO-DEN IBMSA7. ISSN 0018-

8670. URL http://www.
almaden.ibm.com/journal/
sj/391/dillenberger.html.

deOliveira:2004:MEE

[dBdd04] A. A. de Oliveira, T. H.

Braga, M. de, Almeida Maia, and R. da Silva Bigonha. MetaJ: An extensible environment for metaprogramming in Java. J.UCS: Journal of Universal Computer Science, 10(7):872–??, July 28, 2004. CODEN ???? ISSN 0948-6968. URL http:/

/www.jucs.org/jucs_10_7/
metaj_an_extensible_environment.

Dunkel:2004:CJP

[DBH04] Jürgen Dunkel, Ralf Bruns, and Andreas Holitschke. Comparison of JavaServer Pages and XSLT: a software engineering perspective. Software—Practice and Experience, 34(1):1–13, January 2004. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Deitsch:2001:JI

[DC01] Andrew Deitsch and David Czarnecki. Java internationalization. Java series. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2001. ISBN 0-596-00019-7. xvi + 444 pp. LCCN QA76.73.J38 D45 2001.

Depradine:2003:PCD

[DC03a] C. Depradine and P. Chaudhuri. P3: a code and design conventions preprocessor for Java. Software—Practice and Experience, 33(1):61– 76, 2003. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Deters:2003:ADS

[DC03b] Morgan Deters and Ron K. Cytron. Automated discovery of scoped memory regions for real-time Java. ACM SIG-PLAN Notices, 38(2s):132– 142, February 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dann:2009:EAC

[DC09] Wanda Dann and Stephen Cooper. Education Alice 3: concrete to abstract. Communications of the ACM, 52 (8):27–29, August 2009. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Doyle:2004:DIM

[DCA04] Patrick Doyle, Carlos Cavanna, and Tarek S. Abdelrahman. The design and implementation of a modular and extensible Java Virtual Machine. Software—Practice and Experience, 34(3): 287–313, March 2004. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

deBeer:2002:MIR

[dCG^{+}02] R. deBeer, A. Coron, D. Graveron-Demilly, R. Lethmate, S. Nastase, D. vanOrmondt, and F. T. Wajer. MR image reconstruction algorithms for sparse k-space data: a Javabased integration. Magma, 15 (1–3):18–26, 2002. CODEN MAGMEY. ISSN 0968-5243.

deDinechin:2001:JQW

[dD01a] Christophe de Dinechin. Java Q&A: What is Moka? Dr. Dobb’s Journal of Software Tools, 26(7):125–126, 128, 130–131, July 2001. CO-DEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/ftp/2001/
2001_07/jqa0701.txt.

Bois:2001:DEF

[Dd01b] André Rauber Du Bois and Antônio Carlos da Rocha Costa. Distributed execution of functional programs using the JVM. Lecture Notes in Computer Science, 2178:570– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2178/21780570.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2178/21780570.
pdf.

Deitel:2002:CJT

[DD02a] Harvey M. Deitel and Paul J.

Deitel. The Complete Java 2 Training Course. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2002. ISBN 0-13-064931-7. ???? pp. LCCN ???? US$109.99. URL http:/

/www.phptr.com/ptrbooks/
ptr_0130649317.html.

Deitel:2002:JHP

[DD02b] Harvey M. Deitel and Paul J. Deitel. Java: how to program. How to program series. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, fourth edition, 2002. ISBN 0-13-034151-7. lii + 1546 + 8 pp. LCCN QA76.73.J38 D45 2002. CD-ROM contains Java TM 2 SDK, Standard Edition, 1.3.1, Java Media Framework API 2.1.1, Forte for Java, Release 2.0, Community Edition and Java Plug-in HTML Converter 1.3.

Dellwig:2002:J

[DD02c] Elmar Dellwig and Ingo Dellwig. JavaScript. Addison-Wesley nitty gritty programming series. Addison-Wes-ley, Reading, MA, USA, 2002. ISBN 0-201-75875-X (paperback). xiii + 289 pp. LCCN QA76.73.J39 D45 2002. Translation of: JavaScript 1.3 Nitty Gritty.

Deitel:2003:JHP

[DD03] Harvey M. Deitel and Paul J. Deitel. Java: how to program. How to program series. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, fifth edition, 2003. ISBN 0-13-101621-0, 0-13-101623-7 (CD-ROM), 0-13-120236-7. li + 1447 pp. LCCN QA76.73.J38 D45 2003.

Deitel:2007:JHP

[DD07] Paul J. Deitel and Paul J.

Deitel. Java: how to program. How to program series. Pearson Prentice Hall, Upper Saddle River, NJ, USA, seventh edition, 2007. ISBN 0-13-222220-5. xliv + 1596 pp. LCCN QA76.73.J38 D45 2007.

DeMeuter:2004:OOL

[DDDM04] W. DeMeuter, S. Ducasse, T. DHondt, and O. L. Madsen. Object-oriented language engineering for the post-Java era. Lecture Notes in Computer Science, 3013:143–153, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Debbabi:2003:SSC

[DDF^{+}03] M. Debbabi, J. Desharnais, M. Fourati, E. Menif, F. Painchaud, and N. Tawbi. Secure self-certified code for Java. Lecture Notes in Computer Science, 2629:133–151, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dufour:2003:DMJ

[DDHV03] Bruno Dufour, Karel Driesen, Laurie Hendren, and Clark Verbrugge. Dynamic metrics for Java. ACM SIGPLAN Notices, 38(11):149–168, November 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Deitel:2002:AJP

[DDS02] Harvey M. Deitel, Paul J. Deitel, and S. Santry. Advanced Java 2 platform: how to program. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2002. ISBN 0-13-089560-1. xxxviii + 1811 + EULA-13 pp. LCCN QA76.73.J38 D445 2001. CD-ROM contains Java TM 2 Software Development Kit Standard Edition Version 1.3.1 for Windows and Linux (Intel x86); Forte for Java, Release 2.0, Community Edition for All Platforms; BEA WebLodge Server TM, Version 6.0 (Windows/Linux) with Service Pack 1 or 2, 30 Day Trial.

deCarmo:2004:JOA

[deC04] L. deCarmo. Java & The OpenCable Application Platform. Dr. Dobb’s Journal of Software Tools, 29(7):34– 41, 2004. CODEN DDJOEB. ISSN 1044-789X.

Deitel:2008:JFI

[Dei08] Harvey M. Deitel. Java fundamentals. I and II: Video LiveLessons. LiveLessons. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2008. ISBN 0-13-713129-1. LCCN QA76.73.J38. URL http://

proquest.safaribooksonline.
com/9780137131297.  1

streaming video file (16h36m25s).

Drossopoulou:2001:FTJ

[DEJ^{+}01] Sophia Drossopoulou, Susan Eisenbach, Bart Jacobs, Gary T. Leavens, Peter Müller, and Arnd Poetzsch-Heffter. Formal techniques for Java programs. Lecture Notes in Computer Science, 1964:41– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/1964/19640041.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/1964/19640041.
pdf.

Dekel:2000:SIJ

[Dek00] Eliezer Dekel. Special issue on Java on clusters. Journal of Parallel and Distributed Computing, 60(10):1155–1158, October 2000. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jpdc.
2000.1648;   http://www.
idealibrary.com/links/doi/
10.1006/jpdc.2000.1648/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.
2000.1648/ref.

Debbabi:2003:MCA

[DEK^{+}03] M. Debbabi, M. Erhioui, L. Ketari, N. Tawbi, H. Yahyaoui, and S. Zhioua. Method call acceleration in embedded Java virtual machines. Lecture Notes in Computer Science, 2659:750–759, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dekker:2006:LFP

[Dek06] Anthony H. Dekker. Lazy functional programming in Java. ACM SIGPLAN Notices, 41(3):30–39, March 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Drossopoulou:2002:FTJ

[DEL^{+}02] Sophia Drossopoulou, Susan Eisenbach, Gary T. Leavens, Arnd Poetzsch-Heffter, and Erik Poll. Formal techniques for Java-like programs. Lecture Notes in Computer Science, 2548:203–210, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/
                                        service/series/0558/bibs/
                                        2548/25480203.htm;   http:
                                        //link.springer.de/link/
                                        service/series/0558/papers/
                                        2548/25480203.pdf.

DePasquale:2003:UJU

[DeP03a] C. J. DePasquale. Using the JVMPI to understand the behavior of Java classes during the development process. Cmg, 2(??):821–832, 2003. CODEN ????

Depradine:2003:ESE

[Dep03b] C. Depradine. Expert system for extracting syntactic information from Java code. Expert Systems with Applications, 25(2):187–198, 2003. CODEN ESAPEH. ISSN 0957-4174.

Deshpande:2001:CDA

[Des01] Salil Deshpande. CORBA and Distributed Applications Including Java. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2001. ISBN 0-13-349960-X. ???? pp. LCCN ???? US$45.

Deters:2001:SMA

[Det01]  Ralph Deters.  A scal-

able multi-agent system. In ACM [ACM01b], page ?? ISBN 1-58113-359-6. LCCN QA76.9.O35 A26 2001. URL

http://www.philippsen.com/
JGI2001/camerareadyabstracts/
19.html.

Deugo:2000:MJG

[Deu00] Dwight Deugo, editor. More Java Gems. Cambridge University Press, Cambridge, UK, 2000. ISBN 0-521-77477-

2. xix + 484 pp. LCCN QA76 .73 J38 M657 2000. US$40.

Dahlen:2003:AJP

[DF03] Torbjörn Dahlén and Thorbiörn Fritzon. Advanced J2EE Platform Development: Applying Integration Tier Patterns. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2003. ISBN 0-13-044912-1. xv + 195 pp. LCCN QA76.76.J38 D34 2004. US$39.99.

Du:2003:CSE

[DFA03] Wei Du, Renato Ferreira, and Gagan Agrawal. Compiler support for exploiting coarse-grained pipelined parallelism. In ACM [ACM03b], page ?? ISBN 1-58113-695-

1. LCCN ???? URL http:

//www.sc-conference.org/
sc2003/inter_cal/inter_
cal_detail.php?eventid=
10692#2;   http://www.sc-
conference.org/sc2003/paperpdfs/
pap133.pdf.

Duarte:2000:BJA

[DFL00] Carlos H. C. Duarte, Martin Fogarty, and Robert C. Larrabee. Bookshelf: Java application frameworks use case driven object: Modeling with UML: a practical approach: Chaos and complexity in software, challenging the industry and the new science. IEEE Software, 17(5): 120–123, September/October 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:

//dlib.computer.org/so/
books/so2000/pdf/s5120.
pdf.

diFlora:2004:IPL

[dFR04] C. diFlora, M. Ficco, and S. Russo. Indoor positioning for location-aware applications on Java-based mobile devices. Lecture Notes in Computer Science, 3292:383– 393, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DiStefano:2003:CRE

[DFT03] A. DiStefano, M. Fargetta, and E. Tramontana. Computational reflection for embedded Java systems. Lecture Notes in Computer Science, 2889:437–450, 2003. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Deng:2004:TWD

[DFW04] Yuetang Deng, Phyllis Frankl, and Jiong Wang. Testing web database applications. ACM SIGSOFT Software Engineering Notes, 29 (5):1–10, September 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dutheil:2002:BJE

[DG02] J. Dutheil and N. Galtier.

BAOBAB: a Java editor for large phylogenetic trees. Bioinformatics, 18(6):892– 893, June 2002. CODEN ???? ISSN 1367-4803 (print), 1367-4811 (electronic).

Damiani:2008:TSS

[DGGD08] Ferruccio Damiani, Elena Giachino, Paola Giannini, and Sophia Drossopoulou. A type safe state abstraction for coordination in Java-like languages. Acta Informatica, 45(7–8):479–536, December 2008. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0001-5903&volume=45&
issue=7&spage=479.

Domani:2003:TLH

[DGK^{+}03] Tamar Domani, Gal Goldshtein, Elliot K. Kolodner, Ethan Lewis, Erez Petrank, and Dafna Sheinwald. Thread-local heaps for Java. ACM SIGPLAN Notices, 38(2s):183–194, February 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Debbabi:2006:SDC

[DGMY06] Mourad Debbabi, Abdelouahed Gherbi, Azzam Mourad, and Hamdi Yahyaoui. A selective dynamic compiler for embedded Java virtual machines targeting ARM processors. Science of Computer Programming, 59(1–2): 38–63, January 2006. CO-DEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

deBeer:2004:DCS

[dGNv04] R. deBeer, D. Graveron-Demilly, S. Nastase, and D. vanOrmondt. A distributed computing system for magnetic resonance imaging: Java-based processing and binding of XML. Computer Methods and Programs in Biomedicine, 73 (3):221–231, 2004. CO-DEN CMPBEK. ISSN 0169-2607 (print), 1872-7565 (electronic).

Dwyer:2000:APL

[DH00] Matthew B. Dwyer and John Hatcliff. Adapting programming languages technologies for finite-state verification. ACM SIGSOFT Software Engineering Notes, 25(1):46– 49, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Daly:2004:ALS

[DH04a] C. Daly and J. M. Horgan.

An automated learning system for Java programming. IEEE transactions on education, 47(1):10–17, 2004. CO-DEN IEEDAB. ISSN 0018-9359.

Dujmovic:2004:VJW

[DH04b] J. Dujmovic and C. Herder.

Visualization of Java workloads using ternary diagrams. Software Engineering Notes, 29(1):261–265, 2004. CODEN ???? ISSN 0163-5948.

dAmorim:2005:EBR

[dH05] Marcelo d’Amorim and Klaus Havelund. Event-based runtime verification of Java programs. ACM SIGSOFT Software Engineering Notes, 30 (4):1–7, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dagenais:2008:ESA

[DH08] Barthélémy Dagenais and Laurie Hendren. Enabling static analysis for partial Java programs. ACM SIGPLAN Notices, 43(10):313–328, September 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dicken:2000:DLO

[DHMT00] Hans Dicken, Gunther Hipper, and Peter Müßig-Trapp. Datenbanken unter Linux: Oracle 8i, MySQL, Adabas, Informix, Sybase, DB2, PostgreSQL, MiniSQL, Empress; [Tipps zur optimalen Installation und Konfiguration; Backup, Recovery, Ausfallsicherheit; mit PHP und Java ins Web]. MITP-Verlag, Bonn, Germany, 2000. ISBN 3-8266-0555-1. 516 (est.) pp. LCCN ????

Daly:2001:PID

[DHPW01] Charles Daly, Jane Horgan, James Power, and John Waldron. Platform independent dynamic Java Virtual Machine Analysis: the Java Grande Forum benchmark suite. In ACM [ACM01b], pages 106–

115. ISBN 1-58113-359-

6. LCCN QA76.9.O35 A26

2001. URL http://www.
philippsen.com/JGI2001/
camerareadyabstracts/16.
html;   http://www.philippsen.
com/JGI2001/finalpapers/
18500106.ps.

Duncan:2001:LPD

[DHR^{+}01] Ray Duncan, Duncan Harris, Douglas Reilly, Craig Rodrigues, Michael Birken, and Paul S. Person. Letters: Plugin desupport; threading and the .Net framework; CORBA interoperability; game over for Java; totally wired. Dr. Dobb’s Journal of Software Tools, 26(11):10, 12, November 2001. CODEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/.

Drysdale:2005:YRC

[DHRH05] S. Drysdale, J. Hromcik, D. Reed, and R. Hahne. The year in review: Changes and lessons learned in the design and implementation of the AP CS Exam in Java. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37 (1):323–324, 2005. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Daley:2002:FTD

[DHS02] Nigel Daley, Daniel Hoffman, and Paul Strooper. A framework for table driven testing of Java classes. Software— Practice and Experience, 32 (5):465–493, April 25, 2002. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

                                        www3.interscience.wiley.
                                        com/cgi-bin/abstract/93513336/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=93513336&PLACEBO=IE.
                                        pdf.

Drysdale:2003:JMJ

[DHWH03] S. Drysdale, J. Hromcik, M. A. Weiss, and R. Hahne. Java in the morning… Java in the evening… Java in 2004. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):271–272, 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dibble:2002:RTJ

[Dib02]  Peter C. Dibble.  Real-

Time Java Platform Programming. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2002. ISBN 0-13-028261-8. 352 (est.) pp. LCCN QA76.73.J38 D53 2002. US$49.99. URL http:/

                                        /www.phptr.com/ptrbooks/
                                        ptr_0130282618.html;   http:
                                        //www.sun.com/books/catalog/
                                        dibble.

Dice:2001:IFJ

[Dic01] David Dice. Implementing fast Java^{T M} monitors with relaxed-locks. In USENIX Association [USE01c], page ?? ISBN 1-880446-11-1. LCCN QA76.73.J38 J42 2001. URL

http://www.usenix.org/publications/
library/proceedings/jvm01/

dice.html. Sponsored by the USENIX Association.

Dieckmann:2000:SOD

[Die00] Sylvia Dieckmann. A study of the object demographics of large Java applications. Thesis (M.S.), University of California, Santa Barbara, Santa Barbara, CA, USA, 2000.

Diehl:2001:DVW

[Die01] Stephan Diehl. Distributed virtual worlds: foundations and implementation techniques using VRML, Java, and CORBA. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 3-540-67624-4. xii + 166 pp. LCCN QA76.9.D5 D427 2001.

Dill:2000:MCJ

[Dil00] David Dill. Model checking Java programs (abstract only). ACM SIGSOFT Software Engineering Notes, 25 (5):179, September 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

DiMaggio:2004:TJS

[DiM04] L. DiMaggio. Testing Java servlets. Dr. Dobb’s Journal of Software Tools, 29(8):26– 31, 2004. CODEN DDJOEB. ISSN 1044-789X.

Denney:2000:CJC

[DJ00] Ewen Denney and Thomas Jensen. Correctness of Java card method lookup via logical relations. Lecture Notes in Computer Science, 1782:104– ??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/1782/17820104.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/1782/17820104.
pdf.

Dysvik:2001:JEE

[DJ01] B. Dysvik and I. Jonassen. J-Express: exploring gene expression data using Java. Bioinformatics, 17(4):369– 370, April 2001. CODEN ???? ISSN 1367-4803 (print), 1367-4811 (electronic).

Denney:2002:CJC

[DJ02] Ewen Denney and Thomas Jensen. Correctness of Java card method lookup via logical relations. Theoretical Computer Science, 283(2): 305–331, June 2002. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Donsez:2001:TMA

[DJLT01] Didier Donsez, Sébastien Jean, Sylvain Lecomte, and Olivier Thomas. Turning multi-applications smart cards services available from anywhere at anytime: a SOAP /MOM approach in the context of Java cards. Lecture Notes in Computer Science, 2140:83–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2140/21400083.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2140/21400083.
pdf.

Pauw:2002:VEJ

[DJM^{+}02] Wim De Pauw, Erik Jensen, Nick Mitchell, Gary Sevitsky, John Vlissides, and Jeaha Yang. Visualizing the execution of Java programs. Lecture Notes in Computer Science, 2269:151–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2269/22690151.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2269/22690151.
pdf.

Djordjevic:2008:JPM

[Djo08]  Mirela Djordjevic.  Java

projects motivated by student interests. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):321, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Djordjevic:2009:PAC

[Djo09]  Mirela Djordjević.  Pro-

gressive assignment in CS1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):346, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Delsart:2002:JLM

[DJP02] Bertrand Delsart, Vania Joloboff, and Eric Paire. JCOD: a lightweight modular compilation technology for embedded Java. Lecture Notes in Computer Science, 2491:197–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/
                                        service/series/0558/bibs/
                                        2491/24910197.htm;   http:
                                        //link.springer.de/link/
                                        service/series/0558/papers/
                                        2491/24910197.pdf.

Drofenik:2002:IPE

[DK02] U. Drofenik and J. W. Kolar. Interactive power electronics seminar (iPES) — A Web-based introductory power electronics course employing Java-applets. IEEE Power Electronics Specialists Conference, CONF33(VOL 2):443–448, 2002. CODEN ???? ISSN 0275-9306.

DeSouza:2003:JPM

[DK03] J. DeSouza and L. V. Kale.

Jade: a parallel messagedriven Java. Lecture Notes in Computer Science, 2659:760– 772, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Domani:2001:IFG

[DKL^{+}01] Tamar Domani, Elliot K.

Kolodner, Ethan Lewis, Eliot E. Salant, Katherine Barabash, Itai Lahan, Yossi Levanoni, Erez Petrank, and Igor Yanorer. Implementing an on-the-fly garbage collector for Java. ACM SIG-PLAN Notices, 36(1):155– 166, January 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Domani:2000:GFG

[DKP00] Tamar Domani, Elliot K.

Kolodner, and Erez Petrank. A generational on-the-fly garbage collector for Java. ACM SIG-PLAN Notices, 35(5):274– 284, May 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL

                                        http://www.acm.org/pubs/
                                        articles/proceedings/pldi/
                                        349299/p274-domani/p274-
                                        domani.pdf;   http://www.
                                        acm.org/pubs/citations/
                                        proceedings/pldi/349299/
                                        p274-domani/.

Donovan:2004:CJP

[DKTE04] A. Donovan, A. Kiezun, M. S. Tschantz, and M. D. Ernst. Converting Java programs to use generic libraries. ACM SIGPLAN Notices, 39 (10):15–34, 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Doherty:2000:JU

[DL00] Donald Doherty and Rick Leinecker. JavaBeans Unleashed. Howard W. Sams, Indianapolis, IN 46268, USA, 2000. ISBN 0-672-31424-X. xv + 676 pp. LCCN A76.73.J38 D627 2000. US$50.

Deng:2002:JUJ

[DL02] Q. Deng and X. Lu. JOGR — utilizing Java object group relationship to support multimodel parallel programming. Acta Electronica Sinica, 30 (11):1718–1721, 2002. CO-DEN ???? ISSN 0372-2112.

deLeeuw:2005:BRC

[dL05] Jan de Leeuw. Book review: Correspondence Analysis and Data Coding with Java and R. Journal of Statistical Software, 14(BR-5):1– 2, September 2005. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v14/b05.

Drossopoulou:2006:FMD

[DLE06] Sophia Drossopoulou, Giovanni Lagorio, and Susan Eisenbach. A flexible model for dynamic linking in Java and C#. Theoretical Computer Science, 368(1–2):1–29, December 5, 2006. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Deng:2003:RCJ

[DLL03] G. Deng, W. Li, and Z. Li.

Research on CORBA/Javabased Web database application model. Journal — Dalian University of Technology, 43 (1):124–128, 2003. CODEN ???? ISSN 1000-8608.

Dutchyn:2001:MDJ

[DLS^{+}01] Christopher Dutchyn, Paul Lu, Duane Szafron, Steven Bromling, and Wade Holst. Multi-dispatch in the Java Virtual Machine: Design and implementation. In USENIX [USE01a], page ?? ISBN 1-880446-12-X. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/coots01/
dutchyn.html.

deMelo:2004:CJF

[dM04] A. C. deMelo and B. M.

Moutinho. On the composition of Java frameworks control-flows. Information and Software Technology, 46 (5):301–307, 2004. CO-DEN ISOTE7. ISSN 0950-5849 (print), 1873-6025 (electronic).

Drechsler:2007:YSL

[DM07] R. L. Drechsler and J. M. Mocenigo. The Yoix(R) scripting language: a different way of writing Java ^{T M} applications. Software—Practice and Experience, 37(6): 643–667, May 2007. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Dmitriev:2002:LSM

[Dmi02] Mikhail Dmitriev. Languagespecific make technology for the Java programming language. ACM SIGPLAN Notices, 37(11):373–385, November 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dmitriev:2004:PJA

[Dmi04]  M. Dmitriev.  Profiling

Java applications using code hotswapping and dynamic call graph revelation (position paper). Software Engineering Notes, 29(1):139–150, 2004. CODEN ???? ISSN 0163-5948.

Duplantis:2002:VFA

[DMKN02] Willa Duplantis, Eve Mac-Gregor, Maria M. Klawe, and Michele Ng. ’virtual family’: an approach to introducing Java programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(2):40–43, June 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp:

//ftp.math.utah.edu/pub/
mirrors/ftp.ira.uka.de/
bibliography/Misc/DBLP/

2002.bib. Women and Computing.

Dietl:2005:TSC

[DMP05] W. Dietl, P. Muller, and A. PoetzschHeffter. A type system for checking applet isolation in Java card. Lecture Notes in Computer Science, 3362:129–150, 2005. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ducournau:2009:EAO

[DMP09] Roland Ducournau, Floréal Morandat, and Jean Privat. Empirical assessment of object-oriented implementations with multiple inheritance and static typing. ACM SIGPLAN Notices, 44(10):41–60, October 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

deMoor:2008:TID

[dMSAV08] Oege de Moor, Damien Sereni, Pavel Avgustinov, and Mathieu Verbaere. Type inference for datalog and its application to query optimisation. In Lenzerini and Lembo [LL08a], pages 291–300. ISBN 1-60560-932-3. LCCN ????

Dershem:2002:AJL

[DMU02] Herbert L. Dershem, Ryan L. McFall, and Ngozi Uti. Animation of Java linked lists. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34 (1):53–57, March 2002. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Dyer:2006:NPD

[DNR06] Robert Dyer, Harish Narayanappa, and Hridesh Rajan. Nu: preserving design modularity in object code. ACM SIGSOFT Software Engineering Notes, 31(6):1–2, November 2006. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Detlefs:2005:STP

[DNS05] David Detlefs, Greg Nelson, and James B. Saxe. Simplify: a theorem prover for program checking. Journal of the ACM, 52(3):365–473, May 2005. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Dobbing:2001:OSJ

[Dob01a] Brian Dobbing. Overview of the Sun Java Community Process’s Real-Time Expert Group specification of RT-Java: Session Summary. ACM SIGADA Ada Letters, 21(1):18–19, March 2001. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Dobbing:2001:RPH

[Dob01b] Brian Dobbing. The Ravenscar profile for high-integrity Java programs? ACM SIGADA Ada Letters, 21(1): 56–61, March 2001. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Doernhoefer:2006:J

[Doe06] Mark Doernhoefer. JavaScript.

ACM SIGSOFT Software Engineering Notes, 31(4):16–24, July 2006. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

deOliveira:2003:JMT

[dOHS^{+}03a] Jauvane C. de Oliveira, Mojtaba Hosseini, Shervin Shirmohammadi, François Malric, Saeid Nourian, Abdulmotaleb El Saddik, and Nicolas D. Georganas. Java multimedia telecollaboration. IEEE MultiMedia, 10(3):18– 29, July–September 2003. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/csdl.computer.org/comp/
mags/mu/2003/03/u3018abs.
htm;   http://csdl.computer.
org/dl/mags/mu/2003/03/
u3018.htm;   http://csdl.
computer.org/dl/mags/mu/
2003/03/u3018.pdf.

Oliveira:2003:JMT

[dOHS^{+}03b] Jauvane C. de Oliveira, Mojtaba Hosseini, Shervin Shirmohammadi, François Malric, Saeid Nourian, Abdulmotaleb El Saddik, and Nicolas D. Georganas. Java multimedia telecollaboration. IEEE MultiMedia, 10(3):18– 29, July–September 2003. CODEN IEMUE4. ISSN 1070-986X (print), 1941-0166 (electronic). URL http:/

/csdl.computer.org/comp/
mags/mu/2003/03/u3018abs.
htm;   http://csdl.computer.
org/dl/mags/mu/2003/03/
u3018.htm;   http://csdl.
computer.org/dl/mags/mu/
2003/03/u3018.pdf.

Dorobonceanu:2002:CFN

[Dor02] Bogdan Dorobonceanu. Comparing fuzzy numbers. Dr. Dobb’s Journal of Software Tools, 27(12):38, 40, 42, 44– 45, December 2002. CO-DEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/documents/s=
7718/ddj0212e/.

Denti:2005:MPJ

[DOR05] Enrico Denti, Andrea Omicini, and Alessandro Ricci. Multiparadigm Java–Prolog integration in tu Prolog. Science of Computer Programming, 57(2):217–250, August 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Dorin:2007:LR

[Dor07] Philip M. Dorin. Laboratory redux. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(2):84–87, June 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp:

//ftp.math.utah.edu/pub/
mirrors/ftp.ira.uka.de/
bibliography/Misc/DBLP/
2007.bib.

Distefano:2008:JTP

[DP08] Dino Distefano and Matthew J.

Parkinson. jStar: towards practical verification for Java. ACM SIGPLAN Notices, 43(10):213–226, September 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Delbourg:2002:JBC

[DPT^{+}02] D. Delbourg, G. Penillault, T. K. Tuong, M. Decourt, N. Borome, H. Harroch, B. Lessellier, B. Waast, and J. P. Mouffron. A Java-based control system for the Orsay tandem accelerator. Pramana: Journal of Physics, 59 (6):1025–1034, 2002. CODEN PRAMCI. ISSN 0304-4289.

Dray:2000:NPA

[Dra00] Jim Dray. NIST performance analysis of the final round Java AES candidates. In NIST [NIS00], pages 149–

160. ISBN ???? LCCN

                                        ???? URL http://csrc.
                                        nist.gov/encryption/aes/
                                        round2/conf3/aes3conf.htm;
                                         http://csrc.nist.gov/encryption/
                                        aes/round2/conf3/papers/
                                        AES3Proceedings-1.pdf;
                                        http://csrc.nist.gov/encryption/
                                        aes/round2/conf3/papers/
                                        AES3Proceedings-2.pdf;
                                        http://csrc.nist.gov/encryption/
                                        aes/round2/conf3/papers/
                                        AES3Proceedings-3.pdf;
                                        http://csrc.nist.gov/encryption/
                                        aes/round2/conf3/papers/
                                        AES3Proceedings.pdf.

Drossopoulou:2001:AMJ

[Dro01a] Sophia Drossopoulou. An abstract model of Java dynamic linking and loading. Lecture Notes in Computer Science, 2071:53–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.
                                        com/link/service/series/
                                        0558/bibs/2071/20710053.
                                        htm;   http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2071/20710053.
                                        pdf.

Drozdek:2001:DSA

[Dro01b] Adam Drozdek. Data structures and algorithms in Java. Brooks/Cole, Pacific Grove, CA, USA, 2001. ISBN 0-534-37668-1. xvii + 622 pp. LCCN QA76.73.J38 D695 2001.

Delzanno:2002:TAV

[DRV02] Giorgio Delzanno, Jean-François Raskin, and Laurent Van Begin. Towards the automated verification of multithreaded Java programs. Lecture Notes in Computer Science, 2280:173–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2280/22800173.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2280/22800173.
pdf.

Daconta:2000:XDJ

[DS00a] Michael C. Daconta and Al Saganich. XML development with Java 2. Howard W. Sams, Indianapolis, IN 46268, USA, 2000. ISBN 0-672-31653-6. ix + 429 pp. LCCN QA76.76.H94 D32 2000.

DePauw:2000:VRP

[DS00b] Wim De Pauw and Gary Sevitsky. Visualizing reference patterns for solving memory leaks in Java. Concurrency: Practice and Experience, 12(14):1431–1454, December 10, 2000. CO-DEN CPEXEI. ISSN 1040-

                                        3108. URL http://www3.
                                        interscience.wiley.com/
                                        cgi-bin/abstract/76500722/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=76500722&PLACEBO=IE.
                                        pdf.

DiStefano:2000:JKE

[DS00c] Antonella Di Stefano and Corrado Santoro. A Java kernel for embedded systems in distributed process control. IEEE Concurrency, 8(4):55– 63, October/December 2000. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http:

                                        //dlib.computer.org/pd/
                                        books/pd2000/pdf/p4055.
                                        pdf;   http://www.computer.
                                        org/concurrency/pd2000/
                                        p4055abs.htm.

Aires-de-Sousa:2002:JJT

[dS02] J. Aires de Sousa. JA-TOON: Java tools for neural networks. Chemometrics and Intelligent Laboratory Systems, 61(1):167–173, February 28, 2002. CODEN CILSEN. ISSN 0169-7439.

Ding:2004:EJP

[DS04] Haibiao Ding and Mansur H. Samadzadeh. Extraction of Java program fingerprints for software authorship identification. The Journal of systems and software, 72(1): 49–57, June 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Desai:2009:AIC

[DS09] Aniruddha Desai and Jugdutt Singh. Architecture independent characterization of embedded Java workloads. IEEE Computer Architecture Letters, 8(1):29–32, January/ June 2009. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Drejhammar:2003:FJD

[DSBH03] F. Drejhammar, C. Schulte, P. Brand, and S. Haridi. Flow Java: Declarative concurrency for Java. Lecture Notes in Computer Science, 2916:346–360, 2003. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

daSilva:2005:EEJ

[dSC05] A. Faustino da Silva and V. Santos Costa. An experimental evaluation of JAVA JIT technology. J.UCS: Journal of Universal Computer Science, 11(7):1291–1310, ???? 2005. CODEN ???? ISSN 0948-6968. URL http:/

/www.jucs.org/jucs_11_7/
an_experimental_evaluation_
of.

daSilva:2006:OEO

[dSC06] A. F. da Silva and V. S.

Costa. Our experiences with optimizations in Sun’s Java just-in-time compilers.

J.UCS: Journal of Universal Computer Science, 12(7):788– 810, ???? 2006. CODEN ???? ISSN 0948-6968. URL http:

//www.jucs.org/jucs_12_
7/our_experiences_with_
optimizations.

Dietrich:2001:RGU

[DSCU01] Suzanne W. Dietrich, Dan Suceava, Chakrapani Cherukuri, and Susan D. Urban. A reusable graphical user interface for manipulating objectoriented databases using Java and XML. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):362–366, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Danelutto:2002:LSP

[DT02] M. Danelutto and P. Teti. Lithium: a structured parallel programming environment in Java. Lecture Notes in Computer Science, 2330:844– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2330/23300844.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2330/23300844.
pdf.

DeSutter:2004:CJL

[DTD04] B. DeSutter, F. Tip, and J. Dolby. Customization of Java library classes using type constraints and profile information. Lecture Notes in Computer Science, 3086:585– 610, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ducournau:2008:PHA

[Duc08] Roland Ducournau. Perfect hashing as an almost perfect subtype test. ACM Transactions on Programming Languages and Systems, 30(6): 33:1–33:56, October 2008. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Duddy:2006:BRK

[Dud06]  Keith Duddy.  Book re-

view: Kevin Lano, Advanced Systems Design with Java, UML and MDA. Elsevier (2005). ISBN 0-7506-6496-7. £27.99/US$44.95. 378 pp. Softbound. The Computer Journal, 49(4):500–501, July 2006. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/full/49/
4/500;   http://comjnl.
oxfordjournals.org/cgi/
reprint/49/4/500.

Dietrich:2002:JDC

[DUK02] Suzanne W. Dietrich, Susan D. Urban, and Ion Kyriakides. JDBC demonstration courseware using Servlets and Java Server Pages. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):266– 270, March 2002. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Dunn:2002:JR

[Dun02] Douglas Dunn. Java rules. Addison-Wesley, Reading, MA, USA, 2002. ISBN 0-201-70916-3. xxv + 690 pp. LCCN QA76.73.J38 D84 2002.

Durney:2002:EJC

[Dur02] Brian Durney. The essential Java class reference for programmers. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2002. ISBN 0-13-093385-6 (paperback). viii + 168 pp. LCCN QA76.73.J38 D863 2002.

Dobbing:2001:RSA

[DV01] Brian Dobbing and Tullio Vardanega. Report of session: analysis of the J consortium real-time Java proposal. ACM SIGADA Ada Letters, 21(1): 17–18, March 2001. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Draganova:2007:TAW

[DV07] Chrisina Draganova and Vassil Vassilev. Teaching AJAX in Web-centric courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39 (3):311, September 2007. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Distasio:2007:ICS

[DW07] Joseph Distasio and Thomas Way. Inclusive computer science education using a ready-made computer game framework. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):116– 120, September 2007. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Dwelly:2000:JXL

[Dwe00a]  Andrew Dwelly.  Java,

XML, and literate programming. Dr. Dobb’s Journal of Software Tools, 25 (2):62, 64–66, 68, February 2000. CODEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/ftp/2000/
2000_02/litjava.txt;   http:
//www.ddj.com/ftp/2000/
2000_02/litjava.zip.

Dwelly:2000:XRP

[Dwe00b] Andrew Dwelly. XML, reflective pattern matching, and Java. Dr. Dobb’s Journal of Software Tools, 25(6): 46, 49–52, 54, June 2000. CODEN DDJOEB. ISSN 1044-789X. URL http:

                                        //www.ddj.com/ftp/2000/
                                        2000_06/marius05.zip;
                                        http://www.ddj.com/ftp/
                                        2000/2000_06/xmljava.txt.

Dale:2001:IJS

[DWH01] Nell B. Dale, Chip Weems, and Mark R. Headington. Introduction to Java and software design. Jones and Bartlett, Boston, MA, USA, 2001. ISBN 0-7637-1064-4. xxiii + 771 pp. LCCN QA76.73.J38 D34 2001.

Deng:2005:DRE

[DYH05] A. Deng, H. Yu, and S. Hu. Design and realization of embedded system development platform based on Java technology. Information and Control, 34(1):96–103, 2005. CO-DEN IFCNA4. ISSN 1002-0411.

Ding:2003:LJB

[DZHS03] K. Ding, K. Zhou, F. He, and Y. Shen. LDA — A Java-based linkage disequilibrium analyzer. Bioinformatics, 19(16):2147–2148, 2003. CODEN ???? ISSN 1367-4803 (print), 1367-4811 (electronic).

Edwards:2006:JAE

[EA06] James Edwards and Cameron Adams. The JavaScript anthology: 101 essential tips, tricks and hacks. SitePoint Pty. Ltd., Collingwood, VIC, Australia, 2006. ISBN 0-9752402-6-9. xvi + 588 pp. LCCN QA76.73.J39 E392 2006. URL http://

www.oreilly.com/catalog/
9780975240267.

Eaddy:2001:CVJ

[Ead01] Marc Eaddy. C# versus Java. Dr. Dobb’s Journal of Software Tools, 26(2):74, 76, 78, 80, 82, February 2001. CODEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/ftp/2001/
2001_02/csharpj.txt.

Earls:2003:JSM

[Ear03] A. R. Earls. With Java skills, market a complex picture. Application Development Trends, 10(12):28, 2003. CODEN ???? ISSN 1073-9564.

Eberhart:2002:AGJ

[Ebe02] Andreas Eberhart. Automatic generation of Java/ SQL based inference engines from RDF schema and RuleML. Lecture Notes in Computer Science, 2342:102– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/

0558/bibs/2342/23420102.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2342/23420102.
pdf.

Ernest:2005:WMD

[EBG^{+}05] J. C. Ernest, A. S. Bowser, S. Ghule, S. Sudireddy, J. P. Porter, D. A. Talbert, and M. J. Kosa. Weathering MindStorms with drizzle and DIODE in CS0. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3): 353, September 2005. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Eckel:2000:TJ

[Eck00] Bruce Eckel. Thinking in Java. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2000. ISBN 0-13-027363-5. 1127 pp. LCCN QA76.73.J38 E25 2000. US$44.99. URL http:

//www.jowsey.com/books-
music/javabooks.html;
http://www.phptr.com/ptrbooks/
ptr_0130273635.html.

Eckstein:2002:JEB

[Eck02] Robert Eckstein. Java Enterprise best practices. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2002. ISBN 0-596-00384-6. x + 277 pp. LCCN QA76.73.J38 J3662 2002.

Edmondson:2009:PFY

[Edm09] Carol Edmondson. Proglets for first-year programming in Java. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(2):108–112, June 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Edwards:2000:CJC

[Edw00] W. Keith Edwards. Core JINI — The Complete Video Course. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2000. ISBN 0-13-086386-6. ???? pp. LCCN ???? US$129.99. URL http:/

/www.phptr.com/ptrbooks/
ptr_0130863866.html;   http:
//www.sun.com/books/catalog/
edwards2/.

Edwards:2001:CJ

[Edw01] W. Keith Edwards. Core Jini. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2001. ISBN 0-13-089408-7. xliii + 962 pp. LCCN QA76.9.D5 E38 2001. US$54.99. URL http:/

/www.phptr.com/ptrbooks/
ptr_0130894087.html.

Eberhart:2002:JTU

[EF02] Andreas Eberhart and Stefan Fischer. Java tools: using XML, EJB, CORBA, Servlets and SOAP. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 2002.

ISBN 0-471-48666-3. 493 pp. LCCN QA76.73.J38 E24

                                        2002. URL http://www.
                                        i-u.de/schools/eberhart/
                                        book/.

Efford:2000:DIP

[Eff00] Nick Efford. Digital image processing: a practical introduction using Java. Addison-Wesley, Reading, MA, USA, 2000. ISBN 0-201-59623-7. xxiii + 340 + 4 pp. LCCN TA1637 .E35 2000. Includes CD-ROM.

Edelstein:2003:FTM

[EFG^{+}03] Orit Edelstein, Eitan Farchi, Evgeny Goldin, Yarden Nir, Gil Ratsaby, and Shmuel Ur. Framework for testing multi-threaded Java programs. Concurrency and Computation: Practice and Experience, 15(3–5):485–499, March/April 2003. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Emmi:2007:LA

[EFJM07] Michael Emmi, Jeffrey S. Fischer, Ranjit Jhala, and Rupak Majumdar. Lock allocation. ACM SIGPLAN Notices, 42(1):291–296, January 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Edelstein:2001:MJP

[EFN^{+}01] Orit Edelstein, Eitan Farchi, Yarden Nir, Gil Ratsaby, and Shmuel Ur. Multithreaded Java program test generation. In ACM [ACM01b], page 181. ISBN 1-58113-359-6. LCCN QA76.9.O35

A26 2001. URL http://www.
philippsen.com/JGI2001/
camerareadyabstracts/18.
html;   http://www.philippsen.
com/JGI2001/finalpapers/
18500181.ps.

Edelstein:2002:MJP

[EFN^{+}02] O. Edelstein, E. Farchi, Y. Nir, G. Ratsaby, and S. Ur. Multithreaded Java program test generation. IBM Systems Journal, 41(1):111– 125, ???? 2002. CODEN IBMSA7. ISSN 0018-8670.

URL http://www.research.
ibm.com/journal/sj/411/
edelstein.html;   http:
//www.research.ibm.com/
journal/sj/411/edelstein.
pdf.

Elliott:2008:HHS

[EFO08] James Elliott, Ryan Fowler, and Tim OB́rien. Harnessing Hibernate: [a step-by-step guide to Java persistence]. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2008. ISBN 0-596-51772-6. xiv + 363 pp. LCCN ???? US$39.99.

Eeckhout:2003:HJP

[EGD03] Lieven Eeckhout, Andy Georges, and Koen De Bosschere. How Java programs interact with virtual machines at the microarchitectural level. ACM SIG-PLAN Notices, 38(11):169– 186, November 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ertl:2002:VGE

[EGKP02] M. Anton Ertl, David Gregg, Andreas Krall, and Bernd Paysan. Vmgen — a generator of efficient virtual machine interpreters. Software —Practice and Experience, 32(3):265–294, March 2002. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http:/

/www.complang.tuwien.ac.
at/anton/vmgen/;   http:
//www.complang.tuwien.ac.
at/papers/ertl+02.ps.gz;
 http://www3.interscience.
wiley.com/cgi-bin/abstract/
90010508/START;   http:/
/www3.interscience.wiley.
com/cgi-bin/fulltext?ID=
90010508&PLACEBO=IE.pdf.

ElKharashi:2002:JPJ

[EGLZ02] M. W. ElKharashi, F. Gebali, K. F. Li, and F. Zhang. The JAFARDD processor: a Java architecture based on a folding algorithm, with reservation stations, dynamic translation, and dual processing. IEEE Transactions on Consumer Electronics, 48(4): 1004–1015, 2002. CODEN ITCEDA. ISSN 0098-3063.

Escribano:2008:DTJ

[EGST08] Carmen Escribano, Antonio Giraldo, María Asunción Sastre, and Mario Trotta. Digital topology Java applet. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):329, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Egyedi:2001:SFC

[Egy01] T. Egyedi. Strategies for de facto compatibility: Standardization, proprietary and open source approaches to Java. Knowledge, Technology, and Policy, 14(2):113– 128, July 1, 2001. CODEN ???? ISSN 0897-1986.

Eason:2004:PDU

[EH04] R. Eason and G. Heath.

Paintbrush of discovery: Using Java applets to enhance mathematics education. PRIMUS (Problems Resources and Issues in Mathematics Undergraduate Studies), 14(1):79–95, 2004. CO-DEN ???? ISSN 1051-1970.

Ekman:2007:JEJ

[EH07] Torbjörn Ekman and Görel Hedin. The JastAdd extensible Java compiler. ACM SIGPLAN Notices, 42(10):1– 18, October 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Eich:2005:JTY

[Eic05] Brendan Eich. JavaScript at ten years. ACM SIGPLAN Notices, 40(9):129, September 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Eluard:2001:OSJ

[ÉJD01] Marc Éluard, Thomas Jensen, and Ewen Denne. An operational semantics of the Java card firewall. Lecture Notes in Computer Science, 2140:95– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2140/21400095.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2140/21400095.
pdf.

Emmerich:2001:CTJ

[EK01] Wolfgang Emmerich and Nima Kaveh. Component technologies: Java Beans, COM, CORBA, RMI, EJB and the CORBA Component Model. ACM SIGSOFT Software Engineering Notes, 26 (5):311–312, September 2001. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Engelbrecht:2003:TSB

[EK03] R. L. Engelbrecht and D. G.

Kourie. Translating Smalltalk blocks to Java. IEE Proceedings Software, 150(3):203– 211, 2003. CODEN ???? ISSN 1462-5970.

El-Kharashi:2001:ATA

[EKEL01] M. Watheq El-Kharashi, Fayez Elguibaly, and Kin F. Li. Adapting Tomasulo’s algorithm for bytecode folding based Java processors. ACM SIGARCH Computer Architecture News, 29(5):1– 8, December 2001. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Epstein:2000:JQ

[EKM00] David Epstein, Joseph Kiniry, and John Motil. Java Q&A: What is “JJ”? Dr. Dobb’s Journal of Software Tools, 25(4):101–106, April 2000. CODEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/ftp/2000/
2000_04/jqa400.txt.

Elkarablieh:2007:SSA

[EKVM07] Bassem Elkarablieh, Sarfraz Khurshid, Duy Vu, and Kathryn S. McKinley. STARC: static analysis for efficient repair of complex data. ACM SIGPLAN Notices, 42(10):387–404, October 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Eisenbach:2001:SIF

[EL01] Susan Eisenbach and Gary T. Leavens. Special issue: formal techniques for Java programs. Concurrency and Computation: Practice and Experience, 13(13):1121–1123, November 2001. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (elec-

tronic). URL http://www3.
interscience.wiley.com/
cgi-bin/abstract/88011339/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=88011339&PLACEBO=IE.
pdf.

Eckstein:2002:JS

[EL02] Robert Eckstein and Marc Loy. Java Swing. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2002. ISBN 0-596-00408-7. xxiv + 1252 pp. LCCN QA76.73.J38 E27 2002.

Elnagar:2004:GPP

[EL04] A. Elnagar and L. Lulu. A global path planning Javabased system for autonomous mobile robots. Science of Computer Programming, 53 (1):107–122, 2004. CO-DEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Edelson:2009:JC

[EL09] Justin Edelson and Henry Liu. JRuby cookbook. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2009. ISBN 0-596-15506-9. 250 (est.) pp. LCCN QA76.73.R83 E33 2008.

Ellis:2000:TMD

[Ell00] Ainslie Ellis. Toolbook multimedia demonstrations for Java programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3): 181, September 2000. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Elliott:2006:GSH

[Ell06]  James Elliott.  Getting

started with Hibernate 3. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2006. ISBN 0-596-52818-3. LCCN QA76.73.J38 E45 2006eb; QA76.73.J38. URL http://

www.oreilly.com/catalog/
9780596528188.

Eisenbach:2004:FTJ

[ELM^{+}04] S. Eisenbach, G. T. Leavens, P. Muller, A. PoetzschHeffter, and E. Poll. Formal techniques for Java-like programs. Lecture Notes in Computer Science, 3013:62– 71, 2004. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic).

Everitt:2003:JBI

[EM03] R. A. J. Everitt and P. W. McOwan. Java-based Internet biometric authentication system. IEEE Transactions on Pattern Analysis and Machine Intelligence, 25 (9):1166–1171, 2003. CODEN ???? ISSN 0162-8828.

Eisenberg:2004:ELX

[EM04] Andrew Eisenberg and Jim Melton. An early look at XQuery API for Java (XQJ). SIGMOD Record (ACM Special Interest Group on Management of Data), 33(2):105– 111, 2004. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Emurian:2004:PIT

[Emu04] H. H. Emurian. A programmed instruction tutoring system for Java: consideration of learning performance and software self-efficacy. Computers in Human Behavior, 20(3):423–459, 2004. CO-DEN CHBEEQ. ISSN 0747-5632.

English:2000:MNCa

[Eng00] Marie English. Micro news: New benchmark for Unigraphics V15; wireless applications grow; tool set for the Java Card platform; biomechanical discovery affects mobile applications, robots; hard to navigate Web. IEEE Micro, 20(3):3–3, 86–87, May/ June 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2000/pdf/m3003.
pdf.

Englander:2002:JS

[Eng02] Robert Englander. Java and SOAP. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2002. ISBN 0-596-00175-4. xv + 258 pp. LCCN QA76.76.H94 E54 2002. US$39.95. URL http://

safari.oreilly.com/0596001754;
 http://www.oreilly.com/
catalog/9780596001759;
http://www.oreilly.com/
catalog/javasoap.

English:2004:AAG

[Eng04] John English. Automated assessment of GUI programs using JEWL. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):137–141, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

English:2006:CAA

[Eng06] John English. The checkpoint automated assessment system. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):337, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Elmas:2007:GRT

[EQT07] Tayfun Elmas, Shaz Qadeer, and Serdar Tasiran. Goldilocks: a race and transaction-aware Java runtime. ACM SIG-PLAN Notices, 42(6):245– 255, June 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Edwards:2001:JEE

[ER01] W. Keith Edwards and Tom Rodden. Jini Example by Example. Sun BluePrints Program. Sun Microsystems Press, Palo Alto, CA, USA, 2001. ISBN 0-13-033858-3. xii + 569 pp. LCCN QA76.9.D5E383 2001.

                                        URL http://www.sun.com/
                                        books/catalog/edwards4/
                                        index.html.

English:2009:ESP

[ER09] John English and Tammy Rosenthal. Evaluating students’ programs using automated assessment: a case study. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):371, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Elsharnouby:2005:USJ

[ES05a] T. Elsharnouby and A. U. Shankar. Using SeSF Java in teaching introductory network courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):67–71, 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Elsharnouby:2005:UST

[ES05b] Tamer Elsharnouby and A. Udaya Shankar. Using SeSFJava in teaching introductory network courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37 (1):67–71, March 2005. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Evripidou:2006:MMA

[ES06] Paraskevas Evripidou and George Samaras. Metacomputing with mobile agents. International Journal of Parallel Programming, 34 (5):429–458, October 2006. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=34&
issue=5&spage=429.

Saddik:2000:JJA

[ESGS00] Abdulmotaleb El Saddik, Shervin Shirmohammadi, Nicolas D. Georganas, and Ralf Steinmetz. JASMINE: Java Application Sharing in Multiuser INteractive Environments. Lecture Notes in Computer Science, 1905:214– ??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/1905/19050214.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/1905/19050214.
pdf.

Espak:2006:JRB

[Esp06] M. Espák. Japlo: Rulebased programming on Java. J.UCS: Journal of Universal Computer Science, 12 (9):1177–1189, ???? 2006. CODEN ???? ISSN 0948-6968. URL http:/

/www.jucs.org/jucs_12_9/
japlo_rule_based_programming.

Evripidou:2001:PMP

[ESPP01] Paraskevas Evripidou, George Samaras, Christoforos Panayiotou, and Evaggelia Pitoura. The PaCMAn Metacomputer: parallel computing with Java mobile agents. Future Generation Computer Systems, 18(2):265–280, October 2001. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/60/31/33/abstract.
html.

Esquembre:2004:EJS

[Esq04] F. Esquembre. Easy Java simulations: a software tool to create scientific simulations in Java. Computer Physics Communications, 156 (2):199–204, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Eisenbach:2002:EDJ

[ESS02] Susan Eisenbach, Chris Sadler, and Shakil Shaikh. Evolution of distributed Java programs. Lecture Notes in Computer Science, 2370:51– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2370/23700051.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2370/23700051.
pdf.

Erdogan:2004:DEE

[ESS04] N. Erdogan, Y. E. Selcuk, and O. Sahingoz. A distributed execution environment for shared Java objects. Information and Software Technology, 46(7):445– 455, 2004. CODEN ISOTE7. ISSN 0950-5849 (print), 1873-6025 (electronic).

Estell:2001:IWB

[Est01] John K. Estell. IPP: a Web-based interactive programming portfolio. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):149– 153, March 2001. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Estrella:2002:WWG

[Est02] Steven Estrella. The Web wizard’s guide to JavaScript. Addison-Wesley, Reading, MA, USA, 2002. ISBN 0-201-75833-4 (paperback). x + 198 pp. LCCN QA76.73.J39 E88 2002.

Eberhard:2001:EOC

[ET01] John Eberhard and Anand Tripathi. Efficient object caching for distributed Java RMI applications. Lecture Notes in Computer Science, 2218:15–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.
                                        com/link/service/series/
                                        0558/bibs/2218/22180015.
                                        htm;   http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2218/22180015.
                                        pdf.

Emory:2002:JDL

[ET02] David Emory and Roberto Tamassia. JERPA: a distance-learning environment for introductory Java programming courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1): 307–311, March 2002. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Eckerdal:2005:NJP

[ET05] Anna Eckerdal and Michael Thuné. Novice Java programmers’ conceptions of “object” and “class”, and variation theory. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):89–93, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Eberhard:2007:MOC

[ET07] John Eberhard and Anand Tripathi. Mechanisms for object caching in distributed applications using Java RMI. Software—Practice and Experience, 37(8):799–831, July 10, 2007. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ethington:2001:DPS

[Eth01]  Russ Ethington.  Diag-

nosing proxy server problems. Dr. Dobb’s Journal of Software Tools, 26(3): 104, 106, 108, 110, March 2001. CODEN DDJOEB. ISSN 1044-789X. URL

http://www.ddj.com/ftp/
2001/2001_03/proxy.txt;
 http://www.ddj.com/ftp/
2001/2001_03/proxy.zip.

Eubanks:2005:WCJ

[Eub05] Brian D. Eubanks. Wicked cool Java: code bits, opensource libraries, and project ideas. No Starch Press, San Francisco, CA, USA, 2005. ISBN 1-59327-061-5. ???? pp. LCCN QA76.73.J38 E92 2005. URL ftp:/

                                        /uiarchive.cso.uiuc.edu/
                                        pub/etext/gutenberg/;
                                        http://www.loc.gov/catdir/
                                        toc/ecip0518/2005024341.
                                        html.

Eugster:2006:UPJa

[Eug06] Patrick Eugster. Uniform proxies for Java. ACM SIG-PLAN Notices, 41(10):139– 152, October 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Eichelberger:2002:VJP

[EvG02] H. Eichelberger and J. Wolff von Gudenberg. On the visualization of Java programs. Lecture Notes in Computer Science, 2269:295–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.
                                        com/link/service/series/
                                        0558/bibs/2269/22690295.
                                        htm;   http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2269/22690295.
                                        pdf.

Eichelberger:2004:OOP

[EvG04] Holger Eichelberger and Jürgen Wolff von Gudenberg. Object-oriented processing of Java source code. Software —Practice and Experience, 34(12):1157–1185, October 2004. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Erkan:2007:DSV

[EVS07] Ali S. Erkan, T. J. VanSlyke, and Timothy M. Scaffidi. Data structure visualization with L^{A}TEX and Prefuse. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39 (3):301–305, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Eichler:2005:CJT

[EXA^{+}05] M. L. Eichler, P. R. Xavier, R. C. Araujo, R. C. Forte, and J. C. DelPino. Carbopolis: a Java technology-based free software for environmental education installing Carbopolis. Journal of Computers in Mathematics and Science Teaching, 24(1):43–72, 2005. CODEN ???? ISSN 0731-9258.

Fabry:2002:SDE

[Fab02]  Johan Fabry.  Support-

ing development of enterprise JavaBeans through declarative meta programming. Lecture Notes in Computer Science, 2425:280–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.
                                        com/link/service/series/
                                        0558/bibs/2425/24250280.
                                        htm;   http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2425/24250280.
                                        pdf.

Falco:2000:JBX

[Fal00a] Joe Falco. Java-based XML utility for the NIST machine tool data repository. ????, November 2000. 13 pp. Shipping list number 2001-0146-M.

Falco:2000:JXU

[Fal00b] Joe Falco. Java-based XML utility for the NIST machine tool data repository. Gaithersburg, MD, USA, November 2000. 13 pp. Shipping list no.: 2001-0146-M.

Faulkner:2002:JCN

[Fau02]  B. R. Faulkner.  Java

classes for nonprocedural variogram modeling. Computers and Geosciences, 28(3): 387–397, April 2002. CODEN CGEODT, CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Fleissner:2007:EAA

[FB07] Sebastian Fleissner and Elisa L. A. Baniassad. Epi-aspects: aspect-oriented conscientious software. ACM SIGPLAN Notices, 42(10):659–674, October 2007. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Feizabadi:2003:UAS

[FBR^{+}03] S. Feizabadi, W. Beebee, B. Ravindran, P. Li, and M. Rinard. Utility accrual scheduling with realtime Java. Lecture Notes in Computer Science, 2889:550– 563, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Funika:2004:MSD

[FBS04] W. Funika, M. Bubak, and M. Smetek. Monitoring system for distributed Java applications. Lecture Notes in Computer Science, 3038:472– 479, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fong:2000:PLM

[FC00] Philip W. L. Fong and Robert D. Cameron. Proof linking: modular verification of mobile programs in the presence of lazy, dynamic linking. ACM Transactions on Software Engineering and Methodology, 9 (4):379–409, October 2000. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL

http://www.acm.org/pubs/

articles/journals/tosem/
2000-9-4/p379-fong/p379-
fong.pdf;   http://www.
acm.org/pubs/citations/
journals/tosem/2000-9-4/
p379-fong/.

Fong:2001:PLD

[FC01] Philip W. L. Fong and Robert D. Cameron. Proof linking: Distributed verification of Java classfiles in the presence of multiple classloaders. In USENIX Association [USE01c], page ?? ISBN 1-880446-11-1. LCCN QA76.73.J38 J42 2001. URL

http://www.usenix.org/publications/
library/proceedings/jvm01/

fong.html. Sponsored by the USENIX Association.

Farley:2006:JEN

[FC06] Jim Farley and William Crawford. Java Enterprise in a nutshell. In a nutshell. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, third edition, 2006. ISBN 0-596-10142-2 (paperback). xviii + 871 pp. LCCN QA76.73.J38 F375 2006.

Farley:2002:JEN

[FCF02] Jim Farley, William Crawford, and David Flanagan. Java Enterprise in a Nutshell. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2002. ISBN 0-596-00152-

5. xiii + 971 pp. LCCN QA76.73.J38 J366 2002.

US$39.95. URL http://

safari.oreilly.com/0596001525;
http://www.oreilly.com/
catalog/jentnut2.

Fenton:2002:RTC

[FCHE02] F. H. Fenton, E. M. Cherry, H. M. Hastings, and S. J. Evans. Real-time computer simulations of excitable media: Java as a scientific language and as a wrapper for C and Fortran programs. Biosystems (A6E), 64(1):73– 96, January 2002. CODEN BSYMBO. ISSN 0303-2647.

Farzan:2004:FAJ

[FCMR04] A. Farzan, F. Chen, J. Meseguer, and G. Rosu. Formal analysis of Java programs in JavaFAN. Lecture Notes in Computer Science, 3114:501–505, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fukunari:2001:BWJ

[FCW01] Miki Fukunari, Yu-Liang Chi, and Philip M. Wolfe. Best of Websim99: JavaBeanbased simulation with operational procedure table (OPT). Future Generation Computer Systems, 17(5): 513–523, March 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/30/27/abstract.
html.

Forax:2004:RIJ

[FDR04] R. Forax, E. Duris, and G. Roussel. A reflective implementation of Java multimethods. IEEE Transactions on Software Engineering, 30 (12):1055–1071, December 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
1377197.

Felea:2002:EPJ

[FDTL02] Violeta Felea, Nathalie Devesa, Bernard Toursel, and Pierre Lecouffe. Expressing parallelism in Java applications distributed on clusters. Lecture Notes in Computer Science, 2326:249–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2326/23260249.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2326/23260249.
pdf.

Feijs:2001:MNA

[Fei01] Loe M. G. Feijs. Mechanisms for naming: an algebraic approach with an application to Java. Science of Computer Programming, 39(2–3): 149–188, March ??, 2001. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:/

/www.elsevier.nl/gej-ng/
10/39/21/43/23/20/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/39/21/43/23/
20/article.pdf.

Feigenbaum:2004:JRS

[Fei04] B. Feigenbaum. Java reflection & Smalltalk-like method dispatching. Dr. Dobb’s Journal of Software Tools, 29 (7):42–45, 2004. CODEN DDJOEB. ISSN 1044-789X.

Feinberg:2007:VOO

[Fei07] Dave Feinberg. A visual object-oriented programming environment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):140– 144, March 2007. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fekete:2002:TDS

[Fek02]  Alan Fekete.  Teaching

data structures with multiple collection class libraries. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34 (1):396–400, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Fekete:2008:TSD

[Fek08] Alan D. Fekete. Teaching students to develop threadsafe Java classes. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3): 119–123, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Felber:2003:SAP

[Fel03] P. A. Felber. Semi-automatic parallelization of Java applications. Lecture Notes in Computer Science, 2888: 1369–1383, 2003. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Felber:2004:UJX

[Fel04] U. Dietrich Felber. Using Java and XML in interdisciplinary research: A new data-gathering tool for historians working with Euro-ClimHist. Historical Methods, 37(4):174–185, 2004. CODEN ???? ISSN 0161-5440.

Ferguson:2007:CCM

[Fer07] Arron Ferguson. Creating content management systems in Java. Charles River Media, Inc., 403 VFW Drive, PO Box 417, Rockland, MA 02370, USA, 2007. ISBN 1-58450-466-8 (paperback). ???? pp. LCCN TK6680.5 .M38 2004. URL http:/

/www.loc.gov/catdir/toc/
ecip0612/2006012903.html.

Includes CD-ROM.

Feustel:2002:WSJ

[Feu02] Edward Feustel. Writing secure Java applications: a review of Java Security, 2nd ed. IEEE Distributed Systems Online, 3(3), 2002. ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:

//dsonline.computer.org/
0203/departments/bks_a.
htm.

Flanagan:2000:TBR

[FF00] Cormac Flanagan and Stephen N.

Freund. Type-based race detection for Java. ACM SIG-PLAN Notices, 35(5):219– 232, May 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL

http://www.acm.org/pubs/
articles/proceedings/pldi/
349299/p219-flanagan/p219-
flanagan.pdf;   http://
www.acm.org/pubs/citations/
proceedings/pldi/349299/
p219-flanagan/.

Forman:2005:JRA

[FF05] Ira Forman and Nate Forman. Java Reflection in Action. Manning Publications, Greenwich, CT, USA, 2005. ISBN 1-932394-18-4. xxii + 273 pp. LCCN QA76.73.J38 F674 2005. US$44.95. URL http://

www.manning.com/catalog/
view.php?book=forman.

Furr:2008:CTS

[FF08] Michael Furr and Jeffrey S.

Foster. Checking type safety of foreign function calls. ACM Transactions on Programming Languages and Systems, 30(4):18:1–18:63, July 2008. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Flanagan:2009:FEP

[FF09] Cormac Flanagan and Stephen N. Freund. FastTrack: efficient and precise dynamic race detection. ACM SIGPLAN Notices, 44(6):121–133, June 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Farkas:2000:QEC

[FFB^{+}00] Keith I. Farkas, Jason Flinn, Godmar Back, Dirk Grunwald, and Jennifer M. Anderson. Quantifying the energy consumption of a pocket computer and a Java Virtual Machine. ACM SIGMETRICS Perform. Eval. Rev., 28(1): 252–263, June 2000. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Flanagan:2002:JEN

[FFC02] David Flanagan, Jim Farley, and William Crawford. Java Enterprise in a nutshell: a desktop quick reference. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2002. ISBN 0-596-00152-5. xiii + 971 pp. LCCN QA76.73.J38 J366 2002.

Flanagan:2000:JPL

[FFCM00] David Flanagan, Jim Farley, William Crawford, and Kris Magnusson. Java Professional Library. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, November 2000. ISBN 0-596-00107-X. LCCN ???? URL http://

www.oreilly.com/catalog/
9780596001070;   http://
www.oreilly.com/catalog/
javaprolib.

Flanagan:2008:TAS

[FFLQ08] Cormac Flanagan, Stephen N.

Freund, Marina Lifshin, and Shaz Qadeer. Types for atomicity: Static checking and inference for Java. ACM Transactions on Programming Languages and Systems, 30(4): 20:1–20:52, July 2008. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Freeman:2004:HFD

[FFSB04] Eric Freeman, Elisabeth Freeman, Kathy Sierra, and Bert Bates, editors. Head first design patterns. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2004. ISBN 0-596-00712-4, 0-596-00712-4. xxxvi + 638 pp. LCCN QA76.76.D47 H427 2004eb; QA76.76.D47 F74 2004; QA76.76.D47; QA76.76.D47 H427 2004. URL http://

www.oreilly.com/catalog/
9780596007126.

Franciscus:2005:SR

[FG05] George Franciscus and Danilo Gurovich. Struts recipes. Manning Publications, Greenwich, CT, USA, 2005. ISBN 1-932394-24-9 (paperback). xxix + 489 pp. LCCN TK5105.8885.S76 F73 2005.

Frey:2004:JBU

[FGLS04] H. Frey, D. Gorgen, J. K. Lehnert, and P. Sturm. A Java-based uniform workbench for simulating and executing distributed mobile applications. Lecture Notes in Computer Science, 2952:116– 127, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

FigueroadelCid:2000:RFF

[Fig00] Samuel Figueroa del Cid. A Rigorous Framework for Fully Supporting the IEEE Standard for Floating-Point Arithmetic in High-Level Programming Languages. Ph.D. thesis, Department of Computer Science, New York University, New York, NY, USA, January 2000. 345 pp. URL http:

//www.cs.nyu.edu/csweb/
Research/theses.html;
http://www.cs.nyu.edu/csweb/
Research/Theses/figueroa_
sam.html;   http://www.
cs.nyu.edu/csweb/Research/
Theses/figueroa_sam.pdf;

 http://www.cs.nyu.edu/
csweb/Research/Theses/figueroa_
sam.ps.gz;   http://wwwlib.
umi.com/dissertations/fullcit/
9956669;   http://wwwlib.
umi.com/dissertations/preview/
9956669.

Fitzgerald:2007:GAS

[Fit07] Michael Fitzgerald. Google Ajax search API. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2007. ISBN 0-596-52953-8. LCCN TK5105.8885.A52 F58 2007eb; TK5105.8885.A52. URL

http://www.oreilly.com/
catalog/9780596529536.

Fitzgerald:2009:ARN

[Fit09] Sue Fitzgerald. All I really need to know I learned in CS1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):1, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Fahringer:2001:MDP

[FJ01] Thomas Fahringer and Clóvis Seragiotto Júnior. Modeling and detecting performance problems for distributed and parallel programs with JavaPSL. In ACM [ACM01c], page ?? ISBN 1-58113-293-X. LCCN ???? URL http:

//www.sc2001.org/papers/
pap.pap245.pdf.

Fahringer:2005:JNP

[FJ05a] Thomas Fahringer and Alexandru Jugravu. JavaSymphony: a new programming paradigm to control and synchronize locality, parallelism and load balancing for parallel and distributed computing. Concurrency and Computation: Practice and Experience, 17 (7–8):1005–1025, June/July 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Funika:2005:PIJ

[FJ05b] W. Funika and A. Janik. Providing interoperability for Java-oriented monitoring tools with JINEXT. Lecture Notes in Computer Science, 3515:158–165, 2005. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fields:2000:WDJ

[FK00] Duane K. Fields and Mark A. Kolb. Web Development with JavaServer Pages. Manning Publications, Greenwich, CT, USA, 2000. ISBN 1-884777-99-6. xxvi + 554 pp. LCCN TK5105.8885.J38 F54 2000. US$44.95. URL http:/

                                        /www.manning.com/fields;
                                         http://www.manning.com/
                                        Fields/index.html.  In-

cludes CD-ROM.

Friedman:2003:TFT

[FK03] R. Friedman and A. Kama. Transparent fault-tolerant Java Virtual Machine. Proceedings of the Symposium on Reliable Distributed Systems, EDIT22:319–328, 2003. CO-DEN ???? ISSN 1060-9857.

Fitzgerald:2000:MOC

[FKR^{+}00] Robert Fitzgerald, Todd B.

Knoblock, Erik Ruf, Bjarne Steensgaard, and David Tarditi. Marmot: an optimizing compiler for Java. Software—Practice and Experience, 30(3):199–232, March 2000. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

http://www3.interscience.
wiley.com/cgi-bin/abstract/
71001774/START;   http:/
/www3.interscience.wiley.
com/cgi-bin/fulltext?ID=
71001774&PLACEBO=IE.pdf.

Flanagan:2001:HAA

[FL01] C. Flanagan and K. R. M.

Leino. Houdini, an annotation assistant for ESC/Java. Lecture Notes in Computer Science, 2021:500–517, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ferrari:2002:PLM

[FL02] Giulio Ferrari and Dario Laverde, editors. Programming Lego Mindstorms with Java: the ultimate tool for Mindstorms maniacs! Syngress, Rockland, MA, USA, 2002. ISBN 1-928994-55-5. xxii + 441 pp. LCCN ???? US$19.95.

Frickey:2004:CJA

[FL04] Tancred Frickey and Andrei Lupas. CLANS: a Java application for visualizing protein families based on pairwise similarity. Bioinformatics, 20(18):3702–3704, December 12, 2004. CODEN ???? ISSN 1367-4803 (print), 1367-4811 (electronic).

Flanagan:2000:JEN

[Fla00] David Flanagan. Java examples in a nutshell: a tutorial companion to Java in a nutshell. Nutshell handbook. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2000. ISBN 0-596-00039-1. xvi + 566 pp. LCCN QA76.73.J38 F5518 2000. es.

Flanagan:2002:JND

[Fla02a] David Flanagan. Java in a nutshell: a desktop quick reference. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, fourth edition, 2002. ISBN 0-596-00283-1. xxi + 969 pp. Covers Java 1.4.

Flanagan:2002:JPR

[Fla02b] David Flanagan. JavaScript Pocket Reference. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2002. ISBN 0-596-00411-7. 136 (est.) pp. LCCN ???? US$14.95. URL http://

www.oreilly.com/catalog/
jscriptpr2.

Flanagan:2002:JDG

[Fla02c] David Flanagan. JavaScript:

the definitive guide. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, fourth edition, 2002. ISBN 0-596-00048-0. xvii + 916 pp. LCCN QA76.73.J39 F52 2002.

Flanagan:2004:JENa

[Fla04a] David Flanagan. Java examples in a nutshell. A nutshell handbook; The Java series. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, third edition, 2004. ISBN 0-596-00620-9. xiv + 701 pp. LCCN QA76.73.J38 F485 2004; QA76.73.J38 F552 2004eb; QA76.73.J38 F552 2004; QA76.73.J38 F53 2004. URL http://

www.oreilly.com/catalog/
9780596006204.

Flanagan:2004:JENb

[Fla04b] David Flanagan. Java examples in a nutshell: a tutorial companion to Java in a nutshell. In a nutshell. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, third edition, 2004. ISBN 0-596-00585-7. xiv + 701 pp. LCCN QA76.73.J38 F5518 2004. URL http://

www.oreilly.com/catalog/
9780596005856.

Flanagan:2005:JN

[Fla05a] David Flanagan. Java in a nutshell. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, fifth edition, 2005. ISBN 0-596-00773-6. xxiv + 1224 pp. LCCN QA76.73.J38 F553 2005.

Flanagan:2005:JND

[Fla05b] David Flanagan. Java in a nutshell: [a desktop quick reference]. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, fifth edition, 2005. ISBN 0-596-00773-6. xxiv + 1224 pp. LCCN QA76.73.J38 F553 2005. URL http://

                                        www.oreilly.com/catalog/
                                        9780596007737.

Flanagan:2006:JDG

[Fla06] David Flanagan. JavaScript: the definitive guide. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, fifth edition, 2006. ISBN 0-596-10199-6. xxii + 994 pp. LCCN QA76.73.J39 F53 2006eb; QA76.73.J39 F53 2006; QA76.73.J39. URL

                                        http://www.oreilly.com/
                                        catalog/9780596101992.

Fleury:2000:PJS

[Fle00] Ann E. Fleury. Programming in Java: student-constructed rules. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):197–201, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fleury:2001:ERV

[Fle01] Ann E. Fleury. Encapsulation and reuse as viewed by Java students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):189–193, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Flenner:2003:JPU

[Fle03] Robert Flenner. Java P2P unleashed. Howard W. Sams, Indianapolis, IN 46268, USA, 2003. ISBN 0-672-32399-0 (paperback). xv + 732 pp. LCCN QA76.73.J38 J3788 2003.

Findler:2001:BCB

[FLF01] Robert Bruce Findler, Mario Latendresse, and Matthias Felleisen. Behavioral contracts and behavioral subtyping. ACM SIGSOFT Software Engineering Notes, 26 (5):229–236, September 2001. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Flanagan:2002:ESC

[FLL^{+}02] Cormac Flanagan, K. Rustan M. Leino, Mark Lillibridge, Greg Nelson, James B.

Saxe, and Raymie Stata. Extended static checking for Java. ACM SIGPLAN Notices, 37(5):234–245, May 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fisher:2006:JEN

[FLMS06] Marina Fisher, Ray Lai, Laurence Moroney, and Sonu Sharma, editors. Java EE and .NET interoperability: integration strategies, patterns, and best practices. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2006. ISBN 0-13-147223-2 (paperback). xxxvii + 608 pp. LCCN QA76.76.M52 J23

2006. URL http://www.loc.
gov/catdir/toc/ecip066/
2006000235.html;   http:
//www.phptr.com/bookstore/
product.asp?isbn=0131472232.

Fung:2004:JBP

[FLWW04] Chun Che Fung, Jia Bin Li, Kok Wai Wong, and Kit Po Wong. A Java-based parallel platform for the implementation of evolutionary computation for engineering applications. International Journal of Systems Science, 35(13– 14):741–750, November 2004. CODEN IJSYA9. ISSN 0020-7721.

Freund:2003:TSJ

[FM03] Stephen N. Freund and John C. Mitchell. A type system for the Java bytecode language and verifier. Journal of Automated Reasoning, 30(3– 4):271–321, May 2003. CO-DEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A%3A1025011624925.

Fang:2002:JJB

[FMA02] Xiang Fang, John A. Miller, and Jonathan Arnold. J3DV: a Java-based 3D database visualization tool. Software— Practice and Experience, 32 (5):443–463, April 25, 2002. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/93513335/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=93513335&PLACEBO=IE.
pdf.

Flanagan:2000:JEC

[FMHH^{+}00] David Flanagan, Richard Monson-Haefel, Jason Hunter, Scott Oaks, and Jim Farley. The Java enterprise CD bookshelf. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2000. ISBN 1-56592-850-4 (set), 1-56592-483-5 (book). xvi + 604 pp. LCCN QA76.73.J38 J366 1999. Title from disc label Bonus book has title: Java Enterprise in a nutshell : a desktop quick reference / David Flanagan … [et al.]. 1999 Contents: Java enterprise in a nutshell – Java in a nutshell, 3rd ed. – Java foundation classes in a nutshell – Enterprise JavaBeans, 2nd ed. – Java servlet programming – Java security – Java distributed computing.

Fuzitaki:2003:MNL

[FMMd03] C. N. Fuzitaki, P. B. Menezes, J. P. Machado, and S. A. daCosta. Mapping Nautilus language into Java: Towards a specification and programming environment for distributed systems. Lecture Notes in Computer Science, 2809:243–252, 2003. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Farzan:2005:FJC

[FMR05] Azadeh Farzan, José Meseguer, and Grigore Roşu. Formal JVM code analysis in JavaFAN. Technical report, Department of Computer Science, University of Illinois at Urbana-Champaign, Urbana-Champaign, IL, USA, January 19, 2005. URL http://

                                        www.cs.toronto.edu/~azadeh/
                                        papers/amast04.pdf.

Fu:2005:RTJ

[FMRW05] C. Fu, A. Milanova, B. G. Ryder, and D. G. Wonnacott. Robustness testing of Java server applications. IEEE Transactions on Software Engineering, 31(4):292–311, April 2005. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
1435351.

Ford:2004:LOG

[For04a]  Lindsey Ford.  A learn-

ing object generator for programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):268, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ford:2004:AJW

[For04b] Neal Ford. Art of Java Web development: Struts, Tapestry, Commons, Velocity, JUnit, Axis, Cocoon, InternetBeans, Web-Works. Manning Publications, Greenwich, CT, USA, 2004. ISBN 1-932394-06-0. xxx + 593 pp. LCCN QA76.73.J38 F66

2004. URL http://www.loc.
gov/catdir/bios/ipg051/
2004271136.html;   http:
//www.loc.gov/catdir/description/
ipg051/2004271136.html.

Ford:2006:NFJ

[For06] Neal Ford, editor. No fluff, just stuff anthology: the 2006 edition. Pragmatic Bookshelf, Raleigh, NC, USA, 2006. ISBN 0-9776166-6-5 (paperback). xi + 240 pp. LCCN QA76.76.D47 N646 2006. URL http://

                                        www.oreilly.com/catalog/
                                        9780977616664.

Foster:2003:MM

[Fos03]  K. R. Foster.  Mightier

math. IEEE Spectrum, 40(2): 58–59, February 2003. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fujiwara:2004:SAJ

[FOS^{+}04] Y. Fujiwara, S. Okada, T. Suzuki, Y. Ohnishi, and H. Yoshida. Self-adaptive Java production system and its application to a learning assistance system. IEICE Transactions on Information and Systems E Series D, 87 (9):2186–2194, 2004. CODEN ???? ISSN 0916-8532.

Fox:2000:ESIa

[Fox00a] Geoffrey Fox. Editorial: Special issue: ACM 1999 Java Grande Conference. Concurrency: Practice and Experience, 12(6):341–342, May 2000. CODEN CPEXEI. ISSN 1040-3108. URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract/72515722/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=72515722&PLACEBO=IE.
                                        pdf.

Fox:2000:ESIb

[Fox00b] Geoffrey Fox. Editorial: Special issue: ACM 1999 Java Grande Conference. Concurrency: Practice and Experience, 12(7):493–494, May 2000. CODEN CPEXEI. ISSN 1040-3108. URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/72516216/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=72516216&PLACEBO=IE.
pdf.

Fox:2000:ESIc

[Fox00c] Geoffrey Fox. Editorial: Special issue: ACM 1999 Java Grande Conference. Concurrency: Practice and Experience, 12(8):611–612, July 2000. CODEN CPEXEI. ISSN 1040-3108. URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/72516225/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=72516225&PLACEBO=IE.
pdf.

Fox:2000:CAJ

[Fox00d] Joshua Fox. Collaborative applications and the Java Shared Data Toolkit. Dr. Dobb’s Journal of Software Tools, 25(2):32, 34–40, February 2000. CODEN DDJOEB. ISSN 1044-789X. URL

http://www.ddj.com/ftp/
2000/2000_02/jsdt.txt;
http://www.ddj.com/ftp/
2000/2000_02/jsdt.zip.

Fox:2000:JQW

[Fox00e] Joshua Fox. Java Q&A:

So what is SO KEEPALIVE?

Dr. Dobb’s Journal of Software Tools, 25(9):123–126, 128, 130, September 2000. CODEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/ftp/2000/
2000_09/jqa0900.txt;   http:
//www.ddj.com/ftp/2000/
2000_09/jqa0900.zip.

Fox:2001:SIA

[Fox01a] Geoffrey Fox. Special issue: ACM 2000 Java Grande Conference. Concurrency and Computation: Practice and Experience, 13(8–9):643– 644, July/August 2001. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (elec-

tronic). URL http://www3.
interscience.wiley.com/
cgi-bin/abstract/84503221/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=84503221&PLACEBO=IE.
pdf.

Foxwell:2001:PJD

[Fox01b]  H. J. Foxwell.  Profes-

sional Java Data, by Danny Ayres, John Bell, Carl Bettis, Thomas Bishop, Bjarki Holm, Glenn Mitchell II, Kelly Lin Poon, Sean Rhody, Mike Bogavich, Matthew Ferris, Rick Grehan, Tony Loton, Nitin Nanda, and Mark Wilcox. Markup languages: theory & practice, 3(3):382–383, December 10, 2001. CODEN MLTPFG. ISSN 1099-6621.

Foxwell:2001:JXE

[Fox01c] Harry J. Foxwell. Java and XSLT by Eric M. Burke.

Markup languages: theory & practice, 3(4):446– 447, Fall 2001. CODEN MLTPFG. ISSN 1099-6621.

URL http://mitpress.mit.
eduhttp://mitpress.mit.
edu/catalog/item/default.
asp?sid=81FFBAAF-F78B-4799-
8942-1A7B831BFC9A&ttype=
6&tid=9097.

Foxwell:2001:RPJ

[Fox01d] Harry J. Foxwell. Review:

Professional Java Data by Danny Ayres et al. Markup languages: theory & practice, 3(3):382–383, Summer 2001. CODEN MLTPFG. ISSN 1099-6621.

Foxwell:2002:JX

[Fox02] H. J. Foxwell. Java and XSLT. Markup languages: theory & practice, 3(4):446– 447, September 10, 2002. CO-DEN MLTPFG. ISSN 1099-6621.

Fox:2003:CSE

[Fox03a]  Geoffrey Fox.  Comput-

ing in science & engineering: Web computing: Java and Grande applications. IEEE Distributed Systems Online, 4(4), 2003. ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:

//dsonline.computer.org/
0304/d/web_print.htm.

Fox:2003:JGA

[Fox03b]  Geoffrey Fox.  Java and

Grande applications. Computing in Science and Engineering, 5(1):60–62, January/February 2003. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/dl/mags/cs/
2003/01/c1060.htm;   http:
//csdl.computer.org/dl/
mags/cs/2003/01/c1060.pdf.

Fox:2005:SIA

[Fox05] Geoffrey Fox. Special issue: ACM 2002 Java Grande-ISCOPE Conference. Concurrency and Computation: Practice and Experience, 17 (5–6):419–422, April/May 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Fuhrer:2003:MDV

[FP03] P. Fuhrer and J. Pasquier-Rocha. Massively distributed virtual worlds: a framework approach (MaDViWorld: a Java software framework for massively distributed virtual worlds). Lecture Notes in Computer Science, 2604:111– 121, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fuller:2006:CPB

[FPA^{+}06] Ursula Fuller, Arnold Pears, June Amillo, Chris Avram, and Linda Mannila. A computing perspective on the Bologna Process. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(4):

115–131, December 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp:

//ftp.math.utah.edu/pub/
mirrors/ftp.ira.uka.de/
bibliography/Misc/DBLP/
2006.bib.

Forax:2000:RTP

[FR00] Rémi Forax and Gilles Roussel. Recursive types and pattern-matching in Java. Lecture Notes in Computer Science, 1799:147–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/1799/17990147.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/1799/17990147.
pdf.

Felber:2002:ACC

[FR02] Pascal Felber and Michael K.

Reiter. Advanced concurrency control in Java. Concurrency and Computation: Practice and Experience, 14 (4):261–285, April 10, 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/93513491/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=93513491{\&}PLACEBO=
IE.pdf.

Freeby:2001:CDJ

[Fre01] James M. Freeby. A comparison of DHTML and JAVA applets. Thesis (M.S.), University of California, Santa Cruz, Santa Cruz, CA, USA, 2001.

Frens:2004:TTT

[Fre04] J. D. Frens. Taming the tiger: Teaching the next version of Java. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):151–155, 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fredlund:2005:GCP

[Fre05] L. A. Fredlund. Guaranteeing correctness properties of a Java card applet. Electronic Notes in Theoretical Computer Science, 113(??): 217–233, 2005. CODEN ???? ISSN 1571-0661.

Frenzel:2007:ERB

[Fre07] Leif Frenzel. Experience report: building an eclipsebased IDE for Haskell. ACM SIGPLAN Notices, 42(9): 220–222, September 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frenger:2008:HJ

[Fre08] Paul Frenger. Hard Java. ACM SIGPLAN Notices, 43 (5):5–9, May 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fricke:2002:EJO

[Fri02] V. Fricke. Embedded Java and OSGi — new technologies in the car of the future. VDI Berichte, 1728(??): 41–50, 2002. CODEN ???? ISSN 0083-5560.

Fu:2004:TJW

[FRMW04] C. Fu, B. G. Ryder, A. Milanova, and D. Wonnacott. Testing of Java Web services for robustness. Software Engineering Notes, 29(4):23–33, 2004. CODEN ???? ISSN 0163-5948.

Frost:2007:FGC

[Fro07] Daniel Frost. Fourth grade computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):302– 306, March 2007. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Frost:2008:UJL

[Fro08] Daniel Frost. Ucigame, a Java library for games. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1): 310–314, March 2008. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Frye:2003:SGJ

[Fry03] C. Frye. In search of a gentler Java J2EE. Application Development Trends, 10(12):24– 29, 2003. CODEN ???? ISSN 1073-9564.

Fry:2008:VD

[Fry08] Ben Fry. Visualizing data. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2008. ISBN 0-596-51455-7. xiii + 366 pp. LCCN T385 .F79 2008; T385. URL

                                        http://www.oreilly.com/
                                        catalog/9780596514556.

Foster:2003:UNP

[FS03a] W. Foster and S. Simpson. Using native plug-ins with Java. Research Disclosure, 467:453, 2003. CODEN ???? ISSN 0374-4353.

Fukushima:2003:SFS

[FS03b] K. Fukushima and K. Sakurai. A software fingerprinting scheme for Java using classfiles obfuscation. Lecture Notes in Computer Science, 2908:303–316, 2003. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ferrero:2003:RJB

[FSBP03] A. Ferrero, S. Salicone, C. Bonora, and M. Parmigiani. ReMLab: a Javabased remote, didactic measurement laboratory. IEEE Transactions on Instrumentation and Measurement, 52 (3):710–715, 2003. CO-DEN IEIMAO. ISSN 0018-9456 (print), 1557-9662 (electronic).

Factor:2006:PID

[FSS06] Michael Factor, Assaf Schuster, and Konstantin Shagin. A platform-independent distributed runtime for standard multithreaded Java. International Journal of Parallel Programming, 34(2):113–142, April 2006. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=34&
issue=2&spage=113.

Fuentes:2000:TOM

[FT00] Lidia Fuentes and Josée M.

Troya. Towards an open multimedia service framework. ACM Computing Surveys, 32 (1es), March 2000. CO-DEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/surveys/2000-32-
1/p24-fuentes/. Article No.

24.

Felea:2006:DLB

[FT06] Violeta Felea and Bernard Toursel. Dynamic loadbalancing mechanism for distributed Java applications. Concurrency and Computation: Practice and Experience, 18(3):305–331, March 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Felea:2003:CDO

[FTD03] V. Felea, B. Toursel, and N. Devesa. Les collections distribuées: un outil pour la conception d’applications Java parallèles. (French) [Distributed collections: a tool for creation of paralel Java applications]. Technique et science informatiques : TSI, 22 (3):289–314, 2003. CODEN TTSIDJ. ISSN 0752-4072, 0264-7419.

Fischmeister:2001:EST

[FVK01] Sebastian Fischmeister, Giovanni Vigna, and Richard A. Kemmerer. Evaluating the security of three Java-based mobile agent systems. Lecture Notes in Computer Science, 2240:31–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.
                                        com/link/service/series/
                                        0558/bibs/2240/22400031.
                                        htm;   http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2240/22400031.
                                        pdf.

Freiwald:2002:JBC

[FW02] Uwe Freiwald and Jörg R. Weimar. The Java based cellular automata simulation system — JCASim. Future Generation Computer Systems, 18(7):995–1004, August 2002. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Fang:2003:DGO

[FWL03] Weijian Fang, Cho-Li Wang, and Francis C. M. Lau. On the design of global object space for efficient multithreading Java computing on clusters. Parallel Computing, 29(11–12):1563–1587, November/December 2003. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Fiedler:2005:TMT

[FWR^{+}05] Daniel Fiedler, Kristen Walcott, Thomas Richardson, Gregory M. Kapfhammer, Ahmed Amer, and Panos K. Chrysanthis. Towards the measurement of tuple space performance. ACM SIG-METRICS Perform. Eval. Rev., 33(3):51–62, December 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Fahndrich:2007:EOI

[FX07] Manuel Fahndrich and Songtao Xia. Establishing object invariants with delayed types. ACM SIGPLAN Notices, 42(10):337–350, October 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fink:2008:ETV

[FYD^{+}08] Stephen J. Fink, Eran Yahav, Nurit Dor, G. Ramalingam, and Emmanuel Geay. Effective typestate verification in the presence of aliasing. ACM Transactions on Software Engineering and Methodology, 17(2):9:1–9:??, April 2008. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Gannon:2001:JCC

[G^{+}01] Dennis Gannon et al. Panel discussion: Java, C++, C#, and virtual machines for high performance. In ACM [ACM01b], page ?? ISBN 1-58113-359-6. LCCN QA76.9.O35 A26 2001.

Gabarro:2007:WAD

[Gab07] Steven A. Gabarró. Web application design and implementation: Apache 2, PHP5, MySQL, JavaScript, and Linux/Unix. Quantitative software engineering series. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 2007. ISBN 0-471-77391-3 (cloth). xv + 295 pp. LCCN TK5105.8883 .G33

                                        2007. URL http://www.loc.
                                        gov/catdir/toc/ecip0613/
                                        2006014999.html.

Gadde:2003:JCA

[Gad03] S. Gadde. Java class aliasing during run time. Research Disclosure, 470:764, 2003. CODEN RSDSBB. ISSN 0374-4353.

Gagne:2002:JNB

[Gag02] Greg Gagne. To java.net and beyond: teaching networking concepts using the Java networking API. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1): 406–410, March 2002. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Gehtland:2006:PAW

[GAG06] Justin Gehtland, Dion Almaer, and Ben Galbraith. Pragmatic Ajax: a Web 2.0 primer. Pragmatic Bookshelf, Raleigh, NC, USA, 2006. ISBN 0-9766940-8-5. 200 (est.) pp. LCCN ????

Galambos:2001:LDI

[Gal01]  Leo Galambos.  Lemma-

tizer for document information retrieval systems in JAVA. Lecture Notes in Computer Science, 2234:243–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2234/22340243.

htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2234/22340243.
pdf.

Nicholas:2002:CID

[Gal02] Nicholas Galbreath. Cryptography for Internet and database applications: developing secret and public key techniques with Java. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 2002. ISBN 0-471-21029-3. 400 pp. LCCN QA76.9.A25 G35 2002.

Gamess:2000:PTE

[Gam00] Eric Gamess. plapackJava: Towards an efficient Java interface for high performance parallel linear algebra. Information Processing Letters, 75(5):191–197, October 31, 2000. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.elsevier.nl/gej-ng/
10/23/20/64/30/26/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/23/20/64/30/
26/article.pdf.

Gamess:2003:ESP

[Gam03] E. Gamess. Execution of sequential and parallel Java bytecode in a metacomputing system. Parallel Processing Letters, 13(1):53–64, 2003. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Gaona:2000:RDC

[Gao00] Amparo López Gaona. The relevance of design in CS1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32 (2):53–55, June 2000. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Garber:2000:NBC

[Gar00] Lee Garber. News briefs:

Companies power up networking; magnetic RAM research reaching critical stage; Micro Java is a platform under construction; Lotus offers on-the-fly translation. Computer, 33(8):26–28, August 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//dlib.computer.org/co/
books/co2000/pdf/r8026.
pdf.

Garrido:2001:OOD

[Gar01] Jose M. Garrido. Objectoriented discrete-event simulation with Java: a practical introduction. Series in computer systems. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2001. ISBN 0-306-46688-0. xvi + 256 pp. LCCN QA76.73.J38 G365 2001. Includes CD-ROM with simulation models explained in the book.

Guelfi:2003:SED

[GAR03] Nicolas Guelfi, E. Astesiano, and Gianna Reggio, editors. Scientific engineering for distributed Java applications: international workshop, FIDJI 2002, Luxembourg-Kirchberg, Luxembourg, November 28–29, 2002: revised papers, number 2604 in Lecture notes in computer science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. ISBN 3-540-00679-6 (softcover). LCCN QA76.758 .F53 2002. URL

http://link.springer-ny.
com/link/service/series/
0558/tocs/t2604.htm.

Guelfi:2004:SED

[GAR04] Nicolas Guelfi, Egidio Astesiano, and Gianna Reggio, editors. Scientific Engineering of Distributed Java Applications: Third International Workshop, FIDJI 2003, Luxembourg-Kirchberg, Luxembourg, November 27– 28, 2003: Revised Papers, volume 2952 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. CODEN LNCSD9. ISBN 3-540-21091-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758 .F53 2003. URL

http://link.springer-ny.
com/link/service/series/

0558/tocs/t2952.htm;   http:
//www.springerlink.com/
openurl.asp?genre=issue&
issn=0302-9743&volume=2952;
 http://www.springerlink.
com/openurl.asp?genre=volume&
id=doi:10.1007/b95352.

Gardner:2009:DGP

[Gar09] James Gardner. The definitive guide to Pylons: [Pylons is a lightweight web framework emphasizing flexibility and rapid development using standard tools from the Python community ; includes SQLAlchemy, JavaScript, and WSG!]. The expert’s voice in web development; Books for professionals by professionals. Apress, Berkeley, CA, USA, 2009. ISBN 1-59059-934-9 (paperback). xxv + 536 pp. LCCN ???? US$46,99.

Gates:2003:DTT

[Gat03] L. Gates. Development tools and technologies: Java IDEs further coverage of life cycle. Application Development Trends, 10(7):39–42, 2003. CODEN ADTRF4. ISSN 1073-9564.

Grimm:2001:SAC

[GB01] Robert Grimm and Brian N.

Bershad. Separating access control policy, enforcement, and functionality in extensible systems. ACM Transactions on Computer Systems, 19(1):36–70, 2001. CODEN ACSYEC. ISSN 0734-2071 (print), 1557-7333 (electronic). URL

                                        http://www.acm.org/pubs/
                                        articles/journals/tocs/
                                        2001-19-1/p36-grimm/p36-
                                        grimm.pdf;   http://www.
                                        acm.org/pubs/citations/
                                        journals/tocs/2001-19-1/
                                        p36-grimm/.

Gu:2000:EHP

[GBCW00] W. Gu, N. A. Burns, M. T. Collins, and W. Y. P. Wong. The evolution of a high-performing Java virtual machine. IBM Systems Journal, 39(1):135– 150, ???? 2000. CO-DEN IBMSA7. ISSN 0018-

                                        8670. URL http://www.
                                        almaden.ibm.com/journal/
                                        sj/391/gu.html.

Georges:2007:SRJ

[GBE07] Andy Georges, Dries Buytaert, and Lieven Eeckhout. Statistically rigorous Java performance evaluation. ACM SIGPLAN Notices, 42(10):57–76, October 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Georges:2004:MLP

[GBED04] A. Georges, D. Buytaert, L. Eeckhout, and K. DeBosschere. Method-level phase behavior in Java workloads. ACM SIGPLAN Notices, 39 (10):270–287, 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gonzalez-Castano:2001:JCV

[GCARPC^{+}01] F. J. González-Castaño, L. Anido-Rifón, J. M. Pousada-Carballo, P. S. Rodríguez-Hernández, and R. López-Gómez. A Java/CORBA virtual machine architecture for remote execution of optimization solvers in heterogeneous networks. Software—Practice and Experience, 31(1): 1–16, January 2001. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (elec-

tronic). URL http://www3.
interscience.wiley.com/
cgi-bin/abstract/76502341/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=76502341&PLACEBO=IE.
pdf.

Garti:2000:OMP

[GCB^{+}00] Dror Garti, Shem-Tov Cohen, Amnon Barak, Arie Keren, and Ricardo Szmit. Object mobility for performance improvements of parallel Java applications. Journal of Parallel and Distributed Computing, 60(10):1311–1324, October 2000. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jpdc.
2000.1655;   http://www.
idealibrary.com/links/doi/
10.1006/jpdc.2000.1655/
pdf;   http://www.idealibrary.

                                        com/links/doi/10.1006/jpdc.
                                        2000.1655/ref.

Goldovsky:2005:BVN

[GCEO05] Leon Goldovsky, Ildefonso Cases, J. Enright, Anton, and A. Ouzounis, Christos. BioLayoutJava: Versatile network visualisation of structural and functional relationships. Appl Bioinformatics, 4(1):71–74, 2005. CODEN ???? ISSN 1175-5636.

Goldweber:2001:URU

[GCF^{+}01] Michael Goldweber, Clare Congdon, Barry Fagin, Deborah Hwang, and Frank Klassner. The use of robots in the undergraduate curriculum: experience reports. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1): 404–405, March 2001. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gupta:2000:OJP

[GCH00] Manish Gupta, Jong-Deok Choi, and Michael Hind. Optimizing Java programs in the presence of exceptions. Lecture Notes in Computer Science, 1850:422–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.
                                        com/link/service/series/
                                        0558/bibs/1850/18500422.
                                        htm;   http://link.springer-
                                        ny.com/link/service/series/

0558/papers/1850/18500422.
pdf.

Georges:2004:JPR

[GCRD04] A. Georges, M. Christiaens, M. Ronsse, and K. De Bosschere. JaRec: a portable record/replay environment for multi-threaded Java applications. Software—Practice and Experience, 34(6): 523–547, May 2004. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gasperoni:2000:MPJ

[GD00] Franco Gasperoni and Gary Dismukes. Multilanguage programming on the JVM: The Ada 95 benefits. ACM SIGADA Ada Letters, 20 (4):3–28, December 2000. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL http:/

/www.acm.org/sigada/ada_
letters/dec2000/ada-on-

jvm.pdf. Special Issue: Presentations from SIGAda 2000.

Grose:2002:MXJ

[GDB02] Timothy J. Grose, Gary C.

Doney, and Stephen A. Brodsky. Mastering XMI: Java programming with XMI, XML, and UML. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 2002. ISBN 0-471-38429-1. xli + 434 pp. LCCN QA76.73.J38 G76 2002 Bar. Includes CD-ROM.

Gonzalez:2004:WOO

[GDC^{+}04] S. Gonzalez, W. DeMeuter, P. Costanza, S. Ducasse, R. Gabriel, and T. DHondt. 2nd workshop on objectoriented language engineering for the post-Java era: Back to dynamicity. Lecture Notes in Computer Science, 3344:49– 61, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gravvanis:2008:JMB

[GE08] George A. Gravvanis and Victor N. Epitropou. Java multithreading-based parallel approximate arrowtype inverses. Concurrency and Computation: Practice and Experience, 20(10): 1151–1172, July 2008. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Geary:2000:GJV

[Gea00] David Geary. Graphic Java 2, Volume 4, 2D API. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, third edition, 2000. ISBN 0-13-079668-9. ???? pp. LCCN ???? US$50.

Geary:2001:AJP

[Gea01] David M. Geary. Advanced JavaServer pages. Java 2 platform, enterprise edition series. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2001. ISBN 0-13-030704-1 (paperback). xx + 487 pp. LCCN TK5105.8885.J38 G43 2001.

URL http://www.sun.com/
books/catalog/gearyjsp/.

Gschwind:2000:BTA

[GEAS00] Michael Gschwind, Kemal Ebcioğlu, Erik Altman, and Sumedh Sathaye. Binary translation and architecture convergence issues for IBM System/390. In Conference proceedings of the 2000 International Conference on Supercomputing: Santa Fe, New Mexico, May 8–11, 2000 [ICS ’00], pages 336–347. ACM Press, New York, NY 10036, USA, 2000. ISBN 1-58113-270-0. LCCN ???? URL

http://delivery.acm.org/
10.1145/340000/335264/p336-
gschwind.pdf.

Georges:2008:JPE

[GEB08] Andy Georges, Lieven Eeckhout, and Dries Buytaert. Java performance evaluation through rigorous replay compilation. ACM SIGPLAN Notices, 43(10):367–384, September 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Geer:2005:EBD

[Gee05] David Geer. Eclipse becomes the dominant Java IDE. Computer, 38(7): 16–18, July 2005. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        co/2005/07/r7016.pdf.

Gravvanis:2007:PPA

[GEG07] George A. Gravvanis, Victor N. Epitropou, and Konstantinos M. Giannoutakis. On the performance of parallel approximate inverse preconditioning using Java multithreading techniques. Applied Mathematics and Computation, 190(1):255– 270, July 1, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Gregg:2001:IEJ

[GEK01] David Gregg, M. Anton Ertl, and Andreas Krall. Implementing an efficient Java interpreter. Lecture Notes in Computer Science, 2110:613– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.
                                        com/link/service/series/
                                        0558/bibs/2110/21100613.
                                        htm;   http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2110/21100613.
                                        pdf.

Gelderblom:2000:OCS

[Gel00] J. Helene Gelderblom. OOPtutor: a CBL system for introductory object-oriented programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(2):35–38, June 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gengler:2000:JBM

[Gen00] Barbara Gengler. JavaScript ban may happen. Network Security, 2000(1):4, January 2000. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S1353485800866489.

Gestwicki:2007:CGM

[Ges07] Paul V. Gestwicki. Computer games as motivation for design patterns. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):233– 237, March 2007. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gal:2009:TBJ

[GES^{+}09] Andreas Gal, Brendan Eich, Mike Shaver, David Anderson, David Mandelin, Mohammad R. Haghighat, Blake Kaplan, Graydon Hoare, Boris Zbarsky, Jason Orendorff, Jesse Ruderman, Edwin W. Smith, Rick Reitmaier, Michael Bebenita, Mason Chang, and Michael Franz. Trace-based just-in-time type specialization for dynamic languages. ACM SIGPLAN Notices, 44(6): 465–478, June 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gal-Ezer:2009:PSC

[GEVZ09a] Judith Gal-Ezer, Tamar Vilner, and Ela Zur. Has the paradigm shift in CS1 a harmful effect on data structures courses: a case study. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):126–130, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Gal-Ezer:2009:PYP

[GEVZ09b] Judith Gal-Ezer, Tamar Vilner, and Ela Zur. The professor on your PC: a virtual CS1 course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):191–195, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Gabrilovich:2001:JCI

[GF01] Evgeniy Gabrilovich and Lev Finkelstein. JNI–C++ integration made easy. C/C++ Users Journal, 19(1):10–??, January 2001. CODEN CCU-JEX. ISSN 1075-2838.

Greenfieldboyce:2007:TQI

[GF07] David Greenfieldboyce and Jeffrey S. Foster. Type qualifier inference for Java. ACM SIGPLAN Notices, 42 (10):321–336, October 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

GomezMartin:2003:JVE

[GGG03] P. P. GomezMartin, M. A.

GomezMartin, and P. A. GonzalezCalero. Javy: Virtual environment for casebased teaching of Java Virtual Machine. Lecture Notes in Computer Science, 2773:906– 913, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ghosale:2003:IHP

[GGH^{+}03] N. Ghosale, J. A. Green, J. HernandezHerrero, G. G. Huang, and P. S. Parikh. On implementing a highperformance open API with Java. Bell Labs Technical Journal, 7(4):123–138, 2003. CODEN BLTJFD. ISSN 1089-7089.

Gunnels:2001:FFL

[GGHvdG01] John A. Gunnels, Fred G.

Gustavson, Greg M. Henry, and Robert A. van de Geijn. FLAME: Formal Linear Algebra Methods Environment. ACM Transactions on Mathematical Software, 27(4):422– 455, December 2001. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Genaud:2008:EPC

[GGL^{+}08] Stéphane Genaud, Pierre Gançarski, Guillaume Latu, Alexandre Blansché, Choopan Rattanapoka, and Damien Vouriot. Exploitation of a parallel clustering algorithm on commodity hardware with P2P-MPI. The Journal of Supercomputing, 43(1):21–41, January 2008. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

                                        //www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0920-8542&volume=43&
                                        issue=1&spage=21.

Green:2000:JC

[GH00] Marcus Green and David Hecksel. Java certification. Dr. Dobb’s Journal of Software Tools, 25(10):??, October 2000. CODEN DDJOEB. ISSN 1044-789X.

Gagnon:2001:SRF

[GH01] Etienne M. Gagnon and Laurie J. Hendren. SableVM: a research framework for the efficient execution of Java bytecode. In USENIX Association [USE01c], page ?? ISBN 1-880446-11-1. LCCN QA76.73.J38 J42 2001. URL

                                        http://www.usenix.org/publications/
                                        library/proceedings/jvm01/

gagnon.html. Sponsored by the USENIX Association.

Gagnon:2003:EIT

[GH03] E. Gagnon and L. Hendren.

Effective inline-threaded interpretation of Java bytecode using preparation sequences. Lecture Notes in Computer Science, 2622:170–184, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Geary:2004:CJF

[GH04] David M. Geary and Cay S.

Horstmann. Core JavaServer faces. Java 2 platform, enterprise edition series. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2004. ISBN 0-13-146305-5 (paperback). xvii + 658 pp. LCCN TK5105.8885.J38 G433 2004.

Geary:2007:CJF

[GH07] David M. Geary and Cay S.

Horstmann. Core JavaServer faces. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, second edition, 2007. ISBN 0-13-173886-0 (paperback). xx + 723 pp. LCCN TK5105.8885.J38 G433 2007. URL http:/

/www.loc.gov/catdir/toc/
ecip0711/2007006830.html.

Gegg-Harrison:2003:SPCa

[GHBG^{+}03a] Timothy S. Gegg-Harrison, Gary R. Bunce, Rebecca D. Ganetzky, Christina M. Olson, and Joshua D. Wilson. Studying program correctness by constructing contracts. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):129–133, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gegg-Harrison:2003:SPCb

[GHBG^{+}03b] Timothy S. Gegg-Harrison, Gary R. Bunce, Rebecca D. Ganetzky, Christina M. Olson, and Joshua D. Wilson. Studying program correctness in ProVIDE. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3): 262, September 2003. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Glitho:2001:AFU

[GHH01] R. Glitho, R. Hamadi, and R. Huie. Architectural framework for using Java servlets in a SIP environment. Lecture Notes in Computer Science, 2094:707–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.
                                        com/link/service/series/
                                        0558/bibs/2094/20940707.
                                        htm;   http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2094/20940707.
                                        pdf.

Gonzalez:2001:EDT

[GHM^{+}01] Evelio J. González, Alberto F. Hamilton, Lorenzo Moreno, José F. Sigut, and Roberto L. Marichal. Evenet 2000:

Designing and training arbitrary neural networks in Java. Lecture Notes in Computer Science, 2085:104–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2085/20850104.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2085/20850104.
pdf.

Ghosh:2001:JJT

[Gho01] Subdipto Ghosh, editor.

Java/Jini technologies: 21–22 August 2001, Denver, USA, volume 4521 of SPIE proceedings series. SPIE Optical Engineering Press, Bellingham, WA, USA, 2001. ISBN 0-8194-4245-3. LCCN QA76.9.D5 J38 2001.

Ghosh:2004:GJC

[Gho04] Debasish Ghosh. Generics in Java and C++: a comparative model. ACM SIG-PLAN Notices, 39(5):40–47, May 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Greenhouse:2005:OAE

[GHS05] Aaron Greenhouse, T. J.

Halloran, and William L. Scherlis. Observations on the assured evolution of concurrent Java programs. Science of Computer Programming, 58(3):384–411, December 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Gentleman:2000:JD

[GI00] Robert Gentleman and Ross Ihaka. Java and Dylan. Journal of Computational and Graphical Statistics, 9(3):452–??, September 2000. ISSN 1061-8600 (print), 1537-2715 (elec-

                                        tronic). URL http://www.
                                        amstat.org/publications/
                                        jcgs/abstracts00/Gentleman.
                                        htm.

Gibbons:2001:TDJ

[Gib01] Chad Gibbons. Tape devices and Java. Dr. Dobb’s Journal of Software Tools, 26 (2):44, 46, 48–50, February 2001. CODEN DDJOEB. ISSN 1044-789X. URL http:

                                        //www.ddj.com/ftp/2001/
                                        2001_02/javatape.txt;
                                        http://www.ddj.com/ftp/
                                        2001/2001_02/javatape.zip.

Gibson:2009:SRP

[Gib09] J. Paul Gibson. Software reuse and plagiarism: a code of practice. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):55–59, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Giguere:2000:JME

[Gig00] Eric Giguere. Java 2 micro edition: professional developer’s guide. Professional developer’s guide series. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 2000. ISBN 0-471-39065-8 (paperback). xxiii + 295 pp. Also available via the World Wide Web (Restricted to NetLibrary subscribers).

Gill:2000:JVJ

[Gil00a] Dudley W. Gill. Java with Visual J++. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2000. ISBN 0-8493-2048-8. 462 pp. LCCN QA76.73.J38 G56 2000. US$69.95. URL

http://www.crcpress.com/.

Gilorien:2000:DJ

[Gil00b]  Gilorien.  DHTML and

JavaScript. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2000. ISBN 0-13-086334-3. xx + 1142 pp. LCCN QA76.76.H94 G55 2000.

Gilreath:2000:RDP

[Gil00c] William Gilreath. Review:

Distributed Programming with Java. IEEE Distributed Systems Online, 1(1), 2000. ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:

//dsonline.computer.org/
archives/ds100/ds1bookprint.
htm.

Gilreath:2001:JNP

[Gil01] William F. Gilreath. Java Network Programming introduces the fundamentals. IEEE Distributed Systems Online, 2(5), 2001. ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:

                                        //dsonline.computer.org/
                                        0105/departments/bks0105b.
                                        htm.

Gittleman:2000:OCJ

[Git00] Art Gittleman. Objects to components with the Java platform. Scott/Jones, El Granada, CA, USA, 2000. ISBN 1-57676-035-9. xv + 589 pp. LCCN QA76.73.J38 G583 2000.

Gestwicki:2004:JJI

[GJ04] Paul V. Gestwicki and Bharat Jayaraman. JIVE: Java interactive visualization environment. ACM SIGPLAN Notices, 39(10):226–228, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gregersen:2009:DUJ

[GJ09] Allan Raundahl Gregersen and Bo Nørregaard Jørgensen. Dynamic update of Java applications — balancing change flexibility vs programming transparency. Journal of Software Maintenance and Evolution: Research and Practice, 21(2):

81–112, March 2009. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Gosling:2000:JLS

[GJSB00] James Gosling, Bill Joy, Guy L. Steele, and Gilad Bracha. The Java Language Specification. Java series. Addison-Wesley, Reading, MA, USA, second edition, 2000. ISBN 0-201-31008-2. xxv + 505 pp. LCCN QA76.73.J38 G68 2000. URL http://

java.sun.com/people/jag/.

Gosling:2005:JLS

[GJSB05] James Gosling, Bill Joy, Guy L. Steele, and Gilad Bracha. The Java language specification. The Java series. Addison-Wesley, Reading, MA, USA, third edition, 2005. ISBN 0-321-24678-0 (paperback). xxxii + 651 pp.

Gerlach:2003:GPS

[GK03] J. Gerlach and J. Kneis.

Generic programming for scientific computing in C++, Java, and C#. Lecture Notes in Computer Science, 2834:301–310, 2003. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Griffith:2005:MME

[GK05] Rean Griffith and Gail Kaiser.

Manipulating managed execution runtimes to support self-healing systems. ACM SIGSOFT Software Engineering Notes, 30(4):1–7, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gabay:2007:CJR

[GK07] Yarom Gabay and Assaf J. Kfoury. A calculus for Java’s reference objects. ACM SIGPLAN Notices, 42(8):9– 17, August 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ghosh:2008:BFI

[GK08] Sudipto Ghosh and John L. Kelly. Bytecode fault injection for Java software. The Journal of systems and software, 81(11):2034–2043, November 2008. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Godefroid:2008:GBW

[GKL08] Patrice Godefroid, Adam Kiezun, and Michael Y. Levin. Grammar-based whitebox fuzzing. ACM SIG-PLAN Notices, 43(6):206– 215, June 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ghaly:2001:SEA

[GKM01] Ragae Ghaly, Krishna Kothapalli, and Uma Meyyappan. Selecting EJB application servers: Benchmark and test a variety of EJB servers before making a decision. Dr. Dobb’s Journal of Software Tools, 26(9):96, 98, 100–101, September 2001. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/.

Galant:2003:HTN

[GKM03] A. Galant, R. Kutner, and A. Majerowski. Heat transfer, Newton’s law of cooling and the law of entropy increase simulated by the real-time computer experiment in Java. Lecture Notes in Computer Science, 2657:45–53, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gall:2004:BEC

[GKMZ04] M. Gall, R. Kutner, A. Majerowski, and D. Zebrowski. Bose–Einstein condensation studied by the real-time Monte Carlo simulation in the frame of Java applet. Lecture Notes in Computer Science, 3039:673–678, 2004. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gall:2004:PIC

[GKW04] M. Gall, R. Kutner, and W. Wesela. The proof and illustration of the central limit theorem by Brownian numerical experiments in real time within the Java applet. Lecture Notes in Computer Science, 3037:467–474, 2004.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goldwasser:2008:TOO

[GL08] Michael H. Goldwasser and David Letscher. Teaching an object-oriented CS1 -: with Python. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):42–46, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Glass:2006:RCP

[Gla06] Robert L. Glass. Review of COBOL Programmers Swing With Java by E. Reed Doke, Bill C. Hardgrave and Richard A. Johnson, Cambridge University Press, 2005. ACM SIGSOFT Software Engineering Notes, 31(3):44–45, May 2006. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gu:2001:JBP

[GLC01] Yan Gu, Bu-Sung Lee, and Wentong Cai. JBSP: a BSP programming library in Java. Journal of Parallel and Distributed Computing, 61(8):1126–1142, August 1, 2001. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/jpdc.
                                        2001.1735;   http://www.

idealibrary.com/links/doi/
10.1006/jpdc.2001.1735/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.
2001.1735/ref.

Gleim:2002:JPI

[Gle02]  Urs Gleim.  JaRTS: a

portable implementation of real-time core extensions for Java. In USENIX Association [USE02], page ?? ISBN 1-931971-01-3. LCCN QA76.73.J38 J42 2002. URL

http://www.usenix.org/publications/
library/proceedings/javavm02/
gleim.html.

Guha:2002:DII

[GLS02] S. Guha, B. P. Le, and I. Suzuki. Design and implementation of an interactive Java-based tutor. Simulation series, 34(1):9–12, 2002. CO-DEN SMCPAX. ISSN 0735-9276.

Griesemer:2000:CJH

[GM00] Robert Griesemer and Srdjan Mitrovic. A compiler for the Java HotSpot virtual machine. In Böszörményi et al. [BGP00], page ?? ISBN 1-55860-723-4. LCCN QA76.756 .S36 2000. US$39.95.

URL http://www.mkp.com/
books_catalog/catalog.asp?
ISBN=1-55860-723-4.

Gordon:2002:LHQ

[GM02] O. E. Gordon and T. E. Malloy. On-line Homework/ quiz/ exam applet: Freely available Java software for evaluating performance on line. Behavior Research Methods, Instruments, and Computers, 34(2): 241–244, May 1, 2002. CO-DEN BRMCEW. ISSN 0743-3808 (print), 1532-5970 (electronic).

Gruntz:2003:JST

[GM03] D. Gruntz and R. Muller. Java smart ticket demo application scrutinized. Lecture Notes in Computer Science, 2591:121–134, 2003. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gil:2005:MPJ

[GM05a] Joseph (Yossi) Gil and Itay Maman. Micro patterns in Java code. ACM SIGPLAN Notices, 40(10):97–116, October 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Guinness:2005:SMM

[GM05b] David Mc Guinness and Liam Murphy. A simulation model of a multi-server EJB system. ACM SIGSOFT Software Engineering Notes, 30 (4):1–7, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Gutterman:2005:HYS

[GM05c] Z. Gutterman and D. Malkhi. Hold your sessions: An attack on Java session-ID generation. Lecture Notes in Computer Science, 3376:44– 57, 2005. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gil:2008:WIS

[GM08] Joseph Gil and Itay Maman. Whiteoak: introducing structural typing into Java. ACM SIGPLAN Notices, 43(10):73–90, September 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gupta:2000:TSH

[GMM00] Manish Gupta, Samuel P.

Midkiff, and Jose E. Moreira. Tutorial S4: High performance numerical computing in Java: Compiler, language, and application solutions. In ACM [ACM00c], page 13.

URL http://www.sc2000.
org/proceedings/info/fp.
pdf.

Groth:2009:MPD

[GMM09] Paul Groth, Simon Miles, and Luc Moreau. A model of process documentation to determine provenance in mashups. ACM Transactions on Internet Technology (TOIT), 9(1):3:1–3:??, February 2009. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Gustedt:2002:TJP

[GMT02] Jens Gustedt, Ole A. Mæhle, and Jan Arne Telle. The treewidth of Java programs. Lecture Notes in Computer Science, 2409:86–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2409/24090086.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2409/24090086.
pdf.

Goncalves:2002:JMO

[GMW^{+}02] Marcos André Gonçalves, Paul Mather, Jun Wang, Ye Zhou, Ming Luo, Ryan Richardson, Rao Shen, Liang Xu, and Edward A. Fox. Java MARIAN: From an OPAC to a modern digital library system. Lecture Notes in Computer Science, 2476:194– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/
service/series/0558/bibs/
2476/24760194.htm;   http:
//link.springer.de/link/
service/series/0558/papers/
2476/24760194.pdf.

Gore:2001:CAM

[GN01a] Rajeev Goré and Lan Duy Nguyen. CardKt: Automated multi-modal deduction on Java cards for multiapplication security. Lecture Notes in Computer Science, 2041:38–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2041/20410038.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2041/20410038.
pdf.

Gore:2001:CMT

[GN01b] Rajeev Prabhakar Goré and Phuong Thê Nguyên. CardS4: Modal theorem proving on Java smartcards. Lecture Notes in Computer Science, 2140:111–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2140/21400111.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2140/21400111.
pdf.

Gordon:2004:C

[GNB04] Donald Gordon, James Noble, and Robert Biddle. ConstrainedJava. ACM SIG-PLAN Notices, 39(10):33– 34, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Garbervetsky:2005:PIR

[GNYZ05] D. Garbervetsky, C. Nakhli, S. Yovine, and H. Zorgati. Program instrumentation and run-time analysis of scoped memory in Java. Electronic Notes in Theoretical Computer Science, 113(??):105– 121, 2005. CODEN ???? ISSN 1571-0661.

Goeschl:2001:JTT

[Goe01] Siegfried Goeschl. The JUnit++ testing tool. Dr. Dobb’s Journal of Software Tools, 26(2):34, 36–38, February 2001. CODEN DDJOEB. ISSN 1044-789X. URL http:

                                        //www.ddj.com/ftp/2001/
                                        2001_02/junitpp.txt;   http:
                                        //www.ddj.com/ftp/2001/
                                        2001_02/junitpp.zip.

Goldstein:2000:HJC

[Gol00]  Mitch Goldstein.  Hard-

core JFC: Conquering the Swing Architecture. Cambridge University Press, Cambridge, UK, 2000. ISBN 0-521-66489-6. 356 pp. LCCN QA76.73.J38 G59 2001. US$42.95.

Goldman:2001:JQW

[Gol01] Oliver Goldman. Java Q&A: What are checked exceptions? Dr. Dobb’s Journal of Software Tools, 26 (11):111–112, 114, November 2001. CODEN DDJOEB. ISSN 1044-789X. URL http:

                                        //www.ddj.com/ftp/2001/
                                        2001_11/jqa1101.txt.

Goldman:2004:IEB

[Gol04a] K. J. Goldman. An interactive environment for beginning Java programmers. Science of Computer Programming, 53(1):3–24, 2004. CO-DEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Goldman:2004:CFI

[Gol04b] Kenneth J. Goldman. A concepts-first introduction to computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):432– 436, March 2004. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Goodwill:2000:PJJ

[Goo00] James Goodwill. Pure JSP, Java Server pages. Howard W. Sams, Indianapolis, IN 46268, USA, 2000. ISBN 0-672-31902-0. vii + 320 pp. LCCN QA76.73.J38 G664

2000. URL http://www.
mcp.com/detail.cfm?item=

0672319020. Edited by Steve Anglin and Tim Ryan.

Goodman:2001:JB

[Goo01a] Danny Goodman. JavaScript bible. Hungry Minds, Indianapolis, IN, USA, gold edition, 2001. ISBN 0-7645-4718-6 (paperback). xxxv + 1511 pp. LCCN QA76.73.J39 G65 2001b. Accompanying CD-ROM includes a searchable eversion of the book, nearly 300 ready-to-run scripts from the book, printable version of the JavaScript and Browser Object Quick Reference, the full version on WebSpice Objects, a demo of BBEdit, and TextPad shareware Contents: CD-ROM contents: nearly 300 HTML document listings from the book – Supplemental book material – Electronic version of the Quick Reference – A complete, searchable version of the book – Adobe Acrobat Reader – Thirdparty software: Microsoft Internet Explorer 5.5 for Windows – Windows and Macintosh versions of Netscape Navigator 6.0 – WebSpice Objects (full version) – TextPad (Windows only) – BBEdit (Macintosh) (demo).

Goodman:2001:JEB

[Goo01b] Danny Goodman. JavaScript examples bible: the essential companion to JavaScript bible. Hungry Minds, Indianapolis, IN, USA, 2001. ISBN 0-7645-4855-7 (paperback). xix + 525 pp. LCCN QA76.73.J39 .G653 2001. Accompanying CD-ROM includes a searchable eversion of this book and over 300 ready-to-run scripts from the text.

Goodman:2002:DHD

[Goo02a] Danny Goodman. Dynamic HTML: The Definitive Reference: a Comprehensive Resource for HTML, CSS, DOM & JavaScript. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2002. ISBN 0-596-00316-1. xiii + 1401 pp. LCCN QA76.76.H94 G657 2002. US$59.95. URL

http://www.oreilly.com/
catalog/dhtmlref2.

Goodsen:2002:EJT

[Goo02b]  John Goodsen.  Effective

Java testing strategies. Lecture Notes in Computer Science, 2418:275–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/
service/series/0558/bibs/
2418/24180275.htm;   http:
//link.springer.de/link/
service/series/0558/papers/
2418/24180275.pdf.

Goodman:2003:JDC

[Goo03a] Danny Goodman. JavaScript and DHTML cookbook. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2003. ISBN 0-596-00467-2. xvii + 520 pp. LCCN QA76.73.J39 G63 2003. URL http://

www.oreilly.com/catalog/
9780596004675.

Goody:2003:IVJ

[Goo03b] Roy W. Goody. Introduction to Visual J++ (Version 6.0): applications and applets: Java 2 compliant. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, second edition, 2003. ISBN 0-13-048260-

9. xx + 580 pp. LCCN QA76.73.J38 G662 2003.

Goodman:2007:JDC

[Goo07] Danny Goodman. JavaScript and DHTML cookbook. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2007. ISBN 0-596-51408-5 (paperback). xx + 582 pp. LCCN QA76.73.J39 G63 2007eb; QA76.73.J39 G63 2007; QA76.73.J39. URL

                                        http://www.oreilly.com/
                                        catalog/9780596514082.

Gosling:2000:JLR

[Gos00a] James Gosling. JAVA: a language for the real world,

                                        2000. URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/osdi2000/
                                        wiess2000/invitedtalks/
                                        gosling_html. Unpublished

invited talk at First Workshop on Industrial Experiences with Systems Software (WIESS 2000), October 22, 2000, Paradise Point Resort, San Diego, California, USA.

Gosselin:2000:JC

[Gos00b] Don Gosselin. JavaScript: comprehensive. Web warrior series. Course Technology, Cambridge, MA, USA, 2000. ISBN 0-619-01555-1. xvi + 710 + 46 pp. LCCN QA76.73.J39 G682 2000.

Goschl:2003:JXB

[Gös03] Siegfried Göschl. Jelly: An XML-based scripting language. Dr. Dobb’s Journal of Software Tools, 28(7): 66–67, July 2003. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2003/2003_07/jelly.
zip.

Goth:2006:NSN

[Got06] Greg Goth. News: Not in the script — news of Java’s demise is premature. IEEE Distributed Systems Online, 7 (2):??, February 2006. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
ds/2006/02/o2004.pdf.

Gourley:2001:ALB

[Gou01] Don Gourley. An Apache load balancing cluster. SysAdmin, 10(1):8, 10, 12, 14, 16, January 2001. CODEN SYADE7. ISSN 1061-2688. URL http:

//www.samag.com/.

Gousie:2006:RWP

[Gou06] Michael B. Gousie. A robust Web programming and graphics course for nonmajors. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):72–76, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Getov:2001:JCL

[GP01] V. Getov and M. Philippsen. Java communications for large-scale parallel computing. Lecture Notes in Computer Science, 2179:33–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2179/21790033.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2179/21790033.
pdf.

Ghahramani:2003:ISP

[GP03] Bahador Ghahramani and Mark A. Pauley. IT systems perspective: Java in high-performance environments. Computer, 36(9):109– 111, September 2003. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/co/
2003/09/r9109.htm;   http:
//csdl.computer.org/dl/
mags/co/2003/09/r9109.pdf.

GerthVictor:2005:JTD

[GP05] E. Gerth Victor and D. Vize Peter. A Java tool for dynamic web-based 3D visualization of anatomy and overlapping gene or protein expression patterns. Bioinformatics, 21(7):1278–1279, April 01, 2005. CODEN ???? ISSN 1367-4803 (print), 1367-4811 (electronic).

Goetz:2006:JCP

[GPB^{+}06] Brian Goetz, Tim Peierls, Joshua Bloch, Joseph Bowbeer, Doug Lea, and David Holmes. Java Concurrency in Practice. Addison-Wes-ley, Reading, MA, USA, 2006. ISBN 0-321-34960-1 (paperback). xx + 403 pp. LCCN QA76.73.J38 G588

2006. URL http://www.loc.
gov/catdir/toc/ecip0612/
2006012205.html.

Gal:2005:IJB

[GPF05] A. Gal, C. W. Probst, and M. Franz. Integrated Java bytecode verification. Electronic Notes in Theoretical Computer Science, 131:27–38, 2005. CODEN ???? ISSN 1571-0661.

Gal:2008:JBV

[GPF08] Andreas Gal, Christian W.

Probst, and Michael Franz. Java bytecode verification via static single assignment form. ACM Transactions on Programming Languages and Systems, 30(4):21:1– 21:21, July 2008. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Gontmakher:2003:CVJ

[GPS03] A. Gontmakher, S. Polyakov, and A. Schuster. Complexity of verifying Java shared memory execution. Parallel Processing Letters, 13(4):721– 734, 2003. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Gregg:2003:PID

[GPW03] David Gregg, James Power, and John Waldron. Platform independent dynamic Java virtual machine analysis: the Java Grande Forum benchmark suite. Concurrency and Computation: Practice and Experience, 15(3–5):459– 484, March/April 2003. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Gregg:2005:MLC

[GPW05] David Gregg, James Power, and John Waldron. A method-level comparison of the Java Grande and SPEC JVM98 benchmark suites. Concurrency and Computation: Practice and Experience, 17(7–8):757–773, June/July 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Genaud:2007:PMP

[GR07] Stéphane Genaud and Choopan Rattanapoka. P2P–MPI: a peer-to-peer framework for robust execution of message passing parallel programs on Grids. Journal of Grid Computing, 5(1):27–42, March 2007. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=1570-7873&volume=5&
issue=1&spage=27.

Gray:2004:JBA

[Gra04] J. Gray. A Java-based approach for teaching principles of adaptive and evolvable software. Science of Computer Programming, 53 (1):51–69, 2004. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Grissom:2000:PFI

[Gri00] Scott Grissom. A pedagogical framework for introducing Java I/O in CS1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(4):57– 59, December 2000. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Griffith:2002:JXJ

[Gri02a] Arthur Griffith. Java, XML, and JAXP. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 2002. ISBN 0-471-20907-4 (PA-PER/WEBSITE). xii + 222 pp. LCCN QA76.73.J38 G77 2002 Bar.

Grinder:2002:AAC

[Gri02b] Michael T. Grinder. Animating automata: a crossplatform program for teaching finite automata. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1): 63–67, March 2002. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Grinder:2003:PEE

[Gri03] Michael T. Grinder. A preliminary empirical evaluation of the effectiveness of a finite state automaton animator. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35 (1):157–161, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Grimm:2006:BET

[Gri06] Robert Grimm. Better extensibility through modular syntax. ACM SIGPLAN Notices, 41(6):38–51, June 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gries:2008:PAT

[Gri08]  David Gries.  A princi-

pled approach to teaching OO first. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):31–35, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Grosbol:2002:CJC

[Gro02a] P. Grosbol. A compact Java class library for FITS. Astronomical Society of the Pacific Conference Series, 281 (??):136–139, 2002. CODEN ???? ISSN 1050-3390.

Grosso:2002:JR

[Gro02b] William Grosso. Java RMI.

O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2002. ISBN 1-56592-452-5. xxiii + 545 pp. LCCN QA76.73.J38 G76 2002. Designing and building distributed applications.

Grosso:2002:JRD

[Gro02c] William Grosso. Java RMI:

Designing and building distributed applications. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2002. ISBN 1-56592-452-5. xxiii + 545 pp. LCCN QA76.73.J38 G78 2002. URL http://

www.oreilly.com/catalog/
9781565924529. Designing

and building distributed applications.

Guelfi:2005:SED

[GRR05] Nicolas Guelfi, Gianna Reggio, and Alexander Romanovsky, editors. Scientific Engineering of Distributed Java Applications: 4th International Workshop, FIDJI 2004, Luxembourg-Kirchberg, Luxembourg, November 24– 25, 2004. Revised selected papers, volume 3409 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. CODEN LNCSD9. ISBN 3-540-25053-0 (soft cover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.758 .F53 2004. URL http:

//www.springerlink.com/
openurl.asp?genre=issue&
issn=0302-9743&volume=3409;
 http://www.springerlink.
com/openurl.asp?genre=volume&
id=doi:10.1007/b106620.

Geppert:2000:T

[GS00a] L. Geppert and W. Sweet. Technology 2000. IEEE Spectrum, 37(1):26–31, January 2000. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gilreath:2000:BRJ

[GS00b] William F. Gilreath and Benjamin R. Seyfarth. Book reviews: Java Distributed Computing is solid but not flawless; Elements of Network Protocol Design: Network protocol textbook is an asset. IEEE Concurrency, 8 (2):86–87, April/June 2000. CODEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic). URL http:

//dlib.computer.org/pd/

books/pd2000/pdf/p2086.
pdf.

Gontmakher:2000:JCN

[GS00c] Alex Gontmakher and Assaf Schuster. Java consistency: nonoperational characterizations for Java memory behavior. ACM Transactions on Computer Systems, 18(4):333–386, 2000. CODEN ACSYEC. ISSN 0734-2071 (print), 1557-7333 (electronic). URL

http://www.acm.org/pubs/
articles/journals/tocs/
2000-18-4/p333-gontmakher/
p333-gontmakher.pdf;   http:
//www.acm.org/pubs/citations/
journals/tocs/2000-18-4/
p333-gontmakher/.

Garms:2001:PJS

[GS01] Jess Garms and Daniel Somerfield. Professional Java security. Wrox Press, Chicago, IL, USA, 2001. ISBN 1-86100-425-7. vii + 523 pp. LCCN QA76.73.J38 G37 2001.

Gundersen:2004:DSJ

[GS04] Geir Gundersen and Trond Steihaug. Data structures in Java for matrix computations. Concurrency and Computation: Practice and Experience, 16(8):799–815, July 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Geller:2005:TME

[GS05a] Viktor Geller and Christelle Scharff. Traditional and more “exotic” .NET languages: VB .NET, J#, C# and SML .NET. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):406, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Genaim:2005:IFA

[GS05b] S. Genaim and F. Spoto. Information flow analysis for Java bytecode. Lecture Notes in Computer Science, 3385:346–362, 2005. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gestwicki:2008:TDP

[GS08] Paul Gestwicki and Fu-Shing Sun. Teaching design patterns through computer game development. ACM Journal on Educational Resources in Computing (JERIC), 8(1): 2:1–2:??, March 2008. CO-DEN ???? ISSN 1531-4278.

Griffin:2005:EEG

[GSaC05] Paul Griffin, Witawas Srisaan, and J. Morris Chang. An energy efficient garbage collector for Java embedded devices. ACM SIGPLAN Notices, 40(7):230–238, July 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Govindaraju:2000:RER

[GSC^{+}00] Madhusudhan Govindaraju, Aleksander Slominski, Venkatesh Choppella, Randall Bramley, and Dennis Gannon. Requirements for and evaluation of RMI protocols for scientific computing. In ACM [ACM00c], page 76.

URL http://www.sc2000.
org/proceedings/techpapr/
papers/pap261.pdf.

Goh:2006:DBM

[GSHO06] Leng Goh, Yanfeng Shu, Zhiyong Huang, and Chin Ooi. Dynamic buffer management with extensible replacement policies. VLDB Journal: Very Large Data Bases, 15(2):99–120, June 2006. CO-DEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Gsoedl:2000:JQC

[Gso00] Jacob Gsoedl. Java Q&A:

Can you implement COM components using Java? Dr. Dobb’s Journal of Software Tools, 25(6):119–120, 122, 124, 126, June 2000. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2000/2000_06/jqa600.
txt.

Grigorenko:2005:VTG

[GST05] Pavel Grigorenko, Ando Saabas, and Enn Tyugu. Visual tool for generative programming. ACM SIGSOFT Software Engineering Notes, 30(5):249–252, September 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Glossner:2002:JED

[GSV02] C. John Glossner, Michael Schulte, and Stamatis Vassiliadis. A Java-enabled DSP. Lecture Notes in Computer Science, 2268:307–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2268/22680307.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2268/22680307.
pdf;   http://mesa.ece.
wisc.edu/publications/cp_
2002-01.pdf.

Gurevich:2000:IJC

[GSW00] Yuri Gurevich, Wolfram Schulte, and Charles Wallace. Investigating Java concurrency using abstract state machines. Lecture Notes in Computer Science, 1912:151– ??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/1912/19120151.
htm;   http://link.springer-
ny.com/link/service/series/

0558/papers/1912/19120151.
pdf.

Gardner:2008:LHR

[GSWZ08] Philippa A. Gardner, Gareth D.

Smith, Mark J. Wheelhouse, and Uri D. Zarfaty. Local Hoare reasoning about DOM. In Lenzerini and Lembo [LL08a], pages 261– 270. ISBN 1-60560-932-3. LCCN ????

Goodrich:1997:DSA

[GT97] Michael T. Goodrich and Roberto Tamassia. Data Structures and Algorithms in Java. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 1997. ISBN 0-471-19308-9 (paperback). xviii + 738 pp. LCCN QA76.73.J38 G66

1997. URL http://www.loc.
gov/catdir/enhancements/
fy0706/97039794-d.html;
http://www.loc.gov/catdir/
toc/onix01/97039794.html.

Gottleber:2000:MEH

[GT00] Timothy T. Gottleber and Timothy N. Trainor. More excellent HTML with an introduction to JavaScript. Irwin/McGraw Hill, Boston, MA, USA, 2000. ISBN 0-07-233745-1, 0-07-233839-3. xxiii + 550 pp. LCCN QA76.76.H94G69 2000.

Goodrich:2001:DSA

[GT01] Michael T. Goodrich and Roberto Tamassia. Data structures and algorithms in Java. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, second edition, 2001. ISBN 0-471-38367-8 (cloth). xiii + 641 pp. LCCN QA76.76.J38 G66 2001.

Goodrich:2004:DSA

[GT04] Michael T. Goodrich and Roberto Tamassia. Data Structures and Algorithms in Java. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, third edition, 2004. ISBN 0-471-46983-1 (hardcover). xvii + 681 pp. LCCN QA76.73.J38 G66 2004. URL http://

www.loc.gov/catdir/bios/
wiley046/2003071070.html;
 http://www.loc.gov/catdir/
description/wiley041/2003071070.
html;   http://www.loc.
gov/catdir/toc/wiley041/
2003071070.html.

Gehtland:2005:SDN

[GT05] Justin Gehtland and Bruce A. Tate. Spring: a developer’s notebook. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2005. ISBN 0-596-00910-0. xv + 184 pp. LCCN QA76.73.S67 T38 2005.

Goodrich:2006:DSA

[GT06] Michael T. Goodrich and Roberto Tamassia. Data Structures and Algorithms in Java. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, fourth edition, 2006. ISBN 0-471-73884-0. xxii + 696 pp. LCCN QA76.73.J38 G66

2006. URL http://www.loc.
gov/catdir/enhancements/
fy0627/2005282681-b.html;
 http://www.loc.gov/catdir/
enhancements/fy0627/2005282681-
d.html;   http://www.loc.
gov/catdir/enhancements/
fy0627/2005282681-t.html.

Goodrich:2010:DSA

[GT10] Michael T. Goodrich and Roberto Tamassia. Data Structures and Algorithms in Java. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, fifth edition, 2010. ISBN 0-470-38326-7 (hardcover), 0-470-39880-9 (paperback). xxii + 714 pp. LCCN QA76.73.J38 G66 2010.

Guha:2007:CIF

[Guh07] Rajarshi Guha. Chemical informatics functionality in R. Journal of Statistical Software, 18(5):1–16, January 2007. CODEN JSSOBK. ISSN 1548-7660. URL http:

//www.jstatsoft.org/v18/
i05.

Guizzo:2008:GFG

[Gui08] Erico Guizzo. The gameframe guild. IEEE Spectrum, 45(8):44–52, August 2008. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gunton:2001:SSD

[Gun01] Neil Gunton. SOAP: Simplifying distributed development: Putting the Simple Object Access Protocol to work. Dr. Dobb’s Journal of Software Tools, 26 (9):89, 91–92, 95, September 2001. CODEN DDJOEB. ISSN 1044-789X. URL http:

                                        //www.ddj.com/ftp/2001/
                                        2001_09/soap.txt.

Gutz:2000:SSU

[Gut00] Steven Gutz. Up to Speed with Swing: User Interfaces with Java Foundation Classes. Manning Publications, Greenwich, CT, USA, second edition, 2000. ISBN 1-884777-75-9. xxxvi + 536 pp. LCCN QA76.9.U83 G88 2000.

                                        URL http://www.manning.
                                        com/gutz2.

Groce:2002:HMC

[GV02a] Alex Groce and Willem Visser. Heuristic model checking for Java programs. Lecture Notes in Computer Science, 2318:242–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.
                                        com/link/service/series/
                                        0558/bibs/2318/23180242.
                                        htm;   http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2318/23180242.
                                        pdf.

Groce:2002:MCJ

[GV02b] Alex Groce and Willem Visser. Model checking Java programs using structural heuristics. ACM SIGSOFT Software Engineering Notes, 27(4):12–21, July 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Groce:2004:HMC

[GV04] Alex Groce and Willem Visser. Heuristics for model checking Java programs. International Journal on Software Tools for Technology Transfer (STTT), 6 (4):260–276, August 2004. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=1433-2779&volume=6&
issue=4&spage=260.

Gerth:2005:JTD

[GV05] V. E. Gerth and P. D. Vize.

A Java tool for dynamic Web-based 3D visualization of anatomy and overlapping gene or protein expression patterns. Bioinformatics, 21 (7):1278–1279, 2005. CODEN ???? ISSN 1367-4803 (print), 1367-4811 (electronic).

Getov:2001:MCJ

[GvLPF01] Vladimir Getov, Gregor von Laszewski, Michael Philippsen, and Ian Foster. Multiparadigm communications in Java for Grid computing. Communications of the ACM, 44(10):118–125, October 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gourley:2000:BWB

[GW00] Don Gourley and Wei Wu. Building a Web-based Java application server with Apache JServ. SysAdmin, 9(2):59– 65, February 2000. CO-DEN SYADE7. ISSN 1061-

2688. URL http://www.
samag.com/.

Guo:2001:DDS

[GW01] Yike Guo and Patrick Wendel. Developing a distributed scalable Java component server. Future Generation Computer Systems, 17(8): 1051–1057, June 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/35/40/abstract.
html.

Gilliam:2002:PJ

[GW02] Jason Gilliam and R. Allen Wyke. Pure JavaScript. Howard W. Sams, Indianapolis, IN 46268, USA, second edition, 2002. ISBN 0-672-32141-6. liii + 1545 pp. LCCN QA76.73.J39 P87 2002. Rev. ed. of: Pure JavaScript / Jason Gilliam, c1999.

Gebotys:2008:EAW

[GW08] Catherine H. Gebotys and Brian A. White. EM analysis of a wireless Java-based PDA. ACM Transactions on Embedded Computing Systems, 7(4):44:1–44:??, July 2008. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Habibi:2004:JRE

[Hab04] Mehran Habibi. Java Regular Expressions: Taming the java.util.regex Engine. Apress, Berkeley, CA, USA, 2004. ISBN 1-59059-107-0. 255 (est.) pp. LCCN ????

Hachiya:2001:JUM

[Hac01] Shouichi Hachiya. Java use in mobile information devices: Introducing JTRON. IEEE Micro, 21(4):16–21, July/August 2001. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//dlib.computer.org/mi/
books/mi2001/pdf/m4016.
pdf;   m4016abs.htm.

Hagan:2000:UBT

[Hag00a] Dianne Hagan. Using BlueJ to teach Java (poster session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32 (3):188–189, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Haggar:2000:PJP

[Hag00b] Peter Haggar. Practical Java: programming language guide. Addison-Wesley professional computing series. Addison-Wesley, Reading, MA, USA, 2000. ISBN 0-201-61646-

7. xxx + 279 pp. LCCN QA76.73.J38 H34 2000.

Haggar:2002:JQD

[Hag02] Peter Haggar. Java Q&A: Does Java guarantee thread safety? Dr. Dobb’s Journal of Software Tools, 27 (6):91–83, June 2002. CO-DEN DDJOEB. ISSN 1044-789X. URL http:

                                        //www.ddj.com/ftp/2002/
                                        2002_06/jqa0602.txt. Com-

ments on lack of atomicupdate guarantee in Java for objects larger than 32 bits,

                                        such as long and double,

with sample code to exhibit the failure.

Hall:2000:CSJ

[Hal00] Marty Hall. Core Servlets and JavaServer Pages. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2000. ISBN 0-13-089340-4. xxvii + 575 pp. LCCN QA76.73.J38 H345 2000. US$42.99. URL http:

                                        //www.coreservlets.com/;
                                         http://www.phptr.com/ptrbooks/
                                        ptr_0130893404.html.

Hall:2001:MHC

[Hal01a] Marty Hall. Marty Hall’s Core Servlets and JavaServer Pages Training Course: a digital seminar on CD-ROM. Sun BluePrints Program. Sun Microsystems Press, Palo Alto, CA, USA, 2001. ISBN 0-13-093400-3. 308 pp. LCCN ????

Halter:2001:JEE

[Hal01b] Steven L. Halter. JavaSpaces Example by Example. Sun BluePrints Program. Sun Microsystems Press, Palo Alto, CA, USA, 2001. ISBN 0-13-061916-7. xiv + 272 pp. LCCN QA76.76.J38H35

2001. URL http://www.sun.
com/books/catalog/halter2/
index.html.

Hall:2002:MSJ

[Hal02a] Marty Hall. More Servlets and JavaServer Pages. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2002. ISBN 0-13-067614-4. 752 (est.) pp. LCCN QA76.73.J38 H3455 2001. US$49.99. URL http:/

/www.phptr.com/ptrbooks/
ptr_0130676144.html.

Halloway:2002:CDJ

[Hal02b] Stuart Dabbs Halloway. Component development for the Java platform. Develop-Mentor series. Addison-Wes-ley, Reading, MA, USA, 2002. ISBN 0-201-75306-5. xx + 334 pp. LCCN QA76.73.J38 H346 2002.

Harkey:2002:WJP

[HAL02c] Dan Harkey, Shan Appajodu, and Mike Larkin. Wireless Java programming for enterprise applications: mobile devices go corporate. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 2002. ISBN 0-471-21878-2. xxiv + 690 pp. LCCN TK5103.2 .H369 2002.

Halloway:2009:PC

[Hal09] Stuart Dabbs Halloway. Programming Clojure. The pragmatic programmer. Pragmatic Bookshelf, Raleigh, NC, USA, 2009. ISBN 1-934356-33-6 (paperback). xxi + 280 pp. LCCN QA76.62 .H35 2009.

Hammond:2002:PLJ

[Ham02] Michael (Michael T.) Hammond. Programming for linguists: Java technology for language researchers. Blackwell Publishers, Malden, MA, USA, 2002. ISBN 0-631-23041-6, 0-631-23042-4 (paperback). xi + 209 pp. LCCN P98 .H34 2002.

Hamada:2007:WBT

[Ham07] Mohamed Hamada. Webbased tools for active learning in information theory. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39 (1):60–64, March 2007. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hanegan:2001:CCS

[Han01] Kevin Hanegan. Custom CGI scripting with Perl. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 2001. ISBN 0-471-01379-X (e-book), 0-471-39597-8. xxvii + 276 pp. LCCN QA76.73.P22 H36 2001eb.

Han:2005:RCK

[Han05a] Y. Han. Representing configuration knowledge for portand-connector model with Java. Journal of Computer Aided Design and Computer Graphics, 17(3):511– 516, 2005. CODEN ???? ISSN 1003-9775.

Hansen:2005:IJP

[Han05b] S. Hansen. Interpreting Java program runtimes. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):36–40, 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hapner:2002:JMS

[Hap02] Mark Hapner. Java message service API tutorial and reference: messaging for the J2EE platform. Java series. Addison-Wesley, Reading, MA, USA, 2002. ISBN 0-201-78472-6 (paperback). xxi + 510 pp. LCCN QA76.73.J38 J3675 2002 Stacks SEMSTAX-NBKS.

Hardin:2000:RTS

[Har00a] David Hardin. The realtime specification for Java. Dr. Dobb’s Journal of Software Tools, 25(2):78, 80–82, 84, February 2000. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2000/2000_02/rtjava.
txt.

Hardy:2000:JAG

[Har00b]  Vincent J. Hardy.  Java

2D API graphics. Sun Microsystems Press Java series. Sun Microsystems Press, Palo Alto, CA, USA, 2000. ISBN 0-13-014266-2. xli + 509 pp. LCCN T385 .H334 2000.

URL http://www.sun.com/
books/catalog/hardy/. In-

cludes CD-ROM.

Harold:2000:JNP

[Har00c] Elliotte Rusty Harold. Java network programming. Java series. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2000. ISBN 1-56592-870-9. xxvi + 731 pp. LCCN QA76.625 .H367 2000.

Harrison:2000:DWP

[Har00d] Graham (Graham Paul) Harrison. Dynamic Web programming: using Java, JavaScript, and Informix. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2000. ISBN 0-13-086184-7. xviii + 995 pp. LCCN QA76.625.H37 2000.

Hartley:2000:AYM

[Har00e] Stephen J. Hartley. “Alfonse, you have a message!”. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32 (1):60–64, March 2000. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Harms:2001:JSM

[Har01a] David (David Gerhard) Harms.

JSP, servlets, and MySQL. M&T Books, M&T Publishing, Inc., 501 Galveston Drive, Redwood City, CA 94063, USA, 2001. ISBN 0-7645-4787-9. xxiv + 499 pp. LCCN QA76.73.J38 H365 2001. URL ftp:/

/uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
bios/wiley045/2001016954.
html;   http://www.loc.
gov/catdir/description/
wiley038/2001016954.html;
 http://www.loc.gov/catdir/
toc/wiley021/2001016954.
html.

Hartley:2001:AGM

[Har01b] Stephen J. Hartley. “Alfonse, give me a call!”. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):229– 232, March 2001. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Harold:2002:XCB

[Har02] Elliotte Rusty Harold, editor. The XML CD bookshelf: XML in a nutshell. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, version 1.0. edition, 2002. ISBN 0-596-00335-8. 600 (est.) pp. LCCN QA76.76. US$119.95. URL

                                        http://www.oreilly.com/
                                        catalog/9780596003357;
                                        http://www.oreilly.com/
                                        catalog/xmlcdbksf. One

CD-ROM.

Harold:2003:PXJ

[Har03] Elliotte Rusty Harold. Processing XML with Java: a guide to SAX, DOM, JDOM, JAXP, and TrAX. Addison-Wesley, Reading, MA, USA, 2003. ISBN 0-201-77186-1. xxxvii + 1071 pp. LCCN QA76.76.H94 H337 2003 Bar.

Harold:2004:JNP

[Har04] Elliotte Rusty Harold. Java network programming. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, third edition, 2004. ISBN 0-596-00721-3. xxii + 735 pp. LCCN QA76.625; QA76.625 .H367 2004eb; QA76.625 .H367 2004; QA76.625 .H367 2005.

                                        URL http://www.oreilly.
                                        com/catalog/9780596007218.

Harold:2006:J

[Har06] Elliotte Rusty Harold. Java I/O. The Java series. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2006. ISBN 0-596-52750-0. 726 (est.) pp. LCCN ???? US$49.99. URL http://

www.oreilly.com/catalog/
javaio2/.

Hassler:2002:JCP

[Has02] Vesna Hassler. Java card for e-payment applications. Artech House computer security series. Artech House Inc., Norwood, MA, USA, 2002. ISBN 1-58053-291-8. xvii + 362 pp. LCCN QA76.73.J38 J3638 2002.

Hawlitzek:2002:J

[Haw02] Florian Hawlitzek. Java 2.

Addison-Wesley nitty gritty programming series. Addison-Wesley, Reading, MA, USA, 2002. ISBN 0-201-75880-6. 382 pp. LCCN QA76.73.J38 H3913 2002.

Hall:2001:CWP

[HB01] Marty Hall and Larry Brown.

Core Web programming. Sun Microsystems Press Java series. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2001. ISBN 0-13-089793-0. xli + 1398 pp. LCCN QA76.625 .H35

2001. URL http://www.sun.
com/books/catalog/hall2/
index.html.

Hulaas:2008:PTL

[HB08] Jarle Hulaas and Walter Binder. Program transformations for light-weight CPU accounting and control in the Java Virtual Machine. Higher-Order and Symbolic Computation, 21 (1–2):119–146, June 2008. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=1388-3690&volume=21&
issue=1&spage=119.

Hanks:2009:SUP

[HB09] Brian Hanks and Matt Brandt. Successful and unsuccessful problem solving approaches of novice programmers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):24–28, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Hulaas:2004:EJG

[HBD04] J. Hulaas, W. Binder, and G. DiMarzoSerugendo. Enhancing Java Grid computing security with resource control. Lecture Notes in Computer Science, 3270:30– 47, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hubbard:2001:PJB

[HBH01] J. R. (John Rast) Hubbard, Anthony Q. Baxter, and J. R. Hubbard. Programming with Java: based on Schaum’s outline of programming with Java. Schaum’s outline series. McGraw-Hill, New York, NY, USA, 2001. ISBN 0-07-136977-5 (paperback). v + 162 pp. LCCN QA76.76.J38 H817 2001.

Hertz:2002:EFG

[HBM^{+}02] Matthew Hertz, Stephen M.

Blackburn, J. Eliot B. Moss, Kathryn S. McKinley, and Darko Stefanović. Errorfree garbage collection traces: how to cheat and not get caught. ACM SIGMETRICS Perform. Eval. Rev., 30(1): 140–151, June 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hertz:2006:GOL

[HBM^{+}06] Matthew Hertz, Stephen M.

Blackburn, J. Eliot B. Moss, Kathryn S. McKinley, and Darko Stefanović. Generating object lifetime traces with Merlin. ACM Transactions on Programming Languages and Systems, 28(3): 476–516, May 2006. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Harrison:2000:MUD

[HBR00] William Harrison, Charles Barton, and Mukund Raghavachari. Mapping UML designs to Java. ACM SIGPLAN Notices, 35(10):178–187, October 2000. CODEN SINODQ.

ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        proceedings/oops/353171/
                                        p178-harrison/.

Huang:2004:MIV

[HBX^{+}04] S. Huang, R. Baimouratov, P. Xiao, A. Ananthasubramaniam, and W. L. Nowinski. A medical imaging and visualization toolkit in Java [5367-05]. Proceedings of the SPIE — The International Society for Optical Engineering, 5367:48–58, 2004. CODEN PSISDG. ISSN 0277-786X (print), 1996-756X (electronic).

Horstmann:2000:CJV

[HC00] Cay Horstmann and Gary Cornell. Core Java 2, Volume I: Fundamentals. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, fifth edition, 2000. ISBN 0-13-089468-0. 750 (est.) pp. LCCN QA76.73.J38 H6753 2001. US$44.99. URL http:/

                                        /www.phptr.com/ptrbooks/
                                        ptr_0130894680.html.

Horstmann:2001:CJ

[HC01a] Cay S. Horstmann and Gary Cornell. Core Java 2. Sun Microsystems Press Java series. Sun Microsystems Press, Palo Alto, CA, USA, fifth edition, 2001. ISBN 0-13-089468-0 (vol. 1). ???? pp. LCCN QA76.73.J38 H6753 2001.

URL http://www.sun.com/
books/catalog/horstmann8/

index.html. Includes CD-ROM with complete source code examples, J2SE Version 1.3, Forte for Java Community Edition (a complete toolkit for Java application development), and trial versions of other useful tools.

Hunter:2001:JSP

[HC01b] Jason Hunter and William Crawford. Java servlet programming. Java series. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2001. ISBN 0-596-00040-5. xxiii + 753 pp. LCCN QA76.73.J38 H87 2001.

Horstmann:2002:CJV

[HC02] Cay S. Horstmann and Gary Cornell. Core Java 2, Volume II: Advanced Features. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, fifth edition, 2002. ISBN 0-13-092738-4. 1232 (est.) pp. LCCN ???? US$49.99. URL http://www.

phptr.com/ptrbooks/ptr_
0130927384.html;   http:
//www.sun.com/books/catalog/
horstmann9/index.html.

Horstmann:2003:CJV

[HC03] Cay S. Horstmann and Gary Cornell. Core Java 2: Volume 1: Fundamentals. Sun BluePrints Program. Sun Microsystems Press, Palo Alto, CA, USA, sixth edition, 2003.

ISBN 0-13-047177-1. 751 (est.) pp. LCCN QA76.73.J38 H67 2003.

Hendrix:2004:EFP

[HCB04a] T. Dean Hendrix, James H. Cross II, and Larry A. Barowski. An extensible framework for providing dynamic data structure visualizations in a lightweight IDE. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36 (1):387–391, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Huet:2004:HPJ

[HCB04b] Fabrice Huet, Denis Caromel, and Henri E. Bal. A high performance Java middleware with a real application. In ACM [ACM04], page 2. ISBN 0-7695-2153-3. LCCN ????

Hendrix:2000:DVI

[HCMM00] T. Dean Hendrix, James H. Cross II, Saeed Maghsoodloo, and Matthew L. McKinney. Do visualizations improve program comprehensibility? experiments with control structure diagrams for Java. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):382–386, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hatcliff:2001:UBT

[HD01] John Hatcliff and Matthew Dwyer. Using the bandera tool set to model-check properties of concurrent Java software. Lecture Notes in Computer Science, 2154:39–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2154/21540039.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2154/21540039.
pdf.

Hagimont:2002:NFC

[HD02] Daniel Hagimont and Noël De Palma. Non-functional capability-based access control in the Java environment. Lecture Notes in Computer Science, 2425:323–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2425/24250323.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2425/24250323.
pdf.

Henkel:2003:DAS

[HD03a] J. Henkel and A. Diwan.

Discovering algebraic specifications from Java classes. Lecture Notes in Computer Science, 2743:431–456, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hong:2003:RDW

[HD03b] Z. Y. Hong and A. Dong. Research and development of WebGIS based on MapXtreme/Java. Journal — Shenyang Institute of Technology, 22(3):45–48, 2003. CODEN ???? ISSN 1003-1251.

Husted:2003:SAB

[HD03c] Ted Husted and Cedric Dumoulin. Struts in action: Building Web applications with the leading Java framework. Manning, Greenwich, CT, USA, 2003. ISBN 1-930110-50-2. xxxiv + 630 pp. LCCN TK5105.8885.S76 H87

                                        2003. URL http://www.loc.
                                        gov/catdir/bios/ipg051/
                                        2003275314.html;   http:
                                        //www.loc.gov/catdir/description/
                                        ipg051/2003275314.html.

Hartel:2001:PMP

[HdJ01] Pieter H. Hartel and Eduard de Jong. A programming and a modelling perspective on the evaluation of Java card implementations. Lecture Notes in Computer Science, 2041:52–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.
                                        com/link/service/series/
                                        0558/bibs/2041/20410052.
                                        htm;   http://link.springer-

ny.com/link/service/series/
0558/papers/2041/20410052.
pdf.

HuertaYero:2005:JIJ

[HdS^{+}05] E. J. Huerta Yero, F. de Oliveira Lucchese, F. S. Sambatti, M. von Zuben, and M. A. Amaral Henriques. JoiN: The implementation of a Java-based massively parallel grid. Future Generation Computer Systems, 21 (5):791–810, 2005. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Hoepner:2003:JBO

[HE03] P. Hoepner and K. P. Eckert. Java-based open platform for distributed health telematics applications. Studies in Health Technology and Informatics, 96:73–79, 2003. CO-DEN ???? ISSN 0926-9630.

Heckler:2007:BRB

[Hec07] Charles E. Heckler. Book review: Correspondence Analysis and Data Coding with Java and R by Fionn Murtagh. Technometrics, 49(1):100, February 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http:

//www.jstor.org/stable/
25471279.

Hadharan:2000:EEP

[HECR00] R. Hadharan, W. K. Ehrlich, D. Cura, and P. K. Reeser.

End to end performance modeling of Web server architectures. ACM SIGMETRICS Perform. Eval. Rev., 28(2): 57–63, September 2000. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Heffelfinger:2007:JED

[Hef07] David R. Heffelfinger. Java EE 5 development using Glassfish application server. Packt, Birmingham, UK, 2007. ISBN 1-84719-260-2. 408 pp. LCCN ????

Heijl:2001:DXS

[Hei01]  Danny Heijl.  The Del-

phi XML SAX2 component and MSXML 3.0. Dr. Dobb’s Journal of Software Tools, 26(9):42, 46, 48, 50, 52, 54, September 2001. CODEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/ftp/2001/
2001_09/xmlsax2.txt;   http:
//www.ddj.com/ftp/2001/
2001_09/xmlsax2.zip. See

correction [TEM^{+}01].

Heines:2003:EXS

[Hei03a]  J. M. Heines.  Enabling

XML storage from Java applets in a GUI programming course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(2):88–93, 2003. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Heinlein:2003:ATS

[Hei03b] C. Heinlein. Advanced thread synchronization in Java using interaction expressions. Lecture Notes in Computer Science, 2591:345–365, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoffman:2009:SAT

[HEJ09] Kevin J. Hoffman, Patrick Eugster, and Suresh Jagannathan. Semantics-aware trace analysis. ACM SIG-PLAN Notices, 44(6):453– 464, June 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Helmick:2007:IOC

[Hel07a] Michael T. Helmick. Integrated online courseware for computer science courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39 (3):146–150, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Helmick:2007:IBP

[Hel07b] Michael T. Helmick. Interfacebased programming assignments and automatic grading of Java programs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):63– 67, September 2007. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Hepper:2004:JPS

[Hep04] S. Hepper. The Java portlet specification. Die Java Portlet Spezifikation. IT, 46(5): 233–244, 2004. CODEN ???? ISSN 1611-2776.

Hassler:2000:OFA

[HF00] Vesna Hassler and Oliver Fodor. OpenCard Framework application development. Dr. Dobb’s Journal of Software Tools, 25 (2):70, 72, 74–76, February 2000. CODEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/ftp/2000/
2000_02/ocfjava.txt;   http:
//www.ddj.com/ftp/2000/
2000_02/ocfjava.zip.

Harrison:2006:MSP

[HF06] Guy Harrison and Steven Feuerstein. MySQL stored procedure programming: building high-performance web applications with PHP, Perl, Python, Java & .NET. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2006. ISBN 0-596-10089-2. xxiii + 609 pp. LCCN QA76.73.S67 H377 2006eb; QA76.73.S67. URL

http://www.oreilly.com/
catalog/9780596100896.

Hau:2003:SJA

[HFL03] Gary K. W. Hau, Anthony Fong, and Mok Pak Lun. Support of Java API for the jHISC system. ACM SIGARCH Computer Architecture News, 31(4):12– 17, September 2003. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Halloway:2007:RJD

[HG07a] Stuart Dabbs Halloway and Justin Gehtland. Rails for Java developers. Pragmatic programmers; Facets of Ruby series. Pragmatic Bookshelf, Raleigh, NC, USA, 2007. ISBN 0-9776166-9-X. xix + 311 pp. LCCN QA76.73.R83 H35 2007. URL http:/

/www.loc.gov/catdir/toc/
fy0709/2007273022.html;
http://www.oreilly.com/
catalog/9780977616695.

Hirzel:2007:JGJ

[HG07b] Martin Hirzel and Robert Grimm. Jeannie: granting Java Native Interface developers their wishes. ACM SIG-PLAN Notices, 42(10):19– 38, October 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Haase:2008:FRC

[HG08] Chet Haase and Romain Guy. Filthy rich clients: developing animated and graphical effects for desktop Java applications. The Java series. Addison-Wesley, Reading, MA, USA, 2008. ISBN 0-13-241393-0 (paperback). xxvii + 572 pp. LCCN QA76.73.C153 H33

2008. URL http://www.loc.
gov/catdir/toc/ecip0717/
2007019818.html.

Hakala:2001:GAD

[HHK^{+}01] Markku Hakala, Juha Hautamäki, Kai Koskimies, Jukka Paakki, Antti Viljamaa, and Jukka Viljamaa. Generating application development environments for Java frameworks. Lecture Notes in Computer Science, 2186:163–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2186/21860163.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2186/21860163.
pdf.

Hakala:2003:GPB

[HHKS03] M. Hakala, J. Hautamaki, K. Koskimies, and P. Savolainen. Generating pattern-based Web tutorials for Java frameworks. Lecture Notes in Computer Science, 2604:99–110, 2003. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic).

Harder:2004:JUV

[HHM04] R. W. Harder, R. R. Hill, and J. T. Moore. A Java universal vehicle router for routing unmanned aerial vehicles. International Transactions in Operational Research, 11(3): 259–275, May 2004. CODEN ITORF9. ISSN 0969-6016.

Higuera:2004:MMR

[HIBP04] T. Higuera, V. Issarny, M. Banatre, and F. Parain. Memory management for real-time Java: An efficient solution using hardware support. Real-Time Systems, 26(1):63–87, 2004. CODEN RESYE9. ISSN 0922-6443.

Hightower:2003:PPJ

[Hig03] Richard Hightower. Python programming with the Java class libraries: a tutorial for building Web and Enterprise applications with Jython. Addison-Wesley, Reading, MA, USA, 2003. ISBN 0-201-61616-5. xii + 620 pp. LCCN QA76.73.P98 H54 2003.

HigueraToledano:2004:SBS

[Hig04] M. T. HigueraToledano.

Studying the behaviour of the single parent rule in realtime Java. Lecture Notes in Computer Science, 3292:268– 277, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hinke:2002:ICS

[Hin02] John Hinke. Implementing C++ servlet containers. Dr. Dobb’s Journal of Software Tools, 27(4):52, 54–56, April 2002. CODEN DDJOEB. ISSN 1044-789X. URL http:

                                        //www.ddj.com/ftp/2002/
                                        2002_04/cppserv.txt.

Hirsch:2000:CJI

[Hir00] Michael Hirsch. Comparing Java implementations for Linux. Linux Journal, 76: ??, August 2000. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Hirzel:2007:DLO

[Hir07] Martin Hirzel. Data layouts for object-oriented programs. ACM SIGMETRICS Perform. Eval. Rev., 35(1): 265–276, June 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Hitchens:2002:JN

[Hit02] Ron Hitchens. Java NIO. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2002. ISBN 0-596-00288-2. xvii + 282 pp. LCCN QA76.73.J38 H58 2002. US$34.95. URL http:/

                                        /safari.oreilly.com/0596002882;
                                         http://www.oreilly.com/
                                        catalog/javanio.

Hitzer:2003:KIS

[Hit03] E. M. S. Hitzer. Kamiwaai: Interactive 3D sketching with Java based on Cl(4, 1), conformal model of Euclidean space. Advances in Applied Clifford Algebras, 13(1):11– 46, 2003. CODEN ???? ISSN 0188-7009.

Huisman:2000:JPV

[HJ00] Marieke Huisman and Bart Jacobs. Java program verification via a Hoare logic with abrupt termination. Lecture Notes in Computer Science, 1783:284–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/1783/17830284.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/1783/17830284.
pdf.

Holmes:2001:OOP

[HJ01] Barry (Barry J.) Holmes and Daniel T. Joyce. Objectoriented programming with Java. Jones and Bartlett, Boston, MA, USA, second edition, 2001. ISBN 0-7637-1435-6. xviii + 826 pp. LCCN QA76.73.J38 H62 2001.

Hobona:2006:WBV

[HJF06] Gobe Hobona, Philip James, and David Fairbairn. Webbased visualization of 3D geospatial data using Java3D. IEEE Computer Graphics and Applications, 26(4):28– 33, July/August 2006. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Hansen:2000:KTL

[HJL00] Ole Kristian Hansen, Fredrik Johnsen, and Inger Helene Lund. Klient… tjener løsning på web, basert på Apache og MySQL ved hjelp av Java server programmering: Support system, volume 222 of Prosjektarbeid / Høgskolen i Hedmark. HHE, Rena, Norway, 2000. 94–?? pp. LCCN ????

Harrold:2001:RTS

[HJL^{+}01] Mary Jean Harrold, James A. Jones, Tongyu Li, Donglin Liang, and Ashish Gujarathi. Regression test selection for Java software. ACM SIG-PLAN Notices, 36(11):312– 326, November 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Proceedings of the 2001 ACM SIGPLAN Conference on Object Oriented Programming, Systems, Languages and Applications (OOPSLA’01).

Hericko:2003:OSA

[HJR^{+}03] Marjan Hericko, Matjaz B. Juric, Ivan Rozman, Simon Beloglavec, and Ales Zivkovic. Object serialization analysis and comparison in Java and .NET. ACM SIGPLAN Notices, 38(8):44– 54, August 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Huisman:2001:CSC

[HJvdB01] Marieke Huisman, Bart Jacobs, and Joachim van den Berg. A case study in class library verification: Java’s vector class. International Journal on Software Tools for Technology Transfer (STTT), 3(3):332–352, August 2001. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Hammouda:2002:PBJ

[HK02a] I. Hammouda and K. Koskimies.

A pattern-based J2EE application development environment. Nordic Journal of Computing, 9(3):248–??, Fall 2002. CODEN NJCOFR. ISSN 1236-6064. Selected papers of the Tenth Nordic Workshop on Programming Environment Research (NW-PER’2002), August 18–20, 2002.

Hannemann:2002:DPI

[HK02b] Jan Hannemann and Gregor Kiczales. Design pattern implementation in Java and AspectJ. ACM SIGPLAN Notices, 37(11):161–173, November 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hosny:2000:IJB

[HKF00] H. Hosny, O. Khaled, and M. E. Fathalla. ILE: a Javabased environment for CS courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):152–155, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hirayama:2003:FBE

[HKHK03] K. Hirayama, N. Kunieda, Y. Hayashi, and M. Koshiba. FEM-Based electromagnetic wave simulator running on some platforms by use of Java and a commercial tool. IEICE Transactions on Electronics E Series C, 86(11):2191–2198, 2003. CODEN ???? ISSN 0916-8524.

Higo:2008:MBA

[HKI08] Yoshiki Higo, Shinji Kusumoto, and Katsuro Inoue. A metricbased approach to identifying refactoring opportunities for merging code clones in a Java software system. Journal of Software Maintenance and Evolution: Research and Practice, 20(6): 435–461, November 2008. CODEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Harf:2001:APS

[HKK^{+}01] Mait Harf, Kristiina Kindel, Vahur Kotkas, Peep Küngas, and Enn Tyugu. Automated program synthesis for Java programming language. Lecture Notes in Computer Science, 2244:157–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2244/22440157.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2244/22440157.
pdf.

Holmes:2009:IJS

[HKL09] Susan Holmes, Adam Kapelner, and Peter P. Lee. An interactive Java statistical image segmentation system: GemIdent. Journal of Statistical Software, 30(10):??, June 2009. CODEN JSSOBK. ISSN 1548-7660. URL http:

//www.jstatsoft.org/v30/
i10.

Hong:2009:CAT

[HKM^{+}09] Sunghyun Hong, Jin-Chul Kim, Soo-Mook Moon, Jin Woo Shin, Jaemok Lee, Hyeong-Seok Oh, and Hyung-Kyu Choi. Client ahead-of-time compiler for embedded Java platforms. Software—Practice and Experience, 39(3): 259–278, March 10, 2009. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Haneda:2002:LJU

[HKS02] I. Haneda, T. Kiyohara, and N. Shiraishi. Linux/Java — unified platform for the nextgeneration of Zaurus. Sharp Technical Journal, 84:18–21, 2002. CODEN ???? ISSN 0285-0362.

Hong:2007:JCA

[HKS^{+}07] SungHyun Hong, Jin-Chul Kim, Jin Woo Shin, Soo-Mook Moon, Hyeong-Seok Oh, Jaemok Lee, and Hyung-Kyu Choi. Java client aheadof-time compiler for embedded systems. ACM SIG-PLAN Notices, 42(7):63–72, July 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Henry:2000:JQH

[HL00] Ethan Henry and Ed Lycklama. Java Q&A: How do you plug Java memory leaks? Dr. Dobb’s Journal of Software Tools, 25(2):115–119, 121, February 2000. CODEN DDJOEB. ISSN 1044-789X.

                                        URL http://www.ddj.com/
                                        ftp/2000/2000_02/jqa220.
                                        txt.

Hightower:2002:JTE

[HL02a] Richard Hightower and Nicholas Lesiecki. Java tools for extreme programming: mastering open source tools including Ant, JUnit, and Cactus. Java open source library. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 2002. ISBN 0-471-20708-X. xxvii + 516 pp. LCCN QA76.73.J38 H54 2002.

Huang:2002:JCA

[HL02b] B. Huang and H. Lin. A Java/CGI approach to developing a geographic virtual reality toolkit on the Internet. Computers and Geosciences, 28(1):13–19, February 2002. CODEN CGEODT, CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Harrison:2003:NBP

[HL03a] P. G. Harrison and C. M.

Llado. A new blocking problem from Java-based schedulers. Performance evaluation, 51(2–4):229–246, 2003. CODEN PEEVD9. ISSN 0166-5316 (print), 1872-745X (electronic).

Huang:2003:JBD

[HL03b] H. P. Huang and C. H. Lu.

Java-based distance learning environment for electronic instruments. IEEE transactions on education, 46(1):88– 94, 2003. CODEN IEEDAB. ISSN 0018-9359.

Hunt:2003:GJE

[HL03c] John Hunt and Chris Loftus. Guide to J2EE: Enterprise Java. Springer professional computing. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. ISBN 1-85233-704-4. xxv + 672 pp. LCCN QA76.73.J38 H855 2003. US$69.95.

Hayden:2004:INW

[HL04] T. W. Hayden and E. C. Loomis. Integrating new Web/Java technologies: Results from KnowledgeKinetics Version 4.0 Research and Development. Simulation series, 36(1):56–61, 2004. CODEN SMCPAX. ISSN 0735-9276.

Haustein:2006:JDJ

[HL06] Max Haustein and Klaus-Peter Löhr. JAC: declarative Java concurrency. Concurrency and Computation: Practice and Experience, 18 (5):519–546, April 25, 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Herlihy:2006:FFIa

[HLM06] Maurice Herlihy, Victor Luchangco, and Mark Moir. A flexible framework for implementing software transactional memory. ACM SIG-PLAN Notices, 41(10):253– 262, October 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Harner:2009:JJR

[HLT09] E. James Harner, Dajie Luo, and Jun Tan. JavaStat: a Java/R-based statistical computing environment. Computational Statistics, 24(2): 295–302, May 2009. CO-DEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s00180-008-0131-y.

Halter:2000:EJP

[HM00] Steven L. Halter and Steven J.

Munroe. Enterprise Java Performance. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2000. ISBN 0-13-017296-0. xxvii + 402 pp. LCCN QA76.73.J38 H347 2001. URL http:/

/www.java-zone.com/free/
bookrevview.asp?content_
id=482;   http://www.phptr.
com/ptrbooks/ptr_0130172960.
html;   http://www.sun.
com/books/catalog/halter/

.

Hartel:2001:FSJ

[HM01a] Pieter H. Hartel and Luc Moreau. Formalizing the safety of Java, the Java Virtual Machine, and Java card. ACM Computing Surveys, 33 (4):517–558, December 2001. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

Hudson:2001:SCG

[HM01b] Richard Hudson and Eliot Moss. Sapphire: Copying GC without stopping the world. In ACM [ACM01b], pages 48–57. ISBN 1-58113-359-6. LCCN QA76.9.O35 A26

2001. URL http://www.
philippsen.com/JGI2001/

camerareadyabstracts/32.
html;   http://www.philippsen.
com/JGI2001/finalpapers/
18500048.pdf.

Hummel:2002:UVB

[HM02] Joe Hummel and Jean Mehta. Using Visual Basic in the CS curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):283–284, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Heidinger:2004:JMS

[HMD04] F. Heidinger, M. Mathes, and H. Dohmann. Java Messaging Service (JMS)— Einsatz in der Industrieautomation. (German) [Java Messaging Service (JMS)— employment in industrial automation]. Automatisierungstechnische Praxis, 46(5):61– 70, 2004. CODEN ???? ISSN 0178-2320.

Hristova:2003:ICJ

[HMRM03] M. Hristova, A. Misra, M. Rutter, and R. Mercuri. Identifying and correcting Java programming errors for introductory computer science students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):153– 156, 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Heydon:2000:PLJ

[HN00] Allan Heydon and Marc Najork. Performance limitations of the Java core libraries. Concurrency: Practice and Experience, 12(6): 363–373, May 2000. CO-DEN CPEXEI. ISSN 1040-

3108. URL http://www3.
interscience.wiley.com/
cgi-bin/abstract/72515723/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=72515723&PLACEBO=IE.
pdf.

Huang:2003:JGJ

[HNZS03] Y. Huang, T. Ni, L. Zhou, and S. Su. JXP4BIGI: a generalized, Java XML-based approach for biological information gathering and integration. Bioinformatics, 19(18): 2351–2358, 2003. CODEN ???? ISSN 1367-4803 (print), 1367-4811 (electronic).

Higuchi:2003:STS

[HO03] Tomoyuki Higuchi and Atsushi Ohori. A static type system for JVM access control. ACM SIGPLAN Notices, 38(9):227–237, September 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Higuchi:2007:STS

[HO07] Tomoyuki Higuchi and Atsushi Ohori. A static type system for JVM access control. ACM Transactions on Programming Languages and Systems, 29(1): 4:1–4:42, January 2007. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Hohpe:2003:AWO

[Hoh03] G. Hohpe. An asynchronous world: Online, you can get all the documentation you want on the Java Messaging Service API. Software Development, 11(7):32–37, 2003. CODEN ???? ISSN 1070-8588.

Holub:2000:TJT

[Hol00a] Allen Holub. Taming Java Threads. Apress, Berkeley, CA, USA, 2000. ISBN 1-893115-10-0. x + 300 pp. LCCN QA76.73.J38 H635 2000. US$34.95.

Holub:2000:CDJ

[Hol00b] Allen I. Holub. Compiler Design in Java. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2000. ISBN 1-55860-560-6. ???? pp. LCCN ???? US$40.

Holzner:2000:JBB

[Hol00c] Steve Holzner. Java Black Book. Coriolis Group Books, Scottsdale, AZ, USA, 2000. ISBN 1-57610-531-8. xxxi + 1143 pp. LCCN QA76.73.J38 H644 2000. US$50.

Holliday:2004:JAI

[Hol04a] M. A. Holliday. A Java applet for illustrating Internet error control. Mathematics and computer education, 38 (3):326–332, 2004. CODEN MCEDDA. ISSN 0730-8639.

Holloway:2004:JGI

[Hol04b] A. Holloway. James Gosling:

The inventor of Java on programming, luck and his grandmother. Canadian Business, 77(8):74, 2004. CO-DEN ???? ISSN 0008-3100.

Holzner:2004:EC

[Hol04c] Steve Holzner. Eclipse Cookbook. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2004. ISBN 0-596-00710-8. 362 (est.) pp. US$44.95, CAN$65.95, UK£31.95. URL

http://www.oreilly.com/
catalog/eclipseckbk/.

Holzner:2004:E

[Hol04d]  Steven Holzner.  Eclipse.

O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2004. ISBN 0-596-00641-1. xiv + 317 pp. LCCN QA76.73.J38 H63 2004. URL

http://www.oreilly.com/
catalog/9780596006419.

Holzner:2005:ADG

[Hol05] Steven Holzner. Ant: the definitive guide. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2005. ISBN 0-596-00609-8. xv + 316 pp. LCCN QA76.73.J38 T57 2005.

Holmes:2006:RFM

[Hol06] Susan Holmes. Review of Fionn Murtagh’s book: Correspondence Analysis and Data Coding with Java and R. R News: the Newsletter of the R Project, 6(4):41–43, October 2006. CODEN ???? ISSN 1609-3631. URL http:/

                                        /CRAN.R-project.org/doc/
                                        Rnews/.

Hong:2005:CAG

[Hon05] S. Hong. Construction and application of GPS locationbased services system based on JAVA. Journal — Chongqing Jianzhu University, 27(2):116–120, 2005. CODEN ???? ISSN 1006-7329.

Hook:2005:BCP

[Hoo05] David Hook. Beginning cryptography and PKI in Java. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 2005. ISBN 0-7645-9633-0. xxvi + 448 pp. LCCN QA76.9.A25 H645 2005. URL ftp:/

                                        /uiarchive.cso.uiuc.edu/
                                        pub/etext/gutenberg/;
                                        http://www.loc.gov/catdir/
                                        toc/ecip0511/2005011272.
                                        html.

Hubbers:2004:IFV

[HOP04] E. Hubbers, M. Oostdijk, and E. Poll. Implementing a Formally Verifiable Security Protocol in Java Card. Lecture Notes in Computer Science, 2802:213–226, 2004. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Horstmann:2000:CCV

[Hor00a] Cay S. Horstmann. Counterpoint: Cobol versus Java. IEEE Software, 17(2):23, 25, March/April 2000. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:

//dlib.computer.org/so/
books/so2000/pdf/s2022.
pdf.

Horstmann:2000:PCD

[Hor00b] Cay S. Horstmann. Counterpoint: Cobol versus Java. IEEE Software, 17(2):23, 25, March/April 2000. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:

//dlib.computer.org/so/
books/so2000/pdf/s2022.
pdf.

Horwitz:2000:DRT

[Hor00c] Susan Horwitz. Debugging via run-time type checking. ACM SIGSOFT Software Engineering Notes, 25 (1):58, January 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Horstmann:2002:BJ

[Hor02a] Cay S. Horstmann. Big Java. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 2002. ISBN 0-471-40248-6 (cloth). xviii + 1202 pp. LCCN QA76.73.J38 H674 2002. URL http:/

/www.loc.gov/catdir/toc/
onix06/2001045305.html.

Horstmann:2002:BJP

[Hor02b] Cay S. Horstmann. Big Java: programming and practice. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 2002. ISBN 0-471-40248-6 (cloth). ???? pp.

Horstmann:2003:CCJ

[Hor03] Cay S. Horstmann. Computing concepts with Java essentials. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 3rd ed edition, 2003. ISBN 0-471-24371-X (paperback). xv + 846 pp. LCCN QA76 .H665 2003.

Horstmann:2005:BJ

[Hor05] Cay S. Horstmann. Big Java. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, second edition, 2005. ISBN 0-471-69703-6 (paperback). xxxii + 1216 pp. LCCN QA76.73.J38 H674 2005.

Houlding:2000:PSC

[Hou00] David Houlding. Publish and subscribe with CORBA Web events. Dr. Dobb’s Journal of Software Tools, 25(7):88, 90–93, July 2000. CODEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/ftp/2000/
2000_07/pscorba.txt;   http:
//www.ddj.com/ftp/2000/
2000_07/pscorba.zip.

Havelund:2000:MCJ

[HP00] Klaus Havelund and Thomas Pressburger. Model checking JAVA programs using JAVA PathFinder. International Journal on Software Tools for Technology Transfer (STTT), 2(4):366–381, March 2000. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Heinle:2002:DJC

[HP02] Nick Heinle and Bill Pena.

Designing with JavaScript: creating dynamic Web pages. O’Reilly Web studio. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2002. ISBN 1-56592-360-X. xi + 216 pp. LCCN QA76.73.J39 H45 2002.

Hubbers:2004:RAC

[HP04] E. Hubbers and E. Poll. Reasoning about card tears and transactions in Java Card. Lecture Notes in Computer Science, 2984(??):114–128, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hartman:2000:EBC

[HPB^{+}00] John Hartman, Larry Peterson, Andy Bavier, Peter Bigot, Patrick Bridges, Brady Montz, Rob Piltz, Todd Proebsting, and Oliver Spatscheck. Experiences building a communicationoriented JavaOS. Software—Practice and Experience, 30(10):1107–1126, August 2000. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (elec-

tronic). URL http://www3.
interscience.wiley.com/
cgi-bin/abstract/72510350/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=72510350&PLACEBO=IE.
pdf.

Herrmann:2003:BJP

[HPH03] A. D. Herrmann, M. E. Patzkowsky, and S. M. Holland. BIOMODULE: a Java program to help model and interpret the stratigraphic record. Computers and Geosciences, 29(1):99–105, 2003. CODEN CGEODT, CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Hovemeyer:2002:AIJ

[HPS02] David Hovemeyer, William Pugh, and Jaime Spacco. Atomic instructions in Java. Lecture Notes in Computer Science, 2374:133–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2374/23740133.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2374/23740133.
pdf.

HarEl:2000:JCB

[HR00] Zvi Har’El and Zvi Rosberg. Java class broker — A seamless bridge from local to distributed programming. Journal of Parallel and Distributed Computing, 60(10):1223–1237, October 2000. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jpdc.
2000.1651;   http://www.
idealibrary.com/links/doi/
10.1006/jpdc.2000.1651/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.
2000.1651/ref.

Havelund:2004:MJP

[HR04a] K. Havelund and G. Rosu.

Monitoring Java programs with Java PathExplorer. Electronic Notes in Theoretical Computer Science, 55(2): 1–18, January 2004. CODEN ???? ISSN 1571-0661.

Havelund:2004:ORV

[HR04b] K. Havelund and G. Rosu.

An overview of the runtime verification tool Java PathExplorer: Special issue on selected papers from the First International Workshop on Runtime Verification Held in Paris, July 2001 (RV01). Formal Methods in System Design, 24(2):189–215, March 2004. CODEN FMSDE6. ISSN 0925-9856.

Hatcher:2005:CCJ

[HRAB05] Philip Hatcher, Mathew Reno, Gabriel Antoniu, and Luc Bougé. Cluster computing with Java. Computing in Science and Engineering, 7 (2):34–39, March/April 2005. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //csdl.computer.org/dl/
                                        mags/cs/2005/02/c2034.htm;
                                         http://csdl.computer.org/
                                        dl/mags/cs/2005/02/c2034.
                                        pdf.

Henkel:2007:DDJ

[HRD07] J. Henkel, C. Reichenbach, and A. Diwan. Discovering documentation for Java container classes. IEEE Transactions on Software Engineering, 33(8):526–543, August 2007. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=

4267024. See errata [HRD08a].

Henkel:2008:EDD

[HRD08a] J. Henkel, C. Reichenbach, and A. Diwan. Errata for “Discovering Documentation for Java Container Classes” [aug 07 526–543]. IEEE Transactions on Software Engineering, 34(2):303, March/ April 2008. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=

4476755. See [HRD07].

Henkel:2008:DDA

[HRD08b] Johannes Henkel, Christoph Reichenbach, and Amer Diwan. Developing and debugging algebraic specifications for Java classes. ACM Transactions on Software Engineering and Methodology, 17(3): 14:1–14:??, June 2008. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Hibbard:2002:JDO

[HRE^{+}02] William Hibbard, Curtis Rueden, Steve Emmerson, Tom Rink, David Glowacki, Tom Whittaker, Don Murray, David Fulker, and John Anderson. Java distributed objects for numerical visualization in VisAD. Communications of the ACM, 45 (4ve):160–170, April 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:

//www.acm.org/cacm/0402/
portal.acm.org/cacm/ve2002/
p160-hibbard.pdf.  Web

only: not in printed issue.

Hibbard:2005:JDC

[HRE^{+}05] W. Hibbard, C. Rueden, S. Emmerson, T. Rink, D. Glowacki, T. Whittaker, D. Murray, D. Fulker, and J. Anderson. Java distributed components for numerical visualization in VisAD. Communications of the ACM, 48(3):98–104, 2005. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hennen:2000:OJL

[HRM00] Dennis S. Hennen, Suresh Ramachandran, and Sandra A. Mamrak. The Object-JavaScript language. Software—Practice and Experience, 30(14):1571–1585, November 25, 2000. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (elec-

                                        tronic). URL http://www3.
                                        interscience.wiley.com/
                                        cgi-bin/abstract/73001731/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=73001731&PLACEBO=IE.
                                        pdf.

Hancock:2000:SCP

[HS00a] Gareth Hancock and Syngress Media, Inc., editors. Sun Certification Programmer for Java 2 Study Guide, Exam 310-025. McGraw-Hill, New York, NY, USA, 2000. ISBN 0-07-212370-2, 0-07-212372-9 (set). xliv + 625 pp. LCCN QA76.3 .S86 2000. US$44.

Harris:2000:LOO

[HS00b] Jacob Harris and Vivek Sarkar. Lightweight objectoriented shared variables for cluster computing in Java. Journal of Parallel and Distributed Computing, 60(10):1238–1259, October 2000. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jpdc.
2000.1652;   http://www.
idealibrary.com/links/doi/
10.1006/jpdc.2000.1652/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.
2000.1652/ref.

Hardy:2001:CQC

[HS01] Yorick Hardy and W.-H Steeb. Classical and quantum computing with C++ and Java simulations. Birkhäuser Verlag, Basel, Switzerland, 2001. ISBN 3-7643-6610-9, 0-8176-6610-9. xxiii + 589 pp. LCCN QA76.73.C153 H367 2001.

Hou:2002:PEJ

[HS02a] Y. Hou and Y. Sun. Partial evaluation of Java byte code program. Journal — Shanghai Jiaotong University, 36 (9):1301–1303, 2002. CODEN ???? ISSN 1006-2467.

Huh:2002:DJB

[HS02b] Moon Yul Huh and Kwangryeol Song. DAVIS: A Java-based data visualization system. Computational Statistics, 17(3):411– 423, September 2002. CO-DEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s001800200116.

Herzog:2005:PJS

[HS05] A. Herzog and N. Shahmehri. Performance of the Java security manager. Computers and Security, 24(3):192–207, 2005. CODEN CPSEDU. ISSN 0167-4048.

Huang:2008:ESS

[HS08] Shan Shan Huang and Yannis Smaragdakis. Expressive and safe static reflection with MorphJ. ACM SIGPLAN Notices, 43(6):79–89, June 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hsiao:2009:EPP

[HSB09] I-Han Hsiao, Sergey Sosnovsky, and Peter Brusilovsky. Extending parameterized problemtracing questions for Java with personalized guidance. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):392, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Hauswirth:2004:PEU

[HSD04] Matthias Hauswirth, Peter F.

Sweeney, and Amer Diwan. Performance explorer: understanding Java application behavior. ACM SIGPLAN Notices, 39(10):31–32, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hsia:2005:TJC

[HSSC05] J. I. Hsia, E. Simpson, D. Smith, and R. Cartwright. Taming Java for the classroom. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):327–331, 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hsu:2001:CAS

[Hsu01] John Y. Hsu. Computer Architecture: Software Aspects, Coding, Hardware. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2001. ISBN 0-8493-1026-1, 1-351-83604-8, 1-4200-4110-X (e-book). 427 pp. LCCN A76.9.A73 H758 2001. US$89.95, UK£59.99.

Hnetynka:2003:FCN

[HT03] P. Hnetynka and P. Tuma.

Fighting class name clashes in Java component systems. Lecture Notes in Computer Science, 2789:106–109, 2003.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hunt:2004:PUT

[HT04] Andrew Hunt and David Thomas. Pragmatic unit testing: in Java with JUnit, volume 2, [pt. 2] of Pragmatic starter kit series. Pragmatic Bookshelf, Raleigh, NC, USA, 2004. ISBN 0-9745140-1-2. xv + 159 pp. LCCN QA76.76.T48 H862 2004. URL http://

www.oreilly.com/catalog/
9780974514017.

Higuera-Toledano:2006:HSD

[HT06] M. Teresa Higuera-Toledano. Hardware support for detecting illegal references in a multiapplication real-time Java environment. ACM Transactions on Embedded Computing Systems, 5(4):753–772, November 2006. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Hayes:2007:IAA

[HTSW07] Alan Hayes, Pete Thomas, Neil Smith, and Kevin Waugh. An investigation into the automated assessment of the design-code interface. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39 (3):324, September 2007. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Hokao:2003:TDM

[HTY^{+}03] S. Hokao, H. Tanaka, M. Yoshihama, T. Furukawa, and M. Ohchi. TAI-18-5 development of management system for student course records using Java and PostgreSQL. Sice, 2:1693–1698, 2003. CO-DEN ????

Hu:2003:FAA

[Hu03] Chenglie Hu. A framework for applet animations with controls. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(4):90–93, December 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp:

//ftp.math.utah.edu/pub/
mirrors/ftp.ira.uka.de/
bibliography/Misc/DBLP/
2003.bib.

Huang:2003:JJB

[Hua03] Y. Huang. JISGA: a Jinibased service-oriented Grid architecture. International Journal of High Performance Computing Applications, 17 (3):317–327, August 2003. CODEN ???? ISSN 1094-3420 (print), 1741-2846 (electronic).

Hubbard:2001:SOT

[Hub01] J. R. (John Rast) Hubbard.

Schaum’s outline of theory and problems of data structures with Java. Schaum’s outline series. McGraw-Hill, New York, NY, USA, 2001. ISBN 0-07-136128-6 (paperback). viii + 369 pp. LCCN QA76.73.J38 H82 2001.

Hubert:2002:CAB

[Hub02] Richard Hubert. Convergent architecture: building model-driven J2EE systems with UML. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 2002. ISBN 0-471-10560-0 (paperback). xxv + 276 pp. LCCN QA76.9.A73

H8 2001. URL http://www.
convergentarchitecture.
com.

Hughes:2002:HMT

[Hug02] Elliott Hughes. How many trivial getter methods does Java have? ACM SIGPLAN Notices, 37(8):19–24, August 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Huisman:2002:VJA

[Hui02] Marieke Huisman. Verification of Java’s Abstract-Collection class: A case study. Lecture Notes in Computer Science, 2386:175–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/

0558/bibs/2386/23860175.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2386/23860175.
pdf.

Hunt:2000:UPP

[Hun00] John Hunt. The unified process for practitioners: objectoriented design, UML and Java. Practitioner series. ISSN: 1439-9245. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. ISBN 1-85233-275-1. xiii + 280 pp. LCCN QA76.9.O35 H86 2000.

Hunt:2002:JOO

[Hun02] John Hunt. Java and object orientation: an introduction. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2002. ISBN 1-85233-569-6, 3-540-76201-9. xxvii + 484 pp. LCCN QA76.73.J38 H86 2002.

Hunt:2003:LSM

[Hun03a] J. Hunt. Look sharp! Microsoft’s C# column has often been described as a Java killer, and the languages have a lot in common. Application Development Advisor, 7(2):32–35, 2003. CODEN ???? ISSN 1369-4200.

Hunt:2003:UIP

[Hun03b] Kenny Hunt. Using image processing to teach CS1 and CS2. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(4):86–89, December 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp:

//ftp.math.utah.edu/pub/
mirrors/ftp.ira.uka.de/
bibliography/Misc/DBLP/
2003.bib.

Hunt:2005:JFE

[Hun05] K. Hunt. A Java framework for experimentation with steganography. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):282– 286, 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hawblitzel:2002:LFJ

[HvE02] Chris Hawblitzel and Thorsten von Eicken. Luna: a flexible Java protection system. Operating Systems Review, 36(5S):391–403, December 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Herlihy:2000:TTD

[HW00] Maurice Herlihy and Michael P. Warres. A tale of two directories: implementing distributed shared objects in Java. Concurrency: Practice and Experience, 12(7): 555–572, May 2000. CO-DEN CPEXEI. ISSN 1040-

3108. URL http://www3.
interscience.wiley.com/

cgi-bin/abstract/72516215/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=72516215&PLACEBO=IE.
pdf.

Hu:2003:DJV

[HWB03] E. Y. S. Hu, A. J. Wellings, and G. Bernat. Deriving Java Virtual Machine timing models for portable worst-case execution time analysis. Lecture Notes in Computer Science, 2889:411–424, 2003. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hu:2004:XED

[HWB04] E. Y. S. Hu, A. Wellings, and G. Bernat. XRTJ: An extensible distributed highintegrity real-time Java environment. Lecture Notes in Computer Science, 2968:208– 228, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Helmer:2001:AID

[HWM01] Guy Helmer, Johnny Wong, and Subhasri Madaka. Anomalous intrusion detection system for hostile Java applets. The Journal of systems and software, 55(3): 273–286, January 15, 2001. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http:/

/www.elsevier.nl/gej-ng/
10/29/11/54/27/28/abstract.
html;   http://www.elsevier.

                                        nl/gej-ng/10/29/11/54/27/
                                        28/article.pdf.

Hyde:2000:JTP

[Hyd00] Paul Hyde. Java thread programming. Howard W. Sams, Indianapolis, IN 46268, USA, 2000. ISBN 0-672-31585-8. iv + 510 pp. LCCN QA76.73.J38 H93 1999.

Hyun:2005:PDC

[Hyu05] C. M. Hyun. Parallel and distributed components with Java. Lecture Notes in Computer Science, 3482:927–937, 2005. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hua:2005:CJE

[HYX05] Han Hua, Dai Yafei, and Li Xiaoming. CSFS: a Java enabled network file storage system. Concurrency and Computation: Practice and Experience, 17(7–8):991– 1003, June/July 2005. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Huang:2004:FPL

[HZC^{+}04] H. Huang, L. Zhang, Q. Cui, T. Jiang, S. Ma, and Y. Gao. Finding potential ligands for PDZ domains by Tailfit, a JAVA program. Chinese Medical Sciences Journal, 19(2): 97–104, 2004. CODEN ???? ISSN 1001-9294.

Huang:2008:DSL

[HZS08] Shan Shan Huang, David Zook, and Yannis Smaragdakis. Domain-specific languages and program generation with meta-AspectJ. ACM Transactions on Software Engineering and Methodology, 18(2):6:1–6:??, November 2008. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Ibbett:2002:WVC

[Ibb02] Roland N. Ibbett. WWW visualisation of computer architecture simulations. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3): 247, September 2002. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Izatt:2000:ATE

[ICB00] Matthew Izatt, Patrick Chan, and Tim Brecht. Ajents: towards an environment for parallel, distributed and mobile Java applications. Concurrency: Practice and Experience, 12(8):667–685, July 2000. CODEN CPEXEI. ISSN 1040-3108. URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/72516221/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=72516221&PLACEBO=IE.
pdf.

IEEE:2002:STI

[IEE02a] IEEE, editor. SC2002: From Terabytes to Insight. Proceedings of the IEEE ACM SC 2002 Conference, November 16–22, 2002, Baltimore, MD, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2002. ISBN 0-7695-1524-X. LCCN ????

IEEE:2002:WII

[IEE02b]  IEEE, editor.  WWC-

5: 2002 IEEE International Workshop on Workload Characterization: November 25, 2002, Austin, Texas. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2002. ISBN 0-7803-7681-1. LCCN QA76.9.S88; QA76.9.S88 W67 2002eb; Internet. URL

http://ieeexplore.ieee.
org/servlet/opac?punumber=

8689. IEEE Catalog Number 02EX633.

IEEE:2003:LES

[IEE03a]  IEEE, editor.  Learning

from Earth’s shapes and sizes: proceedings / IGARSS 2003, 2003 IEEE International Geoscience and Remote Sensing Symposium, Centre de Congrès Pierre Baudis, Toulouse, France, 21–25 July 2003. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2003. ISBN 0-7803-7929-2 (softbound) 0-7803-7930-6 (CD-ROM). LCCN QE33.2.R4 .I58 2003. 7 volumes.

IEEE:2003:PSR

[IEE03b] IEEE, editor. Proceedings of the Symposium on Reliable Distributed Systems, 22 October 2003, Florence, Italy. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2003. CODEN ???? ISBN 0-7695-1955-

5. ISSN 1060-9857. LCCN QA76.9.D5 S94 2003.

Iyer:2001:JBR

[IH01] K. Viswanathan Iyer and M. Harinath. Java based replicated server objects and their remote method invocations in a distributed environment. Software—Practice and Experience, 31(4):357– 370, April 10, 2001. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (elec-

tronic). URL http://www3.
interscience.wiley.com/
cgi-bin/abstract/77004438/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=77004438&PLACEBO=IE.
pdf.

Ishii:2004:SJS

[II04a] Y. Ishii and T. Ito. A secure Java system with confined types and its application. Record of Electrical and Communication Engineering Conversazione Tohoku University, 72(2):132–133, 2004. CODEN ???? ISSN 0385-7719.

IssiCamy:2004:WPD

[II04b] Lázaro Issi Camy and Joseph Issi Cohen. The Web programmer’s desk reference: a complete cross-reference to HTML, CSS, and JavaScript. No Starch Press, San Francisco, CA, USA, 2004. ISBN 1-59327-011-9. xxxiii + 1068 pp. LCCN TK5105.888 .I572 2004eb. URL http://

                                        www.oreilly.com/catalog/
                                        9781593270117.

Itzstein:2003:IHL

[IJ03] G. S. Itzstein and M. Jasiunas. On implementing high level concurrency in Java. Lecture Notes in Computer Science, 2823:151–165, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Itani:2004:JAL

[IK04] W. Itani and A. Kayssi. J2ME application-layer endto-end security for m-commerce. Journal of Network and Computer Applications, 27(1):13– 32, January 2004. CO-DEN JNCAF3. ISSN 1084-8045 (print), 1095-8592 (electronic).

Icking:2003:JAD

[IKKM03] C. Icking, R. Klein, P. Kollner, and L. Ma. Java applets for the dynamic visualization of Voronoi diagrams. Lecture Notes in Computer Science, 2598:191–205, 2003. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Illmann:2001:TMM

[IKKW01] Torsten Illmann, Tilman Krueger, Frank Kargl, and Michael Weber. Transparent migration of mobile agents using the Java platform debugger architecture. Lecture Notes in Computer Science, 2240:198–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2240/22400198.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2240/22400198.
pdf.

Inagaki:2003:IPS

[IKN03] Tatsushi Inagaki, Hideaki Komatsu, and Toshio Nakatani. Integrated prepass scheduling for a Java Just-In-Time compiler on the IA-64 architecture. In ACM [ACM03a], pages 159–168. ISBN 0-7695-1913-X. ACM Order No.: 530033.

Ishizaki:2000:SDT

[IKY^{+}00a] Kazuaki Ishizaki, Motohiro Kawahito, Toshiaki Yasue, Hideaki Komatsu, and Toshio Nakatani. A study of devirtualization techniques for a Java^{T M} Just-In-Time compiler. ACM SIGPLAN Notices, 35(10):294–310, October 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
proceedings/oops/353171/
p294-ishizaki/.

Ishizaki:2000:DIE

[IKY^{+}00b] Kazuaki Ishizaki, Motohiro Kawahito, Toshiaki Yasue, Mikio Takeuchi, Takeshi Ogasawara, Toshio Suganuma, Tamiya Onodera, Hideaki Komatsu, and Toshio Nakatani. Design, implementation, and evaluation of optimizations in a Java^{T M} Just-In-Time compiler. Concurrency: Practice and Experience, 12(6): 457–475, May 2000. CO-DEN CPEXEI. ISSN 1040-

3108. URL http://www3.
interscience.wiley.com/
cgi-bin/abstract/72515728/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=72515728&PLACEBO=IE.
pdf.

Inoue:2009:HJV

[IN09] Hiroshi Inoue and Toshio Nakatani. How a Java VM can get more from a hardware performance monitor. ACM SIGPLAN Notices, 44(10):137–154, October 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Inghelbrecht:2009:OOD

[Ing09] Yanic Inghelbrecht. Objectoriented design with trace modeler and Trace4J. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3): 375, September 2009. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Ishikawa:2005:JOL

[INM05] A. Ishikawa, E. Nakamura, and J. J. Mahoney. Jurassic oceanic lithosphere beneath the southern Ontong Java Plateau: Evidence from xenoliths in alnoite, Malaita, Solomon Islands. Geology, 33 (5):393–396, 2005. CODEN ???? ISSN 0091-7613.

Igarashi:2001:FJM

[IPW01] Atsushi Igarashi, Benjamin C.

Pierce, and Philip Wadler. Featherweight Java: a minimal core calculus for Java and GJ. ACM Transactions on Programming Languages and Systems, 23(3):396–450, May 2001. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Iosif:2003:TLP

[IS03] R. Iosif and R. Sisto. Temporal logic properties of Java objects. The Journal of systems and software, 68(3):243– 251, 2003. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Inoue:2006:PJO

[ISF06] H. Inoue, D. Stefanovic, and S. Forrest. On the prediction of Java object lifetimes. IEEE Transactions on Computers, 55(7):880–892, July 2006. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&
arnumber=1637403.

Ishimoto:2001:POB

[Ish01]  Kenya Ishimoto.  Prefer-

ence object-based internationalization for distributed application framework in Java. In Unicode Consortium [Uni01], page ?? ISBN ???? LCCN ???? URL

http://www.unicode.org/
iuc/iuc18/papers/a16.ppt.

ISO:2005:IDM

[ISO05]  ISO.  ISO 19005-1:2005,

Document management— Electronic document file format for long-term preservation— Part 1: Use of PDF 1.4 (PDF/A-1). International Organization for Standardization, Geneva, Switzerland, 2005. URL http:/

/www.aiim.org/documents/
standards/ISO_19005-1_(
E).doc;   http://www.aiim.
org/pdf_a/.

ISO:2008:IIId

[ISO08] ISO. ISO/IEC 9075-13:2008 Information technology — Database languages — SQL — Part 13: QL Routines and Types Using the Java TM Programming Language (SQL/JRT). International Organization for Standardization, Geneva, Switzerland, July 17, 2008. ISBN ???? 186 pp. LCCN ???? URL

http://www.iso.org/iso/
iso_catalogue/catalogue_
tc/catalogue_detail.htm?
csnumber=38646.

Ishizaki:2003:ECP

[ITK^{+}03] Kazuaki Ishizaki, Mikio Takeuchi, Kiyokuni Kawachiya, Toshio Suganuma, Osamu Gohda, Tatsushi Inagaki, Akira Koseki, Kazunori Ogata, Motohiro Kawahito, Toshiaki Yasue, Takeshi Ogasawara, Tamiya Onodera, Hideaki Komatsu, and Toshio Nakatani. Effectiveness of cross-platform optimizations for a Java just-in-time compiler. ACM SIGPLAN Notices, 38(11):187–204, November 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Igarashi:2006:VPT

[IV06] Atsushi Igarashi and Mirko Viroli. Variant parametric types: a flexible subtyping scheme for generics. ACM Transactions on Programming Languages and Systems, 28(5):795–847, September 2006. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Igarashi:2007:VPT

[IV07] Atsushi Igarashi and Mirko Viroli. Variant path types for scalable extensibility. ACM SIGPLAN Notices, 42 (10):113–132, October 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ivancsy:2002:HWJ

[Iva02] R. Ivancsy. How to write Java applet? Elektrotechnika, 95 (10):300–302, 2002. CODEN ???? ISSN 0367-0708.

Ive:2003:TER

[Ive03a] Anders Ive. Towards an embedded real-time Java Virtual Machine. Thesis (fil.dr.), Lund University, Department of Computer Science, Lund, Sweden, 2003. viii + 139 pp.

Iverson:2003:MXJ

[Ive03b] Will Iverson. Mac OS X for Java Geeks. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2003. ISBN 0-596-00400-1. xii + 282 pp. LCCN QA76.76.O63 I936 2003. US$39.95 US, CAN$61.95, UK£28.50. URL

                                        http://www.oreilly.com/
                                        catalog/macxjvgks/.

Jepsen:2001:JTS

[JA01] Thomas C. Jepsen and Farooq Anjum. Java in telecommunications: solutions for next generation networks. Wiley series in communications networking and distributed systems. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 2001. ISBN 0-471-49826-2. vii + 193 pp. LCCN QA76.73.J38 J368 2001. URL http:/

/www.loc.gov/catdir/toc/
onix06/2001026305.html.

Jackson:2001:JQW

[Jac01a] Jared Jackson. Java Q&A:

What are resource bundles and what do they have to do with internationalization? Dr. Dobb’s Journal of Software Tools, 26 (9):117–119, September 2001. CODEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/ftp/2001/
2001_09/jqa0901.txt.

Jacobs:2001:FJE

[Jac01b] Bart Jacobs. A formalisation of Java’s exception mechanism. Lecture Notes in Computer Science, 2028:284–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2028/20280284.
htm;   http://link.springer-
ny.com/link/service/series/

                                        0558/papers/2028/20280284.
                                        pdf.

Jacobs:2001:JPV

[Jac01c]  Bart Jacobs.  JavaCard

program verification. Lecture Notes in Computer Science, 2152:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.
                                        com/link/service/series/
                                        0558/bibs/2152/21520001.
                                        htm;   http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2152/21520001.
                                        pdf.

Jacobs:2003:JIT

[Jac03] B. Jacobs. Java’s integral types in PVS. Lecture Notes in Computer Science, 2884:1– 15, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jacobs:2004:WPC

[Jac04a] B. Jacobs. Weakest precondition reasoning for Java programs with JML annotations. Journal of Logic and Algebraic Programming, 58(1–2):61–88, 2004. CODEN ???? ISSN 1567-8326.

Jacobsen:2004:MAI

[Jac04b] Hans-Arno Jacobsen, editor. Middleware 2004: ACM/IFIP/USENIX International Middleware Conference, Toronto, Canada, October 18–22, 2004: proceedings, volume 3231 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. CODEN LNCSD9. ISBN 3-540-23428-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.M54 I33 2004. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.springerlink.
com/openurl.asp?genre=issue&
issn=0302-9743&volume=3231;
 http://www.springerlink.
com/openurl.asp?genre=volume&
id=doi:10.1007/b101561.

Jamil:2001:CBN

[Jam01] T. Jamil. The complex binary number system. IEEE Potentials, 20(5):39–41, December 2001. CODEN IEPTDF. ISSN 0278-6648 (print), 1558-1772 (electronic).

Jipping:2003:UJT

[JBMP03] M. J. Jipping, A. Bugaj, L. Mihalkova, and D. E. Porter. Using Java to teach networking concepts with a programmable network sniffer. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):120–124, 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Jo:2004:CCF

[JC04] J. W. Jo and B. M. Chang. Constructing control flow graph for Java by decoupling exception flow from normal flow. Lecture Notes in Computer Science, 3043:106–113, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jordan:2004:EJT

[JCKS04] Mick Jordan, Grzegorz Czajkowski, Kirill Kouklinski, and Glenn Skinner. Extending a J2EE server with dynamic and flexible resource management. In Jacobsen [Jac04b], pages 439– ?? CODEN LNCSD9. ISBN 3-540-23428-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.M54 I33 2004. URL ftp://

                                        uiarchive.cso.uiuc.edu/
                                        pub/etext/gutenberg/;
                                        http://www.springerlink.
                                        com/openurl.asp?genre=issue&
                                        issn=0302-9743&volume=3231;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=volume&
                                        id=doi:10.1007/b101561.

Jipping:2007:TSJ

[JCOP07] Michael J. Jipping, Cameron Calka, Brian O’Neill, and Christopher R. Padilla. Teaching students Java bytecode using Lego Mindstorms robots. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):170–174, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Jeon:2005:JJB

[JCP^{+}05] Yoon-Seong Jeon, Hwanwon Chung, Sunyoung Park, Inae Hur, Jae-Hak Lee, and Jongsik Chun. jPHYDIT: a Java-based integrated environment for molecular phylogeny of ribosomal RNA sequences. Bioinformatics, 21(14):3171–3173, July 15, 2005. CODEN ???? ISSN 1367-4803 (print), 1367-4811 (electronic).

Jo:2004:UEA

[JCYC04] Jang-Wu Jo, Byeong-Mo Chang, Kwangkeun Yi, and Kwang-Moo Choe. An uncaught exception analysis for Java. The Journal of systems and software, 72(1): 59–69, June 2004. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Jordan:2006:SJT

[JDJ^{+}06] Mick Jordan, Laurent Daynès, Marcin Jarzab, Ciarán Bryce, and Grzegorz Czajkowski. Scaling J2EE ^{T M} application servers with the Multi-tasking Virtual Machine. Software— Practice and Experience, 36 (6):557–580, May 2006. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jennings:2000:JQC

[Jen00a] Mike Jennings. Java Q&A: Can you write NT services in Java? Dr. Dobb’s Journal of Software Tools, 25(3):113–116, 118, March 2000. CODEN DDJOEB. ISSN 1044-789X. URL

                                        http://www.ddj.com/ftp/
                                        2000/2000_03/jqa300.txt;
                                         http://www.ddj.com/ftp/
                                        2000/2000_03/jqa300.zip.

Jennings:2000:JQH

[Jen00b] Mike Jennings. Java Q&A: How do I establish a net connection? Dr. Dobb’s Journal of Software Tools, 25(11): 141–142, 144–147, November 2000. CODEN DDJOEB. ISSN 1044-789X. URL http:

                                        //www.ddj.com/ftp/2000/
                                        2000_11/jqa0011.txt;   http:
                                        //www.ddj.com/ftp/2000/
                                        2000_11/jqa0011.zip.

Jensen:2001:DRT

[Jen01] E. D. Jensen. The distributed real-time specification for Java: an initial proposal. International Journal of Computer Systems Science and Engineering, 16(2): ??, March 2001. CODEN CSSEEI. ISSN 0267-6192.

Jenkins:2002:GJP

[Jen02a] Stephen B. Jenkins. Generating JavaScript from Perl. Dr. Dobb’s Journal of Software Tools, 27(5):59–63, May 2002. CODEN DDJOEB.

ISSN 1044-789X. URL http:

//www.ddj.com/ftp/2002/
2002_05/jscript.txt.

Jennings:2002:JQ

[Jen02b] Mike Jennings. Java Q&A:

How do I use the Java plug-in tool? Dr. Dobb’s Journal of Software Tools, 27(1):93–94, January 2002. CODEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/ftp/2002/
2002_01/jqa0102.zip.

Jugravu:2005:JPM

[JF05] Alexandru Jugravu and Thomas Fahringer. JavaSymphony, a programming model for the Grid. Future Generation Computer Systems, 21 (1):239–246, January 1, 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Jacobi:2006:PJA

[JF06] Jonas Jacobi and John R. Fallows. Pro JSF and Ajax: building rich Internet components. The Expert’s voice in Java. Apress, Berkeley, CA, USA, 2006. ISBN 1-59059-580-7. xxiii + 435 pp. LCCN ????

Jarc:2000:ABI

[JFH00] Duane J. Jarc, Michael B.

Feldman, and Rachelle S. Heller. Assessing the benefits of interactive prediction using Web-based algorithm animation courseware. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):377– 381, March 2000. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Jubin:2000:EJE

[JFt00] Henri Jubin, Jürgen Friedrichs, and the Jalapeño Team. Enterprise JavaBeans by Example. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2000. ISBN 0-13-022475-8. xvii + 223 pp. LCCN QA76.73.J38 J82 2000. US$49.99. URL http:/

/www.phptr.com/ptrbooks/
ptr_0130224758.html.

Jha:2003:JIP

[JH03] S. Jha and M. Hassan. Java implementation of policybased bandwith management. International Journal of Network Management, 13(4): 249–258, 2003. CODEN ???? ISSN 1055-7148.

Johnson:2005:PJD

[JHA^{+}05] Rod Johnson, Juergen Hoeller, Alef Arendsen, Thomas Risberg, and Colin Sampaleanu, editors. Professional Java development with the Spring Framework. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 2005. ISBN 0-7645-7483-3 (paper). xxviii + 644 pp. LCCN QA76.73.J38 P74585 2005.

URL http://www.wiley.
com/WileyCDA/WileyTitle/
productCd-0764574833,descCd-
authorInfo.html.

Jiahai:2004:TWO

[JHJX04] Yang Jiahai, Duan Haixin, Wu Jianping, and Li Xing. Thresholds: Workflow oriented network management: a Web/Java approach. Journal of Network and Systems Management, 12(4):431–439, December 2004. CODEN ???? ISSN 1064-7570.

Jun:2003:CDT

[JHSL03] Q. Jun, L. Hongzhi, L. Shiqi, and C. Lie. A collaborative design tool based on Java 3D and CORBA. Journal - Huazhong University of Science and Technology Nature Science Chinese Edition, 31 (1):25–27, 2003. CODEN ???? ISSN 1671-4512.

Jia:2000:OOS

[Jia00] Xiaoping Jia. Object-oriented software development using Java: principles, patterns, and frameworks. Addison-Wesley, Reading, MA, USA, 2000. ISBN 0-201-35084-X. xvi + 507 pp. LCCN QA76.64.J53 2000.

Jian:2004:DJJ

[Jia04] L. Jian. Distributed Java just-in-time compiler for embedded system. Cognitive Science Research Paper- University of Sussex Csrp, 571:

30–31, 2004. CODEN ???? ISSN 1350-3162.

Jibson:2002:JPU

[JJ02a] Randall W. Jibson and Matthew W. Jibson. Java programs for using Newmark’s method to model slope performance during earthquakes. Denver, CO, USA, version 1.0 edition, 2002. Includes CD-ROM.

Jung:2002:DIS

[JJ02b] Jun-Young Jung and Min-Soo Jung. Design and implementation of small-sized Java Virtual Machine on Java platform Jini. Lecture Notes in Computer Science, 2343:571– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/
service/series/0558/bibs/
2343/23430571.htm;   http:
//link.springer.de/link/
service/series/0558/papers/
2343/23430571.pdf.

Jones:2000:AJC

[JK00] Joel Jones and Samuel Kamin. Annotating Java class files with virtual registers for performance. Concurrency: Practice and Experience, 12 (6):389–406, May 2000. CO-DEN CPEXEI. ISSN 1040-

3108. URL http://www3.
interscience.wiley.com/
cgi-bin/abstract/72515727/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?

ID=72515727&PLACEBO=IE.
pdf.

Juric:2004:JRR

[JKH^{+}04] Matjaz B. Juric, Bostjan Kezmah, Marjan Hericko, Ivan Rozman, and Ivan Vezocnik. Java RMI, RMI tunneling and Web services comparison and performance analysis. ACM SIGPLAN Notices, 39(5):58–65, May 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jung:2005:RTE

[JKJ05] J. Y. Jung, K. S. Kim, and M. S. Jung. Real-time embedded middleware system using Java-native combination model. Transactions — Korean Institute of Electrical Engineers D, 54(3):141–147, 2005. CODEN ???? ISSN 1229-6287.

Jipping:2004:IWW

[JKKL04] M. J. Jipping, A. Kalafut, N. Kooistra, and K. Ludewig. Investigating wired and wireless networks using a Javabased programmable sniffer. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):12–16, 2004. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Jacobs:2003:JPV

[JKW03] B. Jacobs, J. Kiniry, and M. Warnier. Java program verification challenges. Lecture Notes in Computer Science, 2852:202–219, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jacobs:2002:DSD

[JLV02] P. H. M. Jacobs, N. A. Lang, and A. Verbraeck. D-SOL: a distributed Java based discrete event simulation architecture. Winter Simulation Conference, 1(??):793–800, 2002. CODEN WSCPDK. ISSN 0743-1902 (print), 2160-9276 (electronic).

Jaen-Martinez:2000:JME

[JM00] Javier Jaen-Martinez. The Java Management Extensions (JMX): Is your cluster ready for evolution? Journal of Parallel and Distributed Computing, 60(10):1341–1353, October 2000. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

                                        http://www.idealibrary.
                                        com/links/doi/10.1006/jpdc.
                                        2000.1657;   http://www.
                                        idealibrary.com/links/doi/
                                        10.1006/jpdc.2000.1657/
                                        pdf;   http://www.idealibrary.
                                        com/links/doi/10.1006/jpdc.
                                        2000.1657/ref.

Joao:2008:IPOa

[JMK^{+}08a] Jose A. Joao, Onur Mutlu, Hyesoon Kim, Rishi Agarwal, and Yale N. Patt. Improving the performance of objectoriented languages with dynamic predication of indirect jumps. ACM SIGARCH Computer Architecture News, 36(1):80–90, March 2008. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Joao:2008:IPOb

[JMK^{+}08b] Jose A. Joao, Onur Mutlu, Hyesoon Kim, Rishi Agarwal, and Yale N. Patt. Improving the performance of objectoriented languages with dynamic predication of indirect jumps. Operating Systems Review, 42(2):80–90, March 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Joao:2008:IPOc

[JMK^{+}08c] Jose A. Joao, Onur Mutlu, Hyesoon Kim, Rishi Agarwal, and Yale N. Patt. Improving the performance of objectoriented languages with dynamic predication of indirect jumps. ACM SIGPLAN Notices, 43(3):80–90, March 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Joshi:2003:FOJ

[JMM03] Rushikesh K. Joshi, Maureen Mascarenhas, and Yogesh Murarka. Filter objects for Java. Software— Practice and Experience, 33 (6):509–522, May 2003. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Joao:2009:FRC

[JMP09] José A. Joao, Onur Mutlu, and Yale N. Patt. Flexible reference-counting-based hardware acceleration for garbage collection. ACM SIGARCH Computer Architecture News, 37(3):418– 428, June 2009. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Jipping:2002:UJD

[JMS02] Michael J. Jipping, Steve Marlowe, and Alexander Sherstov. Using Java to design and test hardware circuits over a classroom network. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):162–166, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Joisha:2002:EAJ

[JMSG02] Pramod G. Joisha, Samuel P. Midkiff, Mauricio J. Serrano, and Manish Gupta. Efficiently adapting Java binaries in limited memory contexts. International Journal of Parallel Programming, 30(4):257–289, August 2002. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http://

ipsapp009.lwwonline.com/
content/getfile/4773/28/
3/abstract.htm;   http:/
/ipsapp009.lwwonline.com/
content/getfile/4773/28/
3/fulltext.pdf;   http:
//www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=30&
issue=4&spage=257.

Jank:2003:OOI

[JO03] K. Jank and R. Oberhauser.

An object-oriented invocation layer for the Java Message Service. Lecture Notes in Computer Science, 2889:57– 69, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Johnson:2000:DSC

[Joh00a]  Steve Johnson.  A dy-

namic select component for JavaScript. Dr. Dobb’s Journal of Software Tools, 25 (1):36, 38–42, January 2000. CODEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/ftp/2000/
2000_01/dselect.txt;   http:
//www.ddj.com/ftp/2000/
2000_01/dselect.zip.

Johnson:2000:SFP

[Joh00b] Verlyn Johnson. The San Francisco project: business process components and infrastructure. ACM Computing Surveys, 32(1es), March 2000. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/surveys/2000-32-
                                        1/p25-johnson/. Article No.

25.

Johnson:2003:SJA

[Joh03]  P. Johnson.  Scaling up

Java applications on Windows servers. Cmg, 1(??):103– 112, 2003. CODEN ????

Johnson:2006:JT

[Joh06] Stephen Johnson. Java in a teacup. ACM Queue: Tomorrow’s Computing Today, 4(3):36–41, April 2006. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Jolin:2001:JQC

[Jol01] Art Jolin. Java Q&A: Can J2EE manage cache and pool memory? Dr. Dobb’s Journal of Software Tools, 26 (10):117–120, October 2001. CODEN DDJOEB. ISSN 1044-789X. URL http:

                                        //www.ddj.com/ftp/2001/
                                        2001_10/jqa1001.txt;   http:
                                        //www.ddj.com/ftp/2001/
                                        2001_10/jqa1001.zip.

Jones:2002:JMA

[Jon02] M. Tim Jones. Java mobile agents and the Aglets SDK. Dr. Dobb’s Journal of Software Tools, 27(1):42, 44, 46– 48, January 2002. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2002/2002_01/aglet.
txt.

Jorelid:2002:JFT

[Jor02] Lennart Jorelid. J2EE FrontEnd technologies: a programmer’s guide to servlets, JavaServer pages, and Enterprise JavaBeans. Books for professionals by professionals. Apress, Berkeley, CA, USA, 2002. ISBN 1-893115-96-8. xiii + 1112 pp. LCCN TK5105.8885.J38 J67 2002.

Jacobs:2000:MBJ

[JP00] Bart Jacobs and Erik Poll.

A monad for basic Java semantics. Lecture Notes in Computer Science, 1816:150– ??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/1816/18160150.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/1816/18160150.
pdf.

Jacobs:2001:LJM

[JP01] Bart Jacobs and Erik Poll.

A logic for the Java modeling language JML. Lecture Notes in Computer Science, 2029:284–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/

                                        0558/bibs/2029/20290284.
                                        htm;   http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2029/20290284.
                                        pdf.

Jacobs:2003:CMS

[JP03] B. Jacobs and E. Poll. Coalgebras and monads in the semantics of Java. Theoretical Computer Science, 291 (3):329–349, 2003. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Jacobs:2004:JPV

[JP04] B. Jacobs and E. Poll. Java program verification at Nijmegen: Developments and perspective. Lecture Notes in Computer Science, 3233:134– 153, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jung:2008:EEH

[JPB^{+}08] Dong-Heon Jung, JongKuk Park, Sung-Hwan Bae, Jaemok Lee, and Soo-Mook Moon. Efficient exception handling in Java bytecode-to-C aheadof-time compiler for embedded systems. Computer Languages, Systems and Structures, 34(4):170–183, December 2008. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1477842407000176.

Jaworski:2000:JSH

[JPC00] Jamie Jaworski, Paul J. Perrone, and Venkata S. R. R. Chaganti. Java security handbook. Howard W. Sams, Indianapolis, IN 46268, USA, 2000. ISBN 0-672-31602-1. xiv + 534 pp. LCCN QA76.9.A25 J395 2000. Also available via the Internet.

Jovanovic:2005:MDS

[JPJ05] N. Jovanovic, R. Popovic, and Z. Jovanovic. Modeling distributed system entities which communicate asynchronous via message passing in Java. Neural, Parallel and Scientific Computations, 13(1):63–76, 2005. CODEN NPACEM. ISSN 1061-5369.

Jacobs:2008:PMC

[JPS^{+}08] Bart Jacobs, Frank Piessens, Jan Smans, K. Rustan M. Leino, and Wolfram Schulte. A programming model for concurrent object-oriented programs. ACM Transactions on Programming Languages and Systems, 31(1): 1:1–1:48, December 2008. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Joshi:2009:RDP

[JPSN09] Pallavi Joshi, Chang-Seo Park, Koushik Sen, and Mayur Naik. A randomized dynamic program analysis technique for detecting real deadlocks. ACM SIGPLAN Notices, 44(6):110–120, June 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jacob:2002:CAP

[JR02] Matthias Jacob and Keith Randall. Cross-architectural performance portability of a Java Virtual Machine implementation. In USENIX Association [USE02], page ?? ISBN 1-931971-01-3. LCCN QA76.73.J38 J42 2002. URL

                                        http://www.usenix.org/publications/
                                        library/proceedings/javavm02/
                                        jacob.html.

Jordan:2003:JDO

[JR03] David Jordan and Craig Russell. Java Data Objects. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2003. ISBN 0-596-00276-9. xxi + 356 pp. LCCN QA76.73.J38 J66 2003. US$39.95 US, CAN$61.95, UK£28.50. URL http://

                                        www.oreilly.com/catalog/
                                        jvadtaobj/.

Jeffrey:2005:JJF

[JR05] A. Jeffrey and J. Rathke. Java Jr.: Fully abstract trace semantics for a core Java language. Lecture Notes in Computer Science, 3444:423–438, 2005. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jayaraman:2005:KDI

[JRH05] G. Jayaraman, V. P. Ranganath, and J. Hatcliff. Kaveri: Delivering the Indus Java Program Slicer to Eclipse. Lecture Notes in Computer Science, 3442:269– 272, 2005. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Juric:2000:JDO

[JRN00] Matjaz B. Juric, Ivan Rozman, and Simon Nash. Java 2 distributed object middleware performance analysis and optimization. ACM SIGPLAN Notices, 35(8):31– 40, August 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jagannathan:2001:ICS

[JS01] Ramesh Jagannathan and Paolo A. G. Sivilotti. Increasing client-side confidence in remote component implementations. ACM SIGSOFT Software Engineering Notes, 26(5):52–61, September 2001. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Jeong:2004:JBS

[JSSM04] J. Jeong, D. Shin, D. Shin, and K. Moon. Java-based single sign-on library supporting SAML (Security Assertion Markup Language) for distributed Web services.

Lecture Notes in Computer Science, 3007:891–894, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jacobson:2004:ITE

[JT04] N. Jacobson and A. Thornton. It is time to emphasize ArrayLists over Arrays in Java-based first programming courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(4):88–92, 2004. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Juola:2007:PCO

[Juo07] Patrick Juola. Principles of computer organization and assembly language: using the Java Virtual Machine. Pearson/Prentice Hall, Upper Saddle River, NJ, USA, 2007. ISBN 0-13-148683-7 (paperback). xv + 317 pp. LCCN QA76.9.C643 J96 2007. URL http:/

/www.loc.gov/catdir/toc/
ecip073/2006034154.html.

Jacobs:2004:STS

[JV04] P. H. M. Jacobs and A. Verbraeck. Single-threaded specification of process-interaction formalism in Java. Winter Simulation Conference, 2: 1548–1555, 2004. CODEN ???? ISSN 0743-1902 (print), 2160-9276 (electronic).

Jiang:2003:AJM

[JWC03] Y. Jiang, Z. Wang, and T. Chen. Application of Java in multimedia Netware courseware. Journal — Jiangsu Polytechnic University, 15(1):61–64, 2003. CO-DEN ????

Kniesel:2002:CCC

[KA02] Günter Kniesel and Michael Austermann. CC4J — code coverage for Java. Lecture Notes in Computer Science, 2370:155–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2370/23700155.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2370/23700155.
pdf.

Kafura:2000:OOS

[Kaf00]  Dennis Kafura.  Object-

Oriented Software Design and Construction with Java. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2000. ISBN 0-13-011264-X. xv + 656 pp. LCCN QA76.64. K35 2000. US$49.00. URL http:/

/www.phptr.com/ptrbooks/
esm_013011264X.html.

Kagawa:2009:WWB

[Kag09] Koji Kagawa. WAPPEN: a Web-based application framework for programming and its bison/flex plug-in. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3): 383, September 2009. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Kahrel:2006:AIR

[Kah06a]  Peter Kahrel.  Automat-

ing InDesign with regular expressions. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2006. ISBN 0-596-52937-6. LCCN Z253.532.A34; Z253.532.A34 K34 2006eb. URL http://

www.oreilly.com/catalog/
9780596529376.

Kahrel:2006:SIJ

[Kah06b]  Peter Kahrel.  Scripting

InDesign with JavaScript. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2006. ISBN 0-596-52817-5. LCCN Z253.532.A34. URL http://

www.oreilly.com/catalog/
9780596528171.

Kalin:2001:OOP

[Kal01]  Martin Kalin.  Object-

Oriented Programming in Java. The JK computer science and mathematics series. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2001. ISBN 0-13-019859-5. xxvii + 641 pp. LCCN QA76.64 .K36 2001. US$58.00. URL http:/

/www.phptr.com/ptrbooks/
esm_0130198595.html.

Kalinovsky:2004:CJT

[Kal04]  Alex Kalinovsky.  Covert

Java: Techniques for Decompiling, Patching, and Reverse Engineering. Howard W. Sams, Indianapolis, IN 46268, USA, 2004. ISBN 0-672-32638-8. x + 268 pp. LCCN QA76.73.J38 K35 2004. US$29.99.

Kanalakis:2002:WSJ

[Kan02] John M. Kanalakis, Jr.

Web services and Java server pages. Dr. Dobb’s Journal of Software Tools, 27(1):28, 30–34, January 2002. CODEN DDJOEB. ISSN 1044-789X. URL

http://www.ddj.com/ftp/
2002/2002_01/wsjsp.txt;
http://www.ddj.com/ftp/
2002/2002_01/wsjsp.zip.

Keane:2003:DJP

[KAN^{+}03] T. Keane, R. Allen, T. J.

Naughton, J. McInerney, and J. Waldron. Distributed Java platform with programmable MIMD capabilities. Lecture Notes in Computer Science, 2604:122–131, 2003. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kolling:2004:EAB

[KB04a] M. Kolling and D. J. Barnes.

Enhancing apprentice-based learning of Java. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):286– 290, 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kosa:2004:TVC

[KB04b] Martha J. Kosa and Mark A. Boshart. Treemap visualizations for CS2. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3): 259, September 2004. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kreuzinger:2003:RTE

[KBP^{+}03] J. Kreuzinger, U. Brinkschulte, M. Pfeffer, S. Uhrig, and T. Ungerer. Real-time eventhandling and scheduling on a multithreaded Java microcontroller. Microprocessors and Microsystems, 27(1):19– 31, 2003. CODEN MIMID5. ISSN 0141-9331 (print), 1872-9436 (electronic).

Kats:2008:MSB

[KBV08] Lennart C. L. Kats, Martin Bravenboer, and Eelco Visser. Mixing source and bytecode: a case for compilation by normalization. ACM SIGPLAN Notices, 43(10):91–108, September 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Klemm:2007:JIO

[KBVP07] Michael Klemm, Matthias Bezold, Ronald Veldema, and Michael Philippsen. JaMP: an implementation of OpenMP for a Java DSM. Concurrency and Computation: Practice and Experience, 19(18):2333–2352, December 25, 2007. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kim:2000:JBO

[KC00] Austin Kim and Morris Chang. Java bytecode optimization with advanced instruction folding mechanism. Lecture Notes in Computer Science, 1940:268–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/1940/19400268.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/1940/19400268.
pdf.

Kingston:2001:ADS

[KC01] Jeffrey Kingston and Brian Cole. Algorithms and Data Structures: Design, Correctness, Analysis: Java Edition. Addison-Wesley, Reading, MA, USA, 2001. ISBN 0-201-36062-4. ???? pp. LCCN ???? US$44.

Krapf:2003:ESP

[KC03] R. Krapf and L. Carro. Efficient signal processing in embedded Java systems. IEEE International Symposium on Circuits and Systems, 4:IV– 61–IV–64, 2003. CODEN ???? ISSN 0271-4310.

Keeton:2001:SEU

[KCF01] Brian Keeton, Chuck Cavaness, and Geoff Friesen. Special edition using Java 2. Que Corporation, Indianapolis, IN, USA, 2001. ISBN 0-7897-2468-5. xiv + 1126 pp. LCCN QA76.73.J38 K44 2001.

Kazi:2000:TOH

[KCSL00] Iffat H. Kazi, Howard H. Chen, Berdenia Stanley, and David J. Lilja. Techniques for obtaining high performance in Java programs. ACM Computing Surveys, 32(3):213–240, 2000. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic). URL

                                        http://www.acm.org/pubs/
                                        articles/journals/surveys/
                                        2000-32-3/p213-kazi/p213-
                                        kazi.pdf;   http://www.
                                        acm.org/pubs/citations/
                                        journals/surveys/2000-32-
                                        3/p213-kazi/.

Kapitza:2006:FIA

[KDH^{+}06] Rüdiger Kapitza, Jörg Domaschka, Franz J. Hauck, Hans P. Reiser, and Holger Schmidt. FORMI: Integrating adaptive fragmented objects into Java RMI. IEEE Distributed Systems Online, 7 (10):??, October 2006. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
ds/2006/10/ox001.pdf.

Kats:2009:PRF

[KdJNNV09] Lennart C. L. Kats, Maartje de Jonge, Emma Nilsson-Nyman, and Eelco Visser. Providing rapid feedback in generated modular language environments: adding error recovery to scannerless generalized-LR parsing. ACM SIGPLAN Notices, 44(10):445–464, October 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Keschenau:2004:REU

[Kes04]  Martin Keschenau.  Re-

verse engineering of UML specifications from Java programs. ACM SIGPLAN Notices, 39(10):326–327, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kistler:2000:ADM

[KF00] Thomas Kistler and Michael Franz. Automated datamember layout of heap objects to improve memoryhierarchy performance. ACM Transactions on Programming Languages and Systems, 22(3):490–505, 2000. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL

                                        http://www.acm.org/pubs/
                                        articles/journals/toplas/
                                        2000-22-3/p490-kistler/
                                        p490-kistler.pdf;   http:
                                        //www.acm.org/pubs/citations/
                                        journals/toplas/2000-22-
                                        3/p490-kistler/.

Karaorman:2005:JJR

[KF05] M. Karaorman and J. Freeman. jMonitor: Java runtime event specification and monitoring library. Electronic Notes in Theoretical Computer Science, 113(??):181– 200, 2005. CODEN ???? ISSN 1571-0661.

Khondkar:2004:AAI

[KFLN04] P. Khondkar, M. A. Fukase, C. D. Lima, and T. Nakamura. An architecture to allow intrinsic concurrency on Java processors for embedded systems. Information, 7(3): 351–366, 2004. CODEN ???? ISSN 1343-4500.

Khondkar:2004:EEB

[KFN04] P. Khondkar, M. A. Fukase, and T. Nakamura. Enhancing execution bandwidth of Javaembedded processor. Information, 7(2):215–226, 2004. CODEN ???? ISSN 1343-4500.

Kamalov:2005:JAT

[KGH^{+}05] R. Kamalov, M. Guillaud, D. Haskins, A. Harrison, R. Kemp, D. Chiu, M. Follen, and C. MacAulay. A Java application for tissue section image analysis. Computer Methods and Programs in Biomedicine, 77(2):99–113, 2005. CODEN CMPBEK. ISSN 0169-2607 (print), 1872-7565 (electronic).

Keen:2004:JFD

[KGMO04] A. W. Keen, T. Ge, J. T.

Maris, and R. A. Olsson. JR: Flexible distributed programming in an extended Java. ACM Transactions on Programming Languages and Systems, 26(3):578–608, 2004. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Kim:2000:MSB

[KH00] Jin-Soo Kim and Yarsun Hsu.

Memory system behavior of Java programs: methodology and analysis. In ACM [ACM00b], pages xi + 329. ISBN 1-58113-194-1. ISSN 0163-5999 (print), 1557-9484 (electronic). LCCN QA1 .A854724.

Kiczales:2001:AOP

[KH01] Gregor Kiczales and Erik Hilsdale. Aspect-oriented programming. ACM SIG-SOFT Software Engineering Notes, 26(5):313, September 2001. CODEN SFENDP.

ISSN 0163-5948 (print), 1943-

5843 (electronic).

Kielmann:2001:EJH

[KHBB01] Thilo Kielmann, Philip Hatcher, Luc Bougé, and Henri E. Bal. Enabling Java for high-performance computing. Communications of the ACM, 44(10):110–117, October 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Khoo:2009:DJA

[KHFS09] Yit Phang Khoo, Michael Hicks, Jeffrey S. Foster, and Vibha Sazawal. Directing JavaScript with arrows. ACM SIGPLAN Notices, 44(12):49–58, December 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kingsley-Hughes:2001:JE

[KHKH01] Adrian Kingsley-Hughes and Kathie Kingsley-Hughes. JavaScript 1.5 by example. By example series. Que Corporation, Indianapolis, IN, USA, 2001. ISBN 0-7897-2499-5. xii + 299 pp. LCCN QA76.73.J39 K55 2001.

Karlsson:2005:EPD

[KHMW05] M. Karlsson, E. Hagersten, K. E. Moore, and D. A. Wood. Exploring processor design options for Java-based middleware. In Skeie et al. [SY^{+}05], pages 59–68. ISBN 0-7695-2381-1. LCCN QA76.58; QA76.58 .I34 2005. URL

http://ieeexplore.ieee.
org/servlet/opac?punumber=

9965. IEEE Computer Society order number P2381.

Kiczales:2003:ATA

[Kic03] G. Kiczales. AspectJ(tm):

Aspect-oriented programming in Java. Lecture Notes in Computer Science, 2591: 1, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kiczales:2004:CLG

[Kic04] G. Kiczales. Crosscut: a little goes a long way: Attributes in C# and annotations in Java provide a hook for principled macros and metaprogramming in the C family. Lisp has had this technology for years, offering many lessons. Software Development, 12(5):52–53, 2004. CO-DEN ???? ISSN 1070-8588.

Kientzle:2001:JQH

[Kie01] Tim Kientzle. Java Q&A:

How can I speed up my JDBC-based programs? Dr. Dobb’s Journal of Software Tools, 26(12):111–112, December 2001. CO-DEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/ftp/2001/
2001_12/jqa1201.txt;   http:
//www.ddj.com/ftp/2001/
2001_12/jqa1201.zip.

Kientzle:2002:JQH

[Kie02] Tim Kientzle. Java Q&A: How can I make my Java servlets faster? Dr. Dobb’s Journal of Software Tools, 27 (4):97–99, April 2002. CO-DEN DDJOEB. ISSN 1044-789X. URL http:

                                        //www.ddj.com/ftp/2002/
                                        2002_04/jqa0402.txt;   http:
                                        //www.ddj.com/ftp/2002/
                                        2002_04/jqa0402.zip.

Kilgore:2002:OOS

[Kil02]  R. A. Kilgore.  Object-

oriented simulation with Java, Silk and OpenSML1.Net languages. Winter Simulation Conference, 1(??):227–233, 2002. CODEN WSCPDK. ISSN 0743-1902 (print), 2160-9276 (electronic).

Kilburn:2003:MUJ

[Kil03a] W. Kilburn. Microsoft unveils Java converter kit. Application Development Trends, 10 (12):27, 2003. CODEN ???? ISSN 1073-9564.

Kilgore:2003:OOS

[Kil03b]  R. A. Kilgore.  Object-

oriented simulation with SML and Silk in .Net and Java. Winter Simulation Conference, 1(??):218–224, 2003. CODEN ???? ISSN 0743-1902 (print), 2160-9276 (electronic).

Kim:2002:DIM

[Kim02] Jong-Hak Kim. Development of intelligent milling machine using Java tool: research project. Master of science, plan ii, Department of Mechanical Engineering, University of California, Berkeley, Berkeley, CA, USA, 2002.

King:2000:JP

[Kin00] K. N. (Kim N.) King. Java Programming. W. W. Norton & Co., New York, NY, USA, 2000. ISBN 0-393-97437-5. xxiii + 788 pp. LCCN QA76.73.J38 K56 2000. US$40.

Kim:2002:SOC

[KJ02] Do-Woo Kim and Min-Soo Jung. A study on the optimization of class file for Java card platform. Lecture Notes in Computer Science, 2343:563–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/
service/series/0558/bibs/
2343/23430563.htm;   http:
//link.springer.de/link/
service/series/0558/papers/
2343/23430563.pdf.

Kazi:2000:JCS

[KJBH^{+}00] I. H. Kazi, D. P. Jose, B. Ben-Hamida, C. J. Hescott, C. Kwok, J. A. Konstan, D. J. Lilja, and P.- C Yew. JaViz: a client/ server Java profiling tool. IBM Systems Journal, 39(1): 96–117, ???? 2000. CO-DEN IBMSA7. ISSN 0018-

                                        8670. URL http://www.
                                        almaden.ibm.com/journal/
                                        sj/391/kazi.html.

Koch:2000:AFG

[KK00] Michael Koch and Jürgen Koch. Application of frameworks in groupware — the Iris group editor environment. ACM Computing Surveys, 32 (1es), March 2000. CO-DEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/surveys/2000-32-
                                        1/p28-koch/.

Koga:2003:MRT

[KK03a] M. Koga and S. Kawakami. MAI-17-3 real-time remote control system in Java and its application to swing up control of inverted pendulum. Sice, 1:358–361, 2003. CO-DEN ????

Korochkin:2003:EPA

[KK03b] D. Korochkin and S. Korochkin. Experimental performance analysis of the Ada95 and Java parallel program on SMP systems. ACM SIGADA Ada Letters, 23(1):53–56, 2003. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Kaczmarek:2004:SEE

[KK04a] J. Kaczmarek and M. Kucharski. Size and effort estimation for applications written in Java. Information and Software Technology, 46(9):589– 601, 2004. CODEN ISOTE7. ISSN 0950-5849 (print), 1873-6025 (electronic).

Ko:2004:TCG

[KK04b] K. M. Ko and S. G. Kim.

Target code generation using the code expansion technique for Java bytecode. Lecture Notes in Computer Science, 3320:752–755, 2004. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klohs:2005:MRJ

[KK05] K. Klohs and U. Kastens.

Memory requirements of Java bytecode verification on limited devices. Electronic Notes in Theoretical Computer Science, 132(1):95–111, 2005. CODEN ???? ISSN 1571-0661.

Kumar:2009:GCM

[KKG09] Avadhesh Kumar, Rajesh Kumar, and P. S. Grover. Generalized coupling measure for aspect-oriented systems. ACM SIGSOFT Software Engineering Notes, 34(3):1–6, May 2009. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kouh:2004:DJP

[KKJY04] H. J. Kouh, K. T. Kim, S. M.

Jo, and W. H. Yoo. Debugging of Java programs using HDT with program slicing. Lecture Notes in Computer Science, 3046:524–533, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kulkarni:2004:VJS

[KKK04] P. Kulkarni, A. Kembhavi, and S. Kale. VOTable JAVA streaming writer and applications. Astronomical Society of the Pacific Conference Series, 314:346–349, 2004. CODEN ???? ISSN 1050-3390.

Kim:2004:JMRa

[KKL^{+}04] M. Kim, S. Kannan, I. Lee, O. Sokolsky, and M. Viswanathan. Java-MaC — A run-time assurance tool for Java programs. Electronic Notes in Theoretical Computer Science, 55(2):1–18, January 2004. CODEN ???? ISSN 1571-0661.

Kawahito:2006:NIR

[KKM^{+}06] Motohiro Kawahito, Hideaki Komatsu, Takao Moriyama, Hiroshi Inoue, and Toshio Nakatani. A new idiom recognition framework for exploiting hardware-assist instructions. ACM SIGPLAN Notices, 41(11):382–393, November 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kawahito:2000:ENP

[KKN00] Motohiro Kawahito, Hideaki Komatsu, and Toshio Nakatani. Effective null pointer check elimination utilizing hardware trap. ACM SIG-PLAN Notices, 35(11):139– 149, November 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kawahito:2006:ESE

[KKN06] Motohiro Kawahito, Hideaki Komatsu, and Toshio Nakatani. Effective sign extension elimination for Java. ACM Transactions on Programming Languages and Systems, 28(1): 106–133, January 2006. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Kawachiya:2002:LRJ

[KKO02] Kiyokuni Kawachiya, Akira Koseki, and Tamiya Onodera. Lock reservation: Java locks can mostly do without atomic operations. ACM SIG-PLAN Notices, 37(11):130– 141, November 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kumar:2003:PBD

[KKT04] C. Bala Kumar, Paul J. Kline, and Timothy J. Thompson. Bluetooth application programming with the Java APIs. The Morgan Kaufmann series in networking. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2004. ISBN 1-55860-934-2. xxii + 498 pp. LCCN TK5103.3 .K86 2004.

US$44.95. URL ftp://

                                        uiarchive.cso.uiuc.edu/
                                        pub/etext/gutenberg/;
                                        http://www.loc.gov/catdir/
                                        description/els051/2003107476.
                                        html;   http://www.loc.
                                        gov/catdir/toc/els051/2003107476.
                                        html.

Kiciman:2007:APR

[KL07] Emre Kiciman and Benjamin Livshits. AjaxScope: a platform for remotely monitoring the client-side behavior of Web 2.0 applications. Operating Systems Review, 41(6): 17–30, December 2007. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Klebanov:2005:JFN

[Kle05a]  V. Klebanov.  A JMM-

Faithful non-interference calculus for Java. Lecture Notes in Computer Science, 3409:101–111, 2005. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klein:2005:VJB

[Kle05b] G. Klein. Verified Java: Bytecode verification: Verified Java bytecode verification. It, 47(2):107–110, 2005. CODEN ???? ISSN 1611-2776.

Kou:2003:RST

[KLL03] Y. Kou, Z. Liao, and Z. Li. Research on the scalable technologies of network management software based on Java. Mini-Micro Systems, 24(1):17–19, 2003. CODEN XWJXEH. ISSN 1000-1220.

Kumar:2000:SAM

[KLS00] Bharath M. Kumar, R. Lakshminarayanan, and Y. N. Srikant. A study of automatic migration of programs across the Java event models. ACM SIGSOFT Software Engineering Notes, 25(3):24–29, May 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Krishna:2001:SRI

[KM01] Ksheerabdhi Krishna and Michael Montgomery. A simple(r) interface distribution mechanism for Java card. Lecture Notes in Computer Science, 2041:114–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2041/20410114.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2041/20410114.
pdf.

Ko:2002:CBA

[KM02] Ren-Song Ko and Matt W.

Mutka. A component-based approach for adaptive soft real-time Java within heterogeneous environments. Parallel and Distributed Computing Practices, 5(1):89–104, March 2002. CODEN ???? ISSN 1097-2803.

Khurshid:2004:CJI

[KM04a] S. Khurshid and D. Marinov. Checking Java implementation of a naming architecture using Testera. Electronic Notes in Theoretical Computer Science, 55(3):1– 21, January 2004. CODEN ???? ISSN 1571-0661.

Khurshid:2004:TSB

[KM04b] S. Khurshid and D. Marinov. TestEra specificationbased testing of Java programs using SAT. Automated Software Engineering, 11(4): 403–434, 2004. CODEN ???? ISSN 0928-8910.

Kortenkamp:2004:GTW

[KM04c] U. Kortenkamp and D. Materlik. Geometry teaching in wireless classroom environments using Java and J2ME. Science of Computer Programming, 53(1):71–85, 2004. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Koletzke:2007:OJF

[KM07] Peter Koletzke and Duncan Mills. Oracle JDeveloper 10g for forms and PL/SQL developers. McGraw-Hill, New York, NY, USA, 2007. ISBN 0-07-225960-4 (paperback), (paperback). xxx + 562 pp. LCCN QA76.73.J38 K655

                                        2006. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0806/2007298524-b.html;
                                         http://www.loc.gov/catdir/

enhancements/fy0806/2007298524-
d.html;   http://www.loc.
gov/catdir/enhancements/
fy0806/2007298524-t.html.

Kireev:2008:RTJ

[KM08] Alexandre N. Kireev and Olivier J. F. Martin. Realtime Java simulations of multiple interference dielectric filters. Computer Physics Communications, 179(12): 903–907, December 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465508002531.

Kim:2004:VJJ

[KMEA04] S. Kim, S. M. Moon, K. Ebcioglu, and E. Altman. VLaTTe: a Java just-in-time compiler for VLIW with fast scheduling and register allocation. IEICE Transactions on Information and Systems E Series D, 87(7):1712–1720, 2004. CODEN ???? ISSN 0916-8532.

Kimura:2003:IJA

[KMOS03] M. Kimura, M. H. Miki, T. Onoye, and I. Shirakawa. Implementation of Java accelerator for high-performance embedded systems. IEICE Transactions on Fundamentals of Electronics Communications and Computer Sciences E Series A, 86(12): 3079–3088, 2003. CODEN ???? ISSN 0916-8508 (print), 1745-1337 (electronic).

Kamin:2002:ICS

[KMR02] Samuel N. Kamin, M. Dennis Mickunas, and Edward M. Reingold. An introduction to computer science using Java. McGraw-Hill, New York, NY, USA, second edition, 2002. ISBN 0-07-232305-1. xxix + 753 pp. LCCN QA76 .K262 2001.

Kirkegaard:2004:SAX

[KMS04] C. Kirkegaard, A. Moller, and M. I. Schwartzbach. Static analysis of XML transformations in Java. IEEE Transactions on Software Engineering, 30(3):181–192, March 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        1271173.

Kimball:2008:CCW

[KMSB08] Aaron Kimball, Sierra Michels-Slettvet, and Christophe Bisciglia. Cluster computing for Web-scale data processing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):116–120, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Kistijantoro:2003:CRD

[KMSL03] A. Kistijantoro, G. Morgan, S. Shrivastava, and M. Little. Component replication in distributed systems: a case study using enterprise Java Beans. In IEEE [IEE03b], pages 89–98. CODEN ???? ISBN 0-7695-1955-5. ISSN 1060-9857. LCCN QA76.9.D5 S94 2003.

Klein:2006:MCM

[KN06] Gerwin Klein and Tobias Nipkow. A machine-checked model for a Java-like language, virtual machine, and compiler. ACM Transactions on Programming Languages and Systems, 28(4): 619–695, July 2006. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Kumar:2002:DPP

[KNG02] R. V. Kumar, B. L. Narayanan, and R. Govindarajan. Dynamic path profile aided recompilation in a Java justin-time compiler. Lecture Notes in Computer Science, 2552:495–505, 2002. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koved:2001:SCE

[KNN^{+}01] L. Koved, A. Nadalin, N. Nagaratnam, M. Pistoia, and T. Shrader. Security challenges for Enterprise Java in an e-business environment. IBM Systems Journal, 40(1): 130–152, ???? 2001. CODEN IBMSA7. ISSN 0018-8670.

URL http://www.research.

                                        ibm.com/journal/sj/401/
                                        koved.html.

Knoernschild:2002:JDO

[Kno02] Kirk Knoernschild. Java design: objects, UML, and process. Addison-Wesley, Reading, MA, USA, 2002. ISBN 0-201-75044-9 (paperback). xxiv + 277 pp. LCCN QA76.73.J38 K57 2002 Bar.

Knuckles:2001:IIP

[Knu01a] Craig D. Knuckles. Introduction to interactive programming on the Internet: using HTML and JavaScript. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 2001. ISBN 0-471-38366-X (paperback). xviii + 423 pp. LCCN QA76.625 .K58 2001.

Knudsen:2001:WJD

[Knu01b] Jonathan Knudsen. Wireless Java: developing with Java 2, micro edition. Books for professionals by professionals. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 1-893115-50-X. xiii + 226 pp. LCCN QA76.73.J38 K58 2001.

Kloukinas:2003:MTS

[KNY03] C. Kloukinas, C. Nakhli, and S. Yovine. A methodology and tool support for generating scheduled native code for real-time Java applications. Lecture Notes in Computer Science, 2855:274–289, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kambites:2001:OLI

[KOB01] M. E. Kambites, J. Obdržálek, and J. M. Bull. An OpenMP-like interface for parallel programming in Java. Concurrency and Computation: Practice and Experience, 13(8–9):793–814, July/August 2001. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (elec-

tronic). URL http://www3.
interscience.wiley.com/
cgi-bin/abstract/84503220/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=84503220&PLACEBO=IE.
pdf.

Kodaganallur:2004:ILP

[Kod04] Viswanathan Kodaganallur.

Incorporating language processing into Java applications: a JavaCC tutorial. IEEE Software, 21(4):70–77, July/August 2004. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (elec-

tronic). URL http://csdl.
computer.org/comp/mags/
so/2004/04/s4070s.pdf.

Koga:2004:CAT

[Kog04] M. Koga. Computer aided total development of control system — control system design in Java. Systems, Control and Information = Shisutemu Seigyo Joho Gakkai shi, 48 (4):138–143, 2004. CODEN SYCNA9. ISSN 0916-1600.

Konsella:2003:ASJ

[Kon03] S. Konsella. Adapting standard Java GUI APIs for front panel user interfaces on peripheral devices. Research Disclosure, 466:236, 2003. CODEN RSDSBB. ISSN 0374-4353.

Kong:2004:IDI

[Kon04] J. Kong. The implementation of Dicom interface software using Java. Chinese Journal of Biomedical Engineering, 23(1):10–14, 2004. CO-DEN ???? ISSN 0258-8021.

Kawachiya:2008:ARM

[KOO08] Kiyokuni Kawachiya, Kazunori Ogata, and Tamiya Onodera. Analysis and reduction of memory inefficiencies in Java strings. ACM SIGPLAN Notices, 43(10):385–402, September 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kuo:2001:AAJ

[KP01] Dean Kuo and Doug Palmer. Automated analysis of Java message service providers. Lecture Notes in Computer Science, 2218:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.
                                        com/link/service/series/
                                        0558/bibs/2218/22180001.

htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2218/22180001.
pdf.

Kermany:2006:CCI

[KP06] Haim Kermany and Erez Petrank. The Compressor: concurrent, incremental, and parallel compaction. ACM SIGPLAN Notices, 41(6): 354–363, June 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kalibera:2009:CBV

[KPH^{+}09] Tomas Kalibera, Pavel Parizek, Ghaith Haddad, Gary T. Leavens, and Jan Vitek. Challenge benchmarks for verification of real-time programs: (abstract only). ACM SIGPLAN Notices, 44(11):7– 8, November 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Koved:2002:ARA

[KPK02] Larry Koved, Marco Pistoia, and Aaron Kershenbaum. Access rights analysis for Java. ACM SIGPLAN Notices, 37(11):359–372, November 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kavadias:2003:ESS

[KPKL03] C. Kavadias, B. Perrin, V. Kollias, and M. Loupis.

Enhanced SDL subset for the design and implementation of Java-enabled embedded signalling systems. Lecture Notes in Computer Science, 2708:137–149, 2003. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kurtz:2002:EIE

[KPN02] Barry L. Kurtz, Dee Parks, and Eric Nicholson. Effective Internet education: a progress report. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):312– 316, March 2002. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Kaiser:2006:CJC

[KPPÉR06] Claude Kaiser, Jean-François Pradat-Peyre, Sami Évangelista, and Pierre Rousseau. Comparing Java, C# and Ada monitors queuing policies: a case study and its Ada refinement. ACM SIGADA Ada Letters, 26(2):23–37, August 2006. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Kolling:2000:OFJ

[KR00] Michael Kölling and John Rosenberg. Objects first with Java and BlueJ (seminar session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):429, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Knoblock:2001:TES

[KR01a] Todd B. Knoblock and Jakob Rehof. Type elaboration and subtype completion for Java bytecode. ACM Transactions on Programming Languages and Systems, 23(2): 243–272, March 2001. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Kolling:2001:GTO

[KR01b] Michael Kölling and John Rosenberg. Guidelines for teaching object orientation with Java. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):33–36, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kleijnen:2003:OWS

[KR03] Stans Kleijnen and Srikanth Raju. An open Web services architecture. ACM Queue: Tomorrow’s Computing Today, 1(1):38–46, March 2003. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Kreger:2001:JME

[Kre01] H. Kreger. Java Management Extensions for application management. IBM Systems Journal, 40(1):104– 129, ???? 2001. CODEN IBMSA7. ISSN 0018-8670.

URL http://www.research.
ibm.com/journal/sj/401/
kreger.html.

Kroeker:2000:PCL

[Kro00a]  Kirk L. Kroeker.  Prod-

ucts: Controlling the look and feel of Windows; sophisticated project estimation for $99; Web applications for the Internet-challenged; network-based digital asset manager; Linux middleware framework; XML-based information portal; Java development for small computing devices; build installations for Windows 2000. Computer, 33(3):105–107, March 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//dlib.computer.org/co/
books/co2000/pdf/r3105.
pdf;   http://www.stardock.
net/.

Kroeker:2000:PEN

[Kro00b]  Kirk L. Kroeker.  Prod-

ucts: Enterasys Networks’ Ecommerce access platforms; Tascom Software’s ASP editor; Vital’s text editor for program development; RapidStream’s security appliance; Kemma Software’s help desk software; Telelogic’s real-time UML profiling software; ParaSoft extends product support to Windows 2000; Spyglass’ interactive TV software; Metrowerks releases CodeWarrior with PersonalJava support. Computer, 33(6):96–98, June 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//dlib.computer.org/co/
books/co2000/pdf/r6096.
pdf.

Klemm:2001:EJS

[KS01a] Reinhard Klemm and Navjot Singh. Enhancing Java server availability with JAS. Software—Practice and Experience, 31(7):697–716, June 2001. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

http://www3.interscience.
wiley.com/cgi-bin/abstract/
78003146/START;   http:/
/www3.interscience.wiley.
com/cgi-bin/fulltext?ID=
78003146&PLACEBO=IE.pdf.

Kurzyniec:2001:FCL

[KS01b] Dawid Kurzyniec and Vaidy Sunderam. Flexible class loader framework: Sharing Java resources in harness system. Lecture Notes in Computer Science, 2073:375–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.

                                        com/link/service/series/
                                        0558/bibs/2073/20730375.
                                        htm;   http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2073/20730375.
                                        pdf.

Kozen:2002:ECI

[KS02a] Dexter Kozen and Matt Stillerman. Eager class initialization for Java. Lecture Notes in Computer Science, 2469:71–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/
                                        service/series/0558/bibs/
                                        2469/24690071.htm;   http:
                                        //link.springer.de/link/
                                        service/series/0558/papers/
                                        2469/24690071.pdf.

Kurzyniec:2002:MBT

[KS02b] D. Kurzyniec and V. Sunderam. The MICROBE benchmarking toolkit for Java: a component-based approach. Lecture Notes in Computer Science, 2330:709– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.
                                        com/link/service/series/
                                        0558/bibs/2330/23300709.
                                        htm;   http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2330/23300709.
                                        pdf.

Kozlenkov:2004:PRB

[KS04] A. Kozlenkov and M. Schroeder.

PROVA: Rule-based Javascripting for a Bioinformatics Semantic Web. Lecture Notes in Computer Science, 2994:17–30, 2004. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kuehne:2007:CPL

[KS07] Thomas Kuehne and Daniel Schreiber. Can programming be liberated from the two-level style: multi-level programming with Deep-Java. ACM SIGPLAN Notices, 42(10):229–244, October 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kaur:2009:VMC

[KS09] Parminder Kaur and Hardeep Singh. Version management and composition of software components in different phases of software development life cycle. ACM SIG-SOFT Software Engineering Notes, 34(4):1–9, July 2009. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Kautz:2000:LLI

[KSC^{+}00] Frederick Kautz, Dimitrios Souflis, Robert Carbonari, DDJ, Al Gettier, Tina Femia, Andrew D. Todd, Anonymous, Alan Horn, Anonymous, Brad Clawsie, and Samuel Pearlman. Letters:

The S language; IP security; Pnuts pointer; flyby; worker shortage?; design by interface. Dr. Dobb’s Journal of Software Tools, 25(4): 10, 12, April 2000. CODEN DDJOEB. ISSN 1044-789X. See [Ste00].

Kaiya:2004:MDF

[KSK04a] H. Kaiya, K. Sasaki, and K. Kaijiri. A method to develop feasible requirements for Java mobile code application. IEICE Transactions on Information and Systems E Series d, 87(4):811–821, 2004. CODEN ???? ISSN 0916-8532.

Krishna:2004:ERT

[KSK04b] A. Krishna, D. Schmidt, and R. Klefstad. Enhancing Real-Time CORBA via Real-Time Java features. International Conference on Distributed Computing Systems, 24(??):66–73, 2004. CODEN ???? ISSN 1063-6927.

Kassem:2000:DEA

[KT00] Nicholas Kassem and The Enterprise Team. Designing Enterprise applications with the Java 2 Platform. Addison-Wesley, Reading, MA, USA, 2000. ISBN 0-201-70277-0. xxi + 341 pp. LCCN QA76.76.A65 K37 2000. US$35.95.

Kniesel:2001:JAR

[KT01a] Günter Kniesel and Dirk Theisen. JAC — Access right based encapsulation for Java. Software—Practice and Experience, 31(6): 555–576, May 2001. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (elec-

tronic). URL http://www3.
interscience.wiley.com/
cgi-bin/abstract/78003102/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=78003102&PLACEBO=IE.
pdf.

Krall:2001:JLS

[KT01b] A. Krall and P. Tomsich. Java for large-scale scientific computations? Lecture Notes in Computer Science, 2179:228– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2179/21790228.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2179/21790228.
pdf.

Kamina:2004:MDI

[KT04] T. Kamina and T. Tamai.

McJava — A design and implementation of Java with mixin-types. Lecture Notes in Computer Science, 3302:398– 414, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kim:2004:EEJ

[KTV^{+}04] S. Kim, S. Tomar, N. Vijaykrishnan, M. Kandemir, and M. J. Irwin. Energyefficient Java execution using local memory and object co-location. IEE Proceedings. Computers and Digital Techniques, 151(1):33–42, 2004. CODEN ICDTEA. ISSN 1350-2387 (print), 1359-7027 (electronic).

Kuc:2006:ROS

[Kuc06]  Bernard Kuc.  Review

of “Open Source Wireless Java Tools Suite by Robert Virkus,” Apress, 2005, $39.99, ISBN 1-59059-503-3. ACM Queue: Tomorrow’s Computing Today, 4(4):48, May 2006. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). See [Vir05].

Kumaran:2001:JTO

[Kum01] S. Ilango Kumaran. Jini technology: an overview. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2001. ISBN 0-13-033385-9. 358 pp. LCCN QA76.9.D5 K83 2002.

Kumaran:2002:JTO

[Kum02] S. Ilango Kumaran. Jini Technology: An Overview. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2002. ISBN 0-13-033385-9. xix + 358 pp. LCCN QA76.9.D5 K83 2002. US$34.99. URL http:/

                                        /www.phptr.com/ptrbooks/
                                        ptr_0130333859.html.

Kumar:2004:WBT

[Kum04] A. Kumar. Web-based tutors for learning programming in C++/Java. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):266, 2004. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kumar:2005:OTC

[Kum05] Amruth N. Kumar. Online tutors for C++/Java programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):387, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kunkle:2002:WBI

[Kun02] Wanda Kunkle. A Web-based integral evaluator: a demonstration of the successful integration of WebEQ, Maple, and Java. In Anonymous [Ano02i], page ?? ISBN ???? LCCN ????

Kurniawan:2004:JFP

[Kur04] Budi Kurniawan. JavaServer Faces programming. McGraw-Hill/Osborne, New York, NY, USA, 2004. ISBN 0-07-222983-7. xxii + 503 pp. LCCN TK5105.8885.J38 K87

2004. URL http://www.loc.
gov/catdir/description/
mh051/2004303988.html;
http://www.loc.gov/catdir/
toc/mh051/2004303988.html.

Kim:2004:JMRb

[KVK^{+}04] M. Kim, M. Viswanathan, S. Kannan, I. Lee, and O. Sokolsky. Java-MaC: a run-time assurance approach for Java programs. Formal Methods in System Design, 24(2):129–155, March 2004. CODEN FMSDE6. ISSN 0925-9856. Special Issue on Selected Papers from the First International Workshop on Runtime Verification Held in Paris, July 2001 (RV01).

Koffman:2001:SJP

[KW01a] Elliot Koffman and Ursula Wolz. A simple Java package for GUI-like interactivity. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33 (1):11–15, March 2001. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Krintz:2001:UJC

[KW01b] Chandra Krintz and Rich Wolski. Using JavaNws to compare C and Java TCP-Socket performance. Concurrency and Computation: Practice and Experience, 13(8–9):815–839, July/August 2001. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (elec-

tronic). URL http://www3.
interscience.wiley.com/
cgi-bin/abstract/84503224/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?

ID=84503224&PLACEBO=IE.
pdf.

Komodromos:2002:UJD

[KW02] P. I. Komodromos and J. R.

Williams. Utilization of Java and database technology in the development of a combined discrete and finite element multibody dynamics simulator. Geotechnical Special Publication, 117:118–124, 2002. CODEN ???? ISSN 0895-0563.

Klein:2003:VBS

[KW03] G. Klein and M. Wildmoser. Verified bytecode subroutines. Journal of Automated Reasoning, 30(3–4): 363–398, May 2003. CO-DEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A%3A1025095122199.

Kwon:2003:AJP

[KWK03] Jagun Kwon, Andy Wellings, and Steve King. Assessment of the Java programming language for use in high integrity systems. ACM SIGPLAN Notices, 38(4):34–46, April 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kwon:2005:RJH

[KWK05] Jagun Kwon, Andy Wellings, and Steve King. Ravenscar– Java: a high integrity profile for real-time Java. Concurrency and Computation: Practice and Experience, 17 (5–6):681–714, April/May 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kotzmann:2008:DJH

[KWM^{+}08] Thomas Kotzmann, Christian Wimmer, Hanspeter Mössenböck, Thomas Rodriguez, Kenneth Russell, and David Cox. Design of the Java HotSpot^{T M} client compiler for Java 6. ACM Transactions on Architecture and Code Optimization, 5(1):7:1– 7:??, May 2008. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Kurniawan:2004:CSW

[KX04] B. Kurniawan and J. Xue. A comparative study of Web application design models using the Java technologies. Lecture Notes in Computer Science, 3007:711–721, 2004. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kouh:2003:ADJ

[KY03a] H. J. Kouh and W. H. Yoo. Automated debugging in Java programs using HDM. Lecture Notes in Computer Science, 2660:849–858, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kouh:2003:EDS

[KY03b] H. J. Kouh and W. H. Yoo.

The efficient debugging system for locating logical errors in Java programs. Lecture Notes in Computer Science, 2667:684–693, 2003. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lyon:2000:LWS

[LAB^{+}00] Douglas Lyon, Roger T.

Alexander, James M. Bieman, John Viega, Bruce D. Wilner, Roger T. Alexander, James M. Bieman, John Viega, Derek B. Eckert, Ted Lewis, Howard C. Anderson, Thomas M. English, and Timothy C. Lethbridge. Letters: We shape our tools and our tools shape us; Java design flaws create programming stress; funny money’s role in the government’s favorite hobby; rigorous training teaches students how to think. Computer, 33(7):4–7, July 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//dlib.computer.org/co/
books/co2000/pdf/r7004.
pdf.

Labouseur:2009:BBO

[Lab09]  Alan G. Labouseur.  A

browser-based operating systems project: JavaScript adventures in dinosaur slaying.

SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41 (4):71–75, December 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ladd:2001:PEU

[Lad01] Eric Ladd. Platinum edition using XHTML, XML and Java 2. Que Corporation, Indianapolis, IN, USA, platinum edition, 2001. ISBN 0-7897-2473-1. xxiv + 1413 pp. LCCN QA76.76.H94 P585 2001.

Lagorio:2003:TSC

[Lag03]  G. Lagorio.  Towards a

smart compilation manager for Java. Lecture Notes in Computer Science, 2841:302– 315, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lau:2006:OPA

[LAHC06] Jeremy Lau, Matthew Arnold, Michael Hind, and Brad Calder. Online performance auditing: using hot optimizations without getting burned. ACM SIGPLAN Notices, 41(6):239–251, June 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Laird:2001:JQW

[Lai01] Cameron Laird. Java Q&A: When should you script Java with TCL? Dr. Dobb’s Journal of Software Tools, 26(3):133–134, March 2001. CODEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/ftp/2001/
2001_03/jqa0301.txt.

Lai:2003:JPW

[Lai03] Ray Lai. J2EE Platform Web Services. Sun Microsystems Press, Palo Alto, CA, USA, 2003. ISBN 0-13-101402-1. 592 (est.) pp. LCCN TK5105.88813 .L35 2003. US$49.99.

Lai:2008:JIA

[Lai08] Charlie Lai. Java insecurity:

Accounting for subtleties that can compromise code. IEEE Software, 25(1):13–19, January/February 2008. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Lakshman:2002:OJD

[Lak02]  Bulusu Lakshman.  Ora-

cle and Java development. Howard W. Sams, Indianapolis, IN 46268, USA, 2002. ISBN 0-672-32117-3. xv + 438 pp. LCCN QA76.9.D3 L35 2002.

Lobosco:2002:JHP

[LAL02] M. Lobosco, C. Amorim, and O. Loques. Java for highperformance network-based computing: a survey. Concurrency and Computation: Practice and Experience, 14 (1):1–31, January 2002. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (elec-

                                        tronic). URL http://www3.
                                        interscience.wiley.com/
                                        cgi-bin/abstract/91014114/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=91014114&PLACEBO=IE.
                                        pdf.

Lamm:2003:BAV

[Lam03] E. Lamm. Booch’s Ada vs. Liskov’s Java: Two approaches to teaching software design. Lecture Notes in Computer Science, 2655:102– 112, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Langr:2000:EJS

[Lan00] Jeff Langr. Essential Java style: patterns for implementation. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2000. ISBN 0-13-085086-1. xxi + 277 pp. LCCN QA76.13.J38 L357 2000.

Laneve:2002:TSJ

[Lan02] Cosimo Laneve. A type system for JVM threads. Theoretical Computer Science, 290 (1):741–778, October 2002. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Langr:2004:TCS

[Lan04]  J. Langr.  Tech center:

Sweet and simple Java 1.5’s new features-including enhanced for loops, autoboxing, variable arguments, typesafe enum, static import and metadata-make coding easier and more enjoyable to write. Part 3 of 3. Software Development, 12(5):48–51, 2004. CO-DEN ???? ISSN 1070-8588.

Langridge:2005:DUM

[Lan05a] Stuart Langridge. DHTML utopia: modern Web design using JavaScript and DOM. SitePoint, Collingwood, VIC, Australia, 2005. ISBN 0-9579218-9-6. xiv + 318 pp. LCCN TK5105.888 .L36 2005.

Lano:2005:ASD

[Lan05b] K Lano. Advanced systems design with Java, UML, and MDA. Elsevier Butterworth-Heinemann, Amsterdam, The Netherlands, 2005. ISBN 0-7506-6496-7. viii + 378 pp. LCCN QA76.76.D47 L37 2005; QA76.76.D47 L375 2005.

Larsen:2001:JPB

[Lar01] Albert L. Larsen. Java programming: from the beginning. ACM SIGSOFT Software Engineering Notes, 26 (1):97–98, January 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Laszlo:2002:OOP

[Las02] Michael Jay Laszlo. Objectoriented programming featuring graphical applications in Java. Addison-Wesley, Reading, MA, USA, 2002. ISBN 0-201-72627-0 (paperback). xii + 468 pp. LCCN QA76.73.C153 L396 2002.

Lim:2004:IAW

[LAT04] B. Lim, S. R. Ajjarapu, and K. Thummala. Interfacing with Amazon Web services using Java and .NET: a comparative study. Journal of Internet Commerce, 3(4):19–42, 2004. CODEN ???? ISSN 1533-2861.

Laure:2001:OJF

[Lau01]  Erwin Laure.  OpusJava:

A Java framework for distributed high performance computing. Future Generation Computer Systems, 18 (2):235–251, October 2001. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http://

                                        www.elsevier.com/gej-ng/
                                        10/19/19/60/31/31/abstract.
                                        html.

Lau:2003:TSS

[Lau03] F. C. M. Lau. Towards a single system image for high-performance Java. Lecture Notes in Computer Science, 2745:6–7, 2003. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lau:2004:NLJ

[Lau04] H. T. (Hang Tong) Lau. A numerical library in Java for scientists and engineers. Chapman and Hall/CRC, Boca Raton, FL, USA, 2004. ISBN 1-58488-430-4. xxiii + 1063 pp. LCCN QA76.73.J38 L363

2004. URL http://www.loc.
gov/catdir/enhancements/
fy0646/2003055149-d.html.

Lawton:2002:MJM

[Law02]  George Lawton.  Moving

Java into mobile phones. Computer, 35(6):17–20, June 2002. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//csdl.computer.org/dl/
mags/co/2002/06/r6017.htm;
http://csdl.computer.org/
dl/mags/co/2002/06/r6017.
pdf.

Lazic:2007:BRBa

[Laz07] S. E. Lazic. Book review:

Correspondence Analysis and Data Coding with Java and R, by F. Murtagh. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(2):509, March 2007. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (elec-

tronic). URL http://www.
jstor.org/stable/4623181.

Lewis:2000:MPJ

[LB00] Bil Lewis and Daniel J. Berg.

Multithreaded Programming with Java Technology. Sun BluePrints Program. Sun Microsystems Press, Palo Alto, CA, USA, 2000. ISBN 0-13-017007-0. xxv + 461 pp. LCCN QA76.73.J38 L488 2000. US$39.99.

URL http://www.sun.com/
books/catalog/lewis3/index.
html.

Lawhead:2003:LJP

[LBD^{+}03] P. Lawhead, C. Bland, M. Duncan, M. Schep, D. Barnes, and M. Goldweber. Legos, Java, programming assignments and CS1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):47–48, 2003. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Li:2002:RBA

[LBJ02] T. Li, R. Bhargava, and L. K. John. Rehashable BTB: An adaptive branch target buffer to improve the target predictability of Java code. Lecture Notes in Computer Science, 2552:597–608, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Li:2005:ABT

[LBJ05] Tao Li, Ravi Bhargava, and Lizy Kurian John. Adapting branch-target buffer to improve the target predictability of Java code. ACM Transactions on Architecture and Code Optimization, 2(2):109– 130, June 2005. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Langtangen:2000:AST

[LBQ00] Hans Petter Langtangen, Are Magnus Bruaset, and Ewald Quak, editors. Advances in Software Tools for Scientific Computing, volume 10 of Lecture Notes in Computational Science and Engineering. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. CO-DEN LNCSA6. ISBN 3-540-66557-9 (softcover), 3-642-57172-7 (e-book). ISSN 1439-7358. LCCN QA76.6 .A336 2000. URL http:/

/link.springer.com/book/
10.1007/978-3-642-57172-
5;   http://www.springerlink.
com/content/978-3-642-57172-

5. Papers from an International Workshop on Modern Software Tools for Scientific Computing (Sci-Tools’98), Oslo, Norway, September 14–16, 1998.

Laufer:2000:SSC

[LBR00] Konstantin Läufer, Gerald Baumgartner, and Vincent F. Russo. Safe structural conformance for Java. The Computer Journal, 43(6): 469–481, ???? 2000. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

www3.oup.co.uk/computer_
journal/hdb/Volume_43/Issue_
06/430469.sgm.abs.html;
 http://www3.oup.co.uk/
computer_journal/hdb/Volume_

                                        43/Issue_06/pdf/430469.
                                        pdf.

Leavens:2006:PDJ

[LBR06] Gary T. Leavens, Albert L. Baker, and Clyde Ruby. Preliminary design of JML: a behavioral interface specification language for Java. ACM SIGSOFT Software Engineering Notes, 31(3):1–38, May 2006. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Lu:2004:DIM

[LC04] E. J. Lu and Y. Y. Cheng. Design and implementation of a mobile database for Java phones. Computer Standards and Interfaces, 26 (5):401–410, 2004. CO-DEN CSTIEZ. ISSN 0920-5489 (print), 1872-7018 (electronic).

Lee:2005:DDR

[LC05] Yueh-Feng Lee and Ruei-Chuan Chang. Developing dynamic-reconfigurable communication protocol stacks using Java. Software—Practice and Experience, 35(6): 601–620, May 2005. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lublinerman:2009:PPO

[LCC09] Roberto Lublinerman, Swarat Chaudhuri, and Pavol Cerny. Parallel programming with object assemblies. ACM SIG-PLAN Notices, 44(10):61– 80, October 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lim:2005:CCH

[LCFkL05] Sang Boem Lim, Bryan Carpenter, Geoffrey Fox, and Han ku Lee. Collective communication for the HPJava programming language. Concurrency and Computation: Practice and Experience, 17(7–8):867–894, June/July 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lee:2004:HJP

[LCFL04] H. K. Lee, B. Carpenter, G. Fox, and S. B. Lim. HP Java: programming support for high-performance Gridenabled applications. Parallel Algorithms and Applications, 19(2/3):175–193, 2004. CO-DEN PAAPEC. ISSN 1063-7192.

Lin:2003:SRP

[LCHY03] Jin Lin, Tong Chen, Wei-Chung Hsu, and Pen-Chung Yew. Speculative register promotion using advanced load address table (ALAT). In ACM [ACM03a], pages 125– 134. ISBN 0-7695-1913-X. ACM Order No.: 530033.

Li:2004:FRT

[LCS04] S. Q. Li, H. Y. Chen, and Y. X. Su. A framework of reachability testing for Java multithread programs. IEEE International Conference on Systems Man and Cybernetics, 3:2730–2734, 2004. CO-DEN ???? ISSN 1062-922X.

Li:2004:WAS

[LCZ04] H. Li, J. Cai, and R. Zhu. Wireless applications service platform based on XML and Java. Journal- Zhejiang University of Technology, 32(2): 182–185, 2004. CODEN ???? ISSN 1006-4303.

Locke:2003:JTC

[LD03] C. D. Locke and P. C. Dibble. Java technology comes to real-time applications (invited paper). Proceedings of the IEEE, 91(7):1105–1113, 2003. CODEN IEEPAD. ISSN 0018-9219 (print), 1558-2256 (electronic).

Lawhead:2003:RMT

[LDB^{+}03] Pamela B. Lawhead, Michaele E. Duncan, Constance G. Bland, Michael Goldweber, Madeleine Schep, David J. Barnes, and Ralph G. Hollingsworth. A road map for teaching introductory programming using LEGO^{T M} Mindstorms robots. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(2):191–201, June 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp:

//ftp.math.utah.edu/pub/
mirrors/ftp.ira.uka.de/

bibliography/Misc/DBLP/
2003.bib;   ftp://ftp.math.
utah.edu/pub/mirrors/ftp.
ira.uka.de/bibliography/
Techreports/UKent.bib;
http://www.cs.kent.ac.uk/
pubs/2003/1666.

Leavens:2002:FTJ

[LDE^{+}02] Gary T. Leavens, Sophia Drossopoulou, Susan Eisenbach, Arnd Poetzsch-Heffter, and Erik Poll. Formal techniques for Java programs. Lecture Notes in Computer Science, 2323:30– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2323/23230030.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2323/23230030.
pdf.

Lindquist:2004:JCS

[LDM04] T. Lindquist, M. Diarra, and B. Millard. A Java cryptography service provider implementing one-time pad (INIDP04). Proceedings of the Annual Hawaii International Conference on System Sciences, CONF37:189, 2004. CODEN ???? ISSN 1060-3425.

Lea:2000:CPJ

[Lea00a] Douglas Lea. Concurrent programming in Java design principles and patterns. Java series. Addison-Wesley, Reading, MA, USA, second edition, 2000. ISBN 0-201-31009-0. ix + 411 pp. LCCN QA76.73.J38 L4 2000.

Lear:2000:NBY

[Lea00b] Anne C. Lear. News briefs:

Y2K rollover: Few problems, many questions; Sun withdraws Java from standards process; wireless net has more potential than substance. Computer, 33 (2):22–24, February 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

                                        //dlib.computer.org/co/
                                        books/co2000/pdf/r2022.
                                        pdf.

Lea:2002:HEE

[Lea02] Rodger Lea. HAVi: example by example: Java programming for home entertainment devices. Prentice Hall PTR example by example series. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2002. ISBN 0-13-060035-0. xxix + 451 pp. LCCN TK7881.3 .H39 2002.

Lea:2005:JUC

                          [Lea05]  Doug Lea. The java.util.concurrent

synchronizer framework. Science of Computer Programming, 58(3):293–309, December 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Lee:2003:MWS

[Lee03] Arthur H. Lee. A manageable Web software architecture: searching for simplicity. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35 (1):229–233, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lehrbaum:2001:FESi

[Leh01] Rick Lehrbaum. Focus on embedded systems: Embedded Linux at JavaOne. Linux Journal, 89:48–49, September 2001. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Lehrbaum:2002:FESb

[Leh02] Rick Lehrbaum. Focus on embedded systems: Embedded Linux and Java — wave of the future? Linux Journal, 94: 54, 56, February 2002. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Lerner:2001:FEJ

[Ler01a]  Reuven M. Lerner.  At

the forge: Enterprise JavaBeans. Linux Journal, 92: 20–22, 24–25, December 2001. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http:

//noframes.linuxjournal.
com/lj-issues/issue92/article.
php?sid=5497.

Lerner:2001:FJ

[Ler01b] Reuven M. Lerner. At the forge: JavaBeans. Linux Journal, 86:18, 20–22, June 2001. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http:

                                        //noframes.linuxjournal.
                                        com/lj-issues/issue86/4686.

Lerner:2001:FJP

[Ler01c]  Reuven M. Lerner.  At

the forge: JavaServer pages. Linux Journal, 85:16, 18, 20–22, May 2001. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http:/

                                        /noframes.linuxjournal.
                                        com/lj-issues/issue85/4639.

Lerner:2001:FSS

[Ler01d]  Reuven M. Lerner.  At

the forge: Server-side Java with Jakarta-Tomcat. Linux Journal, 84:50, 52–54, 56– 58, April 2001. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Leroy:2001:JBV

[Ler01e] Xavier Leroy. Java bytecode verification: An overview. Lecture Notes in Computer Science, 2102:265–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.
                                        com/link/service/series/
                                        0558/bibs/2102/21020265.

htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2102/21020265.
pdf.

Leroy:2001:CBV

[Ler01f] Xavier Leroy. On-card bytecode verification for Java card. Lecture Notes in Computer Science, 2140:150–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2140/21400150.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2140/21400150.
pdf.

Leroy:2002:BVJ

[Ler02]  Xavier Leroy.  Bytecode

verification on Java smart cards. Software—Practice and Experience, 32(4):319– 340, April 10, 2002. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (elec-

tronic). URL http://www3.
interscience.wiley.com/
cgi-bin/abstract/91016433/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=91016433&PLACEBO=IE.
pdf.

Leroy:2003:JBV

[Ler03] Xavier Leroy. Java bytecode verification: Algorithms and formalizations. Journal of Automated Reasoning, 30(3– 4):235–269, May 2003. CO-DEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1023/A%3A1025055424017.

Leska:2003:LDG

[Les03] C. Leska. Learning to develop GUIs in Java using closed labs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):228, 2003. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lewis:2000:CEJ

[Lew00] Kevin Lewis. Creating Effective JavaHelp. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2000. ISBN 1-56592-719-2. xiii + 171 pp. LCCN QA76.73.J38 L495 2000. US$30.

Loy:2002:JS

[LEW^{+}02] Marc Loy, Robert Eckstein, David Wood, James E. lliott, and Brian Cole. Java Swing. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2002. ISBN 0-596-00408-7. xxiv + 1252 pp. LCCN QA76.73.J38 J3855 2003. US$54.95. URL

                                        http://www.oreilly.com/
                                        catalog/jswing2.

Loy:2003:JS

[LEW^{+}03] Marc Loy, Robert Eckstein, David Wood, James Elliott, and Brian Cole, editors. Java Swing. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2003. ISBN 0-596-00408-7. xxiv + 1252 pp. LCCN QA76.73.J38 J3855 2003. US$54.95. URL

http://www.oreilly.com/
catalog/9780596004088;
http://www.oreilly.com/
catalog/jswing2.

Lex:2002:EVN

[Lex02] M. Lex. Einführung, Vorteil und Nützen eines JAVAbasierten Betriebsmanagementsystems am Beispiel einer Müllverbrennungsanlage. (German) Introduction, advantages, and use of a Javabased business management system with example of a waste incinerator. VDI Berichte, 1711:23–32, 2002. CODEN ???? ISSN 0083-5560.

Lujan:2000:OOO

[LFG00] Mikel Luján, T. L. Freeman, and John R. Gurd. OoLALA: an object oriented analysis and design of numerical linear algebra. ACM SIGPLAN Notices, 35(10):229–252, October 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
proceedings/oops/353171/
p229-lujan/.

Lun:2003:OOP

[LFH03] Mok Pak Lun, Anthony Fong, and Gary K. W. Hau. Object-oriented processor requirements with instruction analysis of Java programs. ACM SIGARCH Computer Architecture News, 31(5):10– 15, December 2003. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Lemos:2009:ITO

[LFM09] Otávio Augusto Lazzarini Lemos, Ivan Gustavo Franchin, and Paulo Cesar Masiero. Integration testing of Object-Oriented and Aspect-Oriented programs: a structural pairwise approach for Java. Science of Computer Programming, 74(10):861–878, August 15, 2009. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Li:2004:MSJ

[LFP04] B. X. Li, X. C. Fan, and J. Pang. A model for slicing JAVA programs hierarchically. Journal of Computer Science and Technology, 19 (6):848–858, 2004. CODEN JCTEEM. ISSN 1000-9000.

Larman:1999:JPI

[LG99] Craig Larman and Rhett Guthrie. Java 2 Performance and Idiom Guide. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999.

ISBN 0-13-014260-3. xvii + 299 pp. LCCN QA76.73.J38 L359 2000. URL http:/

/www.phptr.com/ptrbooks/
ptr_0130142603.html.

Larman:2000:JPI

[LG00a] Craig Larman and Rhett Guthrie. Java 2 Performance and Idiom Guide. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2000. ISBN 0-13-014260-3. xvii + 299 pp. LCCN QA76.73.J38 L359 2000. URL http:/

/www.phptr.com/ptrbooks/
ptr_0130142603.html.

Liskov:2000:PDJ

[LG00b] B. Liskov and John Guttag. Program development in Java: abstraction, specification, and object-oriented design. Addison-Wesley, Reading, MA, USA, 2000. ISBN 0-201-65768-6. xix + 443 pp. LCCN QA76.73.J38 L58 2001.

Lujan:2005:EJA

[LGFM05] Mikel Luján, John R. Gurd, T. L. Freeman, and José Miguel. Elimination of Java array bounds checks in the presence of indirection. Concurrency and Computation: Practice and Experience, 17 (5–6):489–514, April/May 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lorenzen:2002:CCW

[LH02] T. Lorenzen and W. Heilman. CS1 and CS2: Write computer games in Java! SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(4):99– 100, 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lee:2003:RSC

[LH03a] J. S. Lee and P. L. Hsu. Remote supervisory control of the human-in-the-loop system by using Petri nets and Java. IEEE Transactions on Industrial Electronics, 50(3):431– 439, 2003. CODEN ITIED6. ISSN 0278-0046 (print), 1557-9948 (electronic).

Lhotak:2003:SJP

[LH03b] O. Lhotak and L. Hendren. Scaling Java points-to analysis using SPARK. Lecture Notes in Computer Science, 2622:153–169, 2003. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lhotak:2004:JBB

[LH04] Ondřej Lhoták and Laurie Hendren. Jedd: a BDDbased relational extension of Java. ACM SIGPLAN Notices, 39(6):158–169, May 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lhotak:2005:RTE

[LH05] Ondřej Lhoták and Laurie Hendren. Run-time evaluation of opportunities for object inlining in Java. Concurrency and Computation: Practice and Experience, 17 (5–6):515–537, April/May 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lin:2007:SEA

[LH07] Chin-Yang Lin and Ting-Wei Hou. A simple and efficient algorithm for cycle collection. ACM SIGPLAN Notices, 42(3):7–13, March 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lhotak:2008:EBC

[LH08a] Ondřej Lhoták and Laurie Hendren. Evaluating the benefits of contextsensitive points-to analysis using a BDD-based implementation. ACM Transactions on Software Engineering and Methodology, 18(1):3:1– 3:??, September 2008. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Lhotak:2008:RAB

[LH08b] Ondřej Lhoták and Laurie Hendren. Relations as an abstraction for BDD-based program analysis. ACM Transactions on Programming Languages and Systems, 30(4): 19:1–19:63, July 2008. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Lin:2007:SIM

[LHFL07] Jim-Min Lin, Zeng-Wei Hong, Guo-Ming Fang, and Che-Tai Lee. A style for integrating MS-Windows software applications to client-server systems using Java technology. Software—Practice and Experience, 37(4):417–440, April 10, 2007. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lee:2009:DAY

[LHGM09] Byeongcheol Lee, Martin Hirzel, Robert Grimm, and Kathryn S. McKinley. Debug all your code: portable mixed-environment debugging. ACM SIGPLAN Notices, 44(10):207–226, October 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Long:2003:TST

[LHS03] B. Long, D. Hoffman, and P. Strooper. Tool support for testing concurrent Java components. IEEE Transactions on Software Engineering, 29(6):555–566, June 2003. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
1205182.

Lin:2004:OJB

[LHS04a] L. Lin, L. Huang, and Y. Sun. Optimizing Java based Web services by partial evaluation. Lecture Notes in Computer Science, 3032:1071–1074, 2004. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lopez-Herrejon:2004:UIT

[LHS04b] Roberto E. Lopez-Herrejon and Morrie Schulman. Using interactive technology in a short Java course: an experience report. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):203–207, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Li:2002:AIF

[Li02] Bixin Li. Analyzing informationflow in Java program based on slicing technique. ACM SIG-SOFT Software Engineering Notes, 27(5):98–103, September 2002. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Li:2003:JBM

[Li03] Sing Li. A Java-based music player for MP3, Ogg, and WAV. Dr. Dobb’s Journal of Software Tools, 28(2):38, 40, 42, 44–45, February 2003. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/documents/s=
7790/ddj0302e/.

Li:2004:DID

[Li04] C. Li. Design and implementation of database connection — pool based on Java. Journal- Anhui University of Technology, 21(1):53– 56, 2004. CODEN ????

Liang:2000:IJPa

[Lia00a] Y. Daniel Liang. Introduction to Java Programming with JBuilder 3. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2000. ISBN 0-13-086911-2. xxvi + 771 pp. LCCN A76.73.J38 L53 2000. US$65.00. URL http:/

/www.phptr.com/ptrbooks/
esm_0130869112.html.

Liang:2000:IJPb

[Lia00b]  Y. Daniel Liang.  Intro-

duction to Java Programming with Microsoft Visual J++ 6.0. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2000. ISBN 0-13-086912-0. xxviii + 752 pp. LCCN QA76.73.J38 L532 2000. US$65.00.

Liang:2000:RJA

[Lia00c] Y. Daniel Liang. Rapid Java Application Development Using JBuilder 3. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2000. ISBN 0-13-026161-0. xxii + 727 pp.

LCCN A76.73.J38 L534 2000. US$53.

Liang:2001:IJP

[Lia01] Y. Daniel Liang. Introduction to Java Programming. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, third edition, 2001. ISBN 0-13-031997-X. 784 (est.) pp. LCCN A76.73.J38 L52 2001. US$68.00. URL http:/

/www.phptr.com/ptrbooks/
esm_013031997X.html.

Liang:2002:IJP

[Lia02]  Y. Daniel Liang.  Intro-

duction to Java programming with Jbuilder 4. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2002. ISBN 0-13-033364-6. xxxii + 939 pp. LCCN QA76.73.J38 L5313 2002.

Liang:2003:IJP

[Lia03a]  Y. Daniel Liang.  Intro-

duction to Java programming with Sun One Studio 4. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2003. ISBN 0-13-009258-4. xxxv + 962 pp. LCCN QA76.73.J38 L525 2003.

Liao:2003:THM

[Lia03b] L. Liao. Temperature and humidity monitor system based on 1-wire and Java VM technologies. Journal — Sichuan Normal University Natural Science Edition, 26(1):93–96, 2003. CODEN ???? ISSN 1001-8395.

Likos:2004:JBCa

[Lik04a] Johannis Likos. μoνo2πoλυ: Java-based conversion of monotonic to polytonic Greek. In Syropoulos et al. [SBH^{+}04], pages 34–54. CODEN LNCSD9. ISBN 3-540-22801-2 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Z253.3 I58 2004. URL

http://link.springer-ny.
com/link/service/series/
0558/tocs/t3130.htm;   http:
//www.springerlink.com/
openurl.asp?genre=issue&
issn=0302-9743&volume=3130;
 http://www.springerlink.
com/openurl.asp?genre=volume&
id=doi:10.1007/b99374.

Likos:2004:PMJb

[Lik04b] Johannis Likos. monopoly: Java-based conversion of monotonic to polytonic Greek. TUGboat, 25(0):111, August 2004. CODEN ???? ISSN 0896-3207. URL https:/

/tug.org/TUGboat/tb25-0/
likos.pdf.

Lindley:2000:DAJ

[Lin00] Craig A. Lindley. Digital audio with Java. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2000. ISBN 0-13-087676-3. xl + 380 pp. LCCN TK7881.4 .L5415 2000. Includes CD-ROM with two complete audio applications, complete library of JavaBeans user interface controls, and twelve audio processing effects.

Lingsong:2001:EDB

[Lin01]  He Lingsong.  Exchang-

ing data between Java and C/C++ Windows programs. C/C++ Users Journal, 19(1): 34–??, January 2001. CO-DEN CCUJEX. ISSN 1075-2838.

Lin:2003:DEA

[Lin03a] W. Lin. Development of electronic acquisition model for project scheduling (e-AMPS) using Java-XML. NIST Special Publication SP, 989:41– 46, 2003. CODEN ???? ISSN 1048-776X.

Link:2003:UTJ

[Lin03b] Johannes Link. Unit testing in Java: How Tests Drive the Code. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2003. ISBN 1-55860-868-0. xvii + 376 pp. LCCN QA76.76.T48 L55 2003. US$39.95. With contributions by Peter Fröhlich.

Lippman:2001:CD

[Lip01] Stanley B. Lippman. The C# delegate. C/C++ Users Journal, 19(7):42–??, July 2001. CODEN CCUJEX. ISSN 1075-2838.

Litwak:2000:PJ

[Lit00] Kenneth Litwak. Pure Java 2.

Howard W. Sams, Indianapolis, IN 46268, USA, 2000. ISBN 0-672-31654-4. xxii + 790 pp. LCCN QA76.73.J38 L59 2000.

Liu:2003:SIJ

[Liu03] C. Liu. Security issues for Java — based agents. Journal — Jishou University Natural Science, 24(1):26–36, 2003. CODEN ???? ISSN 1007-2985.

Liu:2004:DFA

[Liu04] C. H. Liu. Data flow analysis and testing of Java Server Pages. Compsac, CONF28 (SUPP):114–119, 2004. CO-DEN ???? ISSN 0730-3157.

Liu:2008:UOS

[Liu08] Peter L. Liu. Using opensource robocode as a Java programming assignment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40 (4):63–67, December 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lee:2007:WFJ

[LJ07] Gun Ho Lee and Junsu Jung. Web framework with Java and XML in multi-tiers for productivity. Future Generation Computer Systems, 23 (2):263–268, February 2007. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Lucas:2008:ITJ

[LJ08] Joan M. Lucas and Jonathan Jarvis. Incorporating transformations into JFLAP for enhanced understanding of automata. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):14–18, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Li:2000:UCS

[LJN^{+}00] Tao Li, Lizy Kurian John, Vijaykrishnan Narayanan, Anand Sivasubramaniam, Jyotsna Sabarinathan, and Anupama Murthy. Using complete system simulation to characterize SPECjvm98 benchmarks. In ACM [ACM00a], pages 22–33. ISBN 1-58113-270-0. LCCN QA76.88 .I573 2000. URL

http://www.acm.org/pubs/
contents/proceedings/supercomputing/
335231.

Lawlor:2001:SDP

[LK01] Orion Lawlor and Laxmikant Kale. Supporting dynamic parallel object arrays. In ACM [ACM01b], pages 21–

28. ISBN 1-58113-359-

6. LCCN QA76.9.O35 A26

2001. URL http://www.
philippsen.com/JGI2001/
camerareadyabstracts/7.
html;   http://www.philippsen.
com/JGI2001/finalpapers/
18500021.pdf.

Lee:2003:TIW

[LKL^{+}03] S. M. Lee, O. S. Kwon, J. H.

Lee, C. J. Oh, and S. H.

Ko. TY*SecureWS: An integrated Web service security solution based on Java. Lecture Notes in Computer Science, 2738:186–195, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liu:2006:II

[LKM06] Jed Liu, Aaron Kimball, and Andrew C. Myers. Interruptible iterators. ACM SIG-PLAN Notices, 41(1):283– 294, January 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lewis:2000:JSS

[LL00] John Lewis and William Loftus. Java software solutions: foundations of program design. Addison-Wesley, Reading, MA, USA, second edition, 2000. ISBN 0-201-61271-2. xxviii + 780 pp. LCCN QA76.73.J38 L49 2000.

Lee:2001:IEW

[LL01a] Raymond S. T. Lee and James N. K. Liu. iJADE eMiner — A Web-based mining agent based on intelligent Java agent development environment (iJADE) on Internet shopping. Lecture Notes in Computer Science, 2035:28– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.
                                        com/link/service/series/

0558/bibs/2035/20350028.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2035/20350028.
pdf.

Lewis:2001:JSS

[LL01b] John Lewis and William Loftus. Java Software Solutions: Foundations of Program Design. Addison-Wesley, Reading, MA, USA, second edition, 2001. ISBN 0-201-72597-5. xxviii + 784 pp. LCCN QA76.73.J38 L49 2001.

LewisJohn:2001:JSS

[LL01c] Ph. D. Lewis, John and William Loftus. Java software solutions: foundations of program design. Addison-Wesley Longman, Reading, MA, USA, second edition, 2001. ISBN 0-201-72597-5 (paperback). xxviii + 784 pp. LCCN QA76.73.J38 L49 2001. Accompanying CD-ROM contains Java development tools, source code, and PowerPoint slides from the text.

Luthi:2001:IPC

[LL01d] Johannes Lüthi and Catalina M.

Lladó. Interval parameters for capturing uncertainties in an EJB performance model. ACM SIGMETRICS Perform. Eval. Rev., 29(1):291– 300, June 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Lewis:2003:JSS

[LL03] John Lewis and William Loftus. Java software solutions: foundations of program design. Addison-Wesley, Reading, MA, USA, third edition, 2003. ISBN 0-201-78129-8. xxix + 923 pp. LCCN QA76.73.J38 L49 2003.

Lenzerini:2008:PTS

[LL08a] Maurizio Lenzerini and Domenico Lembo, editors. Proceedings of the Twenty-Seventh ACM SIGMOD-SIGACT-SIGART Symposium on Principles of Database Systems: PODS’08, Vancouver, BC, Canada, June 9–11, 2008. ACM Press, New York, NY 10036, USA, 2008. ISBN 1-60560-932-3. LCCN ????

Liguori:2008:JPG

[LL08b] Robert Liguori and Patricia Liguori. Java pocket guide. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2008. ISBN 0-596-51419-0. xiii + 177 pp. LCCN QA76.73.J38 L545 2008.

Lim:2008:RSS

[LLCF08] Sang Boem Lim, Hanku Lee, Bryan Carpenter, and Geoffrey Fox. Runtime support for scalable programming in Java. The Journal of Supercomputing, 43 (2):165–182, February 2008. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0920-8542&volume=43&
issue=2&spage=165.

Lobosco:2008:ERT

[LLdA08] Marcelo Lobosco, Orlando Loques, and Claudio L. de Amorim. On the effectiveness of runtime techniques to reduce memory sharing overheads in distributed Java implementations. Concurrency and Computation: Practice and Experience, 20(13):1509– 1538, September 10, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Lu:2003:PVP

[LLK03] H. Lu, Y. Li, and F. Kong.

The principle of VRML parser in Java language. Journal — Zhejiang University of Technology, 31(2):139–143, 2003. CODEN ???? ISSN 1006-4303.

Lau:2003:MMT

[LLMK03] T. C. Lau, J. Lu, J. Mylopoulos, and K. Kontogiannis. The migration of multitier E-commerce applications to an enterprise Java environment. Information Systems Frontiers, 5(2):149–160, 2003. CODEN ???? ISSN 1387-3326.

Liu:2008:PBH

[LLS^{+}08] Tiancheng Liu, Ying Li, Andrew Schofield, Matt Hogstrom, Kewei Sun, and Ying Chen. Partition-based heap memory management in an application server. Operating Systems Review, 42 (1):98, January 2008. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Liu:2002:JIA

[LM02] J. Liu and A. C. Myers. JMatch: Iterable abstract pattern matching for Java. Lecture Notes in Computer Science, 2562:110–127, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liu:2004:JPV

[LM04] H. Liu and J. S. Moore. Java program verification via a JVM deep embedding in ACL2. Lecture Notes in Computer Science, 3223:184–200, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lewis:2006:GGD

[LM06] Mark C. Lewis and Berna Massingill. Graphical game development in CS2: a flexible infrastructure for a semester long project. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38 (1):505–509, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lewis:2000:APH

[LMG00] Brian T. Lewis, Bernd Mathiske, and Neal M. Gafter. Architecture of the PEVM: a high-performance orthogonally persistent Java Virtual Machine. Technical Report 2000-93, Sun Microsystems, Palo Alto, CA, USA, October 2000. 19 pp.

Lewis:2001:APH

[LMG01] Brian Lewis, Bernd Mathiske, and Neal Gafter. Architecture of the PEVM: a highperformance orthogonally Persistent Java^{T M} Virtual Machine. Lecture Notes in Computer Science, 2135:18– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2135/21350018.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2135/21350018.
pdf.

Li:2006:PBH

[LMK06] W. Li, S. Mohanty, and K. Kavi. A page-based hybrid (software–hardware) dynamic memory allocator. IEEE Computer Architecture Letters, 5(2):13, February 2006. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Lee:2008:EHS

[LMK08] Seong-Won Lee, Soo-Mook Moon, and Seong-Moo Kim. Enhanced hot spot detection heuristics for embedded Java just-in-time compilers. ACM SIGPLAN Notices, 43(7):13– 22, July 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

LEcuyer:2002:SFS

[LMV02] P. L’Ecuyer, L. Meliani, and J. Vaucher. SSJ: a framework for stochastic simulation in Java. Winter Simulation Conference, 1(??):234–242, 2002. CODEN WSCPDK. ISSN 0743-1902 (print), 2160-9276 (electronic).

Lefranc:2002:CPA

[LN02] S. Lefranc and D. Naccache. Cut-&-Paste attacks with JAVA. Lecture Notes in Computer Science, 2587:1– 15, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:2004:JBN

[LN04] Y. Lee and S. Na. Java bytecode-to-.NET MSIL translator for construction of platform independent information systems. Lecture Notes in Computer Science, 3215:826– 832, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lambert:2000:JFP

[LO00a] Kenneth Alfred Lambert and Martin Osborne. Java: a framework for program design and data structures. Brooks/Cole, Pacific Grove, CA, USA, 2000. ISBN 0-534-36691-0. xxii + 677 pp. LCCN QA76.73.J38 L25 2000. Includes CD-ROM.

Lambert:2000:JCC

[LO00b] Kenneth (Kenneth A.) Lambert and Martin Osborne. Java: complete course in programming and problem solving. South-Western Educational Publishers, Cincinnati, OH, USA, 2000. ISBN 0-538-68707-X (hardbound), 0-538-68711-8 (softcover). x + 468 pp. LCCN QA76.73.J38 L355 2000.

Lambert:2003:FJC

[LO03a] Kenneth Alfred Lambert and Martin Osborne. Fundamentals of Java: comprehensive course. Thomson/Course Technology, Boston, MA, USA, second edition, 2003. ISBN 0-619-05963-X. various pp. LCCN QA76.73.J38 L354 2003.

Lambert:2003:JB

[LO03b] Kenneth Alfred Lambert and Martin Osborne. Java basics. Thomson/Course Technology, Boston, MA, USA, 2003. ISBN 0-619-05967-2. xv + 260 pp. LCCN QA76.73.J38 L354 2002.

Loton:2002:WCM

[Lot02] Tony Loton. Web content mining with Java: techniques for exploiting the World Wide Web. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 2002. ISBN 0-470-84311-X. xviii + 305 pp. LCCN QA76.9.D343 L68 2002.

Louridas:2005:JUT

[Lou05] Panagiotis Louridas. JUnit: Unit testing and coding in tandem. IEEE Software, 22 (4):12–15, July/August 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Leather:2009:RPE

[LOW09] Hugh Leather, Michael O’Boyle, and Bruce Worton. Raced profiles: efficient selection of competing compiler optimizations. ACM SIGPLAN Notices, 44(7):50–59, July 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Launay:2001:EPP

[LP01a] Pascale Launay and Jean-Louis Pazat. Easing parallel programming for clusters with Java. Future Generation Computer Systems, 18(2): 253–263, October 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/

19/19/60/31/32/abstract.
html.

Levanoni:2001:FRC

[LP01b] Yossi Levanoni and Erez Petrank. An on-the-fly reference counting garbage collector for Java. ACM SIG-PLAN Notices, 36(11):367– 380, November 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Proceedings of the 2001 ACM SIGPLAN Conference on Object Oriented Programming, Systems, Languages and Applications (OOPSLA’01).

Landau:2005:FCS

[LP05] Rubin H. Landau and M. J.

Páez. A first course in scientific computing: symbolic, graphical, and numeric problem solving using Maple, Java, Mathematica, and Fortran. Princeton University Press, Princeton, NJ, USA, 2005. ISBN 0-691-12183-4 (hardcover). xxiv + 481 pp. LCCN Q183.9 .L36 2005.

Levanoni:2006:FRC

[LP06] Yossi Levanoni and Erez Petrank. An on-the-fly reference-counting garbage collector for Java. ACM Transactions on Programming Languages and Systems, 28(1):1–69, January 2006. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Liang:2001:EEF

[LPH01] Donglin Liang, Maikel Pennings, and Mary Jean Harrold. Extending and evaluating flow-insensitive and context-insensitive points-to analyses for Java. In ACM [ACM01a], pages 73–79. ISBN 1-58113-413-4. LCCN QA76.758 .A345 2001. URL

http://www.acm.org/pubs/
articles/proceedings/soft/
379605/p73-liang/p73-liang.
pdf;   http://www.acm.org/
pubs/citations/proceedings/
soft/379605/p73-liang/.

Supplement to ACM SIG-PLAN Notices.

Liang:2002:EPS

[LPH02] Donglin Liang, Maikel Pennings, and Mary Jean Harrold. Evaluating the precision of static reference analysis using profiling. ACM SIGSOFT Software Engineering Notes, 27(4):22–32, July 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Liang:2006:EIC

[LPH06] Donglin Liang, Maikel Pennings, and Mary Jean Harrold. Evaluating the impact of context-sensitivity on Andersen’s algorithm for Java programs. ACM SIGSOFT Software Engineering Notes, 31 (1):6–12, January 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Liu:2004:AJI

[LPSY04] H. Liu, Q. Peng, J. Shen, and C. Yan. Algorithms and Java implementation of NASO publish-propagate replication scheme. Mini-Micro Systems, 25(7):1307–1312, 2004. CO-DEN XWJXEH. ISSN 1000-1220.

Leff:2004:AES

[LR04] Avraham Leff and James T.

Rayfield. Alternative edgeserver architectures for Enterprise JavaBeans applications. In Jacobsen [Jac04b], pages 195–?? CODEN LNCSD9. ISBN 3-540-23428-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.M54 I33 2004. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.springerlink.
com/openurl.asp?genre=issue&
issn=0302-9743&volume=3231;
http://www.springerlink.
com/openurl.asp?genre=volume&
id=doi:10.1007/b101561.

Leff:2005:EJC

[LR05] Avraham Leff and James T.

Rayfield. Enterprise JavaBeans caching in clustered environments. Concurrency and Computation: Practice and Experience, 17(7–8): 1027–1051, June/July 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Luxton-Reilly:2009:SFI

[LRD09] Andrew Luxton-Reilly and Paul Denny. A simple framework for interactive games in CS1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):216–220, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Long:2002:BSM

[LRO02] Carol A. Long, V. V. S. Raveendra, and Tope Omitola. Bookshelf: Software manufacturing manager’s handbook; inside Java 2 platform security; so you want to build an operating system. IEEE Software, 19(4):131–133, July/August 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:

//dlib.computer.org/so/
books/so2002/pdf/s4131.

pdf. Review of Software Product Lines: Practices and Patterns by Paul Clements and Linda Northrop, Inside Java 2 Platform Security: Architecture, API Design and Implementation by Li Gong, and Applied Operating Systems Concepts by Avi Silberschatz, Peter B. Galvin, Abraham Silberschatz, and Greg Gagne.

Li:2000:WGW

[LRSW00] M. Li, O. F. Rana, M. S.

Shields, and D. W. Walker. A wrapper generator for wrapping high performance legacy codes as Java/CORBA components. In ACM [ACM00c],

page 51. URL http://www.
sc2000.org/proceedings/
techpapr/papers/pap135.
pdf.

Li:2001:WMB

[LRW01] Maozhen Li, Omer F. Rana, and David W. Walker. Wrapping MPI-based legacy codes as Java/CORBA components. Future Generation Computer Systems, 18(2): 213–223, October 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/60/31/29/abstract.
html.

Lee:2000:JAT

[LS00] Rosanna Lee and Scott Seligman. JNDI API tutorial and reference: building directoryenabled Java applications. Java series. Addison-Wesley, Reading, MA, USA, 2000. ISBN 0-201-70502-8. xxiii + 821 pp. LCCN QA76.76.A63 L44 2000.

Lim:2003:SOI

[LS03] S. Lim and Y. Song. Secure Online Internet Reservation of e-Government Service (ORGS) using Java Card Applications Toolkit (J-CAT). Lecture Notes in Computer Science, 2739:272–277, 2003.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:2004:OPD

[LS04a] H. J. (Hyun Jin) Lee and W. E. Schiesser. Ordinary and partial differential equation routines in C, C++, Fortran, Java^{(}R), Maple^{(}R), and MATLAB^{(}R). Chapman and Hall/CRC, Boca Raton, FL, USA, 2004. ISBN 1-58488-423-1. xiii + 519 pp. LCCN QA371.5.D37 L44 2004.

LopezHerrejon:2004:UIT

[LS04b] R. E. Lopez Herrejon and M. Schulman. Using interactive technology in a short Java course: An experience report. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):203–207, 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Liu:2006:FFCa

[LS06] Yu David Liu and Scott F. Smith. A formal framework for component deployment. ACM SIGPLAN Notices, 41(10):325–344, October 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Liquori:2008:EFJ

[LS08a] Luigi Liquori and Arnaud Spiwack. Extending FeatherTrait Java with Interfaces. Theoretical Computer Science, 398(1–3):243–260, May 28, 2008. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Liquori:2008:FME

[LS08b] Luigi Liquori and Arnaud Spiwack. FeatherTrait: a modest extension of Featherweight Java. ACM Transactions on Programming Languages and Systems, 30(2): 11:1–11:32, March 2008. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Lorenzen:2008:OFU

[LS08c] Torben Lorenzen and Abdul Sattar. Objects first using Alice to introduce object constructs in CS1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(2):62–64, June 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp:

//ftp.math.utah.edu/pub/
mirrors/ftp.ira.uka.de/
bibliography/Misc/DBLP/
2008.bib.

Lind:2002:RPH

[LSK^{+}02] L. Lind, E. Sundvall, D. Karlsson, N. Shahsavar, and H. Ahlfeldt. Requirements and prototyping of a home health care application based on emerging Java technology. International Journal of Medical Informatics, 68(1–3): 129–139, 2002. CODEN ???? ISSN 1386-5056.

League:2002:TPC

[LST02] Christopher League, Zhong Shao, and Valery Trifonov. Type-preserving compilation of Featherweight Java. ACM Transactions on Programming Languages and Systems, 24(2):112–152, March 2002. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

League:2003:PPT

[LST03] C. League, Z. Shao, and V. Trifonov. Precision in practice: a type-preserving Java compiler. Lecture Notes in Computer Science, 2622:106–120, 2003. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Long:2007:MVC

[LSW07] Brad Long, Paul Strooper, and Luke Wildman. A method for verifying concurrent Java components based on an analysis of concurrency failures. Concurrency and Computation: Practice and Experience, 19(3):281– 294, March 10, 2007. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Langmaack:2008:DAI

[LSW08] Hans Langmaack, Andrzej Salwicki, and Marek Warpechowski. A deterministic algorithm for identifying direct superclasses in Java. Fundamenta Informaticae, 85(1– 4):343–357, September 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lee:2002:POO

[LT02] Richard C. Lee and William M.

Tepfenhart. Practical objectoriented development with UML and Java. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2002. ISBN 0-13-067238-6. xxv + 468 pp. LCCN QA76.64 .L448 2002.

Laskowski:2007:BCS

[LTOT07] Eryk Laskowski, Marek Tudruj, Richard Olejnik, and Bernard Toursel. Byte-code scheduling of Java programs with branches for Desktop Grid. Future Generation Computer Systems, 23(8): 977–982, November 2007. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Lujan:2005:SFS

[LUH^{+}05] M. Lujan, A. Usman, P. Hardie, T. L. Freeman, and J. R. Gurd. Storage formats for sparse matrices in Java. Lecture Notes in Computer Science, 3514:364–371, 2005. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Luke:2004:DMT

[Luk04] Sean Luke. Documentation for the Mersenne Twister in Java. Web site., October

2004. URL http://www.cs.
gmu.edu/~sean/research/
mersenne.

Lutz:2000:NBM

[Lut00] Michael J. Lutz. New books: Masters of technology or slaves to the E-grind?; managing apples and oranges; descrambling data on the web; designing system networks; fixing things before things fall apart; Java 2 guidelines. Computer, 33(2):119, February 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//dlib.computer.org/co/
books/co2000/pdf/r2119.

pdf. Includes short review of [LG99].

Lutz:2001:NBIb

[Lut01] Michael J. Lutz. New books: Investigating evidential computing; steering the computing juggernaut; in the mood for Java, Smalltalk, and algorithms?; meeting the computational intelligence challenge; assessing the digital economy. Computer, 34(2):112, February 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//dlib.computer.org/co/
books/co2001/pdf/r2112.
pdf.

Lutz:2002:BAN

[Lut02] Michael J. Lutz. Bookshelf:

Architecture by the numbers [Computer Architecture: A Quantitative Approach]; half a Java tone [Algorithms in Java]; embedded systems simulation [Embedded Systems and Computer Architecture]; brand-name multicasting [Interdomain Multicast Routing: Practical Juniper Networks and Cisco Systems Solutions]. Computer, 35(11): 77, November 2002. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/co/
2002/11/ry077.htm;   http:
//csdl.computer.org/dl/
mags/co/2002/11/ry077.pdf.

Lutz:2003:BBC

[Lut03a] Michael J. Lutz. Bookshelf:

Building C# programs with patterns [C# Design Patterns]; using Star Schema in Oracle [Oracle DBA Guide to Data Warehousing and Star Schemas]; immersive Java text [Java Outside In]; applying AI to the Web [Web Intelligence]; moronic marketing [In Search of Stupidity: Over 20 Years of High-Tech Marketing Disasters]. Computer, 36(9):96– 97, September 2003. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/co/

                                        2003/09/r9096.htm;   http:
                                        //csdl.computer.org/dl/
                                        mags/co/2003/09/r9096.pdf.

Lutz:2003:BSW

[Lut03b] Michael J. Lutz. Bookshelf: Surveying the Web’s best and brightest [Deep Sites: Intelligent Innovation in Contemporary Web Design]; moving voice communications to the Web [Implementing Voice over IP]; developing custom .NET management tools [.NET Systems Management Services]; evolving Java [Applied Evolutionary Algorithms in Java]. Computer, 36(7):86, July 2003. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/dl/mags/co/
                                        2003/07/r7086.htm;   http:
                                        //csdl.computer.org/dl/
                                        mags/co/2003/07/r7086.pdf.

Lutz:2003:BFE

[Lut03c]  Michael J. Lutz.  Book-

shelf: The five essential metrics [Five Core Metrics: The Intelligence Behind Successful Software Management]; hunting bugs in concurrent systems [The SPIN Model Checker: Primer and Reference Model]; revised digital testing tome [Digital Logic Testing and Simulation]; Java for science and engineering [Technical Java: Developing Scientific and Engineering Applications]. Computer, 36 (11):85, November 2003. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/co/
2003/11/ry085.htm;   http:
//csdl.computer.org/dl/
mags/co/2003/11/ry085.pdf.

Liu:2003:RII

[LW03] F. Liu and H. Wang. Realization of an interactive and individual long-distance teaching system based on Java technology. Journal — Beijing University of Chemical Technology Natural Science Edition, 30(2):85–88, 2003. CODEN ????

Liu:2003:IRL

[LY03] Y. Liu and K. Yang. The implementation of remote loading classes on the Java platform. Journal — Hebei Institute of Technology, 25(4): 100–105, 2003. CODEN ???? ISSN 1007-2829.

Lee:2002:AOI

[LYC02] Ji-Hyun Lee, Cheol-Jung Yoo, and Ok-Bae Chang. Analysis of object interaction during the enterprise JavaBeans lifecycle using formal specification technique. ACM SIGPLAN Notices, 37(6):82– 92, June 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lee:2000:RVC

[LYK^{+}00] Junpyo Lee, Byung-Sun Yang, Suhyun Kim, Kemal Ebcioğlu, Erik Altman, Seungil Lee, Yoo C. Chung, Heungbok Lee, Je Hyung Lee, and Soo-Mook Moon. Reducing virtual call overheads in a Java VM just-in-time compiler. ACM SIGARCH Computer Architecture News, 28 (1):21–33, March 2000. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Lykins:2002:SYB

[Lyk02] D. Lykins. Should you bet your business on Java or .NET? E Business Advisor, 20(1):18–24, 2002. CODEN ???? ISSN 1098-8912.

Liu:2004:JBD

[LYL^{+}04] Z. Liu, H. Yu, E. P. Lim, M. Yin, D. H. Goh, Y. L. Theng, and W. K. Ng. A Java-based digital library portal for geography education. Science of Computer Programming, 53(1):87–105, 2004. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Lee:2004:EJE

[LYM04] SeungIl Lee, Byung-Sun Yang, and Soo-Mook Moon. Efficient Java exception handling in just-in-time compilation. Software—Practice and Experience, 34(15):1463– 1480, December 2004. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lyon:2002:SMI

[Lyo02]  Douglas Lyon.  Simulat-

ing multiple inheritance in Java. Concurrency and Computation: Practice and Experience, 14(12):987–1008, October 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (elec-

tronic). URL http://www3.
interscience.wiley.com/
cgi-bin/abstract/98516164/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=98516164{\&}PLACEBO=
IE.pdf.

Li:2004:ACF

[LZ04] Peng Li and Steve Zdancewic.

Advanced control flow in Java card programming. ACM SIGPLAN Notices, 39(7): 165–174, July 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Liu:2003:RDE

[LZZ03] H. Liu, X. Zhou, and B. Zhong. The realization of data exchange between different system platforms using Java and XML under the environment of Lotus Domino. Journal — Sichuan Normal University Natural Science Edition, 26(3):327–330, 2003. CODEN ???? ISSN 1001-

Malks:2000:PJ

[M^{+}00] Dan Malks et al. Professional JSP. Wrox Press, Chicago, IL, USA, 2000. ISBN 1-86100-362-5. xxv + 897 pp. LCCN TK5105.8885.J38 P76 2000. US$59.99. URL http:

                                        //www.wrox.com/Consumer/
                                        Store/Download.asp?ISBN=
                                        1861003625.

Marinacci:2005:SHT

[MA05] Joshua Marinacci and Chris Adamson. Swing Hacks: Tips and Tools for Killer GUIs. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00907-0. xix + 519 pp. LCCN QA76.73.J38 M37 2005. US$9.95, CAN$41.95, UK£20.95.

Macvittie:2005:PAI

[Mac05] D. Macvittie. Product analysis: Imported Java. Network Computing, 16(1):45– 49, 2005. CODEN NCOMEV. ISSN 1046-4468.

Madrigal:2001:FOD

[Mad01] Victor M. Madrigal III. Fundamentals of OOP and data structures in Java. ACM SIGSOFT Software Engineering Notes, 26(1):99, January 2001. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Mahmoud:2002:LWJ

[Mah02] Qusay H. Mahmoud. Learning Wireless Java. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2002. ISBN 0-596-00243-2. xiii + 245 pp. LCCN QA76.73.J38 M323 2002; TK5103.2 .M33 2002. US$34.95. URL http://

safari.oreilly.com/0596002432;
http://www.oreilly.com/
catalog/wirelessjava. Help

for new J2ME developers. Covers Motorola and Palm hardware.

Mahmoud:2004:PEJ

[Mah04a] Q. H. Mahmoud. Practice and experience with Java in education. Science of Computer Programming, 53(1):1– 2, 2004. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Mahmoud:2004:WJA

[Mah04b] Qusay H. Mahmoud. Wireless Java applications development. ;login: the USENIX Association newsletter, 29(2): ??, April 2004. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2004-04/pdfs/Mahmoud.pdf.

Mahemoff:2006:ADP

[Mah06] Michael Mahemoff. Ajax Design Patterns. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2006. ISBN 0-596-10180-5. 352 (est.) pp. LCCN ???? EUR 38.00.

Main:2003:DSO

[Mai03] M. (Michael) Main. Data structures and other objects using Java. Addison-Wesley, Reading, MA, USA, second edition, 2003. ISBN 0-201-74093-1. 808 pp. LCCN QA76.73.J38 M33 2003.

Miller:2003:LTB

[MAJC03] G. R. Miller, P. Arduino, J. Jang, and C. Choi. Localized tensor-based solvers for interactive finite element applications using C++ and Java. Computers and Structures, 81(7):423–437, 2003. CODEN CMSTCJ. ISSN 0045-7949 (print), 1879-2243 (electronic).

Mak:2003:JNC

[Mak03] Ronald Mak. Java Number Cruncher: The Java Programmer’s Guide to Numerical Computing. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2003. ISBN 0-13-046041-9. xi + 464 pp. LCCN QA76.73.J38 M335 2003.

Mamlin:2001:OSX

[Mam01] B. Mamlin. An open-source XML-based Java implementation of the medical gopher order and note writing tool. Journal of Biomedical Informatics, 35(SUPP):826– ??, 2001. CODEN JBIOBL. ISSN 1532-0464.

Manduchi:2001:DJA

[Man01] Gabriele Manduchi. Developing Java applications for a nuclear fusion experiment: a test case for Java applicability in a demanding environment. Software—Practice and Experience, 31(11):1025– 1042, September 2001. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (elec-

tronic). URL http://www3.
interscience.wiley.com/
cgi-bin/abstract/83001949/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=83001949&PLACEBO=IE.
pdf.

Mann:2005:JFA

[Man05] Kito D. Mann. JavaServer Faces in action. Manning Publications, Greenwich, CT, USA, 2005. ISBN 1-932394-12-5. xxxiv + 702 pp. LCCN TK5105.8885.J38 M35

2005. URL http://www.loc.
gov/catdir/enhancements/
fy0626/2005273936-b.html;
 http://www.loc.gov/catdir/
enhancements/fy0626/2005273936-
d.html.

Margulies:2000:UJT

[Mar00] Edwin Margulies. Understanding Java Telephony. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2000. ISBN 1-199-00626-2. ???? pp. LCCN ????

Marco:2001:EJJ

[Mar01a] Lou Marco. EJB and JSP: Java on the edge, unlimited edition. Professional mindware. M&T Books, M&T Publishing, Inc., 501 Galveston Drive, Redwood City, CA 94063, USA, 2001. ISBN 0-7645-4802-6. xxiv + 452 pp. LCCN QA76.73.J38 M3482 2001.

Marti:2001:ZZH

[Mar01b]  Don Marti.  ZapMedia

ZapStation/Harman Kardon DMC 100. Embedded Linux Journal, 2:12–13, March/ April 2001. CODEN ???? ISSN 1534-083X. URL http:

                                        //embedded.linuxjournal.
                                        com/magazine/issue02/4526;
                                         http://www.linuxdevices.
                                        com/articles/AT8585138080.
                                        html.

Marques:2002:BSJ

[Mar02] Paulo Marques. Building secure Java RMI servers. Dr. Dobb’s Journal of Software Tools, 27(11):36, 38, 40–42, 44, November 2002. CODEN DDJOEB. ISSN 1044-789X. URL http://

                                        www.ddj.com/documents/s=
                                        7644/ddj0211d/.

Mares:2005:BRA

[Mar05] Peter Mares. Book review: Art of Java Web Development, by N. Ford. The Computer Journal, 48(2): 253, March 2005. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

www3.oup.co.uk/computer_
journal/free_pdf/bxh071.
pdf;   http://www3.oup.
co.uk/computer_journal/
hdb/Volume_48/Issue_02/
bxh071.sgm.abs.html.

Mason:2000:PCL

[Mas00] Oliver Mason. Programming for corpus linguistics: how to do text analysis with Java. Edinburgh University Press, Edinburgh, Scotland, 2000. ISBN 0-7486-1407-9. 245 pp. LCCN P98 .M29 2000. Specialised linguistic research needs can no longer be met by available software. This book enables the researcher to write programs for text and corpus processing, using the popular and easy to learn Java language.

Masum:2001:BRBa

[Mas01] Hassan Masum. Book review: Data Structures and Algorithms in Java (2nd ed): Michael T Goodrich and Roberto Tamassia. ACM SIGACT News, 32(1):3– 5, March 2001. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [GT97, GT01, GT04, GT06, GT10].

Maurer:2002:CPL

[Mau02] W. Douglas Maurer. The comparative programming languages course: a new chain of development. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1): 336–340, March 2002. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Maly:2001:IHJ

[MAWW^{+}01] R. Maly, H. Abdel-Wahab, C. Wild, C. M. Overstreet, A. Gupta, A. Abdel-Hamid, S. Ghanem, A. Gonzalez, and X. Zhu. IRI-h, a Javabased distance education system: architecture and performance. ACM Journal on Educational Resources in Computing (JERIC), 1(1es): ??, March 2001. CODEN ???? ISSN 1531-4278.

Mahovsky:2003:AJB

[MB03] J. Mahovsky and L. Benedicenti. An architecture for Java-based real-time distributed visualization. IEEE Transactions on Visualization and Computer Graphics, 9(4): 570–579, 2003. CODEN ???? ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Moritz:2005:DFC

[MB05] Sally H. Moritz and Glenn D. Blank. A design-first curriculum for teaching Java in a CS1 course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(2):89–93, June 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp:

//ftp.math.utah.edu/pub/
mirrors/ftp.ira.uka.de/
bibliography/Misc/DBLP/
2005.bib.

Maebe:2006:JSBa

[MBED06] Jonas Maebe, Dries Buytaert, Lieven Eeckhout, and Koen De Bosschere. Javana: a system for building customized Java program analysis tools. ACM SIGPLAN Notices, 41(10):153–168, October 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Marquez:2001:IOP

[MBMZ01] Alonso Marquez, Stephen M.

Blackburn, Gavin Mercer, and John Zigman. Implementing orthogonally persistent Java. Lecture Notes in Computer Science, 2135:247– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2135/21350247.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2135/21350247.
pdf.

Menon:2008:SGL

[MBS^{+}08] Vijay Menon, Steven Balensiefer, Tatiana Shpeisman, Ali-Reza Adl-Tabatabai, Richard L. Hudson, Bratin Saha, and Adam Welc. Single global lock semantics in a weakly atomic STM. ACM SIG-PLAN Notices, 43(5):15–26, May 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mountjoy:2004:WDG

[MC04] Jon Mountjoy and Avinash Chugh. WebLogic: The Definitive Guide. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2004. ISBN 0-596-00432-X. xvii + 828 pp. LCCN TK5105.8885.B4 M68 2004. URL http://

                                        www.oreilly.com/catalog/
                                        9780596004323.

Moon:2006:TMS

[MC06] Sewon Moon and Byeong-Mo Chang. A thread monitoring system for multithreaded Java programs. ACM SIG-PLAN Notices, 41(5):21–29, May 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McCluskey:2000:JPa

[McC00a] Glen McCluskey. Java performance. ;login: the USENIX Association newsletter, 25(1): ??, February 2000. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2000-2/features/javaperformance.
                                        html.

McCluskey:2000:JPb

[McC00b] Glen McCluskey. Java performance. ;login: the USENIX Association newsletter, 25(2): ??, April 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-4/features/java.html.

McCluskey:2000:JPc

[McC00c] Glen McCluskey. Java performance. ;login: the USENIX Association newsletter, 25(3): ??, June 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-6/features/java.html.

McCluskey:2000:JPd

[McC00d] Glen McCluskey. Java performance. ;login: the USENIX Association newsletter, 25(5): ??, August 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-8/features/java.html.

McCluskey:2000:JPe

[McC00e] Glen McCluskey. Java performance. ;login: the USENIX Association newsletter, 25(6): ??, October 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-10/pdfs/javaperformance.
pdf.

McCluskey:2000:JPf

[McC00f] Glen McCluskey. Java performance. ;login: the USENIX Association newsletter, 25(8): ??, December 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-12/pdfs/javaperf.pdf.

McCoy:2000:SP

[McC00g] John H. McCoy. Scripting for Pnuts. Dr. Dobb’s Journal of Software Tools, 25 (1):21–22, 24, 26, January 2000. CODEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/ftp/2000/
2000_01/pnutdemo.zip;
http://www.ddj.com/ftp/
2000/2000_01/pnuts.txt.

McCluskey:2001:JPa

[McC01a] Glen McCluskey. Java performance. ;login: the USENIX Association newsletter, 26(1): ??, February 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2001-02/pdfs/mccluskey.
pdf.

McCluskey:2001:JPb

[McC01b] Glen McCluskey. Java performance. ;login: the USENIX Association newsletter, 26(2): ??, April 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2001-04/pdfs/mccluskey1.
pdf.

Mytkowicz:2009:ICP

[MCD09] Todd Mytkowicz, Devin Coughlin, and Amer Diwan. Inferred call path profiling. ACM SIGPLAN Notices, 44(10):175–190, October 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McFarland:2008:JMM

[McF08] David Sawyer McFarland.

JavaScript: the missing manual. Missing manual. Pogue Press/O’Reilly, Sebastopol, CA, USA; Beijing, China, 2008. ISBN 0-596-51589-8. xiv + 528 pp. LCCN QA76.73.J39 M388 2008; **See. URL http://

proquest.safaribooksonline.
com/9780596515898.

Matthews:2003:MJD

[MCG03a] Mark Matthews, Jim Cole, and Joe Gradecki. MySQL and Java developer’s guide. Java open source library. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 2003. ISBN 0-471-26923-9. xxii + 410 pp. LCCN QA76.3.S67 M38 2003. URL ftp:/

/uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
bios/wiley045/2002155887.
html;   http://www.loc.
gov/catdir/description/
wiley038/2002155887.html;
 http://www.loc.gov/catdir/

                                        toc/wiley031/2002155887.
                                        html.

McGowan:2003:JCA

[McG03b] D. McGowan. Has Java changed anything? the sound and fury of innovation litigation. Minnesota Law Review, 87(6):2039–2062, 2003. CO-DEN ???? ISSN 0026-5535.

McGinnis:2004:DLS

[McG04] L. F. McGinnis. Distributing a large-scale, complex fab simulation using HLA and Java: Issues and lessons. Winter Simulation Conference, 2:1891–1895, 2004. CO-DEN ???? ISSN 0743-1902 (print), 2160-9276 (electronic).

Myles:2005:ETS

[MCHN05] Ginger Myles, Christian Collberg, Zachary Heidepriem, and Armand Navabi. The evaluation of two software watermarking algorithms. Software—Practice and Experience, 35(10):923–938, August 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

McKenzie:2001:JQJ

[McK01]  Nadine McKenzie.  Java

Q&A: Is JavaScript an objectoriented language? Dr. Dobb’s Journal of Software Tools, 26(8):115–116, August 2001. CODEN DDJOEB. ISSN 1044-789X. URL http:

                                        //www.ddj.com/ftp/2001/
                                        2001_08/jqa0801.txt.

McLaughlin:2000:JX

[McL00] Brett McLaughlin. Java and XML. Java series. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2000. ISBN 0-596-00016-2. xvi + 479 pp. LCCN QA76.73.J38 M39 2000.

McLaughlin:2001:JX

[McL01a] Brett McLaughlin. Java and XML. Java series. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2001. ISBN 0-596-00197-5 (paperback). xv + 509 pp. LCCN QA76.73.J38 M39 2001. Also available via the Internet.

McLaughlin:2001:JXE

[McL01b]  Brett McLaughlin.  Java

and XML: Solutions to Real-World Problems. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, September 2001. ISBN 0-596-00197-5. xv + 509 pp. LCCN QA76.73.J38 .M39

2001.  US$39.95.  URL

http://www.oreilly.com/
catalog/9780596001971;
http://www.oreilly.com/
catalog/javaxml2.

McLaughlin:2002:BJE

[McL02a] Brett McLaughlin. Building Java enterprise applications: vol. 1: architecture. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2002. ISBN 0-596-00123-1. xiii + 302 pp. LCCN QA76.73.J38 M38 2002. Designing with EJBs, databases, and directory servers.

McLaughlin:2002:JXD

[McL02b] Brett McLaughlin. Java and XML data binding. Nutshell handbook. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2002. ISBN 0-596-00278-5. xi + 200 pp. LCCN QA76.73.J38 M39 2002 Stacks.

McLaughlin:2006:HRA

[McL06a] Brett McLaughlin. Head rush Ajax. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2006. ISBN 0-596-10225-9. xxxi + 413 pp. LCCN QA76.73.J39 M35 2006eb; QA76.73.J39. URL http://

www.oreilly.com/catalog/
9780596102258.

McLaughlin:2006:JX

[McL06b] Brett McLaughlin. Java and XML. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, third edition, 2006. ISBN 0-596-10149-X. xii + 465 pp. LCCN QA76.73.J38 M39 2006eb; QA76.73.J38; QA76.73.J38 M39 2006. URL

http://www.oreilly.com/
catalog/9780596101497.

McLaughlin:2007:JX

[McL07]  Brett McLaughlin.  Java

and XML. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, third edition, 2007. ISBN 0-596-10149-X (paperback). 465 pp. LCCN QA76.73.J38 M39

2007. URL http://www.loc.
gov/catdir/enhancements/
fy0715/2007273941-d.html;
http://www.loc.gov/catdir/
toc/fy0710/2007273941.html.

Masala:2002:JBG

[MCLC02] S. Masala, D. Chan, H. Lu, and R. Chalaturnyk. A Javabased graphical user interface for a 2-D discrete element program. Geotechnical Special Publication, 117:125–132, 2002. CODEN ???? ISSN 0895-0563.

Marchand:2001:APG

[MCLDP01] Richard Marchand, Mathieu Charbonneau-Lefort, Mathieu Dumberry, and Benoit Pronovost. ARANEA, a program for generating unstructured triangular meshes with a Java Graphics User Interface. Computer Physics Communications, 139(2):172–185, September 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465501001874.

Machover:2000:NPH

[MD00] Carl Machover and John Dill. New products: Hardware: Modeling system for office environment; smart fabric control surface support; head tracker enables VR; programmable 3D/2D input device; digital projectors; portable 3D scene digitizer; slim-line panel PC. software: Visual effects upgrade; mobile CAD; 3D development tool; visual interpretation of financial data; software development for Linux; feature recognition program; Altia Java connection. IEEE Computer Graphics and Applications, 20(6):110–112, November/December 2000. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). URL http:

//dlib.computer.org/cg/
books/cg2000/pdf/g6110.
pdf.

Marrs:2006:JWP

[MD06] Tom Marrs and Scott Davis. JBoss at work: a practical guide. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2006. ISBN 0-596-00734-5. xvi + 287 pp. LCCN TK5105.8885.J43; TK5105.8885.J43 M37 2006eb.

URL http://www.oreilly.
com/catalog/9780596007348.

Martin:2001:ATG

[MdB01] Hugues Martin and Lydie du Bousquet. Automatic test generation for Java-card applets. Lecture Notes in Computer Science, 2041:121–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2041/20410121.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2041/20410121.
pdf.

Moreau:2005:BDR

[MDJ05] Luc Moreau, Peter Dickman, and Richard Jones. Birrell’s distributed reference listing revisited. ACM Transactions on Programming Languages and Systems, 27(6): 1344–1395, November 2005. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Mahmoud:2004:RIC

[MDS04] Q. H. Mahmoud, W. Dobosiewicz, and D. Swayne. Redesigning introductory computer programming with HTML, JavaScript, and Java. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36 (1):120–124, 2004. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Melton:2000:USJ

[ME00a] Jim Melton and Andrew Eisenberg. Understanding SQL and Java Together: a Guide to SQLJ, JDBC, and Related Technologies. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2000. ISBN 1-55860-562-2. xxix + 514 pp. LCCN QA76.73.S67 M43 2000. US$49.95.

URL http://www.mkp.com/
books_catalog/catalog.asp?
ISBN=1-55860-562-2.

Moon:2000:JTC

[ME00b] SooMook Moon and Kemal Ebcioğlu. A justin-time compiler. Computer, 33(3):41, March 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//dlib.computer.org/co/
books/co2000/pdf/r3040.
pdf.

Mehner:2002:JUB

[Meh02] Katharina Mehner. JaVis: a UML-based visualization and debugging environment for concurrent Java programs. Lecture Notes in Computer Science, 2269:163–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2269/22690163.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2269/22690163.
pdf.

Mengant:2000:WJC

[Men00] Jean-Yves Mengant. Writing a Java class to manage RPM package content. Linux Journal, 76:??, August 2000. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Mengant:2003:NBJ

[Men03] J. Y. Mengant. A.NET bridge to a Java Virtual Machine: Java and .NET interoperability, with a little help from C++. C/C++ Users Journal, 21(10):28, 2003. CODEN CCUJEX. ISSN 1075-2838.

Merzbacher:2000:TDM

[Mer00] Matthew Merzbacher. Teaching database management systems with Java. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1): 31–35, March 2000. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Merson:2004:MJR

[Mer04] P. Merson. Managing J2EE risks: If you’re making the leap to distributed application development with Java 2 Enterprise Edition, take heart: You’re smack in the middle of the bell curve. Here’s a handy guide to assessing whether your team has what it takes to succeed with J2EE. Software Development, 12(7):44–47, 2004. CO-DEN ???? ISSN 1070-8588.

Metsker:2001:BPJ

[Met01] Steven John Metsker. Building Parsers With Java. Addison-Wesley, Reading, MA, USA, 2001. ISBN 0-201-71962-2 (paperback). xxv + 371 pp. LCCN QA76.73.J38 M48 2001. US$40.00. Includes CD-ROM.

Metsker:2002:DPJ

[Met02] Steven John Metsker. Design patterns Java workbook. The software patterns series. Addison-Wesley, Reading, MA, USA, 2002. ISBN 0-201-74397-3 (paperback). xviii + 475 pp. LCCN QA76.73.J38 M483 2002.

Meyer:2003:CIC

[Mey03] M. Meyer. Creating .INI classes in Java. Dr. Dobb’s Journal of Software Tools, 28 (8):38–44, 2003. CODEN DDJOEB. ISSN 1044-789X.

Mikheev:2001:CCM

[MF01a] Vitaly V. Mikheev and Stanislav A. Fedoseev. Compilercooperative memory management in Java. Lecture Notes in Computer Science, 2244:511–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2244/22440511.
htm;   http://link.springer-

ny.com/link/service/series/
0558/papers/2244/22440511.
pdf.

Morgenthal:2001:EAI

[MF01b] J. P. (Jeffrey P.) Morgenthal and Bill La Forge. Enterprise Application Integration with XML and Java. The definitive XML series from Charles F. Goldfarb. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2001. ISBN 0-13-085135-3. xix + 504 pp. LCCN QA76.76.H94 M6 2001. US$44.99. URL http:/

/www.phptr.com/ptrbooks/
ptr_0130851353.html. In-

cludes CD-ROM.

Moreno:2003:FDC

[MF03] R. A. Moreno and S. S. Furuie. Framework for the development of the clinical image access service using Java [5033-55]. Proceedings of the SPIE — The International Society for Optical Engineering, 5033:432–439, 2003. CODEN PSISDG. ISSN 0277-786X (print), 1996-756X (electronic).

McLaughlin:2004:JTD

[MF04] Brett McLaughlin and David Flanagan. Java 1.5 Tiger: a developer’s notebook. The developer’s notebook series. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2004. ISBN 0-596-00738-8. xix + 177 pp. LCCN QA76.73.J38 M3897 2004. URL http://

                                        www.oreilly.com/catalog/
                                        9780596007386.

Ma:2007:IAE

[MF07a] Kin-Keung Ma and Jeffrey S. Foster. Inferring aliasing and encapsulation properties for Java. ACM SIGPLAN Notices, 42(10):423–440, October 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Matthews:2007:OSM

[MF07b] Jacob Matthews and Robert Bruce Findler. Operational semantics for multi-language programs. ACM SIGPLAN Notices, 42(1):3–10, January 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Matthews:2009:OSM

[MF09] Jacob Matthews and Robert Bruce Findler. Operational semantics for multi-language programs. ACM Transactions on Programming Languages and Systems, 31(3): 12:1–12:44, April 2009. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

McDirmid:2001:JNA

[MFH01] Sean McDirmid, Matthew Flatt, and Wilson C. Hsieh. Jiazzi: new-age components for old-fashioned Java. ACM SIGPLAN Notices, 36(11): 211–222, November 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Proceedings of the 2001 ACM SIGPLAN Conference on Object Oriented Programming, Systems, Languages and Applications (OOPSLA’01).

Ma:2007:IVM

[MFRW07] Linxiao Ma, John Ferguson, Marc Roper, and Murray Wood. Investigating the viability of mental models held by novice programmers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39 (1):499–503, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Millstein:2009:EMP

[MFRW09] Todd Millstein, Christopher Frost, Jason Ryder, and Alessandro Warth. Expressive and modular predicate dispatch for Java. ACM Transactions on Programming Languages and Systems, 31(2):7:1–7:54, February 2009. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Mikheev:2002:EEL

[MFSL02] Vitaly V. Mikheev, Stanislav A.

Fedoseev, Vladimir V. Sukharev, and Nikita V. Lipsky. Effective enhancement of loop versioning in Java. Lecture Notes in Computer Science, 2304:293–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2304/23040293.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2304/23040293.
pdf.

Meyerovich:2009:FPL

[MGB^{+}09] Leo A. Meyerovich, Arjun Guha, Jacob Baskin, Gregory H. Cooper, Michael Greenberg, Aleks Bromfield, and Shriram Krishnamurthi. Flapjax: a programming language for Ajax applications. ACM SIGPLAN Notices, 44(10):1–20, October 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Menon:2006:VSP

[MGM^{+}06] Vijay S. Menon, Neal Glew, Brian R. Murphy, Andrew McCreight, Tatiana Shpeisman, Ali-Reza Adl-Tabatabai, and Leaf Petersen. A verifiable SSA program representation for aggressive compiler optimization. ACM SIGPLAN Notices, 41(1):397–408, January 2006. CODEN SINODQ.

ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Miyashita:2000:JAV

[MH00a] Kensuke Miyashita and Reiji Hashimoto. A Java applet to visualize algorithms on reconfigurable mesh. Lecture Notes in Computer Science, 1800:137–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/1800/18000137.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/1800/18000137.
pdf.

Monson-Haefel:2000:EJ

[MH00b] Richard Monson-Haefel. Enterprise JavaBeans. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2000. ISBN 1-56592-869-5. xvii + 472 pp. LCCN QA76.73.J38 M65 2000.

Monson-Haefel:2001:EJ

[MH01] Richard Monson-Haefel. Enterprise JavaBeans. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, third edition, 2001. ISBN 0-596-00226-2. xix + 567 pp. LCCN QA76.73.J38 M65 2001. Covers EJB 2.0.

Miecznikowski:2002:DJB

[MH02] Jerome Miecznikowski and Laurie Hendren. Decompiling Java bytecode: Problems, traps and pitfalls. Lecture Notes in Computer Science, 2304:111–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.
                                        com/link/service/series/
                                        0558/bibs/2304/23040111.
                                        htm;   http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2304/23040111.
                                        pdf.

Monson-Haefel:2004:EJ

[MH04] Richard Monson-Haefel. Enterprise JavaBeans. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, fourth edition, 2004. ISBN 0-596-00530-X. xxi + 763 pp. LCCN QA76.73.J38 M65 2004. URL

                                        http://www.oreilly.com/
                                        catalog/9780596005306.

Murtagh:2009:HAO

[MH09] Jeanne Murtagh and Drew Hamilton. How Ada object orientation works. ACM SIGADA Ada Letters, 29(3): 5–6, December 2009. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Monson-Haefel:2006:EJ

[MHB06] Richard Monson-Haefel and Bill Burke. Enterprise JavaBeans. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, fifth edition, 2006. ISBN 0-596-00978-X. 760 (est.) pp. LCCN ????

URL http://www.oreilly.
com/catalog/entjbeans5/.

Monson-Haefel:2001:JMS

[MHC01] Richard Monson-Haefel and David Chappell. Java message service. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2001. ISBN 0-596-00068-5. xvi + 220 pp. LCCN QA76.73.J38 M655 2001.

Menth:2006:TPP

[MHZG06] Michael Menth, Robert Henjes, Christian Zepfel, and Sebastian Gehrsitz. Throughput performance of popular JMS servers. ACM SIGMET-RICS Perform. Eval. Rev., 34 (1):367–368, June 2006. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Matsuoka:2001:TPE

[MI01] Satoshi Matsuoka and Shigeo Itou. Towards performance evaluation of highperformance computing on multiple Java platforms. Future Generation Computer Systems, 18(2):281–291, October 2001. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.

elsevier.com/gej-ng/10/
19/19/60/31/34/abstract.
html.

Midkiff:2001:JCM

[Mid01] Sam Midkiff. A Java compiler for many memory models. In USENIX Association [USE01c], page ?? ISBN 1-880446-11-1. LCCN QA76.73.J38 J42 2001. URL

http://www.usenix.org/publications/
library/proceedings/jvm01/
JVM_wips/S11.pdf. Spon-

sored by the USENIX Association.

Miles:2005:AC

[Mil05] Russ Miles. AspectJ cookbook. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00654-

3. xix + 331 pp. LCCN QA76.64 .M54 2005 .M55 2004; QA76.64 .M55 2004; QA76.64 .M54 2005; QA76.64 .M54 2005eb; QA76.64. URL

http://www.oreilly.com/
catalog/9780596006549.

Miller:2008:BRP

[Mil08]  J. Miller.  Book review:

Practical API Design: Confessions of a Java Framework Architect (1st edition), Jaroslav Tulach, Apress, 2008, $74.99, ISBN: 978-1-4302-0973-7. ACM Queue: Tomorrow’s Computing Today, 6(6):54, October 2008. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Milner:2009:BMJ

[Mil09] Walter W. Milner. A broken metaphor in Java. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(4):76– 77, December 2009. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Milde:2000:EUV

[MJ00] Jan-Torsten Milde and Bernhard Jung. Educational use of VRML and Java in agent-based AI and computer graphics. Future Generation Computer Systems, 17(1):79– 87, September 2000. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/24/34/abstract.
html.

MacAuley:2001:JPR

[MJ01] Christian MacAuley and Paul Jobson. JavaScript programmer’s reference. Osborne/McGraw-Hill, Berkeley, CA, USA, 2001. ISBN 0-07-219296-

8. xiii + 272 pp. LCCN QA76.76.J39 M34 2001.

Muthukumar:2006:YSG

[MJ06] R. M. Muthukumar and D. Janakiram. Yama: a scalable generational garbage collector for Java in multiprocessor systems. IEEE Transactions on Parallel and Distributed Systems, 17(2):148– 159, February 2006. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Montgomery:2001:FIF

[MK01] Michael Montgomery and Ksheerabdhi Krishna. A flexible invocation framework for Java card. Lecture Notes in Computer Science, 2140:188– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2140/21400188.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2140/21400188.
pdf.

Murphy:2006:HJS

[MKF06] Gail C. Murphy, Mik Kersten, and Leah Findlater. How are Java software developers using the Eclipse IDE? IEEE Software, 23(4):76–83, July/August 2006. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Murphy:2008:BTD

[MKKC08] Christian Murphy, Eunhee Kim, Gail Kaiser, and Adam Cannon. Backstop: a tool for debugging runtime errors. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):173–177, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Mohapatra:2006:DDS

[MKM^{+}06] Durga P. Mohapatra, Rajeev Kumar, Rajib Mall, D. S. Kumar, and Mayank Bhasin. Distributed dynamic slicing of Java programs. The Journal of systems and software, 79(12):1661–1678, December 2006. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Murray:2003:EIJ

[MKS^{+}03] K. A. Murray, M. Kolling, N. C. Schaller, J. M. Heines, T. Moore, P. J. Wagner, and J. A. Trono. Experiences with IDEs and Java teaching: What works and what doesn’t. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):215–216, 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Myers:2000:PPU

[ML00] Andrew C. Myers and Barbara Liskov. Protecting privacy using the decentralized label model. ACM Transactions on Software Engineering and Methodology, 9 (4):410–442, October 2000. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL

http://www.acm.org/pubs/

                                        articles/journals/tosem/
                                        2000-9-4/p410-myers/p410-
                                        myers.pdf;   http://www.
                                        acm.org/pubs/citations/
                                        journals/tosem/2000-9-4/
                                        p410-myers/.

Malan:2007:SBC

[ML07] David J. Malan and Henry H. Leitner. Scratch for budding computer scientists. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):223– 227, March 2007. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Makela:2009:CBC

[ML09] Sami Mäkelä and Ville Leppänen. Client-based cohesion metrics for Java programs. Science of Computer Programming, 74(5–6):355– 378, March 1, 2009. CO-DEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Mazumdar:2002:JBC

[MLG02a] S. Mazumdar, J. F. Leathrum, and O. R. Gonzalez. A JAVA based circuit diagram tool for an intelligent questioning and problem solving system. Simulation series, 34(1):3–8, 2002. CODEN SMCPAX. ISSN 0735-9276.

Mikheev:2002:OEJ

[MLG^{+}02b] V. Mikheev, N. Lipsky, D. Gurchenkov, P. Pavlov, V. Sukharev, A. MarkovS. Kuksenko, S. Fedoseev, D. Leskov, and A. Yeryomin. Overview of Excelsior JET, a high performance alternative to Java Virtual Machines. In Proceedings of the third international workshop on Software and performance: WOSP ’02, July 24–26, 2002 Rome, Italy, pages 104–113. ACM Press, New York, NY 10036, USA, 2002.

Meunier:2004:MRT

[MLJH04] J. N. Meunier, F. Lippert, R. Jadhav, and N. Harding. MDA and real-time java: The HIDOORS project. Technical Report — University of Kent at Canterbury Computing Laboratory, 17:89–95, 2004. CODEN ???? ISSN 1368-1036.

Murphy:2008:DGB

[MLM^{+}08] Laurie Murphy, Gary Lewandowski, Renée McCauley, Beth Simon, Lynda Thomas, and Carol Zander. Debugging: the good, the bad, and the quirky — a qualitative analysis of novices’ strategies. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40 (1):163–167, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Mlsna:2004:WPM

[Mls04]  P. A. Mlsna.  Work in

progress: Mastery of digital logic skills through practice using Java applets. Frontiers in Education Conference, CONF34(VOL 1):T1D– 13–T1D14, 2004. CODEN PFECDR. ISSN 0190-5848.

Markidis:2005:IPP

[MLVB05] S. Markidis, G. Lapenta, W. B. VanderHeyden, and Z. Budimlić. Implementation and performance of a particle-in-cell code written in Java. Concurrency and Computation: Practice and Experience, 17(7–8):821– 837, June/July 2005. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Moodley:2004:CMP

[MM04] K. Moodley and H. Murrell. A colour-map plugin for the open source, Java based, image processing package, ImageJ. Computers and Geosciences, 30 (6):609–618, 2004. CODEN CGEODT, CGOSDN. ISSN 0098-3004 (print), 1873-7803 (electronic).

Moreno:2004:PAJ

[MMBAS04] Andrés Moreno, Niko Myller, Mordechai Ben-Ari, and Erkki Sutinen. Program animation in Jeliot 3. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3): 265, September 2004. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Moreira:2000:JPH

[MMG^{+}00a] J. E. Moreira, S. P. Midkiff, M. Gupta, P. V. Artigas, M. Snir, and R. D. Lawrence. Java programming for high-performance numerical computing. IBM Systems Journal, 39(1):21– 56, ???? 2000. CODEN IBMSA7. ISSN 0018-8670.

URL http://www.almaden.
ibm.com/journal/sj/391/
moreira.html.

Moreira:2000:FMJ

[MMG00b] José E. Moreira, Samuel P.

Midkiff, and Manish Gupta. From flop to megaflops: Java for technical computing. ACM Transactions on Programming Languages and Systems, 22(2): 265–295, March 2000. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/toplas/2000-22-
2/p265-moreira/.

Moreira:2001:CTA

[MMG01a] José Moreira, Sam Midkiff, and Manish Gupta. A comparison of three approaches to language, compiler, and library support for multidimensional arrays in Java. In ACM [ACM01b], pages 116–125. ISBN 1-58113-359-6. LCCN QA76.9.O35 A26

                                        2001. URL http://www.
                                        philippsen.com/JGI2001/
                                        camerareadyabstracts/44.
                                        html;   http://www.philippsen.
                                        com/JGI2001/finalpapers/
                                        18500116.ps.

Moreira:2001:NP

[MMG^{+}01b] José E. Moreira, Samuel P. Midkiff, Manish Gupta, Pedro V. Artigas, Peng Wu, and George Almasi. The NINJA project. Communications of the ACM, 44(10): 102–109, October 2001. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Moreira:2002:NJH

[MMG^{+}02] José E. Moreira, Samuel P. Midkiff, Manish Gupta, et al. NINJA: Java for high performance numerical computing. Scientific Programming, 10(1):19–33, 2002. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:/

                                        /iospress.metapress.com/
                                        app/home/contribution.asp%
                                        3Fwasp=9ejnuvwuvby9737jte27%
                                        26referrer=parent%26backto=
                                        issue%2C3%2C9%3Bjournal%
                                        2C2%2C12%3Blinkingpublicationresults%
                                        2C1%2C1.

Moreira:2003:SMA

[MMG03] José E. Moreira, Samuel P. Midkiff, and Manish Gupta. Supporting multidimensional arrays in Java. Concurrency and Computation: Practice and Experience, 15(3–5):317– 340, March/April 2003. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mohapatra:2004:ETD

[MMK04] D. P. Mohapatra, R. Mall, and R. Kumar. An efficient technique for dynamic slicing of concurrent Java programs. Lecture Notes in Computer Science, 3285:255–262, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McCown:2009:WWS

[MMN09] Frank McCown, Catherine C.

Marshall, and Michael L. Nelson. Why Web sites are lost (and how they’re sometimes found). Communications of the ACM, 52(11): 141–145, November 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Marche:2004:KTC

[MMU04] C. Marche, C. Paulin Mohring, and X. Urbain. The KRAKA-TOA tool for certification of JAVA/JAVACARD programs annotated in JML. Journal of Logic and Algebraic Programming, 58(1–2):89–106, 2004. CODEN ???? ISSN 1567-8326.

Massol:2005:MDN

[MOL05] Vincent Massol, Timothy O’Brien, and Michael Kosta Loukides. Maven: a developer’s notebook. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2005. ISBN 0-596-00750-7. xxv + 191 pp. LCCN QA76.76.D47 M37 2005.

Moore:2002:BED

[Moo02] Thomas K. Moore. Bringing the enterprise into a database systems course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):262– 265, March 2002. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Moore:2003:PTA

[Moo03a] J. S. Moore. Proving theorems about Java and the JVM with ACL2. NATO Science Series Sub Series III Computer and Systems Sciences, 191:227–290, 2003. CODEN ???? ISSN 1387-6694.

Moore:2003:SHS

[Moo03b] R. Moore. Sun hasn’t set yet on Java applets. Application Development Trends, 10(1):33–38, 2003. CODEN ADTRF4. ISSN 1073-9564.

Moore:2006:IAO

[Moo06] J. Strother Moore. Inductive assertions and operational semantics. International Journal on Software Tools for Technology Transfer (STTT), 8(4–5):359–371, August 2006. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=1433-2779&volume=8&
issue=4&spage=359.

Morelli:2000:JJJ

[Mor00] R. (Ralph) Morelli. Java, Java, Java!: object-oriented problem solving. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2000. ISBN 0-13-011332-8. xviii + 964 pp. LCCN QA76.64 .M64 2000. CD-ROM includes Java code for all the examples in the text, Java 2 SDK (Standard Edition), Borland JBuilder 3 (University Edition) and Net-Beans. BUY-A-BOOK gift of Brendon O’Brien in honor of Amy S. O’Brien.

Morris:2002:AGJ

[Mor02] D. S. Morris. Automatically grading Java programming assignments via reflection, inheritance, and regular expressions. Frontiers in Education Conference, 1(??): T3G–22–??, 2002. CODEN PFECDR. ISSN 0190-5848.

Morelli:2003:JJJ

[Mor03a] R. (Ralph) Morelli. Java, Java, Java!: object-oriented problem solving. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, second edition, 2003. ISBN 0-13-033370-0, 0-13-060813-0 (CD-ROM). xxx + 862 pp. LCCN QA76.64 .M64 2003. Includes CD-ROM.

Morgan:2003:BRA

[Mor03b] Donald Morgan. Book review: Ant book full of practical advice: a review of Java Development with Ant by Eric Hatcher and Steve Loughran. IEEE Distributed Systems Online, 4(7), 2003. ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:

                                        //dsonline.computer.org/
                                        0307/d/bks_b.htm.

Morrisett:2003:AIC

[Mor03c] Greg Morrisett. Analysis issues for cyclone. ACM SIGSOFT Software Engineering Notes, 28(1):26, January 2003. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Morrison:2008:ACK

[Mor08a] Michael Morrison. Ajax construction kit: building plugand-play Ajax applications. Negus live Linux series. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2008. ISBN 0-13-235008-4 (paperback). xvii + 291 pp. LCCN TK5105.8885.A52 M67 2008.

Morrison:2008:HFJ

[Mor08b] Michael Morrison. Head first JavaScript. Head first series. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2008. ISBN 0-596-52774-8. xxxiv + 615 pp. LCCN QA76.73.J39eb; QA76.73.J39. URL http://

www.oreilly.com/catalog/
9780596527747.

Moller:2004:LCO

[MORW04] M. Moller, E. R. Olderog, H. Rasch, and H. Wehrheim. Linking CSP-OZ with UML and Java: a case study. Lecture Notes in Computer Science, 2999(??):267–286, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moller:2008:IFM

[MORW08] Michael Möller, Ernst-Rüdiger Olderog, Holger Rasch, and Heike Wehrheim. Integrating a formal method into a software engineering process with UML and Java. Formal Aspects of Computing, 20(2): 161–204, March 2008. CO-DEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s00165-007-0042-7.

Moss:2000:JQ

[Mos00] Karl Moss. Java Q&A: How can I measure Java code performance? Dr. Dobb’s Journal of Software Tools, 25 (10):135–136, 138, 140–142, 144, October 2000. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2000/2000_10/jqa010.
txt.

Mostowski:2005:FDS

[Mos05a] W. Mostowski. Formal Development of Safe and Secure Java Card Applets. Doktorsavhandlingar, Chalmers Tekniska Högskola, Göteborg, Sweden, 2005. ??–?? pp.

Mostowski:2005:FVJ

[Mos05b] W. Mostowski. Formalisation and verification of Java Card security properties in dynamic logic. Lecture Notes in Computer Science, 3442:357– 371, 2005. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Muller-Olm:2007:AMA

[MOS07] Markus Müller-Olm and Helmut Seidl. Analysis of modular arithmetic. ACM Transactions on Programming Languages and Systems, 29(5): 29:1–29:27, August 2007. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Manson:2001:CSM

[MP01a] Jeremy Manson and William Pugh. Core semantics of multithreaded Java. In ACM [ACM01b], pages 29–

38. ISBN 1-58113-359-

6. LCCN QA76.9.O35 A26

                                        2001. URL http://www.
                                        philippsen.com/JGI2001/
                                        camerareadyabstracts/42.
                                        html;   http://www.philippsen.
                                        com/JGI2001/finalpapers/
                                        18500029.pdf.

Meijer:2001:TFF

[MP01b] Hans Meijer and Erik Poll.

Towards a full formal specification of the JavaCard API. Lecture Notes in Computer Science, 2140:165–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2140/21400165.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2140/21400165.
pdf.

Moore:2001:EFJ

[MP01c] J. Strother Moore and George M. Porter. An executable formal Java Virtual Machine thread model. In USENIX Association [USE01c], page ?? ISBN 1-880446-11-1. LCCN QA76.73.J38

J42 2001. URL http://www.
usenix.org/publications/
library/proceedings/jvm01/

moore.html. Sponsored by the USENIX Association.

Masri:2005:UDI

[MP05] Wes Masri and Andy Podgurski.

Using dynamic information flow analysis to detect attacks against applications. ACM SIGSOFT Software Engineering Notes, 30(4):1–7, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Manson:2005:JMM

[MPA05] Jeremy Manson, William Pugh, and Sarita V. Adve. The Java memory model. ACM SIGPLAN Notices, 40 (1):378–391, January 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Malabarba:2000:RST

[MPG^{+}00] Scott Malabarba, Raju Pandey, Jeff Gragg, Earl Barr, and J. Fritz Barnes. Runtime support for type-safe dynamic Java classes. Lecture Notes in Computer Science, 1850:337– ??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.
                                        com/link/service/series/
                                        0558/bibs/1850/18500337.
                                        htm;   http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/1850/18500337.
                                        pdf.

Moors:2008:GHK

[MPO08] Adriaan Moors, Frank Piessens, and Martin Odersky. Generics of a higher kind. ACM SIGPLAN Notices, 43(10): 423–438, September 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Muschevici:2008:MDP

[MPTN08] Radu Muschevici, Alex Potanin, Ewan Tempero, and James Noble. Multiple dispatch in practice. ACM SIGPLAN Notices, 43(10):563–582, September 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Malkhi:2000:SEJ

[MR00a] D. Malkhi and M. K. Reiter. Secure execution of Java applets using a remote playground. IEEE Transactions on Software Engineering, 26 (12):1197–1209, December 2000. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
888632.

Mughal:2000:PGJ

[MR00b] Khalid Azim Mughal and Rolf W. Rasmussen. A programmer’s guide to Java certification: a comprehensive primer. Addison-Wes-ley, Reading, MA, USA, 2000. ISBN 0-201-59614-8 (paperback). xxxv + 754 pp. LCCN QA76.3 .M846 1999. US$44.95.

Moreau:2002:MOJ

[MR02] Luc Moreau and Daniel Ribbens. Mobile objects in Java. Scientific Programming, 10(1):91–100, 2002. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:/

/iospress.metapress.com/
app/home/contribution.asp%

                                        3Fwasp=9ejnuvwuvby9737jte27%
                                        26referrer=parent%26backto=
                                        issue%2C9%2C9%3Bjournal%
                                        2C2%2C12%3Blinkingpublicationresults%
                                        2C1%2C1.

Markov:2006:IWD

[MR06] Zdravko Markov and Ingrid Russell. An introduction to the WEKA data mining system. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):367–368, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Marchetto:2009:OST

[MR09] Alessandro Marchetto and Filippo Ricca. From objects to services: toward a stepwise migration approach for Java applications. International Journal on Software Tools for Technology Transfer (STTT), 11(6):427–440, December 2009. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=1433-2779&volume=11&
                                        issue=6&spage=427.

Markow:2006:CST

[MRB06] Tanya Markow, Eugene Ressler, and Jean Blair. Catch that speeding turtle: latching onto fun graphics in CS1. ACM SIGADA Ada Letters, 26(3):29–34, December 2006. CODEN AALEE5.

ISSN 1094-3641 (print), 1557-9476 (electronic).

Millstein:2003:RMB

[MRC03] Todd Millstein, Mark Reay, and Craig Chambers. Relaxed MultiJava: balancing extensibility and modular typechecking. ACM SIGPLAN Notices, 38(11):224–240, November 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Milanova:2002:POS

[MRR02] Ana Milanova, Atanas Rountev, and Barbara G. Ryder. Parameterized object sensitivity for points-to and sideeffect analyses for Java. ACM SIGSOFT Software Engineering Notes, 27(4):1–11, July 2002. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Milanova:2005:POS

[MRR05] Ana Milanova, Atanas Rountev, and Barbara G. Ryder. Parameterized object sensitivity for points-to analysis for Java. ACM Transactions on Software Engineering and Methodology, 14(1): 1–41, January 2005. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Maessen:2000:IJM

[MS00a] Jan-Willem Maessen and Xiaowei Shen. Improving the Java memory model using CRF. ACM SIGPLAN Notices, 35(10):1–12, October 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
proceedings/oops/353171/
p1-maessen/.

Mathiske:2000:APM

[MS00b] Bernd Mathiske and Daniel Schneider. Automatic persistent memory management for the Spotless virtual machine on the Palm connected organizer. SMLI TR 2000-89, Sun Microsystems, Palo Alto, CA, USA, June 2000. 18 pp.

Matena:2001:AEJ

[MS01] Vlada Matena and Beth Stearns. Applying Enterprise JavaBeans: Component-Based Development for the J2EE Platform. Java series. Addison-Wesley, Reading, MA, USA, 2001. ISBN 0-201-70267-3 (paperback). xxv + 436 pp. LCCN QA76.73.J38 M377 2001. US$39.95.

Mitchell:2003:LAL

[MS03] N. Mitchell and G. Sevitsky. LeakBot: An automated and lightweight tool for diagnosing memory leaks in large Java applications. Lecture Notes in Computer Science, 2743:351–377, 2003. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marrero:2005:TFE

[MS05] Will Marrero and Amber Settle. Testing first: emphasizing testing in early programming courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):4–8, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Metzger:2003:MBP

[MSF03] J. Metzger, M. Schillo, and K. Fischer. A multiagentbased peer-to-peer network in Java for distributed spam filtering. Lecture Notes in Computer Science, 2691:616–625, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maessen:2001:PAS

[MSG01] Jan-Willem Maessen, Vivek Sarkar, and David Grove. Program analysis for safety guarantees in a Java virtual machine written in Java. In ACM [ACM01a], pages 62–65. ISBN 1-58113-413-4. LCCN QA76.758 .A345 2001.

URL http://www.acm.org/
pubs/articles/proceedings/
soft/379605/p62-maessen/
p62-maessen.pdf;   http:
//www.acm.org/pubs/citations/
proceedings/soft/379605/
p62-maessen/. Supplement

to ACM SIGPLAN Notices.

Miura:2009:AGI

[MSK09] Motoki Miura, Taro Sugihara, and Susumu Kunifuji. Anchor Garden: an interactive workbench for basic data concept learning in object oriented programming languages. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):141–145, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

McCreight:2007:GFC

[MSLL07] Andrew McCreight, Zhong Shao, Chunxiao Lin, and Long Li. A general framework for certifying garbage collectors and their mutators. ACM SIGPLAN Notices, 42(6):468–479, June 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mattson:2005:PPP

[MSM05] Timothy G. Mattson, Beverly A. Sanders, and Berna Massingill. Patterns for Parallel Programming. Addison-Wesley, Reading, MA, USA, 2005. ISBN 0-321-22811-1 (hardcover). xiii + 355 pp. LCCN QA76.642 .M38

2005. URL http://www.loc.
gov/catdir/toc/ecip0418/
2004013240.html.

Miller:2003:OCP

[MSR03] W. W. Miller, C. Sontag, and J. F. Rose. OPUS: a CORBA pipeline for Java, Python, and Perl applications. Astronomical Society of the Pacific Conference Series, 295:261–264, 2003. CODEN ???? ISSN 1050-3390.

Malik:2009:SCU

[MSR09] Avinash Malik, Zoran Salcic, and Partha S. Roop. SystemJ compilation using the Tandem Virtual Machine approach. ACM Transactions on Design Automation of Electronic Systems, 14(3): 34:1–34:??, May 2009. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Migliardi:2000:DJS

[MSS00] Mauro Migliardi, Simon Schubiger, and Vaidy Sunderam. A distributed JAVA SPACE implementation for HARNESS. Journal of Parallel and Distributed Computing, 60(10):1325–1340, October 2000. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jpdc.
2000.1656;   http://www.
idealibrary.com/links/doi/
10.1006/jpdc.2000.1656/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.
2000.1656/ref.

Murray:2000:PIM

[MSSJ00] Paul Murray, Todd Smith, Suresh Srinivas, and Matthias Jacob. Performance issues for multi-language Java applications. Lecture Notes in Computer Science, 1800:544– ??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.
                                        com/link/service/series/
                                        0558/bibs/1800/18000544.
                                        htm;   http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/1800/18000544.
                                        pdf.

Mathiske:2008:ADF

[MSU08] Bernd Mathiske, Doug Simon, and Dave Ungar. An assembler and disassembler framework for JavaTM programmers. Science of Computer Programming, 70(2–3): 127–148, February 1, 2008. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Moir:2005:CSJ

[MSV05] Mark Moir, Nir Shavit, and Jan Vitek. Concurrency and synchronization in Java programs. Science of Computer Programming, 58(3): 291–292, December 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Melton:2007:ESC

[MT07] Hayden Melton and Ewan Tempero. An empirical study of cycles among classes in Java. Empirical Software Engineering, 12(4):389– 415, August 2007. CO-DEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (elec-

tronic). URL http://link.
springer.com/accesspage/
article/10.1007/s10664-
006-9033-1.

McGovern:2003:JWS

[MTSM03] James McGovern, Sameer Tyagi, Michael Stevens, and Sunil Mathew. Java Web Services Architecture. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2003. ISBN 1-55860-900-8 (paperback). xii + 833 pp. LCCN TK5105.88813 J38 2003. US$59.95. URL ftp:/

/uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
description/els051/2002117799.
html;   http://www.loc.
gov/catdir/toc/els031/2002117799.
html.

Muchow:2002:CJT

[Muc02] John W. Muchow. Core J2ME Technology and MIDP. Sun Microsystems Press Java series. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2002. ISBN 0-13-066911-3. xxiii + 710 pp. LCCN QA76.73.J38 M83 2002. US$49.99. URL http:/

                                        /www.phptr.com/ptrbooks/
                                        ptr_0130669113.html;   http:
                                        //www.sun.com/books/catalog/
                                        muchow/.

Muldner:2000:CJP

[Mul00] Tomasz Muldner. C for Java programmers. Addison-Wes-ley Longman, Reading, MA, USA, 2000. ISBN 0-201-70279-7. various pp. LCCN QA76.73.C15 M853 2000.

Murdock:2000:JYV

[Mur00] Kelly Murdock. JavaScript: your visual blueprint for building dynamic Web pages. I D G Books Worldwide, Indianapolis, IN, USA, 2000. ISBN 0-7645-4730-5. xvii + 285 pp. LCCN QA76.73.J39 M87 2000.

Murtagh:2005:CAD

[Mur05]  Fionn Murtagh.  Corre-

spondence analysis and data coding with Java and R. Computer science and data analysis series. Chapman and Hall/CRC, Boca Raton, FL, USA, 2005. ISBN 1-58488-528-9. 230 pp. LCCN QA76.73.J38 M877

                                        2005. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0648/2005041353-d.html.

Murtagh:2007:SBV

[Mur07] Thomas P. Murtagh. Squint: barely visible library support for CS1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):526–530, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Muir:2009:IGE

[MV09] James A. Muir and Paul C.

Van Oorschot. Internet geolocation: Evasion and counterevasion. ACM Computing Surveys, 42(1):4:1– 4:23, December 2009. CO-DEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

Marin:2007:ICC

[MVM07] Marius Marin, Arie Van Deursen, and Leon Moonen. Identifying Crosscutting Concerns Using Fan-In Analysis. ACM Transactions on Software Engineering and Methodology, 17(1):3:1– 3:37, December 2007. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Maassen:2001:EJR

[MVV^{+}01] Jason Maassen, Rob Van Nieuwpoort, Ronald Veldema, Henri Bal, Thilo Kielmann, Ceriel Jacobs, and Rutger Hofman. Efficient Java RMI for parallel programming. ACM Transactions on Programming Languages and Systems, 23(6):747–775, November 2001. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Munawar:2005:BPB

[MW05] Mohammad A. Munawar and Paul A. S. Ward. Better performance or better manageability? ACM SIGSOFT Software Engineering Notes, 30(4):1–4, July 2005. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ma:2000:JJE

[MWL00] Matchy J. M. Ma, Cho-Li Wang, and Francis C. M. Lau. JESSICA: Java-enabled single-system-image computing architecture. Journal of Parallel and Distributed Computing, 60(10):1194–1222, October 2000. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

http://www.idealibrary.
com/links/doi/10.1006/jpdc.
2000.1650;   http://www.
idealibrary.com/links/doi/
10.1006/jpdc.2000.1650/
pdf;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.
2000.1650/ref.

Morelli:2001:JAH

[MWM01] Ralph Morelli, Ralph Walde, and Gregg Marcuccio. A Java API for historical ciphers: an object-oriented design project. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):307– 311, March 2001. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ma:2004:JTP

[MZ04] H. Ma and C. Zhang. JAVA and threads programming. Hebei Journal of Industrial Science and Technology, 21 (4):40–42, 2004. CODEN ???? ISSN 1008-1534.

Marquez:2000:FPO

[MZB00] Alonso Marquez, John N. Zigman, and Stephen M. Blackburn. Fast portable orthogonally persistent Java^{T M}. Software—Practice and Experience, 30(4):449–479, April 10, 2000. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

http://www3.interscience.
wiley.com/cgi-bin/abstract/
71004143/START;   http:/
/www3.interscience.wiley.
com/cgi-bin/fulltext?ID=
71004143&PLACEBO=IE.pdf.

Neward:2000:SBJ

[N^{+}00] Ted Neward et al. Server-Based Java Programming. Manning Publications, Greenwich, CT, USA, 2000. ISBN 1-884777-71-6. xxxiii + 556 pp. LCCN QA76.73.J38 N49 2000. US$44.

Naik:2007:CMA

[NA07] Mayur Naik and Alex Aiken.

Conditional must not aliasing for static race detection. ACM SIGPLAN Notices, 42(1):327–338, January 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nami:2008:COO

[Nam08] Mohammad Reza Nami. A comparison of object-oriented languages in software engineering. ACM SIGSOFT Software Engineering Notes, 33(4):6:1–6:??, July 2008. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Narasimhan:2005:LSJ

[Nar05] Balasubramanian Narasimhan. Lisp–Stat to Java to R. Journal of Statistical Software, 13(4):1–10, 2005. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/counter.php?
                                        id=118&url=v13/i04&ct=2;
                                         http://www.jstatsoft.org/
                                        counter.php?id=118&url=
                                        v13/i04/v13i04.pdf&ct=1.

Nicoara:2008:CSE

[NAR08] Angela Nicoara, Gustavo Alonso, and Timothy Roscoe. Controlled, systematic, and efficient code replacement for running Java programs. Operating Systems Review, 42 (4):233–246, May 2008. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Nash:2004:EGJ

[Nas04]  Michael P. Nash.  Ex-

plorer’s Guide to Java Open Source Tools. Manning Publications, Greenwich, CT, USA, 2004. ISBN 1-932394-19-2. 300 (est.) pp. US$39.95. URL http://

www.manning.com/catalog/
view.php?book=nash.

NASA:2000:EJU

[Nat00] National Aeronautics and Space Administration. Enhancing Jini For Use Across Non-multicastable Networks: Final Report NASA/CR-2000-210329. Washington, DC, USA, December 6, 2000. Shipping list number 2002-0173-M. Shipping list date: 03/04/2002.

Naumovich:2002:CAC

[Nau02] Gleb Naumovich. A conservative algorithm for computing the flow of permissions in Java programs. ACM SIGSOFT Software Engineering Notes, 27(4):33–43, July 2002. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Naik:2006:ESR

[NAW06] Mayur Naik, Alex Aiken, and John Whaley. Effective static race detection for Java. ACM SIGPLAN Notices, 41(6):308–319, June 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nicholas:2000:OTD

[NB00] Tyrone Nicholas and Jerzy A. Barchanski. Overview of TOS: a distributed educational operating system in Java. Operating Systems Review, 34(1):2–10, January 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Nicholas:2001:TED

[NB01] Tyrone Nicholas and Jerzy A. Barchanski. TOS: an educational distributed operating system in Java. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):312– 316, March 2001. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Naumovich:2004:SAR

[NC04a] Gleb Naumovich and Paolina Centonze. Static analysis of role-based access control in J2EE applications. ACM SIGSOFT Software Engineering Notes, 29(5):1– 10, September 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Nepomuceno-Chamorro:2004:JSM

[NC04b] I. A. Nepomuceno-Chamorro. A Java simulator for membrane computing. J.UCS:

Journal of Universal Computer Science, 10(5):620–??, May 28, 2004. CODEN ???? ISSN 0948-6968. URL http:/

/www.jucs.org/jucs_10_5/
a_java_simulator_for.

Neary:2005:AES

[NC05] Michael O. Neary and Peter Cappello. Advanced eager scheduling for Javabased adaptive parallel computing. Concurrency and Computation: Practice and Experience, 17(7–8):797–819, June/July 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Nystrom:2003:PEC

[NCM03] N. Nystrom, M. R. Clarkson, and A. C. Myers. Polyglot: An extensible compiler framework for Java. Lecture Notes in Computer Science, 2622:138–152, 2003. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nagasaki:2002:GON

[NDS^{+}02] M. Nagasaki, A. Doi, M. Sasaki, C. J. Savoie, H. Matsuno, and S. Miyano. Genomic object net in Java: a platform for biopathway modeling and simulation. Genome Informatics Series, SERS13:252– 253, 2002. CODEN ????

Nimmer:2004:SVD

[NE04] J. W. Nimmer and M. D. Ernst. Static verification of dynamically detected program invariants — integrating Daikon and ESC/Java. Electronic Notes in Theoretical Computer Science, 55(2): 1–22, January 2004. CODEN ???? ISSN 1571-0661.

Nelson:2004:ESC

[Nel04] G. Nelson. Extended static checking for Java. Lecture Notes in Computer Science, 3125:1, 2004. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Newmarch:2000:PGJ

[New00] Jan Newmarch. A programmer’s guide to Jini technology. Books for professionals by professionals. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. ISBN 1-893115-80-1. xxi + 448 pp. LCCN QA76.9.D5 N548 2000.

Newhouse:2001:JAE

[New01] Travis Gregory Newhouse. Java active extensions: a mobile-code mechanism for extending client resources. Thesis (M.S.), University of California, San Diego, San Diego, CA, USA, 2001.

Newman:2004:EJC

[New04] D. J. Newman. Embedded Java controllers. Circuit Cellar, 166:16–21, 2004. CODEN ???? ISSN 1528-0608.

Neward:2005:EEJ

[New05] Ted Neward. Effective Enterprise Java. Addison-Wes-ley, Reading, MA, USA, 2005. ISBN 0-321-13000-6 (paperback). xix + 470 pp. LCCN QA76.73.J38 N48

2005. URL http://www.loc.
gov/catdir/toc/ecip0418/
2004012164.html.

Nino:2002:IPO

[NH02] Jaime Nino and Frederick A. Hosch. An introduction to programming and object-oriented design using JAVA. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 2002. ISBN 0-471-35489-9 (paperback). xxx + 733 pp. LCCN QA76.64 .N57 2002. URL http:/

/www.loc.gov/catdir/toc/
onix05/2001017636.html.

Nakano:2004:AVF

[NHY^{+}04] J. Nakano, M. Y. Huh, Y. Yamamoto, T. Fujiwara, and I. Kobayashi. Adding visualization functions of DAVIS to Jasp: Mixing two Java-based statistical systems. Computational Statistics, 19(1):137– 146, 2004. CODEN CSTAEB. ISSN 0943-4062.

Nilsson:2004:IJC

[NIEH04] A. Nilsson, A. Ive, T. Ekman, and G. Hedin. Implementing Java compilers using ReRAGs. Nordic Journal of Computing, 11(3):213–234, 2004. CODEN NJCOFR. ISSN 1236-6064.

Nikishkov:2003:GCF

[Nik03] G. P. Nikishkov. Generating contours on FEM/BEM higher-order surfaces using Java 3D textures. Advances in Engineering Software, 34 (8):469–476, 2003. CO-DEN AESODT. ISSN 0965-9978 (print), 0141-1195 (electronic).

Nakaike:2006:PBG

[NIKN06] Takuya Nakaike, Tatsushi Inagaki, Hideaki Komatsu, and Toshio Nakatani. Profilebased global live-range splitting. ACM SIGPLAN Notices, 41(6):216–227, June 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nilsen:2005:JSD

[Nil05]  K. Nilsen.  Java sounds

the death knell for C++. Electronic Design, 53(11):22, 2005. CODEN ELODAW. ISSN 0013-4872.

Nipkow:2001:VBV

[Nip01] T. Nipkow. Verified bytecode verifiers. Lecture Notes in Computer Science, 2030:347– 363, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nipkow:2003:JBV

[Nip03] Tobias Nipkow. Java bytecode verification. Journal of Automated Reasoning, 30 (3–4):233, May 2003. CO-DEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.
springer.com/accesspage/
article/10.1023/A%3A1025086804452.

NIST:2000:TAE

[NIS00] NIST, editor. The Third Advanced Encryption Standard Candidate Conference, April 13–14, 2000, New York, NY, USA. National Institute for Standards and Technology, Gaithersburg, MD, USA, 2000. ISBN ???? LCCN

???? URL http://csrc.
nist.gov/encryption/aes/
round2/conf3/aes3conf.htm;
 http://csrc.nist.gov/encryption/
aes/round2/conf3/papers/
AES3Proceedings-1.pdf;
http://csrc.nist.gov/encryption/
aes/round2/conf3/papers/
AES3Proceedings-2.pdf;
http://csrc.nist.gov/encryption/
aes/round2/conf3/papers/
AES3Proceedings-3.pdf;
http://csrc.nist.gov/encryption/
aes/round2/conf3/papers/
AES3Proceedings.pdf.

Nisley:2002:ES

[Nis02a] Ed Nisley. Embedded space:

Java: Stirring the cup. Dr. Dobb’s Journal of Software Tools, 27(2):90, 92–93, February 2002. CODEN DDJOEB.

ISSN 1044-789X. URL http:

                                        //www.ddj.com/.

Nisley:2002:ESJ

[Nis02b] Ed Nisley. Embedded space: Java: The mix. Dr. Dobb’s Journal of Software Tools, 27 (3):83–84, 86, March 2002. CODEN DDJOEB. ISSN 1044-789X. URL http://

                                        www.ddj.com/.

Nisley:2003:ELH

[Nis03] E. Nisley. Ed looks at the history of cryptography and examines what it means for embedded systems developers programming in Java today. Dr. Dobb’s Journal of Software Tools, 28(11):73–75, 2003. CODEN DDJOEB. ISSN 1044-789X.

Niemeyer:2000:LJ

[NK00] Patrick Niemeyer and Jonathan Knudsen. Learning Java. Java series. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2000. ISBN 1-56592-718-4 (paperback). xvi + 706 pp. LCCN QA76.73.J38 N545 2000.

Niemeyer:2002:LJ

[NK02] Pat Niemeyer and Jonathan Knudsen. Learning Java. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2002. ISBN 0-596-00285-8. xvii + 807 pp. LCCN QA76.73.J38 N545 2002.

US$44.95. URL http://

safari.oreilly.com/0596002858;
http://www.oreilly.com/
catalog/learnjava2.

Nilsen:2003:IDI

[NK03] K. Nilsen and A. Klein. Issues in the design and implementation of efficient interfaces between hard and soft real-time Java components. Lecture Notes in Computer Science, 2889:451–465, 2003. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Niemeyer:2005:LJ

[NK05] Patrick Niemeyer and Jonathan Knudsen. Learning Java. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, third edition, 2005. ISBN 0-596-00873-2. xx + 954 pp. LCCN QA76.73.J38 N545 2005.

Nagpurkar:2006:PBV

[NK06] Priya Nagpurkar and Chandra Krintz. Phase-based visualization and analysis of Java programs. Science of Computer Programming, 59(1–2): 64–81, January 2006. CO-DEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Nelisse:2001:OBC

[NKBM01] Arnold Nelisse, Thilo Kielmann, Henri Bal, and Jason Maassen. Object-based collective communication in Java. In ACM [ACM01b], pages 11–20. ISBN 1-58113-359-6. LCCN QA76.9.O35

A26 2001. URL http://www.
philippsen.com/JGI2001/
camerareadyabstracts/26.
html;   http://www.philippsen.
com/JGI2001/finalpapers/
18500011.ps.

Nurvitadhi:2003:DCC

[NLC03] E. Nurvitadhi, W. W. Leung, and C. Cook. Do class comments aid Java program understanding? Frontiers in Education Conference, 1(??): T3C–13–T3C–17, 2003. CO-DEN PFECDR. ISSN 0190-5848.

Neelands:2002:UDJ

[NLFA02] K. Neelands, R. Ledbetter, S. Foti, and E. Alkazemi. Using DHTML and Java to synergistically combine textbased and interactive instructional materials. Journal of Interactive Instruction Development, 15(1):19–26, 2002. CODEN ???? ISSN 1040-0370.

Newhall:2000:PMD

[NM00] Tia Newhall and Barton P. Miller. Performance measurement of dynamically compiled Java executions. Concurrency: Practice and Experience, 12(6):343–362, May 2000. CODEN CPEXEI. ISSN 1040-3108. URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/72515726/
START;   http://www3.interscience.

wiley.com/cgi-bin/fulltext?
ID=72515726&PLACEBO=IE.
pdf.

Newhall:2002:CPC

[NM02] Tia Newhall and Lisa Meeden. A comprehensive project for CS2: combining key data structures and algorithms into an integrated Web browser and search engine. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34 (1):386–390, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Nishiyama:2002:SCA

[NMH^{+}02] A. Nishiyama, K. Miyoshi, T. Hikita, K. Tsukamoto, and M. Tsujigado. A study on CORBA applications for sequence control in Ada95 and Java. IECON Proceedings, 3 (??):2397–2402, 2002. CO-DEN ???? ISSN ????

Nelisse:2003:COB

[NMKB03] Arnold Nelisse, Jason Maassen, Thilo Kielmann, and Henri E. Bal. CCJ: object-based message passing and collective communication in Java. Concurrency and Computation: Practice and Experience, 15 (3–5):341–369, March/April 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Narasimhan:2001:IJR

[NMMS01] N. Narasimhan, L. E. Moser, and P. M. Melliar-Smith. Interceptors for Java Remote Method Invocation. Concurrency and Computation: Practice and Experience, 13(8–9):755–774, July/August 2001. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (elec-

                                        tronic). URL http://www3.
                                        interscience.wiley.com/
                                        cgi-bin/abstract/84503222/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=84503222&PLACEBO=IE.
                                        pdf.

Nikishkov:2003:CCJ

[NNS03] G. P. Nikishkov, Y. G. Nikishkov, and V. V. Savchenko. Comparison of C and Java performance in finite element computations. Computers and Structures, 81(24– 25):2401–2408, 2003. CO-DEN CMSTCJ. ISSN 0045-7949 (print), 1879-2243 (electronic).

Nolan:2004:DJ

[Nol04] Godfrey Nolan. Decompiling Java. Apress, Berkeley, CA, USA, 2004. ISBN 1-59059-265-4. xiii + 264 pp. LCCN QA76.73.J38 N65 2004.

Norman:2000:FEJ

[Nor00] Arthur C. Norman. Further evaluation of Java for symbolic computation. In Traverso [Tra00b], pages 258–

265. ISBN 1-58113-218-

2. LCCN QA76.95.I59 2000.

URL http://www.acm.org/
pubs/articles/proceedings/
issac/345542/p258-norman/
p258-norman.pdf;   http:
//www.acm.org/pubs/citations/
proceedings/issac/345542/
p258-norman/. ACM order

number 505000.

Narasimhan:2001:CBS

[NP01] Srivatsan Narasimhan and Santosh Pande. Compiler based scheduling of Java mobile agents. Lecture Notes in Computer Science, 2017:372– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2017/20170372.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2017/20170372.
pdf.

Noonan:2002:UTF

[NP02] Robert E. Noonan and Richard H. Prosl. Unit testing frameworks. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):232– 236, March 2002. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Niemeyer:2003:EPA

[NP03] Glenn Niemeyer and Jeremy Poteet. Extreme programming with Ant: building and deploying Java applications with JSP, EJB, XSLT, XDoclet, and JUnit. Howard W. Sams, Indianapolis, IN 46268, USA, 2003. ISBN 0-672-32562-4. ix + 444 pp. LCCN QA76.76.D47 N54 2003.

Noguera:2007:AEA

[NP07] Carlos Noguera and Renaud Pawlak. AVal: an extensible attribute-oriented programming validator for Java. Journal of Software Maintenance and Evolution: Research and Practice, 19(4): 253–275, July 2007. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Neary:2001:JJB

[NPRC01] Michael O. Neary, Alan Phipps, Steven Richman, and Peter Cappello. Javelin 2.0: Java-based parallel computing on the Internet. Lecture Notes in Computer Science, 1900:1231–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.
                                        com/link/service/series/
                                        0558/bibs/1900/19001231.
                                        htm;   http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/1900/19001231.
                                        pdf.

Nystrom:2006:JNIa

[NQM06] Nathaniel Nystrom, Xin Qi, and Andrew C. Myers. J&: nested intersection for scalable software composition. ACM SIGPLAN Notices, 41(10):21–36, October 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Null:2005:CIM

[NR05] Linda Null and Karishma Rao. CAMERA: introducing memory concepts via visualization. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):96–100, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Nanda:2006:ISM

[NR06] Mangala Gowri Nanda and S. Ramesh. Interprocedural slicing of multithreaded programs with applications to Java. ACM Transactions on Programming Languages and Systems, 28(6): 1088–1144, November 2006. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Neelakantam:2007:HAR

[NRS^{+}07] Naveen Neelakantam, Ravi Rajwar, Suresh Srinivas, Uma Srinivasan, and Craig Zilles. Hardware atomicity for reliable software speculation.

ACM SIGARCH Computer Architecture News, 35(2): 174–185, May 2007. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Natarajan:2000:PVD

[NRV00] Vijaykumar Natarajan, Stefan Reich, and Bhaskar Vasudevan. Programming with VisiBroker; A Developer’s Guide to VisiBroker for Java. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, second edition, 2000. ISBN 0-471-37682-5. xxviii + 543 pp. LCCN QA76.73.J38 N36 2000. US$50.

Negrino:2001:JWW

[NS01a] Tom Negrino and Dori Smith. JavaScript for the World Wide Web. Visual quickstart guide. Peachpit Press, Inc., 1085 Keith Avenue, Berkeley, CA 94708, USA, fourth edition, 2001. ISBN 0-201-73517-2. xiv + 430 pp. LCCN QA76.73.J39 N44 2001.

Ngo:2001:IJJ

[NS01b] Ton Ngo and Steve Smith. Implementing JNI in Java for Jalapeño. In USENIX Association [USE01c], page ?? ISBN 1-880446-11-1. LCCN QA76.73.J38 J42 2001. URL

http://www.usenix.org/publications/
library/proceedings/jvm01/
JVM_wips/S06.pdf. Spon-

sored by the USENIX Association.

Nickell:2003:TPJ

[NS03] E. Nickell and I. Smith. Three patterns in Java unit testing. Lecture Notes in Computer Science, 2675:170–179, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nakamura:2003:DJF

[NSI03] N. Nakamura, M. Shimono, and K. Ikeda. Development of a Java framework for rapid, low-cost business systems. Sei Technical Review, 55:50–54, 2003. CODEN ???? ISSN 1343-4349.

Nugent:2005:DDV

[NSS^{+}05] Gwen Nugent, Leen-Kiat Soh, Ashok Samal, Suzette Person, and Jeff Lang. Design, development, and validation of a learning object for CS1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37 (3):370, September 2005. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Nakajima:2001:BAE

[NT01] Shin Nakajima and Tetsuo Tamai. Behavioural analysis of the Enterprise JavaBeans^{T M} Component Architecture. Lecture Notes in Computer Science, 2057: 163–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.

springer-ny.com/link/service/
series/0558/bibs/2057/20570163.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2057/20570163.
pdf.

Narayanan:2002:JM

[NW02a] Vijaykrishnan Narayanan and Mario I. Wolczko. Java microarchitectures. Number SECS 679 in The Kluwer international series in engineering and computer science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2002. ISBN 1-4020-7034-9. xiv + 252 pp. LCCN QA76.73.J38 J376 2002.

Newsome:2002:PCD

[NW02b] Matt Newsome and Des Watson. Proxy compilation of dynamically loaded Java classes with MoJo. ACM SIGPLAN Notices, 37(7):204–212, July 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nevison:2003:TOE

[NW03] C. Nevison and B. Wells. Teaching objects early and design patterns in Java using case studies. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):94–98, 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Naftalin:2006:JGC

[NW06] Maurice Naftalin and Philip Wadler. Java Generics and Collections. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2006. ISBN 0-596-52775-6. 294 (est.) pp. LCCN ???? US$34.99.

Naftalin:2007:JGC

[NW07] Maurice Naftalin and Philip Wadler. Java Generics and Collections. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2007. ISBN 0-596-52775-6. xvii + 273 pp. LCCN QA76.73.J38 N325 2007. US$34.99.

URL http://www.loc.gov/
catdir/enhancements/fy0715/
2006287224-d.html;   http:
//www.loc.gov/catdir/toc/
fy0705/2006287224.html;
http://www.oreilly.com/
catalog/9780596527754.

Nyberg:2002:WSW

[Nyb02] Greg Nyberg. WebLogic 6.1 Server Workbook for Enterprise JavaBeans. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, third edition, 2002. ISBN 0-596-00417-6. 256 (est.) pp. LCCN ???? US$24.95. URL http://

www.oreilly.com/catalog/
weblogicworkbook.

Noble:2001:SCJ

[NZ01] Michael S. Noble and Stoyanka Zlateva. Scientific computation with JavaSpaces. Lecture Notes in Computer Science, 2110:657–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2110/21100657.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2110/21100657.
pdf.

NiewiadomskaSzynkiewicz:2003:AJB

[NZM03] E. NiewiadomskaSzynkiewicz, M. Zmuda, and K. Malinowski. Application of a Javabased framework to parallel simulation of large-scale systems. International Journal of Applied Mathematics and Computer Science, 13(4): 537–548, 2003. CODEN ???? ISSN 1641-876X.

Oaks:2001:JS

[Oak01] Scott Oaks. Java security. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2001. ISBN 0-596-00157-6. xvi + 599 pp. LCCN QA76.73.J38 O247 2001. Covers JAAS and JSEE; Writing and deploying secure applications; Covers Java 1.1, Java 2, and JCE 1.2.1.

OBrien:2005:JCC

[O’B05] Timothy M. O’Brien. Jakarta commons cookbook. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00706-X. xix + 377 pp. LCCN QA76.76.A65 O246 2005. URL http://

www.oreilly.com/catalog/
9780596007065.

OBrien:2005:BBW

[OBr05] L. OBrien. Best of both worlds? morphing the productivity of ASP.NET into a deployment environment based on Java, Visual Main-Win for J2EE might serve as the solution for your dualplatform development needs. Software Development, 13(3): 32–35, 2005. CODEN ???? ISSN 1070-8588.

Ochem:2009:GIA

[Och09a] Quentin Ochem. Gem #55:

introduction to Ada /Java interfacing. ACM SIGADA Ada Letters, 29(2):43–45, August 2009. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Ochem:2009:GCA

[Och09b] Quentin Ochem. Gem #56:

creating Ada to Java calls using GNAT-AJIS. ACM SIGADA Ada Letters, 29(2): 46–49, August 2009. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Ochem:2009:GAJa

[Och09c] Quentin Ochem. Gem #57:

Ada /Java cross dispatching. ACM SIGADA Ada Letters, 29(2):50–52, August 2009. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Ochem:2009:GAJb

[Och09d] Quentin Ochem. Gem #58: Ada /Java exception handling. ACM SIGADA Ada Letters, 29(2):53–55, August 2009. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Ochem:2009:MLP

[Och09e]  Quentin Ochem.  Multi-

language programming with Ada. ACM SIGADA Ada Letters, 29(3):19–20, December 2009. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Oestreicher:2001:ECJ

[Oes01]  Dan Oestreicher.  Expe-

rience with a commercial Java implementation of group communication using reliable multicast. Operating Systems Review, 35(4):21–31, October 2001. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Offutt:2000:STA

[Off00] Jeff Offutt. Software testing and analysis of objectoriented software. ACM SIGSOFT Software Engineering Notes, 25(1):68, January 2000. CODEN SFENDP.

ISSN 0163-5948 (print), 1943-5843 (electronic).

Oechsle:2005:DDA

[OG05] Rainer Oechsle and Tim Gottwald. DisASTer (distributed algorithms simulation terrain): a platform for the implementation of distributed algorithms. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):44– 48, September 2005. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Oliver:2001:SEE

[OGA^{+}01] José Oliver, Jordi Guitart, Eduard Ayguadé, Nacho Navarro, and Jordi Torres. Strategies for the efficient exploitation of loop-level parallelism in Java. Concurrency and Computation: Practice and Experience, 13(8–9):663– 680, July/August 2001. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (elec-

tronic). URL http://www3.
interscience.wiley.com/
cgi-bin/abstract/84503218/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=84503218&PLACEBO=IE.
pdf.

Ogasawara:2009:NAM

[Oga09] Takeshi Ogasawara. NUMAaware memory manager with dominant-thread-based copying GC. ACM SIGPLAN Notices, 44(10):377–390, October 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Oaks:2002:JN

[OGT02] Scott Oaks, Li Gong, and Bernard Traversat. JXTA in a Nutshell. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2002. ISBN 0-596-00236-X. xii + 401 pp. LCCN TK5105.525 .O255

2002.  US$34.95.  URL

                                        http://www.oreilly.com/
                                        catalog/jxtaian.

ONeill:2005:IAS

[OHL^{+}05] Ian O’Neill, Philip Hanna, Xingkun Liu, Des Greer, and Michael McTear. Implementing advanced spoken dialogue management in Java. Science of Computer Programming, 54(1):99–124, January 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Oi:2005:DLV

[Oi05] Hitoshi Oi. On the design of the local variable cache in a hardware translation-based Java Virtual Machine. ACM SIGPLAN Notices, 40(7):87– 94, July 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Oi:2006:IFH

[Oi06] Hitoshi Oi. Instruction folding in a hardware-translation based Java Virtual Machine. In ACM [ACM06], pages 139– 146. ISBN 1-59593-302-6. LCCN ???? ACM order number 104060.

Oi:2008:LVA

[Oi08] Hitoshi Oi. Local variable access behavior of a hardwaretranslation based Java virtual machine. The Journal of systems and software, 81(11): 2059–2068, November 2008. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Oiwa:2009:IMS

[Oiw09] Yutaka Oiwa. Implementation of the memory-safe full ANSI-C compiler. ACM SIG-PLAN Notices, 44(6):259– 269, June 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Overbey:2009:RLR

[OJ09] Jeffrey L. Overbey and Ralph E. Johnson. Regrowing a language: refactoring tools allow programming languages to evolve. ACM SIG-PLAN Notices, 44(10):493– 502, October 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Odekirk:2000:TSC

[OJJ00] Elizabeth Odekirk, Dominic Jones, and Peter Jensen. Three semesters of CSO using Java: assignments and experiences. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):144–147, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Olsson:2004:JPL

[OK04] R. A. Olsson and A. W. Keen. The JR programming language: Concurrent programming in an extended Java. Kluwer International Series in Engineering and Computer Science, 774:??, 2004. CO-DEN ???? ISSN 0893-3405.

Onodera:2004:LRJ

[OKK04] T. Onodera, K. Kawachiya, and A. Koseki. Lock reservation for Java reconsidered. Lecture Notes in Computer Science, 3086:560–584, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ogasawara:2001:SEH

[OKN01] Takeshi Ogasawara, Hideaki Komatsu, and Toshio Nakatani. A study of exception handling and its dynamic optimization in Java. ACM SIG-PLAN Notices, 36(11):83–95, November 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Proceedings of the 2001 ACM SIGPLAN Conference on Object Oriented Programming, Systems, Languages and Applications (OOPSLA’01).

Ogata:2002:BFOa

[OKN02a] Kazunori Ogata, Hideaki Komatsu, and Toshio Nakatani. Bytecode fetch optimization for a Java interpreter. ACM SIGARCH Computer Architecture News, 30(5):58– 67, December 2002. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Ogata:2002:BFOb

[OKN02b] Kazunori Ogata, Hideaki Komatsu, and Toshio Nakatani. Bytecode fetch optimization for a Java interpreter. Operating Systems Review, 36(5): 58–67, December 2002. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ogata:2002:BFOc

[OKN02c] Kazunori Ogata, Hideaki Komatsu, and Toshio Nakatani. Bytecode fetch optimization for a Java interpreter. ACM SIGPLAN Notices, 37 (10):58–67, October 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Also published as Computer Architecture News (ACM SIGARCH) 37(5) 58– 67 (2002).

Ogasawara:2004:OPO

[OKN04] Takeshi Ogasawara, Hideaki Komatsu, and Toshio Nakatani. Optimizing precision overhead for x86 processors. Software—Practice and Experience, 34(9):875–893, July 25, 2004. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ogasawara:2006:EED

[OKN06] Takeshi Ogasawara, Hideaki Komatsu, and Toshio Nakatani. EDO: Exception-Directed Optimization in Java. ACM Transactions on Programming Languages and Systems, 28(1):70–105, January 2006. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Orleans:2001:DDA

[OL01] Doug Orleans and Karl Lieberherr. DJ: Dynamic adaptive programming in Java. Lecture Notes in Computer Science, 2192:73–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.
                                        com/link/service/series/
                                        0558/bibs/2192/21920073.
                                        htm;   http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2192/21920073.
                                        pdf.

Olson:2001:BJP

[Ols01]  Quentin Olson.  Brewing

Java at the point of sale. Embedded Linux Journal, 6: 28–31, November/December 2001. CODEN ???? ISSN 1534-083X. URL http:

//embedded.linuxjournal.
com/magazine/issue06/;
http://www.linuxdevices.
com/articles/AT8821374910.
html.

Olsen:2007:AJ

[Ols07] Steven Douglas Olsen. Ajax on Java. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2007. ISBN 0-596-10187-2. xiv + 211 pp. LCCN QA76.73.J38 O58 2007eb; QA76.73.J38. URL http://

www.oreilly.com/catalog/
9780596101879.

Offutt:2004:EMS

[OMK04] Jeff Offutt, Yu-Seung Ma, and Yong-Rae Kwon. An experimental mutation system for Java. ACM SIGSOFT Software Engineering Notes, 29(5):1–4, September 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Omma:2001:BRS

[Omm01] Mikael Omma. Book review: Server Based Java Programming. Parallel and Distributed Computing Practices, 4(4):??, December 2001. CO-DEN ???? ISSN 1097-2803.

Omondi:2003:DIJ

[Omo03] A. R. Omondi. Design and implementation of Java processors. Lecture Notes in Computer Science, 2823:86– 96, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oliva:2008:ALF

[ONRV08] E. Oliva, A. Natali, A. Ricci, and M. Viroli. An adaptation logic framework for Java-based component systems. J.UCS: Journal of Universal Computer Science, 14(13):2158–??, ???? 2008. CODEN ???? ISSN 0948-6968. URL http:

                                        //www.jucs.org/jucs_14_
                                        13/an_adaptation_logic_
                                        framework.

Ogata:2006:RCIa

[OOK^{+}06] Kazunori Ogata, Tamiya Onodera, Kiyokuni Kawachiya, Hideaki Komatsu, and Toshio Nakatani. Replay compilation: improving debuggability of a just-in-time compiler. ACM SIGPLAN Notices, 41(10):241–252, October 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ozaki:2007:MOV

[OOM^{+}07] K. Ozaki, T. Ogita, S. Miyajima, S. Oishi, and S. M. Rump. A method of obtaining verified solutions for linear systems suited for Java. Journal of Computational and Applied Mathematics, 199(2): 337–344, February 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042705007697.

Ohira:2005:ACP

[OOOiM05] Masao Ohira, Naoki Ohsugi, Tetsuya Ohoka, and Ken ichi Matsumoto. Accelerating cross-project knowledge collaboration using collaborative filtering and social networks. ACM SIGSOFT Software Engineering Notes, 30(4):1–5, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Owens:2002:JIW

[OPS^{+}02] Barbara Boucher Owens, Rich Pattis, Chris Stephenson, Jack Rehder, and Dean Sanders. JAVA IDEs: why and how we use what we do. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34 (1):213–214, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Oechsle:2002:JAP

[OS02] Rainer Oechsle and Thomas Schmitt. JAVAVIS: Automatic program visualization with object and sequence diagrams using the Java debug interface (JDI). Lecture Notes in Computer Science, 2269:176–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2269/22690176.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2269/22690176.
pdf.

Orso:2004:SRT

[OSH04] Alessandro Orso, Nanjuan Shi, and Mary Jean Harrold. Scaling regression testing to large software systems. ACM SIGSOFT Software Engineering Notes, 29(6):241– 251, November 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ogawa:2000:OOE

[OSM^{+}00] Hirotaka Ogawa, Kouya Shimura, Satoshi Matsuoka, Fuyuhiko Maruyama, Yukihiko Sohda, and Yasunori Kimura. OpenJIT: An openended, reflective JIT compiler framework for Java. Lecture Notes in Computer Science, 1850:362–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/

0558/bibs/1850/18500362.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/1850/18500362.
pdf.

Ourosoff:2002:PTJ

[Our02] Nick Ourosoff. Technical opinion: Primitive types in Java considered harmful. Communications of the ACM, 45(8):105–106, August 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Oaks:2000:JDQ

[OW00] Scott Oaks and Henry Wong.

Jini: a desktop quick reference. In a nutshell. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2000. ISBN 1-56592-759-1 (paperback). xiii + 400 pp. LCCN QA76.9.C55 O25 2000. es.

Oaks:2004:JT

[OW04] Scott Oaks and Henry Wong.

Java threads. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, third edition, 2004. ISBN 0-596-00782-5. xvi + 340 pp. LCCN QA76.73.J38 O25 2004; QA76.73.J38 O25 2004eb; QA76.73.J38. URL

http://www.oreilly.com/
catalog/9780596007829.

Owen:2004:JJE

[OWR04] T. Owen, I. Wakeman, and J. Rathke. JPolicy: a Java extension for dynamic access control. Technical report, University of Sussex Computer Science Department, 2004.

Pedrick:1998:PVC

[P^{+}98] Doug Pedrick et al. Programming with VisiBroker (CORBA & JDBC). John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, February 1998. ISBN 0-471-23901-1. xvi + 435 pp. LCCN QA76.73.J38 P79 1998. US$45. URL http://

www.wiley.com/compbooks/
catalog/23901-1.htm.

Palmer:2002:JEH

[Pal02] Grant Palmer. Java event handling. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2002. ISBN 0-13-041802-1. xix + 590 pp. LCCN QA76.73.J38 P33 2002. Accompanying CD-ROM includes all of the examples described in the book. It contains not only the Java source code but also the associated HTML and JPEG files.

Panda:2004:WDA

[Pan04] D. Panda. What the database administrator needs to know about Java. Elektron: Journal of the South African Institute of Electrical Engineers, 21(7):74–77, 2004. CODEN ???? ISSN 1026-9193.

Pandey:2009:EWR

[Pan09] R. K. Pandey. Exploiting web resources for teaching/ learning best software design tips. ACM SIGSOFT Software Engineering Notes, 34 (6):1–7, November 2009. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Paprzycki:2000:BRJ

[Pap00] M. Paprzycki. Book review: Java distributed computing is solid but not flawless. IEEE Concurrency, 8 (2):86, April/June 2000. CO-DEN IECMFX. ISSN 1092-3063 (print), 1558-0849 (electronic).

Papanikolaou:2005:BRBb

[Pap05] Nikolaos Papanikolaou. Book review: Classical and Quantum Computing with C++ and Java Simulations, by Yorick Hardy and Willi-Hans Steeb, Birkhäuser Verlag, 2001. ACM SIGACT News, 36(3):5–9, September 2005. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). URL http://

doi.acm.org/10.1145/1086649.

1086654. See [HS01].

Parson:2000:UJR

[Par00] Dale E. Parson. Using Java reflection to automate extension language parsing. ACM SIGPLAN Notices, 35(1):67– 80, January 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pardi:2004:PCD

[Par04a]  W. Pardi Jr.  Program-

ming concurrent and distributed algorithms in Java. IEEE Distributed Systems Online, 5(11):5, November 2004. CODEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic). URL

http://ieeexplore.ieee.
org/iel5/8968/30052/01377092.
pdf?isnumber=30052&prod=
JNL&arnumber=1377092&arSt=
+5&ared=+5&arAuthor=+Pardi%
2C+W.%2C+Jr.;   http://
ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=30052&
arnumber=1377092&count=
5&index=4.

Parlante:2004:NAG

[Par04b] N. Parlante. Nifty assignments: Graphics and Java. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(2):26–27, 2004. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Parlante:2004:GJ

[Par04c]  Nick Parlante.  Graphics

and Java. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(2):26–27, June 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp:

//ftp.math.utah.edu/pub/
mirrors/ftp.ira.uka.de/
bibliography/Misc/DBLP/
2004.bib.

Parlante:2004:N

[Par04d]  Nick Parlante.  Niftiness.

SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36 (4):26–27, December 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp:

//ftp.math.utah.edu/pub/
mirrors/ftp.ira.uka.de/
bibliography/Misc/DBLP/
2004.bib.

Parsons:2005:JAM

[Par05] D. Parsons. Java architectures for mobilised enterprise systems. In Proceedings of the Annual Hawaii International Conference on System Sciences, volume CONF38, page 298. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2005. CODEN ???? ISSN 1060-3425.

Pascarello:2004:JYV

[Pas04] Eric Pascarello. JavaScript:

your visual blueprint for building dynamic Web pages. Read less, learn more. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, second edition, 2004. ISBN 0-7645-7497-3 (paperback). xiii + 321 pp. LCCN TK5105.888 .P375

2004. URL http://www.loc.
gov/catdir/enhancements/
fy0620/2004112337-b.html;
 http://www.loc.gov/catdir/
enhancements/fy0620/2004112337-
d.html;   http://www.loc.
gov/catdir/toc/wiley051/
2004112337.html.

Paulson:2001:NBRb

[Pau01] Linda Dailey Paulson. News briefs: Researchers work on improving memory: Tracking your E-mail: Handheld-to-handheld: Fighting over Java; future searches; location, location, location; Michigan goes to cybercourt. Computer, 34(7):19–21, July 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//dlib.computer.org/co/
books/co2001/pdf/r7019.
pdf.

Paulson:2003:NBR

[Pau03] Linda Dailey Paulson. News briefs: Researchers automate the digital animation process; new compression technology would improve online video quality; Sun tries to guide Mobile Java; new search engines keep humming along. Computer, 36 (4):24–26, April 2003. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/co/
2003/04/r4024.htm;   http:
//csdl.computer.org/dl/
mags/co/2003/04/r4024.pdf.

Pausch:2008:ADM

[Pau08] Randy Pausch. Alice: a dying man’s passion. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):1, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Payne:2004:PJB

[Pay04] W. Payne. Periscope Java is becoming the platform of choice for developing a new generation of medical devices. British Journal of Healthcare Computing and Information Management, 21(7):14, 2004. CODEN ????

Peterson:2006:OCI

[PB06] Laurence I. Peterson and Dale Benham. Overview of the cybertech-ITEST project: an initiative to attract and prepare under-represented students for tomorrow’s careers in the computing sciences. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38 (1):453–455, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Parkinson:2008:SLA

[PB08] Matthew J. Parkinson and Gavin M. Bierman. Separation logic, abstraction and inheritance. ACM SIGPLAN Notices, 43(1):75–86, January 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Philippsen:2001:JHP

[PBG^{+}01] Michael Philippsen, Ronald F. Boisvert, Valdimir S. Getov, Roldan Pozo, José Moreira, Dennis Gannon, and Geoffrey C. Fox. JavaGrande — high performance computing with Java. Lecture Notes in Computer Science, 1947:20– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/1947/19470020.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/1947/19470020.
pdf.

Pugla:2003:JPD

[PC03] S. Pugla and S. Chawla. A JAVA program for designing synchronously tuned extended cavity diode laser in Littrow configuration. Indian Journal of Pure and Applied Physics, 41(11):879–883, 2003. CODEN ???? ISSN 0019-5596.

Parker:2004:PAC

[PC04] D. Parker and D. Cleary. A P2P approach to ClassLoading in Java. Lecture Notes in Computer Science, 2872:144– 149, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pullen:2008:DAL

[PC08] J. Mark Pullen and Jim X.

Chen. Distributed application launching for high quality graphics in synchronous distance education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3): 204–208, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Pidd:2000:UJD

[PCC00] M. Pidd, R. A. Cassel, and Correspondence. Using Java to develop discrete event simulations. OR: the journal of the Operational Research Society, 51(4):405–412, April 1, 2000. CODEN OPRQAK. ISSN 0160-5682 (print), 1476-9360 (electronic).

Pollet:2001:DSD

[PCC01] Isabelle Pollet, Baudouin Le Charlier, and Agostino Cortesi. Distinctness and sharing domains for static analysis of Java programs. Lecture Notes in Computer Science, 2072:77–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2072/20720077.
htm;   http://link.springer-
ny.com/link/service/series/

                                        0558/papers/2072/20720077.
                                        pdf.

Pacios:2002:JBG

[PDCL02] L. Pacios, A. DeLaPena, R. Carrasco, and F. Lapayese. Java-based gas inlet control system for the Stellarator TJ-II. VTT Symposium, 220 (??):243, 2002. CODEN ???? ISSN 0357-9387.

Pasareanu:2001:FFC

[PDV01] Corina S. Pasareanu, Matthew B. Dwyer, and Willem Visser. Finding feasible counterexamples when model checking abstracted Java programs. Lecture Notes in Computer Science, 2031:284–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.
                                        com/link/service/series/
                                        0558/bibs/2031/20310284.
                                        htm;   http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2031/20310284.
                                        pdf.

Paul:2006:CJN

[PE06] Nathanael Paul and David Evans. Comparing Java and .NET security: Lessons learned and missed. Computers and Security, 25(5): 338–350, July 2006. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0167404806000290.

Pekowsky:2000:JP

[Pek00] Larne Pekowsky. JavaServer Pages. Addison-Wesley, Reading, MA, USA, 2000. ISBN 0-201-70421-8. 300 (est.) pp. LCCN K5105.8885.J38 P45 2000. US$39.95. URL

http://cseng.aw.com/bookdetail.
qry?ISBN=0-201-70421-8.

Includes CD-ROM.

Pellizzari:2003:CPJ

[Pel03] L. Pellizzari. Cresce la presenza di Java nelle applicazioni embedded. Elettronica Oggi, 322:56–57, 2003. CO-DEN ELOGDA. ISSN 0391-6391.

Perry:2001:OND

[Per01] Tekla S. Perry. Operating in neutral [digital music legislation]. IEEE Spectrum, 38(10): 52–57, October 2001. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Perry:2002:JME

[Per02] J. Steven (James Steven)

Perry. Java Management Extensions. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2002. ISBN 0-596-00245-9. x + 300 pp. LCCN QA76.73.J38 P472 2002. US$34.95. URL http://

safari.oreilly.com/0596002459;
 http://www.oreilly.com/
catalog/javamngext.

Perry:2004:JSJ

[Per04] Bruce W. Perry. Java servlet and JSP cookbook. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2004. ISBN 0-596-00572-5. xix + 723 pp. LCCN QA76.73.J38 P469 2004. URL http://

                                        www.oreilly.com/catalog/
                                        9780596005726.

Perry:2006:AH

[Per06] Bruce W. Perry. Ajax hacks. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2006. ISBN 0-596-10169-4. xxi + 414 pp. LCCN QA76.73.J39 P47 2006eb; QA76.73.J39. URL

                                        http://www.oreilly.com/
                                        catalog/9780596101695.

Petitpierre:2003:JTC

[Pet03] C. Petitpierre. Java threads can be very useful building blocks. Lecture Notes in Computer Science, 2604:204, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Petullo:2005:DGA

[Pet05] Mike Petullo. Developing GNOME applications with Java. Linux Journal, 2005 (135):??, July 2005. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Petro:2006:RMJ

[Pet06] Andrew Petro. Review of More Java Gems by Dwight Deugo, Cambridge University Press, 2000, ISBN: 0-521-77477-2. ACM SIGSOFT Software Engineering Notes, 31(3):42, May 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Pew:2000:WPJ

[Pew00] John A. Pew. Web Publishing with Java: The Complete Video Course. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2000. ISBN 0-13-085842-0. ???? pp. LCCN ???? US$50.

Plante:2005:SJI

[PFJ05] I. L. Plante, A. FilaliMouhim, and J. P. JayGerin. Simul-Rad: a Java interface for a Monte-Carlo simulation code to visualize in 3D the early stages of water radiolysis. Radiation Physics and Chemistry, 72(2–3):173–180, 2005. CODEN ???? ISSN 0969-806X.

Prinz:2005:JBD

[PFS05] M. Prinz, G. Fischer, and E. Schuster. The JAVA-based DICOM query interface DicoSE. International Journal of Medical Informatics, 74(2– 4):325–333, 2005. CODEN ???? ISSN 1386-5056.

Philippsen:2000:CNJ

[PG00] Michael Philippsen and Edwin Günthner. Complex numbers for Java. Concurrency: Practice and Experience, 12 (6):477–491, May 2000. CO-DEN CPEXEI. ISSN 1040-

3108. URL http://www3.
interscience.wiley.com/
cgi-bin/abstract/72515730/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=72515730&PLACEBO=IE.
pdf.

Pinilla:2003:UJT

[PG03a] R. Pinilla and M. Gil. ULT: a Java threads model for platform independent execution. Operating Systems Review, 37(4):48–62, 2003. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Pinilla:2003:JPI

[PG03b] Ruben Pinilla and Marisa Gil. JVM: platform independent vs. performance dependent. Operating Systems Review, 37 (2):44–56, April 2003. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

PerezLopez:2005:JBL

[PGM^{+}05] S. A. Perez Lopez, M. J. Gonzalez Brana, J. Marcos Acevedo, M. D. Valdes, and E. Mandado. Java-based learning of algorithms for VLSI physical design automation. International Journal of Engineering Education, 21 (2):306–317, 2005. CODEN IEEDEF. ISSN 0949-149X.

Pandey:2000:PFG

[PH00a] Raju Pandey and Brant Hashii. Providing fine-grained access control for Java programs via binary editing. Concurrency: Practice and Experience, 12(14):1405– 1430, December 10, 2000. CODEN CPEXEI. ISSN 1040-3108. URL http://

www3.interscience.wiley.
com/cgi-bin/abstract/76500721/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=76500721&PLACEBO=IE.
pdf.

Perelman-Hall:2000:JQ

[PH00b] David Perelman-Hall. Java Q&A: How do I build a find? Dr. Dobb’s Journal of Software Tools, 25(7): 119–120, 122, 124, 126, July 2000. CODEN DDJOEB. ISSN 1044-789X. URL

http://www.ddj.com/ftp/
2000/2000_07/jqa700.txt;
 http://www.ddj.com/ftp/
2000/2000_07/jqa700.zip.

Philippsen:2000:LOJ

[PH00c] Michael Philippsen and Bernhard Haumacher. Locality optimization in JavaParty by means of static type analysis. Concurrency: Practice and Experience, 12(8): 613–628, July 2000. CO-DEN CPEXEI. ISSN 1040-

                                        3108. URL http://www3.
                                        interscience.wiley.com/
                                        cgi-bin/abstract/72516226/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=72516226&PLACEBO=IE.
                                        pdf.

Pike:2002:BTA

[PH02] Geoff Pike and Paul N. Hilfinger. Better tiling and array contraction for compiling scientific programs. In IEEE [IEE02a], page ?? ISBN 0-7695-1524-X. LCCN

                                        ???? URL http://www.sc-
                                        2002.org/paperpdfs/pap.
                                        pap260.pdf.

Paterson:2003:TJU

[PH03] J. H. Paterson and J. Haddow. Teaching Java: Using an object-oriented database and the BlueJ IDE. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):273, 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Paterson:2004:AOP

[PH04] J. H. Paterson and J. Haddow. Approaches to object persistence in Java projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):256, 2004. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Paterson:2005:UBI

[PHBM05] James H. Paterson, John Haddow, Miriam Birch, and Alex Monaghan. Using the BlueJ IDE in a data structures course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3): 349, September 2005. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Parrish:2001:IAV

[PHM^{+}01] Allen Parrish, Joe Hollingsworth, Peter Maurer, Benjamin Shults, and Bruce Weide. Identifying an appropriate view of software components for undergraduate education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33 (1):394–395, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Philippsen:2000:MES

[PHN00] Michael Philippsen, Bernhard Haumacher, and Christian Nester. More efficient serialization and RMI for Java. Concurrency: Practice and Experience, 12(7): 495–518, May 2000. CO-DEN CPEXEI. ISSN 1040-

3108. URL http://www3.
interscience.wiley.com/
cgi-bin/abstract/72516213/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?

                                        ID=72516213&PLACEBO=IE.
                                        pdf.

Pizlo:2007:HRT

[PHV07] Filip Pizlo, Antony L. Hosking, and Jan Vitek. Hierarchical real-time garbage collection. ACM SIGPLAN Notices, 42(7):123–133, July 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pilone:2004:EVE

[Pil04] M. Pilone. The Eclipse Visual Editor for Java. Dr. Dobb’s Journal of Software Tools, 29 (10):67–71, 2004. CODEN DDJOEB. ISSN 1044-789X.

Pilgrim:2005:GH

[Pil05]  Mark Pilgrim.  Grease-

monkey hacks. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2005. ISBN 0-596-10165-1. 352 pp. LCCN TK5105.882eb; TK5105.882. URL http://

                                        www.oreilly.com/catalog/
                                        9780596101657.

Pipka:2003:TDW

[Pip03] J. U. Pipka. Test-driven Web application development in Java. Lecture Notes in Computer Science, 2591:378– 393, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Piroumian:2002:WJP

[Pir02] Vartan Piroumian. Wireless J2ME platform programming. Sun BluePrints Program. Sun Microsystems Press, Palo Alto, CA, USA, 2002. ISBN 0-13-044914-8. xxv + 374 pp. LCCN QA76.73.J38 P575 2002.

URL http://www.sun.com/
books/catalog/piroumian/.

Pillay:2005:ISC

[PJ05] Nelishia Pillay and Vikash R.

Jugoo. An investigation into student characteristics affecting novice programming performance. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(4):107–110, December 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp:

//ftp.math.utah.edu/pub/
mirrors/ftp.ira.uka.de/
bibliography/Misc/DBLP/
2005.bib.

Proulx:2009:UTJ

[PJ09] Viera K. Proulx and Weston Jossey. Unit testing in Java. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):349, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Pree:2000:FSL

[PK00] Wolfgang Pree and Kai Koskimies. Framelets — small and loosely coupled frameworks. ACM Computing Surveys, 32(1es):6:1– 6:??, March 2000. CO-DEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/surveys/2000-32-
1/p6-pree/.

Pelrine:2001:MED

[PKC01] Joseph Pelrine, Alan Knight, and Adrian Cho. Mastering ENVY/Developer, volume 22 of Advances in object technology. Cambridge University Press, Cambridge, UK, 2001. ISBN 0-521-66650-3. LCCN QA76.64 .P46 2001. US$39.95.

Paal:2002:CDC

[PKF02] Stefan Paal, Reiner Kammüller, and Bernd Freisleben. Customizable deployment, composition, and hosting of distributed Java applications. Lecture Notes in Computer Science, 2519:845–865, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/
service/series/0558/bibs/
2519/25190845.htm;   http:
//link.springer.de/link/
service/series/0558/papers/
2519/25190845.pdf.

Paal:2003:JCD

[PKF03] S. Paal, R. Kammuller, and B. Freisleben. Java class deployment using class collections. Lecture Notes in Computer Science, 2591:135– 151, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pancake:2001:HPJ

[PL01a] Cherri Pancake and Christian Lengauer. High-performance Java. Communications of the ACM, 44(10):98–101, October 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Park:2001:RRJ

[PL01b] Jung Gyu Park and Arthur H.

Lee. Removing reflection from Java programs using partial evaluation. Lecture Notes in Computer Science, 2192:274– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2192/21920274.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2192/21920274.
pdf.

Payne:2003:PJT

[PL03] Damon Payne and Ed Lyons.

Professional Java tools: Real World Ant, JUnit, CVS, Cactus, Bugzilla, Maven, Jmeter and XDoclet. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 2003. ISBN 0-7645-4389-X, 1-86100-799-X. 450 (est.) pp. LCCN ????

Pollet:2005:TCS

[PL05] I. Pollet and B. LeCharlier. Towards a complete static analyser for Java: an abstract interpretation framework and its implementation. Electronic Notes in Theoretical Computer Science, 131: 85–98, 2005. CODEN ???? ISSN 1571-0661.

Plauger:2000:SCC

[Pla00] P. J. Plauger. Standard C/C++: Java standard time. C/C++ Users Journal, 18(1): ??, January 2000. CODEN CCUJEX. ISSN 1075-2838.

Pleumann:2002:MP

[Ple02] Jörg Pleumann. The Mail4Me Project. Dr. Dobb’s Journal of Software Tools, 27(6): 36, 38, 40, 42, 44, June 2002. CODEN DDJOEB. ISSN 1044-789X. URL http:

                                        //www.ddj.com/ftp/2002/
                                        2002_06/mail4me.txt.

Pohl:2000:JDE

[PM00] Ira Pohl and Charlie McDowell. Java by dissection: the essentials of Java programming. Addison-Wesley, Reading, MA, USA, 2000. ISBN 0-201-61248-8. xv + 509 pp. LCCN QA76.73.J38 P66 2000.

Pitt:2001:JRR

[PM01a] Esmond Pitt and Kathy Mc-Niff. Java.RMI: the remote method invocation guide. Pearson Education, Harlow, England, 2001. ISBN 0-201-70043-3. xxxi + 284 pp. LCCN QA76.73.J38 P58 2001.

Pohl:2001:JDU

[PM01b] Ira Pohl and Charlie McDowell. Java by Dissection: Update with C Primer. Addison-Wesley, Reading, MA, USA, 2001. ISBN 0-201-72596-7. ???? pp. LCCN ???? US$68.

Potanin:2006:GOGa

[PNCB06] Alex Potanin, James Noble, Dave Clarke, and Robert Biddle. Generic ownership for generic Java. ACM SIG-PLAN Notices, 41(10):311– 324, October 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pistoia:2004:EJS

[PNKN04] Marco Pistoia, Nataraj Nagaratnam, Larry Koved, and Anthony Nadalin. Enterprise Java Security: building secure J2EE applications. Addison-Wesley, Reading, MA, USA, 2004. ISBN 0-321-11889-8. xxiii + 581 pp. LCCN QA76.73.J3E58 2004.

Pollock:2001:JBG

[Pol01] John Pollock. JavaScript: a beginner’s guide. Osborne/McGraw-Hill, Berkeley, CA, USA, 2001. ISBN 0-07-213140-

3. xx + 603 pp. LCCN QA76.73.J39 P65 2001.

Pont:2003:CCL

[Pon03] Mick Pont. Calling C library routines from Java. Dr. Dobb’s Journal of Software Tools, 28(7):28, 32, 34– 36, 38, July 2003. CODEN DDJOEB. ISSN 1044-789X.

                                        URL http://www.ddj.com/
                                        ftp/2003/2003_07/c_java.
                                        txt;   http://www.ddj.com/
                                        ftp/2003/2003_07/c_java.
                                        zip.

Potratz:2004:PCB

[Pot04]  E. Potratz.  A practical

comparison between Java and Ada in implementing a real-time embedded system. ACM SIGADA Ada Letters, 24(1):71–83, 2004. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Potter:2008:CJC

[Pot08]  Bruce Potter.  Control-

ling JavaScript with Caja. Network Security, 2008(1): 7–8, January 2008. CO-DEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1353485808700060.

Powers:2007:LJ

[Pow07] Shelley Powers. Learning JavaScript. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2007. ISBN 0-596-52746-2. xiv + 335 pp. LCCN QA76.73.J39 P7 2007eb; QA76.73.J39. URL

http://www.oreilly.com/
catalog/9780596527464.

Park:2002:SJP

[PP02a] J. G. Park and M. S. Park.

Specializing Java programs in a distributed environment. Journal of Information Science and Engineering, 18(5): 787–802, 2002. CODEN JI-NEEY. ISSN 1016-2364.

Park:2002:ASJ

[PP02b] Jung Gyu Park and Myong-Soon Park. Automatic specialization of Java programs in the distributed environment. Lecture Notes in Computer Science, 2344:757–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/
service/series/0558/bibs/
2344/23440757.htm;   http:
//link.springer.de/link/
service/series/0558/papers/
2344/23440757.pdf.

Prodan:2002:CJC

[PP02c] A. Prodan and R. Prodan. A collection of Java class libraries for stochastic modeling and simulation. Lecture Notes in Computer Science, 2329:1040–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/

                                        0558/bibs/2329/23291040.
                                        htm;   http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2329/23291040.
                                        pdf.

Parikh:2003:JMW

[PPJ03] A. Parikh, R. Pradhan, and S. Jiang. J2EE meets Web services J2EE was tearing up the charts when Web services appeared on the scene, and the Java community has reacted quickly. Application Development Trends, 10(10): 45–49, 2003. CODEN ???? ISSN 1073-9564.

Pominville:2001:FOJ

[PQVR^{+}01] Patrice Pominville, Feng Qian, Raja Vallée-Rai, Laurie Hendren, and Clark Verbrugge. A framework for optimizing Java using attributes. Lecture Notes in Computer Science, 2027:334– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.
                                        com/link/service/series/
                                        0558/bibs/2027/20270334.
                                        htm;   http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2027/20270334.
                                        pdf.

Pedroni:2002:JE

[PR02] Samuele Pedroni and Noel Rappin. Jython Essentials. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2002.

ISBN 0-596-00247-5. xx + 277 pp. LCCN QA76.73.J38 P43 2002. US$24.95. URL

http://www.oreilly.com/
catalog/jythoness.

Pegueroles:2003:ESM

[PR03] J. Pegueroles and F. RicoNovella. Enabling secure multicast using a new Java LKH rekeying tool. Lecture Notes in Computer Science, 2722:293–294, 2003. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Proulx:2004:JIT

[PR04] V. K. Proulx and R. Rasala.

Java IO and testing made simple. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):161–165, 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Prasad:2003:OSJ

[Pra03] G. Prasad. Open Source Java:

Fortune 500 systems at twoguys-in-a-garage prices. Cutter IT Journal, 16(5):10–15, 2003. CODEN ???? ISSN 1522-7383.

Pratter:2008:SGJ

[Pra08] Frederick M. Pratter. SAS graphics for Java: Examples using SAS AppDev Studio and the Output Delivery System. The American Statistician, 62(4):359, November 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Permandla:2007:TSP

[PRB07] Pratibha Permandla, Michael Roberson, and Chandrasekhar Boyapati. A type system for preventing data races and deadlocks in the Java Virtual Machine language: 1. ACM SIGPLAN Notices, 42(7):10, July 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Prechelt:2000:ECS

[Pre00a] Lutz Prechelt. An empirical comparison of seven programming languages. Computer, 33(10):23–29, October 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//dlib.computer.org/co/
books/co2000/pdf/rx023.
pdf;   http://www.computer.
org/computer/co2000/rx023abs.
htm.

Preiss:2000:DSA

[Pre00b] Bruno R. Preiss. Data structures and algorithms with object-oriented design patterns in Java. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 2000. ISBN 0-471-34613-6 (cloth). xvii + 635 pp. LCCN QA76.64 .P744 2000.

Prechelt:2003:SLG

[Pre03] L. Prechelt. Are scripting languages any good? A validation of Perl, Python, Rexx, and Tcl against C, C++, and Java. Advances in Computers, 57:207–271, 2003. CO-DEN ???? ISSN 0065-2458.

Price:2001:JPO

[Pri01]  Jason Price.  Java pro-

gramming with Oracle SQLJ. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2001. ISBN 0-596-00087-1 (paperback). xiii + 381 pp. LCCN QA76.73.J38 P74 2001.

Prochazka:2001:ATE

[Pro01] Marek Prochazka. Advanced transactions in Enterprise Java Beans. Lecture Notes in Computer Science, 1999:215– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/1999/19990215.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/1999/19990215.
pdf.

Proulx:2002:OBG

[PRR02] Viera K. Proulx, Jeff Raab, and Richard Rasala. Objects from the beginning — with GUIs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):65–69, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Powell:2001:JCR

[PS01] Thomas A. Powell and Fritz Schneider. JavaScript: the complete reference. McGraw-Hill, New York, NY, USA, 2001. ISBN 0-07-219127-9. xxiv + 1079 pp. LCCN QA76.73.J39 P69 2001.

Pugh:2003:MJH

[PS03] W. Pugh and J. Spacco. MP Java: High-performance message passing in Java using Java.nio. Lecture Notes in Computer Science, 2958:323– 339, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pawlak:2001:JFS

[PSDF01] Renaud Pawlak, Lionel Seinturier, Laurence Duchien, and Gérard Florin. JAC: a flexible solution for aspectoriented programming in Java. Lecture Notes in Computer Science, 2192:1– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.
                                        com/link/service/series/
                                        0558/bibs/2192/21920001.
                                        htm;   http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2192/21920001.
                                        pdf.

Pratikakis:2004:TPJ

[PSH04] P. Pratikakis, J. Spacco, and M. Hicks. Transparent proxies for Java futures. ACM SIGPLAN Notices, 39(10): 206–223, 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pang:2001:PSR

[PSM01a] James Pang, Gholamali Shoja, and Eric Manning. Providing soft real-time QoS guarantees for Java threads. In ACM [ACM01b], pages 39–46. ISBN 1-58113-359-

6. LCCN QA76.9.O35 A26

2001. URL http://www.
philippsen.com/JGI2001/
camerareadyabstracts/21.
html;   http://www.philippsen.
com/JGI2001/finalpapers/
18500039.pdf.

Pang:2001:SSR

[PSM01b] James C. Pang, Gholamali C.

Shoja, and Eric G. Manning. Supporting soft real-time tasks and QoS on the Java platform. Lecture Notes in Computer Science, 2026:86– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2026/20260086.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2026/20260086.
pdf.

Pang:2003:PSR

[PSM03] James C. Pang, Gholamali C. Shoja, and Eric G. Manning. Providing soft real-time quality of service guarantees for Java threads. Concurrency and Computation: Practice and Experience, 15(3–5):521– 538, March/April 2003. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Praehofer:2001:BWC

[PSS01] Herbert Praehofer, Johannes Sametinger, and Alois Stritzinger. Best of Websim99: Concepts and architecture of a simulation framework based on the JavaBeans component model. Future Generation Computer Systems, 17(5): 539–559, March 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        19/19/45/30/29/abstract.
                                        html.

Perez:2007:RJI

[PSW07] Carlos Javier Pérez, Hansgeorg Schwibbe, and Petra Weidner. RAGE: a Javaimplemented visual random generator. Journal of Statistical Software, 17(10):1– 10, January 2007. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v17/i10.

Padala:2007:ACV

[PSZ^{+}07] Pradeep Padala, Kang G.

Shin, Xiaoyun Zhu, Mustafa Uysal, Zhikui Wang, Sharad Singhal, Arif Merchant, and Kenneth Salem. Adaptive control of virtualized resources in utility computing environments. Operating Systems Review, 41(3): 289–302, June 2007. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Prechelt:2001:IMI

[PT01] Lutz Prechelt and Rainer Typke. An interface for melody input. ACM Transactions on Computer-Human Interaction, 8(2):133–149, 2001. CODEN ATCIF4. ISSN 1073-0516 (print), 1557-7325 (electronic). URL

http://www.acm.org/pubs/
articles/journals/tochi/
2001-8-2/p133-prechelt/
p133-prechelt.pdf;   http:
//www.acm.org/pubs/citations/
journals/tochi/2001-8-2/
p133-prechelt/.

Papadimitriou:2009:JIS

[PT09a] Stergios Papadimitriou and Konstantinos Terzidis. jLab: Integrating a scripting interpreter with Java technology for flexible and efficient scientific computation. Computer Languages, Systems and Structures, 35 (3):217–240, October 2009. CODEN ???? ISSN 1477-8424 (print), 1873-6866 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S147784240800002X.

Pucella:2009:HST

[PT09b] Riccardo Pucella and Jesse A. Tov. Haskell session types with (almost) no class. ACM SIGPLAN Notices, 44(2):25– 36, February 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Papadimitriou:2009:SSJ

[PTML09] Stergios Papadimitriou, Konstantinos Terzidis, Seferina Mavroudi, and Spiridon Likothanassis. Scientific scripting for the Java platform with jLab. Computing in Science and Engineering, 11 (4):50–60, July/August 2009. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Pothier:2007:SOD

[PTP07] Guillaume Pothier, Éric Tanter, and José Piquer. Scalable omniscient debugging. ACM SIGPLAN Notices, 42 (10):535–552, October 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pfeffer:2004:RTG

[PUF^{+}04] M. Pfeffer, T. Ungerer, S. Fuhrmann, J. Kreuzinger, and U. Brinkschulte. Realtime garbage collection for a multithreaded Java microcontroller. Real-Time Systems, 26(1):89–106, 2004. CODEN RESYE9. ISSN 0922-6443.

Pugh:2000:JMM

[Pug00] William Pugh. The Java memory model is fatally flawed. Concurrency: Practice and Experience, 12(6): 445–455, May 2000. CO-DEN CPEXEI. ISSN 1040-

3108. URL http://www3.
interscience.wiley.com/
cgi-bin/abstract/72515725/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=72515725&PLACEBO=IE.
pdf.

Palacz:2003:JST

[PV03a] K. Palacz and J. Vitek. Java subtype tests in real-time. Lecture Notes in Computer Science, 2743:378–404, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pedersen:2003:JPS

[PV03b] K. S. Pedersen and B. Vinter.

Java PastSet: a structured distributed shared memory system. IEE Proceedings Software, 150(2):147–154, 2003. CODEN ???? ISSN 1462-5970.

Pasareanu:2004:VJP

[PV04] C. S. Pasareanu and W. Visser.

Verification of Java programs using symbolic execution and invariant generation. Lecture Notes in Computer Science, 2989:164–181, 2004. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pickett:2006:SSF

[PV06] Christopher J. F. Pickett and Clark Verbrugge. SableSpMT: a software framework for analysing speculative multithreading in Java. ACM SIGSOFT Software Engineering Notes, 31(1):59– 66, January 2006. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Prokopski:2008:APC

[PV08] Gregory B. Prokopski and Clark Verbrugge. Analyzing the performance of codecopying virtual machines. ACM SIGPLAN Notices, 43(10):403–422, September 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Paleczny:2001:JHS

[PVC01] Michael Paleczny, Christopher Vick, and Cliff Click. The Java HotSpot^{T M} server compiler. In USENIX Association [USE01c], page ?? ISBN 1-880446-11-1. LCCN QA76.73.J38 J42 2001. URL

http://www.usenix.org/publications/
library/proceedings/jvm01/

paleczny.html. Sponsored

by the USENIX Association.

Poll:2001:FSJ

[PvdBJ01] Erik Poll, Joachim van den Berg, and Bart Jacobs. Formal specification of the JavaCard API in JML: the APDU class. Computer Networks (Amsterdam, Netherlands: 1999), 36(4):407– 421, July 16, 2001. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL http://

www.elsevier.nl/gej-ng/
10/15/22/61/28/28/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/15/22/61/28/
28/article.pdf.

Pearce:2007:PA

[PWBK07] David J. Pearce, Matthew Webster, Robert Berry, and Paul H. J. Kelly. Profiling with AspectJ. Software —Practice and Experience, 37 (7):747–777, June 2007. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Pooley:2000:DDM

[PWC00] R. Pooley, P. Wilcox, and Correspondence. Distributing decision making using Java simulation across the World Wide Web. OR: the journal of the Operational Research Society, 51(4):395–404, April 1, 2000. CODEN OPRQAK. ISSN 0160-5682 (print), 1476-9360 (electronic).

Pike:2000:CCC

[PWH00] Scott M. Pike, Bruce W. Weide, and Joseph E. Hollingsworth. Checkmate: cornering C++ dynamic memory errors with checked pointers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1): 352–356, March 2000. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pietrzak:2004:ABS

[PWN04] B. Pietrzak, B. Walter, and J. R. Nawrocki. Automatic bad-smell detection in Java programs. Foundations of Computing and Decision Sciences, 29(1/4):149–165, 2004. CODEN ???? ISSN 0867-6356.

Parson:2000:JNI

[PZ00] Dale Parson and Zhenyu Zhu. Java Native Interface idioms for C++ class hierarchies. Software—Practice and Experience, 30(15):1641– 1660, December 2000. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (elec-

                                        tronic). URL http://www3.
                                        interscience.wiley.com/
                                        cgi-bin/abstract/73501271/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=73501271&PLACEBO=IE.
                                        pdf.

Qian:2000:FSJ

[QGC00] Zhenyu Qian, Allen Goldberg, and Alessandro Coglio. A formal specification of Java^{T M} class loading. ACM SIGPLAN Notices, 35(10): 325–336, October 2000. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
proceedings/oops/353171/
p325-qian/.

Qian:2003:ARB

[QH03] Feng Qian and Laurie Hendren. An adaptive, regionbased allocator for Java. ACM SIGPLAN Notices, 38 (2s):233–244, February 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Qian:2002:CAA

[QHV02] Feng Qian, Laurie Hendren, and Clark Verbrugge. A comprehensive approach to array bounds check elimination for Java. Lecture Notes in Computer Science, 2304:325– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2304/23040325.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2304/23040325.
pdf.

Qian:2000:SFI

[Qia00]  Zhenyu Qian.  Standard

fixpoint iteration for Java bytecode verification. ACM Transactions on Programming Languages and Systems, 22(4):638–672, 2000. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/toplas/2000-22-
                                        4/p638-qian/.

Qi:2009:MTS

[QM09a] Xin Qi and Andrew C. Myers. Masked types for sound object initialization. ACM SIGPLAN Notices, 44(1):53– 65, January 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Qi:2009:SCB

[QM09b] Xin Qi and Andrew C. Myers. Sharing classes between families. ACM SIGPLAN Notices, 44(6):281–292, June 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Quigley:2003:PLJ

[Qui03] C. L. Quigley. A programming logic for Java bytecode programs. Lecture Notes in Computer Science, 2758:41– 54, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rellermeyer:2007:CSP

[RA07] Jan S. Rellermeyer and Gustavo Alonso. Concierge: a service platform for resourceconstrained devices. Operating Systems Review, 41(3): 245–258, June 2007. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Rutherford:2002:REJ

[RAC^{+}02] Matthew J. Rutherford, Kenneth Anderson, Antonio Carzaniga, Dennis Heimbigner, and Alexander L. Wolf. Reconfiguration in the enterprise JavaBean component model. Lecture Notes in Computer Science, 2370:67– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2370/23700067.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2370/23700067.
pdf.

Ruiz:2004:FRD

[RAC^{+}04] J. L. Ruiz, J. L. Arciniegas, R. Ceron, J. Bermejo, and J. C. Duenas. A framework for resolution of deployment dependencies in Javaenabled service gateways. Lecture Notes in Computer Science, 2952:1–11, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Radenski:2006:PFL

[Rad06] Atanas Radenski. “Python first”: a lab-based digital introduction to computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):197–201, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Roman:2002:MEJ

[RAJ02] Ed Roman, Scott Ambler, and Tyler Jewell. Mastering Enterprise JavaBeans. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, second edition, 2002. ISBN 0-471-41711-4 (PA-PER/WEBSITE). xxxii + 639 pp. LCCN QA76.73.J38 R658 2002.

Raner:2002:LJV

[Ran02] Mirko Raner. A lightweight Java Virtual Machine for a stack-based microprocessor. In USENIX Association [USE02], page ?? ISBN 1-931971-01-3. LCCN QA76.73.J38 J42 2002. URL

                                        http://www.usenix.org/publications/
                                        library/proceedings/javavm02/
                                        raner.html.

Rana:2003:WJP

[Ran03] Rana. Wood: Java programming for spatial sciences. Environment and Planning B, 30 (3):478, 2003. CODEN ???? ISSN 0265-8135.

Rao:2000:UJa

[Rao00a] Prithvi Rao. Using Java.

;login: the USENIX Association newsletter, 25(1): ??, February 2000. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
java/usingjava20.html.

Rao:2000:UJb

[Rao00b] Prithvi Rao. Using Java.

;login: the USENIX Association newsletter, 25(2): ??, April 2000. CODEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
java/usingjava21.html.

Rao:2000:UJc

[Rao00c] Prithvi Rao. Using Java.

;login: the USENIX Association newsletter, 25(3): ??, June 2000. CODEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
java/usingjava22.html.

Rao:2000:UJd

[Rao00d] Prithvi Rao. Using Java.

;login: the USENIX Association newsletter, 25(5): ??, August 2000. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
java/usingjava22.html.

Rao:2000:UJf

[Rao00e] Prithvi Rao. Using Java.

;login: the USENIX Association newsletter, 25(6): ??, October 2000. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
java/usingjava24.pdf.

Rao:2000:UJg

[Rao00f] Prithvi Rao. Using Java. ;login: the USENIX Association newsletter, 25(8): ??, December 2000. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
java/usingjava25.pdf.

Rao:2001:UCJa

[Rao01a] Prithvi Rao. Using CORBA with Java. ;login: the USENIX Association newsletter, 26(1):??, February 2001. CODEN LOGNEM. ISSN 1044-6397. URL http://

www.usenix.org/publications/
java/usingjava26.pdf.

Rao:2001:UCJb

[Rao01b] Prithvi Rao. Using CORBA with Java. ;login: the USENIX Association newsletter, 26(3):??, June 2001. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
java/usingjava27.pdf.

Rao:2002:JQ

[Rao02]  Surlu Rao.  Java Q&A:

How do message-driver beans work in EJB 2.0? Dr. Dobb’s Journal of Software Tools, 27(2):94–96, February 2002. CODEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/ftp/2002/
2002_02/jqa0202.txt;   http:
//www.ddj.com/ftp/2002/
2002_02/jqa0202.zip.

Rapaport:2003:TPJ

[Rap03] L. Rapaport. Theory and practice: The Java content API. Transform Magazine, 12 (1):48, 2003. CODEN ????

Rasala:2000:TFY

[Ras00] Richard Rasala. Toolkits in first year computer science: a pedagogical imperative. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):185–191, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rasala:2003:EOV

[Ras03] Richard Rasala. Embryonic object versus mature object: object-oriented style and pedagogical theme. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):89– 93, September 2003. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Russell:2001:HSA

[RB01] Kenneth Russell and Lars Bak. The HotSpot^{T M} serviceability agent: An outof-process high-level debugger for a Java^{T M} Virtual Machine. In USENIX Association [USE01c], page ?? ISBN 1-880446-11-1. LCCN QA76.73.J38 J42 2001. URL

http://www.usenix.org/publications/
library/proceedings/jvm01/
russell.html. Sponsored by

the USENIX Association.

Rodziewicz:2004:OAJ

[RB04] P. Rodziewicz and B. Bell. Overview and architecture of the Java integration framework, hybrid scheduler, and Web-enabled LIMS. Journal — Association for Laboratory Automation, 9(6):411– 420, 2004. CODEN ???? ISSN 1535-5535.

Roberts:2005:AJT

[RBC^{+}05] E. Roberts, K. Bruce, R. Cutler, J. H. Cross, S. Grissom, K. Klee, S. Rodger, F. Trees, I. Utting, and F. Yellin. The ACM Java Task Force: Status report. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):46–47, 2005. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Roberts:2006:AJT

[RBC^{+}06] Eric Roberts, Kim Bruce, James H. Cross II, Robb Cutler, Scott Grissom, Karl Klee, Susan Rodger, Fran Trees, Ian Utting, and Frank Yellin. The ACM Java Task Force: final report. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):131–132, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Roth:2001:EJA

[RC01] Volker Roth and Vania Conan. Encrypting Java archives and its application to mobile agent security. Lecture Notes in Computer Science, 1991:229–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/1991/19910229.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/1991/19910229.
pdf.

Reis:2004:TPI

[RC04] Charles Reis and Robert Cartwright. Taming a professional IDE for the classroom. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):156–160, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Riley:2001:HPJ

[RCB01] Christopher Riley, Siddhartha Chatterjee, and Rupak Biswas. High-performance Java codes for computational fluid dynamics. In ACM [ACM01b], pages 143–152. ISBN 1-58113-359-6. LCCN QA76.9.O35

A26 2001. URL http://www.
philippsen.com/JGI2001/
camerareadyabstracts/5.
html;   http://www.philippsen.
com/JGI2001/finalpapers/
18500143.pdf.

Riley:2003:HPJ

[RCB03] Christopher J. Riley, Siddhartha Chatterjee, and Rupak Biswas. Highperformance Java codes for computational fluid dynamics. Concurrency and Computation: Practice and Experience, 15(3–5):395–415, March/April 2003. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Romero:2002:VAR

[RCdBL02] Pablo Romero, Richard Cox, Benedict du Boulay, and Rudi Lutz. Visual attention and representation switching during Java program debugging: a study using the restricted focus viewer. Lecture Notes in Computer Science, 2317:221– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2317/23170221.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2317/23170221.
pdf.

Ren:2006:IFC

[RCR06] Xiaoxia Ren, O. C. Chesley, and B. G. Ryder. Identifying failure causes in Java programs: An application of change impact analysis. IEEE Transactions on Software Engineering, 32(9):718–732, September 2006. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
1707669.

Russell:2006:ESRa

[RD06] Kenneth Russell and David Detlefs. Eliminating synchronizationrelated atomic operations with biased locking and bulk rebiasing. ACM SIGPLAN Notices, 41(10):263–272, October 2006. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Reis:2007:BVD

[RDW^{+}07] Charles Reis, John Dunagan, Helen J. Wang, Opher Dubrovsky, and Saher Esmeir. BrowserShield: Vulnerabilitydriven filtering of dynamic HTML. ACM Transactions on the Web (TWEB), 1(3): 11:1–11:??, September 2007. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Renaud:2001:JRJ

[RE01] Karen Renaud and Huw Evans. JavaCloak: Reflecting on Java typing for class reuse using proxies. Lecture Notes in Computer Science, 2192:278–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2192/21920278.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2192/21920278.
pdf.

Reddy:2001:FJP

[Red01] Achut Reddy. Fast Java: Performance Tuning Guide. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2001. ISBN 0-13-081245-5. ???? pp. LCCN ???? US$45.

Reese:2000:DPJ

[Ree00] George Reese. Database programming with JDBC and Java. Java series. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2000. ISBN 1-56592-616-1. xvii + 328 pp. LCCN QA76.9.D26 R44 2000.

Reed:2001:RCJ

[Ree01] David Reed. Rethinking CS0 with JavaScript. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):100– 104, March 2001. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Reed:2002:DAJ

[Ree02] Paul R. Reed. Developing applications with Java and UML. Addison-Wesley, Reading, MA, USA, 2002. ISBN 0-201-70252-5. xxvi + 463 pp. LCCN QA76.73.J38 R44 2002.

Reese:2003:JDB

[Ree03] George Reese. Java database best practices. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2003. ISBN 0-596-00522-9. xvi + 267 pp. LCCN QA76.73 .J38 R445 2003. URL http://

www.oreilly.com/catalog/
9780596005221.

Reges:2000:CRJ

[Reg00]  Stuart Reges.  Conserva-

tively radical Java in CS1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32 (1):85–89, March 2000. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Reges:2002:CCR

[Reg02a]  S. Reges.  Can C# re-

place Java in CS1 and CS2? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):4–8, 2002. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Reges:2002:SFI

[Reg02b] Stuart Reges. SIGCSE 2002 forum: industry in the ivory tower. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):367, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Reges:2006:BBC

[Reg06] Stuart Reges. Back to basics in CS1 and CS2. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):293– 297, March 2006. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Reilly:2000:JQH

[Rei00a] David Reilly. Java Q&A: How do I user servlets for state and session management? Dr. Dobb’s Journal of Software Tools, 25(5):111–115, May 2000. CODEN DDJOEB. ISSN 1044-789X. URL

                                        http://www.ddj.com/ftp/
                                        2000/2000_05/jqa500.txt;
                                         http://www.ddj.com/ftp/
                                        2000/2000_05/jqa500.zip.

Reinholtz:2000:JWF

[Rei00b] Kirk Reinholtz. Java will be faster than C++. ACM SIGPLAN Notices, 35(2):25– 28, 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Reinholtz:2000:TCJ

[Rei00c]  Kirk Reinholtz.  Techni-

cal correspondence: Java will be faster than C++. ACM SIGPLAN Notices, 35 (2):25–28, February 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://ftp.
informatik.rwth-aachen.
de/dblp/db/indices/a-tree/
r/Reinholtz:Kirk.html.

Reiss:2003:JVJ

[Rei03] S. P. Reiss. JIVE: Visualizing Java in action-demonstration description. International Conference on Software Engineering, 25(??):820–821, 2003. CODEN ???? ISSN 0270-5257.

Reiss:2005:DDV

[Rei05] Steven P. Reiss. Dynamic detection and visualization of software phases. ACM SIG-SOFT Software Engineering Notes, 30(4):1–6, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Rempt:2001:SJP

[Rem01] Boudewijn Rempt. Scripting with Java and Python: Building a Python console window in a Java application. Dr. Dobb’s Journal of Software Tools, 26(10):56, 60–61, October 2001. CODEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/ftp/2001/
2001_10/pyconsol.txt;
http://www.ddj.com/ftp/
2001/2001_10/pyconsol.zip.

Renaud:2000:HNI

[Ren00]  Karen Renaud.  HER-

CULE: Non-invasively tracking Java TM componentbased application activity. Lecture Notes in Computer Science, 1850:447–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/1850/18500447.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/1850/18500447.
pdf.

Renaud:2002:ESG

[Ren02] K. Renaud. Experience with statically-generated proxies for facilitating Java runtime specialisation. IEE Proceedings Software, 149(6):169– 178, 2002. CODEN ???? ISSN 1462-5970.

Requet:2003:BME

[Req03] Antoine Requet. A B model for ensuring soundness of a large subset of the Java Card virtual machine. Science of Computer Programming, 46(3):283–306, March 2003. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Radenski:2008:JGC

[RFZ08] Atanas Radenski, Jeff Furlong, and Vladimir Zanev. The Java 5 generics compromise orthogonality to keep compatibility. The Journal of systems and software, 81(11): 2069–2078, November 2008. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Rousselle:2000:PSJ

[RG00] Philip Rousselle and Daniel Greff. Publish, subscribe, and the JMS API. Dr. Dobb’s Journal of Software Tools, 25(7):44, 46, 48, 51, July 2000. CODEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/ftp/2000/
2000_07/pubsub.zip.

Richards:2005:JDN

[RG05] Norman Richards and Sam Griffith. JBoss: a developer’s notebook. The developer’s notebook series. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-10007-8. xix + 149 pp. LCCN TK5105.8885.J43 R53 2005.

Ruiz:2007:JLC

[RGN07] Irene Luque Ruiz and Miguel Ángel Gómez-Nieto. A Java library for the calculation of molecular descriptors. In Simos and Maroulis [SM07], pages 1347–1350. ISBN 0-7354-0476-3 (set), 0-7354-0477-1 (vol. 1), 0-7354-0478-X (vol. 2). LCCN Q183.9 2007.

URL http://proceedings.
aip.org/getpdf/servlet/
GetPDFServlet?filetype=
pdf&amp;   id=APCPCS000963000002001347000001&
amp; idtype=cvips. Two

volumes.

Ranganath:2004:PIR

[RH04] V. P. Ranganath and J. Hatcliff. Pruning interference and ready dependence for slicing concurrent Java programs. Lecture Notes in Computer Science, 2985:39– 56, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ranganath:2007:SCJ

[RH07] Venkatesh Prasad Ranganath and John Hatcliff. Slicing concurrent Java programs using Indus and Kaveri. International Journal on Software Tools for Technology Transfer (STTT), 9(5–6):489–504, October 2007. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=1433-2779&volume=9&
issue=5&spage=489.

Roberson:2008:ESM

[RHDB08] Michael Roberson, Melanie Harries, Paul T. Darga, and Chandrasekhar Boyapati. Efficient software model checking of soundness of type systems. ACM SIGPLAN Notices, 43(10):493–504, September 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rajan:2002:CPJ

[RHR02] A. S. Rajan, Shiwen Hu, and J. Rubio. Cache performance in Java virtual machines: a study of constituent phases. In IEEE [IEE02b], pages 81–90. ISBN 0-7803-7681-1. LCCN QA76.9.S88; QA76.9.S88 W67 2002eb; Internet. URL

http://ieeexplore.ieee.
org/servlet/opac?punumber=

8689. IEEE Catalog Number 02EX633.

Richter:2000:IYA

[Ric00] Mathias W. Richter. Iava:

yet another interpreter for scripting within the Java platform. Software—Practice and Experience, 30(2): 81–106, February 2000. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (elec-

tronic). URL http://www3.
interscience.wiley.com/
cgi-bin/abstract/69503461/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=69503461&PLACEBO=IE.
pdf.

Riccardi:2001:PDS

[Ric01] Greg Riccardi. Principles of database systems with Internet and Java applications. Addison-Wesley, Reading, MA, USA, 2001. ISBN 0-201-61247-X. xxi + 518 pp. LCCN QA76.9.D3 R52 2001.

Richardson:2006:PAD

[Ric06a]  Chris Richardson.  PO-

JOs in Action: Developing Enterprise Applications with Lightweight Frameworks. Manning Publications, Greenwich, CT, USA, 2006. ISBN 1-932394-58-3 (paperback). xxxii + 560 pp. LCCN ???? US$44.95 (paperback), US$22.50 (elec-

                                        tronic). URL http://java.
                                        sun.com/developer/Books/
                                        javaprogramming/pojos/;
                                         http://www.manning.com/
                                        books/crichardson.

Richardson:2006:UEJ

[Ric06b] Chris Richardson. Untangling Enterprise Java. ACM Queue: Tomorrow’s Computing Today, 4(5):36–44, June 2006. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Riley:2002:OJI

[Ril02] David D. Riley. The object of Java: introduction to programming using software engineering principles. Addison-Wesley, Reading, MA, USA, 2002. ISBN 0-201-71585-6 (paperback). xxvi + 673 pp. LCCN QA76.73.J38 R54 2002.

Riley:2003:OJI

[Ril03] David D. Riley. The object of Java: introduction to Java using software engineering principles. Addison-Wes-ley, Reading, MA, USA, 2003. ISBN 0-321-12173-2 (paperback). xxviii + 722 pp. LCCN QA76.73.J38 R54 2003.

Riordan:2002:TIL

[Rio02] Denis Riordan. Towards an integrated learning laboratory environment for firstyear computer science students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(4):112–116, December 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp:

//ftp.math.utah.edu/pub/
mirrors/ftp.ira.uka.de/
bibliography/Misc/DBLP/
2002.bib.

Rodriguez:2003:DSM

[RJFG03] J. I. Rodriguez, J. M.

Jimenez, F. J. Funes, and J. GarciadeJalon. Dynamic simulation of multibody systems on Internet using CORBA, Java and XML. Multibody System Dynamics, 10(2):177–199, 2003. CODEN ???? ISSN 1384-5640.

Rozman:2006:QQA

[RJGH06] Ivan Rozman, Matjaz B.

Juric, Izidor Golob, and Marjan Hericko. Qualitative and quantitative analysis and comparison of Java distributed architectures. Software—Practice and Experience, 36(14):1543–1562, November 25, 2006. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rayside:2002:EJL

[RK02] Derek Rayside and Kostas Kontogiannis. Extracting Java library subsets for deployment on embedded systems. Science of Computer Programming, 45(2–3):245– 270, November/December 2002. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Rountev:2004:SDA

[RKG04] A. Rountev, S. Kagan, and M. Gibas. Static and dynamic analysis of call chains in Java. Software Engineering Notes, 29(4):1–11, 2004. CO-DEN ???? ISSN 0163-5948.

Rojiani:2003:WBJ

[RKK03] Rojiani, Kim, and Kapania. Web-based Java applets for teaching engineering mechanics. Computers in education journal, 13(4):59–67, 2003. CODEN CEJOE7. ISSN 1069-3769.

Rukoz:2000:SJT

[RLR00] M. Rukoz, C. Leon, and M. Rivas. SIMA: a Java tool for constructing image processing applications on a heterogeneous network. Parallel and Distributed Computing Practices, 3(1):??, ???? 2000. CODEN ???? ISSN 1097-

2803. URL http://www.cs.
okstate.edu/~pdcp/vols/
vol03/vol03no1abs.html#
rukoz.

Robillard:2000:DRJ

[RM00] Martin P. Robillard and Gail C. Murphy. Designing robust Java programs with exceptions. ACM SIGSOFT Software Engineering Notes, 25(6):2–10, November 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ramirez:2004:CBS

[RM04] R. Ramirez and J. Martinez. Constraint-based synchronization and verification of distributed Java programs. Lecture Notes in Computer Science, 3132:473–474, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rafieymehr:2007:JVD

[RM07a] Ali Rafieymehr and Richard McKeever. Java visual debugger. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(2):75–79, June 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp:

//ftp.math.utah.edu/pub/
mirrors/ftp.ira.uka.de/

                                        bibliography/Misc/DBLP/
                                        2007.bib.

Robillard:2007:RCS

[RM07b] Martin P. Robillard and Gail C. Murphy. Representing concerns in source code. ACM Transactions on Software Engineering and Methodology, 16(1):3:1–3:??, February 2007. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Reyes:2008:GDJ

[RM08] Miguel Reyes and Águeda Mata. Games developed in Java for teaching “combinatorial game theory”. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3): 359, September 2008. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Richards:2009:JMS

[RMHC09] Mark Richards, Richard Monson-Haefel, and David Chappell. Java Message Service. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2009. ISBN 0-596-52204-5. xx + 305 pp. LCCN ???? US$39.99.

Rountev:2001:PAJ

[RMR01] Atanas Rountev, Ana Milanova, and Barbara G. Ryder. Points-to analysis for Java using annotated constraints. ACM SIGPLAN Notices, 36(11):43–55, November 2001. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Proceedings of the 2001 ACM SIGPLAN Conference on Object Oriented Programming, Systems, Languages and Applications (OOPSLA’01).

Rountev:2003:FCA

[RMR03] A. Rountev, A. Milanova, and B. G. Ryder. Fragment class analysis for testing of polymorphism in Java software. International Conference on Software Engineering, 25(??): 210–220, 2003. CODEN ???? ISSN 0270-5257.

Rountev:2004:FCA

[RMR04] A. Rountev, A. Milanova, and B. G. Ryder. Fragment class analysis for testing of polymorphism in Java software. IEEE Transactions on Software Engineering, 30(6):372–387, June 2004. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
1321060.

Robbins:2000:EBB

[Rob00a] Steven Robbins. Experimentation with bounded buffer synchronization. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):330– 334, March 2000. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Robbins:2000:RLJ

[Rob00b] Steven Robbins. Remote logging in Java using Jeli: a facility to enhance development of accessible educational software. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):114–118, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Robbins:2001:SPE

[Rob01a] Steven Robbins. Starving philosophers: experimentation with monitor synchronization. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):317–321, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Roberts:2001:OM

[Rob01b] Eric Roberts. An overview of MiniJava. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):1–5, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Robison:2001:ICE

[Rob01c] Arch Robison. Impact of compiler economics on program optimization. In ACM [ACM01b], pages 1–

10. ISBN 1-58113-359-

6. LCCN QA76.9.O35 A26

2001. URL http://www.
philippsen.com/JGI2001/
camerareadyabstracts/13.
html;   http://www.philippsen.
com/JGI2001/finalpapers/
001.ps.

Robbins:2002:EPI

[Rob02] Steven Robbins. Exploration of process interaction in operating systems: a pipe-fork simulator. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):351–355, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Robbins:2003:URL

[Rob03] Steven Robbins. Using remote logging for teaching concurrency. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):177–181, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Robbins:2004:DHS

[Rob04a] Steven Robbins. A disk head scheduling simulator.

SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36 (1):325–329, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Roberts:2004:RSU

[Rob04b] E. Roberts. Resources to support the use of Java in introductory computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36 (1):233–234, 2004. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Roberts:2004:DCL

[Rob04c] Eric Roberts. The dream of a common language: the search for simplicity and stability in computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):115–119, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Roberts:2006:ITS

[Rob06] Eric Roberts. An interactive tutorial system for Java. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38 (1):334–338, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Robbins:2007:JES

[Rob07a] Steven Robbins. A Java execution simulator. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):536– 540, March 2007. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Roberts:2007:RAP

[Rob07b] Eric Roberts. Resurrecting the applet paradigm. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39 (1):521–525, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rockwell:2001:XXJ

[Roc01]  Westy Rockwell.  XML,

XSLT, Java and JSP: a case study in developing a Web application. New Riders Publishing, Carmel, IN, USA, 2001. ISBN 0-7357-1089-9. xviii + 746 pp. LCCN QA76.76.H94 R62 2000.

Rodrigues:2001:BIA

[Rod01] Lawrence H. Rodrigues.

Building imaging applications with Java technology: using AWT Imaging, Java 2D, and Java Advanced Imaging (JAI). Addison-Wesley, Reading, MA, USA, 2001. ISBN 0-201-70074-3 (paperback). xx + 857 pp. LCCN QA76.73.J38 R65 2001.

Roelofs:2000:JCC

[Roe00] Brian Roelofs. Java custom class loaders. Dr. Dobb’s Journal of Software Tools, 25 (6):74, 76, 78, 80, 82, June 2000. CODEN DDJOEB. ISSN 1044-789X. URL http:

                                        //www.ddj.com/ftp/2000/
                                        2000_06/javaload.txt;
                                        http://www.ddj.com/ftp/
                                        2000/2000_06/javaload.zip.

Rogatkin:2003:JNI

[Rog03] D. Rogatkin. JAVA NIO & the iTunes database. Dr. Dobb’s Journal of Software Tools, 28(12):36–47, 2003. CODEN DDJOEB. ISSN 1044-789X.

Rojas:2000:SKZ

[Roj00]  Raul Rojas.  Simulating

Konrad Zuse’s computers. Dr. Dobb’s Journal of Software Tools, 25(9):64, 66–69, September 2000. CODEN DDJOEB. ISSN 1044-789X.

                                        URL http://www.ddj.com/
                                        ftp/2000/2000_09/z3.zip.

Rolfe:2005:LPS

[Rol05] Timothy J. Rolfe. List processing: sort again, naturally. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(2):46–48, June 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp:

                                        //ftp.math.utah.edu/pub/
                                        mirrors/ftp.ira.uka.de/

bibliography/Misc/DBLP/
2005.bib.

Rolfe:2008:PFO

[Rol08a]  Timothy J. Rolfe.  Per-

verse and foolish oft I strayed. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(2):52–55, June 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp:

//ftp.math.utah.edu/pub/
mirrors/ftp.ira.uka.de/
bibliography/Misc/DBLP/
2008.bib.

Rolfe:2008:SMA

[Rol08b] Timothy J. Rolfe. A specimen MPI application: N-queens in parallel. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(4):42–45, December 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ronthal:2001:WJI

[Ron01] Adam Ronthal. Why Java isn’t appropriate for everything. SysAdmin, 10(7): 52, 54–55, July 2001. CO-DEN SYADE7. ISSN 1061-

2688. URL http://www.
samag.com/.

Roseman:2000:PTJ

[Ros00] Mark Roseman. Proxy Tk:

a Java applet user interface toolkit for Tcl. In USENIX [USE00b], page ??

ISBN 1-880446-24-3. LCCN

????  URL http://db.
usenix.org/publications/
library/proceedings/tcl2k/
roseman.html.

Rose:2002:OJM

[Ros02a] J. F. Rose. The OPUS Java managers. Astronomical Society of the Pacific Conference Series, 281(??):445–448, 2002. CODEN ???? ISSN 1050-3390.

Ross:2002:GST

[Ros02b] J. M. Ross. Guiding students through programming puzzles: Value and examples of Java game assignments. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(4):94–98, 2002. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rose:2003:LBV

[Ros03] Eva Rose. Lightweight bytecode verification. Journal of Automated Reasoning, 31(3–4):303–334, November 2003. CODEN JA-REEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/B%3AJARS.0000021015.
15794.82.

Rossling:2006:TPI

[Röß06] Guido Rößling. Translator: a package for internationalization for Java-based applications and GUIs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3): 312, September 2006. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Roth:2002:JSA

[Rot02] Volker Roth. Java security architecture and extensions. Dr. Dobb’s Journal of Software Tools, 27(4):34, 36–38, April 2002. CODEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/ftp/2002/
2002_04/jca.txt;   http:
//www.ddj.com/ftp/2002/
2002_04/jca.zip.

Roth:2005:SVE

[Rot05] A. Roth. Specification and verification of encapsulation in Java programs. Lecture Notes in Computer Science, 3535:195–210, 2005. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roumani:2002:DGL

[Rou02a] Hamzeh Roumani. Design guidelines for the lab component of objects-first CS1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34 (1):222–226, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Rousselle:2002:IJP

[Rou02b]  Philip Rousselle.  Imple-

menting the JMS publish/ subscribe API. Dr. Dobb’s Journal of Software Tools, 27 (4):28, 30–32, April 2002. CODEN DDJOEB. ISSN 1044-789X. URL http://

                                        www.ddj.com/.

Rajaravivarma:2003:WIO

[RP03a] R. Rajaravivarma and I. Pevac. When to introduce objects in teaching Java. Southeastern Symposium on System Theory, 35(??):437–440, 2003. CODEN ???? ISSN 0094-2898.

Ryan:2003:MDC

[RP03b] C. Ryan and S. Perry. Mob-JeX: a declaratively configurable Java based framework for resource aware object mobility. Lecture Notes in Computer Science, 2889:11– 13, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Raymond:2006:PQR

[RP06] Scott Raymond and Sergio Pereira. Prototype quick reference. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2006. ISBN 0-596-51018-7. LCCN QA76.73.J38 R39 2006eb; QA76.73.J38. URL

                                        http://www.oreilly.com/
                                        catalog/9780596510183.

Roy:2009:LPF

[RPB^{+}09] Indrajit Roy, Donald E.

Porter, Michael D. Bond, Kathryn S. McKinley, and Emmett Witchel. Laminar: practical fine-grained decentralized information flow control. ACM SIGPLAN Notices, 44(6):63–74, June 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rodriguez:2004:ETJ

[RPJ04] N. Rodriguez, L. J. Pulido, and J. P. Jessel. Enhancing a telerobotics Java tool with augmented reality. Lecture Notes in Computer Science, 3061:9–18, 2004. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rossi:2007:JJL

[RPP07] G. Rossi, E. Panegai, and E. Poleo. JSetL: a Java library for supporting declarative programming in Java. Software—Practice and Experience, 37(2):115–149, February 2007. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Rose:2001:JAP

[RR01] Eva Rose and Kristoffer Høgsbro Rose. Java access protection through typing. Concurrency and Computation: Practice and Experience, 13(13):1125– 1132, November 2001. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (elec-

tronic). URL http://www3.
interscience.wiley.com/
cgi-bin/abstract/88011340/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=88011340&PLACEBO=IE.
pdf.

Reilly:2002:JNP

[RR02] David Reilly and Michael Reilly. Java network programming and distributed computing. Addison-Wesley, Reading, MA, USA, 2002. ISBN 0-201-71037-4. xvii + 464 pp. LCCN QA76.73.J38 R45 2002.

Raab:2000:PPT

[RRP00] Jeff Raab, Richard Rasala, and Viera K. Proulx. Pedagogical power tools for teaching Java. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):156–159, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rasala:2001:JPT

[RRP01] Richard Rasala, Jeff Raab, and Viera K. Proulx. Java power tools: model software for teaching object-oriented design. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):297–301, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rasala:2002:SMD

[RRP02] Richard Rasala, Jeff Raab, and Viera K. Proulx. The SIGCSE 2001 Maze Demonstration program. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):287– 291, March 2002. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Ramirez:2000:DCJ

[RS00a] Rafael Ramirez and Andrew E. Santosa. Declarative concurrency in Java. Lecture Notes in Computer Science, 1800:332–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/1800/18000332.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/1800/18000332.
pdf.

Rossbach:2000:JSS

[RS00b] Peter Rossbach and Hendrik Schreiber. Java Server and Servlets: Building Portable Web Applications. Addison-Wesley, Reading, MA, USA, 2000. ISBN 0-201-67491-2.

???? pp. LCCN QA76.73.J38 R68 2000. US$45.

Rummler:2001:EJF

[RS01] Andreas Rummler and Gerd Scarbata. eaLib — A Java framework for implementation of evolutionary algorithms. Lecture Notes in Computer Science, 2206:92– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2206/22060092.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2206/22060092.
pdf.

Rainsberger:2005:JRP

[RS05] J. B. (Joe B.) Rainsberger and Scott Stirling. JUnit Recipes: Practical Methods for Programmer Testing. Manning Publications, Greenwich, CT, USA, 2005. ISBN 1-932394-23-0. xxx + 721 pp. LCCN QA76.64 .R35 2005. US$49.95. URL

http://www.manning.com/
rainsberger.

Ritley:2001:DEP

[RSD01] K. A. Ritley, M. Schlestein, and H. Dosch. DataScan: An extensible program for image analysis in Java. Computer Physics Communications, 137(2):300–311, June 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465501001552.

Ramirez:2001:IDC

[RSH01] Rafael Ramirez, Andrew E.

Santosa, and Lee Wei Hong. Implementing declarative concurrency in Java. Lecture Notes in Computer Science, 1900:700–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/1900/19000700.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/1900/19000700.
pdf.

Reimer:2004:SSA

[RSS^{+}04] Darrell Reimer, Edith Schonberg, Kavitha Srinivas, Harini Srinivasan, Bowen Alpern, Robert D. Johnson, Aaron Kershenbaum, and Larry Koved. SABER: smart analysis based error reduction. ACM SIGSOFT Software Engineering Notes, 29 (4):243–251, July 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ren:2004:CTC

[RST^{+}04] X. Ren, F. Shah, F. Tip, B. G. Ryder, and O. Chesley. Chianti: a tool for change impact analysis of Java programs. ACM SIG-PLAN Notices, 39(10):432– 448, 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Revetria:2002:UJA

[RT02] R. Revetria and F. Tonelli. Using Java applet-based application to improve practices in environmental production monitoring. Simulation series, 34(1):13–16, 2002. CO-DEN SMCPAX. ISSN 0735-9276.

Radhakrishnan:2000:AIE

[RTJ00] Ramesh Radhakrishnan, Deependra Talla, and Lizy Kurian John. Allowing for ILP in an embedded Java processor. ACM SIGARCH Computer Architecture News, 28 (2):294–305, May 2000. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Riggs:2001:PWD

[RTVH01] Roger Riggs, Antero Taivalsaari, Mark VandenBrink, and Jim Holliday. Programming wireless devices with the Java 2 platform, micro edition: J2ME Connected Limited Device Configuration (CLDC), Mobil Information Device Profile (MIDP). Java series. Addison-Wesley, Reading, MA, USA, 2001. ISBN 0-201-74627-1. xxiii + 345 pp. LCCN TK5103.2 .R54 2001.

Ruf:2000:ESR

[Ruf00] Erik Ruf. Effective synchronization removal for Java. ACM SIGPLAN Notices, 35(5):208–218, May 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL

http://www.acm.org/pubs/
articles/proceedings/pldi/
349299/p208-ruf/p208-ruf.
pdf;   http://www.acm.org/
pubs/citations/proceedings/
pldi/349299/p208-ruf/.

Rumpe:2001:BNP

[Rum01] Bernhard Rumpe. ≪Java≫OCL based on new presentation of the OCL-syntax. Lecture Notes in Computer Science, 2263:189–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2263/22630189.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2263/22630189.
pdf.

Rajsbaum:2005:OOA

[RV05] Sergio Rajsbaum and Elisa Viso. Object-oriented algorithm analysis and design with Java. Science of Computer Programming, 54(1): 25–47, January 2005. CO-DEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Radhakrishnan:2001:JRS

[RVJ^{+}01] R. Radhakrishnan, N. Vijaykrishnan, L. K. John, A. Sivasubramaniam, J. Rubio, and J. Sabarinathan. Java runtime systems: characterization and architectural implications. IEEE Transactions on Computers, 50(2): 131–146, February 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Rosenschein:2004:WPP

[RVZ04] J. S. Rosenschein, T. Vilner, and E. Zur. Work in progress: Programming knowledge: Does it affect success in the course introduction to computer science using Java? Frontiers in Education Conference, CONF34 (VOL 1):T2H–3–T2H4, 2004. CODEN PFECDR. ISSN 0190-5848.

Rauch:2003:FJT

[RW03a] N. Rauch and B. Wolff. Formalizing Java’s two’scomplement integral type in Isabelle/HOL. Electronic Notes in Theoretical Computer Science, 80:1–18, 2003. CODEN ???? ISSN 1571-0661.

Rudys:2003:EJR

[RW03b] A. Rudys and D. S. Wallach. Enforcing Java run-time properties using bytecode rewriting. Lecture Notes in Computer Science, 2609:185– 200, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ryan:2004:AAT

[RW04] C. Ryan and C. Westhorpe. Application adaptation through transparent and portable object mobility in Java. Lecture Notes in Computer Science, 3291: 1262–1284, 2004. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rosa:2003:SPC

[RWC^{+}03] L. S. Rosa, F. R. Wagner, L. Carro, A. S. Carissimi, and A. I. Reis. Scheduling policy costs on a JAVA microcontroller. Lecture Notes in Computer Science, 2889:520– 533, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reus:2001:HCV

[RWH01] Bernhard Reus, Martin Wirsing, and Rolf Hennicker. A Hoare calculus for verifying Java realizations of OCLconstrained design models. Lecture Notes in Computer Science, 2029:300–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2029/20290300.
htm;   http://link.springer-

                                        ny.com/link/service/series/
                                        0558/papers/2029/20290300.
                                        pdf.

Rahimi:2007:PPA

[RWL07] Shahram Rahimi, Michael Wainer, and Delano Lewis. A performance and programming analysis of Java communication mechanisms in a distributed environment. Scalable Computing: Practice and Experience, 8(2): 209–226, June 2007. CO-DEN ???? ISSN 1895-

                                        1767. URL http://www.
                                        scpe.org/vols/vol08/no2/
                                        SCPE_8_2_07.pdf;   http:
                                        //www.scpe.org/vols/vol08/
                                        no2/SCPE_8_2_07.zip.

Rataj:2009:TJP

[RWZ09] Artur Rataj, Bożena Woźna, and Andrzej Zbrzezny. A translator of Java programs to TADDs. Fundamenta Informaticae, 93(1–3):305– 324, January 2009. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Rui:2003:CMW

[RYD^{+}03] X. Rui, C. Yang, P. Dong, Q. Cheng, and Y. Bai. Coal mine WebGIS developing with Java. International Geoscience and Remote Sensing Symposium, 4:IV:2659– IV:2661, 2003. CODEN ????

Raje:2001:CSD

[RZW01] Rajeev R. Raje, Ming Zhong, and Tongyu Wang. Case study: a distributed concurrent system with AspectJ. ACM SIGAPP Applied Computing Review, 9 (2):17–23, July 2001. CO-DEN ???? ISSN 1559-6915 (print), 1931-0161 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
512000.512004.

BCS:2004:HTJ

[S.04a] B. C. S. Hot topic: Java 1.5.

The Computer Bulletin, 46 (5):32, September 2004. CO-DEN COBUAH. ISSN 0010-4531 (print), 1464-357X (electronic).

S:2004:HTJ

[S.04b] B. C. S. Hot topic: Java 1.5.

The Computer Bulletin, 46 (5):32, September 2004. CO-DEN COBUAH. ISSN 0010-4531 (print), 1464-357X (electronic).

Saini:2002:JMD

[SA02] S. P. Saini and F. Ahmad.

Java model of DSA (Digital Signature Algorithm). IETE Technical Review, 19(4):189– 194, 2002. CODEN ITREEI. ISSN 0256-4602.

Spoonhower:2006:ESP

[SAB^{+}06] Daniel Spoonhower, Joshua Auerbach, David F. Bacon, Perry Cheng, and David Grove. Eventrons: a safe programming construct for highfrequency hard real-time applications. ACM SIGPLAN Notices, 41(6):283–294, June 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shankar:2008:JLD

[SAB08] Ajeet Shankar, Matthew Arnold, and Rastislav Bodik. Jolt: lightweight dynamic analysis and removal of object churn. ACM SIGPLAN Notices, 43(10):127–142, September 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Safonov:2002:VVJ

[Saf02] V. O. (Vladimir Olegovich) Safonov. Vvedenie v Javatekhnologiiu. “Nauka”, Sankt-Peterburg, Russia, 2002. ISBN 5-02-024985-8. 187 pp. LCCN ????

SerraSagrista:2003:JFE

[SAFG03] J. SerraSagrista, F. Auli, C. Fernandez, and F. Garcia. A Java framework for evaluating still image coders applied to remote sensing applications. In IEEE [IEE03a], pages VI:3595–VI:3597. ISBN 0-7803-7929-2 (softbound) 0-7803-7930-6 (CD-ROM). LCCN QE33.2.R4 .I58 2003. 7 volumes.

Sahni:2000:DSA

[Sah00] Sartaj Sahni. Data structures, algorithms, and applications in Java. McGraw-Hill, New York, NY, USA, 2000. ISBN 0-07-109217-X. xxvi + 833 pp. LCCN QA76.73.J38 S24 2000.

Sahu:2001:JSP

[Sah01] Maneesh Sahu. Java Server Pages from scratch. From scratch programming series. Que Corporation, Indianapolis, IN, USA, 2001. ISBN 0-7897-2459-6. ix + 404 pp. LCCN QA76.73.J38 S25 2001. CD-ROM includes: evaluation software for IBM’s Websphere Studio 3.0 entry edition, IBM’s VisualAge for Java 3.0 entry edition, IBM’s DB2 Universal Database professional edition v6.1, Tomcat Reference Implementation v3.1, NetObject’s Fusion demo version, NetStudio 2000, Nokia WAP Toolkit product information.

Saha:2002:RLP

[Sah02a]  Prasenjit Saha.  Rambu-

tan: Literate programming in Java. TUGboat, 23(3/4):304– 309, 2002. ISSN 0896-3207.

Saha:TB23-3-304

[Sah02b]  Prasenjit Saha.  Rambu-

tan: Literate programming in Java. TUGboat, 23(3/4):304– 309, 2002. ISSN 0896-3207.

Sakamura:2001:EMJ

[Sak01] Ken Sakamura. EIC message: a Java-enabled evolution. IEEE Micro, 21(4): 2–3, 5, July/August 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //dlib.computer.org/mi/
                                        books/mi2001/pdf/m4002.
                                        pdf;   m4002abs.htm.

Saldanha:2004:JTE

[Sal04]  Alok J. Saldanha.  Java

Treeview — extensible visualization of microarray data. Bioinformatics, 20(17):3246– 3248, November 22, 2004. CODEN ???? ISSN 1367-4803 (print), 1367-4811 (electronic).

Sally:2006:EJG

[Sal06] Gene Sally. Embedded Java with GCJ. Linux Journal, 2006(145):??, May 2006. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Samet:2004:OBI

[Sam04] Hanan Samet. Object-based and image-based object representations. ACM Computing Surveys, 36(2):159–217, June 2004. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

Sanders:2000:UAA

[San00] Beverly A. Sanders. Using atomic await commands to develop concurrent programs in Java. Software — Concepts and Tools, 19(4):168– 174, October 2000. CO-DEN SCOTE5. ISSN 0945-8115 (print), 1432-2188 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s003789900004.

Sanden:2002:RTP

[San02a] Bo I. Sandén. Real-time programming safety in Java and Ada. Ada User Journal, 23 (2):105–??, June 2002. CO-DEN AUJOET. ISSN 1381-6551.

Santoro:2002:JTT

[San02b] K. Santoro. 2002-21-0002 Java telematics technology — building and leveraging a market ecosystem. SAE Conference Proceedings, 381:13– 18, 2002. CODEN ???? ISSN ????

Sanden:2003:RTP

[San03] B. Sanden. Real-time programming safety in Java and Ada. ACM SIGADA Ada Letters, 23(2):32–46, 2003. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Sanden:2004:CJT

[San04a] B. Sanden. Coping with Java threads: Java works for many kinds of concurrent software, but it was not designed for safety-critical realtime applications and does not protect the programmer from the pitfalls associated with multithreading. Computer, 37(4):20–27, 2004. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sandya:2004:JJL

[San04b] S. M. Sandya. Jazzing up JVMs with off-line profile data: does it pay? ACM SIGPLAN Notices, 39(8):72– 80, August 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sarra:2003:SSP

[Sar03] Scott A. Sarra. The spectral signal processing suite. ACM Transactions on Mathematical Software, 29(2):195–217, June 2003. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Spanias:2003:AJD

[SASZ03] A. Spanias, K. I. U. Ahmed, A. Papandreou Suppappola, and M. Zaman. Assessment of the Java-DSP (J-DSP) on-line laboratory software. Frontiers in Education Conference, 1(??):T2E– 16–T2E–21, 2003. CODEN PFECDR. ISSN 0190-5848.

Sato:2002:SJL

[Sat02] Y. Sato. A study of Java language for effective thread migration. Record of Electrical and Communication Engineering Conversazione Tohoku University, 71(1):597– 598, 2002. CODEN ???? ISSN 0385-7719.

Satoh:2004:CNP

[Sat04] I. Satoh. Configurable network processing for mobile agents on the Internet. Cluster Computing, 7(1):73–83, January 2004. CODEN ???? ISSN 1386-7857.

Savitch:2001:JIC

[Sav01] Walter J. Savitch. Java:

An Introduction to Computer Science and Programming. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, second edition, 2001. ISBN 0-13-031697-0. xxxiv + 1049 pp. LCCN QA76.73.J38 S27 2001. US$68.00. URL

http://cw.prenhall.com/
bookbind/pubbooks/savitch;
 http://www.phptr.com/ptrbooks/
esm_0130316970.html. In-

cludes CD-ROM with text’s programs, icons and TextPad. Study Guide and Slides available on the World Wide Web.

Sekkaki:2001:DAM

[SAWW01] A. Sekkaki, L. M. Cáceres Alvarez, W. Tatsuya Watanabe, and B. Westphall. Development of accounting management based service environment in Tina, Java and Corba architectures. Lecture Notes in Computer Science, 2094:438–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.
                                        com/link/service/series/
                                        0558/bibs/2094/20940438.
                                        htm;   http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2094/20940438.
                                        pdf.

Sirer:2000:UPG

[SB00] Emin Guün Sirer and Brian N. Bershad. Using production grammars in software testing. ACM SIGPLAN Notices, 35(1):1–13, 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sierra:2003:HFE

[SB03a] Kathy Sierra and Bert Bates. Head first EJB. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2003. ISBN 0-596-00571-7. xxx + 700 pp. LCCN QA76.3 .S468 2003. URL http://

                                        www.oreilly.com/catalog/
                                        9780596005719.

Sierra:2003:HFJ

[SB03b] Kathy Sierra and Bert Bates. Head first Java. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2003. ISBN 0-596-00465-6. xxx + 619 pp. LCCN QA76.73.J38 S535 2003. URL http://

                                        www.oreilly.com/catalog/
                                        9780596004651.

Sierra:2005:HFJ

[SB05] Kathy Sierra and Bert Bates.

Head first Java. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, second edition, 2005. ISBN 0-596-00920-8 (paperback). xxxii + 688 pp. LCCN QA76.73.J38 S535 2005.

Sam-Bodden:2006:BPN

[SB06a] Brian Sam-Bodden. Beginning POJOs: from novice to professional. Apress, Berkeley, CA, USA, 2006. ISBN 1-59059-596-3 (paperback). LCCN QA76.73.J38 S36 2006eb; QA76.73 .J38 INTERNET. URL http:

//ezproxy.lib.ucalgary.
ca:2048/login?url=http:
//library.books24x7.com/
library.asp?bookid=14653.

Sridharan:2006:RBC

[SB06b] Manu Sridharan and Rastislav Bodík. Refinement-based context-sensitive points-to analysis for Java. ACM SIG-PLAN Notices, 41(6):387– 400, June 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shankar:2007:DAI

[SB07] Ajeet Shankar and Rastislav Bodík. DITTO: automatic incrementalization of data structure invariant checks (in Java). ACM SIGPLAN Notices, 42(6):310–319, June 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Stuer:2001:PSA

[SBA01] Gunther Stuer, Jan Broeckhove, and Frans Arickx. Performance and stability analysis of a message oriented reliable multicast for distributed virtual environments in Java. Lecture Notes in Computer Science, 2073:423– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2073/20730423.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2073/20730423.
pdf.

Saleh:2001:ADC

[SBAD01] K. Saleh, A. A. Boujarwah, and J. Al-Dallal. Anomaly detection in concurrent Java programs using dynamic data flow analysis. Information and Software Technology, 43 (15):973–981, December 23, 2001. CODEN ISOTE7. ISSN 0950-5849 (print), 1873-6025 (electronic).

Schuppan:2005:JIR

[SBB05] V. Schuppan, M. Baur, and A. Biere. JVM independent replay in Java. Electronic Notes in Theoretical Computer Science, 113(??):

85–104, 2005. CODEN ???? ISSN 1571-0661.

Schultz:2003:CJL

[SBCK03] Ulrik Pagh Schultz, Kim Burgaard, Flemming Gram Christensen, and Jørgen Lindskov Knudsen. Compiling Java for low-end embedded systems. ACM SIGPLAN Notices, 38(7):42–50, July 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Syropoulos:2004:TXD

[SBH^{+}04] Apostolos Syropoulos, Karl Berry, Yannis Haralambous, Baden Hughes, Steven Peter, and John Plaice, editors. TEX, XML, and Digital Typography: International Conference on TEX, XML, and Digital Typography, held jointly with the 25th Annual Meeting of the TEX Users Group, TUG 2004, Xanthi, Greece, August 30–September 3, 2004: Proceedings, volume 3130 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. CODEN LNCSD9. ISBN 3-540-22801-2 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Z253.3 I58 2004. URL

http://link.springer-ny.
com/link/service/series/
0558/tocs/t3130.htm;   http:
//www.springerlink.com/

                                        openurl.asp?genre=issue&
                                        issn=0302-9743&volume=3130;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=volume&
                                        id=doi:10.1007/b99374.

Serrano:2000:QQS

[SBMG00] Mauricio Serrano, Rajesh Bordawekar, Sam Midkiff, and Manish Gupta. Quicksilver: a quasi-static compiler for Java. ACM SIGPLAN Notices, 35(10):66–82, October 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        proceedings/oops/353171/
                                        p66-serrano/.

Smith:2001:PJG

[SBO01] L. A. Smith, J. M. Bull, and J. Obdrzálek. A parallel Java Grande benchmark suite. In ACM [ACM01c], page ?? ISBN 1-58113-293-X. LCCN ???? URL http:

                                        //www.sc2001.org/papers/
                                        pap.pap158.pdf.

Sanchez:2001:JWC

[SC01a] Julio Sanchez and Maria P. Canton. Java 2 weekend crash course. I D G Books Worldwide, Indianapolis, IN, USA, 2001. ISBN 0-7645-4768-2. xxiv + 427 pp. LCCN QA76.73.J38 S25 2001.

Strohmeier:2001:SSC

[SC01b] Alfred Strohmeier and Stanislav Chachkov. A side-by-side comparison of exception handling in Ada and Java. ACM SIGADA Ada Letters, 21(3): 41–56, September 2001. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Sanchez:2002:JPE

[SC02a] Julio Sanchez and Maria P.

Canton. Java Programming for Engineers. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2002. ISBN 0-8493-0810-0. xvi + 338 pp. LCCN QA76.76.J38 S26 2002. US$59.95, UK£41.99.

Skotiniotis:2002:EIM

[SC02b] T. Skotiniotis and J. Chang.

Estimating internal memory fragmentation for Java programs. The Journal of systems and software, 64(3):235– 246, 2002. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Sotomayor:2005:GTP

[SC05] Borja Sotomayor and Lisa Childers. Globus Toolkit 4: Programming Java Services. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2005. ISBN 0-12-369404-3. 320 (est.) pp. LCCN ???? US$39.95.

Sasitorn:2007:CNS

[SC07] James Sasitorn and Robert Cartwright. Component NEXTGEN: a sound and expressive component framework for Java. ACM SIG-PLAN Notices, 42(10):153– 170, October 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Smith:2008:JTI

[SC08] Daniel Smith and Robert Cartwright. Java type inference is broken: can we fix it? ACM SIGPLAN Notices, 43(10):505–524, September 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shafi:2009:NPM

[SCB09] Aamir Shafi, Bryan Carpenter, and Mark Baker. Nested parallelism for multicore HPC systems using Java. Journal of Parallel and Distributed Computing, 69(6): 532–545, June 2009. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Shafi:2009:CSJ

[SCBH09] Aamir Shafi, Bryan Carpenter, Mark Baker, and Aftab Hussain. A comparative study of Java and C performance in two large-scale parallel applications. Concurrency and Computation: Practice and Experience, 21(15):1882–1906, October 2009. CODEN CCPEBO.

ISSN 1532-0626 (print), 1532-0634 (electronic).

Shi:2008:VMS

[SCEG08] Yunhe Shi, Kevin Casey, M. Anton Ertl, and David Gregg. Virtual machine showdown: Stack versus registers. ACM Transactions on Architecture and Code Optimization, 4(4):2:1–2:??, January 2008. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Steven:2000:JCR

[SCFP00] John Steven, Pravir Chandra, Bob Fleck, and Andy Podgurski. jRapture: a capture/replay tool for observationbased testing. ACM SIG-SOFT Software Engineering Notes, 25(5):158–167, September 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Schaub:2000:TJG

[Sch00a] Stephen Schaub. Teaching Java with graphics in CS1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32 (2):71–73, June 2000. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Schussler:2000:BPS

[Sch00b] H. Schussler. BAPI Programming for SAP R/3 with Java. Addison-Wesley, Reading, MA, USA, 2000. ISBN 0-201-67474-2. ???? pp. LCCN ????

Schildt:2001:JCR

[Sch01]  Herbert Schildt.  Java

2: the complete reference. Osborne/McGraw-Hill, Berkeley, CA, USA, fourth edition, 2001. ISBN 0-07-213084-9. xxvi + 1077 pp. LCCN QA76.73.J38 S3283 2001.

Schreiner:2002:JTT

[Sch02] W. Schreiner. A Java toolkit for teaching distributed algorithms. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):111–115, 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Schilling:2003:SHM

[Sch03a] Jonathan L. Schilling. The simplest heuristics may be the best in Java JIT compilers. ACM SIGPLAN Notices, 38(2):36–46, February 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schmid:2003:UEJ

[Sch03b]  H. A. Schmid.  On the

use of Enterprise Java Beans 2.0 local interfaces. Lecture Notes in Computer Science, 2596:144–156, 2003. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schoeberl:2003:JJO

[Sch03c] M. Schoeberl. JOP: a Java optimized processor. Lecture Notes in Computer Science, 2889:346–359, 2003. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schirmer:2004:AJP

[Sch04a] Norbert Schirmer. Analysing the Java package/access concepts in Isabelle/HOL. Concurrency and Computation: Practice and Experience, 16 (7):689–706, June 2004. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Schoeberl:2004:JTF

[Sch04b] M. Schoeberl. Java technology in an FPGA. Lecture Notes in Computer Science, 3203:917–921, 2004. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schoeberl:2004:TPI

[Sch04c] M. Schoeberl. A time predictable instruction cache for a Java processor. Lecture Notes in Computer Science, 3292:371–382, 2004. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schrijvers:2004:JGJ

[Sch04d] T. Schrijvers. JmmSolve: a generative Java memory model implemented in Prolog and CHR. Lecture Notes in Computer Science, 3132:475– 476, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Su:2005:CBJ

[SCH05] D. Su, Z. Chen, and L. Huo. Communication between Java and other advance language based on JNI. Journal — Guangxi University Natural Science Edition, 30(1):6–9, 2005. CODEN ???? ISSN 1001-7445.

Sciore:2007:SSJ

[Sci07] Edward Sciore. SimpleDB: a simple Java-based multiuser syst for teaching database internals. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):561–565, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sheard:2008:GSA

[SCL^{+}08] Judy Sheard, Angela Carbone, Raymond Lister, Beth Simon, Errol Thompson, and Jacqueline L. Whalley. Going SOLO to assess novice programmers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):209–213, September 2008. CODEN SIGSD3.

ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Stahl:2004:DTD

[SCLV04] R. Stahl, F. Catthoor, R. Lauwereins, and D. Verkest. Design-time data-access analysis for parallel Java programs with shared-memory communication model. Lecture Notes in Computer Science, 3149:206–213, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Scott:2002:MMI

[Sco02] Kirk Scott. MISC: the minimal instruction set computer. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34 (3):223, September 2002. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Scott:2003:TGI

[Sco03] K. Scott. Teaching graphical interface programming in Java with the Game of Wari. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):254, 2003. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Shelly:2001:JPI

[SCS01] Gary B. Shelly, Thomas J.

Cashman, and Joy Starks.

Java programming: introductory concepts and techniques. Shelly Cashman series. Course Technology, Cambridge, MA, USA, 2001. ISBN 0-7895-5959-5 (paperback). various pp. LCCN QA76.73.J38 S49 2001.

Su:2008:SOE

[SCWL08] Wei Su, Chuan Cai, Paul S. Wang, and Lian Li. A solution for online entering and editing mathematical formulas (abstract only). ACM Communications in Computer Algebra, 42(1–2):87–88, March/June 2008. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Sarkar:2001:HPS

[SD01a] Vivek Sarkar and Julian Dolby. High-performance scalable Java virtual machines. Lecture Notes in Computer Science, 2228:151–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2228/22280151.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2228/22280151.
pdf.

Seymour:2001:ATF

[SD01b] Keith Seymour and Jack Dongarra. Automatic translation of Fortran to JVM bytecode. In ACM [ACM01b], pages 126–133. ISBN 1-58113-359-

6. LCCN QA76.9.O35 A26

2001. URL http://www.
netlib.org/utk/people/JackDongarra/
PAPERS/f2jreport.pdf;
http://www.philippsen.com/
JGI2001/camerareadyabstracts/
51.html;   http://www.philippsen.
com/JGI2001/finalpapers/
18500126.ps.

Sanders:2003:JTI

[SD03a] Dean Sanders and Brian Dorn. Jeroo: a tool for introducing object-oriented programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):201–204, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Seymour:2003:ATF

[SD03b] Keith Seymour and Jack Dongarra. Automatic translation of Fortran to JVM bytecode. Concurrency and Computation: Practice and Experience, 15(3–5):207–222, March/April 2003. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL http:

//www.netlib.org/netlib/
utk/people/JackDongarra/
PAPERS/f2jreport.pdf.

Sun:2004:JBA

[SD04] H. Sun and R. V. Davuluri. Java-based application framework for visualization of gene regulatory region annotations. Bioinformatics, 20 (5):727–734, March 22, 2004. CODEN ???? ISSN 1367-4803 (print), 1367-4811 (electronic).

Schonberg:2008:PAS

[SD08] E. Schonberg and R. Dewar. A principled approach to software engineering education, or Java considered harmful. Ada User Journal, 29(3):200– ??, September 2008. CODEN AUJOET. ISSN 1381-6551.

Schmietendorf:2000:MBA

[SDF00] Andreas Schmietendorf, Reiner Dumke, and Erik Foltin. Metrics based asset assessment. ACM SIGSOFT Software Engineering Notes, 25(4):51–55, July 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sanchez:2004:JMB

[SDPM04] J. Sanchez, S. Dormido, R. Pastor, and F. Morilla. A Java/ Matlab-based environment for remote control system laboratories: Illustrated with an inverted pendulum. IEEE transactions on education, 47(3):321–329, 2004. CODEN IEEDAB. ISSN 0018-9359.

Sweedyk:2005:CGC

[SdSK05] Elizabeth Sweedyk, Marianne deLaet, Michael C. Slattery, and James Kuffner. Computer games and CS education: why and how. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):256– 257, March 2005. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Selcuk:2004:JEJ

[SE04] Y. E. Selcuk and N. Erdogan. JAWIRO: Enhancing Java with roles. Lecture Notes in Computer Science, 3280:927–934, 2004. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Seaman:2002:JQH

[Sea02] Mark Seaman. Java Q&A:

How can I generate Java code for tables in my database? Dr. Dobb’s Journal of Software Tools, 27(8):??, August 2002. CODEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/ftp/2002/
2002_08/jqa0802.txt;   http:
//www.ddj.com/ftp/2002/
2002_08/jqa0802.zip.

Sedgewick:2003:AJ

[Sed03] Robert Sedgewick. Algorithms in Java. Addison-Wes-ley, Reading, MA, USA, third edition, 2003. ISBN 0-201-36120-5. 737 pp. LCCN QA76.73.J38 S4 2003. Parts 1-4.

Schafer:2008:SER

[SEdM08] Max Schäfer, Torbjörn Ekman, and Oege de Moor.

Sound and extensible renaming for Java. ACM SIGPLAN Notices, 43(10):277–294, September 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Seegmiller:2004:PRO

[See04] R. Seegmiller. Product review: Optimal Java: a review of Compuware’s OptimalJ 3.1. Application Development Trends, 11(3):48–49, 2004. CODEN ???? ISSN 1073-9564.

Shirmohammadi:2003:JJT

[SEGS03] S. Shirmohammadi, A. El-Saddik, N. D. Georganas, and R. Steinmetz. JASMINE: a Java tool for multimedia collaboration on the Internet. Multimedia Tools and Applications, 19(1):5–28, 2003. CODEN MTAPFB. ISSN 1380-7501.

Seidman:2009:AFI

[Sei09] Robert H. Seidman. Alice first: 3D interactive game programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3): 345, September 2009. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Sellin:2003:MAJ

[Sel03]  R. Sellin.  Mobile At-

traktivität mit Java-Games.

Comtec, 81(1):14–19, 2003. CODEN ???? ISSN 1420-3715.

Sen:2008:RDR

[Sen08] Koushik Sen. Race directed random testing of concurrent programs. ACM SIGPLAN Notices, 43(6):11–21, June 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sestak:2000:JPP

[Ses00]  John Sestak.  Java Pro-

gramming Projects. South-Western Educational Publishing, Cincinnati, OH, USA, 2000. ISBN 0-538-69404-1. ???? pp. LCCN QA76.73.J38 S43 2001. US$20.

Sestoft:2002:JP

[Ses02] Peter Sestoft. Java Precisely. MIT Press, Cambridge, MA, USA, 2002. ISBN 0-262-69276-7. vii + 118 pp. LCCN QA76.73.J38 S435 2002. US$14.95, UK£9.95.

Sestoft:2005:JP

[Ses05] Peter Sestoft. Java precisely.

MIT Press, Cambridge, MA, USA, second edition, 2005. ISBN 0-262-69325-9 (paperback). 168 (est.) pp. LCCN QA76.73.J38 S435 2005.

Sestoft:2008:PLC

[Ses08] Peter Sestoft. Programming language concepts for software developers. ACM SIG-PLAN Notices, 43(11):119– 123, November 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Setzer:2003:JFP

[Set03] A. Setzer. Java as a functional programming language. Lecture Notes in Computer Science, 2646:279–298, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sarkar:2001:EDA

[SF01] Vivek Sarkar and Stephen Fink. Efficient dependence analysis for Java arrays. Lecture Notes in Computer Science, 2150:273–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2150/21500273.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2150/21500273.
pdf.

Sridharan:2007:TS

[SFB07] Manu Sridharan, Stephen J. Fink, and Rastislav Bodik. Thin slicing. ACM SIGPLAN Notices, 42(6):112–122, June 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Simon:2007:DAN

[SFM^{+}07] Beth Simon, Sue Fitzgerald, Renée McCauley, Susan M. Haller, John Hamer, Brian Hanks, Michael T. Helmick, Jan Erik Moström, Judy Sheard, and Lynda Thomas. Debugging assistance for novices: a video repository. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(4):137–151, December 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp:

//ftp.math.utah.edu/pub/
mirrors/ftp.ira.uka.de/
bibliography/Misc/DBLP/
2007.bib.

Shah:2001:JSD

[SFMH01] Mehul A. Shah, Michael J.

Franklin, Samuel Madden, and Joseph M. Hellerstein. Java support for dataintensive systems: experiences building the telegraph dataflow system. SIG-MOD, 30(4):103–114, December 2001.

Sivaram:2003:XJO

[SFP03] A. Sivaram, D. Fan, and J. Pryce. XML and Java for Open ATE programming environment. In International Test Conference, pages 793– 801. ????, ????, 2003. CO-DEN ???? ISSN 0743-1686.

Schneider:2000:ICS

[SG00] G. Michael Schneider and Judith Gersting. Invitation to Computer Science: Java Version. Brooks/Cole, Pacific Grove, CA, USA, 2000. ISBN 0-534-37488-3. xxvii + 734 pp. LCCN QA76. .S3596 2000. US$56. Contributing author: Sara Baase.

Shen:2002:JBD

[SG02] J. Shen and G. Gu. Javabased design and implementation of the XML parser. Mini-Micro Systems, 23(12): 1449–1452, 2002. CODEN XWJXEH. ISSN 1000-1220.

Sunkpho:2003:JIF

[SG03] J. Sunkpho and J. H. Garrett. Java inspection framework: Developing field inspection support systems for civil systems inspection. Journal of Computing in Civil Engineering, 17(4):209–218, 2003. CO-DEN JCCEE5. ISSN 0887-3801.

Shuf:2002:CPL

[SGF^{+}02] Yefim Shuf, Manish Gupta, Hubertus Franke, Andrew Appel, and Jaswinder Pal Singh. Creating and preserving locality of Java applications at allocation and garbage collection times. ACM SIGPLAN Notices, 37 (11):13–25, November 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sharma:2009:DAC

[SGK09] Arun Sharma, P. S. Grover, and Rajesh Kumar. Dependency analysis for componentbased software systems. ACM SIGSOFT Software Engineering Notes, 34(4):1–6, July 2009. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sridharan:2005:DDP

[SGSB05] Manu Sridharan, Denis Gopan, Lexin Shan, and Rastislav Bodík. Demand-driven points-to analysis for Java. ACM SIGPLAN Notices, 40 (10):59–76, October 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sage:2004:JTS

[SGV04] S. Sage, G. Grandjean, and J. Verly. Java Tomography System (JaTS), a seismic tomography software using Fresnel volumes, a fast marching eikonal solver and a probabilistic reconstruction method: Conclusive synthetic test cases. Lecture Notes in Earth Sciences, 104:226–235, 2004. CODEN ???? ISSN 0930-0317.

Shegalov:2001:XEW

[SGW01] German Shegalov, Michael Gillmann, and Gerhard Weikum. XML-enabled workflow management for eservices across heterogeneous platforms. VLDB Journal: Very Large Data Bases, 10(1):91–103, August 2001. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic). URL http:

//link.springer.de/link/
service/journals/00778/
bibs/1010001/10100091.htm;
 http://link.springer.de/
link/service/journals/00778/
papers/1010001/10100091.
pdf.

Saiedian:2003:CEG

[SH03] H. Saiedian and S. Hill. A comparative evaluation of generic programming in Java and C++. Software —Practice and Experience, 33(2):121–142, 2003. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Schmalenbach:2004:JVM

[SH04a] C. Schmalenbach and C. Hofig. The Java Virtual Machine profiler interface. Dr. Dobb’s Journal of Software Tools, 29 (7):28–33, 2004. CODEN DDJOEB. ISSN 1044-789X.

Snook:2004:ECC

[SH04b] C. F. Snook and R. Harrison. Experimental comparison of the comprehensibility of a Z specification and its implementation in Java. Information and Software Technology, 46(14):955–971, 2004. CODEN ISOTE7. ISSN 0950-5849 (print), 1873-6025 (electronic).

Subramaniam:2006:PAD

[SH06] Venkat Subramaniam and Andy Hunt. Practices of an Agile Developer: Working in the Real World. Pragmatic Bookshelf, Raleigh, NC, USA, 2006. ISBN 0-9745140-8-X. xiv + 189 pp. LCCN QA76.76.D47. US$29.95; CAN $41.95; UK£20.95.

Shankari:2000:HCN

[Sha00a] K. Shankari. How to connect non-Java devices to a Jini network? Thesis (M.S.), University of California, Santa Cruz, Santa Cruz, CA, USA, 2000.

Shannon:2000:JPE

[Sha00b] Bill Shannon. Java 2 platform, enterprise edition: platform and component specifications. Java series. Addison-Wesley, Reading, MA, USA, 2000. ISBN 0-201-70456-0. xxxiv + 748 pp. LCCN QA76.73.J38 J35945 2000.

Shaofeng:2001:RJR

[Sha01] Wang Shaofeng. The role of Java RMI in designing workflow management system. ACM SIGSOFT Software Engineering Notes, 26 (2):49–52, March 2001. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Shay:2002:MMC

[Sha02] William Shay. A multiplatform/multilanguage client/ server project. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1): 401–405, March 2002. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Shaofeng:2004:MJB

[Sha04] Wang Shaofeng. A method of Java-based electronic document publishing system. Electronic Library, 22(4):351–356, April 1, 2004. CODEN EL-LIDZ. ISSN 0264-0473.

Stefanovic:2003:OFG

[SHB^{+}03] Darko Stefanović, Matthew Hertz, Stephen M. Blackburn, Kathryn S. McKinley, and J. Eliot B. Moss. Older-first garbage collection in practice: Evaluation in a Java virtual machine. ACM SIGPLAN Notices, 38(2s):25–36, February 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shelly:2001:JCC

[She01a] Gary B. Shelly. JavaScript: complete concepts and techniques. Shelly Cashman series. Course Technology, Cambridge, MA, USA, second edition, 2001. ISBN 0-7895-6233-2. various pp. LCCN QA76.73.J38 S457 2001.

Sheong:2001:BDF

[She01b] Chang Sau Sheong. Building dynamic fail-over Java servers. Dr. Dobb’s Journal of Software Tools, 26(11):46, 48, 50, 52–54, 57, November 2001. CODEN DDJOEB. ISSN 1044-789X. URL

http://www.ddj.com/ftp/
2001/2001_11/xander.txt;
http://www.ddj.com/ftp/
2001/2001_11/xander.zip.

Sherer:2003:RTS

[She03]  Matt Sherer.  Real-time

signal analysis and Real-Time Linux: Part 1. Dr. Dobb’s Journal of Software Tools, 28(7):62–65, July 2003. CODEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/ftp/2003/
2003_07/rtlp1.txt.

Steeb:2004:PSS

[SHHS04] W.-H. Steeb, Yorick Hardy, Alexandre Hardy, and Ruedi Stoop. Problems and solutions in scientific computing: with C++ and Java simulations. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2004. ISBN 981-256-112-9, 981-256-125-0 (paperback). ix + 420 pp. LCCN QA297 .P743 2004.

Shirazi:2000:JPT

[Shi00] Jack Shirazi. Java performance tuning. Java series. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2000. ISBN 0-596-00015-4. xiv + 426 pp. LCCN QA76.73.J38 S47 2000.

Shippy:2003:PGT

[Shi03a] P. J. Shippy. Porting the Gnat tasking runtime system to the Java Virtual Machine. Literary Review, 47(2):119– 120, 2003. CODEN ???? ISSN 0024-4589.

Shirazi:2003:JPT

[Shi03b] Jack Shirazi. Java performance tuning. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2003. ISBN 0-596-00377-3. xv + 570 pp. LCCN QA76.73.J38 S47 2003.

Steinbeck:2003:CDK

[SHK^{+}03] C. Steinbeck, Y. Han, S. Kuhn, O. Horlacher, E. Luttmann, and E. Willighagen. The Chemistry Development Kit (CDK): An open-source Java library for chemo- and bioinformatics. Journal of Chemical Information and Computer Sciences, 43(2):493– 500, 2003. CODEN ???? ISSN 0095-2338.

Subramanian:2009:DSU

[SHM09] Suriya Subramanian, Michael Hicks, and Kathryn S. McKinley. Dynamic software updates: a VM-centric approach. ACM SIGPLAN Notices, 44(6):1–12, June 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sundaresan:2000:PVM

[SHR^{+}00] Vijay Sundaresan, Laurie Hendren, Chrislain Razafimahefa, Raja Vallée-Rai, Patrick Lam, Etienne Gagnon, and Charles Godin. Practical virtual method call resolution for Java. ACM SIGPLAN Notices, 35(10):264–280, October 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
proceedings/oops/353171/
p264-sundaresan/.

Saito:2009:STC

[SI09] Chieri Saito and Atsushi Igarashi. Self type constructors. ACM SIGPLAN Notices, 44(10):263–282, October 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Siberz:2000:CCJ

[Sib00] Ken Siberz. From C/C++ to Java: fast transition: July 19–21, 2000, Engineering 819.315. Los Angeles, CA, USA, 2000. various pp. Cover title.

Sigg:2004:MDJ

[Sig04] R. Sigg. Mobile Dienste mit Java realisieren. (German)

[mobile services with a Java implementation]. Comtec, 82 (2):28–30, 2004. CODEN ???? ISSN 1420-3715.

Sigglekow:2005:JSC

[Sig05] Bill Sigglekow. Jakarta Struts cookbook. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00771-X. xvi + 507 pp. LCCN QA76.73.J38 S445 2005.

Sikora:2003:JPG

[Sik03] Zbigniew M. Sikora. Java: practical guide for programmers. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2003. ISBN 1-55860-909-1. xii + 179 pp. LCCN QA76.73.J38 S54 2003.

Simmons:2004:HJ

[Sim04a] Robert Simmons, Jr. Hardcore Java. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2004. ISBN 0-596-00568-7. xvii + 324 pp. LCCN QA76.73.J38 S553 2004. URL http://

www.oreilly.com/catalog/
9780596005689.

Simmons:2004:HJS

[Sim04b] Robert Simmons, Jr. Hardcore Java: Secrets of the Java masters. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2004. ISBN 0-596-00568-7. xvii + 324 pp. LCCN QA76.73.J38 S553 2004.

Sintes:2000:XSC

[Sin00]  Tony Sintes.  XML and

software configuration. Dr. Dobb’s Journal of Software Tools, 25(7):56, 58–62, July 2000. CODEN DDJOEB. ISSN 1044-789X. URL

http://www.ddj.com/ftp/
2000/2000_07/xmlcfg.txt;
http://www.ddj.com/ftp/
2000/2000_07/xmlcfg.zip.

Sivasubramanian:2002:JCM

[Siv02] Madhumathi Sivasubramanian. Java compiler modification for multiple return types. Thesis (m.s.), University of California, Santa Cruz, Santa Cruz, CA, USA, 2002.

Siveroni:2004:OSJ

[Siv04] I. A. Siveroni. Operational semantics of the Java Card Virtual Machine. Journal of Logic and Algebraic Programming, 58(1–2):3–25, 2004. CODEN ???? ISSN 1567-8326.

Shaofeng:2001:FDW

[SJ01] Wang Shaofeng and Sun Jiaguang. A framework design of workflow management system with Java RMI. ACM SIGPLAN Notices, 36(9):86– 93, September 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sucurovic:2005:JCX

[SJ05] S. Sucurovic and Z. Jovanovic. Java cryptography & X.509 authentication. Dr. Dobb’s Journal of Software Tools, 30(2):40–42, 2005. CO-DEN DDJOEB. ISSN 1044-789X.

Saraswat:2003:JIT

[SJG03] V. Saraswat, R. Jagadeesan, and V. Gupta. jcc: Integrating timed default concurrent constraint programming into Java. Lecture Notes in Computer Science, 2902:156– 170, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shelekhov:2000:DFA

[SK00] Vladimir I. Shelekhov and Sergey V. Kuksenko. Data flow analysis of Java programs in the presence of exceptions. Lecture Notes in Computer Science, 1755:389– ??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/1755/17550389.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/1755/17550389.
pdf.

Shimizu:2004:JOL

[SK04] Naohiko Shimizu and Chiaki Kon. Java object look aside buffer for embedded applications. ACM SIGARCH Computer Architecture News, 32 (3):43–49, June 2004. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Singer:2008:DAJ

[SK08] Jeremy Singer and Chris Kirkham. Dynamic analysis of Java program concepts for visualization and profiling. Science of Computer Programming, 70(2–3):111– 126, February 1, 2008. CO-DEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Skansholm:2000:JB

[Ska00] Jan Skansholm. Java from the beginning. Addison-Wes-ley, Reading, MA, USA, 2000. ISBN 0-201-39812-5 (paperback). xiv + 540 pp. LCCN QA76.73.J38 S593 2000.

Schwarz:2009:DFP

[SKC09] E. M. Schwarz, J. S. Kapernick, and M. F. Cowlishaw. Decimal floating-point support on the IBM System z10 processor. IBM Journal of Research and Development, 53(1):4:1–4:10, January/February 2009. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http:

//www.research.ibm.com/
journal/rd/531/schwarz.
pdf.

Skinner:2007:UA

[Ski07] Jesse Skinner. Unobtrusive Ajax. O’Reilly shortcuts. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2007. ISBN 0-596-51024-1. LCCN TK5105.8885.A52. URL

http://www.oreilly.com/
catalog/9780596510244.

Systa:2001:SER

[SKM01] Tarja Systä, Kai Koskimies, and Hausi Müller. Shimba

— an environment for reverse engineering Java software systems. Software—Practice and Experience, 31(4):371– 394, April 10, 2001. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (elec-

tronic). URL http://www3.
interscience.wiley.com/
cgi-bin/abstract/77004439/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=77004439&PLACEBO=IE.
pdf.

Sung:2002:CPE

[SKP^{+}02] Minyoung Sung, Soyoung Kim, Sangsoo Park, Naehyuck Chang, and Heonshik Shin. Comparative performance evaluation of Java threads for embedded applications: Linux Thread vs. Green Thread. Information Processing Letters, 84(4):221– 225, November 30, 2002. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shaham:2001:HPS

[SKS01a] Ran Shaham, Elliot K.

Kolodner, and Mooly Sagiv. Heap profiling for spaceefficient Java. ACM SIG-PLAN Notices, 36(5):104– 113, May 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shaham:2001:EGJ

[SKS01b] Ran Shaham, Elliot K.

Kolodner, and Mooly Sagiv. On the effectiveness of GC in Java. ACM SIGPLAN Notices, 36(1):12–17, January 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shaham:2003:EIH

[SKS03] Ran Shaham, Elliot K.

Kolodner, and Mooly Sagiv. Estimating the impact of heap liveness information on space consumption in Java. ACM SIGPLAN Notices, 38(2s):171–182, February 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Stubblebine:2008:RAK

[SKS08] Toni Stubblebine, Peter Klicman, and Lars Schulten. Reguläre Ausdrücke — kurz and gut [für Perl, Ruby, PHP, C#, Python, Java and .NET]. O’Reillys Taschenbibliothek. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2008. ISBN 3-89721-535-7. 131 pp. LCCN ???? EUR 9.90.

Sterbenz:2000:PAC

[SL00] Andreas Sterbenz and Peter Lipp. Performance of the AES candidate algorithms in Java. In NIST [NIS00], pages 161–

168. ISBN ???? LCCN

???? URL http://csrc.
nist.gov/encryption/aes/
round2/conf3/aes3conf.htm;
 http://csrc.nist.gov/encryption/
aes/round2/conf3/papers/
AES3Proceedings-1.pdf;
http://csrc.nist.gov/encryption/
aes/round2/conf3/papers/
AES3Proceedings-2.pdf;
http://csrc.nist.gov/encryption/
aes/round2/conf3/papers/
AES3Proceedings-3.pdf;
http://csrc.nist.gov/encryption/
aes/round2/conf3/papers/
AES3Proceedings.pdf.

Stoller:2001:TMC

[SL01] Scott D. Stoller and Yanhong A. Liu. Transformations for model checking distributed Java programs. Lecture Notes in Computer Science, 2057:192–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2057/20570192.
htm;   http://link.springer-
ny.com/link/service/series/

0558/papers/2057/20570192.
pdf.

Sung:2004:JBC

[SL04] M. Y. Sung and D. H. Lee.

A Java-based collaborative authoring system for multimedia presentation. Lecture Notes in Computer Science, 3332:96–103, 2004. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sattar:2006:DSM

[SL06] Abdul Sattar and Torben Lorenzen. Develop a shopping mart Web application. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38 (4):68–70, December 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp:

//ftp.math.utah.edu/pub/
mirrors/ftp.ira.uka.de/
bibliography/Misc/DBLP/
2006.bib.

Sattar:2007:DCJ

[SL07] Abdul Sattar and Torben Lorenzen. Develop a compiler in Java for a compiler design course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(2):80–82, June 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp:

//ftp.math.utah.edu/pub/
mirrors/ftp.ira.uka.de/

bibliography/Misc/DBLP/
2007.bib.

Slack:2000:PPS

[Sla00] James M. (James Michael) Slack. Programming and problem solving with Java. Brooks/Cole, Pacific Grove, CA, USA, 2000. ISBN 0-534-37486-7. xii + 1137 pp. LCCN QA76.73.J38 S595 2000. US$61.

Schneck:2002:LCP

[SLB^{+}02] Paul Schneck, Jeffrey Lotspiech, Christopher Bird, Jeffrey R. Carter, Eric A. Weiss, Neville Holmes, and Mario L. Epp. Letters: Content protection alternatives; software development viewpoints; Java questions; seven great blunders; technological musings. Computer, 35(10): 8–??, October 2002. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/co/
2002/10/rx008.htm;   http:
//csdl.computer.org/dl/
mags/co/2002/10/rx008.pdf.

Schultz:2003:APS

[SLC03a] Ulrik P. Schultz, Julia L. Lawall, and Charles Consel. Automatic program specialization for Java. ACM Transactions on Programming Languages and Systems, 25(4): 452–499, July 2003. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Srisaan:2003:AMP

[SLC03b] W. Srisaan, C. T. D. Lo, and J. M. Chang. Active memory processor: a hardware garbage collector for real-time Java embedded devices. IEEE Transactions on Mobile Computing, 2(2):89– 101, 2003. CODEN ???? ISSN 1536-1233.

Sanchez:2002:FTU

[SLPO02] Pedro Sánchez, Patricio Letelier, Juan A. Pastor, and Juan A. Ortega. A framework to translate UML class generalization into Java code. Lecture Notes in Computer Science, 2425:173–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2425/24250173.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2425/24250173.
pdf.

Scherer:2009:SSQ

[SLS09] William N. Scherer, III, Doug Lea, and Michael L. Scott. Scalable synchronous queues. Communications of the ACM, 52(5):100–111, May 2009. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sanchez:2001:BWA

[SM01a] Miguel Sánchez and Pietro Manzoni. Best of Websim99: ANEJOS: a Java based simulator for ad hoc networks. Future Generation Computer Systems, 17(5): 573–583, March 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/45/30/31/abstract.
html.

Shende:2001:IAT

[SM01b] Sameer Shende and Allen Malony. Integration and application of the TAU performance system in parallel Java environments. In ACM [ACM01b], pages 87–

96. ISBN 1-58113-359-

6. LCCN QA76.9.O35 A26

2001. URL http://www.
philippsen.com/JGI2001/
camerareadyabstracts/39.
html;   http://www.philippsen.
com/JGI2001/finalpapers/
18500087.ps.

Shudo:2001:AME

[SM01c] Kazuyuki Shudo and Yoichi Muraoka. Asynchronous migration of execution context in Java Virtual Machines. Future Generation Computer Systems, 18(2): 225–233, October 2001. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
19/19/60/31/30/abstract.
html.

Stanchfield:2001:EVJ

[SM01d] Scott Stanchfield and Isabelle Mauny. Effective VisualAge for Java, version 3: includes coverage of versions 3.02 and 3.5. Wiley Computer Publishers, New York, NY, USA, 2001. ISBN 0-471-31730-6 (paperback). xxi + 936 pp. LCCN QA76.73.J38 S76 2001.

Stelting:2002:AJP

[SM02a] Stephen Stelting and Olav Maassen. Applied Java Patterns. Sun Microsystems Press Java series. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2002. ISBN 0-13-093538-7. xxi + 574 pp. LCCN QA76.73.J38 S845 2002. US$44.99. URL http://www.

phptr.com/ptrbooks/ptr_
0130935387.html;   http:
//www.sun.com/books/catalog/
stelting/.

Surdeanu:2002:DPA

[SM02b] Mihai Surdeanu and Dan Moldovan. Design and performance analysis of a distributed Java Virtual Machine. IEEE Transactions on Parallel and Distributed Systems, 13(6):611–627, June 2002. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http:

//dlib.computer.org/td/
books/td2002/pdf/l0611.
pdf;   http://www.computer.

                                        org/tpds/td2002/l0611abs.
                                        htm.

Shende:2003:IAT

[SM03a] Sameer Shende and Allen D. Malony. Integration and application of TAU in parallel Java environments. Concurrency and Computation: Practice and Experience, 15 (3–5):501–519, March/April 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Spain-McDuffie:2003:JCT

[SM03b] Tina Spain-McDuffie. JavaScript concepts and techniques: programming interactive Web sites. Franklin, Beedle and Associates, Wilsonville, OR, USA, 2003. ISBN 1-887902-69-4. xxii + 730 pp. LCCN QA73.73.J39 S73 2003 Computer. CD-ROM contains all scripts in accompanying book.

Schroder:2004:GEH

[SM04a] L. Schroder and T. Mossakowski. Generic exception handling and the Java monad. Lecture Notes in Computer Science, 3116:443–459, 2004. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stubblebine:2004:SHD

[SM04b] Tony Stubblebine and Junko Mishima. Seiki hyogen desukutoppu rifarensu: regular expressions for Perl, C, PHP, Python, Java, and .NET. Orairi Japan, Tokyo, Japan, 2004. ISBN 4-87311-170-6. vi + 96 pp. LCCN ????

Simos:2007:CMS

[SM07] Theodore E. Simos and George Maroulis, editors. Computation in Modern Science and Engineering: Proceedings of the International Conference on Computational Methods in Science and Engineering 2007 (IC-CMSE 2007), Corfu, Greece, 25–30 September 2007, volume 2A. American Institute of Physics, Woodbury, NY, USA, 2007. ISBN 0-7354-0476-3 (set), 0-7354-0477-1 (vol. 1), 0-7354-0478-X (vol. 2). LCCN Q183.9 2007. Two volumes.

Small:2007:DER

[Sma07] Margot Small. Design error and reusability. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(2):185–187, June 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp:

//ftp.math.utah.edu/pub/
mirrors/ftp.ira.uka.de/
bibliography/Misc/DBLP/
2007.bib.

Smart:2008:JPT

[Sma08] John Ferguson Smart. Java power tools. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2008. ISBN 0-596-52793-4. xxxv + 871 pp. LCCN QA76.73.J38 S63 2008.

Shpeisman:2007:EIO

[SMAT^{+}07] Tatiana Shpeisman, Vijay Menon, Ali-Reza Adl-Tabatabai, Steven Balensiefer, Dan Grossman, Richard L. Hudson, Katherine F. Moore, and Bratin Saha. Enforcing isolation and ordering in STM. ACM SIGPLAN Notices, 42(6):78–88, June 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Saougkos:2007:RJB

[SMBZ07] D. Saougkos, G. Manis, K. Blekas, and A. V. Zarras. Revisiting Java bytecode compression for embedded and mobile computing environments. IEEE Transactions on Software Engineering, 33(7):478–495, July 2007. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

                                        http://ieeexplore.ieee.
                                        org/stamp/stamp.jsp?arnumber=
                                        4227829.

Sadjadi:2004:TJT

[SMCS04] S. M. Sadjadi, P. K. McKinley, B. H. C. Cheng, and R. E. K. Stirewalt. TRAP/J: Transparent generation of adaptable Java programs. Lecture Notes in Computer Science, 3291:1243– 1261, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schneider:2001:APM

[SMES01] Daniel Schneider, Bernd Mathiske, Matthias Ernst, and Matthew Seidl. Automatic persistent memory management for the Spotless Java^{T M} Virtual Machine on the Palm Connected Organizer. In USENIX Association [USE01c], page ?? ISBN 1-880446-11-1. LCCN QA76.73.J38 J42 2001. URL

http://www.usenix.org/publications/
library/proceedings/jvm01/
schneider.html. Sponsored

by the USENIX Association.

Smiley:2001:LPJ

[Smi01a] John Smiley. Learn to program with Java. Osborne/McGraw-Hill, Berkeley, CA, USA, 2001. ISBN 0-07-213189-

6. xii + 608 pp. LCCN QA76.73.J38 S598 2002.

Smith:2001:JQH

[Smi01b] Terry Smith. Java Q&A:

How can I simplify Java data access objects. Dr. Dobb’s Journal of Software Tools, 26 (5):135–136, 138–139, May 2001. CODEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/ftp/2001/
2001_05/jqa0501.txt;   http:
//www.ddj.com/ftp/2001/
2001_05/jqa0501.zip.

S:2002:SPI

[SMK02] Venugopal K. S., Geetha Manjunath, and Venkatesh Krishnan. sEc: a portable interpreter optimizing technique for embedded Java Virtual Machine. In USENIX Association [USE02], page ?? ISBN 1-931971-01-3. LCCN QA76.73.J38 J42 2002. URL

http://www.usenix.org/publications/
library/proceedings/javavm02/
venugopal.html.

Schroeder:2006:VTO

[SML06] Will Schroeder, Ken Martin, and Bill Lorensen. The visualization toolkit: an objectoriented approach to 3D graphics [visualize data in 3D — medical, engineering or scientific; build your own applications with C++, Tcl, Java or Python; includes source code for VTK (supports UNIX, Windows and Mac)]. Kitware, Clifton Park, NY, fourth edition, 2006. ISBN 1-930934-19-X. xvi + 512 pp. LCCN ????

Silva:2000:HPC

[SMS00] Luís Moura Silva, Paulo Martins, and João Gabriel Silva. Heterogeneous parallel computing using Java and WMPI. Concurrency: Practice and Experience, 12(11):1077–1091, September 2000. CO-DEN CPEXEI. ISSN 1040-

3108. URL http://www3.
interscience.wiley.com/

cgi-bin/abstract/76000189/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=76000189&PLACEBO=IE.
pdf.

Sooriamurthi:2004:JET

[SMS^{+}04] R. Sooriamurthi, S. Menzel, A. Sengupta, K. Moor, S. Stamm, and K. Borner. Java engagement for teacher training: an experience report. Frontiers in Education Conference, CONF34 (VOL 1):T2D–17, 2004. CO-DEN PFECDR. ISSN 0190-5848.

Schneider:2008:DOE

[SMSAT08] Florian T. Schneider, Vijay Menon, Tatiana Shpeisman, and Ali-Reza Adl-Tabatabai. Dynamic optimization for efficient strong atomicity. ACM SIGPLAN Notices, 43(10):181–194, September 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shen:2009:SHP

[SMTZ09] Xipeng Shen, Feng Mao, Kai Tian, and Eddy Zheng Zhang. The study and handling of program inputs in the selection of garbage collectors. Operating Systems Review, 43 (3):48–61, July 2009. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Sewell:2007:OET

[SNO^{+}07] Peter Sewell, Francesco Zappa Nardelli, Scott Owens, Gilles Peskine, Thomas Ridge, Susmit Sarkar, and Rok Strniša. Ott: effective tool support for the working semanticist. ACM SIGPLAN Notices, 42(9):1–12, September 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sohda:2001:IPS

[SNOM01] Yukihiko Sohda, Hidemoto Nakada, Hirotaka Ogawa, and Satoshi Matsuoka. Implementation of portable software DSM in Java. In ACM [ACM01b], pages 163–

172. ISBN 1-58113-359-

6. LCCN QA76.9.O35 A26

                                        2001. URL http://www.
                                        philippsen.com/JGI2001/
                                        camerareadyabstracts/23.
                                        html;   http://www.philippsen.
                                        com/JGI2001/finalpapers/
                                        18500163.pdf.

Schildt:2000:JPR

[SO00] Herbert Schildt and Joseph O’Neil. Java 2 Programmer’s Reference. IDG Books, San Mateo, CA, USA, 2000. ISBN 0-07-212354-0. xv + 447 pp. LCCN QA76.73.J38 S332 2000. US$20.

Snoep:2002:JWS

[SO02] J. L. Snoep and B. G. Olivier. Java Web Simulation (JWS); A Web based database of kinetic models. Molecular Biology Reports, 29(1–2):259–263, 2002. CODEN ???? ISSN 0301-4851.

Sojka:2003:AP

[Soj03a]  Petr Sojka.  Animations

in PDF. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):263, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sojka:2003:ITM

[Soj03b] Petr Sojka. Interactive teaching materials in PDF using JavaScript. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):275, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Suganuma:2004:EJJ

[SOK^{+}04] T. Suganuma, T. Ogasawara, K. Kawachiya, M. Takeuchi, K. Ishizaki, A. Koseki, T. Inagaki, T. Yasue, M. Kawahito, and T. Onodera. Evolution of a Java just-in-time compiler for IA-32 platforms. IBM Journal of Research and Development, 48(5/6):767–796, 2004. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Sooriamurthi:2001:PJE

[Soo01]  Raja Sooriamurthi.  Pre-

lude to the Java event model (poster session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3): 193, September 2001. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sooriamurthi:2009:IAD

[Soo09] Raja Sooriamurthi. Introducing abstraction and decomposition to novice programmers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):196–200, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Suganuma:2000:OIJ

[SOT^{+}00] T. Suganuma, T. Ogasawara, M. Takeuchi, T. Yasue, M. Kawahito, K. Ishizaki, H. Komatsu, and T. Nakatani. Overview of the IBM Java Just-in-Time Compiler. IBM Systems Journal, 39(1):175– 193, ???? 2000. CO-DEN IBMSA7. ISSN 0018-

8670. URL http://www.
almaden.ibm.com/journal/
sj/391/suganuma.html.

Stevenson:2003:IOE

[SP03] D. E. Stevenson and A. T. Phillips. Implementing object equivalence in Java using the template method design pattern. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):278–282, 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Shapiro:2001:FJR

[SPB01] Joel A. Shapiro, Warren B.

Powell, and David Bernstein. A flexible Java representation for uncertainty in online operations-research models. INFORMS Journal on Computing, 13(1):29–55, Winter 2001. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:/

/pubsonline.informs.org/
doi/abs/10.1287/ijoc.13.
1.29.9749.

Smiley:2009:SES

[SPBE09] David Smiley, Eric (David Eric)

Pugh, James Brady, and Jerome Eteve. Solr 1.4 Enterprise Search Server: enhance your search with faceted navigation, result highlighting, fuzzy queries, ranked scoring, and more. Packt Publishing, Birmingham, UK, 2009. ISBN 1-84719-588-1. vii + 317 pp. LCCN TK5105.885.S65 S65 2009eb.

Speegle:2002:JPG

[Spe02] Gregory D. (Gregory David)

Speegle. JDBC: practical guide for Java programmers. Morgan Kaufmann practical guides series. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2002. ISBN 1-55860-736-6. xiii + 113 pp. LCCN QA76.73.J38 S64 2002.

Schneider:2007:OOD

[SPG07] Florian T. Schneider, Mathias Payer, and Thomas R. Gross. Online optimizations driven by hardware performance monitoring. ACM SIGPLAN Notices, 42(6): 373–382, June 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Spring:2007:SHT

[SPGV07] Jesper H. Spring, Jean Privat, Rachid Guerraoui, and Jan Vitek. StreamFlex: highthroughput stream programming in Java. ACM SIG-PLAN Notices, 42(10):211– 228, October 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Spielman:2003:JPG

[Spi03a] Sue Spielman. JSTL: Practical Guide for JSP Programmers. Morgan Kaufmann practical guides series. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2003. ISBN 0-12-656755-7. xix + 229 pp. LCCN TK5105.8885.J38 S65 2004. US$19.95. URL ftp:/

                                        /uiarchive.cso.uiuc.edu/
                                        pub/etext/gutenberg/;
                                        http://www.loc.gov/catdir/
                                        description/els031/2003107479.
                                        html.

Spielman:2003:SFP

[Spi03b] Sue Spielman. The Struts Framework: Practical Guide for Java Programmers. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2003. ISBN 1-55860-862-1. xiv + 137 pp. LCCN TK5105.8885.S76 S65 2003. US$19.95.

Spinellis:2005:JMS

[Spi05]  Diomidis Spinellis.  Java

makes scripting languages irrelevant? IEEE Software, 22(3):70–71, May/June 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Stahl:2003:PAI

[SPR^{+}03] R. Stahl, R. Pasko, L. Rijnders, D. Verkest, S. Vernalde, R. Lauwereins, and F. Catthoor. Performance analysis for identification of (sub-)task-level parallelism in Java. Lecture Notes in Computer Science, 2826:313–328, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Scime:2002:LIS

[SPS^{+}02] Anthony Scime, Bob Pedigo, Curtis Sjolander, Bob Colwell, Dick Brodine, Ted Lewis, Bruce D. Wilner, Gary McGraw, Jay Dearien, Neville Holmes, and Peter Stern. Letters: Information system engineering; quantum questions; is Murphy’s Law wrong?; Java questions for Ted; software security concerns; insightful observations on computers and society. Computer, 35 (6):6–9, June 2002. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/co/
2002/06/r6006.htm;   http:
//csdl.computer.org/dl/
mags/co/2002/06/r6006.pdf.

Stromer:2005:JHJ

[SQG^{+}05] Julie N. Stromer, Gerald T. Quon, Paul M. K. Gordon, Andrei L. Turinsky, and Christoph W. Sensen. Jabiru: Harnessing Java 3D behaviors for device and display portability. IEEE Computer Graphics and Applications, 25 (2):70–80, March/April 2005. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Salcianu:2005:PSE

[SR05] A. Salcianu and M. Rinard. Purity and side effect analysis for Java programs. Lecture Notes in Computer Science, 3385:199–215, 2005. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sharp:2006:SAO

[SR06] M. Sharp and A. Rountev. Static analysis of object references in RMI-based Java software. IEEE Transactions on Software Engineering, 32(9):664–681, September 2006. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL

http://ieeexplore.ieee.
org/stamp/stamp.jsp?arnumber=
1707666.

Sowizral:2000:JAS

[SRD00] Henry A. Sowizral, Kevin Rushforth, and Michael Deering. The Java 3D API specification. Java series. Addison-Wesley, Reading, MA, USA, second edition, 2000. ISBN 0-201-71041-2 (paperback/CD-ROM). xxii + 653 pp. LCCN T385 .S6575 2000.

Sun:2008:JBH

[SRJS08] J.-Z. Sun, J. Riekki, M. Jurmu, and J. Sauvola. Javabased HTTP input channel for heterogeneous wireless networks. International Journal of Computer Applications, 30(2):97–104, 2008. ISSN 1206-212X (print), 1925-7074 (electronic). URL https://

www.tandfonline.com/doi/
full/10.1080/1206212X.2008.
11441887.

Shields:2000:JCB

[SRW^{+}00] Matthew S. Shields, Omer F.

Rana, David W. Walker, Maozhen Li, and David Golby. A Java/CORBAbased visual program composition environment for PSEs. Concurrency: Practice and Experience, 12(8):

687–704, July 2000. CO-DEN CPEXEI. ISSN 1040-

                                        3108. URL http://www3.
                                        interscience.wiley.com/
                                        cgi-bin/abstract/72516220/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=72516220&PLACEBO=IE.
                                        pdf.

Stark:2000:PBV

[SS00a] R. F. Stärk and J. Schmid. The problem of bytecode verification in current implementations of the JVM. Technical report, Department of Computer Science, ETH Zürich, Zürich, Switzerland, 2000.

Steflik:2000:AJN

[SS00b] Richard Steflik and Prashant Sridharan. Advanced Java Networking. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2000. ISBN 0-13-084466-7. xxiv + 399 pp. LCCN QA76.73.J38 S83 2000. US$45. Includes CD-ROM.

Serpette:2002:CSJ

[SS02] Bernard Paul Serpette and Manuel Serrano. Compiling scheme to JVM bytecode: a performance study. ACM SIGPLAN Notices, 37 (9):259–270, September 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Stark:2003:CBV

[SS03] Robert F. Stärk and Joachim Schmid. Completeness of a bytecode verifier and a certifying Java-to-JVM compiler. Journal of Automated Reasoning, 30(3–4): 323–361, May 2003. CO-DEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A%3A1025003423108.

Shalev:2006:PLS

[SS06] Ori Shalev and Nir Shavit.

Predictive log-synchronization. Operating Systems Review, 40 (4):305–315, October 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Settle:2007:DLS

[SS07] A. Settle and C. Settle. Distance learning and student satisfaction in Java programming courses. J.UCS: Journal of Universal Computer Science, 13(9):1270–1286, ???? 2007. CODEN ???? ISSN 0948-6968. URL http:

//www.jucs.org/jucs_13_
9/distance_learning_and_
student.

Singh:2008:DRM

[SS08] Paramvir Singh and Hardeep Singh. DynaMetrics: a runtime metric-based analysis tool for object-oriented software systems. ACM SIG-SOFT Software Engineering Notes, 33(6):1–6, November 2008. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Strom:2003:UJT

[SSA03] O. Strom, K. Svarstad, and E. J. Aas. On the utilization of Java technology in embedded systems. Design Automation for Embedded Systems, 8(1):87–106, 2003. CO-DEN ???? ISSN 0929-5585.

Stark:2001:JJV

[SSB01] Robert F. Stärk, Joachim Schmid, and Egon Börger. Java and the Java Virtual Machine: definition, verification, validation. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 3-540-42088-6. x + 381 pp. LCCN QA76.73.J38 S785 2001. US$49.95. Includes CD-ROM with the entire text of the book and numerous examples and exercises.

Shaylor:2003:JVM

[SSB03] Nik Shaylor, Douglas N. Simon, and William R. Bush. A Java virtual machine architecture for very small devices. ACM SIGPLAN Notices, 38(7):34–41, July 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shi:2000:MAS

[SSC00] Hongchi Shi, Yi Shang, and Su-Shing Chen. A multiagent system for computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):1– 4, September 2000. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sammapun:2003:FJM

[SSD^{+}03] U. Sammapun, R. Sharykin, M. DeLap, M. Kim, and S. Zdancewic. Formalizing Java-MaC. Electronic Notes in Theoretical Computer Science, 89(2):1–20, 2003. CO-DEN ???? ISSN 1571-0661.

Suwimonteerabuth:2005:JJB

[SSE05] D. Suwimonteerabuth, S. Schwoon, and J. Esparza. jMoped: a Java bytecode checker based on Moped. Lecture Notes in Computer Science, 3440:541– 545, 2005. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shuf:2001:CMB

[SSGS01] Yefim Shuf, Mauricio J. Serrano, Manish Gupta, and Jaswinder Pal Singh. Characterizing the memory behavior of Java workloads: a structured view and opportunities for optimizations. In ACM [ACM01d], pages 194– 205. ISBN 1-58113-334-0.

LCCN QA76.9.E94; Internet. ACM order number 488010.

Suppi:2002:PDP

[SSL02] R. Suppi, M. Solsona, and E. Luque. Parasite: Distributing processing using Java applets (research note). Lecture Notes in Computer Science, 2400:598–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2400/24000598.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2400/24000598.
pdf.

Sakabe:2003:JOT

[SSM03] Y. Sakabe, M. Soshi, and A. Miyaji. Java obfuscation with a theoretical basis for building secure mobile agents. Lecture Notes in Computer Science, 2828:89– 103, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shudo:2004:CEC

[SSM04] K. Shudo, S. Sekiguchi, and Y. Muraoka. Cost-effective compilation techniques for Java just-in-time compilers. Systems and computers in Japan, 35(12):10–24, 2004. CODEN SCJAEP. ISSN 0882-1666 (print), 1520-684X (electronic).

Strnisa:2007:JMS

[SSP07] Rok Strniša, Peter Sewell, and Matthew Parkinson. The Java module system: core design and semantic definition. ACM SIGPLAN Notices, 42(10):499–514, October 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Soldar:2002:UWS

[SSS02] G. Soldar, B. Spencer, and D. Smith. Using Web services and Java distributed components to develop distributed applications. Applied Informatics, 2:388–394, 2002. CO-DEN ???? ISSN 1027-2666.

Soomro:2005:DDH

[SSS05] H. K. Soomro, S. A. A. Shah, and A. A. G. Shaikh. Development of dynamic hashing key generation method for Java based Sindhi programming environment. Mehran University Research Journal of Engineering and Technology, 24(2):125–130, 2005. CO-DEN ???? ISSN 0254-7821.

Skalka:2005:TES

[SSV05] C. Skalka, S. Smith, and D. VanHorn. A type and effect system for flexible abstract interpretation of Java. Electronic Notes in Theoretical Computer Science, 131: 111–124, 2005. CODEN ???? ISSN 1571-0661.

Snelting:2000:UCH

[ST00a] Gregor Snelting and Frank Tip. Understanding class hierarchies using concept analysis. ACM Transactions on Programming Languages and Systems, 22(3):540–582, 2000. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL

                                        http://www.acm.org/pubs/
                                        articles/journals/toplas/
                                        2000-22-3/p540-snelting/
                                        p540-snelting.pdf;   http:
                                        //www.acm.org/pubs/citations/
                                        journals/toplas/2000-22-
                                        3/p540-snelting/.

Sweeney:2000:ELB

[ST00b] Peter F. Sweeney and Frank Tip. Extracting library-based object-oriented applications. ACM SIGSOFT Software Engineering Notes, 25(6):98– 107, November 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Schrefl:2004:URJ

[ST04] Michael Schrefl and Thomas Thalhammer. Using roles in Java. Software—Practice and Experience, 34(5): 449–464, April 25, 2004. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Spivak:2006:SPT

[ST06] Michal Spivak and Sivan Toledo. Storing a persistent transactional object heap on flash memory. ACM SIG-PLAN Notices, 41(7):22–33, July 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Song:2009:ESL

[ST09] Myoungkyu Song and Eli Tilevich. Enhancing sourcelevel programming tools with an awareness of transparent program transformations. ACM SIGPLAN Notices, 44(10):301–320, October 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Stankovic:2000:OJS

[Sta00] Nenad Stankovic. An open Java system for SPMD programming. Concurrency: Practice and Experience, 12(11):1051–1076, September 2000. CO-DEN CPEXEI. ISSN 1040-

3108. URL http://www3.
interscience.wiley.com/
cgi-bin/abstract/76000192/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=76000192&PLACEBO=IE.
pdf.

Stankovski:2001:AIJ

[Sta01]  Zarko Stankovski.  AR-

LEQUIN: An integrated Java application. In ACM [ACM01b], page 183. ISBN 1-58113-359-6. LCCN QA76.9.O35

                                        A26 2001. URL http://www.
                                        philippsen.com/JGI2001/
                                        camerareadyabstracts/47.
                                        html;   http://www.philippsen.
                                        com/JGI2001/finalpapers/
                                        18500183.dvi.

Stallman:2004:FSJ

[Sta04a] Richard Stallman. Free but shackled — the Java trap. World-Wide Web document, April 12, 2004. URL http:/

                                        /www.gnu.org/philosophy/
                                        java-trap.html.

Stark:2004:FSC

[Sta04b] Eugene W. Stark. Formally specifying CARA in Java. International Journal on Software Tools for Technology Transfer (STTT), 5(4):331– 350, May 2004. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Serfass:2008:SSP

[STB08] Doug Serfass, Kalyan S. Thiparthi, and Coskun Bayrak. Search, sort, pagination, association and (other) CRUD: a comparison of ActiveScaffold and MyFaces tomahawk. ACM SIGSOFT Software Engineering Notes, 33 (4):1:1–1:??, July 2008. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Stevens:2000:CPP

[Ste00] Al Stevens. C programming: The S programming language. Dr. Dobb’s Journal of Software Tools, 25 (2):110–113, February 2000. CODEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/ftp/2000/
2000_02/cprog220.txt. See

comment on another S language [KSC^{+}00].

Steele:2001:NMN

[Ste01] Guy Steele. New models for numerical computing in the Java programming language. In ACM [ACM01b], page ?? ISBN 1-58113-359-6. LCCN QA76.9.O35 A26 2001.

Stenzel:2004:FVC

[Ste04] K. Stenzel. A formally verified calculus for full Java card. Lecture Notes in Computer Science, 3116:491–505, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stelting:2005:RJE

[Ste05] Stephen Stelting. Robust Java: exception handling, testing, and debugging. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2005. ISBN 0-13-100852-8. xxii + 375 pp. LCCN QA76.73.J38 S846 2005.

Steyer:2008:JDI

[Ste08a]  Ralph Steyer.  JavaFX:

dynamische und interaktive Java-Applikationen mit JavaFX. Programmer‘s choice. Addison-Wesley, Reading, MA, USA, 2008. ISBN 3-8273-2615-X. 363 pp. LCCN

????  URL http://www.
gbv.de/dms/ilmenau/toc/
55624642X.PDF.

Steyer:2008:JHC

[Ste08b] Ralph Steyer, editor. JavaFX: Einstieg für Anspruchsvolle - 7 Stunden Video-Training; Lehr-Programm. Addison-Wesley, Reading, MA, USA, 2008. ISBN 3-8273-6144-3. LCCN ???? EUR 39.95, SFR 65.00. URL http://

deposit.d-nb.de/cgi-bin/
dokserv?id=3091158&prov=
M&dok\_var=1&dok\_ext=htm.

1 DVD (audio und rom).

Story:TB22-4-265

[Sto01a] D. P. Story. execJS: A new technique for introducing discardable JavaScript into a PDF file from a L^{A}TEX source. TUGboat, 22(4):265–268, December 2001. ISSN 0896-3207.

Story:TB22-3-161

[Sto01b] Donald P. Story. Techniques of introducing document-level JavaScript into a PDF file from a L^{A}TEX source. TUGboat, 22(3):161–167, September 2001. ISSN 0896-3207.

Stoller:2002:DPO

[Sto02a] Scott D. Stoller. Domain partitioning for open reactive systems. ACM SIGSOFT Software Engineering Notes, 27(4):44–54, July 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Stoller:2002:MCM

[Sto02b] Scott D. Stoller. Modelchecking multi-threaded distributed Java programs. International Journal on Software Tools for Technology Transfer (STTT), 4(1):71–91, October 2002. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic).

Strunk:2001:JQJ

[Str01]  Elisabeth Strunk.  Java

Q&A: Java & NT authentication. Dr. Dobb’s Journal of Software Tools, 26 (2):145–146, 148, February 2001. CODEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/ftp/2001/
2001_02/jqa0201.txt.

Strecker:2002:FVJ

[Str02] Martin Strecker. Formal verification of a Java compiler in Isabelle. Lecture Notes in Computer Science, 2392:63– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2392/23920063.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2392/23920063.
pdf.

Studer:2001:CFF

[Stu01] Thomas Studer. Constructive foundations for featherweight Java. Lecture Notes in Computer Science, 2183:202– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2183/21830202.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2183/21830202.
pdf.

Stubblebine:2007:REP

[Stu07] Tony Stubblebine. Regular expression pocket reference: Regular expressions for Perl, Ruby, PHP, Python, C, Java, and .NET. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2007. ISBN 0-596-51427-1 (paperback). vii + 117 pp. LCCN QA76.9.T48 S78 2007. URL http:/

/www.loc.gov/catdir/toc/
fy0802/2007281074.html.

Sage:2003:TIP

[SU03] D. Sage and M. Unser. Teaching image-processing programming in Java. IEEE signal processing magazine, 20(6):43–52, 2003. CO-DEN ISPRE6. ISSN 1053-5888 (print), 1558-0792 (electronic).

Subramaniam:2008:PST

[Sub08] Venkat Subramaniam. Programming Scala: tackle multicore complexity on the JVM. Pragmatic Bookshelf, Raleigh, NC, USA, 2008. ISBN 1-934356-31-X (paperback). x + 221 pp. LCCN QA76.73.J38 S83 2008.

Sung:2001:DSL

[Sun01] K.-Y. Sung. Design of a specification language and realtime APIs for easy expression of soft real-time constraints with Java. Lecture Notes in Computer Science, 2093:530– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2093/20930530.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2093/20930530.
pdf.

Sun:2002:BJP

[Sun02] Sun Microsystems. BigDecimal (Java 2 Platform SE v1.4.0). Sun Microsystems, Mountain View, CA, USA, 2002. 17 pp. URL http://

java.sun/com/products.

Suokas:2004:JHS

[Suo04]  J. Suokas.  JAVA ALS

HANDY-STANDARD? Telekompraxis, BAND81(NUMB 11/12): 22–23, 2004. CODEN ???? ISSN 0938-5460.

Suri:2001:SCR

[Sur01] Niranjan Suri. State capture and resource control for Java: The design and implementation of the Aroma Virtual Machine. In USENIX Association [USE01c], page ?? ISBN 1-880446-11-1. LCCN QA76.73.J38 J42 2001. URL

http://www.usenix.org/publications/
library/proceedings/jvm01/
JVM_wips/S15.pdf. Spon-

sored by the USENIX Association.

Surveyer:2004:SAO

[Sur04a] J. Surveyer. Sun adds to opensource Java IDE roster: a review of NetBeans Java IDE. Application Development Trends, 11(9):48, 2004. CODEN ???? ISSN 1073-9564.

Surveyer:2004:SJS

[Sur04b] J. Surveyer. Sun Java Studio Creator: The VB of Java? Application Development Trends, 11(11):46–47, 2004. CODEN ???? ISSN 1073-9564.

Silveira:2002:DDI

[SV02] Ricardo Azambuja Silveira and Rosa Maria Vicari. Developing distributed intelligent learning environment with JADE — Java agents for distance education framework. Lecture Notes in Computer Science, 2363:105–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2363/23630105.
htm;   http://link.springer-
ny.com/link/service/series/

0558/papers/2363/23630105.
pdf.

Santone:2005:LAT

[SV05] A. Santone and G. Vaglini.

A local approach for temporal model checking of Java bytecode. Journal of Computer and System Sciences, 70(2):258–281, 2005. CO-DEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic).

Sips:2001:JSC

[SvR01] H. J. Sips and K. van Reeuwijk. Java for scientific computation: Prospects and problems. Lecture Notes in Computer Science, 2179:236– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2179/21790236.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2179/21790236.
pdf.

Shacham:2009:CAS

[SVY09] Ohad Shacham, Martin Vechev, and Eran Yahav. Chameleon: adaptive selection of collections. ACM SIG-PLAN Notices, 44(6):408– 418, June 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Siebert:2001:DEJ

[SW01] Fridtjof Siebert and Andy Walter. Deterministic execution of Java’s primitive bytecode operations. In USENIX Association [USE01c], page ?? ISBN 1-880446-11-1. LCCN QA76.73.J38

                                        J42 2001. URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/jvm01/
                                        siebert.html. Sponsored by

the USENIX Association.

Su:2006:ECI

[SW06] Zhendong Su and Gary Wassermann. The essence of command injection attacks in Web applications. ACM SIG-PLAN Notices, 41(1):372– 382, January 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Swaine:2001:PPA

[Swa01a] Michael Swaine. Programming paradigms: Alternatives to Java. Dr. Dobb’s Journal of Software Tools, 26(2):135– 137, February 2001. CODEN DDJOEB. ISSN 1044-789X.

                                        URL http://www.ddj.com/.

Swan:2001:JJC

[Swa01b] Tom Swan. Java 2 just click! solutions. Hungry Minds, 909 Third Avenue, New York, NY 10022, USA, 2001. ISBN 0-7645-4823-9. xix + 672 pp. LCCN QA76.73.J38 S92 2001. Accompanying CD-ROM contains source code to over 150 sample Java 2 programs, Microsoft Internet Explorer, Netscape Communicator for Windows and Linux and the author’s hyperlinked indexes.

Sward:2007:UAS

[Swa07] Ricky E. Sward. Using Ada in a service-oriented architecture. ACM SIGADA Ada Letters, 27(3):63–68, December 2007. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Sweeney:2006:NMP

[Swe06] Tim Sweeney. The next mainstream programming language: a game developer’s perspective. ACM SIGPLAN Notices, 41(1):269, January 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shao:2004:RPF

[SY04] J. L. Shao and Y. Q. Ye. Realization of programming functions using Java in a VRML world. Computers and Applied Chemistry, 21(4):641– 644, 2004. CODEN ???? ISSN 1001-4160.

Skeie:2005:PIC

[SY^{+}05] Tor Skeie, Chu-Sing Yang, et al., editors. Proceedings: 2005 International Conference on Parallel Processing Workshops, 14–17 June 2005, Oslo, Norway. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2005. ISBN 0-7695-2381-1. LCCN QA76.58; QA76.58 .I34 2005. URL

http://ieeexplore.ieee.
org/servlet/opac?punumber=

9965. IEEE Computer Society order number P2381.

Shah:2005:SET

[SYAS05] Vijay Shah, Nicholas Younan, Torey Alford, and Anthony Skjellum. A spectral estimation toolkit for Java applications. Science of Computer Programming, 54(1): 125–142, January 2005. CO-DEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Suganuma:2001:DOF

[SYK^{+}01] Toshio Suganuma, Toshiaki Yasue, Motohiro Kawahito, Hideaki Komatsu, and Toshio Nakatani. A dynamic optimization framework for a Java just-in-time compiler. ACM SIGPLAN Notices, 36(11):180–195, November 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Proceedings of the 2001 ACM SIGPLAN Conference on Object Oriented Programming, Systems, Languages and Applications (OOPSLA’01).

Suganuma:2005:DED

[SYK^{+}05] Toshio Suganuma, Toshiaki Yasue, Motohiro Kawahito, Hideaki Komatsu, and Toshio Nakatani. Design and evaluation of dynamic optimizations for a Java just-in-time compiler. ACM Transactions on Programming Languages and Systems, 27(4):732–785, July 2005. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Suganuma:2002:ESM

[SYN02] Toshio Suganuma, Toshiaki Yasue, and Toshio Nakatani. An empirical study of method in-lining for a Java just-in-time compiler. In USENIX Association [USE02], page ?? ISBN 1-931971-01-3. LCCN QA76.73.J38 J42 2002. URL

http://www.usenix.org/publications/
library/proceedings/javavm02/
suganuma.html.

Suganuma:2003:RBC

[SYN03] Toshio Suganuma, Toshiaki Yasue, and Toshio Nakatani. A region-based compilation technique for a Java just-in-time compiler. ACM SIG-PLAN Notices, 38(5):312– 323, May 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Suganuma:2006:RBC

[SYN06] Toshio Suganuma, Toshiaki Yasue, and Toshio Nakatani. A region-based compilation technique for dynamic compilers. ACM Transactions on Programming Languages and Systems, 28(1):134–174, January 2006. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Stankovic:2000:EJI

[SZ00] Nenad Stankovic and Kang Zhang. An evaluation of Java implementations of messagepassing. Software—Practice and Experience, 30(7): 741–763, June 2000. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (elec-

                                        tronic). URL http://www3.
                                        interscience.wiley.com/
                                        cgi-bin/abstract/72001830/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=72001830&PLACEBO=IE.
                                        pdf.

Tellis:2004:IMC

[TA04] W. M. Tellis and K. P. Andriole. Integrating multiple clinical information systems using the Java Message Service Framework. Journal of Digital Imaging, 17(2):80–86, 2004. CODEN JDIMEW. ISSN 0897-1889.

Titzer:2007:ESA

[TABP07] Ben L. Titzer, Joshua Auerbach, David F. Bacon, and Jens Palsberg. The ExoVM system for automatic VM and application reduction. ACM SIGPLAN Notices, 42(6):352–362, June 2007. CODEN SINODQ.

ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tamura:2000:DWP

[Tam00] Randall A. Tamura. Domino 5 Web Programming With XML, Java and JavaScript. Que Corporation, Indianapolis, IN, USA, 2000. ISBN 0-7897-2275-5. xvi + 912 pp. LCCN QA76.625 .T36 2000. US$50.

Tang:2007:PRI

[Tan07] Audrey Tang. Perl 6: reconciling the irreconcilable. ACM SIGPLAN Notices, 42 (1):1, January 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tate:2002:BJ

[Tat02] Bruce Tate. Bitter Java.

Manning, Greenwich, CT, USA, 2002. ISBN 1-930110-43-X. xxviii + 339 pp. LCCN QA76.73.J38 T38 2002.

Tate:2005:BJ

[Tat05] Bruce Tate. Beyond Java.

O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2005. ISBN 0-596-10094-9. xii + 185 pp. LCCN QA76.73.J38 T38 2005; QA76.73.J38 T38 2005eb; QA76.73.J38. URL

http://www.oreilly.com/
catalog/9780596100940.

Titchkosky:2003:PCD

[TAW03] Lance Titchkosky, Martin Arlitt, and Carey Williamson. A performance comparison of dynamic Web technologies. ACM SIGMETRICS Perform. Eval. Rev., 31(3):2– 11, December 2003. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Taylor:2002:JJC

[Tay02] Art Taylor. JSP and Java: the complete guide to Website development. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2002. ISBN 0-13-091813-X. xxii + 518 pp. LCCN QA76.73.J38 T39 2002.

Tempero:2000:SMI

[TB00a] Ewan Tempero and Robert Biddle. Simulating multiple inheritance in Java. The Journal of systems and software, 55(1):87–100, November 5, 2000. CODEN JS-SODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL http://

www.elsevier.nl/gej-ng/
10/29/11/54/28/32/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/29/11/54/28/
32/article.pdf.

Turner:2000:HJP

[TB00b] Shane E. Turner and Karl Barksdale. HTML and JavaScript: programming concepts. International Thomson Publishing, London, New York, Boston, Madrid, ldots, 2000. ISBN 0-538-68822-X. xi + 154 pp. LCCN ????

Tilly:2002:ADG

[TB02] Jesse E. Tilly and Eric M.

Burke. Ant: The Definitive Guide. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2002. ISBN 0-596-00184-3. xviii + 269 pp. LCCN QA76.73.J38 T57 2002. US$34.95. URL http:/

/safari.oreilly.com/0596001843;
http://www.oreilly.com/
catalog/anttdg.

Tyman:2009:ABS

[TBM09] Damon Tyman, Nirupama Bulusu, and Jens Mache. An activity-based sensor networks course for undergraduates with sun spot devices. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41 (1):34–38, March 2009. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Tanter:2001:RTO

[TBSN01] Éric Tanter, Noury M. N.

Bouraqadi-Saâdani, and Jacques Noyé. Reflex — towards an open reflective extension of Java. Lecture Notes in Computer Science, 2192:25– ??, 2001. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2192/21920025.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2192/21920025.
pdf.

Tan:2003:JAC

[TC03] F. L. Tan and P. S. K. Chua. Java applet on computation of thermodynamic properties of steam and R134a refrigerant. Computer Applications in Engineering Education, 11 (4):211–225, 2003. CODEN CAPEED. ISSN 1061-3773.

Tsang:2004:OPB

[TC04] A. C. W. Tsang and N. Chan. An online problem-based model for the learning of Java. Journal of Electronic Commerce in Organizations, 2(2): 55–64, 2004. CODEN ???? ISSN 1539-2937.

Ton:2001:EJB

[TCC01] Lee-Ren Ton, Lung-Chung Chang, and Chung-Ping Chung. Exploiting Java bytecode parallelism by enhanced POC folding model (research note). Lecture Notes in Computer Science, 1900:994–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/1900/19000994.
htm;   http://link.springer-

ny.com/link/service/series/
0558/papers/1900/19000994.
pdf.

Ton:2002:APS

[TCC02] L.-R. Ton, L.-C. Chang, and C.-P. Chung. An analytical POC stack operations folding for continuous and discontinuous Java bytecodes. Journal of Systems Architecture, 48(1):1–16, September 2002. CODEN JSARFB. ISSN 1383-7621 (print), 1873-6165 (electronic).

Tigli:2003:WRA

[TCF^{+}03] J. Y. Tigli, D. Cheung, J. Fuchet, G. Joulie, and F. Grillon. Wcomp: Rapid application development toolkit for wearable computer based on Java. In IEEE International Conference on Systems Man and Cybernetics, volume 5, pages 4198– 4203. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2003. CODEN ???? ISSN 1062-922X.

Tucker:2000:LEP

[TCM^{+}00] Andrew Tucker, Edoardo Comar, Scott Meyers, Yves Piguet, Kevin Ruland, Greg Hadaller, Jonathan Erickson, Mike Zhilin, and Todd Stephan. Letters: Editor preferences; Java enums; labor union harassment; smart pointer update; traveling salesman; granting block grant; porting to CE; analyzing algorithms. Dr. Dobb’s Journal of Software Tools, 25 (1):10, 12, January 2000. CO-DEN DDJOEB. ISSN 1044-789X.

Ton:2002:DOF

[TCSC02] L.-R. Ton, L.-C. Chang, J.- J. Shann, and C.-P. Chung. Design of an optimal folding mechanism for Java processors. Microprocessors and Microsystems, 26(8):341–352, November 10, 2002. CO-DEN MIMID5. ISSN 0141-9331 (print), 1872-9436 (electronic).

Ton:2004:SHC

[TCSC04] Lee-Ren Ton, Lung-Chung Chang, Jyh-Jiun Shann, and Chung-Ping Chung. A software/hardware cooperated stack operations folding model for Java processors. The Journal of systems and software, 72(3):377–387, August 2004. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Thiruvathukal:2000:JNW

[TDB00] George K. Thiruvathukal, Phillip M. Dickens, and Shahzad Bhatti. Java on networks of workstations (JavaNOW): a parallel computing framework inspired by Linda and the Message Passing Interface (MPI). Concurrency: Practice and Experience, 12(11):1093–1116, September 2000. CO-DEN CPEXEI. ISSN 1040-

3108. URL http://www3.
interscience.wiley.com/
cgi-bin/abstract/76000187/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=76000187&PLACEBO=IE.
pdf.

Taveira:2003:ARM

[Tddd03] Wendell Figueiredo Taveira, Marco Tulio de Oliveira Valente, Mariza Andrade da Silva Bigonha, and Roberto da Silva Bigonha. Asynchronous remote method invocation in Java. J.UCS: Journal of Universal Computer Science, 9(8):761–??, August 28, 2003. CODEN ???? ISSN 0948-6968.

URL http://www.jucs.org/
jucs_9_8/asynchronous_remote_
method_invocation.

Tan:2004:EEE

[TE04] Roy Patrick Tan and Stephen H.

Edwards. Experiences evaluating the effectiveness of JML-JUnit testing. ACM SIGSOFT Software Engineering Notes, 29(5):1–4, September 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tschantz:2005:JAR

[TE05] Matthew S. Tschantz and Michael D. Ernst. Javari: adding reference immutability to Java. ACM SIG-PLAN Notices, 40(10):211– 230, October 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Todd:2001:LSS

[TEM^{+}01] Andrew W. Todd, Jonathan Erickson, Nadine McKenzie, Chris Cleeland, Richard Huang, Ragae Ghaly, and The Editors. Letters: Shared source and shared secrets; JavaScript fix; CORBA interoperability; EJB application servers update; correction [“The Delphi XML SAX2 Component and MSXML 3.0”]. Dr. Dobb’s Journal of Software Tools, 26(10):10, 12, October 2001. CODEN DDJOEB. ISSN 1044-789X.

                                        URL http://www.ddj.com/.

See [Hei01].

Tennebo:2000:EJE

[Ten00] Frode Tennebø. Elegance of Java and the efficiency of C++ — it’s Ada. Linux Journal, 80:174–176, December 2000. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Thornton:2008:SSW

[TETPQ08] Matthew Thornton, Stephen H. Edwards, Roy P. Tan, and Manuel A. Pérez-Quiñones. Supporting student-written tests of GUI programs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40 (1):537–541, March 2008.

CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Tran:2004:TCB

[TFL^{+}04] F. D. Tran, J. P. Fassino, O. Lobry, J. Pulou, and N. Rivierre. Toward a component-based embedded Java-oriented operating system. Lecture Notes in Computer Science, 3292:256–267, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tate:2004:BFL

[TG04] Bruce Tate and Justin Gehtland. Better, Faster, Lighter Java. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2004. ISBN 0-596-00676-4. xvi + 243 pp. LCCN QA76.73.J38 T36 2004. US$34.95, CAN$50.95, UK£24.95. URL http://

www.oreilly.com/catalog/
bfljava/.

Talpin:2004:HRT

[TGB^{+}04] J. P. Talpin, A. Gamatie, D. Berner, B. LeDez, and P. LeGuernic. Hard realtime implementation of embedded software in Java. Lecture Notes in Computer Science, 2952:33–47, 2004. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thomas:2008:DHF

[TGCF08] Gaël Thomas, Nicolas Geoffray, Charles Clément, and Bertil Folliot. Designing highly flexible virtual machines: the JnJVM experience. Software—Practice and Experience, 38(15):1643– 1675, December ??, 2008. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Tate:2005:SDN

[TGL05] Bruce Tate, Justin Gehtland, and Michael Kosta Loukides. Spring: a developer’s notebook. A developer’s notebook. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2005. ISBN 0-596-00910-0. xv + 184 pp. LCCN QA76.73.S67 T38 2005.

Tan:2000:PEN

[TGO00] Kian-Lee Tan, Cheng Hian Goh, and Beng Chin Ooi. Progressive evaluation of nested aggregate queries. VLDB Journal: Very Large Data Bases, 9(3):261–278, December 2000. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X (electronic).

Tamassia:2001:JDS

[TGV^{+}01] Roberto Tamassia, Michael T. Goodrich, Luca Vismara, Mark Handy, Galina Shubina, Robert Cohen, Benoît Hudson, Ryan S. Baker, Natasha Gelfand, and Ulrik Brandes. JDSL: The data structures library in Java. Dr. Dobb’s Journal of Software Tools, 26(4):21– 22, 24, 26, 28, 30–31, April 2001. CODEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/ftp/2001/
2001_04/jdsl.txt.

Tozawa:2002:FAC

[TH02] A. Tozawa and M. Hagiya.

Formalization and analysis of class loading in Java. Higher-Order and Symbolic Computation, 15(1):7–55, March 2002. CODEN LSCOEX. ISSN 1388-3690 (print), 2212-0793 (electronic).

Thau:2000:BJ

[Tha00] Dave Thau. The Book of JavaScript: a practical guide to interactive Web pages. No Starch Press, San Francisco, CA, USA, 2000. ISBN 1-886411-36-0. xix + 397 pp. LCCN QA76.73.J39 T37 2000.

Thau:2006:BJP

[Tha06] Thau. The book of JavaScript:

a practical guide to interactive Web pages. No Starch Press, San Francisco, CA, USA, second edition, 2006. ISBN 1-59327-106-9. xxx + 490 pp. LCCN QA76.73.J39 T37 2006eb; QA76.73.J39 T37 2007. URL http://

www.oreilly.com/catalog/
9781593271060.

Thiruvathukal:2002:JMA

[Thi02] George K. Thiruvathukal. Java at middle age: Enabling Java for computational science. Computing in Science and Engineering, 4(1):74– 84, January/February 2002. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

                                        computer.org/cise/cs2001/
                                        c1074abs.htm;   http://
                                        dlib.computer.org/cs/books/
                                        cs2002/pdf/c1074.pdf.

Tikir:2003:RDS

[THL03] Mustafa M. Tikir, Jeffrey K. Hollingsworth, and Guei-Yuan Lueh. Recompilation for debugging support in a JIT-compiler. ACM SIG-SOFT Software Engineering Notes, 28(1):10–17, January 2003. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Trost:2003:JEB

[THMT03] E. Trost, H. Hackl, M. Maurer, and Z. Trajanoski. Java editor for biological pathways. Bioinformatics, 19(6): 786–787, 2003. CODEN ???? ISSN 1367-4803 (print), 1367-4811 (electronic).

Thomas:2003:OXC

[Tho03] Michael D. Thomas. Oracle XSQL: combining SQL, Oracle text, XSLT, and Java to publish dynamic Web content. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 2003. ISBN 0-471-27120-9 (paper/website). xix + 572 pp. LCCN QA76.73.S67 T48 2003. URL http://

www.loc.gov/catdir/bios/
wiley045/2002014913.html;
 http://www.loc.gov/catdir/
description/wiley038/2002014913.
html;   http://www.loc.
gov/catdir/toc/wiley031/
2002014913.html.

Timpe:2003:GCJ

[Tim03] R. Timpe. GPIB control with Java on Linux or Windows. Evaluation Engineering, 42(6):18–23, 2003. CO-DEN ???? ISSN 0014-3316.

Tost:2000:UJC

[TJ00] A. Tost and V. M. Johnson. Using JavaBeans components as accessors to Enterprise JavaBeans components. IBM Systems Journal, 39(2): 293–300, ???? 2000. CODEN IBMSA7. ISSN 0018-8670.

URL http://www.research.
ibm.com/journal/sj/392/
tost.html.

Tan:2007:IIL

[TM07] Gang Tan and Greg Morrisett. Ilea: inter-language analysis across Java and C. ACM SIGPLAN Notices, 42(10):39–56, October 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Trofin:2008:SVC

[TM08] Mircea Trofin and John Murphy. Static verification of component composition in contextual composition frameworks. International Journal on Software Tools for Technology Transfer (STTT), 10(3):247–261, June 2008. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http:

                                        //www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=1433-2779&volume=10&
                                        issue=3&spage=247.

Tarau:2005:SDE

[TMF05] Paul Tarau, Rada Mihalcea, and Elizabeth Figa. Semantic document engineering with WordNet and PageRank. In ACM, editor, Proceedings of the 2005 ACM Symposium on Applied computing, pages 782–786. ACM Press, New York, NY 10036, USA, 2005. ISBN 1-58113-964-0.

Thomas:2003:FJJ

[TMG03] V. Thomas, A. McMullen, and L. Graba. FT-Java: a Java-based framework for fault-tolerant distributed software. Lecture Notes in Computer Science, 2889:899– 911, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thomas:2005:BFJ

[TOG^{+}05] G. Thomas, F. Ogel, A. Galland, B. Folliot, and I. Piumarta. Building a flexible Java runtime upon a flexible compiler. International Journal of Computer Applications, 27(1):27–34, 2005. CODEN ???? ISSN 1206-212X (print), 1925-7074 (electronic). URL https://

www.tandfonline.com/doi/
full/10.1080/1206212X.2005.
11441753.

Tonella:2004:ETC

[Ton04] Paolo Tonella. Evolutionary testing of classes. ACM SIG-SOFT Software Engineering Notes, 29(4):119–128, July 2004. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Topley:2000:CSA

[Top00] Kim Topley. Core Swing: Advanced Programming. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2000. ISBN 0-13-083292-8. xxvii + 930 pp. LCCN A76.9.U83 T66 2000.

Topley:2002:CJJ

[Top02a] Kim Topley. Core JFC: Java foundation classes. Prentice Hall PTR core series. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, second edition, 2002. ISBN 0-13-090581-X. xli + 1655 pp. LCCN QA76.73.J38 T67 2002. US$49.99. URL http:/

/www.phptr.com/ptrbooks/
ptr_013090581X.html. In-

cludes CD-ROM.

Topley:2002:JND

[Top02b] Kim Topley. J2ME in a nutshell: a desktop quick reference. Nutshell handbook. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2002. ISBN 0-596-00253-X. xv + 450 pp. LCCN QA76.73.J38 T67 2002.

Topley:2003:JWS

[Top03] Kim Topley. Java Web services in a nutshell. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2003. ISBN 0-596-00399-4. xviii + 642 pp. LCCN TK5105.88813 .T67 2003. URL http://

                                        www.oreilly.com/catalog/
                                        9780596003999.

Torres:2001:DSD

[Tor01] Mario A. Torres. Developing scalable distributed applications: a generic model implemented in Java. Dr. Dobb’s Journal of Software Tools, 26 (9):21–22, 24–25, September 2001. CODEN DDJOEB. ISSN 1044-789X. URL http:

                                        //www.ddj.com/ftp/2001/
                                        2001_09/scalable.txt.

Teodorescu:2001:UJC

[TP01] Radu Teodorescu and Raju Pandey. Using JIT compilation and configurable runtime systems for efficient deployment of Java programs on ubiquitous devices. Lecture Notes in Computer Science, 2201:76–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2201/22010076.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2201/22010076.
pdf.

Tonella:2002:CSC

[TP02] P. Tonella and A. Potrich.

Cjj: a subset of C++ compliant with Java. Science of Computer Programming, 42 (2):229–271, February 2002. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Tseng:2008:PPD

[TP08] I-Lun Tseng and Adam Postula. Partitioning parameterized 45-degree polygons with constraint programming. ACM Transactions on Design Automation of Electronic Systems, 13(3):52:1–52:??, July 2008. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Tripp:2009:TET

[TPF^{+}09] Omer Tripp, Marco Pistoia, Stephen J. Fink, Manu Sridharan, and Omri Weisman. TAJ: effective taint analysis of web applications. ACM SIGPLAN Notices, 44(6):87– 97, June 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Travers:2000:JQW

[Tra00a] Michael Travers. Java Q&A: What is interactive scripting? Dr. Dobb’s Journal of Software Tools, 25(1):103, 105– 111, January 2000. CODEN DDJOEB. ISSN 1044-789X.

                                        URL http://www.ddj.com/
                                        ftp/2000/2000_01/jqa120.
                                        txt.

Traverso:2000:IAU

[Tra00b] Carlo Traverso, editor. IS-SAC 2000: 7–9 August 2000, University of St. Andrews, Scotland: proceedings of the 2000 International Symposium on Symbolic and Algebraic Computation. ACM Press, New York, NY 10036, USA, 2000. ISBN 1-58113-218-2. LCCN QA76.95.I59 2000. URL

                                        http://www.acm.org/pubs/
                                        contents/proceedings/issac/

345542/. ACM order number

505000.

Tremblett:2000:IJP

[Tre00]  Paul Tremblett.  Instant

JavaServer Pages. Osborne/McGraw-Hill, Berkeley, CA, USA, 2000. ISBN 0-07-212601-

9. xiii + 507 pp. LCCN TK5105.8885.J38 T74 2000. US$49.99. URL http://www.

                                        instantjsp.com/.

Tremblett:2001:IEJ

[Tre01] Paul Tremblett. Instant Enterprise JavaBeans. McGraw-Hill, New York, NY, USA, 2001. ISBN 0-07-212943-3. xxi + 550 pp. LCCN QA76.73.J38 T74 2001.

Tremblett:2002:JUR

[Tre02a] Paul Tremblett. Java and UDDI registries. Dr. Dobb’s Journal of Software Tools, 27(9):34, 37–40, September 2002. CODEN DDJOEB. ISSN 1044-789X. URL

http://www.ddj.com/ftp/
2002/2002_09/juddi.txt;
 http://www.ddj.com/ftp/
2002/2002_09/juddi.zip;
 http://www.ddj.com/ftp/
2002/2002_09/uddi.war.

Tremblett:2002:PTJ

[Tre02b] Paul Tremblett. Programmer’s toolchest: Java’s Generic Connection Framework: Wireless networking for mobile devices. Dr. Dobb’s Journal of Software Tools, 27(4):75–78, 80, April 2002. CODEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/ftp/2002/
2002_04/javagcf.txt;   http:
//www.ddj.com/ftp/2002/
2002_04/javagcf.zip.

Trentini:2002:JBF

[Tre02c] Andrea Trentini. A Javabased framework to support computer-assisted creation of structured XML documents. ACM SIGAPP Applied Computing Review, 10 (1):48–53, April 2002. CO-DEN ???? ISSN 1559-6915 (print), 1931-0161 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/abs/10.1145/
                                        568235.568245.

Tremblett:2003:ISS

[Tre03] P. Tremblett. Incorporating server-side Java 2D into your Web application components is an easy way to deliver dynamic data. Dr. Dobb’s Journal of Software Tools, 28 (10):42–49, 2003. CODEN DDJOEB. ISSN 1044-789X.

Tremblett:2004:JME

[Tre04] P. Tremblett. Java management extensions. Dr. Dobb’s Journal of Software Tools, 29 (7):14–21, 2004. CODEN DDJOEB. ISSN 1044-789X.

Tree:2005:NBC

[Tre05] Tree. NoodleGlue: Bridging C/C++ and Java. C/C++ Users Journal, 23(7):6–13, 2005. CODEN CCUJEX. ISSN 1075-2838.

Trofin:2004:FRRa

[Tro04a] Mircea Trofin. A framework for removing redundant context management services in Enterprise JavaBeans application servers. ACM SIG-PLAN Notices, 39(10):58– 59, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Trofin:2004:FRRb

[Tro04b] Mircea Trofin. A framework for removing redundant context management services in enterprise JavaBeans application servers. ACM SIG-PLAN Notices, 39(10):312– 313, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tatibouet:2003:JCC

[TRVH03] B. Tatibouet, A. Requet, J. C. Voisinet, and A. Hammad. Java card code generation from B specifications. Lecture Notes in Computer Science, 2885:306–318, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

TenEyck:2001:JBM

[TS01] James TenEyck and G. Sampath. A Java-based model of resource sharing among independent users on the Internet. Lecture Notes in Computer Science, 2110:124–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2110/21100124.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2110/21100124.
pdf.

Tilevich:2002:JOA

[TS02] Eli Tilevich and Yannis Smaragdakis. J-orchestra: Automatic Java application partitioning. Lecture Notes in Computer Science, 2374:178– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.
                                        com/link/service/series/
                                        0558/bibs/2374/23740178.
                                        htm;   http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2374/23740178.
                                        pdf.

Tilevich:2004:PED

[TS04] Eli Tilevich and Yannis Smaragdakis. Portable and efficient distributed threads for Java. Lecture Notes in Computer Science, 3231:478– 492, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tilevich:2009:JOE

[TS09] Eli Tilevich and Yannis Smaragdakis. J-Orchestra: Enhancing Java programs with distribution capabilities. ACM Transactions on Software Engineering and Methodology, 19(1):1:1–1:??, August 2009. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Tatsubori:2001:BTD

[TSCI01] Michiaki Tatsubori, Toshiyuki Sasaki, Shigeru Chiba, and Kozo Itano. A bytecode translator for distributed execution of “legacy” Java software. Lecture Notes in Computer Science, 2072:236–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2072/20720236.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2072/20720236.
pdf.

Tanter:2002:AJS

[TSDNP02] Éric Tanter, Marc Ségura-Devillechaise, Jacques Noyé, and José Piquer. Altering Java semantics via bytecode manipulation. Lecture Notes in Computer Science, 2487:283–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/
service/series/0558/bibs/
2487/24870283.htm;   http:
//link.springer.de/link/
service/series/0558/papers/
2487/24870283.pdf.

Tip:2002:PET

[TSL^{+}02] Frank Tip, Peter F. Sweeney, Chris Laffra, Aldo Eisma, and David Streeter. Practical extraction techniques for Java. ACM Transactions on Programming Languages and Systems, 24(6): 625–666, November 2002. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Tip:2003:ELB

[TSL03] F. Tip, P. F. Sweeney, and C. Laffra. Extracting library-based Java applications. Communications of the ACM, 46(8):35–40, 2003. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Tangermann:2004:EIF

[TSL^{+}04] M. Tangermann, C. Schwab, A. Luder, L. Ferrarini, and C. Veber. Encapsulation of IEC 61499 function blocks using real-time Java according to the RTSJ. Lecture Notes in Computer Science, 3292:346– 358, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tyagi:2001:MSM

[TT01] Satyam Tyagi and Paul Tarau. A most specific method finding algorithm for reflection based dynamic Prologto-Java interfaces. Lecture Notes in Computer Science, 1990:322–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.
                                        com/link/service/series/
                                        0558/bibs/1990/19900322.
                                        htm;   http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/1990/19900322.
                                        pdf.

Tansey:2008:ARI

[TT08] Wesley Tansey and Eli Tilevich. Annotation refactoring: inferring upgrade transformations for legacy applications. ACM SIGPLAN Notices, 43(10):295–312, September 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Taboada:2003:PME

[TTD03] G. L. Taboada, J. Tourino, and R. Doallo. Performance modeling and evaluation of Java message-passing primitives on a cluster. Lecture Notes in Computer Science, 2840:29–36, 2003. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tanter:2008:FMA

[TTPN08] Éric Tanter, Rodolfo Toledo, Guillaume Pothier, and Jacques Noyé. Flexible metaprogramming and AOP in Java. Science of Computer Programming, 72(1–2): 22–30, June 1, 2008. CO-DEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Tatlock:2008:DTR

[TTS^{+}08] Zachary Tatlock, Chris Tucker, David Shuffelton, Ranjit Jhala, and Sorin Lerner. Deep typechecking and refactoring. ACM SIGPLAN Notices, 43(10):37–52, September 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tuisku:2004:WJE

[Tui04] M. Tuisku. Wireless Javaenabled MIDP devices as peers in a Grid infrastructure. Lecture Notes in Computer Science, 2970:273–281, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tulachan:2002:DEC

[Tul02] Pravin V. Tulachan. Developing EJB 2.0 Components. Sun BluePrints Program. Sun Microsystems Press, Palo Alto, CA, USA, 2002. ISBN 0-13-034863-5. xxx + 622 pp. LCCN QA76.73.J38T85 2002.

                                        URL http://www.sun.com/
                                        books/catalog/tulachan/.

Tulach:2008:PAD

[Tul08] Jaroslav Tulach. Practical API design: confessions of a Java Framework architect. Apress, Berkeley, CA, USA, 2008. ISBN 1-4302-0973-9. LCCN ????

Tavares:2008:GIO

[TV08] Andre L. C. Tavares and Marco Tulio Valente. A gentle introduction to OSGi. ACM SIGSOFT Software Engineering Notes, 33(5):8:1– 8:??, September 2008. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tyagi:2003:CJD

[TVMB03] Sameer Tyagi, Michael Vorburger, Keiron McCammon, and Heiko Bobzin. Core Java Data Objects. The Sun Microsystems Press Java series. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2003. ISBN 0-13-140731-7. xxx + 491 pp. LCCN QA76.73.J38 C663 2004. US$49.99.

Tanaka:2004:DCR

[TYS04] S. Tanaka, T. Yamada, and S. Shiraishi. Dynamic code repositioning for Java. IEICE Transactions on Information and Systems E Series D, 87 (7):1737–1742, 2004. CODEN ???? ISSN 0916-8532.

Turner:2001:JTV

[TZ01] Joseph A. Turner and Joseph L.

Zachary. Javiva: a tool for visualizing and validating student-written Java programs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):45–49, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Umphress:2004:BJI

[UCJ^{+}04] David A. Umphress, James H.

Cross II, Jhilmil Jain, Nischita Meda, and Larry A. Barowski. Bringing J2ME industry practice into the classroom. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):301–305, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Unkel:2008:AIS

[UL08] Christopher Unkel and Monica S. Lam. Automatic inference of stationary fields: a generalization of Java’s final fields. ACM SIGPLAN Notices, 43(1):183–195, January 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Umar:2002:ERT

[Uma02] S. Umar. Embedded real-time Java in an MPU. Circuit Cellar, 148:12–22, 2002. CODEN ???? ISSN 1528-0608.

UC:2001:EIU

[Uni01] Unicode Consortium, editor. Eighteenth International Unicode Conference (IUC18) Unicode and the Web: the Global Connection, April 24–27, 2001, Hong Kong. The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

5082, E-mail: unicode-inc@
unicode.org, 2001. ISBN

???? LCCN ???? URL

http://www.unicode.org/
iuc/iuc18.

USFS:2002:JGI

[Uni02] United States Forest Service. JMFA — A graphically interactive Java program that fits microfibril angle X-ray diffraction data. Research Note FPL-RN-0283, U.S. Department of Agriculture, Washington, DC, USA, 2002. Shipping List number 2002-0228-M. Shipping List Date: 5-Apr-2002.

USGS:2003:JPU

[Uni03] United States Geological Survey. Java programs for using Newmark’s method and simplified decoupled analysis to model slope performance during earthquakes. Open File Report 03-005, U.S. Geological Survey, Denver, CO, USA, 2003. Includes CD-ROM. Shipping List number 2003-0014-E. Shipping List Date: 20-Mar-2003.

Urbanek:2009:HTS

[Urb09] Simon Urbanek. How to talk to strangers: ways to leverage connectivity between R, Java and Objective C. Computational Statistics, 24(2): 303–311, May 2009. CO-DEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s00180-008-0132-x.

USENIX:2000:UAT

[USE00a]  USENIX, editor.  2000

USENIX Annual Technical Conference: San Diego, CA, USA, June 18–23, 2000. USENIX Association, Berkeley, CA, USA, 2000. ISBN 1-880446-22-7. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/usenix2000.

USENIX:2000:PUT

[USE00b] USENIX, editor. Proceedings of the 7th USENIX Tcl/Tk Conference (Tcl/2k): February 14–18, 2000, Austin, Texas, USA. USENIX Association, Berkeley, CA, USA, 2000. ISBN 1-880446-24-3. LCCN ????

URL http://db.usenix.
org/publications/library/
proceedings/tcl2k/.

USENIX:2000:PFSb

[USE00c] USENIX, editor. Proceedings of the Fourth Symposium on Operating Systems Design and Implementation (OSDI 2000), October 23–25, 2000, San Diego, California, USA. USENIX Association, Berkeley, CA, USA, 2000. ISBN 1-880446-16-2. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/osdi2000/

.

USENIX:2000:PNU

[USE00d] USENIX, editor. Proceedings of the Ninth USENIX Security Symposium, August 14–17, 2000, Denver, Colorado. USENIX Association, Berkeley, CA, USA, 2000. ISBN 1-880446-18-9. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/sec2000.

USENIX:2001:PUC

[USE01a] USENIX, editor. Proceedings of the 6th USENIX Conference on Object-Oriented Technologies and Systems, January 29–February 2, 2001, San Antonio, Texas, USA. USENIX Association, Berkeley, CA, USA, 2001. ISBN 1-880446-12-X. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/coots01/

.

USENIX:2001:UJV

[USE01b]  USENIX, editor.  Usenix

Java Virtual Machine Research and Technology Symposium (JVM ’01). USENIX Association, Berkeley, CA, USA, April 2001.

USENIX:2001:PJV

[USE01c] USENIX Association, editor. Proceedings of the Java Virtual Machine Research and Technology Sy[m]posium (JVM ’01): April 23–24, 2001, Monterey, California, USA. USENIX Association, Berkeley, CA, USA, 2001. ISBN 1-880446-11-1. LCCN QA76.73.J38 J42 2001.

URL http://www.usenix.
org/events/jvm01/. Spon-

sored by the USENIX Association.

USENIX:2002:PJV

[USE02] USENIX Association, editor. Proceedings of the 2nd Java Virtual Machine Research and Technology Symposium (JVM ’02) : August 1–2, 2002, San Francisco, California, USA. USENIX Association, Berkeley, CA, USA, 2002. ISBN 1-931971-01-3. LCCN QA76.73.J38

                                        J42 2002. URL http://www.
                                        usenix.org/publications/
                                        library/proceedings.

Utting:2006:PIT

[Utt06] Ian Utting. Problems in the initial teaching of programming using Java: the case for replacing J2SE with J2ME. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38 (3):193–196, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Vermeulen:2000:EJS

[VAB^{+}00] Allan Vermeulen, Scott W. Ambler, Greg Bumgardner, Eldon Metz, Trevor Misfeldt, Jim Shur, and Patrick Thompson. The Elements of Java Style. Cambridge University Press, Cambridge, UK, 2000. ISBN 0-521-77768-2. xiii + 128 pp. LCCN QA76.73.J38 E44 2000. US$12.95.

VanCamp:2004:TNS

[Van04] S. VanCamp. Tech-nically speaking: Sun Microsystems is banking on its Java platform to restore a measure of respectability. Brandweek, 45 (30):14–19, 2004. CODEN ???? ISSN 1064-4318.

Vaughan:2003:IME

[Vau03a] J. Vaughan. Improvements mark evolving Java app servers. Application Development Trends, 10(2):31–34, 2003. CODEN ADTRF4. ISSN 1073-9564.

VaughanNichols:2003:BUJ

[Vau03b] S. J. VaughanNichols. The battle over the universal Java IDE. Computer, 36(4):21– 23, 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Villazon:2001:PRR

[VB01a] Alex Villazón and Walter Binder. Portable resource reification in Javabased mobile agent systems. Lecture Notes in Computer Science, 2240:213–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2240/22400213.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2240/22400213.
pdf.

Vitek:2001:CTJ

[VB01b] Jan Vitek and Boris Bokowski. Confined types in Java. Software—Practice and Experience, 31(6):507–532, May 2001. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

http://www3.interscience.
wiley.com/cgi-bin/abstract/
78003101/START;   http:/
/www3.interscience.wiley.
com/cgi-bin/fulltext?ID=
78003101&PLACEBO=IE.pdf.

vanDoorn:2000:SJV

[vD00] Leendert van Doorn. A secure Java^{T M} Virtual Machine. In USENIX [USE00d], page ?? ISBN 1-880446-18-9. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/sec2000/
vandoorn.html.

vonDincklage:2004:CJC

[vD04] D. vonDincklage and A. Diwan. Converting Java classes to use generics. ACM SIG-PLAN Notices, 39(10):1–14, 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

vandenBercken:2000:JXP

[vdBDS00] Jochen van den Bercken, Jens-Peter Dittrich, and Bernhard Seeger. javax.XXL: a prototype for a library of query processing algorithms. In Chen et al. [CNB00], page 588. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1

.A87. URL http://www.
acm.org/pubs/citations/
proceedings/mod/342009/
p588-van_den_bercken/.

vandenBerg:2001:LCJ

[vdBJ01] Joachim van den Berg and Bart Jacobs. The LOOP compiler for Java and JML. Lecture Notes in Computer Science, 2031:299–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2031/20310299.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2031/20310299.
pdf.

vandenBerg:2001:FSV

[vdBJP01] Joachim van den Berg, Bart Jacobs, and Erik Poll. Formal specification and verification of JavaCard’s application identifier class. Lecture Notes in Computer Science, 2041:137–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2041/20410137.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2041/20410137.
pdf.

vanderLinden:2002:JJ

[vdL02] Peter van der Linden. Just Java 2. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, fifth edition, 2002. ISBN 0-13-032072-2. xxxv + 1098 pp. LCCN QA76.73.J38 V3625 2002b. US$49.99. URL http://www.

                                        phptr.com/ptrbooks/ptr_
                                        0130320722.html;   http:
                                        //www.sun.com/books/catalog/
                                        vanderlinden6/index.html.

Vincenzi:2006:EST

[VDMW06] A. M. R. Vincenzi, M. E. Delamaro, J. C. Maldonado, and W. E. Wong. Establishing structural testing criteria for Java bytecode. Software —Practice and Experience, 36(14):1513–1541, November 25, 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

VanderHeyden:2001:CJC

[VDPC01] William VanderHeyden, Edward Dendy, and Nely Padial-Collins. CartaBlanca — A Java-component-based systems simulation tool for coupled non-linear physics on unstructured grids. In ACM [ACM01b], pages 134–

142. ISBN 1-58113-359-

6. LCCN QA76.9.O35 A26

                                        2001. URL http://www.
                                        philippsen.com/JGI2001/
                                        camerareadyabstracts/3.
                                        html;   http://www.philippsen.
                                        com/JGI2001/finalpapers/
                                        18500134.pdf.

VanderHeyden:2003:CPJ

[VDPC03] W. B. VanderHeyden, E. D.

Dendy, and N. T. Padial-Collins. CartaBlanca — a pure-Java, component-based systems simulation tool for coupled nonlinear physics on unstructured grids — an update. Concurrency and Computation: Practice and Experience, 15(3–5):431–458, March/April 2003. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pol:2002:FSJ

[vdPE02] Jaco van de Pol and Miguel Valero Espada. Formal specification of JavaSpaces^{TM} Architecture using μCRL. Lecture Notes in Computer Science, 2315:274–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2315/23150274.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2315/23150274.
pdf.

vanderSpek:2005:SER

[vdSPP05] P. van der Spek, N. Plat, and C. Pronk. Syntax error repair for a Java-based parser generator. ACM SIGPLAN Notices, 40(4):47–50, April 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Venstermans:2006:BVB

[VED06] Kris Venstermans, Lieven Eeckhout, and Koen De Bosschere. 64-bit versus 32-bit Virtual Machines for Java. Software—Practice and Experience, 36(1):1–26, January 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Venstermans:2007:JOH

[VED07] Kris Venstermans, Lieven Eeckhout, and Koen De Bosschere. Java object header elimination for reduced memory consumption in 64-bit virtual machines. ACM Transactions on Architecture and Code Optimization, 4(3): 17:1–17:??, September 2007. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Veldhuizen:2001:JWY

[Vel01] Todd L. Veldhuizen. Just when you thought your little language was safe: “expression templates” in Java. Lecture Notes in Computer Science, 2177:188–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.
                                        com/link/service/series/
                                        0558/bibs/2177/21770188.
                                        htm;   http://link.springer-
                                        ny.com/link/service/series/

0558/papers/2177/21770188.
pdf.

Veldema:2001:ROJ

[VHBB01] Ronald Veldema, Rutger Hofman, Raoul Bhoedjang, and Henri Bal. Runtime optimizations for a Java DSM implementation. In ACM [ACM01b], pages 153–

162. ISBN 1-58113-359-

6. LCCN QA76.9.O35 A26

2001. URL http://www.
philippsen.com/JGI2001/
camerareadyabstracts/41.
html;   http://www.philippsen.
com/JGI2001/finalpapers/
18500153.ps.

Veldema:2003:RTO

[VHBB03] R. Veldema, R. F. H. Hofman, R. A. F. Bhoedjang, and H. E. Bal. Run-time optimizations for a Java DSM implementation. Concurrency and Computation: Practice and Experience, 15(3–5):299– 316, March/April 2003. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Vincent:2001:AIB

[VHL01] Regis Vincent, Bryan Horling, and Victor Lesser. An agent infrastructure to build and evaluate multi-agent systems: The Java agent framework and multi-agent system simulator. Lecture Notes in Computer Science, 1887:102– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.
                                        com/link/service/series/
                                        0558/bibs/1887/18870102.
                                        htm;   http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/1887/18870102.
                                        pdf.

vanHeiningen:2008:BMD

[vHMB08] Willem van Heiningen, Steve MacDonald, and Tim Brecht. Babylon: middleware for distributed, parallel, and mobile Java applications. Concurrency and Computation: Practice and Experience, 20 (10):1195–1224, July 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Vieregger:2003:PRP

[Vie03]  C. Vieregger.  Product

roundup: a peck of Java portlets. Software Development, 11(3):27–29, 2003. CO-DEN ???? ISSN 1070-8588.

Vilar:2000:JQW

[Vil00] Alexandre Sieira Vilar. Java Q&A: What’s object pooling all about? Dr. Dobb’s Journal of Software Tools, 25(8):123–124, 126, 128, 130, August 2000. CODEN DDJOEB. ISSN 1044-789X. URL

                                        http://www.ddj.com/ftp/
                                        2000/2000_08/jqa800.txt;
                                         http://www.ddj.com/ftp/
                                        2000/2000_08/jqa800.zip.

Villalon:2008:HDD

[Vil08]  Elena Villalon.  High-

dimensionality data reduction with Java. Computing in Science and Engineering, 10(5): 64–69, September/October 2008. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Velazquez-Iturbide:2008:SAS

[VIPCUF08] J. Ángel Velázquez-Iturbide, Antonio Pérez-Carrasco, and Jaime Urquiza-Fuentes. SRec: an animation system of recursion for algorithm courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40 (3):225–229, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Viroli:2003:TPA

[Vir03] M. Viroli. A type-passing approach for the implementation of parametric methods in Java. The Computer Journal, 46(3):263–294, 2003. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Virkus:2005:PJP

[Vir05] Robert Virkus. Pro J2ME Polish: open source wireless Java tools suite. Apress, Berkeley, CA, USA, 2005. ISBN 1-59059-503-3. xxi + 450 pp. LCCN QA76.73.J38

V57 2005. URL http://www.
loc.gov/catdir/enhancements/
fy0663/2005016571-d.html;
 http://www.springerlink.
com/content/p7059524m76vj817/

.

Veldema:2001:OJS

[VKB01] Ronald S. Veldema, Thilo Kielmann, and Henri E. Bal. Optimizing Java-specific overheads: Java at the speed of C? Lecture Notes in Computer Science, 2110:685–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2110/21100685.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2110/21100685.
pdf.

Vijaykrishnan:2001:EBJ

[VKK^{+}01] N. Vijaykrishnan, M. Kandemir, S. Kim, S. Tomar, A. Sivasubramaniam, and M. J. Irwin. Energy behavior of Java applications from the memory perspective. In USENIX Association [USE01c], page ?? ISBN 1-880446-11-1. LCCN QA76.73.J38 J42 2001. URL

http://www.usenix.org/publications/
library/proceedings/jvm01/
vijaykrishnan.html. Spon-

sored by the USENIX Association.

Viswanathan:2000:JVM

[VL00] D. Viswanathan and S. Liang.

Java Virtual Machine Profiler Interface. IBM Systems Journal, 39(1):82–95, ???? 2000. CODEN IBMSA7. ISSN 0018-8670.

URL http://www.research.
ibm.com/journal/sj/391/
viswanathan.html.

vonLaszewski:2001:JCG

[vLFGL01] Gregor von Laszewski, Ian Foster, Jarek Gawor, and Peter Lane. A Java commodity grid kit. Concurrency and Computation: Practice and Experience, 13(8–9):645– 662, July/August 2001. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (elec-

tronic). URL http://www3.
interscience.wiley.com/
cgi-bin/abstract/84503216/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=84503216&PLACEBO=IE.
pdf.

vonLaszewski:2002:FJC

[vLGL^{+}02] Gregor von Laszewski, Jarek Gawor, Peter Lane, Nell Rehn, and Mike Russell. Features of the Java Commodity Grid Kit. Concurrency and Computation: Practice and Experience, 14(13–15):1045– 1055, November/December 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

vonLaszewski:2005:WCJ

[vLH05] Gregor von Laszewski and Mike Hategan. Workflow concepts of the Java CoG Kit. Journal of Grid Computing, 3(3–4):239–258, September 2005. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=1570-7873&volume=3&
issue=3&spage=239.

VanCappellen:2009:XXJ

[VLMO09] Marc Van Cappellen, Zhen Hua Liu, Jim Melton, and Maxim Orgiyan. XQJ: XQuery Java API is completed. SIGMOD Record (ACM Special Interest Group on Management of Data), 38(4):7–13, December 2009. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

vonLaszewski:2001:GBA

[vLSM01] Gregor von Laszewski, Kazuyuki Shudo, and Yoichi Muraoka. Grid-based asynchronous migration of execution context in Java virtual machines. Lecture Notes in Computer Science, 1900:22– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/1900/19000022.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/1900/19000022.
pdf.

Viega:2000:SSJ

[VMMF00] John Viega, Tom Mutdosch, Gary McGraw, and Edward W. Felten. Statically scanning Java code: Finding security vulnerabilities. IEEE Software, 17(5):68– 74, September/October 2000. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:

//computer.org/software/
so/s5068abs.htm;   http:
//dlib.computer.org/so/
books/so2000/pdf/s5068.
pdf.

vandenBrand:2005:GES

[vMV05] M. vandenBrand, P. E.

Moreau, and J. Vinju. Generator of efficient strongly typed abstract syntax trees in Java. IEE Proceedings Software, 152(2):70–78, 2005. CODEN ???? ISSN 1462-5970.

Vincenzi:2005:CTJ

[VMWD05] A. M. Vincenzi, J. C. Maldonado, W. E. Wong, and M. E. Delamaro. Coverage testing of Java programs and components. Science of Computer Programming, 56 (1–2):211–230, 2005. CO-DEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Viroli:2000:PPJ

[VN00] Mirko Viroli and Antonio Natali. Parametric polymorphism in Java: an approach to translation based on reflective features. ACM SIG-PLAN Notices, 35(10):146– 165, October 2000. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
proceedings/oops/353171/
p146-viroli/.

Vaughan-Nichols:2003:TNB

[VN03] Steven J. Vaughan-Nichols. Technology news: The battle over the universal Java IDE. Computer, 36(4): 21–23, April 2003. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/co/
2003/04/r4021.htm;   http:
//csdl.computer.org/dl/
mags/co/2003/04/r4021.pdf.

vanNieuwpoort:2001:SEP

[vNKB01] Rob V. van Nieuwpoort, Thilo Kielmann, and Henri E. Bal. Satin: Efficient parallel divide-and-conquer in Java. Lecture Notes in Computer Science, 1900:690–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/1900/19000690.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/1900/19000690.
pdf.

vanNieuwpoort:2005:SSE

[vNMKB05] Rob van Nieuwpoort, Jason Maassen, Thilo Kielmann, and Henri E. Bal. Satin: Simple and efficient Java-based Grid programming. Scalable Computing: Practice and Experience, 6 (3):19–32, September 2005. CODEN ???? ISSN 1895-

1767. URL http://www.
scpe.org/vols/vol06/no3/
SCPE_6_3_03.pdf;   http:
//www.scpe.org/vols/vol06/
no3/SCPE_6_3_03.zip.

vanNieuwpoort:2005:IFE

[vNMW^{+}05] Rob V. van Nieuwpoort, Jason Maassen, Gosia Wrzesińska, Rutger F. H. Hofman, Ceriel J. H. Jacobs, Thilo Kielmann, and Henri E. Bal. Ibis: a flexible and efficient Java-based Grid programming environment. Concurrency and Computation: Practice and Experience, 17 (7–8):1079–1107, June/July 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

vonOheimb:2001:HLJ

[vO01] David von Oheimb. Hoare logic for Java in Isabelle/ HOL. Concurrency and Computation: Practice and Experience, 13(13):1173–1214, November 2001. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (elec-

tronic). URL http://www3.
interscience.wiley.com/

                                        cgi-bin/abstract/88011338/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=88011338&PLACEBO=IE.
                                        pdf.

Vogels:2003:HNC

[Vog03] Werner Vogels. HPC.NET — are CLI-based virtual machines suitable for high performance computing? In ACM [ACM03b], page ?? ISBN 1-58113-695-1. LCCN ???? URL http:/

                                        /www.sc-conference.org/
                                        sc2003/inter_cal/inter_
                                        cal_detail.php?eventid=
                                        10710#2;   http://www.sc-
                                        conference.org/sc2003/paperpdfs/
                                        pap251.pdf.

Oheimb:2002:HLN

[vON02a] David von Oheimb and Tobias Nipkow. Hoare logic for NanoJava: Auxiliary variables, side effects, and virtual methods revisited. Lecture Notes in Computer Science, 2391:89–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.
                                        com/link/service/series/
                                        0558/bibs/2391/23910089.
                                        htm;   http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2391/23910089.
                                        pdf.

vonOheimb:2002:HLN

[vON02b] David von Oheimb and Tobias Nipkow. Hoare logic for NanoJava: Auxiliary variables, side effects, and virtual methods revisited. Lecture Notes in Computer Science, 2391:89–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2391/23910089.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2391/23910089.
pdf.

Vormoor:2001:QEI

[Vor01] Oliver Vormoor. Quick and easy interactive molecular dynamics using Java3D. Computing in Science and Engineering, 3(5):98–104, September/October 2001. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

computer.org/cise/cs2001/
c5098abs.htm;   http://
dlib.computer.org/cs/books/
cs2001/pdf/c5098.pdf.

Vivanco:2005:SCJ

[VP05] Rodrigo A. Vivanco and Nicolino J. Pizzi. Scientific computing with Java and C++: a case study using functional magnetic resonance neuroimages. Software—Practice and Experience, 35(3):237–254, March 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Visser:2004:TIG

[VPK04] W. Visser, C. S. Pasareanu, and S. Khurshid. Test input generation with Java PathFinder. Software Engineering Notes, 29(4):97–107, 2004. CODEN ???? ISSN 0163-5948.

Vrba:2003:JBA

[Vrb03] P. Vrba. Java-based agent platform evaluation. Lecture Notes in Computer Science, 2744:47–58, 2003. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

vanReeuwijk:2001:SEJ

[vRKS01] Cees van Reeuwijk, Frits Kuijlman, and Henk J. Sips. Spar: an extension of Java for scientific computation. In ACM [ACM01b], pages 58–67. ISBN 1-58113-359-

6. LCCN QA76.9.O35 A26

                                        2001. URL http://www.
                                        philippsen.com/JGI2001/
                                        camerareadyabstracts/11.
                                        html;   http://www.philippsen.
                                        com/JGI2001/finalpapers/
                                        18500058.dvi.

vanReeuwijk:2003:SSE

[vRKS03] C. van Reeuwijk, F. Kuijlman, and H. J. Sips. Spar: a set of extensions to Java for scientific computation. Concurrency and Computation: Practice and Experience, 15 (3–5):277–297, March/April 2003. CODEN CCPEBO.

ISSN 1532-0626 (print), 1532-0634 (electronic).

vanReeuwijk:2005:ATJ

[vRS05] C. van Reeuwijk and H. J.

Sips. Adding tuples to Java: a study in lightweight data structures. Concurrency and Computation: Practice and Experience, 17(5–6):423– 438, April/May 2005. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Vollmar:2006:MEO

[VS06] Kenneth Vollmar and Pete Sanderson. MARS: an education-oriented MIPS assembly language simulator. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38 (1):239–243, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Vakali:2001:JBM

[VT01] A. I. Vakali and E. D.

Terzi. A Java-based model for I/O scheduling in tertiary storage subsystems. International Journal of Computer Applications, 23(1): 45–50, 2001. ISSN 1206-212X (print), 1925-7074 (electronic). URL https://

www.tandfonline.com/doi/
full/10.1080/1206212X.2001.
11441633.

Vaziri:2006:ASC

[VTD06] Mandana Vaziri, Frank Tip, and Julian Dolby. Associating synchronization constraints with data in an objectoriented language. ACM SIG-PLAN Notices, 41(1):334– 345, January 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

vanTonder:2008:JLD

[vTNC08] Martin van Tonder, Kevin Naude, and Charmain Cilliers. Jenuity: a lightweight development environment for intermediate level programming courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):58– 62, September 2008. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Vandewoude:2002:JID

[VUPB02] Y. Vandewoude, D. Urting, K. Pelckmans, and Y. Berbers. A Java-interface to digital cameras. Applied Informatics, 2:113–118, 2002. CODEN ???? ISSN 1027-2666.

VahaSipila:2005:BCC

[VV05] A. VahaSipila and T. Virtanen. BT-Crowds: Crowdsstyle anonymity with Bluetooth and Java. In Proceedings of the Annual Hawaii International Conference on System Sciences, volume CONF38, page 320. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2005. CODEN ???? ISSN 1060-3425.

VanDenBossche:2005:OCI

[VVG^{+}05] B. VanDenBossche, K. Van-Boxstael, N. Goeminne, F. Gielen, and P. Demeester. An OSGi-compatible implementation of a Java resource monitor [5684-20]. Proceedings- Spie the International Society for Optical Engineering, 5684:181–189, 2005. CODEN ???? ISSN 0277-786X (print), 1996-756X (electronic).

Vieira:2004:LEH

[VVV04] Luiz Filipe M. Vieira, Marcos Augusto M. Vieira, and Newton J. Vieira. Language emulator, a helpful toolkit in the learning process of computer theory. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):135–139, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

VanHoof:2005:MES

[VWS^{+}05] V. VanHoof, A. Wormek, S. Schleutermann, T. Schumacher, O. Lothaire, and C. Trendelenburg. Medical expert systems developed in j.MD, a Java based expert system shell application in clinical laboratories. Studies in Health Technology and Informatics, 107(1):89– 93, 2005. CODEN ???? ISSN 0926-9630.

Vilner:2007:FCC

[VZGE07] Tamar Vilner, Ela Zur, and Judith Gal-Ezer. Fundamental concepts of CS1: procedural vs. object oriented paradigm — a case study. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39 (3):171–175, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Wahli:2004:WSJ

[W^{+}04] Ueli Wahli et al., editors. WebSphere Studio 5.1.2 JavaServer Faces and Service Data Objects. IBM redbooks. IBM International Technical Support Organization, ????, July 2004. ISBN 0-7384-9041-5. xxiv + 566 pp. LCCN TK5105.8885.W43 W4322 2004. Publication number SG24-6361-00.

Waldo:2001:JS

[WA01] Jim Waldo and Ken Arnold. The Jini specifications. Jini technology series. Addison-Wesley, Reading, MA, USA, second edition, 2001. ISBN 0-201-72617-3. xxvii + 645 pp. LCCN QA76.9.D5 J56 2001. Rev. ed of: The Jini specification / Ken Arnold … [et al]. c1999.

Williams:2004:WLC

[WA04] T. Williams and J. Anderson.

Write less code and deliver Java apps faster with Eclipse 3.0. Electronic Design, 52 (21):71–72, 2004. CODEN ELODAW. ISSN 0013-4872.

Webb:2004:LJB

[WAB^{+}04] S. C. Webb, A. Attwood, T. Brooks, T. Freeman, P. Gardner, C. Pritchard, D. Williams, P. Underhill, M. A. Strivens, and A. Greenfield. LIMaS: the JAVA-based application and database for microarray experiment tracking. Mammalian Genome, 15 (9):740–747, 2004. CODEN ???? ISSN 0938-8990.

Walnes:2003:JOS

[WACBL03] Joe Walnes, Ara Abrahamian, Mike Cannon-Brookes, and Patrick A. Lightbody. Java Open source programming: with XDoclet, JUnit, WebWork, Hibernate (Java Open Source Library). John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 2003. ISBN 0-471-46362-0. xx + 459 pp. LCCN QA76.73.J38 J3785

Wadler:2000:GGJ

[Wad00] Philip Wadler. GJ: A Generic Java. Dr. Dobb’s Journal of Software Tools, 25(2):23–26, 28, February 2000. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2000/2000_02/gj.txt.

Wallach:2000:SSM

[WAF00] Dan S. Wallach, Andrew W. Appel, and Edward W. Felten. SAFKASI: a security mechanism for languagebased systems. ACM Transactions on Software Engineering and Methodology, 9 (4):341–378, October 2000. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL

http://www.acm.org/pubs/
articles/journals/tosem/
2000-9-4/p341-wallach/p341-
wallach.pdf;   http://www.
acm.org/pubs/citations/
journals/tosem/2000-9-4/
p341-wallach/.

Welch:2002:CNJ

[WAF02] P. H. Welch, J. R. Aldous, and J. Foster. CSP networking for Java (JCSP.net). Lecture Notes in Computer Science, 2330:695–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2330/23300695.
htm;   http://link.springer-
ny.com/link/service/series/

0558/papers/2330/23300695.
pdf.

Walsh:2002:MJA

[Wal02a] Aaron E. Walsh. The MPEG-4 Java API and MPEGlets. Dr. Dobb’s Journal of Software Tools, 27(4):40, 42– 45, April 2002. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2002/2002_04/jmpeg.
txt;   http://www.ddj.com/
ftp/2002/2002_04/jmpeg.
zip.

Walsh:2002:USG

[Wal02b] Aaron E. Walsh. Understanding scene graphs. Dr. Dobb’s Journal of Software Tools, 27(7):17, 20–26, July 2002. CODEN DDJOEB. ISSN 1044-789X. URL

http://www.ddj.com/ftp/
2002/2002_07/scene.txt;
 http://www.ddj.com/ftp/
2002/2002_07/scene.zip.

Walsh:2003:CJG

[Wal03a] A. E. Walsh. Creating Java Grid services. Dr. Dobb’s Journal of Software Tools, 28 (9):18–23, 2003. CODEN DDJOEB. ISSN 1044-789X.

Walsh:2003:JWS

[Wal03b] Aaron E. Walsh. J2Ee 1.4 Web services. Dr. Dobb’s Journal of Software Tools, 28 (4):34, 36, 38, 39, April 2003. CODEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/documents/s=
7827/ddj0304d/.

Walsh:2003:JP

[Wal03c] Aaron E. Walsh. JavaServer Pages 2.0. Dr. Dobb’s Journal of Software Tools, 28(7):48, 50, 52–57, July 2003. CODEN DDJOEB. ISSN 1044-789X.

                                        URL http://www.ddj.com/
                                        ftp/2003/2003_07/jsp20.
                                        zip.

Wampler:2002:EOO

[Wam02] Bruce E. Wampler. The essence of object-oriented programming with Java and UML. Addison-Wesley, Reading, MA, USA, 2002. ISBN 0-201-73410-9. xix + 290 pp. LCCN QA76.64 .W359 2002.

Wang:2002:UJH

[Wan02a]  A. J. A. Wang.  Us-

ing Java Hartstone benchmark in a real-time systems course. Frontiers in Education Conference, 3(??): S2F–25–??, 2002. CODEN PFECDR. ISSN 0190-5848.

Wang:2002:CSP

[Wan02b] G. Wang. Ch solves portability headaches. IEEE Spectrum, 39(6):59, June 2002. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wang:2003:BAD

[Wan03a] M. Wang. E-business application development with Java technology and Oracle: The Fortune Invest Inc. case. Journal of Information Systems Education, 14(3):293– 300, 2003. CODEN ???? ISSN 1055-3096.

Wang:2003:JOO

[Wan03b] Paul S. Wang. Java with object-oriented programming. Thomson/Brooks Cole, Pacific Grove, CA, USA, second edition, 2003. ISBN 0-534-39276-8. xiv + 575 pp. LCCN QA76.73.J38 W364 2003.

Wang:2003:MLJ

[Wan03c] W. Wang. On memory leaks in Java programs. Journal — Zhejiang University of Technology, 31(4):405–409, 2003. CODEN ???? ISSN 1006-4303.

Wang:2004:UJL

[Wan04] H. Wang. Use of Java language on the Web of real-time observed tide data. Marine Science Bulletin, 23(2):51–55, 2004. CODEN ???? ISSN 1001-6392.

Wang:2005:MDT

[Wan05] W. Wang. Method of data transformation between applications in Java. Journal — Anhui University of Technology, 22(1):60–62, 2005. CO-DEN ???? ISSN 1671-7872.

Warnes:2002:HJL

[War02] Gregory R. Warnes. HYDRA:

a Java library for Markov chain Monte Carlo. Journal of Statistical Software, 7 (4):1–32, 2002. CODEN JS-SOBK. ISSN ???? URL

http://software.biostat.
washington.edu/statsoft/
MCMC/Hydra;   http://software.
biostat.washington.edu/
statsoft/MCMC/Hydra/INSTALL.
html;   http://www.jstatsoft.
org/v07/i04;   http://www.
jstatsoft.org/v07/i04/Description.
pdf;   http://www.jstatsoft.
org/v07/i04/Hydra_1.0.0.
jar;   http://www.jstatsoft.
org/v07/i04/Hydra_1.0.0.
tgz;   http://www.jstatsoft.
org/v07/i04/updates;   http:
//www.jstatsoft.org/v07/
i04/UserGuide.pdf.

Watari:2002:FTU

[Wat02] S. Watari. A forecast tool using JavaScript for predicting arrival time of interplanetary disturbances to the Earth. Journal — Communications Research Laboratory, 49(4): 47–54, 2002. CODEN ???? ISSN 0914-9260.

Wayne:2003:CNK

[Way03] R. Wayne. Curiosity never killed the programmer: PE Explorer helps you delve into the nitty-gritty inside Windows files, Browsersoft’s eQ! Foundation provides a basis to build your Java on, and Red Gate’s ANTS profiler offers some much-needed common sense aimed at .NET. Software Development, 11(5): 17–20, 2003. CODEN ???? ISSN 1070-8588.

Wayne:2005:PYB

[Way05] R. Wayne. Put your best Java forward: Use the power of Java and the agility of a Web app with Canoo’s UltraLight-Client, deconstruct sobriquets with Language Analysis Systems’ Name Parser, and craft and edit with Effexis’s SDE. Software Development, 13(3): 22–26, 2005. CODEN ???? ISSN 1070-8588.

Watt:2000:PLP

[WB00] David A. (David Anthony)

Watt and Deryck F. Brown. Programming language processors in Java: compilers and interpreters. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2000. ISBN 0-13-025786-9 (case). xvi + 436 pp. LCCN QA76.73.J38 W385 2000.

Watt:2001:JCI

[WB01] David A. (David Anthony)

Watt and Deryck F. Brown. Java collections: an introduction to abstract data types, data structures, and algorithms. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 2001. ISBN 0-471-89978-X (paperback). 549 pp. LCCN QA76.73.J38 W384 2001.

Walls:2005:SA

[WB05] Craig Walls and Ryan Breidenbach. Spring in action. Manning Publications, Greenwich, CT, USA, 2005. ISBN 1-932394-35-4 (paperback). xxviii + 444 pp. LCCN QA76.73.J38 W342

                                        2005. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0626/2005276355-b.html;
                                         http://www.loc.gov/catdir/
                                        enhancements/fy0626/2005276355-
                                        d.html.

Walls:2008:SA

[WB08] Craig Walls and Ryan Breidenbach. Spring in action. Manning Publications, Greenwich, CT, USA, second edition, 2008. ISBN (paperback), 1-933988-13-4 (paperback). xxxiv + 730 pp. LCCN QA76.73.J38 W342 2008.

Winter:2006:TPC

[WBF^{+}06] Victor L. Winter, Jason Beranek, Fares Fraij, Steve Roach, and Greg Wickstrom. A transformational perspective into the core of an abstract class loader for the SSP. ACM Transactions on Embedded Computing Systems, 5(4):773–818, November 2006. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Wu:2005:PTT

[WBGM05] Xiaoqing Wu, Barrett R. Bryant, Jeff Gray, and Marjan Mernik. Pattern transformation for two-dimensional separation of concerns. ACM SIGSOFT Software Engineering Notes, 30(4):1–5, July 2005. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Weis:2001:SYH

[WBL01] Rüdiger Weis, Bastiaan Bakker, and Stefan Lucks. Security on your hand: Secure filesystems with a “noncryptographic” JAVA-ring. Lecture Notes in Computer Science, 2041:151–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2041/20410151.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2041/20410151.
pdf.

Walsh:2001:CW

[WBS01] Aaron E. Walsh and Mikael Bourges-Sevenier. Core Web 3D. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2001. ISBN 0-13-085728-9. l + 1088 + 16 pp. LCCN T385 .W364

2001. URL http://www.
phptr.com/ptrbooks/ptr_
0130857289.html;   http:
//www.virtuworlds.com/3DeZine_
01/features/art_coreweb3d.
html.

Welsh:2000:ARS

[WC00a] Matt Welsh and David Culler.

Achieving robust, scalable cluster I/O in Java. Lecture Notes in Computer Science, 1915:16–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.
                                        com/link/service/series/
                                        0558/bibs/1915/19150016.
                                        htm;   http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/1915/19150016.
                                        pdf.

Welsh:2000:JEE

[WC00b] Matt Welsh and David Culler. Jaguar: enabling efficient communication and I/O in Java. Concurrency: Practice and Experience, 12(7): 519–538, May 2000. CO-DEN CPEXEI. ISSN 1040-

                                        3108. URL http://www3.
                                        interscience.wiley.com/
                                        cgi-bin/abstract/72516211/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=72516211&PLACEBO=IE.
                                        pdf.

Wei:2008:JDP

[WC08] Wen Hsiang Wei and Guan Jhih Chen. JavaStatSoft: design patterns and features. Computational Statistics, 23(2): 235–251, April 2008. CO-DEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s00180-007-0092-6.

Wells:2004:LIJ

[WCC04] G. C. Wells, A. G. Chalmers, and P. G. Clayton. Linda implementations in Java for concurrent systems. Concurrency and Computation: Practice and Experience, 16(10):1005– 1022, August 25, 2004. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Wei:2005:SOJ

[WCCL05] Pu-Chen Wei, Chung-Hsin Chen, Cheng-Wei Chen, and Jenq-Kuen Lee. Support and optimization of Java RMI over a Bluetooth environment. Concurrency and Computation: Practice and Experience, 17(7–8):967– 989, June/July 2005. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Weerawarana:2001:BML

[WCD^{+}01] Sanjiva Weerawarana, Francisco Curbera, Matthew J. Duftler, David A. Epstein, and Joseph Kesselman. Bean Markup Language: a composition language for JavaBeans components. In USENIX [USE01a], page ?? ISBN 1-880446-12-X. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/coots01/
weerawarana.html.

Wyman:2007:ZZI

[WCK^{+}07] L. W. Wyman, J. Castaño, J. P. Kubala, R. J. Maddison, B. R. Pierce, and R. R. Rogers. zAAPs and zIIPs: Increasing the strategic value of System z. IBM Journal of Research and Development, 51 (1/2):77–86, January /March 2007. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http:

                                        //www.research.ibm.com/
                                        journal/rd/511/wyman.html.

Walsh:2000:JB

[WCS00] Aaron Walsh, Justin Couch, and Daniel H. Steinberg. Java 2 Bible. IDG Books, San Mateo, CA, USA, 2000. ISBN 0-7645-4632-5. xxx + 905 pp. LCCN QA76.73.J38 W3575 2000. US$40.

Weltman:2000:LPJ

[WD00] Rob Weltman and Tony Dahbura. LDAP programming with Java. Addison-Wesley, Reading, MA, USA, 2000. ISBN 0-201-65758-9. xxi + 692 pp. LCCN QA76.73.J38 W47 2000.

Willrich:2002:MAH

[WDSD02] R. Willrich, P. De Saqui-Sannes, P. Senac, and M. Diaz. Multimedia authoring with hierarchical timed stream Petri nets and Java. Multimedia Tools and Applications, 16(1–2):7–27, January 2002. CODEN MTAPFB. ISSN 1380-7501.

Wear:2000:JSW

[Wea00] William Wear. Java Security On the Web. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2000. ISBN 0-13-010793-X. ???? pp. LCCN ???? US$50.

Weaver:2004:ECS

[Wea04] Alfred C. Weaver. Electronic commerce software laboratory. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):453–456, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Weaver:2007:JSD

[Wea07] James L. Weaver. JavaFX script: dynamic Java scripting for rich Internet/clientside applications. Apress, Berkeley, CA, USA, 2007. ISBN 1-59059-945-4. xi + 201 pp. LCCN ????

Weisser:2001:PCL

[Wei01] M. Weisser. Programming for corpus linguistics: How to do text analysis with Java. Literary and Linguistic Computing, 16(4):498–500, December 2001. CODEN LLCOEI. ISSN 0268-1145.

Weiss:2002:DSP

[Wei02a] Mark Allen Weiss. Data structures and problem solving using Java. Addison-Wes-ley, Reading, MA, USA, second edition, 2002. ISBN 0-201-74835-5. xxxiv + 886 pp. LCCN QA76.73.J38 W45 2002.

Weissinger:2002:DJC

[Wei02b] A. Keyton Weissinger. Developing JSP custom tag libraries. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2002. ISBN 0-596-00155-X. ???? pp. LCCN QA76.73.J38 W46 2002. URL

http://www.oreilly.com/
catalog/9780596001551.

Weiss:2004:JCE

[Wei04] Jason Weiss. Java Cryptography Extensions: Practical Guide for Programmers. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2004. ISBN 0-12-742751-1. xv + 158 pp. LCCN QA76.73.J38 W445 2004. US$21.95.

Welch:2002:POD

[Wel02] P. H. Welch. Process oriented design for Java: Concurrency for all. Lecture Notes in Computer Science, 2330:687– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2330/23300687.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2330/23300687.
pdf.

Wellings:2003:JAR

[Wel03] A. J. Wellings. Is Java augmented with the RTSJ a better real-time … technology than Ada 95? ACM SIGADA Ada Letters, 23(4):16–21, 2003. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wellings:2004:CRT

[Wel04] Andrew J. Wellings. Concurrent and real-time programming in Java. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 2004. ISBN 0-470-84437-X (paperback). xiv + 431 pp. LCCN QA76.73.J38 W465

2004. URL http://www.loc.
gov/catdir/enhancements/
fy0616/2004008522-b.html;
http://www.loc.gov/catdir/
enhancements/fy0616/2004008522-
d.html;   http://www.loc.
gov/catdir/toc/ecip0417/
2004008522.html.

Wells:2006:NIL

[Wel06] George C. Wells. New and improved: Linda in Java. Science of Computer Programming, 59(1–2):82–96, January 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Wenderholm:2005:EJB

[Wen05]  E. Wenderholm.  Eclpss:

a Java-based framework for parallel ecosystem simulation and modeling. Environmental Modelling and Software, 20(9):1081–1100, 2005. CO-DEN ???? ISSN 1364-8152.

Witten:2000:DMP

[WF00] Ian H. Witten and Eibe Frank. Data Mining: Practical Machine Learning Tools and Techniques with Java implementations. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2000. ISBN 1-55860-552-5. xxv + 371 pp. LCCN QA76.9.D343 W58 2000. US$49.95.

URL http://www.mkp.com/
books%5Fcatalog/catalog.
asp?ISBN=1-55860-489-8#
Contents.

Witten:2002:DMP

[WF02] Ian H. Witten and Eibe Frank. Data mining: practical machine learning tools and techniques with Java implementations. SIGMOD, 31 (1):76–77, March 2002.

Washizaki:2004:SSJ

[WF04] Hironori Washizaki and Yoshiaki Fukazawa. A search system for Java programs by using extracted JavaBeans components. ACM SIG-PLAN Notices, 39(10):200– 201, October 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wawersich:2003:SAJ

[WFGK03] C. Wawersich, M. Felser, M. Golm, and J. Kleinoder. The security architecture of the Java operating system JX — A security architecture for distributed parallel computing. Lecture Notes in Computer Science, 2834:85– 95, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Waldron:2001:IQH

[WG01] John Waldron and David Gregg. Identification and quantification of hotspots in Java Grande programs. Lecture Notes in Computer Science, 2110:701–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2110/21100701.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2110/21100701.
pdf.

Walsh:2002:JAJ

[WG02] Aaron E. Walsh and Doug Gehringer. Java 3D API jump-start. Prentice Hall PTR jump-start series. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2002. ISBN 0-13-034076-6. xviii + 245 + 16 pp. LCCN QA76.73.J38 W35

2002. URL http://www.sun.
com/books/catalog/walsh/.

Weaver:2009:PJP

[WGC09] James L. Weaver, Weiqi Gao, and Stephen Chin. Pro JavaFX Platform: Script, Desktop and Mobile RIA with Java Technology. Apress, Berkeley, CA, USA, 2009. ISBN 1-4302-1875-4. xxiii + 543 pp. LCCN ????

Wassermann:2007:SCD

[WGSD07] Gary Wassermann, Carl Gould, Zhendong Su, and Premkumar Devanbu. Static checking of dynamically generated queries in database applications. ACM Transactions on Software Engineering and Methodology, 16(4):14:1– 14:??, September 2007. CO-DEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Woo:2004:AAJ

[WGW04] J. Woo, J. L. Gaudiot, and A. L. Wendelborn. Alias analysis in Java with referenceset representation for highperformance computing. International Journal of Parallel Programming, 32(1):39– 76, 2004. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Whitlock:2001:FPE

[WH01] David Whitlock and Antony L. Hosking. A framework for persistence-enabled optimization of Java object stores. Lecture Notes in Computer Science, 2135:4– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2135/21350004.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2135/21350004.
pdf.

Welch:2001:SVD

[WHBS01] P. H. Welch, G. H. Hilderink, A. W. P. Bakkers, and G. S. Stiles. Safe and verifiable design of concurrent Java programs. International Journal of Computer Applications, 23(3):159–165, 2001. ISSN 1206-212X (print), 1925-7074 (electronic). URL https://

www.tandfonline.com/doi/
full/10.1080/1206212X.2001.
11441647.

Whitbread:2003:DJS

[Whi03a] Martin Whitbread. DCT Java solution for wireless devices and Renesas’ latest 32bit RISC processor. Embedded System Engineering, 11 (6):36–37, 2003. CODEN ???? ISSN 0969-8825.

White:2003:UTL

[Whi03b] Tom White. Using threadlocal variables in Java. Dr. Dobb’s Journal of Software Tools, 28(7):42, 44–46, July 2003. CODEN DDJOEB. ISSN 1044-789X. URL

http://www.ddj.com/ftp/
2003/2003_07/thread.txt;
http://www.ddj.com/ftp/
2003/2003_07/thread.zip.

Wissink:2001:PSA

[WHKS01] Andrew Wissink, Richard Hornung, Scott Kohn, and Steve Smith. Parallel structured adaptive mesh refinement applications using the SAMRAI library. In ACM [ACM01b], page 184. ISBN 1-58113-359-

6. LCCN QA76.9.O35 A26

2001. URL http://www.
philippsen.com/JGI2001/
camerareadyabstracts/57.
html;   http://www.philippsen.
com/JGI2001/finalpapers/
18500184.ps.

Wirthlin:2001:SRH

[WHW01] Michael J. Wirthlin, Brad L. Hutchings, and Carl Worth. Synthesizing RTL hardware from Java byte codes. Lecture Notes in Computer Science, 2147:123–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2147/21470123.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2147/21470123.
pdf.

Wick:2003:OOR

[Wic03] M. R. Wick. An objectoriented refactoring of Huffman encoding using the Java collections framework. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35 (1):283–287, 2003. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wiedermann:2008:IQE

[WIC08] Ben Wiedermann, Ali Ibrahim, and William R. Cook. Interprocedural query extraction for transparent persistence. ACM SIGPLAN Notices, 43(10):19–36, September 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Williams:2000:TII

[Wil00a] Al Williams. The TINI Internet interface. Dr. Dobb’s Journal of Software Tools, 25 (10):82, 84, 86, 88, October 2000. CODEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/ftp/2000/
2000_10/tini.txt.

Wilson:2000:PBA

[Wil00b] Gregory V. Wilson. Programmer’s bookshelf: All this, and frequent flyer miles, too! Dr. Dobb’s Journal of Software Tools, 25(2):135–136, February 2000. CODEN DDJOEB. ISSN 1044-789X.

Wilson:2000:PBC

[Wil00c] Gregory V. Wilson. Programmer’s bookshelf: Classics old and new. Dr. Dobb’s Journal of Software Tools, 25(11): 159–160, November 2000. CODEN DDJOEB. ISSN 1044-789X.

Wilson:2000:PBS

[Wil00d] Gregory V. Wilson. Programmer’s bookshelf: Stirring the gene pool. Dr. Dobb’s Journal of Software Tools, 25(5): 127, 130, May 2000. CODEN DDJOEB. ISSN 1044-789X.

Williams:2001:JWT

[Wil01a] Al Williams. Java and the Waba toolkit. Dr. Dobb’s Journal of Software Tools, 26 (2):92, 94, 96, 98, February 2001. CODEN DDJOEB. ISSN 1044-789X. URL http:

                                        //www.ddj.com/ftp/2001/
                                        2001_02/jwaba.txt.

Wilson:2001:PBT

[Wil01b] Gregory V. Wilson. Programmer’s bookshelf: Time warps. Dr. Dobb’s Journal of Software Tools, 26(2):159– 160, February 2001. CODEN DDJOEB. ISSN 1044-789X.

                                        URL http://www.ddj.com/.

Wildmoser:2002:SJB

[Wil02] M. Wildmoser. Subroutines and Java bytecode verification. Master’s thesis, Institut für Informatik, Technische Universität München, Munich, Germany, 2002.

Wilson:2003:PB

[Wil03a] Gregory V. Wilson. Programmer’s bookshelf. Dr. Dobb’s Journal of Software Tools, 28 (7):75, July 2003. CODEN DDJOEB. ISSN 1044-789X.

                                        URL http://www.ddj.com/.

Wilson:2003:PBF

[Wil03b] Gregory V. Wilson. Programmer’s bookshelf: Facts, fallacies, and numeric computing. Dr. Dobb’s Journal of Software Tools, 28(6):76, June 2003. CODEN DDJOEB.

ISSN 1044-789X. URL http:

//www.ddj.com/documents/
s=8213/ddj0306j/.

Wilson:2003:PBP

[Wil03c] Gregory V. Wilson. Programmer’s bookshelf: Parkinson’s other law. Dr. Dobb’s Journal of Software Tools, 28(4):83, April 2003. CO-DEN DDJOEB. ISSN 1044-789X. URL http://

www.ddj.com/documents/s=
7827/ddj0304k/.

Wilson:2003:PBO

[Wil03d] Gregory V. Wilson. Programmer’s bookshelf: The observation of trifles. Dr. Dobb’s Journal of Software Tools, 28(1):73, January 2003. CODEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/documents/
s=7758/ddj0301j/.

Williams:2004:MAJ

[Wil04a] A. Williams. Mixing ActiveX with Java. Dr. Dobb’s Journal of Software Tools, 29(7):64– 70, 2004. CODEN DDJOEB. ISSN 1044-789X.

Willsey:2004:BLD

[Wil04b] J. W. Willsey. Browsersoft lets developers write a little Java. Application Development Trends, 11(10):46–47, 2004. CODEN ???? ISSN 1073-9564.

Wilson:2005:DCS

[Wil05] Greg Wilson. Data crunching: solve everyday problems using Java, Python and more. The pragmatic programmers. Pragmatic Bookshelf, Raleigh, NC, USA, 2005. ISBN 0-9745140-7-1. viii + 193 pp. LCCN QA63 .W55 2005.

Williams:2006:LRD

[Wil06] Rob Williams. Languages for RTS development — C, Ada and Java. Real-Time Systems Development, ??(??): 341–357, ???? 2006. CODEN ???? ISSN ???? URL http:

//www.sciencedirect.com/
science/article/pii/B9780750664714500189.

Wincelberg:2001:JQH

[Win01]  David Wincelberg.  Java

Q&A: How do you use the

javax.crypto package? Dr.

Dobb’s Journal of Software Tools, 26(4):139–140, April 2001. CODEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/ftp/2001/
2001_04/jqa0401.txt;   http:
//www.ddj.com/ftp/2001/
2001_04/jqa0401.zip.

Winkler:2002:SVU

[Win02] Jürgen F. H. Winkler. A safe variant of the unsafe integer arithmetic of Java^{T M}. Software—Practice and Experience, 32(7): 669–701, June 2002. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (elec-

tronic). URL http://www3.
interscience.wiley.com/
cgi-bin/abstract/94515736/

START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=94515736&PLACEBO=IE.
pdf.

Winkler:2004:CCJ

[Win04] J. F. H. Winkler. C#: a competitor of Java? It, 46(2):85– 93, 2004. CODEN ???? ISSN 1611-2776.

Wise:2006:GJD

[Wis06] Jon Wise. GoJava: a Java development tool for beginners. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38 (3):359, September 2006. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wittenberg:2000:PTC

[Wit00] Craig H. Wittenberg. Progress in testing component-based software (abstract only). ACM SIGSOFT Software Engineering Notes, 25(5): 178, September 2000. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wittmer:2005:EPC

[Wit05] K. Wittmer. EPerl: Perl, C++, and Java. C/C++ Users Journal, 23(7):22–25, 2005. CODEN CCUJEX. ISSN 1075-2838.

Welc:2005:SFJ

[WJH05] Adam Welc, Suresh Jagannathan, and Antony Hosking. Safe futures for Java.

ACM SIGPLAN Notices, 40 (10):439–453, October 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Welc:2006:RTJ

[WJH06] Adam Welc, Suresh Jagannathan, and Antony L. Hosking. Revocation techniques for Java concurrency. Concurrency and Computation: Practice and Experience, 18(12):1613–1656, October 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Winiecki:2002:NJB

[WK02] W. Winiecki and M. Karkowski. A new Java-based software environment for distributed measuring systems design. IEEE Transactions on Instrumentation and Measurement, 51(6):1340–1346, 2002. CO-DEN IEIMAO. ISSN 0018-9456 (print), 1557-9662 (electronic).

Wegiel:2008:MCVa

[WK08a] Michal Wegiel and Chandra Krintz. The mapping collector: virtual memory support for generational, parallel, and concurrent compaction. ACM SIGARCH Computer Architecture News, 36(1):91– 102, March 2008. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Wegiel:2008:MCVb

[WK08b] Michal Wegiel and Chandra Krintz. The Mapping Collector: virtual memory support for generational, parallel, and concurrent compaction. Operating Systems Review, 42(2): 91–102, March 2008. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wegiel:2008:MCVc

[WK08c] Michal Wegiel and Chandra Krintz. The mapping collector: virtual memory support for generational, parallel, and concurrent compaction. ACM SIGPLAN Notices, 43(3):91–102, March 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wegiel:2008:XTS

[WK08d] Michal Wegiel and Chandra Krintz. XMem: type-safe, transparent, shared memory for cross-runtime communication and coordination. ACM SIGPLAN Notices, 43(6): 327–338, June 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wegiel:2009:DPC

[WK09] Michal Wegiel and Chandra Krintz. Dynamic prediction of collection yield for managed runtimes. ACM SIG-PLAN Notices, 44(3):289– 300, March 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wyatt:2002:ISI

[WKB02] Jason N. Wyatt, Martha J. Kosa, and Mark A. Boshart. Implementing student ideas in CS2: a simple IDE. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34 (3):224, September 2002. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wen:2004:IDE

[WL04] Y. Wen and X. Liu. Intelligence data extraction based on Java XML and Web. Journal- Sichuan University Natural Science Edition, 41 (2):318–321, 2004. CODEN ???? ISSN 0490-6756.

Wang:2003:DIE

[WLW^{+}03] J. Wang, T. Lin, J. Wang, G. Han, and H. Zhao. Design and implementation of an embedded real-time Java OS. Journal — China Institute of Communications, 24 (8):78–87, 2003. CODEN ???? ISSN 1000-436X.

Walker:2000:ICE

[WM00a] Robert J. Walker and Gail C. Murphy. Implicit context: easing software evolution and reuse. ACM SIGSOFT Software Engineering Notes, 25 (6):69–78, November 2000. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Whelan:2000:MVA

[WM00b] Paul F. Whelan and Derek Molloy. Machine vision algorithms in Java: techniques and implementation. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. ISBN 1-85233-218-2. xvii + 284 pp. LCCN TA1634 .W54 2000.

Weaver:2004:BJN

[WMC04] James L. Weaver, Kevin Mukhar, and Jim Crume. Beginning J2EE 1.4: From Novice to Professional: The Essentials of the Platform. Apress, Berkeley, CA, USA, 2004. ISBN 1-59059-341-3. 700 (est.) pp. LCCN ????

Whaley:2002:AEO

[WML02] John Whaley, Michael C.

Martin, and Monica S. Lam. Automatic extraction of object-oriented component interfaces. ACM SIG-SOFT Software Engineering Notes, 27(4):218–228, July 2002. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wutka:2004:STY

[WMM04] Mark Wutka, Alan Moffet, and Kunal Mittal. Sams teach yourself Javaserver Pages 2.0 in 24 hours, complete starter kit. Howard W. Sams, Indianapolis, IN 46268, USA, 2004. ISBN 0-672-32597-7. xvi + 527 pp. LCCN TK5105.8885.J38 W88 2004.

Wakelin:2005:CTI

[WMRT^{+}05] J. Wakelin, P. Murray-Rust, S. Tyrrell, Y. Zhang, H. S. Rzepa, and A. Garcia. CML tools and information flow in atomic scale simulations. Molecular simulation, 31(5): 315–322, April 2005. CODEN MOSIEA. ISSN 0892-7022.

Winston:2001:J

[WN01] Patrick Henry Winston and Sundar Narasimhan. On to Java 2. Addison-Wesley, Reading, MA, USA, third edition, 2001. ISBN 0-201-72593-2 (paperback). xi + 451 pp. LCCN QA76.73.J38 W56 2001.

Wicentowski:2005:UIP

[WN05] Richard Wicentowski and Tia Newhall. Using image processing projects to teach CS1 topics. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):287–291, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Weimer:2008:ESP

[WN08] Westley Weimer and George C. Necula. Exceptional situations and program reliability. ACM Transactions on Programming Languages and Systems, 30(2): 8:1–8:51, March 2008. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Wolf:2001:ACH

[Wol01a] Armin Wolf. Adaptive constraint handling with CHR in Java. Lecture Notes in Computer Science, 2239:256– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2239/22390256.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2239/22390256.
pdf.

Wolz:2001:TDP

[Wol01b] Ursula Wolz. Teaching design and project management with Lego RCX robots. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33 (1):95–99, March 2001. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wolle:2003:KAS

[Wol03a] B. Wolle. Kommentar zum artikel “Statische Analyse von Java-Anwendungen… ”. Wirtschaftsinformatik, 45(6): 663, 2003. CODEN WIINE9. ISSN 0937-6429.

Wolle:2003:SAJ

[Wol03b] B. Wolle. Statische Analyse von Java — Anwendungen Eignen sich Lines-of-Code-Metrik und Halstead-Lange? Softwarewartung, Grossenmasse, LOC-Metrik, Halstead-Metrik, Java-Systeme. (German) [Static analysis of Java — are lines-of-code metrics and Halstead lengths suitable? software maintenance, large mass, LOC metrik, Halstead metric, Java systems]. Wirtschaftsinformatik, 45(1):29–40, 2003. CODEN WIINE9. ISSN 0937-6429.

Wolfe:2004:TJJ

[Wol04] Alexander Wolfe. Toolkit: Java is jumpin’. ACM Queue: Tomorrow’s Computing Today, 1(10):16–19, February 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Wong:2003:JPC

[Won03a] W. Wong. Java processor chip stacks JVM, SRAM, Flash. Electronic Design, 51(6):32, 2003. CODEN ELODAW. ISSN 0013-4872.

Wong:2003:SJ

[Won03b] W. Wong. Silent Java. Electronic Design, 51(8):61, 2003. CODEN ELODAW. ISSN 0013-4872.

Wong:2004:JPN

[Won04] W. Wong. Java 5.0 perks up with new language constructs. Electronic Design, 52 (26):44–45, 2004. CODEN ELODAW. ISSN 0013-4872.

Wong:2005:RTJ

[Won05] W. Wong. Real-time Java, CORBA ORB perk up at ESC 2005. Electronic Design, 53(9):36–37, 2005. CODEN ELODAW. ISSN 0013-4872.

Wootton:2001:JPR

[Woo01] Cliff Wootton. JavaScript programmer’s reference. Programmer to programmer. Wrox Press, Chicago, IL, USA, 2001. ISBN 1-86100-459-1 (paperback). xix + 973 pp. LCCN QA76.76.J39 W66 2001. Available also in CD-ROM format.

Wood:2002:JPS

[Woo02] Jo Wood. Java programming for spatial sciences. Taylor and Francis, New York, NY, USA, 2002. ISBN 0-415-26098-1 (paperback), 0-415-26097-3 (cloth). 320 pp. LCCN QA76.73.J38 W6615 2002.

Woods:2003:MJB

[Woo03] R. P. Woods. Multitracer: a Java-based tool for anatomic delineation of grayscale volumetric images. NeuroImage, 19(4):1829–1834, 2003. CO-DEN NEIMEF. ISSN 1053-

Woodward:2004:XPS

[Woo04] W. Woodward. XML persistence self-describing serialized Java objects. Research Disclosure, 484:1134, 2004. CO-DEN ???? ISSN 0374-4353.

Woo:2005:SAJ

[Woo05] Jongwook Woo. Static analysis for Java with alias representation reference-set in high-performance computing. Scalable Computing: Practice and Experience, 6(1): 125–139, March 2005. CO-DEN ???? ISSN 1895-1767.

URL http://www.scpe.org/
vols/vol06/SCPE_6_1_10.
pdf;   http://www.scpe.
org/vols/vol06/SCPE_6_1_
10.zip.

Workman:2002:CMT

[Wor02]  David A. Workman.  A

class and method taxonomy for object-oriented programs. ACM SIGSOFT Software Engineering Notes, 27 (2):53–58, March 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Wiener:2000:FOD

[WP00a] Richard Wiener and Lewis Pinson. Fundamentals of OOP and Data Structures in Java. Cambridge University Press, Cambridge, UK, 2000. ISBN 0-521-66220-6. xv + 463 pp. LCCN QA76.73.J38 W53 2000. US$50.

Wu:2000:CPG

[WP00b] Peng Wu and David Padua.

Containers on the parallelization of general-purpose Java programs. International Journal of Parallel Programming, 28(6):589–605, December 2000. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0885-7458&volume=28&
issue=6&spage=589.

Wellings:2003:EEP

[WP03] A. J. Wellings and P. Puschner.

Evaluating the expressive power of the real-time specification for Java. Real-Time Systems, 24(3):319–359, 2003. CODEN RESYE9. ISSN 0922-6443.

Weatherly:2004:EPI

[WP04] R. M. Weatherly and E. H.

Page. Efficient process interaction simulation in Java: Implementing co-routines within a single Java thread. Winter Simulation Conference, 2: 1437–1443, 2004. CODEN ???? ISSN 0743-1902 (print), 2160-9276 (electronic).

Willis:2008:CIJ

[WPN08] Darren Willis, David J.

Pearce, and James Noble. Caching and incrementalisation in the Java Query Language. ACM SIGPLAN Notices, 43(10):1–18, September 2008. CODEN SINODQ.

ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Winder:2000:DJS

[WR00] Russ Winder and Graham Roberts. Developing Java Software. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, second edition, 2000. ISBN 0-471-60696-0. xvi + 1004 pp. LCCN QA76.73.J38 W557 2000. US$50.

Wang:2008:DSJ

[WR08] Tao Wang and Abhik Roychoudhury. Dynamic slicing on Java bytecode traces. ACM Transactions on Programming Languages and Systems, 30(2):10:1–10:49, March 2008. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Wraxall:2001:JQH

[Wra01] Jason Wraxall. Java Q&A: How can I extend the Swing JComboBox? Dr. Dobb’s Journal of Software Tools, 26(1):181–182, 184, 186, January 2001. CO-DEN DDJOEB. ISSN 1044-789X. URL http:

                                        //www.ddj.com/ftp/2001/
                                        2001_01/jqa0101.txt;   http:
                                        //www.ddj.com/ftp/2001/
                                        2001_01/jqa0101.zip.

Wright:2003:JES

[Wri03] William Wright. J2Me and embedded systems. Dr. Dobb’s Journal of Software Tools, 28(3):54, 56, 58, March 2003. CODEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/documents/
s=7826/ddj0303g/.

Walls:2004:XA

[WRO04] Craig Walls, Norman Richards, and Rickard Oberg. XDoclet in Action. Manning Publications, Greenwich, CT, USA, 2004. ISBN 1-932394-05-2. xxxii + 591 pp. LCCN QA76.73.J38 W343 2004. US$44.95. URL ftp:/

/uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
bios/ipg051/2004271287.
html;   http://www.loc.
gov/catdir/description/
ipg051/2004271287.html.

Wang:2001:FDW

[WS01a] Shaofeng Wang and Jiaguang Sun. A framework design for workflow management system with Java RMI. ACM SIG-PLAN Notices, 36(9):86–93, September 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wang:2001:PCB

[WS01b] Tiejun Wang and Scott F.

Smith. Precise constraintbased type inference for Java. Lecture Notes in Computer Science, 2072:99–??, 2001. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.
                                        com/link/service/series/
                                        0558/bibs/2072/20720099.
                                        htm;   http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2072/20720099.
                                        pdf.

Welch:2001:KUB

[WS01c] Ian Welch and Robert J. Stroud. Kava — using byte code rewriting to add behavioural reflection to Java. In USENIX [USE01a], page ?? ISBN 1-880446-12-X. LCCN

                                        ????  URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/coots01/
                                        welch.html.

Warth:2006:SSOa

[WSM06] Alessandro Warth, Milan Stanojević, and Todd Millstein. Statically scoped object adaptation with expanders. ACM SIGPLAN Notices, 41(10):37–56, October 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wick:2002:UEC

[WSP02] Michael R. Wick, Daniel E. Stevenson, and Andrew T. Phillips. Using an environment chain model to teach inheritance in C++. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):297– 301, March 2002. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Wang:2003:IJM

[WSVX03] L. Wang, R. Sams, M. Verner, and F. Xi. Integrating Java 3D model and sensor data for remote monitoring and control. Robotics and Computer Integrated Manufacturing, 19 (1–2):13–19, 2003. CODEN ???? ISSN 0736-5845.

Weyns:2003:SDE

[WTV03] D. Weyns, E. Truyen, and P. Verbaeten. Serialization of distributed execution-state in Java. Lecture Notes in Computer Science, 2591:41– 61, 2003. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weyns:2005:SDT

[WTV05] Danny Weyns, Eddy Truyen, and Pierre Verbaeten. Serialization of distributed threads in Java. Scalable Computing: Practice and Experience, 6(1):81–98, March 2005. CO-DEN ???? ISSN 1895-1767.

URL http://www.scpe.org/
vols/vol06/SCPE_6_1_07.
pdf;   http://www.scpe.
org/vols/vol06/SCPE_6_1_
07.zip.

Wu:2001:IOO

[Wu01] C. Thomas Wu. An introduction to object-oriented programming with Java. Mc-Graw-Hill, New York, NY, USA, second edition, 2001. ISBN 0-07-239684-9 (softcover). xxviii + 863 pp. LCCN QA76.64. W78 2000.

Wu:2005:TGA

[Wu05]  Mingshen Wu.  Teaching

graph algorithms using online Java package IAPPGA. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37 (4):64–68, December 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp:

                                        //ftp.math.utah.edu/pub/
                                        mirrors/ftp.ira.uka.de/
                                        bibliography/Misc/DBLP/
                                        2005.bib.

Wutka:2000:SEU

[Wut00] Mark Wutka. Special edition using Java Server Pages and servlets. Que Corporation, Indianapolis, IN, USA, 2000. ISBN 0-7897-2441-3. xi + 754 pp. LCCN QA76.73.J38 W87 2000.

Weis:2000:HMD

[WVE^{+}00] Rüdiger Weis, Jürgen Vogel, Wolfgang Effelsberg, Werner Geyer, and Stefan Lucks. How to make a digital whiteboard secure — using Java cards for multimedia applications. Lecture Notes in Computer Science, 1905:187– ??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/1905/19050187.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/1905/19050187.
pdf.

Weir:2005:DTJ

[WVMN05] George R. S. Weir, Tamar Vilner, António José Mendes, and Marie Nordström. Difficulties teaching Java in CS1 and how we aim to solve them. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):344–345, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

White:2006:JJF

[WW06] Timothy M. White and Thomas P. Way. jFAST: a Java finite automata simulator. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):384–388, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wang:2009:AHC

[WW09] Yi-Hsien Wang and I-Chen Wu. Achieving high and consistent rendering performance of Java AWT/Swing on multiple platforms. Software— Practice and Experience, 39 (7):701–736, May ??, 2009. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wang:2007:PAS

[WWJ07] Yi-Hsien Wang, I-Chen Wu, and Jyh-Yaw Jiang. A portable AWT/Swing architecture for Java game development. Software—Practice and Experience, 37(7): 727–745, June 2007. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wright:2006:IJV

[WWMG06] Greg Wright, Mario Wolczko, Phil McGachey, and Erika Gunadi. Introspection of a Java Virtual Machine under simulation. Technical Report TR-2006-159, Sun Microsystems, Inc., Menlo Park, CA, USA, September 2006. URL

                                        http://research.sun.com/
                                        techrep/2006/smli_tr-2006-
                                        159.pdf.

Wang:2002:JEC

[WWSL02] L. Wang, B. Wong, W. Shen, and S. Lang. A Java 3Denabled cyber workspace. Communications of the ACM, 45(11):45–49, November 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wang:2005:JBG

[WXW^{+}05] Y. Wang, Y. Xue, J. Wang, C. Wu, Y. Hu, Y. Luo, S. Zhong, J. Tang, and G. Cai. Java-based Grid service spread and implementation in remote sensing applications. Lecture Notes in Computer Science, 3516:496– 503, 2005. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Xu:2009:GFP

[XAM^{+}09] Guoqing Xu, Matthew Arnold, Nick Mitchell, Atanas Rountev, and Gary Sevitsky. Go with the flow: profiling copies to find runtime bloat. ACM SIGPLAN Notices, 44(6): 419–430, June 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Xiao:2007:HIB

[XAN07] Gaoyu Xiao, Aamer Aziz, and Wieslaw L. Nowinski. Hybrid image-based collision detection in Java 3D. Software —Practice and Experience, 37 (9):963–982, July 25, 2007. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Xu:2001:DAR

[XC01] Baowen Xu and Zhenqiang Chen. Dependence analysis for recursive Java programs. ACM SIGPLAN Notices, 36(12):70–76, December 2001. CODEN SINODQ.

ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Xu:2009:SCC

[XJC09] Dana N. Xu, Simon Peyton Jones, and Koen Claessen. Static contract checking for Haskell. ACM SIGPLAN Notices, 44(1):41–52, January 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Xu:2003:MEJ

[XLG03] Baomin Xu, Weimin Lian, and Qiang Gao. Migration of enterprise JavaBeans with ProActive Interposition Objects. ACM SIGPLAN Notices, 38(8):22–28, August 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Xu:2006:CCT

[XM06] Li Xu and Fred G. Martin. Chirp on crickets: teaching compilers using an embedded robot controller. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38 (1):82–86, March 2006. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Xu:2006:PMP

[XOWM06] Jing Xu, Alexandre Oufimtsev, Murray Woodside, and Liam Murphy. Performance modeling and prediction of enterprise JavaBeans with layered queuing network templates. ACM SIGSOFT Software Engineering Notes, 31(2):5:1–5:??, March 2006. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Xiang:2004:RWG

[XP04] N. Xiang and H. Peng. A research of WEB GIS based on Java and XML. Journal of the Hebei Institute of Technology, 26(3):46–48, 2004. CODEN ???? ISSN 1007-2829.

Xian:2008:CAS

[XSaJ08a] Feng Xian, Witawas Srisa-an, and Hong Jiang. Contentionaware scheduler: unlocking execution parallelism in multithreaded Java programs. ACM SIGPLAN Notices, 43(10):163–180, September 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Xian:2008:GCJ

[XSaJ08b] Feng Xian, Witawas Srisa-an, and Hong Jiang. Garbage collection: Java application servers’ Achilles heel. Science of Computer Programming, 70(2–3):89–110, February 1, 2008. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Xinogalos:2007:TJB

[XSD07] Stelios Xinogalos, Maya Satratzemi, and Vassilios Dagdilelis. Teaching Java with BlueJ: a two-year experience. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):345, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Xu:2004:MAO

[XX04] B. Xu and A. Xu. Mobile active objects in Java. Advances in Engineering Software, 35 (3–4):231–235, 2004. CO-DEN AESODT. ISSN 0965-9978 (print), 0141-1195 (electronic).

Xu:2005:NER

[XX05] K. Xu and L. Xin. Novel 32bit embedded reducedinstruction-set-computer with Java extension. Mini-Micro Systems, 26(1):90–95, 2005. CODEN XWJXEH. ISSN 1000-1220.

Xu:2005:OPJ

[XYC05] Zhigen Xu, Yusong Yan, and Jim X. Chen. OpenGL programming in Java. Computing in Science and Engineering, 7(1):51–55, January/February 2005. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://csdl.
computer.org/dl/mags/cs/
2005/01/c1051.htm;   http:
//csdl.computer.org/dl/
mags/cs/2005/01/c1051.pdf.

Xu:2003:MLP

[XZ03] Y. Xu and G. Zhang. MPLS loop prevention mechanism and its Java realization. Mini-Micro Systems, 24(4): 706–708, 2003. CODEN XWJXEH. ISSN 1000-1220.

Yang:2007:DPP

[YAA07] X. Yang, A. Akram, and R. J.

Allan. Developing portals/ portlets using Enterprise JavaBeans for Grid users. Concurrency and Computation: Practice and Experience, 19 (12):1633–1641, August 25, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Yahav:2001:VSP

[Yah01]  Eran Yahav.  Verifying

safety properties of concurrent Java programs using 3valued logic. ACM SIG-PLAN Notices, 36(3):27– 40, March 2001. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL

http://www.acm.org/pubs/
articles/proceedings/plan/
360204/p27-yahav/p27-yahav.
pdf;   http://www.acm.org/
pubs/citations/proceedings/
plan/360204/p27-yahav/.

Yamamoto:2004:NGM

[Yam04]  T. Yamamoto.  Next-

generation mobile tool by Java for mobile phone “sMobile”. Sumitomo Heavy Industries Technical Review, 155: 39–42, 2004. CODEN ????

Yan:2002:RCC

[Yan02] C. Yan. Race condition and concurrency safety of multithreaded object-oriented programming in Java. IEEE International Conference on Systems Man and Cybernetics, 6:??, 2002. CODEN ???? ISSN 1062-922X.

Yang:2003:WPT

[Yan03] C. D. Yang. Work in progress: Teaching wireless networking and security with Java 2 Micro Edition (J2ME). Frontiers in Education Conference, 1(??):T2C–7, 2003. CO-DEN PFECDR. ISSN 0190-5848.

Yan:2005:EPC

[Yan05] Lu Yan. Enable pervasive computing with Java. IEEE Distributed Systems Online, 6(6):??, June 2005. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/comp/mags/
                                        ds/2005/06/o6004.pdf.

Yuniar:2002:KFJ

[YAW02] A. T. Yuniar, Andreas, and T. Walter. ,,Kodok” — Froschzucht auf Java. Datz, 55(12):18–21, 2002. CODEN ???? ISSN 0723-4066.

Yiyu:2009:IFS

[YCFX09] Tan Yiyu, Yau Chihang, Anthony S. Fong, and Yang Xiaojian. An instruction folding solution for a Java processor. International Journal of Computer Systems Science and Engineering, 24(3): ??, May 2009. CODEN CSSEEI. ISSN 0267-6192.

Yu:2007:JIB

[YCIS07] Dachuan Yu, Ajay Chander, Nayeem Islam, and Igor Serikov. JavaScript instrumentation for browser security. ACM SIGPLAN Notices, 42(1):237–249, January 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yero:2005:JIJ

[YdOLS^{+}05] Eduardo Javier Huerta Yero, Fabiano de Oliveira Lucchese, Francisco Sérgio Sambatti, Miriam von Zuben, and Marco Aurélio Amaral Henriques. JoiN: The implementation of a Java-based massively parallel grid. Future Generation Computer Systems, 21(5):791–810, May 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Yang:2004:TWO

[YDWL04] Jiahai Yang, Haixin Duan, Jianping Wu, and Xing Li.

Thresholds: Workflow oriented network management: a Web/Java approach. Journal of Network and Systems Management, 12(4):431–439, December 2004. CODEN ???? ISSN 1064-7570.

Yilmaz:2004:IDC

[YE04] G. Yilmaz and N. Erdogan. Integrating distributed composite objects into Java environment. Lecture Notes in Computer Science, 3261:322– 331, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yero:2001:JOO

[YHGL01] Eduardo J. H. Yero, Marco A. A. Henriques, Javier R. Garcïa, and Alina C. Leyva. JOINT: An object oriented message passing interface for parallel programming in Java. Lecture Notes in Computer Science, 2110:637–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2110/21100637.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2110/21100637.
pdf.

Ye:2001:WBP

[YHL01] R. Ye, W.-J. Hsu, and Z.-H. Liu. Web-based parallel simulation of AGVs using Java and JINI. Lecture Notes in Computer Science, 2127:379– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2127/21270379.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2127/21270379.
pdf.

Yeo:2004:JBW

[YHL04] C. K. Yeo, S. C. Hui, and B. S. Lee. Java-based, WAP-Enabled unified messaging system. International Journal of Computer Applications, 26(4):191–198, 2004. CODEN IJCAFW. ISSN 1206-212X (print), 1925-7074 (electronic). URL https://

www.tandfonline.com/doi/
full/10.1080/1206212X.2004.
11441741.

Yeung:2003:OJR

[YK03] K. C. Yeung and P. H. J.

Kelly. Optimising Java RMI programs by communication restructuring. Lecture Notes in Computer Science, 2672:324–343, 2003. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yavuz-Kahveci:2002:SVS

[YKB02] Tuba Yavuz-Kahveci and Tevfik Bultan. Specification, verification, and synthesis of concurrency control components. ACM SIGSOFT Software Engineering Notes, 27 (4):169–179, July 2002. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Yanagiuchi:2002:LJI

[YKS^{+}02] S. Yanagiuchi, T. Kiyohara, N. Shiraishi, K. Mori, and M. Ohkita. Linux/Java implemented personal mobile tool SL-A300. Sharp Technical Journal, 84:31–32, 2002. CODEN ???? ISSN 0285-0362.

Yang:2003:UPC

[YL03] D. R. Yang and J. H. Lee. Undergraduate process control education using Java applet modules. Proceedings of the American Control Conference, 1(??):694–699, 2003. CODEN PRACEO. ISSN 0743-1619.

Yang:2007:ERM

[YLL^{+}07] Byung-Sun Yang, Junpyo Lee, SeungIl Lee, Seongbae Park, Yoo C. Chung, Suhyun Kim, Kemal Ebcioglu, Erik Altman, and Soo-Mook Moon. Efficient register mapping and allocation in LaTTe, an open-source Java just-in-time compiler. IEEE Transactions on Parallel and Distributed Systems, 18(1):57– 69, January 2007. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Yu:2005:MXD

[YLM^{+}05] Yijun Yu, Jianguo Lu, John Mylopoulos, Weiwei Sun, Jing-Hao Xue, and Erik H. D’Hollander. Making XML document markup international. Software—Practice and Experience, 35(1): 1–14, January 2005. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Yu:2004:EJO

[YLW04] Z. C. H. Yu, F. C. M. Lau, and C. L. Wang. Exploiting Java objects behavior for memory management and optimizations. Lecture Notes in Computer Science, 3302:437– 452, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yu:2008:OCL

[YLW08] Zoe C. H. Yu, Francis C. M.

Lau, and Cho-Li Wang. Object co-location and memory reuse for Java programs. ACM Transactions on Architecture and Code Optimization, 4(4):4:1–4:??, January 2008. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).

Yang:2005:LMJ

[YME05] Byung-Sun Yang, Soo-Mook Moon, and Kemal Ebcioğlu. Lightweight monitors for the Java virtual machine. Software—Practice and Experience, 35(3):281–299, March 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Yiyu:2005:JPM

[YMP^{+}05] T. Yiyu, L. K. Man, L. M. Pak, Y. W. Shing, and S. F. Anthony. 10.4-4: a Java processor for mobile devices. IEEE International Conference on Consumer Electronics, 23RD:437–438, 2005. CO-DEN ???? ISSN 0747-668X.

Young:2002:EXJ

[You02]  David Young.  Enhydra

XMLC Java presentation development. Howard W. Sams, Indianapolis, IN 46268, USA, 2002. ISBN 0-672-32211-

0. xiii + 483 pp. LCCN QA76.76.A65 Y68 2002.

Yutaka:2000:EJV

[YTY00] Oiwa Yutaka, Kenjiro Taura, and Akinori Yonezawa. Extending Java virtual machine with integer-reference conversion. Concurrency: Practice and Experience, 12(6): 407–422, May 2000. CO-DEN CPEXEI. ISSN 1040-

                                        3108. URL http://www3.
                                        interscience.wiley.com/
                                        cgi-bin/abstract/72515729/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=72515729&PLACEBO=IE.
                                        pdf.

Yuan:2002:JQH

[Yua02]  Michael J. Yuan.  Java

Q&A: How do I map SQL database tables to Java objects? Dr. Dobb’s Journal of Software Tools, 27(3): 89–90, 92–94, March 2002. CODEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/ftp/2002/
2002_03/jqa0302.txt;   http:
//www.ddj.com/ftp/2002/
2002_03/jqa0302.zip.

Yuan:2003:EJD

[Yua03] Michael Juntao Yuan. Enterprise J2ME: Developing Mobile Java Applications. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2003. ISBN 0-13-140530-6. xxviii + 452 pp. LCCN QA76.73.J38 Y83 2004. US$49.99.

Yuan:2004:JCH

[Yua04] M. J. Yuan. Java center:

From headless to rich UI: The OSGi framework isn’t just for embedded gateway devices — this versatile set of open specifications is universally useful. Software Development, 12(5): 28–29, 2004. CODEN ???? ISSN 1070-8588.

Yusuf:2004:EMU

[Yus04] Kareem Yusuf. Enterprise Messaging Using JMS and IBM WebSphere. An IBM Press book. IBM Press, 2004. ISBN 0-13-146863-4. xvi + 330 pp. LCCN QA76.73.J38 Y87 2004. URL http:

//java.sun.com/products/
jms/docs.html.

Yanhong:2003:EID

[YWZ03] G. Yanhong, L. Wenfeng, and W. Zhijian. Exploding image database based on B/S using ISAPI and Java. Journal — Wuhan Transportation University, 27(3):388–390, 2003. CODEN ???? ISSN 1006-2823.

Zou:2009:PFT

[ZABL09] Jia Zou, Joshua Auerbach, David F. Bacon, and Edward A. Lee. PTIDES on flexible task graph: real-time embedded system building from theory to practice. ACM SIG-PLAN Notices, 44(7):31–40, July 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zamulin:2003:ABF

[Zam03a] A. Zamulin. An ASM-based formal model of a Java program. Programming and Computer Software; translation of Programmirovaniye (Moscow, USSR) Plenum, 29 (3):130–139, 2003. CO-DEN PCSODA. ISSN 0361-7688 (print), 1608-3261 (electronic).

Zamulin:2003:FSJ

[Zam03b] A. V. Zamulin. Formal semantics of Java expressions and statements. Programming and Computer Software; translation of Programmirovaniye (Moscow, USSR) Plenum, 29(5):259–270, 2003.

CODEN PCSODA. ISSN 0361-7688 (print), 1608-3261 (electronic).

Zaraysky:2002:OJP

[Zar02] Gregory Zaraysky. Optimization of Java programs for embedded systems. Thesis (m.s.), University of California, Santa Cruz, Santa Cruz, CA, USA, 2002.

Zhuang:2003:DBA

[ZAVT03] H. Zhuang, J. Annese, D. J.

Valentino, and A. W. Toga. Displaying brain atlases using a portable Java application: the anatomist [5029-90]. Proceedings of the SPIE — The International Society for Optical Engineering, 5029:790–796, 2003. CODEN PSISDG. ISSN 0277-786X (print), 1996-756X (electronic).

Zhao:2004:GJB

[ZCQS04] Y. Zhao, W. Chen, Y. Qiu, and J. Shi. GVis: a Javabased architecture for Grid enabled interactive visualization. Lecture Notes in Computer Science, 3252:704–711, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zakhour:2006:JTS

[ZCR^{+}06] Sharon Zakhour, Mary Campione, Jacob Royal, Isaac Rabinovitch, Tom Risser, and Mark Hoeber, editors. The Java tutorial: a short course on the basics. The Java series. Addison-Wesley, Reading, MA, USA, fourth edition, 2006. ISBN 0-321-33420-5 (paperback). xxv + 637 pp. LCCN QA76.73.J38 C365

2006. URL http://www.loc.
gov/catdir/toc/ecip0620/
2006028544.html.

Zendra:2002:STC

[ZD02] Olivier Zendra and Karel Driesen. Stress-testing control structures for dynamic dispatch in Java. In USENIX Association [USE02], page ?? ISBN 1-931971-01-3. LCCN QA76.73.J38 J42 2002. URL

http://www.usenix.org/publications/
library/proceedings/javavm02/
zendra.html.

Zdrnja:2009:ATM

[Zdr09] Bojan Zdrnja. Attack trends: Malicious JavaScript insertion through ARP poisoning attacks. IEEE Security & Privacy, 7(3):72–74, May/ June 2009. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Zeadally:2000:IPQ

[Zea00a] S. Zeadally. Implementation and performance of QoSaware Java applications over ATM networks. The Computer Journal, 43(4):266– 273, ???? 2000. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

www3.oup.co.uk/computer_
journal/hdb/Volume_43/Issue_

04/430266.sgm.abs.html;
 http://www3.oup.co.uk/
computer_journal/hdb/Volume_
43/Issue_04/pdf/430266.
pdf.

Zeadally:2000:PEJ

[Zea00b] S. Zeadally. Performance evaluation of a Java-based networking Application Programming Interface (API). Information Processing Letters, 75(5):199–209, October 31, 2000. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.elsevier.nl/gej-ng/
10/23/20/64/30/27/abstract.
html;   http://www.elsevier.
nl/gej-ng/10/23/20/64/30/
27/article.pdf.

ZenilC:2002:GJP

[Zen02]  Héctor Zenil C.  Good

JavaServer pages introduction: a review of Web Development with JavaServer Pages, 2nd edition by Duane K. Fields and Mark A. Kolb. IEEE Distributed Systems Online, 3(6), 2002. ISSN 1541-4922 (print), 1558-1683 (electronic). URL http:

//dsonline.computer.org/
0206/departments/bks_c.
htm.

Zaks:2000:SCJ

[ZFA00] Ayal Zaks, Vitaly Feldman, and Nava Aizikowitz. Sealed calls in Java packages. ACM SIGPLAN Notices, 35(10):83–92, October 2000. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        proceedings/oops/353171/
                                        p83-zaks/.

Zhen:2004:IBS

[ZFK04] Z. Zhen, B. Fei, and L. Kejun. The implementation of 128 bit strong encryption for SSL by using Java applet. Journal — Huazhong University of Science and Technology Nature Science Chinese Edition, 32(4):74–76, 2004. CODEN ???? ISSN 1671-4512.

Zhang:2004:CAD

[ZG04] M. T. Zhang and K. Goldberg. A computer-aided design tool in Java for planar gripper design. Journal of Computing and Information Science in Engineering, 4(1): 43–48, 2004. CODEN ???? ISSN 1530-9827.

Zhang:2003:IJP

[ZGB03] A. Zhang, J. C. Good, and G. B. Berriman. An interactive Java plotting package for astronomy. Astronomical Society of the Pacific Conference Series, 295(??):461–464, 2003. CODEN ???? ISSN 1050-3390.

Zhao:2005:DMC

[Zha05] J. Zhao. A dependence model for concurrency in Java programs. Information, 8(1):

111–126, 2005. CODEN ???? ISSN 1343-4500.

Zuo:2004:FJD

[ZHC04] T. Zuo, J. Han, and P. Chen.

Formalizing Java dynamic loading in HOL. Lecture Notes in Computer Science, 3223:287–304, 2004. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhu:2003:IJC

[Zhu03] B. Zhu. Integrate JAVA with CORBA to implement the middleware of the distributed systems of heterogeneity. Acta Electronica Sinica, 31(9):1313–1316, 2003. CODEN ???? ISSN 0372-2112.

Zhuk:2004:IRA

[Zhu04]  Jeff Zhuk.  Integration-

ready architecture and design: software engineering with XML, Java, .NET, wireless, speech, and knowledge technologies. Cambridge University Press, Cambridge, UK, 2004. ISBN 0-521-52583-7 (paperback). xxx + 609 pp. LCCN QA76.758 .Z48 2004. URL ftp:/

/uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.loc.gov/catdir/
description/cam041/2003065381.
html;   http://www.loc.
gov/catdir/toc/cam041/2003065381.
html.

Zachary:2003:EVA

[ZJ03] Joseph L. Zachary and Peter A. Jensen. Exploiting value-added content in an online course: introducing programming concepts via HTML and JavaScript. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35 (1):396–400, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Zhang:2004:ACU

[ZK04a] Lingli Zhang and Chandra Krintz. Adaptive code unloading for resourceconstrained JVMs. ACM SIGPLAN Notices, 39(7): 155–164, July 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zheng:2004:JBH

[ZK04b] H. Zheng and J. M. Keith. JAVA-based heat transfer visualization tools. Chemical Engineering Education, 38(4): 282–285, 2004. CODEN ???? ISSN 0009-2479.

Zeller:2005:EOS

[ZK05] Andreas Zeller and Jens Krinke. Essential open source toolset: programming with Eclipse, JUnit, CVS, Bugzilla, Ant, Tcl/Tk and more. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 2005. ISBN 0-470-84445-0 (paperback). xii + 392 pp. LCCN QA76.76.D47 Z45 2005. URL http:/

/www.loc.gov/catdir/toc/
ecip053/2004026271.html.

Zhang:2009:ISE

[ZK09] Lingli Zhang and Chandra Krintz. As-if-serial exception handling semantics for Java futures. Science of Computer Programming, 74(5–6): 314–332, March 1, 2009. CO-DEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Zee:2008:FFV

[ZKR08] Karen Zee, Viktor Kuncak, and Martin Rinard. Full functional verification of linked data structures. ACM SIG-PLAN Notices, 43(6):349– 361, June 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zee:2009:IPL

[ZKR09] Karen Zee, Viktor Kuncak, and Martin C. Rinard. An integrated proof language for imperative programs. ACM SIGPLAN Notices, 44(6): 338–351, June 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhang:2005:ROP

[ZL05] S. Zhang and H. Li. Research of object persistence method based on Java technologies. Mini-Micro Systems, 26(2):264–267, 2005. CODEN XWJXEH. ISSN 1000-1220.

Zhang:2008:VTB

[ZLG08] Hua Zhang, Joohan Lee, and Ratan Guha. VCluster: a thread-based Java middleware for SMP and heterogeneous clusters with thread migration support. Software— Practice and Experience, 38 (10):1049–1071, August ??, 2008. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zhang:2003:DIJ

[ZP03] C. H. Zhang and H. K. Pung. The design and implementation of a Jini/Java-based A/V stream control and management. Multimedia Systems, 9 (4):315–326, 2003. CODEN MUSYEW. ISSN 0942-4962.

Zhao:2003:LCF

[ZPV03] Tian Zhao, Jens Palsber, and Jan Vite. Lightweight confinement for featherweight Java. ACM SIGPLAN Notices, 38(11):135–148, November 2003. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhang:2007:ACA

[ZR07] Weilei Zhang and Barbara G. Ryder. Automatic construction of accurate application call graph with library call abstraction for Java. Journal of Software Maintenance and Evolution: Research and Practice, 19(4): 231–252, July 2007. CO-DEN JSMECT. ISSN 1532-060X (print), 1532-0618 (electronic).

Zhang:2001:HJAb

[ZS01a] Xiaolan Zhang and Margo Seltzer. HBench:Java: an application-specific benchmarking framework for Java Virtual Machines. Concurrency and Computation: Practice and Experience, 13 (8–9):775–792, July/August 2001. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract/84503219/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=84503219&PLACEBO=IE.
pdf.

Zhang:2001:HJAa

[ZS01b] Xiaolan Zhang and Margo Seltzer. HBench:JGC — an application-specific benchmark suite for evaluating JVM garbage collector performance. In USENIX [USE01a], page ?? ISBN 1-880446-12-X. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/coots01/
zhang.html.

Zhuang:2006:AEA

[ZSCC06] Xiaotong Zhuang, Mauricio J. Serrano, Harold W. Cain, and Jong-Deok Choi. Accurate, efficient, and adaptive calling context profiling. ACM SIGPLAN Notices, 41(6):263–271, June 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhao:2009:AWL

[ZSZ^{+}09] Yi Zhao, Jin Shi, Kai Zheng, Haichuan Wang, Haibo Lin, and Ling Shao. Allocation wall: a limiting factor of Java applications on emerging multi-core platforms. ACM SIGPLAN Notices, 44(10):361–376, October 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zhou:2002:GCA

[ZT02] Q. Zhou and P. Tarau. Garbage collection algorithms for Java-based Prolog engines. Lecture Notes in Computer Science, 2562:304–319, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zukowski:2001:JC

[Zuk01] John Zukowski. Java collections. Apress, Berkeley, CA, USA, 2001. ISBN 1-893115-92-5. x + 422 pp. LCCN QA76.73.J38 Z84 2001.

Zuse:2003:KAS

[Zus03] H. Zuse. Kommentar zum Artikel “Statische Analyse von Java-Anwendungen… ”. (German) [commentary on the article “Statistical Analysis of the use of Java”]. Wirtschaftsinformatik, 45(6): 662, 2003. CODEN WIINE9. ISSN 0937-6429.

Zbrzezny:2008:TVJ

[ZW08] Andrzej Zbrzezny and Bożena Woźna. Towards verification of Java programs in perICS. Fundamenta Informaticae, 85(1–4):533–548, September 2008. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Zhu:2003:LTJ

[ZWL03] W. Zhu, C. L. Wang, and F. Lau. Lightweight transparent Java thread migration for distributed JVM. In Proceedings of the International Conference on Parallel Processing, pages 465–472. ????, ????, 2003. CODEN ???? ISSN 0190-3918.

ZhongQun:2005:DRM

[ZX05] W. ZhongQun and X. XiaoDong. Dynamic relocation model and algorithm for application components based on Java. Journal — Nanjing University Natural Sciences Edition, 41(2):180–188, 2005. CODEN ???? ISSN 0469-5097. Check author names: looks like confusion of family and personal names??

Zhao:2002:UJB

[ZXNH02] J. Zhao, L. Xiang, K. Nishimi, and T. Harada. Understanding Java bytecode programs using Kafer. Applied Informatics, 2:108–112, 2002. CO-DEN ???? ISSN 1027-2666.

Zheng:2003:JCB

[ZYC03] Q. Zheng, Z. Yao, and G. Chen. A Java–CORBA based design and implementation for remote debugger on clusters. Journal — University of Science and Technology of China, 33(3):351–358, 2003. CODEN ???? ISSN ????

Zhang:2006:JEJ

[ZYZ06] Bao-Yin Zhang, Guang-Wen Yang, and Wei-Min Zheng. Jcluster: an efficient Java parallel environment on a large-scale heterogeneous cluster. Concurrency and Computation: Practice and Experience, 18(12):1541– 1557, October 2006. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).